A Complete Bibliography of Publications in Notes and Records of the Royal Society of

London

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

08 February 2024 Version 1.29

Title word cross-reference

+ [1811]. $100.00 [1667]. $14.95 [1639, 1611]. $168.00 [1636]. $19.95 [1639]. $22.95 [1670]. $24 [1766]. $24.00 [1769]. $24.95 [1593]. $28.50 [1290]. $29.00 [1801]. $29.95 [1763, 1229]. 3 [1330]. $32.50 [1300]. $35.00 [1102]. $39.95 [1138]. $40.00 [1162]. $50 [1615]. $50.00 [1408]. $65 [1383]. $79.95 [1432]. _{2} [1190, 1811, 2775]. β [2531].

-decay [2531]. -inch [1330].

İhsan [2598].

0 [1068, 1272, 1656, 1088, 1249, 1235, 1595, 1071, 1636, 1146, 1248, 1729, 1638, 1594, 1251, 1455, 1550, 1798, 1528, 1388, 1763, 1689, 1542, 1702, 1785, 1433, 1411, 1361, 1508, 1428, 1673, 1520, 1592, 1671, 1655, 1637, 1731, 1782, 1653, 1589, 1229, 1654, 1458, 1591, 1119, 1746, 1780, 1121, 1085, 1141, 1230, 1429, 1386, 1593, 1705, 1457, 1547, 1685, 1688, 1706, 1387, 1707, 1704, 1213, 1505, 1120, 1231, 1434, 1271, 1384, 1714, 1316, 1613, 1767, 1427, 1484, 1207, 1233, 1270, 1548, 1635, 1730, 1784, 1800, 1634, 1687, 1667, 1786, 1250, 1355, 1748, 1099, 1551, 1334, 1487, 1633, 1709, 1232, 1504, 1672]. 0 [1541, 1629, 1529, 1549, 1435, 1471, 1489, 1545, 1615, 1710, 1749, 1611, 1143, 1319, 1598, 1768, 1617, 1408, 1432, 1473, 1488, 1546, 1526, 1507, 1632, 1769, 1750, 1801, 1087, 1425, 1436, 1485, 1453, 1631, 1669, 1713, 1802, 1543, 1459, 1544, 1732, 1234, 1358, 1385, 1383, 1711, 1614, 1674, 1183, 1522, 1382, 1144]. 0-00-215784-5 [1271]. 0-00-217350-6 [1068]. 0-00-255946-3 [1629]. 0-00-257012-2 [1638]. 0-00-257149-8 [1702]. 0-00-272340-9 [1068]. 0-00-710775-7 [1704]. 0-08-042679-4 [1714]. 0-11-290060-7 [1212]. 0-14-011973-6 [1123]. 0-14-022732-6 [1086]. 0-14-026405-1 [1595]. 0-14-043528-X [1433]. 0-19-217779-6 [1145]. 0-19-262145-9 [1299]. 0-19-286202-2 [1551]. 0-19-286203-0 [1549]. 0-19-506695-2 [1383]. 0-19-508463-2 [1427]. 0-19-512346-8 [1656]. 0-19-514365-5 [1673]. 0-19-515376-6 [1713]. 0-19-553274-0 [1205]. 0-19-818369-0 [1548]. 0-19-820657-7 [1435]. 0-19-850031-9 [1429]. 0-19-850154-4 [1488]. 0-19-850364-4 [1543]. 0-19-850531-0 [1542]. 0-19-850594-9 [1711]. 0-19-850614-7 [1632]. 0-19-850697-X [1598]. 0-19-850841-7 [1780]. 0-19-851053-5 [1667]. 0-19-851538-3 [1685]. 0-19-851744-0 [1335]. 0-19-852049-2 [1187]. 0-19-852602-4 [1746]. 0-19-853948-7 [1362]. 0-19-855379-X [1339]. 0-19-856392-2 [1272]. 0-19-856559-3 [1633]. 0-19-858180-7 [1122]. 0-224-03047-7 [1336]. 0-224-04202-5 [1313]. 0-224-04212-2 [1709]. 0-224-05028-1 [1470]. 0-224-05962-9 [1634]. 0-226-28443-3 [1528]. 0-226-87055-3 [1213]. 0-226-87375-7 [1801]. 0-233-98681-2 [1183]. 0-233-98778-9 [1234]. 0-253-33264-8 [1138, 1276]. 0-262-15031-X [1138]. 0-297-82043-5 [1138]. 0-300-04177-2 [1083]. 0-300-05542-0 [1355]. 0-300-05705-9 [1184]. 0-300-05820-9 [1297]. 0-300-07383-6 [1520]. 0-316-64847-7 [1594]. 0-316-85925-7 [1802]. 0-316-85989-3 [1731]. 0-333-45723-4 [1143]. 0-333-51041-0 [1233]. 0-333-54695-4 [1228]. 0-333-55195-8 [1146]. 0-333-90735-3 [1669]. 0-335-15175-2 [1071]. 0-340-55861-X [1207]. 0-3403-3 [1161]. 0-349-10452-2 [1254]. 0-393-31991-1 [1611]. 0-415-20825-4 [1614]. 0-415-27753-1 [1674]. 0-470-86088-X [1749]. 0-471-37790-2 [1593]. 0-471-96502-2 [1434]. 0-473-05700-X [1615]. 0-485-11433-X [1252]. 0-485-12145-X [1473]. 0-521-23503-0 [1069]. 0-521-25586-4 [1117]. 0-521-25587-2 [1069]. 0-521-25588-0 [1069]. 0-521-25589-9 [1069]. 0-521-25590-2 [1117]. 0-521-25591-0 [1117]. 0-521-25625-9 [1118]. 0-521-26173-2 [1101]. 0-521-30694-9 [1120]. 0-521-33499-3 [1085]. 0-521-33537-X [1142]. 0-521-35055-7 [1069]. 0-521-35434-X [1088]. 0-521-36261-X [1099]. 0-521-38249-1 [1206]. 0-521-39414-7 [1232]. 0-521-40077-5 [1269]. 0-521-40099-6 [1179]. 0-521-40503-3 [1157].

0-521-41700-7 [1232]. 0-521-43182-4 [1338]. 0-521-44051-3 [1409]. 0-521-44205-2 [1250]. 0-521-44456-X [1507]. 0-521-44913-8 [1550]. 0-521-45077-2 [1337]. 0-521-46569-9 [1591]. 0-521-47078-1 [1384]. 0-521-49636-5 [1526]. 0-521-52483-0 [1784]. 0-521-55069-6 [1487]. 0-521-55226-5 [1651]. 0-521-55815-8 [1387]. 0-521-55827-1 [1651]. 0-521-55909-X [1353]. 0-521-56100-0 [1425]. 0-521-56102-7 [1471]. 0-521-56137-X [1379]. 0-521-56185-X [1388]. 0-521-56212-0 [1359]. 0-521-56218-X [1356]. 0-521-56220-1 [1357]. 0-521-56221-X [1382]. 0-521-56223-6 [1382]. 0-521-56224-4 [1432]. 0-521-56309-7 [1386]. 0-521-56669-X [1382]. 0-521-56672-X [1294]. 0-521-56796-3 [1425]. 0-521-57212-6 [1436]. 0-521-57369-6 [1505]. 0-521-59032-9 [1458]. 0-521-59278-X [1457]. 0-521-62145-3 [1504]. 0-521-64066-0 [1501]. 0-521-65177-8 [1706]. 0-521-65191-3 [1688]. 0-521-65696-6 [1706]. 0-521-66310-5 [1782]. 0-521-66359-8 [1522]. 0-521-66790-9 [1631]. 0-521-77229-X [1687]. 0-521-80855-3 [1768]. 0-521-81526-6 [1729]. 0-521-89310-0 [1687]. 0-565-00993-1 [1068]. 0-565-01106-5 [1248]. 0-565-01109-X [1248]. 0-565-01144-8 [1248]. 0-571-16744-6 [1249]. 0-571-19647-0 [1730]. 0-631-16816-8 [1210]. 0-631-16818-4 [1211]. 0-631-17295-5 [1181]. 0-631-17906-2 [1231]. 0-660-18757-4 [1732]. 0-674-57623-3 [1138]. 0-691-01099-4 [1428]. 0-691-01206-7 [1617]. 0-691-01678-X [1485]. 0-691-03745-0 [1411]. 0-691-11445-5 [1800]. 0-7043-3954-4 [1138]. 0-7123-0267-0 [1235]. 0-7139-9152-6 [1592]. 0-7139-9277-8 [1529]. 0-7167-2243-7 [1229]. 0-7181-3641-1 [1298]. 0-7188-3017-2 [1710]. 0-7190-3025-0 [1087]. 0-7195-5282-6 [1253]. 0-7472-7579-3 [1655]. 0-7484-0435-X [1410]. 0-7484-0720-0 [1489]. 0-7503-0100-7 [1163]. 0-7503-0130-9 [1164]. 0-7503-0147-3 [1334]. 0-7503-0391-3 [1406]. 0-7503-0453-7 [1459]. 0-7503-0501-0 [1672]. 0-7503-0571-1. [1546]. 0-7503-0763-3 [1689]. 0-7503-0866-4 [1705]. 0-7509-0327-9 [1275]. 0-7509-2184-6 [1544]. 0-7509-2667-8 [1671]. 0-7546-0079-3 [1654]. 0-7546-0749-6 [1767]. 0-7546-0960-X [1750]. 0-7546-3408-6 [1785]. 0-7923-5744-2 [1636]. 0-8018-2488-5 [1119]. 0-8018-3960-2 [1119]. 0-8018-4801-6 [1319]. 0-8020-2667-2 [1102]. 0-8135-1255-7 [1138]. 0-8173-0735-4 [1358]. 0-85036-470-1 [1484]. 0-85115-321-6 [1230]. 0-85115-506-5 [1082]. 0-85115-523-5 [1116]. 0-85115-553-7 [1453]. 0-85115-594-4 [1333]. 0-85115-631-2 [1314]. 0-85115-706-8 [1455]. 0-85115-721-1 [1547]. 0-85115-864-1 [1707]. 0-85214-052-5 [1317]. 0-85274-317-3 [1190]. 0-85323-027-7 [1121]. 0-85484-063-X [1331]. 0-86078-347-2 [1316]. 0-86078-400-2 [1316]. 0-86078-669-2 [1637]. 0-86078-908-X [1786]. 0-86193-241-2 [1613]. 0-86311-555-1 [1340]. 0-86332-918-7 [1332]. 0-86332-930-6 [1315]. 0-86341-251-3 [1545]. 0-87169-220-1 [1408]. 0-87169-924-9 [1769]. 0-87413-379-3 [1186]. 0-88135-294-2 [1763]. 0-89263-314-X [1300]. 0-901805-65-3 [1295]. 0-905193-05-9 [1798]. 0-905193-06-7 [1798]. 0-905355-29-6 [1318]. 0-905927-43-5 [1144]. 0-905958-91-8 [1160].

0-906450-09-8 [1270]. 0-906795-43-5 [1097]. 0-906795-45-1 [1068]. 0-932826-23-7 [1141]. 0-935702-27-X [1290]. 0-946589-27-5 [1182]. 0-947643-61-3 [1293]. 0-948737-14-X [1185]. 0-9516758-0-X [1251]. 0-9523270-3-1 [1385]. 000 [1380]. 0007 [1506]. 0007-0874 [1506]. 0346-7465 [1291]. 0874 [1506].

1 [1209, 1996, 2304, 1616, 1724, 1595, 1572, 637, 1508, 1766, 1653, 1589, 1639, 1524, 1159, 1525, 1547, 1707, 1661, 1384, 1484, 1635, 1405, 1752, 1541, 1783, 1527, 1539, 1611, 1408, 1596, 1502, 1708, 1430, 1385, 1109, 1674, 1431]. 1-55054-616-3 [1639]. 1-84413-082-7 [1783]. 1-85083-016-9 [1140]. 1-85196-040-6 [1296]. 1-85196-085-6 [1292]. 1-85310-270-9 [1188]. 1-85315-116-5 [1098]. 1-85315-162-9 [1180]. 1-85315-264-1 [1431]. 1-85514-291-0 [1635]. 1-85578-021-6 [1255]. 1-85619-066-8 [1159]. 1-85756-132-5 [1317]. 1-85821-642-7 [1502]. 1-85984-854-0 [1541]. 1-86064-228-4 [1430]. 1-86066-172-6 [1616]. 1-86094-204-0 [1653]. 1-86094-390-X [1766]. 1-86125-028-2 [1572]. 1-86176-187-2 [1708]. 1-86189-059-1 [1596]. 1-86197-400-0 [1714]. 1-86207-285-X [1539]. 1-870325-72-9 [1527]. 1-870948-36-X [1100]. 1-873240-48-1 [1209]. 1-873694-02-4 [1360]. 1-900357-08-9 [1524]. 1-900405-04-0 [1508]. 1-902459-02-4 [1525]. 1-902686-02-0 [1589]. 1. [1546]. 11 [615, 363]. 111 [475]. 13.95 [999]. 14 [616]. 140/$280 [1406]. 16 [2887, 707]. 16. [1508]. 1654/5 [2477]. 1660 [1727, 363]. 1660s [2393]. 1662 [1514]. 1663 [363]. 1678 [1182]. 1680 [1651]. 1685 [1085]. 1690 [149]. 17 [124, 604, 708, 580]. 1700 [1270]. 1703 [1315]. 1727 [1157]. 1736 [1501]. 1750 [1232]. 1758 [1905]. 1778 [2773]. 1790s [2588]. 1792 [1271]. 1792-1872 [2925]. 1795 [1233]. 18 [831]. 1800 [1207]. 1820 [1068, 1251, 1232, 1653]. 1823 [1183]. 1830 [1923]. 1830s [2592]. 1837 [1248, 2812]. 1841 [1252]. 1844 [1069]. 1859 [1359]. 1860 [749, 1273, 2330]. 1860s [2344]. 1862 [1118, 1458]. 1876 [1201]. 1880s [2344]. 1882 [1069]. 18th [1547, 1398]. 18th-century [1398]. 19 [361, 537, 397]. 1915 [1757]. 1918 [1138]. 1918/19 [805]. 1929 [1710]. 1930s [2331]. 1938 [14, 15]. 1939 [1435]. 1942 [2210]. 1945 [82, 92]. 1947 [114, 141, 124]. 1948 [142]. 1948-1949 [154]. 1952 [244]. 1953 [268]. 1956 [313, 308]. 1958 [349]. 1959 [361]. 1960 [399, 402, 397, 403, 398, 405, 400, 409, 411]. 1961 [457]. 1963 [490]. 1967 [574, 577, 580]. 1968 [579, 615, 604, 578, 616, 591]. 1969 [637, 624]. 1970 [661]. 1970s [2192]. 1972 [707, 708, 722]. 1973 [745, 752]. 1974 [777, 764, 753]. 1975 [795, 776]. 1976 [821]. 1977 [831, 820]. 1979 [1138]. 1980s [2410]. 1985 [2680]. 1987 [2908, 2923]. 1989 [1073]. 1990s [2588]. 1991 [1135, 1228]. 1992 [1177, 1208]. 1993 [1226]. 1994 [1267]. 1995 [1311]. 1996 [1351]. 1997 [1405]. 1998 [1452]. 1999 [1500, 1633]. 19th [1263, 1132, 1473]. 19th-century [1263, 1132].

2 [1272, 1725, 1071, 1636, 1638, 1572, 1763, 1637, 1119, 1593, 1231, 1683, 1434, 1316, 1613, 1427, 1250, 1551, 1709, 1710, 1708, 1670, 1383, 1183]. 20 [577, 578, 457]. 20.00 [1162]. 2000 [1571, 1573, 1587, 1596]. 2001 [1640]. 2002 [1690]. 2003 [1751, 1752]. 2004 [1803]. 2005 [1832, 1859, 1849]. 2006 [1910, 1899]. 2007 [1946, 1950]. 2008 [1996]. 2009 [2062]. 2010 [2071, 2070, 2069, 2114]. 2011 [2149]. 2012 [2196]. 2013 [2261]. 2014 [2304]. 2015 [2339]. 2016 [2384, 2363]. 2017 [2392, 2442]. 2018 [2421]. 2019 [2472]. 2022 [2562]. 2023 [2609]. 21 [313, 574, 753]. 23 [399, 2071, 142, 268, 591]. 24 [2210]. 250th [2623, 2550, 1056]. 28 [579, 363]. 29 [2261, 1135, 1351, 1690, 776]. 2nd [1270].

3 [1727, 1456, 1726, 1528, 1689, 1103, 1655, 1731, 1639, 1654, 1085, 1685, 1688, 1213, 1832, 1334, 1633, 1504, 1629, 1545, 1768, 1669]. 3-540-19587-4 [1103]. 3-906768-37-6 [1727]. 3-inch [1352]. 30 [1946, 2062, 2196, 2339, 2363, 2392, 1177, 1226, 1311, 661, 777, 752, 821, 308, 1452, 1500, 1573, 1640, 1803, 1859, 2149, 1899, 2114, 722, 2330]. 300th [531]. 31 [764]. 319 [1597]. 32 [1506]. 32.50/US$45.00 [1485]. 35/US$59.95 [1471]. 350th [2070, 1323]. 37 [963, 941]. 39 [975]. 3cm [1852].

4 [1103, 1428, 1746, 1525, 1705, 92, 1714, 1950, 1143, 1432, 1488, 1543, 1430, 1732, 1358, 1614]. 40 [867]. 40/$80 [1572]. 4th [1271].

5 [1673, 1782, 1271, 1800, 1667, 1526, 1144]. 52 [805]. 54 [1361].

6 [1727, 244, 1475, 1616, 1249, 1729, 1204, 1785, 574, 1520, 1592, 1230, 1706, 1505, 575, 1767, 1487, 1319, 1436, 1713, 1544, 1474].

7 [1594, 1798, 1229, 1780, 1121, 1141, 1386, 1704, 1232, 1783, 1435, 1471, 1617, 1632, 1801, 1502, 1453, 1802, 1459].

8 [1656, 1146, 1455, 1550, 1702, 1765, 1671, 1159, 1387, 1270, 1529, 1522]. 81 [1361, 1748]. 81-224-0501-0 [1361]. 81-7484-054-0 [1748]. 81-85324-00-X [1096]. 84j [941]. 86h [963]. 86i [975]. 88 [1456, 1385]. 88-222-4460-5 [1379]. 88-7794-091-3 [1456].

9 [1068, 2385, 2070, 2262, 1524, 1458, 1591, 1429, 1120, 1634, 1527, 1712, 1769, 1631, 820, 1234, 1711, 1923]. 9.95 [1056]. 90 [1712]. 90-6984-357-9 [1712]. 90th [616]. 91-554-3070-8 [1291]. 950 [1475]. 950-9446-24-6 [1475]. 977 [1670]. 977-424-677-2 [1670]. 978 [2385, 2262, 1950, 1923]. 978-0-7546-0131-9 [2262]. 978-1-118-62077-9 [2385]. 978-1-85196-858-9 [1923]. 978-1-85196-970-8 [2151]. 978-1-86239-214-4 [1950]. 981 [1765, 1474]. 981-02-2692-6 [1474]. 981-02-2876-7 [1474]. 981-238-233-X [1765]. 981-238-234-8 [1765].

= [2631].

A. [1209, 2210, 1231, 1714, 1316, 938, 827, 923, 1543, 2137, 1382]. a.m [402, 403, 411]. Ångström [938]. Aaron [1405, 1452, 1500]. Abacus [1254]. abandonment [1109]. Abbey [579, 615, 831, 708, 789, 1832, 580, 776]. Abbott [981]. Aberdeen [749, 727, 282]. Abir [1138]. Abir-Am [1138]. abolished [2354]. abortive [946]. Abraham [1187, 1411, 2684, 1445, 386, 1632]. abridgement [2800]. abridgements [2530]. abroad [2226, 1969]. absorption [938, 2529]. Abstracts [1571]. abundance [2284]. AC [1640, 1690, 1752, 1803, 1859]. Academia [1849]. academic [2599, 2552, 1757, 2310]. academician [1924]. Académie [1793]. Academies [106, 144, 164, 180, 2558, 129, 119, 105, 128, 876, 88, 89]. Academy [3, 186, 1960, 821, 1096, 1948, 1129, 2577, 538, 2558]. Accademia [1379, 1513]. accompt [363]. Account [2745, 2779, 2873, 2597, 492, 1100, 1303, 1502, 663, 1597, 2467]. accumulation [2430]. Achievement [2705, 1139, 2761, 1290, 1188, 1524]. achromatic [1327]. Acoustics [905]. acquaintance [781]. acquisition [2299]. Across [1621, 2316]. Act [1514]. activist [1846]. activities [75, 112, 1024, 2300, 325, 631, 2444]. activity [2605]. Adam [1163, 1190, 1164, 1635]. Adams [1332, 2729, 1332, 1916]. Adcock [2494, 2493]. Addendum [1277, 188, 488, 463, 1236, 2263, 990, 713, 536, 2427, 233]. Addiscombe [873]. addition [559]. Additions [1075, 2200, 1212, 1648, 1581]. Address [579, 615, 1946, 1996, 2062, 2196, 2261, 2304, 2339, 2363, 2392, 1135, 1177, 1226, 1267, 1311, 142, 577, 92, 398, 707, 831, 708, 1351, 1405, 1452, 1500, 1573, 1715, 789, 1640, 1690, 1752, 1832, 1803, 1859, 2149, 1899, 2114, 1569, 580, 591, 776, 820, 795, 1201, 2421, 268]. Addresses [421]. Adelaide [579, 982]. adept [2822, 1485]. administering [2432]. Administration [2693, 1356, 1294]. administrative [104]. administrator [549]. Admiral [1251, 2819]. Admiralty [1608]. Admission [80, 1112, 1154]. Adriaen [2477]. Adrian [831, 1298, 1161, 1253, 1329, 1484]. Advancement [2330, 435, 1239, 2449]. Adventurer [2721, 2724, 1231, 1382]. Adventures [2440, 2855, 1713]. Advisory [850]. aerial [447, 1396]. aeronautical [403, 1894, 1890]. aeronautics [1646, 1662]. aeroplane [1646]. aeroplanes [1662]. aether [559, 596]. aetiology [2332]. Affair [1689, 1280, 2833, 1529]. Affairs [112]. affect [2033]. Africa [272]. African [1673, 1673, 2711, 2310]. After [2676, 1506, 1357, 1072, 1777, 1084, 1015, 1241, 1519, 1083, 1683]. afterlives [2485]. aftermath [1366, 1394, 1415]. Afterword [2486]. Again [2553]. Against [2873, 1261, 1597]. Age [1406, 1508, 2707, 37, 267, 430, 1098, 1487, 1329, 2900, 1998, 2821, 2527, 2887, 1800]. agency [2465, 2295]. Agent [834]. Ages [1379, 2715]. Agnes [1349]. ago [42, 33]. agricultural [2399]. Agronomic [2547]. ahead [2345]. Ahmed [1670]. Ainslie [1585]. Ainsworth [1143]. Air [2049, 832, 2171, 462, 1080]. aircraft [1095, 1330, 1352, 2137]. Aires [1475, 2157]. Airlife [1188]. airs [1188]. Airy [1131, 548, 778, 2576]. Akademie [1712]. al [1069]. Alabama [1358]. Alan [1275, 1434, 1179, 1772, 1272, 1429]. Albert [1249, 397, 398, 2735, 1289]. alchemical [1259, 2822, 1860, 1485, 1843]. Alchemist [1485, 2456]. alchemists [215]. Aldabra [759]. Alder [1731]. Aldershot [1785, 1637, 1654, 1316, 1767, 1786, 1750]. Alecto [1068]. Alexander [1466, 1299, 2238, 950, 213, 628, 1930, 393]. Alfred [1807, 2286, 1119, 867, 2561, 2776, 2777, 346, 351, 94]. Algebra [2846, 1746, 298, 1536]. alia [1314]. Alic [1138]. Alison [1255]. alive [1933]. aliveness [2151]. alkahest [2093, 2317, 1974]. Allahabad [2485]. Allen [1529]. Allgemeine [2308]. Almanack [1958]. Aloes [1306]. alone [1231]. Alorna [2584]. Alps [1621, 1416]. altitude [2296]. alum [1470]. alum-maker [1470]. Am [1138]. Amateur [2069]. Amazonian [2561]. Ambassador [361]. Ambiguous [2135]. Ambrose [281, 283]. America [2150, 2381, 249, 812]. American [1408, 1769, 1670, 1358, 66, 186, 809, 287, 868, 969, 2710, 237, 2464, 1358, 2549]. ammonoid [1060]. among [2576, 832]. amongst [2582]. Amory [349]. Amoy [2300]. Ampère [1711, 2888, 2739, 1357]. amplifier [1329]. Amsterdam [1712, 14, 210]. Analogies [1126]. analysing [2444]. Analysis [1411, 2684, 1883, 2416, 1993, 802, 2488, 2458, 1698, 2467, 2016]. Analytica [1993]. anatomical [2443]. anatomist [1297]. anatomized [1292]. Anatomy [1248, 2812, 2283, 2183]. ancestors [2104]. ancestry [1720]. Ancient [2458, 1324, 1343]. Anderson [1212, 2262, 1056]. André [1234, 662, 1183, 2739, 1169, 1357]. André-Marie [2739, 1357]. Andreas [245]. Andrew [1779, 1705, 1801, 1651]. Andrewes [2332]. Andrews [2299, 1695]. Andrewy [715]. anecdotal [2564, 2563]. Angel [2513]. Angelo [782]. Anglo [809, 1062, 1218, 287, 959]. Angry [2816, 1542]. Angus [743, 912]. animal [2317, 909, 567]. Animals [2672, 2186, 872, 1551]. animate [2603]. Anita [1273]. Anna [1380]. Annali [1523]. Anne [1484, 2585]. Anniversaries [88, 105, 128, 167, 1679, 1693, 1737]. Anniversary [2623, 2070, 1267, 2640, 1803, 2139, 2421, 1323, 2550, 531, 1056, 4, 293, 349, 553, 1946, 1996, 2062, 2196, 2261, 2304, 2339, 2363, 2392, 1135, 1177, 1226, 1311, 77, 661, 777, 752, 279, 637, 359, 113, 697, 821, 344, 308, 323, 443, 1351, 1405, 1452, 1500, 1573, 801, 23, 1640, 1690, 1752, 1859, 148, 465, 2149, 224, 1899, 2114, 490, 253, 266, 614, 194, 429, 722, 171, 576, 530, 510, 29]. Anno [1523]. Annotated [2807, 1149]. annotations [997, 460]. Annual [2908, 16, 1228, 921]. anonymous [2456]. Ant [2500]. antagonism [2418]. Antarctic [311, 303, 2209, 1499]. Anthologizing [1836]. Anthony [1387]. anthropocentrism [1516]. Anthropological [15]. anthropologist [2046]. anthropology [2302]. Anthropomorphic [1550]. Anti [2338, 1330, 1352, 2137]. anti-aircraft [1330, 1352, 2137]. Anti-reductionism [2338]. antibiotic [2485]. Antimony [2779, 1502, 1502]. Antipodean [1684]. antiquarianism [2414]. Antiquity [2622, 2807, 1138]. antiscientist [767]. antisepsis [2219, 2224]. antiseptic [2221]. antitoxin [131]. Antoine [1763, 2639, 2693, 1356, 1294, 155, 1914]. Antoine-Francois [1914]. Antoni [1760, 1055, 210]. Antony [1275]. any [2580]. apes [2104]. Apothecaries [269]. apothecary [2464]. Appeal [2260]. appearances [2268]. appetite [2595]. apple [211, 233]. Appleton [1375]. application [1908]. Applications [2662, 1189]. Applied [203, 2028, 2403, 966]. applying [2563]. appointment [725, 982, 2208, 610]. appraisal [546]. Appreciation [458, 1400, 659]. apprenticeship [1464]. Approach [2736, 1614]. approved [1170]. Aprahamian [1541]. April [745, 142, 2609, 268, 776, 1757]. Arabella [2383]. Arber [1349]. Arbuthnot [2380]. Arch [2575]. Archaeological [2643, 1269]. Archaeology [2248, 2731, 1707, 2247]. Archbishop [2070]. Archibald [1611, 2634, 1299]. archipelago [1537]. Architect [1107, 1704, 139]. Architectural [1505]. Architecture [2844, 1505]. archive [1463, 1259, 2505, 1574, 1948]. Archives [2177, 869, 2905, 260, 2270, 2317, 295, 1762, 1868, 2329, 170, 1453]. Ardaseer [1171]. areas [1394]. arena [2408]. Argentina [1475]. arguments [2414]. arising [1415]. Aristocratic [1408]. Arkwright [627]. Armillary [2891]. Arms [444, 8]. Army [2506]. Arnott [901]. arranged [416]. arrest [138]. arrives [1852]. Art [2809, 2145, 2853, 2537, 1525, 1040, 2329, 1522]. artefact [2099]. Artefacts [2103]. artful [2078]. Arthur [1064, 1294, 2609, 1419, 1826, 1356, 1864, 2338]. Articles [734, 1288, 262, 275, 291, 305, 331, 340, 358, 367, 441, 452, 474, 487, 498, 508, 517, 528, 540, 551, 570, 588, 601, 619, 634, 646, 657, 671, 683, 704, 718, 733, 747, 761, 772, 785, 799, 807, 818, 829, 842, 852, 861]. artist [2610, 1817]. artistic [2443]. artists [1374]. Arts [186, 2795, 1918, 1129]. Arundel [1393]. Arzamas [2887, 1508]. Arzamas-16 [2887]. Arzamas-16. [1508]. ascent [2176, 288]. Ashbourne [1385]. Ashcroft [1629]. Ashgate [2262, 1785, 1637, 1654, 1767, 1786, 1750]. Ashgate/Variorum [1637]. Ashmole [390]. Ashton [61]. Asiatic [85, 86]. Aspects [1317, 299, 1007, 318, 2925, 2709]. Asperger [1916]. Aspetti [2709, 1379]. aspiring [2822, 1485]. ASS [1615]. Assembly [439]. assessment [1200]. assistance [1932]. Assistant [450, 716, 1150, 567, 280, 237]. Assistants [1930, 1722]. Associated [1186]. associates [2600]. Association [876, 1759, 1317, 162, 2330, 435, 1272, 564, 58, 1547, 45, 1596, 803]. Astbury [891]. astigmatism [548]. Astray [1639]. astrological [1956]. Astronome [213]. Astronomer [1455, 1406, 1689, 2729, 2743, 2774, 1334, 1660, 314, 1495, 1146, 731, 2701, 1332, 1246, 908, 2924, 1099]. astronomers [1986, 1480]. Astronomical [1455, 1380, 1722, 2576, 2572, 1196, 1334]. Astronomy [2646, 2664, 1132, 2553, 2408, 407, 1164, 2500, 2081, 2548, 1189]. Astrophysics [2666, 2198, 1189]. Asylum [2593]. Atena [2544]. Atheism [823]. Athenaeum [2382, 2360]. Athlone [1252, 1473]. Atiyah [1267, 1135]. Atlantic [2466]. atmosphere [2010]. Atmospheres [2665, 1073]. atmospheric [2529, 1519]. atom [1187]. Atomic [1475, 2784, 1174, 2092, 2282]. atómica [1475]. Atómico [2784, 1475]. atomism [2709, 326]. atomists [2282]. Atoms [2660, 2839, 1769, 1488]. attempts [921]. attended [424, 2070]. Attending [2184]. attenuated [2262]. attitudes [2311, 1264, 2380]. Attract [1428]. attraction [596]. Attractions [2696, 1428]. attributable [562]. attributed [583, 756, 2073, 854, 881]. atypical [2332]. Aubrey [738, 710]. audience [1581]. August [2680, 15]. Augustus [1269, 2643, 1882, 1421]. aula [142]. aurora [2526]. Austin [1672, 2282]. Australian [1960]. Author [1389, 1438, 1553, 1618, 1676, 1734, 1788, 469, 1479]. authoritarian [2552]. authority [2195, 2594]. authors [547]. authorship [2494, 2530]. Autobiographical [2873, 1597]. Autobiographies [2902, 1798]. Autobiography [2627, 2901, 1484, 639]. autograph [2475]. autographs [1085]. automatic [561]. autonomy [2361]. autopsic [2329]. Auzout [698]. Avebury [2246]. avoid [1906]. Award [236, 794, 2082, 1716]. awarded [2577]. awarding [2158]. Away [2495, 1394]. Ayres [2151]. Ayrton [1112]. Azarkh [1508].

B [1068, 1293, 637, 1119, 778, 1425, 1265, 902, 1430, 411, 1183, 1455]. B. [178]. B.Sc [758]. b/w [1455]. BA [2039]. Babbage [1594, 1296, 2627, 71, 2200, 1665, 2849, 1109, 1078, 1296, 1664]. Back [10, 52, 242, 332, 342, 348, 369, 541, 589, 603, 620, 636, 672, 685, 705, 720, 735, 750, 762, 774, 903, 914, 924, 934, 942, 954, 964, 976, 983, 992, 1000, 1010, 1022, 1034, 1042, 1058, 1074, 1090, 1104, 1124, 1147, 1166, 1192, 1214, 1237, 1256, 1278, 1301, 1341, 1363, 1390, 1412, 1439, 1460, 1476, 1490, 1510, 1530, 1554, 1575, 1600, 1619, 1641, 1657, 1677, 1691, 1717, 1735, 1753, 1770, 1789, 1804, 1820, 1839, 1861, 1876, 1888, 1900, 1911, 1926, 1938, 1952, 1963, 1976, 1989, 2002, 2012, 2023]. Back [2054, 2083, 2096, 2105, 2116, 2129, 2140, 2153, 2164, 2178, 2187, 2201, 2215, 2227, 2240, 2250, 2264, 2276, 2287, 2595]. Background [2653, 653, 2301, 805]. Bacon [1379, 1206, 2709, 2603, 2785, 2580]. Baconianism [2397]. bacteria [1919]. bacteriophage [2481, 2485, 2484, 2482]. bacteriophage-typing [2481]. bacteriophages [2483, 2480]. bacterium [2048]. Bailey [1138, 2197]. Baillie [2489]. Baker [1198, 755]. Bakerian [1110, 755]. balance [1698]. Balancing [1957]. Balearic [1880]. Ball [385]. Ballads [448]. Balle [1603, 363]. Balliol [654]. balloon [2473]. Baltimore [1119, 1319]. Bang [1549]. Bangs [1671]. Bank [1068, 1178]. Banks [1068, 1293, 675, 1653, 1337, 273, 1487, 1923, 1068, 39, 84, 2878, 768, 2656, 2657, 85, 1444, 2661, 1918, 60, 1369, 2448, 2706, 2707, 741, 757, 2804, 2805, 2806, 2450, 610, 2446, 2449]. Banksian [2447]. Banquet [401, 400]. baptism [793]. Barbados [988]. Barbara [1528]. Barber [1208]. Barkan [1507]. barometer [522]. Barometry [630]. Barrett [1069]. Barrington [997, 2111]. Barrow [1634, 2893, 1324, 1343, 1120]. Bart [273]. Bartle [1317]. Based [2630, 2488]. Baseline [1875, 2411]. Basil [1182, 1211, 1749]. basis [802, 1910]. Basso [1456]. bat [2546]. Batavia [899]. Batchelor [2515]. Bateson [2060, 1264, 234]. Bathyscaphe [2780, 1714]. battery [2007]. Battling [2198]. Baur [1264]. Bausch [1849]. Bawden [707]. Bayes [756]. Bayliss [2075]. Bc [1880]. Bc-Salvador [1880]. be [2463, 2354, 1442]. Beagle [1069, 2683, 1591, 1359, 2910]. Beam [1299]. Beardsley [1708]. bearing [74]. Beast [2594]. Beat [2790, 1362]. Beattie [1639]. Beautiful [2676, 1777]. became [2162]. become [1552]. becoming [1922]. Beddoes [564, 2004, 824, 2005, 2007, 2011]. Bedell [1209, 2786]. Bedoyere [1314]. bee [1240]. bee-house [1240]. Beer [537, 1548]. Beete [837]. Before [2893, 2077, 1724, 2571, 1110, 2548, 2336, 2527, 1120]. beginning [1997]. begins [2472]. beguiling [1872]. behalf [1212]. behaviour [1957]. Behind [2268, 1697]. beings [1181]. Belfast [2389]. Belgium [2369]. Belief [2872, 1353]. Beliefs [2696, 1428]. Believe [1614]. Bell [286, 2560]. Bellenden [1328]. Ben [1922]. Bence [726]. Bence-Jones [726]. Benedetto [2422]. Benedino [1379]. benefactor [1988]. Bengal [85, 86, 1817]. Benjamin [493, 1748, 1443, 1083, 1924, 1151, 1219, 2755, 811, 701, 2773, 724, 12, 231, 562, 1083]. Bennet [1445]. Bennett [1777, 2609]. bequest [34]. Beretta [1763, 1291]. Berlin [1103, 2618]. Bermuda [988, 24]. Bern [1727]. Bernal [2384, 2069, 2472, 1175, 2442, 1541, 1153, 2850]. Bernard [1190, 1706, 563, 1701, 2385, 1146, 315, 922]. Bernheim [2368]. Bertram [758]. Bertrand [1084, 2019, 1063]. Berzelius [151]. Betancourt [2586]. Bethe [1474, 2880]. better [2027]. Between [2368, 782, 980, 1793, 2122, 229, 1347, 856, 676, 2443, 1343, 1880, 2300, 679, 825, 805, 538, 2570]. Beveridge [409]. Beverly [1774, 1463]. Bevis [908]. Bewcastle [364]. Beyond [1061, 2818, 2403, 2588, 1006, 1191, 1436]. Bhabha [2861, 1340]. Bhatnagar [1361, 2912]. Bibliographical [2656, 1068]. Bibliographies [1068, 1097]. Bibliography [262, 275, 291, 305, 331, 340, 358, 367, 441, 452, 474, 487, 498, 508, 517, 528, 540, 551, 570, 588, 601, 619, 634, 646, 657, 671, 683, 704, 718, 733, 734, 747, 761, 772, 785, 799, 807, 818, 829, 842, 852, 861, 2771, 2807, 1097, 2469, 2200]. bibliometric [2016]. Bicentenary [809, 605, 86, 623, 1242, 546, 558]. Biddell [1131, 548]. bifocals [701]. Big [1671, 1549]. bills [833]. Bindman [1255]. biochemistry [940, 1808]. biodiversity [1787]. Biographical [61, 2656, 63, 62, 2807, 2840, 1068, 1400, 1648, 1101]. biographies [2088, 1885, 2438]. Biography [2650, 2654, 1275, 1213, 2808, 2856, 2862, 2494, 1313, 1159, 1087]. Biological [2177, 47, 162, 1997, 268, 1884, 1741, 2798, 2338, 87]. biologie [187]. Biologist [1300, 1205, 1846, 2923]. Biology [2573, 2798, 2089, 187, 1699, 2483, 268, 1538, 2319]. Biopress [1185]. biosphere [2499]. biotechnology [2028, 2535]. Birch [470, 2272, 744]. bird [849]. bird-stuffer [849]. birdman [1209]. birds [2601, 2521, 2786, 2439, 1209]. Birmingham [1525, 301, 546]. Birth [2623, 838, 2675, 1673, 2711, 1704, 511, 501, 723, 1323, 86, 1654, 1056, 494]. Birthday [2885, 616, 130, 1163]. Bishop [1126]. Bits [1594, 1611]. Black [2667, 1189, 2869, 1611, 2624, 2262]. Blacker [1199, 2524]. Blackett [764, 1469, 1483, 1451, 1178]. Blackwell [2385, 1181, 1294, 1211]. Blackwells [1231, 1210]. Blagden [1965, 1909, 185, 1969, 2381, 968, 127, 2424, 2449, 199]. Blair [2100]. Blanc [288]. Blazing [1748]. blights [1919]. Blind [2765, 2766, 1549]. Blink [2674, 1593]. Bliss [1656]. blockade [1698]. blood [462]. Bloomington [1138]. Blows [2744, 1290]. blue [2471, 2009]. blunder [1094]. Bm [1880]. Bm-Sloane [1880]. Board [36, 2111, 2552]. Boas [1070, 1667, 1157, 2302]. boat [1608]. bodies [89]. body [2607, 2458]. Boffin [2646, 1164]. bogus [1582]. Bohr [1187, 2813, 1031]. Bologna [1456, 886, 1622, 1621]. Boltzmann [1488, 2660]. Bomb [1508, 1174]. bombers [1330, 1352]. Bombs [1611]. Bombyx [1262]. bond [1955]. Bondeson [1430]. Bones [2599]. Book [1182, 1209, 1293, 2385, 1272, 1728, 1727, 1612, 1630, 1747, 1764, 1781, 1799, 1338, 2629, 1656, 1456, 1475, 1616, 1088, 1235, 1300, 1317, 1185, 1595, 1071, 1181, 1636, 1146, 1097, 1729, 1410, 1212, 1638, 1572, 1594, 1184, 1251, 1318, 1273, 1455, 1550, 1798, 1145, 1275, 1299, 1528, 978, 1388, 1763, 1381, 1380, 1407, 1406, 1523, 1689, 1542, 1703, 1702, 1765, 1785, 1433, 1411, 1361, 1103, 1508, 1428, 1163, 1673, 1521, 1520, 1592, 1671, 1655, 1637, 1731, 1782, 1766, 1653, 1188, 1100, 1590, 1589, 1639, 1524, 1229, 149, 1654, 1458, 1190, 1591, 1746, 1780, 1121, 1085, 1141, 1180, 1206, 1230, 1292, 1337, 1314]. Book [1429, 1525, 1386, 1297, 1593, 1705, 1457, 1547, 1686, 1685, 1688, 1706, 1387, 1707, 1704, 1213, 1505, 1120, 1231, 1434, 1271, 1332, 1384, 1714, 1613, 1767, 1427, 999, 1160, 1070, 1101, 1208, 1140, 1298, 1313, 1359, 1207, 1233, 1270, 1336, 1506, 1548, 1635, 1730, 1784, 1356, 1098, 1800, 1634, 1687, 1164, 1360, 1335, 1668, 1667, 1252, 1786, 1122, 1138, 1210, 1250, 1294, 1333, 1315, 1355, 1748, 1073, 1099, 1102, 1551, 1334, 1487, 1189, 1633, 1950, 1142, 1709, 1232, 1255, 1504, 1672, 1541, 1629, 1783, 1527, 1529, 1549, 1354, 1353, 1540, 1539, 1357, 1435, 1472, 1471, 1489, 1545, 1615]. Book [1710, 1749, 1611, 1123, 1161, 1143, 1211, 1319, 1598, 1768, 1187, 1617, 1056, 1408, 1432, 1473, 1488, 1546, 1526, 1507, 1596, 1632, 1712, 1769, 1750, 1801, 1087, 1186, 1426, 1425, 1409, 1436, 1486, 1485, 1503, 1502, 1708, 1454, 1453, 1631, 1669, 1670, 1713, 1802, 1086, 1543, 1459, 1597, 1430, 1362, 1162, 1544, 1474, 1732, 1234, 1291, 1358, 1385, 1383, 1276, 1652, 1651, 1711, 1614, 1674, 1431, 1296, 1183, 1522, 1382, 1340, 1253, 1144, 1339, 1295, 1083, 1923, 1254, 2475, 1905, 2175, 1893, 2530]. book-maker [2175]. Books [1182, 734, 1616, 1595, 1639, 1714, 1541, 1290, 1539, 1089, 1123, 1596, 1086, 1385, 1340, 262, 275, 291, 305, 331, 340, 358, 367, 441, 452, 474, 487, 498, 508, 517, 528, 540, 551, 570, 588, 601, 619, 634, 646, 657, 671, 683, 704, 718, 733, 747, 761, 772, 785, 799, 807, 818, 829, 842, 852, 861, 738, 1165, 2016]. Boole [297, 298]. borealis [2526]. Boris [1732]. Born [1675, 1288, 1650, 1643, 2061]. Borneo [2286]. Boscovich [326]. Bose [1421]. Botanic [1293]. botanical [711, 2610, 610]. botanist [1422, 216, 1660]. Botany [2513, 2584, 2833]. both [1919]. Boubée [2590]. bound [2377]. Boundaries [1524, 2366, 2491]. boundary [1467]. bounds [2358]. Bourgeois [2679, 1784]. Boutelou [2547]. Boveri [2064]. Bovier [315]. Bowden [1269]. Bowdich [2293]. Bowler [1211]. Bowring [1317, 2925, 2630]. box [83]. boy [1751]. Boyd [1707, 1111]. Boydell [1455, 1082, 1230, 1314, 1547, 1707, 1613, 1116, 1333, 1453]. Boyle [2124, 381, 1292, 209, 227, 246, 270, 283, 338, 480, 513, 1250, 1425, 1485, 1292, 336, 365, 661, 2685, 462, 534, 446, 1002, 1292, 2459, 337, 2748, 2904, 1604, 1797, 2574, 209, 246, 270, 283, 1250, 1281, 1259, 2822, 2127, 2456, 2748].

Bradley [1919, 1196, 641, 921]. Brae [741]. Bragg [982, 2128, 2852, 1114, 2112, 1825, 1144]. Braikenridge [949]. Brain [2687, 2849, 2017, 1594, 2564, 2588, 1595]. brains [2103]. Bramcote [1508]. breadfruit [2610]. Breaking [1138]. Breeches [2698, 1783]. Bremen [1667]. Brenda [1541, 1702]. Brewer [1613, 2325]. Brian [1141, 1140, 2391, 1235, 1672, 2460]. bridge [792]. bridge-building [792]. Bridgewater [547]. bridging [2018]. Brief [2580, 2273]. brigand [1080]. Brigands [2137]. Bright [1049]. Brighter [1038]. Brimelow [777]. Bringing [1970, 1856]. Bristol [1185, 1689, 1163, 1190, 1705, 1164, 1334, 1546, 1459]. Britain [1592, 1786, 1232, 2195, 1338, 2489, 2360, 2884, 229, 512, 557, 2712, 1095, 2408, 2750, 2907, 1050, 1333, 850, 2275, 967, 2126, 2591, 937, 2361, 812, 2820, 2411, 989, 1473, 1051, 2841, 2847, 2342, 1980, 2874, 1750]. Britanniae [2467]. British [1068, 1235, 1212, 1270, 1506, 1635, 1487, 1522, 1923, 2510, 112, 868, 810, 2640, 1759, 839, 2147, 2436, 2707, 2337, 2274, 857, 1020, 2730, 972, 1330, 1352, 1243, 2410, 940, 2591, 2239, 1916, 1740, 563, 1933, 2330, 2467, 435, 2424]. Broad [1205]. Broad-minded [1205]. broadcast [244, 419]. broadcasting [2410]. Brock [1432]. Brockedon [677, 665]. Brodie [493]. Brompton [577]. bronze [2550]. Brood [2630, 1317]. Brooke [1400, 959]. Brother [2745]. Brothers [2547]. Brouncker [383, 1534, 1535]. Brown [1594, 1671, 1731, 1164, 1802, 1544]. Browne [1313, 1709]. Brownrigg [814, 2074]. Bruce [393, 729]. Bruding [1255]. Brunel [1253, 2859]. Bruno [450]. Brunswick [1138]. Bryan [2282, 1517]. Bryden [1212]. Bubble [2423]. Buckland [791]. Buckley [2383]. Buenos [1475, 2157]. Buffon [2186]. Build [2849, 1594]. building [699, 792]. buildings [2380]. Bulldog [1298]. Bullock [752]. Burchard [461]. Burdon [1007]. Burgess [1551, 2672]. Burkhardt [1069, 1117, 1458, 1359]. Burlington [1637]. Burnet [1205, 2838]. burning [1038, 1578]. business [2615, 1037]. Bust [1268]. busts [617]. butterflies [2467]. Byrom [524, 2727, 1336].

C [1248, 2390, 1782, 1229, 1213, 1332, 716, 1733, 409, 1056, 1597, 387, 613, 2405, 2427, 1386, 2469]. C. [402, 472, 782, 1766, 781, 1935, 2860]. C.B.E [268]. C.H [764]. C.J [1068]. C.K [1300]. C.M.G [758]. C6727 [469]. ca [1965, 1603, 2309, 2490, 1880, 2462, 2307, 2316, 2434]. Cabinet [2537, 2642, 1430, 850]. Caedmon [1318]. Cafaraella [1456]. Cahn [1714]. Cailletet [2236]. Cairo [1670]. Caithness [1527]. calculus [2556, 1263, 2077]. Calcutta [1866, 1863]. Calendar [2670, 1331, 1823, 1069, 1822]. California [1290, 2503, 2500]. call [822]. called [666, 1219]. caloric [1294]. Cambridge [1338, 1088, 1269, 1069, 1117, 1729, 1550, 1798, 1388, 1379, 1501, 795, 1782, 1458, 1591, 1085, 1066, 1206, 1337, 1386, 1457, 1688, 1706, 1387, 1505, 1120, 1332, 1384, 1179, 1101, 1118, 1359, 1784, 1356, 1687, 1138, 1157, 1250, 1099, 1487, 1142, 1232, 1504, 1353, 1357, 1471, 1710, 1768, 1432, 1526, 1507, 1801, 1425, 1409, 1436, 1502, 1631, 1651, 1522, 1382, 1329, 2286, 2886, 232, 1580, 2729, 1332, 2911, 2454, 2792, 248, 457, 1856, 1936, 2864, 1078]. cameralism [2403]. Camm [1188, 2705]. campaign [2582, 1625]. Campbell [628, 1780, 827, 923, 1615, 1296]. Campbell-Kelly [1780, 1296]. Can [2346, 2468, 2580, 1442]. Canada [1639, 1732, 2762, 1102]. Canadian [2490, 423]. cancer [2064, 2332]. Cannell [1252]. Cannizzaro [535]. can’t [1697]. Canterbury [2070]. Canton [1763]. capability [1943]. Cape [1470, 1313, 1336, 1634, 1709]. capital [2430]. Captain [1068, 2745, 605, 622, 898, 608, 606, 609, 607]. captive [2571]. Cardwell [1786, 1087, 1786]. Career [2639, 2752, 1983, 2440, 1196, 1763, 1660, 479, 1704, 2534, 1118, 1761, 1937, 1111, 1893, 2336, 1317, 2925]. Careers [2620, 1138, 2033]. cares [2362]. Cargoes [1589]. Carl [985]. Carlo [1488]. Carlton [1204, 574, 575, 2066, 555]. Carnot [626]. Caroe [999]. Caroline [1798, 1798, 2902, 2742, 1986, 2295]. Carracci [2541]. carriages [1624]. cars [1624]. Carter [1068, 1249]. Cartwright [1504]. Carver [1865]. case [2090, 839, 2223, 676, 2476, 2170, 2444, 1943, 2453, 2371, 2523]. Casimir [637]. Cassiano [1792]. Cassidy [1229]. Cassini [1380]. Castellani [729]. Castelli [2422]. Casting [2443]. Castleford [2324]. Catalogue [2623, 2638, 2891, 170, 995, 1455, 1056, 1291]. Cataloguing [1248, 1779, 1762, 947]. catalyst [1812]. catalysts [1450]. Cataract [1917]. catastrophism [2617]. Catching [1534, 1535]. cathedral [2538, 2070, 417]. Cathode [2682, 1459, 655]. Catholics [2418]. Cauchy [1457, 2843]. cause [150, 149, 863, 2048]. Cautley [874]. Cave [2101]. Cavendish [1449, 1137, 2604, 2757, 260, 1819, 1408, 1812, 2271]. caves [2247]. cavity [1852, 1775]. Cayley [1826, 1646, 1662, 1419, 447]. Cdr [745]. Celebrating [2623, 1056]. Celebration [723, 2640, 2908, 313, 616, 1143]. Celebrations [189, 110, 424, 688, 397, 165, 82, 115, 623, 151, 558]. cell [2135, 2612]. cells [2135, 2612]. Cemetery [2544]. Censoring [2382]. Centenary [1571, 1402, 1568, 254, 2908, 638, 1627, 115, 501, 1131, 151, 1173, 1143, 688, 457]. Centennial [795]. Central [2177, 795, 2400, 50]. Centre [869]. Centuries [2631, 1456, 2809, 2824, 1589, 1493, 2194, 2135, 2341, 1967, 1522, 2513]. Century [2622, 786, 2091, 1814, 2725, 2783, 2807, 1617, 2507, 2270, 2317, 1995, 2576, 2283, 2360, 1482, 2308, 839, 2542, 2553, 2884, 1622, 1971, 792, 2281, 1621, 1794, 1793, 2121, 1428, 1445, 2272, 1520, 2584, 1220, 1942, 2448, 2615, 2557, 1660, 2337, 2714, 2503, 676, 2366, 766, 2731, 2034, 1707, 1343, 582, 2535, 2029, 2400, 2905, 1797, 2594, 1263, 2397, 1506, 2767, 2344, 2914, 2441, 230, 1786, 1108, 1333, 2488, 1132, 1398, 2213, 2033, 2429, 2433, 1830, 2444, 2399, 1860, 834, 2324, 2181, 1632, 2426, 1453, 2570, 1305, 1543, 2396, 1933, 804, 2027, 2424, 2449, 2369, 2453, 1645, 2590]. Century [2510, 886, 2696, 2547, 2709, 1707, 2750, 2388, 1473, 921, 1427, 1138, 1617, 1547, 1070]. ceramic [2488]. Cercignani [1488]. cerebral [1420]. Ceremonies [413]. ceremony [398]. certification [1394]. Cetacean [2431]. Châtelet [1925, 1902, 1581, 1903]. Chadwick [236, 1174, 1247, 1227]. Chaillu [1243]. Chain [2730, 1635]. Chair [2319]. Chairman [1046]. challenge [1833]. challenger [989]. challenging [2194]. chamber [1984]. Chambers [1163, 1653, 1923]. Chameleon [2727, 1336]. Chance [2738, 2774, 1146, 1179]. Chancellor [349]. Chandra [1213, 2862]. Chandrasekhar [1335, 1073, 1189, 1213, 2862]. Chandrasekhara [1740]. Change [2750, 1333, 840]. changed [2350, 1714, 2915, 1749]. Changing [1714, 2489, 802, 1393]. Channel [2316]. Chaotic [1021]. Chapters [2744]. Character [1544, 2388]. Charcot [2368]. charitable [246]. Charles [317, 130, 244, 1965, 1069, 1117, 1594, 2885, 185, 585, 1163, 2616, 1458, 968, 1748, 458, 1925, 131, 1296, 199, 372, 1727, 107, 350, 299, 2627, 355, 2879, 794, 2644, 1069, 2649, 2650, 1729, 2881, 2882, 212, 2683, 1969, 2381, 1591, 1723, 866, 1199, 1810, 2301, 2755, 1313, 2910, 1761, 2088, 1309, 71, 1287, 822, 2501, 825, 1868, 1709, 1774, 858, 2157, 2444, 2263, 2245, 1211, 1902, 769, 1945, 780, 1713, 2849, 1109, 2359, 1664, 2424, 2449, 431, 345, 1850, 1433, 1359]. Charlotte [2587]. Charnley [1103, 2866]. Charter [455, 456, 2051]. Charting [2673, 1429]. Chatham [1708]. Chatto [2151, 1296, 1923]. Chauvinism [963, 932]. checklist [1165, 325]. Cheek [2301]. Chelsea [1197, 1306]. Chemical [203, 2645, 912, 535, 2839, 2540, 795, 960, 2768, 652, 1108, 2488, 2799, 2801, 858, 979, 1432, 1769, 902, 1812, 1955, 115]. Chemist [814, 629, 1386, 1432, 1526, 2176, 987, 1071, 1606, 1696, 959, 1723, 2065, 1122, 1210, 1904, 2831, 448, 859, 1846, 2870, 2325]. Chemistry [567, 2712, 1232, 2799, 2491, 2052, 1920, 873, 815, 405, 2125, 2005, 2087, 2033, 1526, 1856, 2832, 2006, 410, 962, 203, 725, 2768, 2801, 973]. Cherwell [1029]. Chicago [1268, 1528, 1213, 1073, 1189, 1801]. Chicheley [2113]. Chichester [2385, 1749]. chief [1297]. chiefly [856]. Children [1595]. Chile [2426]. China [2192, 1066, 2300, 2768, 2441, 2797, 2798, 2800, 2801, 2492]. Chinese [2510, 2059, 2515, 2600]. Chisholm [1710, 1310, 1930]. choice [1681]. cholera [2146]. Christ [1744, 70]. Christa [1408]. Christchurch [1615]. Christiaan [1053]. Christian [1942]. Christie [820]. Christine [1102]. Christmas [2332]. Christoph [958, 975]. Christopher [561, 1704, 2752, 325, 1205, 2440, 1494, 2589, 597, 2110]. chromatography [1156]. chromosomes [402]. chronological [1165]. Chronology [2462, 2474]. Church [577, 1864, 591, 820, 1744, 707]. chymical [2058, 2316]. chymie [1682]. Ciancio [1331]. Cientificas [189]. ciii [1292]. Cimento [1513]. circa [2574]. Circle [1273, 1138, 2791, 1886, 2584, 1398, 1534]. Circumscribing [1886]. circumstances [2208]. cirripedes [911, 866]. citations [2431, 2124]. City [2676, 1394, 1415, 1189, 1366]. City. [1777]. Civil [568, 2353]. civilisation [1066, 2768, 2797, 2798, 2800, 2801]. civilized [1049]. claim [2092]. Clarendon [1299, 1429, 1122, 1435, 1187]. class [1234]. Classical [1824, 2709, 902]. classico [1379, 2709]. Classics [1433]. Claude [1083, 922]. Clément [626]. clepsydrae [919, 941]. Clerk [749, 838, 884, 2896, 2897, 2898, 2899, 1118, 727, 2915, 1471, 1749]. Clifford [1710, 2669, 998]. Clift [250]. CLII [1379]. Climatology [912]. clinch [2551]. Clinical [2826, 1160]. clock [2589, 1908]. Clockmaker [2677, 1385, 642, 1895]. Clockmaking [1050]. clocks [919, 941]. cloister [2538]. Close [2507, 1584]. closing [795]. clouds [1190]. clover [2398]. Club [616, 1136, 524, 568, 666, 2360, 567, 632]. Clubs [739]. Clyde [1397]. Co [112, 1317, 1594, 2122, 2687, 1229, 1802, 1595, 1811]. Co-evolution [2687, 1595]. Co-operation [112]. Coat [2770, 444, 1802]. coats [2292]. Cock [963]. Cockcroft [580]. coda [2574]. Coffee [1221]. cognition [2606]. Cogwheel [2849, 1594]. Cohen [1616, 1706, 1701]. coin [1956]. Colchester [1182]. Cold [2037, 446]. Coleridge [725, 1286]. Colin [1654, 1207, 1071, 628, 1526]. collaboration [1347, 2409, 2110, 2492]. collaborations [2293, 1534, 1535, 2524]. colleagues [1758]. Collected [2755, 1786, 1068, 1748]. Collecting [1430, 1923, 1971, 2300, 2465, 2466, 2444]. Collection [2638, 833, 2286, 2520, 2541, 2246, 2450, 1272, 2796, 248, 1291, 2524, 1898]. Collections [1212, 2740, 1565, 2517, 2518, 2587]. collective [2545]. collectors [2078]. College [1209, 568, 238, 1766, 1653, 725, 405, 248, 1474, 411, 844, 426, 946, 2208, 445, 480]. Colliding [2667, 1189]. Collins [1068, 1260, 2511]. Colombaria [1379]. Colonel [874]. Colonial [2513, 197, 969, 1661, 1683, 1920]. colonialism [2466]. Colonizing [1834]. Colony [2534]. colour [1826]. colours [1605, 2087]. Columba [591, 820]. combustion [1984]. Come [2728, 1983, 1685]. Comenius [74]. Comes [2402, 1406]. Comet [1883]. Comets [2809, 1019, 1522]. Comfort [2580]. Coming [2651, 1329, 1714]. Comité [1599]. commemorate [312, 1910]. commemorating [2225]. Commemoration [79, 141, 16, 1248, 2717, 219, 1131, 152]. Commemorative [1787, 605, 313]. comment [1945]. commentary [2880, 2920, 1546, 1701, 1474]. comments [973, 2042, 2041, 663, 962]. Commerce [2342, 1218, 2466]. Commitee [92]. Committee [112, 1212, 1959, 40, 147, 1917, 172, 24, 485, 850, 2161]. Common [2668, 1335, 2907, 2468, 1242, 1416, 1750]. Commonplace [978]. Commons [78]. Commonwealth [710, 1661, 1683, 2038, 1672]. Communicating [2754, 1186, 2072]. Communication [1360, 1464, 2349]. Communications [1572, 2379]. Communities [2347]. community [1993, 2583, 2380, 1740, 2310]. Companion [2886, 1706, 2772, 2385]. Companions [2385]. Company [1474, 1276, 416, 2745]. Comparative [1248, 2812]. Comparison [1288, 1264]. comparisons [2570]. compass [816]. Competition [2319]. Compiled [1212, 1334, 1546, 997]. complaints [1415]. Complementary [1870]. complete [1850]. Complex [1457, 2488, 2843, 1757]. Composition [2839, 1324, 1343, 43, 1769]. Compositori [1456]. Compounds [2779, 1502]. comprehensibility [2087]. comprehensive [2467]. Compton [930]. Computer [2849, 1594]. Comune [1380]. Conan [989]. concept [931, 503, 2135, 770, 2047]. conception [1049]. concerned [1089]. concerning [1509]. concise [2469]. conclude [2358]. conduction [405]. conductor [740]. Conference [535, 2923, 100, 134, 2903]. Conferences [165]. confessional [2418]. confident [1253]. Conflicting [2727]. confluence [2338]. Confucius [1844]. confused [1370]. confusions [1533]. Congratulatory [421]. Congress [114, 203, 14, 15]. Congreve [1981]. Connecticut [68]. connection [1312]. connections [1577, 1719, 2509, 2443]. conoid [1395]. Conrad [728]. Conscience [466]. Consejo [189]. consequences [2417]. Conservation [2573, 2502, 2158, 2147, 938, 980, 1787, 2543]. Conservative [2311]. Conserved [1880]. consideration [1701]. considerations [1644]. Consort [445]. conspectus [889]. conspiracy [1526, 2832]. Constable [1525, 1525, 2853]. constant [2212]. constructed [2323]. construction [2575, 2473, 611]. Constructivism [1550, 2713]. Consultant [1483, 1469]. consulting [2195]. consults [434]. consumption [2464]. contacts [1740]. contained [1919]. Contemporary [869, 1762, 1021]. contention [2599]. contested [2610, 2491]. Context [2894, 1471, 2818, 1436, 1721, 1636, 1343, 960, 2417, 2170, 1740, 1920, 1436]. Contexts [2715, 2541, 1379]. contextual [1644]. Continental [1719, 1017]. continua [1324, 1343]. continue [2346]. continuing [2226]. contraction [432]. contribution [931, 2248, 462, 318, 1982, 827, 923, 1420, 611]. Contributions [2640, 566, 792, 2063, 460, 2610, 1583, 1024, 2220, 1040, 230, 909]. Contributors [1263]. control [890]. controversial [2133]. Controversy [653, 2283, 729, 2326, 698, 1290, 2455, 1526, 2832]. convergence [910]. conversation [2597, 2616]. Conversazione [76, 123, 333, 399, 501, 531]. Conversaziones [17, 44, 158, 218, 243, 265, 278, 294, 309, 334, 477, 500, 519, 543, 590, 621, 648, 673, 709, 737, 765, 788, 813, 454, 329]. Convocation [2071]. Conway [1551]. Cook [1068, 622, 1429, 1772, 605, 604, 608, 623, 606, 609, 610, 607]. cookbook [2125]. Cooke [1273, 2787]. Cookman [1593]. Cookson [1654]. cooling [1724, 1725, 1726]. Cooper [1777]. cooperation [1096]. Copenhagen [15]. Copernicus [723]. Copley [271, 794, 1128, 2158, 483, 1895]. copperplate [1068]. Copperplates [2145]. COPUS [2039]. copy [2210, 1149, 1949]. Coram [2505]. Cornish [2528]. cornucopia [1212]. Cornwall [1251]. Corporation [2745]. corpore [597]. corpuscular [1002]. correct [2610]. Correcting [2319]. Correction [2000, 1733, 1925, 548, 923]. Corrections [2001, 1075]. correlation [805]. Correspondence [89, 2630, 2670, 2701, 1331, 968, 1334, 2624, 1464, 1069, 1117, 2881, 1444, 2390, 2262, 1406, 1462, 1689, 1792, 2122, 2272, 1458, 918, 1075, 980, 2906, 1178, 274, 338, 825, 1545, 1791, 560, 2579, 1150]. correspondents [1052]. corresponding [1605]. Corrigenda [571]. Corrigendum [110, 475, 973, 1191, 975]. corrosive [2443]. Cosmologist [2744]. Cotes [2160]. Cotterill [681]. Cotton [2542]. Couch [2729, 1332, 1916]. Coulson [1955]. Council [439, 963, 92, 43, 101, 425, 415, 932]. Count [1443, 2647, 284, 1544]. counterfeiters [482]. counterpart [2408]. countries [90]. Country [814, 1435, 273]. County [415, 967]. course [2476]. Court [950]. covenant [2431]. covers [1273, 1816, 1273]. Cowper [1982]. Cox [2460]. Coxe [469]. Crafting [2323]. Crafts [1450]. craftsman [2256]. Craven [1385]. created [2546]. Creation [2672, 1411, 2684, 1551, 2820, 1592, 1457, 1931, 1080, 2843]. creativity [2233]. Creators [1388]. Creature [2483]. credentials [2336]. Credibility [2312]. credible [2337]. Creed [2781, 1637]. Crichton [1466, 950]. Crick [1123]. cricket [2570]. Crime [2371]. Crimean [2374]. crisis [2472]. criterion [928]. Critical [1410, 2692, 1353, 1811]. criticism [2405]. critics [184]. Croarken [1780]. Crocodile [1377]. Crofton [1039]. Croll [2148]. Cromwell [613, 177, 1891]. Cronica [1475]. Crookes [1917, 2051, 960, 2094]. Croone [389, 432]. Crosbie [1101, 1473]. Crosey [2456]. Crosey-Crucian [2456]. Crosland [1784]. Cross [364]. Crossing [1524, 1467]. crossroads [2586]. Crown [1047]. crucial [2104]. Crucian [2456]. Crucible [2672, 1551]. crucis [595]. Cruickshank [1339]. crust [2454]. crustaceans [1942]. crystal [1081]. crystal-structure [1081]. Crystallographic [1759]. Crystallography [1339, 1609]. crystals [406]. CS765 [1786]. Ctesiphon [2575]. cubes [257]. cubi [2322]. cubico [1395]. cuckoo [742]. Cultural [2636, 2841, 2472, 1548, 2458, 812, 1473]. Culture [2706, 2712, 1232, 2875, 1062, 2537, 1941, 2611, 1933, 1645, 1337, 1688]. cultures [1987, 2303, 2570, 2449]. Cumberland [1613, 2815]. cuneiform [2458]. Cuninghame [2300]. Cunning [1251, 2819]. Curiosities [2642, 1430, 1304, 1430]. curiosity [2384, 1582, 1025, 1645]. curiosity-driven [2384]. Curiosorum} [1849]. Curious [1777, 2753, 2466, 1502, 2384, 2518, 1025, 1958]. Currents [2669, 1710]. curriculum [1078]. Cursetjee [1171]. Cuvier [2293]. Cyril [577, 578, 1410]. Czynski [2371].

D [1293, 349, 1329, 402, 1766, 280, 1175, 1634, 1252, 1541, 1319, 1598, 2850, 1153, 1339]. D. [1572, 1507]. D.C.L [1269]. D.D [1062, 582, 57, 38]. D.J [1212]. D.Mus [792]. D.Sc [492, 268]. D.T [1085]. Dôme [137]. Dadswell [1767]. Dahl [1459]. Daines [997]. dal [1792]. Dale [615, 795, 796, 2136]. Dalitz [1474]. Dallinger [1497]. Dalton [354, 957, 2282, 1681, 1371, 1446, 558]. Daltonian [2282]. Dame [1983]. Damned [1251, 2819, 1183]. damps [2074]. dangerous [2370]. D’Angleterre [150]. Daniel [2510, 1062, 469]. Daniell [725]. Danube [1355]. Dark [2782, 1702]. Darkness [2691, 1766]. Darlington [402]. Darrigol [1711]. Darwin [537, 1437, 347, 352, 1069, 1117, 1729, 1655, 1524, 2710, 1458, 1591, 1221, 1313, 1359, 1709, 1211, 1925, 1358, 333, 350, 2879, 794, 2644, 1069, 2649, 2650, 1729, 2881, 2882, 1307, 1418, 142, 2683, 2889, 506, 2689, 1537, 1655, 2060, 849, 2080, 1458, 1591, 779, 866, 2733, 190, 2301, 1298, 2910, 1027, 1284, 1404, 2761, 1624, 2045, 2088, 911, 939, 825, 1868, 2157, 1892, 2444, 2551, 2263, 2245, 1161, 1806, 1902, 1890, 1945, 2259, 460, 2425, 1348, 345, 536, 584, 1850, 1896, 1906, 1729]. Darwinian [2021]. Darwins [2151, 1417]. data [2432, 2351, 2458, 2478]. database [2152]. Date [838]. dated [583]. Dauben [1411]. Daubeny [858]. David [1616, 1141, 1207, 1210, 820, 1144, 875, 1388, 2607, 1229, 2556, 1386, 1707, 2172, 1149, 649, 1504, 1578]. Davies [1182, 1950, 280]. Davis [1635, 1489, 1425]. Davy [1386, 1210, 970, 484, 2404, 2448, 2763, 2764, 2007, 448, 1399, 1980]. Dawkins [1111]. Dawn [1409, 2836, 2151]. Dawson [1110]. Day [1113, 690, 1546]. Days [599, 2646, 511, 1164]. DCL [2643]. Deacon [1595]. Deadly [2632, 2633, 1708]. Dealing [734, 262, 275, 291, 305, 331, 340, 358, 367, 441, 452, 474, 487, 498, 508, 517, 528, 540, 551, 570, 588, 601, 619, 634, 646, 657, 671, 683, 704, 718, 733, 747, 761, 772, 785, 799, 807, 818, 829, 842, 852, 861]. dealt [1919]. Dean [1982]. Death [223, 1418, 2443, 2315, 2505]. Debate [1338, 2718, 2874, 2368, 1501]. debt [427, 2182]. Debunking [1425]. decay [2531]. deceitful [2009]. December [349, 1996, 2304, 637, 92, 1405, 1752, 2562]. Decided [2788, 1145]. decision [1584]. Decisive [2809, 1522]. decline [2109]. Deconstructing [1857]. decoration [575]. Dedication [202, 1832, 12]. deep [2578, 2101]. defeat [1823]. Defence [2538, 2606, 1080, 147]. Defining [1338, 2283, 2756, 2874, 2484, 1338, 1596]. definitive [2088]. Degree [413]. Dekker [1633]. del [1513, 1475, 1380, 2444]. delegation [669, 320]. delegations [538]. Delhi [114, 1748]. Delisle [1579]. delivered [402, 403, 405, 753, 409, 411]. della [1523]. dell’Atomismo [1379, 2709]. Demarcating [2366]. demise [1398]. democracy [2035, 2005]. demographic [2472]. demolition [2538]. Demon [1319]. demonstration [2416]. Denis [1295, 118]. Denmark [2400]. Dennis [1350]. Dental [2091]. dentistry [2360]. department [2432, 1808]. depictions [2571]. Deprogramming [2397]. derangement [864]. Derby [1385, 2677]. Derbyshire [1385, 1385]. Derek [1585, 2256, 1099]. dermatological [2458]. Desaguliers [2476, 2380]. Descartes [183, 503, 2606, 2315]. described [1060]. description [1168]. descriptive [995]. desiderata [2397]. Design [2738, 1179, 676, 404, 1624]. Designer [1188, 2705, 549]. designing [2110]. designs [2578]. Desmond [1298, 1161, 1729, 1524]. despatch [2406]. destruction [1757]. details [1151, 2589]. detection [2159]. determination [1081, 891]. detractors [2533]. Deutsch [1234, 1183]. developing [2307, 891]. Development [1361, 2912, 1995, 973, 1917, 2884, 2063, 1741, 2598, 2406, 1468, 1786, 2247, 1519, 2619, 642, 1943, 1825, 962, 1078]. deviation [816]. Devil [2690, 1298]. devils [2009, 2502]. Devons [1934]. Devonshire [1317]. Dhar [2390]. diagnosis [1892, 1945]. Dial [2511]. Dialectical [2061]. dialogue [2488, 1955, 2822]. diamagnetism [910]. diamonds [2053]. diaphragms [2612]. Diaries [2172, 1315, 2803, 524, 1604]. diary [1422, 2683, 185, 1498, 2424, 1069]. Dictionaries [2875, 1688]. Dictionary [1388, 2792, 2807]. did [2358, 822, 2551, 1906, 1506]. Difference [1109]. differences [970]. Different [2790, 1362]. differential [1561]. difficult [1306]. Diffraction [2680, 1061, 1339]. Diffusion [786, 766]. Digby [388]. digitization [2523]. dim [1261]. Dimension [810]. Dimensionality [2508]. Diment [1068]. Dingley [1107]. Dinner [349, 638, 490, 4, 293, 349, 553, 77, 661, 777, 752, 279, 637, 359, 113, 697, 821, 344, 308, 323, 443, 801, 23, 616, 148, 465, 224, 253, 266, 614, 194, 429, 722, 171, 576, 530, 510, 29]. diphteria [131]. Dirac [1139]. director [2577]. Directory [1212, 2740].

Dirty [2331]. discharge [1891]. Disciple [2690, 1298]. discourse [2370]. discovered [1604, 1259]. discoverer [1723]. Discoveries [1268, 2556, 1247, 1330, 1352, 2913, 1081, 1178, 1136, 1227, 1312, 1712, 1155, 1114]. Discovering [2169]. discovers [971]. Discovery [2678, 1411, 2686, 2514, 2824, 2829, 1060, 2283, 1036, 2158, 1421, 1589, 2280, 1760, 960, 515, 548, 1857, 2001, 1985, 31, 2020, 1489, 1123, 1806, 2439, 1956, 1916, 981, 1384, 1123]. discussion [910, 2042, 2041]. disease [2270, 2489, 2048, 909]. diseases [230]. disenchantment [2551]. dispensary [986]. Dispenser [1180]. displaced [2539]. Displacing [1769]. display [2541]. Displayed [1355]. Displaying [2357]. disposal [246]. dispute [2417]. disputed [2526]. disputes [1415]. disruptive [2351]. dissemination [2335]. dissension [1275]. Dissensions [2359]. Dissent [2623, 2019, 1056]. dissenter [2456]. Dissenting [1818]. dissolute [2009]. Dissolving [1275]. distance [1401]. distinction [1684]. Distinctions [695]. Distinguished [1431]. distribution [247]. Distrust [2346]. Dive [1330]. Dive-bombers [1330]. divergence [910]. Diversifying [2480]. Diversions [1944]. Divine [2035]. division [2576, 1310]. Dixon [212]. Djerassi [1766]. DMZ [2499]. DNA [1702, 2782, 1598, 2865]. Do [2021, 2343, 2331]. Dobell [998]. Dobinson [940]. Dobson [1794]. doctor [1077, 1418]. documentary [1111]. documents [2093, 2317, 1259]. Does [2033]. Dogramaci [2598]. Doing [2172]. Domb [1410]. Døme [137]. Domenico [1380]. domestic [2361, 2543]. dominant [1255]. Donald [2793, 1786, 1087]. donation [1393]. done [2489]. Donovan [1356, 1294]. don’t [1882]. Dorchester [349, 490, 661, 777, 752, 637, 821, 308, 722]. Dordrecht [1636]. Dorinda [1138]. Doron [1594]. Dorothy [1312, 2597, 916, 2699, 1539]. dots [1755]. Double [2511, 2388]. doubts [2282]. Douglas [586]. down [2382, 1359, 2246]. Doyle [989]. Dr [1185, 1188, 1497, 1802, 1162, 1077, 450, 259, 814, 1318, 375, 821, 1751, 1129, 1221, 1881, 1910, 2770, 824, 2005, 899, 1832, 2329, 160, 1865]. Dr. [2489, 2695, 2793, 2825, 2834]. draft [559, 1265]. dragons [2380]. drain [2017]. Drake [1383]. Dramatic [2691, 1766]. Draper [2330]. drawer [1209]. drawing [2073, 2589]. Drawings [2207, 2329, 677]. dreams [2211]. driven [2384]. Dropped [1709]. Drops [1008, 994]. drugs [2464]. Drum [2790, 1362]. Drummond [1287]. Dryden [468]. dualism [2606]. Dublin [2092]. Duclos [2316]. Due [2631, 1456]. Duillier [149]. Duke [217, 324]. Dumaresq [1062]. Dumb [1506]. Duncan [1458]. Durham [1502, 1982]. During [2513, 2526, 2510, 2576, 1219, 2559, 2430, 1698, 2592, 2492]. Dust [1685]. Dutch [1384, 2124, 2477, 1052, 2829, 1873]. dye [2325]. dynamic [2439]. Dynamical [2680, 1339]. Dynamicists [2515]. dynamics [1335, 1021].

E. [1329, 2122, 2762, 730, 1489, 1080, 1102]. Earl [1271, 2827, 393, 1023, 285, 2827, 1037]. Earlier [2816, 1244, 1542]. earliest [173, 1149]. Earls [1842]. Early [1338, 2646, 1384, 907, 2747, 2749, 2779, 2512, 956, 55, 2867, 2874, 1739, 511, 183, 1464, 2283, 931, 2248, 832, 1071, 2281, 1794, 2063, 492, 512, 1347, 2509, 2406, 2584, 2448, 2337, 905, 1024, 2503, 2556, 1082, 2072, 2452, 1095, 2206, 2443, 1699, 2535, 2595, 2220, 2534, 802, 2198, 960, 1118, 1586, 2594, 1263, 195, 2123, 2088, 72, 1164, 757, 1126, 2464, 855, 927, 966, 1885, 2611, 1974, 2299, 2455, 160, 2239, 1416, 2831, 1502, 184, 176, 2136, 117, 258, 2543, 1375, 362, 2453, 2602, 2492, 990]. Earth [2248, 1248, 1950, 1768, 2734, 1897, 1610, 1685, 2020, 2493, 2854, 1648]. Earthly [2694, 1383]. Earthquake [2549, 2527, 1791, 1795]. East [2745, 1066, 2441]. Eastern [1361]. eccentric [1606]. Echoes [2775, 1190]. Eckersley [1178]. eclipse [2533]. Ecology [159]. economics [2615, 2319]. economies [2400, 2396]. économistes [2583]. economy [2239]. ecosystem [2502]. ed [1068, 1069, 1117, 1455, 1591, 1119, 1453, 1296, 1083]. Eddington [1883]. Edinburgh [217, 2593, 805, 1502, 1947, 2301, 2593, 631, 664]. EDITA [1712]. Edited [1182, 1293, 1300, 1317, 1636, 1248, 1433, 1508, 1163, 1653, 1188, 1141, 1292, 1331, 1705, 1547, 1706, 1387, 1120, 1434, 1208, 1359, 1506, 1360, 1786, 1116, 1138, 1250, 1541, 1545, 1143, 1228, 1750, 1425, 1631, 1162, 1144, 1339, 1295, 1314, 1207, 1334, 1255, 1056]. Editing [1663]. edition [2477, 446, 1270, 247]. Editions [1068, 1589, 1599]. Editor [716, 64, 1729, 2144, 1322, 18]. Editorial [687, 809, 894, 2026, 1928, 1940, 1954, 1966, 2000, 1979, 1991, 2015, 2044, 2057, 2068, 2086, 2108, 2119, 2132, 2168, 2143, 2156, 1320, 2298, 2314, 2321, 2328, 2334, 2356, 1838, 1791, 2505, 2474]. editors [1978, 754]. editorship [659]. Editrice [1456]. Edmond [898, 739, 640, 312, 1239, 978, 1106, 2673, 1610, 1869, 1429]. Edmund [313, 1370]. Edna [1655]. eds [1293, 2262, 1782, 1458, 1780, 1276]. Education [2547, 2552, 722, 319, 1376, 1566, 826, 1063, 2400, 825, 1625, 268, 2136]. educator [930]. Edward [1071, 875, 1379, 1309, 1480, 2831, 1425, 1375, 157, 161, 815, 2700, 2432, 1384, 1242, 2170, 1370, 1526, 2832, 742, 712, 650, 1183, 2453]. Edwardian [1528, 2708]. Edwards [1209, 1517, 2786]. Edwin [1185, 2695]. effect [1402, 1036, 971, 1498, 952, 680]. effects [302]. Egg [2810, 1470]. Egyptologist [2616]. Eighteenth [2725, 2731, 834, 2308, 2542, 1622, 1971, 1621, 1793, 1428, 1445, 2272, 1220, 1942, 1660, 676, 1707, 582, 2400, 2397, 1506, 2914, 230, 1108, 2399, 1305, 1543, 2396, 1967, 2424, 2449, 1645, 2696, 1707, 2513, 921]. Eighteenth-century [2731, 2542, 1971, 1428, 1445, 1942, 1660, 1707, 582, 2400, 230, 2399, 1543, 2396, 2449, 2696, 1707]. eighth [1228]. Eightieth [2885, 1163]. Einstein [1249, 2888, 2407, 2735, 2275, 1627, 1289, 1822, 1824, 2495, 1711]. Eire [1212, 2740]. elaboratory [2537]. Elder [2745]. Eleanor [1273]. Elected [1597]. Election [217, 2133, 178, 289, 37, 267, 430]. electric [492]. Electrical [2195, 2675, 1545, 1400, 795, 1654, 756, 1179, 2008, 1891]. electrician [1445, 1924]. electricity [1421, 2532]. electrochemistry [434]. Electrodynamics [2888, 1711, 1711]. electromagnetic [1200, 2531, 525, 1172]. Electron [2682, 2686, 1459, 1995, 1741, 770, 2619, 952, 1489, 1459]. electron-releasing [952]. electronic [1329, 1830, 1574]. Elements [2839, 2863, 1769, 1208]. Elephant [1377]. Elias [390]. Eliot [1079]. elite [2147]. elixir [1347]. Elizabeth [307, 122]. Elizabethan [1973]. Ellen [1383]. Elliott [1293]. Ellis [834, 1287]. Elly [1633]. Eloge [2609]. Elsevier [1714]. elsewhere [1358]. Elton [2501]. Embrace [2708, 1528]. embryo [891]. emergence [2135, 2236]. Emigration [2600]. Émilie [1925, 1902]. eminently [1273]. Emma [1655, 2733]. Emperor [1653, 2238]. Empire [100, 2506, 2707, 2513, 1101, 1487, 2840, 2845, 1142, 2444, 2447, 1248, 1661, 1683, 1923]. employment [1219]. Encounter [2636, 679, 1548]. encounters [2599]. encyclopaediae [2403]. Encyclopaedic [2875, 1688]. End [1689, 2308, 1532, 1794, 2008, 823]. endangered [2502]. Endeavour} [1547, 1285, 2496, 610]. Enderby [1163]. energia [1475]. energies [409]. Energy [1475, 1473, 2840, 2841, 2158, 2147, 938, 980, 1698, 1101]. engagement [2302]. Engaging [2353]. Engine [1109]. Engineer [2675, 1654, 2375, 1400, 874, 2065, 625, 1253]. Engineering [2675, 1959, 2324, 2859, 1253, 1295, 839, 403, 427, 2575, 549, 1654]. Engineers [2688, 1545, 2549, 1295, 1374, 568]. engines [663]. England [1613, 1999, 2376, 1696, 1428, 1445, 2696, 1520, 1637, 2714, 826, 676, 2731, 1707, 1844, 1263, 632, 2781, 2307, 2815, 1526, 2832, 131, 2543, 2507]. English [1182, 2631, 1456, 1185, 1337, 1651, 296, 1165, 2706, 2008, 2814, 2016, 2867, 2183]. Englishman [1969]. engraved [2169]. engravings [1068]. enhanced [2063, 2001, 1985, 2162]. Enigmatic [1702]. Enlarging [2358]. Enlightened [2581]. enlightening [1337]. Enlightenment [1592, 2712, 1337, 1688, 1232, 2875, 2460, 1930, 1592, 1018, 1592, 2698, 2706, 2008, 1783, 2399, 2820]. enough [1049]. enquiries [2879]. Enquiry [2685, 1425]. Enright [1141]. Ent [1792]. Entail [1371]. enterprise [2374, 2582]. enterprises [2429]. Enterprising [1528]. entitled [1919]. entomological [2500]. Entrepreneurial [1432]. environment [1919]. environments [2502, 2498]. envisaged [1823]. epic [2383, 2799]. epidemic [2146]. epigrams [2475]. Epilogue [394, 356]. episode [966]. Episodes [2810, 1470]. epistemic [2489]. Equations [2846, 1746]. Equatorial [642]. era [1420, 2447]. Erasmus [1437, 1729, 1418, 1221, 347, 352, 2889, 1524, 1027, 1284, 1404, 2761, 1624]. Eric [1334, 1406, 1689, 1802]. Ernest [599, 1098, 695, 1698]. Errant [2859, 1253]. Errata [749, 786, 1675, 463, 941, 1748]. Erratum [537, 1437, 1008, 2128]. Error [2621]. errors [1299, 2548]. Erwin [1264]. erythematosus [1348]. Erzincan [2549]. eschewing [1054]. Essai [1682]. Essay [1249, 1269, 1069, 1117, 1248, 2262, 1379, 1501, 1777, 1119, 1082, 1066, 1331, 1470, 1096, 1588, 2082, 1179, 1118, 1484, 1644, 1116, 1157, 1290, 1205, 1228, 1701, 652]. Essayes [2049]. Essays [1542, 2722, 1316, 346, 2816, 2834, 351, 1162]. Essence [1433]. Est [503]. establish [2582]. established [2311, 2491]. Establishing [2747, 1082]. establishment [1128, 2494, 824]. estate [246]. estrangeirado [2257].

Estudio [1475]. eternity [2234]. Ethics [1353, 2872]. Ethnological [15]. Ethologist [1767]. Eton [336, 365]. Ettehadieh [2488]. Euclid [1710]. Eugenics [1673, 2908, 1228, 2090, 1264, 1403, 1423, 2711, 1673]. Eunice [2457, 2529]. EUR [1712]. Eurobook [1207]. Europe [1207, 1355, 2537, 2037, 1566, 2895, 1924, 2400, 2905, 2365, 2848, 1496, 1453]. European [1547, 1585, 2558, 1831, 2914, 2552, 2488, 2103]. Evaluating [2476]. Evangelist [1186]. Evelyn [391, 1314, 174, 2892, 531]. Evensong [2070]. event [1519]. Eventful [1251, 2819]. events [970, 2193]. ever [1518]. Everest [1130, 1527, 2842]. everything [2915, 1749]. evidence [1239, 2594, 1284, 2104, 1217, 1111]. evidenced [2448]. Evolution [2704, 1211, 1631, 1595, 2687, 404, 1313, 1404, 2484, 2045, 2246, 1890, 1913, 1016, 1437, 1638]. evolutionary [2383, 2052, 2908, 1143]. evolutionist [2088, 2103]. Evolve [1595]. evolves [1117]. evolving [2383]. Ewald [2680, 1339, 2680]. Ewing [867]. examination [2681, 854, 1767]. examples [2123]. Excavator [1269]. exceedingly [2502]. excellence} [1213]. Excellency [361]. exception [2040]. Exchange [2905, 538, 2126, 1453]. Exchanged [1880]. exchanges [2591]. Exchequer [349]. Excursions [1446]. Executive [92]. Exeter [268]. Exhibition [2623, 312, 578, 1056]. exhibitions [418, 2429]. existed [1518]. Exodus [1616]. exotic [1565]. expansive [626]. Expectations [2396, 2398]. Expedition [2635, 2674, 303, 311, 1639, 1593]. expeditionary [2557]. expeditions [47, 2034, 2429]. Experience [2747, 839, 1082, 137]. experiences [2292]. Experiment [2719, 2720, 1157, 2557, 2595, 137, 1811, 1494, 2361, 2006, 655, 1181]. experimenta [2317]. Experimental [2720, 1157, 2220, 2476]. experimentalism [1513]. Experimentation [2049, 2399, 2337]. Experiments [2049, 1819, 1724, 1725, 1726, 1181, 2528, 896, 2529, 2500, 909]. Experimentum [595]. expert [2293]. expertise [2195, 2598, 2346]. Experts [2549]. Explaining [1674, 2111]. explanation [596]. Explicit [928]. Exploration [1673, 2711, 2824, 2845, 2436, 1547, 2914, 1142, 1416, 2592, 1589]. Exploring [2199, 2072]. explosives [2260]. extension [1039]. Extinguished [2540]. Extraordinary [640, 2647, 1269, 1318, 1122, 1544, 2870, 1122]. Extremes [2625, 1629]. eyes [2246, 2565]. Eyewitness [1638, 2704, 2527].

F [1572, 725, 996, 1195, 409, 1459]. F. [1626, 1733]. F.R.C.S [1400]. F.R.S [1269, 1400, 1097, 1638, 1572, 1251, 491, 1163, 1188, 1331, 975, 701, 68, 1228, 393]. F.R.S. [401, 579, 615, 1077, 32, 236, 307, 622, 995, 987, 1062, 1107, 1218, 1325, 1466, 1447, 385, 271, 287, 299, 213, 883, 930, 347, 352, 837, 901, 814, 561, 640, 988, 1517, 387, 711, 1288, 986, 833, 1185, 875, 874, 1328, 612, 628, 738, 564, 613, 1131, 758, 792, 1039, 2670, 675, 726, 1606, 375, 629, 898, 1518, 873, 188, 108, 140, 391, 492, 1007, 791, 1445, 469, 2695, 86, 715, 815, 1100, 900, 731, 997, 710, 381, 388, 177, 743, 912, 959, 1583, 1607, 353, 676, 1199, 470, 1127, 577, 1497, 968, 764, 831, 582, 756]. F.R.S.

[481, 659, 1282, 1373, 1175, 708, 390, 958, 1221, 1484, 598, 493, 2771, 515, 548, 701, 778, 644, 445, 1309, 209, 227, 246, 270, 283, 338, 386, 480, 513, 1533, 889, 392, 616, 449, 315, 324, 155, 1169, 940, 1092, 770, 858, 1349, 364, 346, 906, 389, 979, 834, 989, 122, 505, 384, 769, 374, 742, 560, 639, 642, 1647, 1479, 859, 1261, 700, 887, 563, 1130, 580, 377, 1153, 950, 1076, 57, 436, 776, 662, 680, 55, 562, 611, 755, 354, 981, 1111, 2856, 650, 1283, 908, 1403, 1423, 326, 677, 665, 909, 38]. F.R.S.E. [586]. F.S.A. [1269, 1400]. Faber [1249, 1730]. Fabri [1061]. fabrication [566]. Faces [2168, 1635, 2167]. facets [2413]. facial [2566]. Facilities [147, 238]. Facsimiles [1085]. Fact [2688, 1295]. facts [479]. Fading [1656]. failure [2387, 2391]. failures [2359]. Fairchild [59]. fairy [2332]. Faith [1636]. Faithful [2206]. Falconer [1489]. fall [1297, 2481]. Fallodon [668]. Fame [2539]. Fames [1138, 1276]. familiar [2361]. Family [1675, 2630, 1983, 1650, 60, 1986, 1374, 1217, 1153]. famous [558]. Fantasy [2688, 1295]. far [2441]. Fara [1428, 1783, 1669]. Faraday [526, 1546, 1200, 1308, 2890, 691, 910, 971, 2906, 1663, 1545, 434, 848, 1546, 1665, 1158, 1274, 1399, 714, 1545]. Farewell [1320]. Farmer [2224]. Farnham [2262]. Farquharson [1579]. Farrand [1140]. fasc [1523]. Fate [2635, 2674, 2111, 1639, 1593, 1393, 2161]. Father [1385, 1646, 1662, 967, 447]. Fathers [1727]. fatigue [961]. Fatio [149]. faunal [2411]. Faust [1319, 2732]. Favourite [1317]. Featherstonhaugh [287]. Features [2756, 575, 2483, 1596]. February [2210, 820]. feeling [2553]. feet [1119, 2559]. Feingold [1120]. Felicitation [2834, 1162]. Fellow [1251, 1171, 1910, 2819, 1603]. Fellows [1270, 262, 275, 291, 305, 331, 340, 358, 2070, 1225, 2091, 832, 566, 173, 229, 512, 1347, 1760, 427, 2029, 2749, 1270, 1266, 1745, 822, 1625, 197, 272, 471, 677]. Fellowship [80, 2071, 1661, 1683, 9, 43, 1133]. female [2582, 2588, 2581]. Feminist [2588]. femmes [2583]. Femtoworld [1670]. Fennica [13]. Ferdinando [1355, 2848]. Ferguson [1495]. fermentation [2222]. fern [2286, 2468]. Ferranti [492]. Ferry [1251, 1539]. fertilization [1262]. Festival [2071]. fevers [2369]. février [149]. Feynman [2790, 1362]. FF.R.S. [245]. fibres [2612]. field [1610, 676, 2046]. Fielding [583]. Fields [2636, 1548]. Fighter [1188, 2705]. figural [2073]. figurations [2073]. Figure [1356, 2493, 1648]. Filling [1271]. film [1605]. films [416]. Filosofia [1379, 2709]. fin [2368, 2371, 1887]. fin-de-siècle [2368, 2371, 1887]. final [1915, 2042, 2041, 1892, 1647, 1945]. finances [335, 20, 35]. financial [783, 2423]. find [2468]. Fine [1270]. Finest [2705, 1188]. Fire [2809, 2110, 2074, 1366, 1394, 1415, 2676, 1777, 1522]. fire-damps [2074]. Firenze [1379, 1523]. Firmly [1614]. First [2701, 1767, 400, 1334, 1627, 2849, 1234, 1068, 1884, 484, 1594, 1918, 1455, 1406, 1689, 1366, 2477, 1960, 731, 2210, 19, 455, 535, 1434, 1797, 1742, 1171, 34, 1154, 2917, 2530, 247, 2577, 2237, 1919, 2048, 131, 2341, 2473, 2467, 2027, 2924, 1253, 1645, 288, 2526, 1150]. First-class [1234]. Fish [2507]. Fisher [537, 1410, 1183, 506, 1223, 805, 1265, 2319, 536]. Fit [2121, 2307]. FitzGerald [2296]. Fitzroy [1739, 2197]. Five [1757]. fixation [1424].

Flamsteed [1455, 1406, 1689, 731, 1334, 2701, 1334, 2924, 1455]. Flash [2682, 1459]. Flat [2569]. Flattening [1897]. Fleck [591]. Fled [2789, 1616]. Fleeming [2675, 1654]. fleet [1794]. Fleming [281]. Fletcher [2018, 2260]. flights [136]. Floating [2810, 1470]. Flood [1780]. Florence [1523, 2560]. Florey [579, 1802, 1072, 1468, 2770]. Florida [834]. Florilegium [1068]. flourish [2545]. flourishes [2441]. flower [1068]. flowers [48]. Fluid [2515, 919, 941]. fluids [2913, 1712]. fluxionum [2556]. flying [1352]. Foam [2869, 1611]. Focus [1275, 2856]. Foes [1316]. Folio [1878, 1879]. Folkes [274, 2505]. Follow [2432]. Fontenelle [315]. food [2452]. Foote [2457, 2529]. footnote [522]. footsteps [1537]. Forbes [1406, 1689, 1334]. Force [1636, 2147, 1815]. Forces [1428]. Ford [1235, 1140, 1611]. Fordyce [1606]. Foreign [2596, 1478, 1150, 1435, 88, 89, 1462, 209, 270]. forensic [959]. Foreword [1410, 396]. forgotten [951, 2481, 2482, 2441, 235, 327, 1536]. form [2169, 891]. Formal [574, 397, 398]. Formation [2905, 1453]. formative [2032]. Former [554]. Forms [2343]. formula [1172]. formulae [2317]. Fortey [1638]. fortunes [802]. fossil [2122]. fossils [1560, 2207]. Foster [491, 1496]. Fothergill [899]. Foucault [1768, 1768, 2854]. found [2270, 2317, 2122]. Foundation [1208, 1162, 857, 972, 371, 7, 1910]. Foundations [1379, 2715]. founded [917]. Founder [755, 86]. Founders [512]. Founding [1727, 1721, 2150, 1866, 865, 445, 435]. Fountaine [2294]. Four [696, 1826]. four-colour [1826]. Fourcroy [328]. Fourier [1851, 1081]. Fourth [2827, 2908]. Fowler [951, 1173]. Fox [1269, 2643, 2528]. Fozard [1188]. fractographic [1346]. Fragment [2748, 2318, 1292, 6]. frameworks [1039]. francaises [187]. France [229, 1915, 2366, 1208, 2611, 2590]. Frances [1455, 1406, 1689, 1334, 1230, 1629]. Francis [1410, 1379, 1673, 2711, 2909, 1541, 1489, 1123, 1228, 61, 2171, 2709, 2603, 1206, 1280, 2534, 2163, 2785, 2184, 2580]. Francois [2490, 1914]. Frank [2885, 1163, 1545, 1750, 1457]. Frankland [1526, 1071, 2831, 2832]. Franklin [1702, 1639, 1593, 1716, 701, 1708, 236, 2632, 2633, 2635, 1702, 2674, 1951, 1924, 811, 2773, 2782, 1708, 12, 231, 740, 1083]. Franz [1733, 1700]. FRCP [1866]. Fred [1290]. Frederick [1069, 1117, 1275, 1458, 707, 1359, 457, 95, 2630, 931, 1127, 235, 2856]. Free [973, 2685, 1548, 1425, 2619, 962]. free-electron [2619]. freedoms [2591]. Freeman [1229]. Freind [1925, 1902, 1904]. French [73, 54, 187, 229, 137, 256, 1975, 2530, 2473, 1967, 1581]. Frequency [1172, 1378]. frequent [1441]. Frescoes [1863, 1864]. Fresh [957, 2507, 1284]. freshened [227]. Freud [2302]. Friday [2421]. Friedel [1450]. Friend [2466, 1810, 1904]. Friends [271, 1292, 2748, 1316, 2857, 2574, 1730, 714]. Friendship [1972, 644, 642]. friendships [2038]. fringe [1061]. frogs [2185]. Front [1, 25, 26, 97, 146, 193, 208, 222, 252, 277, 306, 343, 395, 499, 518, 573, 647, 660, 686, 706, 721, 736, 751, 763, 775, 808, 871, 879, 885, 893, 904, 915, 925, 935, 943, 955, 965, 977, 984, 993, 1001, 1011, 1012, 1035, 1043, 1044, 1059, 1091, 1105, 1125, 1148, 1167, 1193, 1215, 1238, 1257, 1279, 1302, 1321, 1364, 1391, 1413, 1440, 1461, 1477, 1491, 1511, 1531, 1555, 1576, 1601, 1620, 1642, 1658, 1678, 1692, 1718, 1736, 1754, 1756, 1790, 1805, 1821, 1840, 1862, 1877, 1889, 1901]. Front [1912, 1927, 1939, 1953, 1964, 1977, 1990, 2003, 2014, 2025, 2056, 2085, 2098, 2107, 2118, 2131, 2142, 2155, 2166, 2180, 2189, 2203, 2217, 2229, 2242, 2252, 2266, 2278, 2289]. frontiers [2038]. Frontispiece [2013, 2024, 2055, 2067, 2043, 2084, 2097, 2106, 2117, 2130, 2141, 2154, 2165, 2179, 2188, 2202, 2216, 2228, 2241, 2251, 2265, 2277, 2288, 2297, 2305, 2313, 2320, 2327, 2333, 2340, 2355, 2364, 2372, 2378, 2386, 2412, 2394, 2395, 2401, 2420, 2428, 2437, 2445, 2479, 2451, 2461, 2470, 2516, 2487, 2497, 2504, 2536, 2562, 2525, 1326]. Frozen [2635, 1639]. FRS [2643, 1965, 1983, 2091, 1907, 1807, 2124, 2885, 2390, 1696, 1552, 1794, 1866, 1826, 1942, 1723, 2546, 1882, 2745, 2300, 2214, 2909, 1910, 1761, 1864, 1640, 1690, 1752, 1803, 1859, 1774, 1891, 2819, 1772, 1899, 1958, 1919, 1708, 1827, 1740, 1680, 2325]. Fruitful [2524]. fruits [1742]. Fryer [856]. FSA [2643]. Fuego [2444]. Full [1943, 1068]. Full-text [1943]. Function [1457, 2843]. functions [2343]. Fund [172]. fundamental [840, 1536]. Fundator [372]. Funding [2053, 2552]. fungi [2535]. Further [581, 918, 651, 612]. Fusion [2853, 1525]. Future [1730, 2857, 1609, 2352, 1833, 1737, 1841, 1757, 1730]. futures [1587].

G [1293, 963, 637, 2262, 1299, 1406, 1689, 1163, 1096, 1384, 1714, 938, 778, 1138, 1334, 825, 1143, 1056, 1276, 891, 1340]. G. [287, 1323, 951, 1671, 1607, 938]. G.B.E [615, 436]. G.I. [1544]. Gabor [1350]. Gabriel [2087]. Gaddum [1921]. Gagarin [438]. Galápagos [1537]. Galenization [2490]. Galenizing [2490]. Galilean [1345]. Galileo [502, 501, 2422, 1873, 2134]. Gallery [1635, 2919, 1596, 1632, 2537, 2541, 604]. Gallica [664]. Galton [1673, 2909, 1228, 2711]. galvanism [2008]. gap [921, 1906]. Garden [2376, 2503, 1197]. Gardening [2547, 1919]. Gardens [1293]. Gardiner [1907, 1902, 1925]. Gardner [2616]. Garland [1069]. Garrod [1299, 2634]. Garth [986]. Gas [2716, 972, 1234, 1095, 2047]. Gascoigne [1337, 1487]. Gases [1712, 484, 2529, 1695]. Gaspar [140]. Gassendi [2606]. Gatekeeper [2645, 1501, 1432]. gatekeepers [2346]. Gates [1528]. gather’d [2463]. gauge [561]. Gaunt [1064]. Gauss [937]. Gavin [537]. Gay [970, 2679, 1784]. Gay-Lussac [970, 2679]. Geburtstag [1849]. Gedde [1240]. Geiger [1639]. Gemelli [1379]. Gender [2584]. gendered [2581]. gene [1893]. General [439, 1269, 2089, 2533, 1233, 2643]. Generation [2728, 1685]. generations [696, 980, 2767, 1617]. generators [676]. Genes [2865]. genesis [850]. genetics [1264, 902]. Geneva [1416]. Genevese [714]. genial [967]. Genius [2647, 2694, 2697, 2733, 2758, 2759, 2919, 1471, 1383, 1254, 1961, 1588, 1632, 1158, 1655, 1669, 1544]. Genomes [2865]. genre [2489]. gentle [1339]. Gentleman [1988, 814, 2262]. Gentleman-benefactor [1988]. Geographical [14, 2434]. geographies [2433]. Geography [2248]. geological [1308, 791, 900, 1094, 1950, 2734]. Geologist [2710, 929, 2148, 2616, 1358]. Geologists [2022]. Geology [2235, 1433, 2778, 2810, 1385, 2879, 873, 1132, 769, 2389, 865, 1470]. Geomagnetic [2631, 1456, 2435]. Geomagnetici [2631, 1456]. Geometry [2417, 1335]. Geons [2869, 1611]. Geophysical [310]. Georg [975, 958]. George [1646, 1662, 223, 1272, 745, 1317, 447, 1706, 548, 2796, 94, 1130, 1209, 2576, 2653, 1131, 2303, 1606, 1696, 1792, 2405, 2427, 2060, 2210, 2534, 1242, 1715, 1252, 2786, 770, 2515, 1079, 297]. Georges [2293]. Georgian [1141, 2489, 2579, 2389]. Georgina [1539]. geoscience [2492]. Gerhard [2922, 1732]. German [1948, 868, 1616, 2550, 2370, 1698, 2011, 2456]. Germany [1616, 2403, 1852, 2789, 2399, 2224, 2371]. Gerstner [1358]. Gertrude [902]. Gesellschaft [28]. Giambattista [228]. Gibbon [161, 1242]. Gideon [200]. Giessen [1931]. Gift [66, 2789, 91, 1616, 2160, 1958]. Gifts [422]. Gilbert [2681, 997, 1767, 2138]. Giles [1524, 2555]. Gillham [1673]. Gillian [1654, 1548]. Gillies [216]. Ginseng [926, 2490]. Gio [1380]. Given [1803, 1946, 1996, 2062, 2070, 2196, 2261, 2304, 2339, 2363, 2392, 1135, 1177, 1226, 1267, 1311, 313, 142, 1351, 1405, 1452, 1500, 1573, 1715, 1640, 1690, 1752, 1832, 1859, 2149, 1899, 2114]. glaciers [678]. Glasgow [1358]. Glass [2780, 1714, 1917]. Gleanings [22, 36]. Glimpses [271]. Global [2878, 1803, 1293, 2219, 2310]. Globes [2891, 1633]. Glorious [1051]. Goddard [375]. Godfrey [271, 283, 1023]. Gods [2736, 1614]. Goethe [700]. golden [1420, 2398]. Goldsmith [578]. Golinski [1550, 1232]. Good [2737, 1815, 1674]. Goodman [1207]. Gordon [1950]. gorillas [1243]. Göttingen [142]. Gouk [1520]. gout [2009]. governance [2591]. Government [890]. Governor [68, 988, 836]. Gowland [235, 457]. Grace [1310]. Graddon [716]. Grand [513, 2576, 1586]. Grander [2752, 1704]. Grandfather [944]. grandson [944]. grangerized [1949]. Grant [1379, 1742]. Granta [1539]. granted [627]. granting [455]. Granville [2329]. Graphical [2589]. grass [2468]. Graunt [495]. Gravesande [2336]. Gravitation [1053, 790]. gravitational [596]. Gray [866]. greasepaint [1608]. Great [2688, 1777, 2907, 1541, 1357, 1750, 1544, 1980, 2868, 1295, 1727, 1585, 1844, 409, 1992, 1366, 1394, 1415, 2676, 2558, 2209, 1795, 2110]. Greate [2846, 1746, 2548]. greatest [1696]. greatness [1873]. Green [1252, 1816, 1919, 2653]. Greenwich [789, 1246, 1633, 604, 1406, 1722, 2891, 1968]. Gregory [55, 2556, 1149, 649, 2491, 2299, 1578]. Greig [1447]. Gresham [844, 426, 2453]. Grew [711, 1097, 2771, 2239, 907, 2614]. Grey [668]. Greystone [1639]. Griess [2325]. Griffith [280, 2506]. Griffiths [1234]. Grimaldi [1061]. Grosseteste [2512]. Grosvenor [401, 400]. Grote [1242]. ground [1394]. Group [2076, 952]. Grove [273]. grown [2903]. growth [9]. Grub [2534]. Gruber [1248]. Gründung [1849]. Guests [401]. Guglielmo [2018]. Guicciardini [1501]. Guide [2656, 1068, 295, 1701]. Guild [1332, 1315]. Guildhall [1905]. Guillaume [2182]. Guillemard [63]. Gulliver [1326]. gunnery [2137]. gunpowder [2799, 2455]. Guns [1748]. Guy [1314, 549]. Guyton [563]. Gwendy [999]. Gwillim [2601].

H [143, 1069, 1636, 637, 2122, 1229, 232, 1190, 1607, 1332, 982, 2301, 1208, 730, 1432, 1713, 2128, 1474, 1431, 1339, 2775]. H. [178, 122, 2112, 1474]. H.-B. [178]. H.M.S [1591]. H.M.S. [2910]. Haber [2375]. Hadfield [981]. Haistwell [650]. Hakim [1162]. Haldane [1007, 1265, 902]. Hale [1077]. Hales [833, 57, 968]. half [2034, 1645]. Hall [1181, 403, 1231, 1316, 1667, 889, 1157, 409, 1543, 2539, 1382, 1608, 2071, 2113, 397, 578, 398, 1070, 2076, 1382]. Halle [1948]. Halley [640, 313, 898, 1429, 739, 433, 312, 313, 1239, 978, 1106, 1194, 2673, 1610, 1869, 314, 2234, 823]. Halliburton [940]. Hamdard [1162]. Hamilton [1417, 1827]. Hanbury [1164, 2510]. Hanckwitz [283]. Hancox [1336]. Hand [2430, 1040]. Hand-in-hand [2430]. Handle [2580]. Hans [255, 641, 256, 887, 257, 1474, 2880]. Happened [2855, 446, 1919, 1713]. Hardback [1319, 2385, 1616, 1248, 1638, 1594, 1550, 2262, 1388, 1689, 1702, 1765, 1428, 1673, 1520, 1592, 1671, 1655, 1637, 1731, 1782, 1766, 1653, 1589, 1591, 1429, 1470, 1593, 1457, 1685, 1688, 1706, 1707, 1704, 1505, 1434, 1384, 1714, 1613, 1427, 1160, 1484, 1730, 1098, 1800, 1634, 1687, 1786, 1487, 1633, 1950, 1709, 2151, 1672, 1529, 1435, 1471, 1545, 1615, 1749, 1598, 1768, 1617, 1432, 1488, 1507, 1632, 1750, 1425, 1436, 1485, 1502, 1708, 1670, 1459, 1544, 1474, 1732, 1385, 1651, 1711, 1614, 1431, 1382, 1923, 1410, 1406, 1359, 1356, 1357, 1409, 1362]. Hardbound [1332]. Hardwicke [1398]. Hardy [1607]. harem [1649]. Harley [1182]. Harlow [2500]. harm [2564]. Harman [1118, 1471]. Harold [1068, 659, 708, 1068, 1155]. HarperCollins [1638, 1702, 1704, 1271, 1629]. Harriot [1746, 1746, 225, 1994, 2846]. Harris [1572, 1685]. Harrison [1332, 1143, 2111, 2793, 483, 95, 1895, 891]. Hart [1635]. Hart-Davis [1635]. Hartley [659, 708, 571, 753]. Hartlib [2315, 258]. Harvard [1138]. Harvey [1458, 337, 1881]. Harvill [1068]. Hastings [94]. Hatchett [1723]. hath [2463]. Hatherly [268]. Hatipoglu [2598]. Hauksbee [2171]. Haven [1184, 1520, 1297, 1355, 1083]. Hawkes [1293]. Hawking [1782, 2911]. Haycock [1707]. Haynes [1319]. hdbk [1180, 1206, 1252]. head [2593, 336]. Headline [1655]. Heads [1785]. Healey [1655]. health [2573, 2219, 2353]. heat [2529, 2532]. Heath [1875]. Heathcoat [349]. Heathcoat-Amory [349]. heaven [2010, 1189]. Heavens [2673, 1834, 1429]. Heberden [1098, 2900]. Hebræi [2322]. Hebrew [2322]. Heidelberg [1103]. Heinrich [2599, 1773]. Heisenberg [2659, 2596, 1229]. Held [2623, 142, 457, 1056, 2923, 1898]. Helden [1300]. Hele [1293, 1729, 1524]. Helge [1617]. Helmholtz [2158, 2147]. helmontii [2270]. Helsinki [1724, 1725, 1726]. Hempstead [1654]. Henri [1975]. Henry [615, 1269, 2643, 795, 945, 1497, 793, 1028, 2754, 1201, 1150, 796, 1921, 1186, 1079, 437, 1464, 387, 1181, 1328, 792, 726, 138, 583, 2710, 470, 479, 581, 1075, 2719, 2723, 2726, 1685, 907, 2301, 1198, 2123, 1773, 1667, 2315, 1819, 1812, 642, 549, 2136, 1358, 755, 1382, 1037]. Henshaw [906, 1347, 2574]. Her [1783, 2597, 2582, 307, 311, 397, 574, 122]. Heralds [1858]. Herapath [484]. Herbaria [1880]. Herbarium [2286]. Herbert [951, 1864]. Heredity [402, 2593, 234]. Heritage [2622, 557, 1138]. Hero [1763, 1885]. heroes [2473]. Herries [1950]. Herschel [1184, 1798, 1647, 971, 2902, 2742, 1986, 980, 846, 847, 848, 2835, 1109, 1078, 2269, 2295]. Hertha [1112]. Hertz [655]. Herzberg [1732, 2922]. Hevelius [698, 1037]. Hicks [2496]. Hidden [2621]. Higgins [2092, 2282]. high [1449, 613, 2296, 1352]. high-altitude [2296]. high-voltage [1449]. higher [2400]. Highfield [1249]. Highness [217, 122]. Hilger [1163, 1190, 1164]. Hill [386, 2210, 1220, 2137]. Hills [1786]. him [316]. Himself [1292, 2748, 2574]. Hinde [1614, 1674]. Hindley [651, 642]. Hinshelwood [577, 578, 1758]. Hinxton [2076]. Hip [2866, 1103]. Historia [1515]. historian [1325]. Historians [2254, 1291]. historic [557]. Historical [1068, 1248, 1410, 2692, 2722, 1947, 1316, 1613, 1442, 1239, 928, 2099, 2575, 2458, 2053, 2515, 1903]. historie [967]. Histories [2402, 2485]. historiography [2480, 2486]. Historiques [1599]. historischen [1849]. History [2622, 1068, 2623, 734, 1675, 2402, 1823, 2638, 1235, 2069, 2648, 1248, 1212, 2883, 2655, 2658, 1798, 1763, 1666, 2682, 1566, 1883, 2703, 1831, 2713, 1867, 1331, 1270, 1506, 2911, 2772, 2779, 2365, 2787, 1949, 2817, 2824, 1056, 1425, 1502, 744, 2839, 2841, 1459, 1291, 2580, 1711, 2524, 2384, 524, 1845, 2385, 262, 275, 291, 305, 331, 340, 358, 367, 441, 452, 474, 487, 498, 508, 517, 528, 540, 551, 570, 588, 601, 619, 634, 646, 657, 671, 683, 704, 718, 733, 747, 761, 772, 785, 799, 807, 818, 829, 842]. history [852, 861, 931, 1650, 1941, 2362, 1273, 1550, 1165, 1204, 2436, 1376, 2272, 179, 1671, 1782, 1589, 1848, 1607, 1780, 2520, 2186, 2498, 2573, 676, 2587, 1095, 1810, 1948, 2594, 2104, 2767, 6, 2441, 1138, 1333, 2575, 522, 966, 1089, 1617, 2052, 2181, 2316, 2450, 1473, 1769, 7, 166, 2426, 2474, 1557, 1305, 2486, 2273, 258, 1041, 2134, 2183, 191, 205, 220, 1504, 999, 446, 1003]. hitherto [585, 479]. Hitler [1616, 2789]. HMS [1285]. HMSO [1212]. Hobbes [175, 226, 2417]. Hodder [1207]. Hodgkin [916, 405, 1179, 1539, 2597, 2699, 1312]. Hofmann [1357]. Hogben [1484, 2901, 2319]. Hokusai [1992, 2209]. Holbrook [1212]. Hole [2502]. Holes [2667, 1189, 1611, 2869]. Holland [676, 996]. Hollingdale [1086]. Holy [577]. Home [1251, 2744, 2819, 2226, 574, 556, 2501, 1290]. homeopathy [2441]. Homeric [2022]. homes [103, 554, 1791]. Homoeopathy [2441]. Hon [349, 661, 792, 400, 722]. Hon. [997]. Hong [1589]. honorable [1219, 2745]. Honorary [413]. Honour [2834, 622, 12, 1162]. Honourable [2070, 338, 381]. Honours [311, 695]. Hooke [382, 1720, 1777, 2903, 1116, 1315, 1925, 1383, 1879, 882, 2637, 1720, 1744, 1366, 1394, 1415, 2676, 2169, 2694, 534, 1777, 1814, 1760, 196, 1061, 1905, 1395, 1346, 1743, 2073, 2751, 2753, 2207, 2323, 1116, 1315, 1832, 1908, 1217, 2803, 881, 1003, 531, 2612, 1914, 1878, 1902, 139, 803, 1060, 2618, 2110, 1777]. Hooker [354, 2551]. Hopkins [1119, 1201, 235, 457, 1319]. Hopkinson [758]. Horace [2208]. horizons [410]. Horizontal [2511]. Horology [196]. Horrocks [56]. Hoskin [1798]. hospital [70]. Hotel [349, 490, 308]. Hounslow [1875]. House [401, 1204, 574, 575, 2513, 2066, 1750, 555, 1240, 273, 2488, 1774, 78, 2745, 1359, 400, 2246, 2430, 116]. Houseflies [2535]. houses [433]. Housewife [2560]. Howard [1173, 700, 1072, 1468, 2331]. Howse [1099]. Hoyle [1290]. Huddart [1100]. Huemul [2784]. Hugh [2100, 1251]. Hughes [931, 875, 1180]. Human [1304, 2687, 2818, 1436, 2103, 1595, 462, 2080, 2573, 2458, 2246, 2006, 1846]. Humanism [1387]. Humanist [2901, 1484]. Humanity [2816]. humans [2048]. Hume [2607]. Humphrey [987, 859]. Humphries [1068]. Humphry [448, 484, 2404, 2448, 1386, 2763, 2764, 1210, 1980]. Humuel [1475]. hundred [883, 1884, 1434, 42, 33, 2917]. Hungarian [2577]. Hunt [91, 2123]. Hunter [1777, 1082, 1292, 1270, 1116, 1250, 1333, 1425, 1453, 909, 1262, 1129, 2127]. Hunterian [1248, 2812]. Hunting [2127]. hurrah [2546, 2104]. Hurricane [1188, 2705]. Hutton [2755, 1748, 2359]. Huxley [1300, 2908, 1186, 2923, 2089, 2122, 2690, 2382, 945, 793, 1028, 2754, 1201, 1298, 1143]. Huygens [1054, 1053, 1873]. Hyacinthe [1108]. hybrid [2007]. Hydrodynamic [2662, 1189]. Hydrogen [2663, 1073]. Hydrologist [640]. Hydromagnetic [2662, 2664, 1189]. hydrostatical [2405, 2427]. Hypatia [2622, 1138]. Hyperarctic [2199]. Hypnosis [2367, 2371]. hypnotic [2370]. hypnotism [2368, 2366, 2365, 2369]. hypnotists [2367]. hypotheses [2194, 1054]. hypothesis [596]. Hysteria [2593].

I. [1005, 1701]. I.G.Y. [311]. Ia [2546]. Ian [1388, 1502]. iatrochemistry [2605]. IBSN [1528]. Ice [2674, 1593, 1593]. iceberg [939]. Iceland [2197]. Icon [1656]. iconic [2393]. Iconography [2639, 1763]. I’d [1542, 2816]. Ida [2284]. Ideas [2905, 2909, 1548, 2765, 2766, 1228, 1027, 937, 1453]. identified [1479, 2456]. identities [1370, 2570]. identity [1092]. ideologies [2591]. Iguanodon [200]. II [1119, 1293, 1662, 1448, 1776, 2897, 1343, 514, 2494, 1178, 1775, 35, 227, 822, 2530, 36, 1535, 431, 372]. III [1272, 2796, 1048, 2899, 246]. Illinois [1268]. Illness [880, 1307, 1892, 1945, 1224, 1348]. ills [863]. illuminated [1249]. illus [1505, 1295]. illust [1355]. Illustrata [2411]. illustrate [2048]. Illustrated [2793, 927]. Illustrating [1875, 1420, 1565]. Illustration [1023, 2703, 1965, 2190, 2204, 2218, 2231, 2243, 2267, 2279, 1235]. illustrations [1271, 1298, 1313, 1359, 2253]. illustrator [2614]. illustrators [1283]. illustrious [2922, 1732]. im [1849]. Image [1669, 2393, 1557, 468]. Images [2703, 2205, 1871, 1235]. Imaging [2639, 1763]. Immaterial [1181]. impact [2575]. impairments [2566]. impeccable [2336]. Imperial [1766, 1653, 1474, 1466, 445, 950]. Imperialism [2845, 2059, 1142]. impiety [225]. importance [2248, 2884, 2303, 1786]. Imprimatur [1069]. imprisonment [138]. improved [1624]. Improvements [1919]. improving [1263]. impulses [1179]. Imre [2577]. inaugural [649, 753]. inaugurate [1757]. inauguration [310]. Inborn [1299]. inch [1330, 1352]. inclinatory [1170]. including [1003, 2822]. indeterminate [681]. Index [51, 111, 133, 154, 182, 207, 241, 264, 292, 322, 341, 368, 428, 442, 464, 489, 509, 529, 552, 572, 602, 635, 658, 684, 719, 748, 773, 800, 819, 843, 862, 878, 892, 913, 933, 953, 974, 991, 1009, 1033, 1057, 1389, 1438, 1553, 1618, 1676, 1734, 1788, 472, 1455, 338, 1830, 744, 560]. India [2745, 1469, 1483, 87, 164, 1392, 1007, 856, 2430, 1920, 1254]. Indian [1096, 114, 1361, 959, 1096, 1171, 1312]. Indiana [1138, 1276]. Indies [238]. Individuality [2634, 1299]. inducements [2414]. induction [1200, 525]. Indus [2506]. Industrial [1930, 2028, 2029, 2324, 961, 1757]. industrialist [2224]. Industry [972, 2021, 2018, 58, 2031]. inertia [503]. infectious [2048, 909]. inferior [1794]. Infinity [2758, 2759, 1254]. Inflammation [2222]. Influence [2644, 2894, 2515, 2606, 2274, 644, 258, 1211, 1636]. Influences [1818, 1004, 1032, 1175]. Informal [2655, 999]. Information [121, 134, 169, 453, 476, 542, 786, 787, 830, 853, 2230, 1688, 2424, 766, 651]. infuriating [1961]. inheritance [805]. initiative [2558]. injuries [2593]. injury [2564]. ink [2415]. innovation [2031]. innovations [2195]. innovator [901]. innumerate [1216]. inquiry [1003]. Insatiable [1271, 2827]. Insects [411, 2184]. Insights [2566, 1383, 2521, 2488]. Inspiration [1655]. Inspirational [2733, 2515, 1655]. inspired [1811]. Instauration [2868, 1727]. Institute [2623, 164, 1406, 1689, 1331, 1705, 1334, 1672, 1228, 1056, 1546, 12, 1459, 1597, 2577, 1622, 2573]. Institution [2749, 999, 2907, 1208, 1545, 1750, 1980, 802, 643, 2655, 2303, 402, 403, 824, 1443, 2128, 2112]. Institutional [1379, 2715, 2408, 2170, 2307]. institutions [237, 167]. Instrument [2787, 2434, 2256, 676, 642, 712, 1273]. Instrumentation [1609]. Instruments [2631, 1456, 2740, 2440, 1212, 557, 2557, 2337, 2435, 2299, 2592]. Integration [1457, 2027]. intellect [864]. Intellectual [1379, 2715, 2750, 1333, 809, 2249, 594, 593, 592]. intelligent [1983, 836]. inter [2558, 1314, 890]. inter-academy [2558]. inter-war [890]. Interaction [2515, 676]. interactions [409]. interdisciplinarity [2610, 2438]. Interest [832, 1607]. interesting [235]. Interests [2817, 1447, 299, 1138]. interference [1605]. International [2526, 439, 2177, 963, 92, 416, 310, 1339, 1920, 2492, 876, 203, 14, 932, 15, 535, 101, 2600]. internationalism [963, 932]. Internet [1604]. interplay [2406]. interpretations [1994]. interpreters [2206]. interrupted [1967]. intertwined [2573]. Intervenor [1483, 1469]. interwar [2552, 2578]. Intimate [2620, 1138]. intrigue [2555]. Introduction [2537, 1248, 1410, 2517, 2692, 2498, 2563, 2480, 2446, 1085, 1292, 1331, 1081, 1045, 2396, 909, 2602, 1567]. Introductory [2042, 2041, 1014]. Invar [1806]. invented [1156]. Invention [1184, 1360, 2514, 2835, 2846, 1646, 613, 1746, 2159, 701, 2921, 1327]. Inventive [1777, 2751]. Inventor [2716, 1234, 1027]. Inverse [2508]. Investigaciones [189]. Investigating [1203]. investigation [1894, 407]. Investments [2076]. Invisible [2648, 1930, 1276, 1467, 1671, 480]. Inwood [1777]. IO [264, 2546]. IoM [1572]. Ion [2663, 1073]. ionosphere [1375]. Iran [2488]. Ireland [481, 514, 1855, 2916, 1705]. Irish [1175]. iron [2431, 58]. Irreversibility [1488]. Isaac [1085, 1120, 854, 248, 1038, 125, 1023, 1004, 1823, 1636, 2281, 482, 99, 2477, 2893, 1032, 2475, 2721, 2724, 1396, 2725, 2358, 1231, 2322, 863, 1258, 1871, 520, 2554, 2802, 2058, 2474, 1543, 2257, 2618, 494, 1382, 1843]. Isaaci [2556].

Isambard [2859, 1253]. ISBN [1068, 1182, 1209, 1293, 2385, 1272, 1727, 1338, 1656, 1456, 1475, 1616, 1088, 1249, 1235, 1300, 1269, 1317, 1069, 1117, 1185, 1595, 1071, 1181, 1636, 1146, 1097, 1248, 1729, 1410, 1212, 1638, 1572, 1594, 1184, 1251, 1318, 1455, 1550, 1798, 2262, 1145, 1275, 1299, 1388, 1763, 1379, 1406, 1501, 1689, 1542, 1702, 1765, 1785, 1433, 1411, 1361, 1103, 1508, 1428, 1163, 1673, 1520, 1592, 1671, 1655, 1637, 1731, 1782, 1766, 1653, 1188, 1100, 1589, 1639, 1524, 1229, 1654, 1458, 1190, 1591, 1119, 1746, 1780, 1121, 1085, 1082, 1141, 1159, 1180, 1206, 1230, 1292, 1337, 1314, 1429, 1331, 1470, 1525, 1386, 1297, 1593]. ISBN [1705, 1457, 1547, 1685, 1688, 1706, 1096, 1387, 1707, 1704, 1213, 1505, 1120, 1231, 1434, 1271, 1332, 1384, 1714, 1316, 1613, 1767, 1179, 1427, 1160, 1101, 1118, 1140, 1298, 1313, 1359, 1484, 1207, 1233, 1270, 1336, 1548, 1635, 1730, 1784, 1356, 1098, 1800, 1634, 1687, 1164, 1360, 1335, 1667, 1252, 1786, 1122, 1116, 1138, 1157, 1210, 1250, 1294, 1333, 1315, 1355, 1748, 1099, 1102, 1551, 1334, 1487, 1633, 1950, 1142, 1709, 1232, 1255, 2151, 1504, 1672, 1541, 1629, 1783, 1527, 1529, 1290, 1549, 1353, 1539, 1357, 1435, 1471, 1489, 1545, 1615, 1710, 1749, 1205, 1611, 1123, 1161]. ISBN [1143, 1211, 1228, 1319, 1598, 1768, 1187, 1617, 1408, 1432, 1473, 1488, 1546, 1526, 1507, 1596, 1632, 1712, 1769, 1750, 1801, 1087, 1186, 1425, 1409, 1436, 1485, 1502, 1708, 1453, 1631, 1669, 1670, 1713, 1802, 1086, 1543, 1459, 1430, 1362, 1544, 1474, 1732, 1234, 1291, 1358, 1385, 1383, 1276, 1651, 1711, 1614, 1674, 1431, 1296, 1183, 1522, 1382, 1340, 1253, 1144, 1339, 1295, 1083, 1923, 1254]. island [1465]. Isle [1308]. isomers [840]. Isotopes [840]. Israel [1660]. ISSN [1506, 1291]. issue [1365, 1414, 1492, 1512, 1771, 2396]. Issues [1141, 1441]. Italia [1456, 2631]. Italian [203, 2709, 855, 1791]. Italy [1456, 2631, 1456, 2368, 1523, 920, 2255, 1345]. itinerant [1480]. itinerary [2561]. IV [1117, 1487, 270]. Ivory [1518, 1648]. ix [1071, 1207, 1270, 1138, 1710, 1056, 1430, 1544, 439].

J [1293, 1475, 1235, 1572, 1184, 1433, 2682, 2686, 1163, 1119, 1583, 1141, 1180, 2494, 1434, 938, 1140, 1360, 716, 2493, 730, 1489, 1545, 1211, 1750, 1186, 1409, 1631, 1459, 316, 2398, 326, 1339]. J. [1626, 2682, 2686, 725, 1583, 866, 280, 1332, 1175, 1709, 1527, 1489, 1265, 902, 2850, 1153, 316, 1431, 1522, 1541]. J.M [1522]. Jack [1198, 1435]. Jackson [987, 1585]. Jacob [1248, 2336, 1942]. Jacobite [1904]. Jacqueline [1746]. Jagadis [1421]. Jagdish [1362]. Jamaica [1465]. Jamaican [628]. James [236, 622, 1636, 1160, 606, 1774, 1471, 1545, 1749, 1161, 2826, 1750, 32, 749, 1218, 1325, 883, 838, 612, 2654, 564, 1196, 2463, 1518, 1648, 884, 2148, 623, 1174, 2896, 2897, 2898, 2899, 2238, 707, 1247, 1880, 2125, 1373, 2300, 2462, 2469, 2468, 624, 727, 1221, 652, 2915, 2465, 2464, 324, 2466, 1227, 1357, 1598, 979, 2299, 625, 2609, 1495, 1695, 2047, 1087, 663, 565, 55, 562, 611, 2467, 921]. Jan [1232, 1550, 2185, 1430]. Jane [1272]. Janet [1313]. January [114, 2680]. Janus [1317]. Japan [669, 1787]. Japanese [1941]. Jaques [1092]. Jardine [1777, 1704, 1777]. Jarrow [2122]. Jean [2262, 1145, 1616, 1108, 662]. Jeffreys [1155]. Jenkin [1654, 681, 2675]. Jenner [157, 742, 1183, 2700, 18]. Jenny [1730]. Jensen [1186]. Jeremiah [212, 56]. Jewels [781]. Jewish [1660]. Jews [159]. Jim [1777, 2609]. Joan [1880]. Joana [2586]. Jodrell [1178]. Johann [461, 140, 2599, 1773, 1849, 1519, 2325]. Johanna [1712]. John [1182, 2630, 2632, 2633, 2091, 1317, 373, 1907, 1248, 1318, 1388, 1406, 1689, 2674, 1188, 1639, 731, 2616, 1337, 1314, 1331, 1593, 1271, 1332, 999, 2214, 1336, 899, 1126, 1334, 1217, 1921, 327, 1925, 2825, 1708, 374, 1647, 460, 580, 1109, 2260, 1253, 1144, 2925, 524, 174, 2111, 2414, 1720, 2248, 1516, 484, 1240, 738, 2318, 957, 2670, 2249, 2244, 1262, 1785, 2677, 873, 214, 261, 300, 391, 1103, 1007, 1049, 2892, 2415, 900, 2701, 1493, 2381, 710, 1220, 856, 216, 495, 971, 2282, 2018, 2144, 1260, 1525, 2727, 2729, 1397, 2255, 2357, 2258, 2457, 980, 1064, 1644]. John [2182, 1634, 1687, 63, 68, 1533, 1355, 1514, 1334, 1487, 2476, 1774, 2184, 2157, 2246, 1353, 2263, 2245, 2247, 1615, 1611, 2388, 483, 1902, 2419, 834, 2418, 2413, 2827, 848, 1446, 1904, 1436, 560, 639, 642, 1479, 1916, 712, 985, 1895, 558, 2853, 1234, 1385, 2511, 2330, 908, 2866, 909, 2872, 1467, 1481, 38]. Johns [1119, 1319, 1201]. Johnson [1748, 259, 2322, 767]. Johnstone [770]. Joining [1968]. joint [1978, 1908]. jointed [1039]. jointly [203]. Jon [1613]. Jonas [1230, 2871]. Jonathan [375, 1458, 1470, 1313, 1336, 1634, 1709]. Jones [2262, 726, 86, 859, 768]. Jordanova [1596]. Joseph [1068, 1293, 39, 2878, 768, 2656, 2657, 85, 1444, 2661, 2262, 675, 1653, 1100, 1066, 1373, 1298, 273, 1487, 1161, 1056, 2834, 1162, 1068, 2623, 2624, 32, 837, 1918, 1411, 245, 1369, 2448, 2706, 2707, 2490, 1337, 2220, 2745, 1282, 1221, 741, 757, 301, 2800, 2804, 2805, 2806, 2222, 2237, 2450, 2048, 1261, 1076, 354, 1172, 2446, 2449, 2221, 1923]. Joseph-Francois [2490]. Joshua [701]. Josiah [1417, 1233, 2823, 471]. Joule [1087, 2158, 2654, 653, 2047, 2532]. Journal [1506, 2615, 2496, 1479, 2342, 2590]. journalistic [1094]. Journals [1978, 2309, 2121, 997, 2346, 2347]. Journey [2860, 2561, 1096]. journeys [2444]. Joy [1458, 1336]. Joyce [1317]. joynt [2465]. joynt-stock [2465]. JSTOR [1574]. Jubilee [281, 1266]. Jubilees [1378]. judgement [2551]. Judith [1068]. Jukes [837]. Julian [1206, 2908, 1143, 1823, 1300, 2089, 2923]. July [401, 124, 399, 745, 604, 402, 397, 403, 398, 405, 400, 409, 411, 14, 2773]. June [2071, 2070, 1248, 1506, 2812, 2330]. Jung [2302]. Jungnickel [1408]. Junior [68]. Jupiter [1561]. Juretschke [1339]. Just [2523]. Justice [2260]. justified [1775]. justifying [2559]. Justus [2645, 1432].

K. [1731]. K.B.E [591]. K.C.B [1251, 491, 580]. K.C.H [1251]. K.G [668]. K.G. [217]. K.M [1251]. Kaiser [28]. Kaiser-Wilhelm [28]. Kamerlingh [1712, 2913]. Kameshwar [1213]. Kanagawa [1992, 2209]. Kangaroo [675]. Kanigel [1254]. Kantian [2608]. Kantianism [2403]. Kapitza [1030]. Karachi [1162]. Karl [2233, 1223, 1800, 2918, 1529, 2821, 1265, 1387]. Karlsruhe [535]. Kármán [2600]. Kashmir [2559]. Kass [1138, 1276]. Kass-Simon [1138, 1276]. Kathleen [1983]. Kato [1339]. Kay [1934]. Kaye [702, 1845]. KCB [2819]. KCH [2819]. KCSI [1866]. Keeping [1933]. Keir [565]. Keith [2515]. Keller [1163]. Kelly [1780, 1296]. Kelvin [1863, 1837, 1865, 2840, 1101]. Kenelm [388]. Kenneth [1611]. Kepler [2548]. Ker [1328]. Kettering [1346]. Kettering-stone [1346]. Kettle [2065]. Kevin [1782]. Kew [1293]. Keynes [1069, 1071, 1591, 1143, 1228, 2193, 248]. Kiessling [1519]. Kilkenny [2122]. Kincardine [102, 393]. Kind [1023, 2466, 1992]. Kindred [2708, 1528]. kinetic [484]. kinetics [1836]. King [223, 1293, 1272, 1729, 1524, 1102, 725, 2796, 248, 122, 372]. King-Hele [1293, 1729, 1524]. Kingdom [1212, 2859, 1253, 2740, 2450]. kitchen [2452]. Klug [1351, 1573]. Kluwer [1636]. KM [2819]. Kneller [1023]. Knew [1777, 2751, 2758, 2759, 1254, 2159]. Knight [1386, 1210, 2859, 1253]. Knight-Errant [2859]. Knole [2539]. know [2568]. Knowledge [1338, 1589, 2706, 2713, 1337, 2549, 1651, 2867, 2874, 1993, 2586, 1550, 2072, 2393, 2123, 2323, 2581, 2476, 2399, 45, 2430, 2396, 2400, 1651]. knowledge-making [2399]. known [835, 1204]. Knox [1782]. Koestler [2338]. Kohn [1910]. Kolb [1549]. Kong [1589]. Koninklijke [1712]. Kontext [1849]. Korean [2499]. Kragh [1617]. Krakatoa [1519]. Krishna [1361]. Krishnan [1498]. Krona [1291]. KT [1690, 1752, 2114, 1946, 1996, 2062, 1794, 1640, 1803, 1899].

L [1456, 1318, 405, 1208, 1786, 1950, 1545, 1750, 1087, 2128, 2112]. L. [996, 1195, 1485]. L.B [1119]. Laboratories [654, 2292, 268]. Laboratory [1571, 1137, 1934, 1191, 890, 260, 1449, 472, 1684, 1006, 302, 1349, 2006, 1933, 1329, 1937, 2076]. laborinth [2548]. labour [2576]. labourers [1722]. Laby [1845]. L’Académie [532]. Ladies [1342]. Lady [2782, 2817, 1702, 1138, 2601, 2582]. Lafitau [2490]. Lagerkvist [1765]. Lakeland [1446]. Lalande [835]. lamb [1344, 2208]. Lancastrian [1071, 2831]. Lancelot [2901, 1484, 2319]. landed [21]. landscape [2499, 2552]. landscaping [2503]. Lane [1269, 2643, 1529]. Lang [1727, 1163]. Langevin [2159, 152]. Language [2687, 1595, 2100, 1943]. Languages [1595]. Lankester [1185, 2695]. Laplace [1874]. Large [2603, 1466]. Larmor [1172]. Larry [1184, 1232, 1409]. Laser [2855, 1713]. Lasers [1713, 2619]. last [699, 215, 1648, 1869, 2350, 1895]. lasting [2606]. Late [2622, 2750, 1198, 1333, 2195, 1793, 2366, 2400, 2441, 632, 1138, 2488, 2449, 2369, 2389, 2507]. late-Georgian [2389]. late-Victorian [632]. Later [2745, 2533]. latest [1067]. Latimer [1317]. Latin [446]. launch [1896]. laureate [1740, 1350, 2136, 1598]. laureates [2597]. laurels [1310]. Laurence [380]. Laurent [2639, 1763, 155]. Laurentius [1849]. Lavoisier [1356, 1294, 155, 2585, 2639, 1763, 2693, 328, 1356]. Law [2018, 1613, 2158, 286, 840, 2369, 2508, 1507]. Lawrence [1056, 2852, 1144, 1825]. Lax [1802]. lay [2367]. layman [836]. Lead [1639]. Leader [1096]. leading [1286]. leaf [2398]. learned [1918, 2309]. Learning [2720, 2219, 1157, 1844, 2387]. learns [2322]. lease [1225]. leave [1697]. Lebenskraft [2608]. L’Écho [2590]. Lecture [1437, 2069, 402, 405, 624, 1787, 753, 1587, 409, 411, 2384, 313, 791, 403, 2472, 1028, 1910, 649, 1173, 1815, 1833, 826, 2, 754, 1055, 1751, 1404, 1040, 2442, 457, 1110, 985, 755, 872]. Lectures [1248, 2812, 643, 858, 769, 1003]. lectureships [873]. lecturing [2408]. Lee [2293]. Leeds [1867]. Leeuwenhoek [1140, 897, 2702, 1760, 1055, 1140, 1052, 210]. Leeuwenhoekiana [998]. LeFanu [1097]. left [1854]. Legacy [2702, 1434, 2909, 1228, 2852, 1907, 2226, 2225, 1493, 1580, 1140, 1013, 2413, 1144, 1140]. Legend [887, 2319, 1203]. Leghorn [1603]. Legibus [1613, 2815]. Leibniz [1323]. Leiden [2336]. Lemur [2571]. Lemurs [2571]. Léon [1768, 2854]. Leonard [1866]. Leonardo [1777, 1315, 2637, 2803, 2285, 2439]. Leopoldina [1948, 1849]. Lepidoptera [2467]. lepidopterist [2294]. Lesley [1406, 1689, 1334]. Leslie [1597]. lessons [2367]. Letter [64, 2422, 2272, 18, 84, 2540, 2406, 526, 907, 899, 1773, 1868, 2315, 1914, 2388, 160, 680, 1037]. Lettere [1379]. Letters [2510, 271, 1030, 1288, 2882, 1444, 2661, 175, 782, 1653, 2896, 2897, 2899, 1359, 2127, 345, 350, 85, 2381, 2448, 2528, 779, 190, 1118, 728, 757, 796, 1545, 1080, 2618, 199, 584, 2579, 1956]. Levant [272]. Levelt [1712]. Lewes [1315, 1079]. Lewis [63, 549, 1930]. liberal [2249]. liberated [90]. Librarian [143, 702, 2352, 450]. Libraries [534, 1841]. Library [1235, 1905, 2152, 2450, 174, 2415, 731, 302, 1842, 1860, 22, 36, 5, 995, 2542, 472, 1393, 1847]. Libyan [48]. Lichfield [1284]. Lichtenberg [958, 975]. Liebig [1432, 2645, 1931]. Lieutenant [1269, 2643, 2496]. Lieutenant-General [1269, 2643]. Life [2402, 2625, 2627, 2632, 2633, 2637, 2641, 1317, 2643, 1185, 2647, 1248, 1729, 2653, 1251, 2659, 1145, 2695, 2699, 1229, 1777, 2711, 2716, 1180, 1292, 2727, 2728, 2729, 1271, 2744, 2745, 2746, 2751, 2472, 2753, 2462, 2758, 2759, 2908, 1336, 1355, 2788, 2790, 2794, 2804, 2805, 2806, 1205, 2819, 2827, 2127, 2505, 1708, 2838, 2848, 1362, 2854, 1234, 2860, 2869, 1296, 1254, 1225, 1656, 1088, 1269, 2597, 228, 1970, 1455, 2889, 1007, 1673, 2893, 1100, 1524, 1493, 2604, 2603, 1660, 19, 1685, 2443, 1096, 1120, 2494, 1332, 1854, 2907, 1151, 1284, 2761, 1715, 1761, 1800, 2541, 209, 227]. life [246, 270, 283, 480, 513, 2915, 889, 2589, 2493, 1541, 1349, 1290, 1749, 1611, 2821, 1143, 1768, 2413, 1750, 1265, 2922, 2850, 950, 1544, 1732, 2924, 2602, 2925, 2876, 1629, 1068, 1539, 2748]. Light [1702, 1855, 1489, 2860, 970, 1184, 957, 1096, 2301, 1533, 1495, 1274, 864, 2592, 1327]. Lighting [2716, 1234]. Lightman [2385]. lightning [740]. likeness [1326]. Lilliputian [1995]. limelight [2495]. liminal [2526]. Limit [2506]. limits [2491, 2608, 2110]. Lincoln [1547]. Lincolnshire [2569, 2579]. Lindemann [1005, 1006, 1191]. line [2100]. lineage [887]. linearity [2209]. lines [910]. linguistic [2059]. Link [1813, 2730, 1635, 2104]. linked [2162]. links [1286]. Linnaeus [1813, 2059, 2571, 985]. Lionel [1403, 1423]. lions [1608]. Lippmann [2087]. liquefaction [2236, 1695]. Liquids [1712]. Lisa [1777, 1704]. Lisbon [2257, 2527]. Lise [2531]. list [1993, 1204]. listens [2597]. Lister [1182, 1283, 2223, 2226, 2225, 2220, 2219, 2182, 2814, 2222, 2237, 2058, 2145, 2316, 2048, 2224, 2221]. Lists [2910, 1591]. Lit [1380, 1379, 1523]. Literary [1288, 2342]. Literatur [2308]. Literatur-Zeitung [2308]. Literature [2732, 1065, 1350, 909, 1319]. Little [1594, 1731, 1802, 2502, 835]. liturgy [2419]. live [2566]. lively [2483]. Liverpool [1121, 1247, 2828]. Lives [2620, 1249, 2735, 1543, 2088, 1138, 1774]. Living [2708, 1179, 2387, 1528]. Ll [1293]. LL.D [268, 792]. Lloyd [2381]. local [872]. localization [2209, 1420]. localized [2162]. Localizing [2598]. Locke [1353, 523, 462, 1049, 1281, 1353, 2872]. Lodge [1121, 1360, 2311, 2828, 2921]. Lodwick [2163]. logic [2326, 298]. Londa [1138]. London [1068, 1209, 1616, 1235, 1317, 1069, 1595, 1146, 1212, 1638, 1572, 1594, 1184, 963, 1455, 1702, 1433, 1103, 1520, 1592, 1655, 1731, 1766, 1653, 1100, 1589, 1524, 1777, 1119, 1159, 1180, 1292, 1331, 1470, 1297, 1547, 1704, 1271, 1714, 999, 1140, 975, 2908, 1298, 1313, 1233, 1336, 1730, 1098, 1634, 1138, 1315, 1355, 941, 1950, 1709, 2151, 1541, 1629, 1783, 1529, 1539, 1489, 1545, 1123, 1161, 1143, 1228, 1319, 1056, 1473, 1186, 1708, 1669, 1234, 1614, 1674, 1296, 1183, 1340, 1923, 415, 2637, 1999, 2089, 1918, 835, 1621, 1366, 1394, 1415, 2676, 2121, 884, 2225, 1777, 1932, 725, 1024]. London [936, 2282, 594, 577, 2734, 98, 593, 2172, 2753, 1171, 2409, 1112, 1154, 2464, 996, 1195, 371, 1136, 2803, 966, 409, 1973, 769, 744, 197, 2319, 591, 820, 804, 2110, 2449, 2453, 2548, 74, 1777]. Londoner [1401, 1194]. long [1401]. long-distance [1401]. Longitude [36, 835, 1561, 483, 2111]. Lonsdale [1983]. Lopez [1083]. Lord [1946, 2062, 661, 1101, 400, 1803, 1899, 579, 1996, 764, 831, 598, 1640, 1690, 1752, 1832, 1859, 2114, 1865, 2840, 93, 591, 2273, 2260]. loss [2092]. Lost [2093, 2674, 1593, 2748, 2463, 1292, 2255, 2213, 2822, 1974]. Louis [930, 922, 2236, 563]. Love [1638, 1184]. Lovell [1146, 1190]. lover [1810]. low [2236]. low-temperature [2236]. Lowe [1119]. Loyal [908]. Loyalties [2727]. Lt [745]. Ltd [1185, 1595, 1798, 1275, 1785, 1332, 1714, 1144]. Lubbock [327, 460, 2248, 2249, 2244, 2246, 2263, 2245, 2247]. Luc [662]. Luca [1331]. Lucas [1293]. Lucasian [1782, 1782, 2911]. Lucretius [503]. Lucy [2669, 1710]. Ludlow [1946, 1996, 2062, 1899, 2114]. Ludmilla [1596]. Ludus [2270]. Ludwig [2660, 1488]. Luigi [1355, 2848]. Luke [700]. lumière [1054]. Lunar [1730, 2857, 2548, 546, 565]. Lunaticks [1437, 1404]. Luncheon [415]. Lupus [1348]. Lush [745]. Lussac [1784, 970, 2679]. Lutterworth [1710]. lxv [1545]. lxvi [1689]. Lydall [710]. Lydia [1505]. Lyell [355, 1433, 769, 1433, 2616, 190]. lynchpin [2484]. Lynn [1255]. Lyons [1660]. Lyrical [448]. lysogeny [2484].

M [1267, 1637, 1458, 1780, 1505, 1271, 1332, 1101, 1118, 1800, 400, 1252, 1102, 1471, 122, 1485, 1431, 1296, 1522, 1144]. M. [1177, 1226, 1311, 1456, 1039, 1780, 1710]. M.A [524, 758]. M.B.E [745]. M.D [1185, 1400, 1097, 726, 1606, 629, 2695, 900, 2771, 858, 908, 909]. M.D. [564, 216]. M.P [722]. MA [1763]. Mabel [2296]. Macclesfield [1842, 1847]. MacCulloch [873, 900]. mace [423]. MacFarlane [2838, 628, 1714, 1205]. machines [1700, 1733]. Maclaurin [949]. MacLeod [1637]. Macmillan [1146, 1233, 1143, 1228, 1669]. Mad [2678, 1077, 1774, 1123]. Madame [2585]. Maddock [2075]. Maddox [1702]. Made [1999, 2816, 2857, 1542, 856, 2238, 1730, 131, 2473]. madness [1957]. Madras [2601]. Magazine [2311]. Magellan [1108]. Magic [1181, 2714, 2719, 1520, 2109, 1956, 1956]. magisterial [2317]. Magnetic [1170, 2696, 2337, 1610, 1428, 2557, 2432, 2369]. magnetism [937, 2316]. Magnetization [2138]. magnetizers [2367]. magnetron [1852, 1775]. magnets [1628]. Magnificent [2861, 1340]. magnitude [2564]. Mahon [1749]. Maidstone [200]. main [575]. Mainly [2630]. Majesty [307, 311, 397, 574, 223, 2006]. Major [438]. maker [1470, 1233, 2175]. Makers [1273, 2741, 1730, 2787, 712, 1086]. Making [1763, 2697, 2345, 2713, 2810, 1470, 2399, 2058, 2224, 1669, 2369, 1550]. Malady [161]. Malcolm [1182]. males [1255]. Mall [1667]. Mallet [1795]. Malta [988]. Malthus [353]. Man [1293, 2634, 2644, 2660, 1765, 1777, 2716, 2751, 2753, 2758, 2759, 1336, 2776, 2777, 689, 1357, 1211, 1768, 2842, 2854, 2866, 1254, 1182, 1103, 1119, 2019, 1027, 2555, 1287, 2915, 1527, 1749, 327, 625, 1488, 1874, 1234, 1283, 2269, 1299, 2833]. manage [2281]. management [2424]. Managing [2391, 2387]. Manchester [1087, 599, 2884, 1934, 1786]. Mancosu [1427]. manganese [981]. manifest [2094]. Manor [116, 895]. Mantell [200]. manufactures [715]. manuscript [833, 2556, 2532]. Manuscripts [1828, 544, 995, 295, 2540, 566, 557, 2415, 1085, 756, 2255, 2802, 127, 2474]. Many [2168, 1768, 2004, 2167, 1336, 1871, 2088, 2413, 1038, 1906]. map [1826, 900, 1397, 1359]. Mapping [1579]. maps [1313, 1355, 1994]. March [579, 244, 578, 707, 1832]. Marco [1763, 1291]. Marconi [2018]. Margaret [1388, 1138, 1631, 2274, 2294, 2271]. Margarette [1547]. Maria [1334, 2586]. Marie [1070, 1157, 2585, 2739, 1667, 1357]. Marie-Anne [2585]. Marilyn [1138]. Marine [162, 1997, 1739, 187, 512]. mariner [845]. Mario [1475]. Mariscotti [1475]. Maritime [604, 2586, 1455, 2891, 1547, 1633]. Mark [1269, 1705, 501, 455, 531]. Marperger [2398]. Marquess [299]. Marquise [2584, 1581, 1903]. marriage [1983, 122]. Marsden [1739, 599]. Marshall [889]. Marsigli [1355, 996, 1195, 1355, 2848]. Marston [579]. Martin [1206, 1714, 1182, 2172, 2182, 274, 724, 2814, 2058, 2316, 2505, 820, 1283]. Martinus [676, 644]. Martyn [1708]. martyr [431]. Martyrs [2473]. Marum [676, 644]. marvellous [533]. Mary [1185, 1212, 2582, 2572, 1307]. Masers [1713]. Maskelyne [515, 1099, 1722, 2743]. Mason [1209, 212]. mass [1757]. Massachusetts [1138]. Master [1188, 2705, 336]. mastering [2554]. Masters [2864, 1801]. MATERIA [2510]. Material [1572, 1714, 1880, 2537, 1905]. Materialism [2608, 2061, 2389]. materialistic [2606]. Materials [2651, 2443, 1714]. mathematica [783]. Mathematical [2883, 2667, 2671, 1411, 2684, 2718, 1213, 2783, 1189, 1501, 1583, 1780, 1063, 2144, 1427, 2416, 1801, 712, 2260, 2864, 1254, 1324]. Mathematician [2653, 1252, 1518, 379, 494]. mathematicians [1017]. Mathematics [1230, 2741, 1427, 2783, 1189, 1801, 2016, 1041, 2871, 2121, 1607, 2418, 326, 2911, 1782, 1086]. Matter [10, 1, 25, 52, 26, 97, 146, 193, 208, 222, 242, 252, 277, 306, 332, 342, 348, 343, 369, 395, 499, 518, 541, 573, 589, 603, 620, 636, 647, 672, 685, 660, 686, 705, 706, 720, 735, 721, 736, 750, 762, 751, 763, 774, 775, 808, 871, 879, 885, 893, 903, 904, 914, 924, 915, 925, 934, 942, 935, 943, 954, 964, 955, 965, 976, 983, 977, 984, 992, 1000, 993, 1001, 1010, 1011, 1022, 1034, 1012, 1035, 1042, 1058, 1043, 1044, 1074, 1090, 1059, 1091, 1104, 1124, 1105]. Matter [1125, 1147, 1166, 1148, 1167, 1192, 1214, 1193, 1215, 1237, 1256, 1238, 1257, 1278, 1301, 1279, 1302, 1341, 1321, 1363, 1390, 1364, 1391, 1412, 1439, 1413, 1440, 1460, 1476, 1490, 1461, 1477, 1491, 1510, 1530, 1554, 1511, 1531, 1555, 1575, 1600, 1619, 1576, 1601, 1620, 1641, 1657, 1677, 1642, 1658, 1678, 1691, 1717, 1735, 1692, 1718, 1736, 1753, 1770, 1789, 1754, 1756, 1790, 1804, 1820, 1839, 1805, 1821, 1840, 1861, 1876, 1888, 1862, 1877, 1889, 1900, 1911, 1926, 1901, 1912, 1927, 1938, 1952, 1963, 1976, 1939, 1953, 1964]. Matter [1977, 1989, 2002, 2012, 2023, 1990, 2003, 2014, 2025, 2054, 2083, 2096, 2056, 2085, 2098, 2105, 2116, 2129, 2140, 2107, 2118, 2131, 2142, 2153, 2164, 2178, 2187, 2155, 2166, 2180, 2189, 2201, 2215, 2227, 2240, 2203, 2217, 2229, 2242, 2250, 2264, 2276, 2287, 2252, 2266, 2278, 2289, 2512, 2605, 2603, 1506, 2316, 2284, 2602]. matters [2442]. Matthew [2617]. Maull [1898]. Maurice [1784]. maverick [1251]. Max [1675, 1288, 1650, 1643, 1542, 142, 2061, 1846]. Maxwell [749, 1118, 838, 681, 1039, 884, 1020, 2896, 2897, 2898, 2899, 1118, 727, 2915, 1471, 1749, 1385]. May [1248, 753, 2475, 1690, 2812]. Mayer [2158, 653]. Mayfield [1385]. mazes [2500]. McCallum [1502]. McCartney [1705]. McConnell [1273]. McCormmach [1408]. McGuire [571]. McHenry [1119]. McKie [586]. MD [1774, 1696]. mean [2343]. meaning [2546, 2397]. meaningfully [2563]. meanings [1516, 1393, 2426]. Means [2877, 1687]. Measure [2621, 1731, 921]. Measured [1777, 2753]. measurement [1040, 2559]. measurements [1605]. Measuring [1731, 613]. Mecca [1291]. mechanic [882]. mechanics [1583, 919, 941, 1021]. méchanique [1682]. mechanisms [2064]. Medal [236, 794, 1128, 2158, 1176, 2258]. Medalist [1895]. medallions [471]. Medallist [483]. medallists [186]. Medals [960, 667]. Medawar [1616, 2069, 1145, 2442, 2384, 2233, 2326, 1145, 2472, 2788, 2391]. Media [2174]. mediated [1955]. MEDICA [2510]. Medical [2642, 2367, 2756, 2779, 1587, 2582, 2292, 1860, 1596, 1502, 1933, 1430, 2027, 909]. Medicine [2623, 2641, 1212, 1866, 2722, 1331, 1160, 1098, 2867, 1431, 2868, 1727, 2028, 1794, 1848, 2573, 1180, 1947, 2005, 2483, 2181, 1056, 1051, 1651, 2027, 2623, 1316, 2793, 1863, 1431, 1656]. medicines [1884]. medieval [2256]. mee [2413]. meet [2017]. Meeting [114, 1571, 1267, 1987, 1803, 2382, 92, 764, 1946, 1996, 2062, 2196, 2261, 2304, 2339, 2363, 2392, 1135, 1177, 1226, 1311, 1351, 1405, 1452, 1500, 1573, 1640, 1690, 1752, 1859, 2149, 1899, 2114]. Meetings [2345, 165, 30, 195, 160]. Mehra [1362]. Meitner [2531]. Mellon [1779]. Meloni [1456]. Mem [91]. members [92]. Membership [2596, 928, 802, 2033]. Memoir [2887, 874, 1508]. Mémoire [149]. Memoirs [1408]. memorandum [1905]. Memorial [579, 615, 202, 2680, 764, 580, 2851, 591, 776, 142, 577, 707, 831, 708, 1832, 1895, 820, 1222, 1141]. Memories [1934, 436]. memory [622, 803, 1903]. Men [2857, 1431, 1516, 229, 256, 1286, 1730, 667, 2101, 2474, 558]. menace [868]. Mencke [461]. Mendel [537, 506, 234, 536]. Mendelian [805]. Mennim [1273]. Menorca [2586]. mental [863]. Mentors [1903]. Mercator [995]. Merchant [1107, 1603, 1325]. Merck [2599]. Mercury [863]. Mere [1524]. Merlin [1484]. Merope [1649]. Merret [1494]. Mesopotamian [2458]. Message [122, 73]. metabolism [1299]. metalline [613, 257]. Metals [1572, 2822]. Metamorphosis [411]. metaphysical [2879, 2148, 2608]. meteorites [1915, 2238]. meteoritic [2238]. Meteorologist [1525]. meteorology [1739, 1982, 230, 2433]. Meteors [2809, 1522]. method [1002, 2370, 2087]. methodological [2283]. methodology [2444]. Methods [2718, 1501, 1081]. Methodus [2556]. methyl [952]. Miall [930]. Michael [1410, 1798, 491, 1508, 1777, 1082, 1292, 1298, 1116, 1250, 1333, 1545, 1161, 1546, 1425, 1453, 1656, 1308, 2890, 526, 2906, 1270, 2127, 1158]. Michette [1434]. Michigan [1141]. Microbes [1765, 1275]. Microbiology [2769, 2028, 1765, 2222]. Micrographia [2169, 531]. microorganisms [1760]. Microscope [1384, 2829, 1559, 1777]. Microscopes [1755, 1586]. microscopic [1942]. Microscopists [1384]. microscopy [1995, 1741]. microworld [2323]. mid [839, 2553, 2272, 2029, 1933]. mid-eighteenth-century [2272]. mid-nineteenth [839, 2553]. mid-twentieth-century [1933]. Middle [1379, 2715, 2247]. Middlemiss [107]. Middlesex [1144]. Middleton [2440]. middling [2465]. midlife [2472]. migraine [778]. Migrants [1778]. Miles [470]. Military [1483, 1469, 1757, 2799]. Mill [1290]. Millar [1388, 32]. Millennium [1532, 1556]. miller [1252]. Millport [1997]. Mills [2324]. Milne [1080, 2495]. Milo [1143, 1228]. Milton [1071]. Mind [2551, 2837, 2565, 1906, 2607, 2099, 2102, 1957, 1040, 1138, 2101, 1495, 49, 2260]. minded [1205]. minds [2103]. mine [2528]. Mineral [2316]. minerallogenesis [2316]. Minister [400]. Ministry [1206]. Mint [504, 2554]. Miracle [2770, 1802, 1372, 1802]. mirror [1038, 1578]. Mirrors [2137, 2058]. misadventures [2423]. Miscellanea [1993]. Miscellaneous [1828]. Miss [1798]. missing [1225, 2104]. Mistakes [1731]. Mixed [1801]. mobility [2566, 2600]. Model [1539, 2354, 2307, 2048]. modem [1138]. Modern [2628, 1410, 1388, 1379, 2692, 2715, 1704, 2820, 2837, 1474, 2867, 1501, 2598, 1592, 2061, 2452, 2443, 1324, 1343, 1126, 2464, 2611, 2239, 1507, 268, 2224, 2543, 2602]. Modernity [1353]. modes [2553]. Mohammed [1162]. Moivre [1993]. Molecular [1628, 2027, 2064, 1699, 1538, 1846, 891]. molecularity [1467, 1481]. Molecules [406, 1670, 1970, 938]. Moll [1431]. Molyneux [481]. Monceaux [1169]. Monday [403]. Monde [2590]. money [2309, 2554]. Mont [288]. Montagu [1271, 2827, 2582, 1810]. Montpellier [2182]. monument [2110]. Monuments [1586]. Moon [1994, 2453]. Moore [1088, 1230, 1161, 2871]. Moral [1613, 2414, 2358]. Morality [2737, 1674, 1674]. morals [2383]. Moray [392, 1958]. morbus [2146].

Mordechai [1120]. Morehouse [1883]. Morgagni [228]. Morgan [91, 261, 1882]. mori [1262]. Moritz [2456]. Morphology [2749]. Morrell [1435]. Morris [1551]. mortality [833]. Mortimer [613, 177]. Morton [1272]. Morveau [563]. Moseley [792]. moss [2468]. most [1518, 2468, 1270, 2070]. Mother [307, 1030]. mothers [1983]. Mould [2770, 1802]. Moulin [1092]. Moulton [2018, 2260]. Mount [2500]. Mountain [2842, 1780, 1527]. mountaineer [859]. mounting [642]. mourning [2225]. move [2591]. movement [872]. Moxon [1282]. Mozart [1115]. MR [1732, 716, 702, 1296]. Mr. [143]. MR0684478 [941]. MR0782423 [963]. MR0782424 [975]. MRCS [1794]. Mrs [496]. MS [1905]. MSS. [232]. Much [1777, 2751, 2413]. Muddle [960]. multiplex [1851]. mummy [2329]. Murchison [1142, 2845]. Murdin [1406, 1689, 1334]. Murdoch [1234, 2716]. Muriel [902]. Murray [999, 1253]. muscular [432]. muses [1132]. Museum [1068, 1272, 1248, 1212, 1455, 2891, 1589, 1547, 1633, 1056, 1295, 1923, 2522, 2519, 2095, 604, 857, 1243]. Museums [2095, 2435, 2436]. Music [1520, 2714, 846, 1608, 1520]. Musical [72]. musicke [2115]. must [2488]. mutualism [2383]. My [1026, 1882, 781, 362]. mycology [2570]. Myopia [515]. mysteries [1336]. Mystery [1520]. myth [1739, 2065, 2269]. myths [1295].

N [1811, 1339]. N. [2390, 1271]. name [2610, 2546]. Names [1635, 744]. naming [675]. Napoleon [18]. Napoleonic [2430]. Napoli [201]. narcissism [2404]. narrating [2292]. narratives [2542, 2483]. Nathanael [1844]. Nathaniel [1480]. National [112, 1212, 1455, 2573, 821, 1762, 1547, 1096, 1635, 1633, 890, 1596, 1750, 3, 164, 203, 1571, 604, 2891, 2552, 1937, 2435]. NATO [1853]. Naturae [1849, 1613, 2815]. Natural [1068, 1338, 2049, 1248, 2671, 495, 2713, 2714, 2718, 1206, 1429, 1613, 2393, 2785, 664, 1232, 2824, 2426, 2847, 2524, 2874, 1325, 2017, 2414, 882, 1516, 1807, 1550, 1501, 1520, 815, 1589, 2520, 2186, 2587, 2898, 2594, 325, 967, 1471, 2181, 2316, 2450, 2183]. natural-product [815]. Naturalia [2181]. Naturalism [2858, 1785]. Naturalist [2681, 1591, 2830, 2175, 929, 1942, 1767, 834, 2519, 1297]. Nature [1550, 2685, 2894, 2708, 1425, 1636, 1528, 948, 1851, 2280, 2373, 2087, 1955, 655, 2543, 1481, 1425, 2312, 638]. natures [1129]. Naval [1251, 1794, 1271]. navies [1048]. navigation [447]. navigator [608]. Nazi [1616, 2789]. NCUACS [1762]. Neapolitan [1795]. nearly [1156]. Nebula [1943]. Nederlandse [1712]. need [2021, 2347]. Needham [1066, 2800, 2834, 1162]. needles [1170]. Negative [2663, 1073]. neglect [2405]. neglected [1291]. negro [849]. Nehemiah [711, 1097, 907, 2771, 2614]. Neil [1653, 901, 1923]. Neile [384]. Neither [1324, 1343]. nel [1379, 2709]. nella [1379, 2709]. Neptune [1332, 1916]. Nérée [2590]. Nernst [1507, 2628]. nervous [795, 405]. Net [1742, 1829]. Netherlands [1050, 1051]. Network [2510, 2585, 2582, 632, 1623]. networking [2359]. networks [2588, 2587, 2581, 2465]. neurodiversity [2563]. neurophysiology [318].

Neutoni [2556]. never [2211, 2258, 2501]. Nevil [2743, 515, 1099]. Newly [1259, 1604]. News [106, 144, 164, 180, 129, 119, 1716]. newspaper [1887]. Newton [125, 1501, 1782, 1085, 1120, 1231, 1208, 941, 1543, 1382, 1701, 467, 894, 1023, 1004, 1823, 2160, 1018, 2668, 2281, 503, 1168, 2886, 1106, 482, 99, 880, 2691, 2477, 1326, 2697, 1766, 691, 2893, 534, 1032, 2894, 2475, 559, 2718, 2144, 2556, 754, 918, 1061, 2721, 1260, 1303, 2724, 1396, 2725, 1580, 1605, 1020, 596, 1706, 2358, 1316, 1742, 2322, 863, 2416, 1258, 1957, 1013, 1871, 1149, 2911, 2193, 520, 1335, 1368, 2554, 545, 211, 233, 854, 896, 919, 2530, 247, 248, 2802, 2423, 1885, 1822, 2058, 2474, 1054, 1669, 1038, 1578, 947, 1053, 864, 1543, 2618, 2863, 466]. Newton [494, 533, 1067, 1041, 1382, 1829, 1843, 1581, 1706, 1636]. Newtonian [1232, 1026, 1582, 927, 2847]. Newtonianism [1015, 1208, 2476]. next [2232]. Niccolo [1501]. Nicholas [1673, 1353]. Nichols [1315]. Nicolaas [1297]. Nicolas [149, 626]. Nicolaus [723, 995]. Nicolson [1138, 364]. Niels [2813, 1187]. Night [515]. Nikolaeva [496]. Nikolaeva-Tereshkova [496]. Nineteenth [2622, 2513, 2429, 2807, 2576, 2283, 2091, 2360, 839, 2553, 2884, 792, 1794, 2121, 2584, 2448, 2557, 2337, 2366, 1070, 2594, 2344, 2441, 2135, 1786, 1138, 2488, 2433, 1830, 2444, 2426, 2341, 2369, 2590, 2510, 2547]. Nineteenth-century [2429, 2283, 2360, 2884, 2121, 2584, 2448, 2557, 2366, 2594, 2344, 2441, 1786, 2488, 2433, 1830, 2444, 2426, 2369]. ninetieth [130]. niobium [1723]. nitre [1396]. nitrogen [1424]. nitrous [2010]. No [1248, 2837, 1109, 963, 2292, 975, 1138, 941]. Noakes [1782]. Nobel [1765, 1920, 2597, 1809, 2769, 2514, 1584, 1740, 1350, 2136, 1757, 2876]. Noble [2211]. Noddack [2284]. noise [2526]. nomination [2390]. nominators [1920]. nominees [1920]. Non [2638, 1411, 1291, 2021, 2573, 2531, 1916]. non-Darwinian [2021]. non-discovery [1916]. non-human [2573]. Non-Printed [1291, 2638]. non-radiative [2531]. Non-Standard [1411]. Nonstandard [1411, 2684]. nor [1324, 1343]. Norfolk [1393, 1623]. norms [2399]. North [1119, 2776, 2777, 237, 272]. Northern [1526]. Northwood [1144]. Norton [1101, 1611]. nos [1204]. Note [894, 612, 963, 249, 949, 319, 329, 2520, 1605, 274, 805, 127, 320]. note-taking [2520]. notebook [1259]. Notebooks [1409, 2836, 2879, 1069, 472]. Notes [61, 366, 537, 703, 963, 1204, 1591, 98, 575, 226, 975, 2910, 63, 62, 941, 1733, 7, 166, 27, 41, 53, 65, 96, 109, 120, 132, 145, 168, 181, 192, 206, 221, 240, 251, 263, 276, 290, 304, 321, 330, 357, 440, 451, 473, 486, 497, 507, 516, 527, 539, 550, 569, 587, 600, 618, 633, 645, 656, 670, 682, 717, 732, 746, 760, 771, 784, 798, 806, 817, 828, 841, 851, 860, 870, 877, 1308, 581, 339, 153, 11, 1843, 475, 51, 111, 133, 154, 182, 207, 241, 264, 292, 322]. Notes [341, 368, 442, 464, 489, 509, 529, 552, 572, 602, 635, 658, 684, 719, 748, 773, 800, 819, 843, 862, 878, 892, 913, 933, 953, 974, 991, 1009, 1033, 1057, 1389, 1438, 1553, 1618, 1676, 1734, 1788, 2191, 2210, 2082, 659, 1811, 1737, 2396]. Nothing [2629, 1634]. Notice [191]. Notices [179, 205, 220, 239]. Notion [2685, 1425]. Nottingham [1508]. Nova [2322]. November [141, 361, 1571, 1946, 2062, 2196, 2261, 2339, 2363, 2392, 1135, 1177, 1226, 1267, 1311, 661, 777, 752, 313, 574, 821, 577, 308, 1351, 1452, 1500, 1573, 363, 616, 1640, 1690, 1803, 1859, 457, 2149, 2421, 1899, 2114, 490, 722, 82]. nowadays [1516]. NPL [1568, 1570]. NRC [1732]. Ntral [1003]. Nuclear [1724, 1725, 1726, 1508, 690, 696, 404, 409, 2887]. Number [442, 1029, 475, 1535]. Nuncius [1523]. Nurse [2196, 2261, 2304, 2339, 2149].

O [1177, 1226, 1267, 1311]. O.B.E [1163]. O.M [579, 615, 916, 585, 577, 764, 831, 1405, 1452, 1500, 457, 580, 776]. O.M. [317, 398]. oak [1314]. OBE [2885]. Object [2438]. objective [1327]. objects [2436, 2007]. obliging [2579, 2518]. oblivion [2284]. O’Brian [1068]. Observation [2361, 2507, 1755, 2527]. observational [230]. Observations [2069, 833, 856, 495, 2453, 2527, 2474]. Observator [1455, 1334]. observatories [2576, 628, 1968]. Observatory [1722, 433, 2576, 2299, 125, 789, 1246, 790, 2500]. Observing [2257]. obsession [1971, 1340]. Obsessions [2861]. obstacle [2484]. occasion [421, 422, 2070, 789, 122]. occupants [1204]. oceanic [2454, 45]. oceanographer [951]. Oceanographical [24]. Oceans [1547, 2914]. October [615, 764, 831, 708, 2773, 1849, 2609, 580]. Odyssey [2621, 2684, 1731, 1411]. off [1992, 2209]. Office [2177]. officers [69, 78]. official [424]. Ogilvie [1138]. O’Hear [1387]. oil [2541]. Old [2630, 1317, 1251, 2311, 2038, 1533, 1745, 864]. old-established [2311]. Oldenburg [387, 1667, 1464, 138, 479, 581, 1075, 2726, 907, 1773, 1037, 1956]. oldest [1950, 2020]. Oldfield [2546]. Olenus [1638]. Oliver [2311, 1121, 2828, 2921, 1360]. Olivier [1711]. Ollard [1159]. Olschki [1379, 1523]. Olson [1522]. OM [1946, 1996, 2062, 1640, 1690, 1752, 1803, 1859, 2114]. omissions [2530]. One [2745, 42, 1823, 1349, 1469, 1366, 1330, 1403, 1467]. online [2523, 1850, 1896]. Onnes [2913, 1712]. Onslow [902]. Open [2636, 1071, 1207, 1548]. Opening [604, 574, 1569, 397, 398, 1201, 268, 2492]. opening-up [2492]. Opera [56]. operation [112, 2237]. operator [283]. Ophthalmoscope [71]. opinion [910]. Opposite [1428]. optical [559]. Opticks [2358, 2530, 1054]. Optics [2646, 1519, 1164]. Optiks [559]. oratory [793]. orbits [1368]. Ordnance [900]. organic [2033, 410]. organism [2048]. organization [1513, 922]. Organizations [2607]. organized [203]. organs [2311]. orientation [1724, 1725, 1726]. origen [1475]. Origin [1655, 1913, 460, 2425, 329, 1851, 2546, 2282, 1178, 2285, 1893]. original [2422, 1085, 2800]. Origines [532]. Origins [2837, 2858, 1785, 2080, 1176, 594, 593, 1138, 371, 1424, 592, 989, 434, 2619, 126, 74]. Orkney [741]. Ornithological [2521]. Orthodoxy [2750, 1333]. Osborne [1470]. O’Shaughnessy [1400, 959]. Osler [1656, 1631, 2641]. Other [1710, 2844, 89, 106, 144, 164, 180, 295, 1679, 1394, 129, 119, 1947, 1505, 756, 2578, 302, 1052, 105, 128, 167, 2009, 909, 1029]. otherwise [1516]. Ottawa [1732]. Our [1455, 1462, 2765, 2766, 1049, 1334, 1416]. Outram [1138]. Outrunning [1255]. Outsider [2873, 1317, 1597]. Outstanding [2752, 1704]. Outward [2377]. owed [2263, 2245]. Owen [1248, 1297, 859, 1639, 2717, 250, 2830]. Owens [2208]. Own [1546, 1349, 1495]. OxDNB [1745]. Oxford [1272, 1656, 1181, 1145, 1299, 1542, 1673, 1746, 1780, 1429, 1685, 1231, 1714, 1427, 1208, 1548, 1335, 1667, 1122, 1210, 1294, 1640, 1752, 1803, 1859, 1551, 1633, 1549, 1435, 1205, 1211, 1598, 1187, 1488, 1632, 1713, 1543, 1362, 1383, 1711, 1339, 433, 1244, 1482, 1626, 1724, 1725, 1726, 654, 1744, 791, 865, 1848, 2407, 710, 594, 593, 649, 1690, 2795, 858, 1808, 160, 2052, 2330]. oxide [2010]. oxygen [1294, 2236].

P [1456, 1185, 2680, 2122, 1118, 2331, 400, 1409, 378, 1339, 537, 2836]. P. [2680, 1427, 1471, 2398, 1339]. p.m [405, 409]. P.R.A. [701]. P.R.S. [1405, 1452, 1500, 274, 273, 383]. PA [1769]. Pacific [1547, 2914]. page [475]. pages [1547]. Paget [1160, 1160, 2826]. paintings [578]. Pais [1187, 1632]. Pak.Rs [1162]. Pakistan [1162]. Palace [1864, 1863]. palaeolithic [31, 2247]. palaeontological [2599]. palaeontologist [874]. palaeontology [1560]. Pall [1667]. Palmyra [2335]. Pan [545]. Pandora [2698, 1783]. Paper [1627, 2801, 1200, 1248, 2522, 2541, 805, 1536, 2519]. Paperback [1318, 1331, 1340, 1068, 1595, 1550, 1765, 1508, 1766, 1639, 1525, 1706, 1548, 1635, 1784, 1687, 1551, 1783, 1549, 1611, 1319, 1546, 1769, 1801, 1425, 1713, 1276, 1651, 1674, 1382, 1253, 1295, 1360, 1353]. Papers [2623, 2896, 2897, 2899, 2214, 2773, 1189, 970, 2111, 1455, 1648, 30, 102, 1118, 1073, 796, 2920, 1921, 1089, 1974, 1056, 1474, 2074, 1083, 1909]. Papilionoidea [2467]. Papilionum [2467]. Papin [118]. parabolical [1395]. paradise [1328]. Paradox [1488]. Paragon [1749]. parallax [921]. Paris [141, 970, 1969, 532, 1968, 2449]. parish [833]. Park [2578]. Parker [1182]. Parkin [1613]. Parkinson [2610, 899]. Parsons [202]. Part [1469, 1483, 1724, 1725, 1726, 1366, 1394, 1415, 1661, 1683, 2494, 1330, 1352, 20, 35, 209, 227, 246, 270, 283, 480, 513, 2493, 2799, 1403, 1423, 1467, 1481]. participation [2344]. Particular [2410]. partition [1156]. Partnership [2742, 1798]. Parts [1768, 1068]. Pasachoff [1522]. Pasadena [2500]. Passages [2627, 1296, 1296]. Passion [2865, 2916, 1598]. Past [1609, 1435, 1235, 1841]. Pasteur [922]. Paston [1347]. patent [627, 1350]. patient [2564, 2563]. patients [2223, 2009]. Patriarch [1380]. Patricia [1138, 1276, 1428, 1783, 1669]. Patrick [1068, 1360, 1469, 1483, 1451, 2617]. patron [1810, 397, 574]. Patronage [2448]. Patronus [372]. Patsy [1358]. Paul [1068, 2070, 2261, 2304, 2339, 1249, 1069, 1097, 417, 2149, 2236, 384, 2196, 2680]. Pauling [1955]. Paulze [2585]. Paulze-Lavoisier [2585]. pay [2354]. pay-to-publish [2354]. Payen [1914]. pbk [1180]. PD [1360]. Peace [2738, 1179]. Peacock [1078]. Pearls [1110]. Pearson [1696, 1800, 1223, 2821, 1265]. pedagogical [2398]. Pedestrian [1318]. peer [2312, 2545]. peers [2293]. Peierls [1474, 2920]. Pell [1533]. Penelope [1520]. Penguin [1595, 1433, 1592, 1529, 1123, 1086]. Penicillin [1962, 2770, 1802, 1072, 1468]. Penjerrick [2528]. Pennsylvania [12]. Penny [1709]. Penrose [1403, 1423]. Pentland [1502]. People [2765, 2766, 2436, 1698]. Pepys [1680, 478, 79, 219, 16, 971, 1159, 2808]. perception [803]. Perceptions [2688, 1232, 1295]. Percussion [2138]. perfect [2047]. performance [2221]. perils [2308, 2005]. Perinaldo [1380]. period [2552, 902, 2492, 480]. periodical [2308, 2343, 2307, 1830]. periodicals [2352, 2306, 2341]. periphery [2338]. perish [2452]. Permission [1023]. Perrault [2186]. Perrin [757, 152]. Persia [856]. Persian [2488]. Persist [2736, 1614]. Person [2466, 1518]. Personal [2646, 2678, 1411, 2684, 689, 1995, 1699, 1164, 1123, 436]. personality [1258]. Perspective [2878, 1995, 2352, 2053, 1211, 2515, 1293]. Perspectives [1188, 2705, 2725, 1455, 2924, 1543]. Perutz [1542, 1846]. pesanteur [150, 149]. Pesticides [2376]. Peter [1727, 2233, 1145, 1360, 2391, 1211, 1546, 2325, 1121, 2788, 1281, 2151, 2456]. Peterhouse [1538]. Petersburg [1721]. Petiver [2463, 1880, 2462, 2469, 2468, 2465, 2464, 2466, 2467]. Petroleum [416]. Petty [376]. Pfauntsch [1434]. Phage [2486]. pharmacological [1910]. Pharmacology [1921]. Phelp [757]. Phenomena [2692, 1410]. phenomenon [1941]. Philadelphia [1071, 1705, 1408, 1769, 1459]. Philanthropist [1107]. philanthropy [2615]. Philately [1134, 1448]. Philip [1150]. Phillip [1248]. Phillips [526, 1138, 1144]. Philomaths [2269]. Philosopher [2627, 2890, 1429, 1387, 1296, 882, 1546, 2336, 1296]. Philosophers [1706, 1018, 666]. Philosophical [66, 2311, 160, 1408, 1769, 1919, 2262, 1822, 511, 1115, 2335, 969, 2309, 72, 2478, 531, 1967]. Philosophicall [881]. Philosophie [1208, 2863]. Philosophy [2671, 2718, 1387, 2783, 2785, 1157, 1232, 2918, 2813, 1187, 1425, 2847, 2414, 1501, 715, 2709, 1867, 1607, 1002, 2898, 2358, 325, 2442, 664, 1471, 2419, 2418, 2338, 2527, 1206, 1427, 2476]. Phlogiston [1769, 1294]. Phocaena [2170]. Phoenix [1748]. phonetic [2388]. Photographic [1409, 1898]. photographing [2087]. Photographs [1975]. Photography [1409, 2835, 2836, 1184]. photosynthetic [1919]. Phrenology [2858, 2389, 1785]. Physic [1197]. Physical [1571, 2628, 2414, 1121, 1937, 890, 2828, 1244, 863, 2087, 1507, 2331]. Physician [814, 629, 1466, 1606, 2900, 1098, 1904, 950, 2175]. Physicians [1209, 1794, 2009, 568]. Physicist [2653, 1252, 1824, 1213]. Physicists [2916, 696, 679, 1632, 1705]. Physics [1005, 2662, 2664, 1406, 1689, 1835, 1020, 1705, 690, 1334, 1189, 1672, 2514, 2813, 1749, 1187, 1617, 1408, 1546, 2841, 1459, 1597, 1474, 2869, 1482, 1972, 1583, 405, 2767, 1136, 1473, 2864, 40, 1584, 1920, 1340, 1611, 1801, 2061]. physiocratic [2583]. Physiological [1329]. Physiologists [2640, 1936, 1255]. Physiology [2640, 2220, 989, 1079]. Piccadilly [707]. Pickering [1292, 2151, 1296, 1923]. Picking [2112, 2303, 2128]. Picture [2152, 156, 2606]. Pictures [1235]. Picturing [2123]. piece [2104]. piezoelectricity [2159]. Pigott [1480]. Pilgrim [2]. pilot [170, 2492]. Piltdown [2833]. Pimlico [1783]. pin [1039]. pin-jointed [1039]. pink [2053]. Pinner [1766]. Pioneer [2681, 492, 1409, 815, 582, 1767, 2482, 598, 31, 940, 1339]. Pioneering [1107, 1670, 987, 1942]. Pioneers [2730, 1264, 2769, 1474, 1635, 1733, 1700, 1765]. Pipes [545]. Piscium [1515]. Pitt [1269, 2643, 1269, 2643]. Place [1783, 2583, 2365, 268, 2399, 1709]. Places [2514, 2569]. placing [2008]. Plague [2471]. plain [624]. plan [174, 2489, 946]. Planck [2212, 142]. Plane [2667, 1189, 927]. Plant [1880, 1971, 404, 2151]. Planta [245]. Planting [1919]. Plants [2150, 1565, 1880, 1068, 2059, 2603, 2468, 2151, 1919]. Plasma [2662, 1189]. plasmon [2162]. Plateau [2492]. Plates [1875, 1887, 1819, 1860, 1455]. platinum [566, 2212]. PLATO [1853]. Playfair [664]. Pleistocene [2578]. Plot [967]. Plymouth [1317]. pneumatic [2006, 824]. Pnina [1138]. poem [117]. Poet [1210, 986, 448]. poetic [1261]. Poetry [1705]. Point [1410, 2692, 2060, 1811]. pointed [740]. Poison [2376]. poisoning [863]. Poisson [2194]. Polar [2526, 2674, 2436, 1593]. Pole [792]. polemic [2606]. Polite [2706, 1337]. Political [2009, 495, 2052, 1740]. politician [1924]. Politics [2815, 67, 2133, 1613, 2417, 664, 2850, 1541]. Polity [2813, 2419, 1187]. Polkinghorne [1436]. pollution [832, 2376]. Polymath [1707]. polymerization [1450]. Polynesia [1817]. Popham [1251]. Popkin [1636]. Popper [1387, 2233, 2918]. Popular [1710, 2553, 2410, 2342]. popularization [2233, 2045, 2590]. popularizer [901, 2572]. Popularizers [2344]. pores [2612]. Portents [1522]. Porter [1592, 1458, 1715, 1800, 2303]. portfolio [2541]. portion [2366, 1919]. Portland [1255, 1417]. Portrait [39, 2630, 916, 1881, 1635, 2919, 1596, 1632, 1898, 32, 612, 1369, 1869, 2275, 471, 562]. Portraits [163, 2756, 1221, 1596, 1522, 81, 2169, 617, 2066, 2139, 2213]. Portsmouth [1023]. Portugal [2235, 2584, 2552, 2126]. Portuguese [2146]. possession [295, 854]. post [2406, 1219, 1345, 2608, 2308]. post-Galilean [1345]. post-Kantian [2608]. Posthuma [56]. potential [2351]. Pour [137, 1208]. Powell [409]. Power [2698, 2763, 2764, 492, 404, 1709, 1783, 2324, 1386, 1210]. Pozzo [1792]. Pp [1068, 1182, 1209, 1293, 2385, 1272, 1727, 1338, 1656, 1456, 1475, 1616, 1088, 1235, 1300, 1269, 1069, 1117, 1185, 1595, 1071, 1181, 1636, 1146, 1097, 1248, 1729, 1410, 1212, 1638, 1572, 1594, 1184, 1251, 1318, 1455, 1550, 1798, 2262, 1145, 1299, 1528, 1388, 1763, 1380, 1379, 1523, 1501, 1689, 1542, 1702, 1765, 1785, 1433, 1361, 1103, 1508, 1428, 1163, 1673, 1520, 1592, 1671, 1655, 1637, 1731, 1782, 1766, 1653, 1188, 1100, 1589, 1639, 1524, 1654, 1458, 1190, 1591, 1119, 1746, 1780, 1121, 1085, 1082, 1141, 1159, 1180, 1206, 1292, 1337, 1314, 1429, 1331, 1470, 1525, 1386, 1297, 1593, 1705, 1457, 1547]. Pp [1685, 1688, 1706, 1096, 1387, 1707, 1704, 1213, 1505, 1120, 1231, 1434, 1271, 1332, 1384, 1714, 1613, 1767, 1179, 1427, 1160, 1101, 1118, 1140, 1298, 1313, 1359, 1484, 1207, 1270, 1336, 1548, 1635, 1730, 1784, 1098, 1634, 1687, 1360, 1667, 1252, 1786, 1122, 1116, 1138, 1157, 1210, 1250, 1294, 1333, 1315, 1355, 1748, 1073, 1099, 1102, 1551, 1334, 1487, 1189, 1633, 1142, 1709, 1232, 2151, 1504, 1672, 1541, 1629, 1783, 1527, 1529, 1290, 1549, 1353, 1539, 1357, 1435, 1471, 1489, 1545, 1615, 1710, 1749, 1205, 1611, 1123, 1161, 1143, 1211, 1228, 1598, 1768, 1187, 1617, 1408, 1432]. Pp [1473, 1488, 1546, 1526, 1507, 1596, 1632, 1712, 1769, 1750, 1801, 1087, 1186, 1425, 1409, 1436, 1485, 1502, 1708, 1453, 1631, 1669, 1670, 1713, 1802, 1086, 1543, 1459, 1597, 1430, 1362, 1162, 1544, 1474, 1732, 1234, 1291, 1358, 1385, 1383, 1276, 1651, 1711, 1614, 1674, 1431, 1296, 1183, 1522, 1382, 1144, 1339, 1295, 1083, 1254, 1273, 1275, 1229, 1316, 999, 1233, 1950, 1056, 1923]. Pp.417 [1319]. Pp.viii [1800]. Pp.xi [1230]. Pp.xlvii [1406]. ppbk [1270]. Practical [1230, 1429, 1919, 2081, 2871, 1196, 989]. Practice [1934, 2783, 1651, 2867, 2520, 2206, 1427, 2464, 1973, 2058]. Practices [2696, 2547, 1428, 2452, 2331]. Prafulla [2390]. praise [558]. Pre [1273, 524, 2453]. Pre-eminently [1273]. pre-history [524]. pre-telescopic [2453]. precarious [2440]. precise [1273]. precision [2916, 2559]. Prediction [753]. Predictions [1563]. Preface [1322, 370, 448]. Preference [2788, 1145]. prehistoric [2099]. Prehistory [1759, 1628, 2246, 1093]. preliminary [1085, 2802, 127]. Première [416]. preparations [2443]. prescient [2064]. Present [2779, 1609, 1841, 1502]. Presentation [85, 2416, 1910, 2295]. presentations [605]. présenté [149]. presented [421, 422, 738]. preservation [544, 2537, 2542, 2452]. Preserve [2452]. Preserved [1212, 2740, 2539]. Preserving [1762, 2543, 2538]. presidency [289]. President [821, 73, 244, 1946, 1996, 2062, 2196, 2261, 2304, 2339, 2363, 2392, 1135, 1177, 1226, 1267, 1311, 1444, 396, 1810, 308, 360, 398, 1351, 1405, 1452, 1500, 1573, 1640, 1690, 1752, 1803, 1859, 2149, 1216, 1899, 2114]. Presidential [81, 2133]. presidents [2150, 824, 2793, 1431]. Press [1272, 1338, 1656, 1088, 1300, 1269, 1069, 1117, 1071, 1729, 1184, 1251, 1455, 1550, 1145, 1299, 1528, 1388, 1379, 1501, 1542, 1411, 1508, 1428, 1673, 1520, 1592, 1782, 1766, 1653, 1100, 1458, 1591, 1119, 1746, 1780, 1121, 1085, 1082, 1066, 1141, 1206, 1230, 1337, 1314, 1429, 1525, 1386, 1297, 1457, 1547, 1685, 1688, 1706, 1387, 1707, 1213, 1505, 1120, 1384, 1179, 1427, 1101, 1118, 1140, 1359, 1484, 1548, 1784, 1800, 1687, 1335, 1667, 1252, 1122, 1116, 1138, 1157, 1250, 1333, 1355, 1748, 1073, 1099, 1102, 1551, 1487, 1189, 1633, 1142, 1255, 1504, 1529, 1549, 1353, 1357, 1435, 1471, 1710, 1205, 1319]. Press [1598, 1768, 1187, 1617, 1432, 1473, 1488, 1526, 1507, 1632, 1801, 1087, 1425, 1409, 1436, 1485, 1502, 1453, 1631, 1670, 1713, 1543, 1362, 1474, 1732, 1358, 1383, 1276, 1651, 1711, 1431, 1522, 1382, 1339, 1083, 1356, 2146]. Presses [1186]. Previous [1232]. Previously [2124]. Price [582, 1295, 2509, 2256]. Priestley [32, 1056, 1261, 2623, 936, 301, 249, 302, 1222]. primate [2573]. Prime [400]. primeval [2102]. primitive [2101]. Prince [994, 1008, 1316, 445]. Princess [122]. Princeton [1411, 1428, 1800, 1617, 1485]. Principe [1485]. Principia [1501, 1335, 523, 247, 1018, 2668, 2671, 1106, 2477, 1017, 783, 2718, 1020, 1019, 1013, 1149, 2802, 1021, 1016, 1093, 1581, 1701]. principle [626]. Principles [2671, 1433, 2778, 234]. print [2121, 2345, 2307]. Printed [1291, 2638, 2016]. Printers [956, 990]. printing [2801, 888]. printmakers [2599]. prior [2571]. priority [2283, 2198, 2457]. prisms [896]. prison [328]. Private [1272, 1249, 2735, 2796, 479, 2408]. Prize [1809, 2514, 1584, 483, 1765, 2769, 1787, 2577, 1920, 1757, 2876]. Prize-awarded [2577]. Prize-winners [1757]. probable [863]. probing [2393]. Problem [521, 1504, 159, 156, 2540, 1826, 864, 1306]. Problems [2664, 1803, 1189, 2918, 907, 408, 1535, 1387]. Proby [874]. procedures [1003]. Proceedings [2908, 1228, 2923]. process [2375]. Proctor [2572]. prodigy [1254]. produced [2541]. producing [2581]. product [815]. production [927, 2260, 2532]. Professional [2873, 2599, 1597]. professionalizing [907]. Professor [402, 433, 461, 1907, 637, 758, 725, 586, 409]. Professors [2911, 1626, 1782, 1782, 664]. professorship [2092, 2336]. Profile [1714]. Profiles [2793]. profit [2399]. profits [2308]. Programme [2177, 966]. programmes [419]. Progress [1637, 1560, 1463, 45, 1724, 1725, 1726]. progressive [2100]. Project [1813, 1666, 2524, 2523, 1779]. Projects [2398]. promise [2535]. Promoter [1487, 1062]. Promoting [1002, 2720, 1157]. proper [744]. property [21]. Proposed [401, 559, 400]. prosecution [1003]. prospective [1742]. Protestant [1126]. Proteus [1534, 1535]. Prout [629, 630]. Proved [2854, 1768]. proves [971]. provided [1919]. provincial [1445, 2579, 1283]. Proving [2337]. provision [2434]. PRS [1946, 1996, 2062, 1832, 2114]. psychoanalytic [2302]. psychological [2011]. pt [1769, 2768]. Public [1338, 2039, 2712, 1356, 2796, 2847, 2874, 2345, 1047, 1232, 1895, 1272]. Publication [2580, 1068, 783, 2556, 1604, 2351, 1644, 235, 531, 674, 191]. Publications [1248, 1798, 1763, 1360, 1887, 1615, 1819, 420, 1509, 1875, 2124, 585, 179, 2345, 1533, 166, 205, 220, 239]. publicity [2308]. publick [1219]. Publish [2545, 2354]. Published [1212, 999, 1056, 969, 663, 1253]. Publishers [1318, 1271, 1430]. Publishing [1317, 1069, 1275, 1406, 1689, 1785, 1671, 1188, 1334, 1672, 1527, 1708, 1459, 1544, 1474, 2310, 2615, 2350, 2348, 2050, 2344, 2170, 2307, 1906]. pump [2171]. purchased [2161]. purchases [1847]. Pure [1196]. Purefoy [2296]. Puritan [1727]. purposes [1049, 2907, 1750]. purse [231]. Pursglove [1508]. pursuing [1054]. Pursuit [2678, 1347, 948, 1123]. pursuits [2586]. Pushing [1629]. put [1442, 1481, 2523]. putrefaction [2222]. Putting [1635, 2595]. Puy [137]. Pyke [1616].

Q [1272, 1409, 2836]. Qinghai [2492]. Quakers [1367, 804]. qualifications [928]. Quantifying [1410]. quantitative [1741]. Quantum [2646, 2767, 2869, 1164, 1611, 1755, 1617, 1955, 503]. quater [501]. quater-centenary [501]. Queen [307, 397, 2727, 1336, 311, 2587]. Quest [2849, 1594, 2598, 2530, 2822, 1485]. question [2457]. questions [2414, 1481]. Quiller [1100]. quincentenary [723]. Quinn [1925].

R [1293, 745, 1572, 2390, 472, 1163, 1458, 1780, 1141, 1714, 1506, 2762, 1102, 1527, 1357, 1056, 378, 1474]. R&D [2029]. R. [1638, 1223, 1591, 1164, 805, 1502, 1265, 2319, 326]. R.G.W [1212]. R.N [622, 745, 898]. R.S. [91, 917, 707, 591]. Radar [2646, 2775, 1776, 1164, 1190]. Radiative [2663, 2531, 1073]. Radical [973, 2630, 1317, 962]. Radicals [2008]. Radio [2646, 1360, 1378, 1944, 407, 1164, 2454, 2921, 1360]. radio-frequency [1378]. radioactive [840]. Radioactivity [2648, 1671]. radiometers [2094]. radium [2273, 2161]. Rae [1318, 2825]. Rahn [1773]. rain [561]. rain-gauge [561]. Rainbow [894, 895]. Ralph [1173]. Ramakrishnan [2363, 2392]. Raman [2001, 1740, 1036, 2063, 1096, 1498, 1985, 2162, 2860]. Ramanujan [1268, 2758, 2759, 1224, 1254]. Ramsay [254]. Ranging [1675, 1650, 1825]. Raphson [1076]. Rasit [2598]. ratio [2194]. rational [1981]. Rattansi [571]. Rawlins [2555]. Rawlinson [1141, 1221, 2851]. ray [1609, 2680, 2560, 1339, 1516, 2390, 1493, 2255, 2182, 1514, 2184, 985]. Raymond [1463]. Rays [2648, 2682, 1671, 2917, 1459, 655, 1434]. Re [2681, 2533, 1767]. Re-examination [2681, 1767]. re-told [2533]. Reactions [2730, 1635, 176]. read [2565]. Reader [2668, 2580, 1335]. readers [2561]. Readership [865]. Reading [2718, 30, 1501, 2565]. reagents [2540]. Reaktion [1596]. Real [2877, 2388, 1687]. realized [2403]. realm [1316, 2293]. reappearance [2422]. Rear [1251, 2819]. Rear-Admiral [2819]. reason [2100, 2900, 1098, 1261]. reasoning [2103]. Rebuilding [2676, 1777, 1415]. Received [2685, 1425]. receiving [2210]. Reception [124, 543, 622, 788, 455, 937, 1896]. Receptions [412, 477, 500]. recipients [1176]. Recollection. [2001]. Recollections [1451, 693, 689, 1227, 902, 1585, 1776, 1699]. recommemorated [1315]. Recommendation [40]. reconceptualizing [2134]. Reconsidered [2904, 1280, 2010, 1250]. reconstruction [2256]. Record [2760, 46, 1138]. Recording [1306]. Records [475, 537, 963, 975, 941, 1733, 2836, 544, 472, 557, 98, 2575, 1409, 51, 111, 133, 154, 182, 207, 241, 264, 292, 322, 341, 368, 442, 464, 489, 509, 529, 552, 572, 602, 635, 658, 684, 719, 748, 773, 800, 819, 843, 862, 878, 892, 913, 933, 953, 974, 991, 1009, 1033, 1057, 1389, 1438, 1553, 1618, 1676, 1734, 1788, 2191, 2210, 2082, 659, 1811, 1737, 2396]. red [1347]. redevelopment [1796]. rediscovered [2093, 2255, 2315, 1914]. Rediscovery [1974, 2145, 2335, 2286]. redivived [1116]. reductionism [2338]. Reed [825]. reef [1907]. reeling [2768]. Rees [1946, 1996, 2062]. referee [1200, 2121]. refereeing [2331, 805]. Reflecting [2087, 1168, 854, 611]. Reflections [1361, 693, 2852, 1114, 1144]. Reflexive [2459]. Reform [2091, 1206, 2785, 2868, 1727, 1823, 2360, 2492]. reformation [1126]. Reformer [901]. regain [1328]. regime [2552]. regional [2369]. Register [50]. registering [883]. registers [833]. Regnault [1975]. regulation [679].

Reid [988, 1248]. Reilly [1233]. reimagining [2342]. reinvented [2539]. Reissued [1382]. rejection [484, 980]. relating [179, 160, 166, 191, 205, 220, 239]. relation [524, 666, 503]. Relations [2722, 1316, 2192, 809, 1065, 229, 700, 563, 55]. Relationship [641, 568, 1721, 2223]. Relationships [1793]. relatives [805]. Relativistic [2666, 1189]. relativity [797, 2533]. releasing [952]. Relics [1399]. Relief [172]. Religion [1181, 2894, 2719, 2731, 2736, 1353, 2815, 1636, 2318, 2235, 1707, 1613, 2417, 2410, 664, 2134, 1614]. Religious [1379, 2715, 2456]. relocating [2538]. Remarkable [2770, 2518, 1942, 1684, 1802]. Remarks [335, 1046, 1014]. remember [316]. remembered [2294]. remembering [2463, 1172]. reminiscence [1137]. Reminiscences [599, 1268, 2738, 1247, 1179, 1330, 1352, 1081, 1178, 1136, 1227, 1312, 1155, 1114, 362, 1289, 1029]. René [183]. renewal [2028]. renewing [2192]. Report [28, 1778, 1666, 1779, 1796, 1762, 1797, 2360, 1195, 1919, 1745]. REPORT. [1744, 1743, 1742]. Reports [165, 922, 674, 2121, 888]. repository [947]. representation [2553]. Representations [2732, 1319]. representatives [90, 424]. Reproduced [1023]. Republic [1384, 2477, 2829, 2579, 2473]. reputation [225]. Requisite [1003]. Rescaling [2349]. Research [147, 963, 932, 759, 485, 2076, 753, 2384, 238, 2360, 1196, 2296, 907, 1931, 2353, 922, 2577, 2236, 2391, 2563, 2032, 1933, 1375, 909]. researchers [2561]. researches [462, 1942, 2010, 1890]. researching [1943]. reserve [2499]. residuary [246]. resolved [1370]. resonating [938]. resource [1463]. respectable [1077]. response [1978]. Restless [2694, 1383]. Restoration [1230, 1613, 2871, 1180, 1844, 1230, 2815]. restored [1746]. results [2533, 610]. Rethinking [1883, 2811, 2446, 1631]. Retirement [143, 659, 716, 702]. return [1879, 2600]. returns [2046]. Reuleaux [1733, 1700]. Rev. [731, 1217]. Reverend [2070, 373, 968, 1497, 825, 374]. Review [2385, 1272, 1235, 1300, 1317, 1595, 1594, 1318, 1273, 2262, 1275, 1299, 1501, 1592, 1766, 1188, 1589, 1229, 149, 1591, 1230, 1314, 1593, 1096, 1213, 1231, 1271, 999, 1070, 1298, 1313, 1233, 1270, 1138, 1294, 1950, 1232, 1319, 1598, 1056, 1596, 1597, 1234, 1276, 1296, 1295, 1923, 1727, 523, 2312, 1656, 1456, 1475, 1616, 1269, 1636, 1729, 1410, 1638, 1572, 1455, 1550, 1798, 1528, 1388, 1763, 1380, 1379, 1406, 1523, 1689, 1542, 1702, 1765, 1785, 1433, 928, 1411, 1361, 1508, 1428, 1163, 1673, 1520, 1671, 1655, 1637, 1731, 1782, 1653, 1639, 1524, 1777, 1654, 1458, 1746, 1780, 1429, 1470, 1525, 1386]. review [1705, 1457, 1547, 1685, 1688, 1706, 1387, 1707, 1704, 1505, 1588, 1434, 1384, 1714, 1613, 1767, 1179, 1427, 1359, 1484, 1506, 1548, 1635, 1730, 1784, 1356, 1800, 1634, 1687, 1360, 1667, 1786, 1355, 1748, 1551, 1487, 1633, 1709, 1504, 1672, 1541, 1629, 1783, 1527, 1529, 1290, 1549, 1353, 1539, 1357, 1435, 1471, 1489, 1545, 1615, 1710, 1749, 1205, 1611, 1768, 1617, 2545, 1408, 1432, 1473, 1488, 1546, 1526, 1507, 1632, 1712, 1769, 1750, 1801, 1425, 1409, 1436, 1485, 1502, 1708, 1453, 1631, 1669, 1670, 1713, 1802, 1543, 1459, 1430, 1362, 1544, 1474, 1732, 1358, 1385, 1383, 1651]. review [1711, 1614, 1674, 1431, 1522, 1382, 1701, 2331]. Reviews [1209, 1249, 1069, 1146, 1097, 1251, 1085, 1332, 1252, 1116, 1157, 1250, 1333, 1334, 1189, 1142, 1187, 1086, 1253, 1339, 1083, 1182, 1293, 1728, 1612, 1630, 1747, 1764, 1781, 1799, 1338, 1088, 1117, 1185, 1071, 1181, 1248, 1212, 1184, 1145, 1381, 1407, 1703, 1103, 1521, 1100, 1590, 1190, 1119, 1121, 1082, 1066, 1141, 1180, 1206, 1292, 1337, 1331, 1297, 1686, 1120, 1160, 1101, 1118, 1208, 1140, 1207, 1336, 1098, 1164, 1335, 1668, 1122, 1210, 1073, 1099, 1102, 1255, 1354, 1540, 1472, 1123, 1161, 1143, 1211, 1228, 1087, 1186, 1426, 1486, 1503, 1454, 1162, 1291, 1652, 1183, 1340, 1144, 1254]. revisions [1212, 1581]. Revisited [2728, 1802, 2158, 1694, 2302, 2234, 1228, 1685]. revival [596]. revivified [1743]. revolt [2035]. Revolution [1342, 2693, 2707, 2723, 2905, 2811, 1051, 1453, 1631, 1930, 2281, 1379, 2102, 1108, 1420, 1812, 1356, 1294, 1487, 1382, 1631]. revolutionary [2381]. revolving [1368]. reward [1128, 667]. reworking [2557]. Reynolds [701, 1129]. Rhetoric [1232, 2847]. rhetorician [1467, 1481]. Ricci [1456]. Rice [1300, 1589, 2923]. Richard [1077, 1338, 1616, 1069, 1636, 1782, 1613, 1786, 1315, 1362, 436, 2049, 1248, 1794, 2509, 641, 948, 2717, 1159, 250, 526, 1297, 1688, 582, 1198, 2123, 2207, 62, 2790, 2815, 2830, 1919, 2851, 1183, 1141]. Richards [1318]. Riera [1880]. Right [2070, 1262, 2035, 373]. Righting [2760, 1711, 1138]. rights [1846]. riots [301]. Rise [2672, 2895, 1297, 1207, 1232, 2826, 2847, 1160, 2481, 520, 1551, 1801, 2864]. rivalries [1739]. river [2248]. riverbank [2385]. Rivers [1269, 2643, 2485, 2643]. ro [191]. Road [1251, 577, 2027]. Roam [1548]. Robert [1203, 1318, 1744, 1347, 1777, 381, 353, 1743, 1208, 392, 1315, 1925, 1958, 1485, 776, 1383, 981, 1254, 1739, 382, 1107, 882, 2637, 336, 365, 1720, 2124, 2597, 2262, 1366, 1394, 1415, 2676, 2685, 2607, 2694, 1795, 1603, 1777, 1760, 743, 2528, 912, 196, 1292, 1905, 2494, 1346, 337, 2903, 2748, 2904, 1604, 1797, 2073, 2574, 2751, 2753, 2214, 2207, 2323, 209, 227, 246, 270, 283, 338, 480, 513, 1122, 1116, 967, 1908, 1733, 1700, 2493, 1217, 2803, 881, 1003, 2612, 2512, 1914, 2822, 1902, 95, 2411, 139, 2127, 1769, 902, 2375, 803, 1060, 2618, 2110, 1614, 1674, 2870]. Robert [2456, 1250, 1142, 1425]. Roberts [1160]. Robertson [2331]. Robin [1233]. Robins [1151, 1748, 1219, 2755]. Robinson [143, 1203, 1411, 1122, 776, 437, 2597, 2684, 1122, 902, 2870]. Robinsons [1156]. Robison [664]. Robson [1780]. Rockall [136]. rocket [1330, 1352]. rockets [1981]. Rockingham [299]. Rocks [2389, 2020]. Rocky [1549]. Roderick [1142, 2845]. Rodger [1271]. Rodolph [188, 108]. rods [2212]. Roentgen [1434]. Roger [1470, 1249, 2160]. Rogers [1866, 1358, 2710]. rogue [2209]. Roland [1649]. Role [786, 2448, 1539, 2036, 766, 1931, 2484, 2032, 1737, 2031, 1825, 2592, 1041]. roles [1271]. Roman [1659, 1792]. Romances [1710]. Romanes [2060]. Romantic [2447]. Romantic-era [2447]. romantics [2008]. Rome [1738]. Ronald [204]. Rondelet [2182]. Röntgen [82, 110, 728]. Rooke [380]. room [1416, 867]. rooms [467, 92]. roots [2472, 1858]. Rosalind [1716, 1702, 1951, 2782]. Roslynn [1319]. Ross [891]. Rossignol [2375].

Rotates [2854, 1768]. rotation [1561]. Rothschild [2326]. Round [1633]. Routledge [1614, 1674]. Rowan [2070, 1827]. Rowlands [1121, 1360]. Roy [963, 975, 941, 805, 1875, 1592, 1637]. Royal [1209, 1293, 568, 666, 739, 80, 425, 786, 2071, 2070, 1721, 259, 2091, 832, 1560, 1572, 1622, 566, 1455, 1558, 1689, 491, 173, 229, 512, 472, 1347, 1796, 1603, 969, 149, 1760, 1082, 1180, 427, 280, 1613, 999, 1160, 1070, 606, 1582, 1270, 1506, 1171, 1716, 1098, 1667, 1157, 1334, 822, 1463, 531, 1133, 1791, 1623, 122, 1056, 1750, 159, 197, 272, 471, 1538, 2128, 1680, 258, 1431, 1645, 431, 844, 2093, 2270, 2317, 524, 568, 666, 1134, 1448, 876, 2623, 1272, 1909, 478, 46, 50, 100, 90, 170, 217, 262, 275, 291, 305]. Royal [303, 331, 340, 358, 367, 441, 421, 422, 424, 452, 474, 487, 498, 508, 517, 528, 540, 551, 570, 588, 601, 619, 634, 646, 657, 671, 683, 704, 718, 733, 734, 747, 761, 772, 785, 799, 807, 818, 829, 842, 852, 861, 2071, 926, 1304, 1344, 1579, 183, 269, 296, 1509, 2017, 1128, 1875, 2360, 295, 335, 2039, 2069, 1917, 484, 1170, 1240, 738, 1560, 2326, 768, 643, 1367, 2655, 1285, 85, 886, 313, 28, 1165, 1245, 2596, 2303, 567, 2030, 1406, 1564, 1778, 1738, 212, 816, 845, 2037, 504, 1722, 928, 1793]. Royal [78, 402, 30, 178, 2121, 583, 397, 2272, 86, 1369, 2040, 1499, 1562, 731, 2701, 1559, 1220, 2345, 2615, 289, 905, 1465, 783, 1024, 857, 403, 594, 766, 920, 945, 1082, 2720, 1303, 2726, 426, 456, 2072, 58, 2034, 2206, 1882, 37, 92, 98, 267, 430, 593, 398, 405, 764, 2459, 556, 2595, 1661, 1683, 2152, 481, 514, 2029, 2035, 2172, 802, 907, 946, 2747, 2749, 2109, 2357, 2050, 972, 1742, 2258, 1282, 2907, 1586, 2593, 2550, 195, 1129, 1881, 1570, 2409, 2193, 2123, 1345, 789, 1246, 21, 1828, 2038, 400, 554, 616, 1112, 1154, 1150, 1195, 790, 850, 371]. Royal [2478, 855, 724, 1398, 854, 2793, 1887, 1868, 1949, 1959, 2213, 2033, 890, 937, 966, 1052, 1393, 423, 753, 1216, 1089, 2032, 1259, 45, 1819, 1860, 592, 1974, 834, 2329, 956, 990, 2455, 36, 104, 2505, 1847, 836, 7, 538, 744, 70, 947, 2579, 2031, 1305, 610, 1197, 126, 184, 176, 117, 1443, 2112, 320, 2161, 55, 804, 1980, 2257, 2618, 444, 2434, 411, 282, 2271, 435, 2924, 1898, 74, 1843, 220, 2527]. Royale [150]. Rt [661, 400, 349, 722]. Rudolf [1474]. Ruestow [1384]. Ruffini [782]. Ruhemann [1202]. rule [2040]. rules [1835]. Rum [1529, 2833]. Rumford [1544, 186, 2647, 284, 1443]. runes [364]. running [2036]. Rupert [1181, 1231, 1316, 1543, 1382, 994, 1008, 1316]. Rupke [1297]. rush [2468]. Russel [1807, 2286, 2561, 346, 351]. Russell [1071, 1207, 1408, 1526, 783, 1084, 2019, 1063]. Russia [2238, 1579, 1721, 1149, 320]. Russian [950, 1062, 1218, 1466, 2409]. Rutgers [1138]. Rutherford [989, 599, 141, 688, 692, 2652, 1113, 691, 1934, 2159, 693, 690, 694, 695, 689, 1031, 1615, 93, 2273].

S [1626, 1361, 782, 1190, 821, 1213, 781, 2912, 1335, 1189, 1087, 1265, 902, 1935, 2775]. S. [1361, 2912, 1073, 378, 2862, 2598]. S.J [326]. S.R.L [1475]. S.R.S. [948]. Saale [1948]. Sabbagh [1529]. Sabine [2432]. Sackler [1463]. Sackville [2539]. Sadi [626]. safer [2219]. Sage [1541, 1210, 1682]. Said [1162]. Saint [1319]. Salt [227]. Saltcoats [591]. Salvador [1880, 1880]. sample [2488]. Samuda [2257]. Samuel [986, 836, 478, 929, 725, 1934, 2322, 911, 2315, 767, 2316, 1680, 258]. Sanderson [1007]. Sandwich [1271, 2827]. Sangam [1340, 2485]. sanitary [743]. Sarah [2293]. Sarton [2210]. satire [583]. satirical [117]. Saturday [411]. Saturni [597]. Saunders [2793]. Saussure [178]. Savant [2590]. savants [1243]. Savilian [433]. Savoy [1416]. Saxony [1800]. Scale [2752, 1704, 1586]. scattering [1985, 2001]. scenarios [2608]. Schaaf [1184, 1409]. Schaffer [1116, 1942]. schemes [2388]. Scheuchzer [140]. Schiebinger [1138]. Schneewind [1119]. scholar [2256]. scholars [2599, 434]. scholarship [2310]. Schonland [2626, 1672]. School [1866, 1863, 2319, 1931, 2913, 2541, 1712, 1153, 779, 584]. schooldays [2136]. Schoolmaster [1625]. Schoolmaster-Fellows [1625]. Schools [485]. Schrödinger [1088, 2596, 2794]. Schunck [815]. Schuster [730]. Schweinfurt [1849]. Science [2620, 2622, 2623, 2095, 114, 319, 1338, 67, 2628, 2636, 2638, 2639, 2039, 1235, 1185, 2248, 839, 1212, 2651, 2659, 1622, 1798, 1388, 1763, 1379, 1392, 1542, 1666, 1361, 1113, 2693, 1637, 1766, 2698, 2506, 2703, 2168, 710, 1229, 1814, 2707, 2712, 2713, 2714, 826, 2715, 2722, 1386, 2730, 2731, 1387, 2348, 2738, 2740, 1714, 2747, 2750, 1316, 1179, 1427, 2754, 2760, 1270, 1506, 2762, 2763, 2764, 2912, 2768, 1356, 1855, 1800, 2772, 2914, 1687, 2552, 1345, 2781, 2379, 1138, 1210, 1294, 1803, 2790, 2795, 2796, 2797, 2798, 2801, 1830, 2807, 2809, 1290, 2919, 2815, 2816, 2818, 1870, 1056]. Science [1436, 1816, 2834, 2837, 2341, 2841, 2847, 1362, 1544, 2853, 2854, 2860, 2868, 435, 2871, 1144, 1295, 2874, 2877, 2590, 2192, 2385, 1272, 1727, 1988, 1062, 2017, 901, 1065, 2440, 1513, 1626, 1300, 1907, 2028, 2403, 2233, 2884, 963, 2362, 1239, 792, 2249, 2244, 932, 2036, 1602, 229, 512, 1376, 1566, 1833, 1520, 2291, 1848, 2167, 2448, 1932, 743, 2528, 2557, 1867, 2432, 2588, 2018, 2274, 403, 1082, 1206, 1337, 1947, 1096, 1707, 1810, 1047, 256, 2907, 1948, 2290, 1286, 1152, 1548, 1635, 2115, 1761, 1040, 598, 632, 667, 2554, 1786, 1126, 2442, 679, 1102, 1132, 1487, 2410, 1232]. science [890, 2800, 2591, 966, 1541, 1629, 1783, 2444, 2491, 2608, 1393, 2361, 2101, 1625, 1768, 625, 2145, 2569, 848, 1473, 1632, 1750, 1186, 1557, 2338, 2922, 2850, 1162, 2447, 1291, 1283, 1873, 2137, 2923, 2434, 1442, 2424, 2134, 2095, 1300, 1212, 1960, 1831, 1066, 1230, 1613, 1333, 1487, 1232, 1435, 1186, 722, 1276, 2330, 1550, 1525, 1714, 1138, 2151, 1507, 1732, 1522, 1547, 1056]. Sciences [164, 1248, 532, 15, 1096, 1948, 1244, 2610, 2429, 2027, 3, 186, 1793, 821, 2577, 538]. Scientiarum [13]. Scientific [1483, 106, 100, 112, 164, 439, 1288, 1235, 2647, 1212, 2658, 1342, 1765, 2678, 1653, 2703, 2613, 2895, 869, 2723, 2896, 2897, 2899, 1688, 92, 2736, 2901, 2740, 2513, 2749, 2905, 1586, 2756, 1207, 1627, 2591, 2811, 2817, 1453, 1631, 2845, 101, 1474, 2549, 2858, 1382, 2875, 2192, 1469, 544, 144, 180, 669, 1447, 299, 1464, 1983, 285, 1203, 2499, 2502, 2124, 2352, 2597, 2147, 2326, 1941, 1379, 1785, 1376, 2509, 2225, 129, 557, 2598, 856, 2306, 119, 2498, 2366, 1094, 2587, 1762, 2346, 2459, 2534, 1854, 2172, 802, 907, 980, 1118, 325, 1484, 2409, 1800, 2387, 237, 1138]. scientific [922, 2126, 1142, 2170, 2307, 1504, 1003, 2293, 2349, 2920, 2821, 1123, 1842, 2611, 2299, 2380, 1973, 1596, 1740, 2519, 2473, 1544, 1943, 2342, 2359, 468, 1614, 2449, 134, 172, 485, 850]. Scientifiques [1599]. Scientist [2120, 2626, 2652, 2675, 2677, 2679, 2681, 1361, 1654, 2732, 1356, 1142, 1672, 1539, 1319, 2845, 1385, 2855, 287, 930, 1767, 1784, 520, 1975, 1615, 1713]. Scientists [1616, 2887, 1762, 2789, 2792, 2816, 1596, 73, 54, 2017, 1508, 969, 1070, 2409, 2410, 563, 1608, 1388, 1542]. Scienza [1523]. scienzato [1380]. Scienze [1379]. Scotia [2411]. Scotland [2225, 900, 2435, 591, 820]. scotoscope [1025]. Scott [317, 780, 931, 1593]. Scottish [162, 950, 55]. Scratching [1785]. screens [2433]. Scribbles [881]. script [2388]. Scriptum [1589]. scurvy [609]. Se [503]. Sea [2423, 1608]. seafaring [2440]. Seal [136]. Seaman [2743, 1099]. search [525, 1943, 891]. Seas [2673, 1231, 1562, 2450, 1429]. Sebastian [492]. Secoli [2631, 1456]. Second [2526, 299, 2606, 2034, 393]. Secord [1433]. secrecy [1775]. Secret [2784, 1485, 1470, 2589]. Secretary [491, 280, 1150, 1863, 722, 87, 237]. Secreto [2784, 1475]. Seeds [1379, 1205, 2838]. Seeing [1190]. seem [2475]. seen [1061, 2146]. Segovia [2538]. Seicento [1380, 1379, 2709]. seismic [2454]. Selborne [1767, 2681]. Select [2638, 1291]. Selected [2880, 2920, 1073, 1189, 1474]. Selection [2882, 2661, 2890, 547, 1807, 1653, 1359, 1546]. selections [2852, 1144]. Selenelion [1914]. Self [1253, 883, 2269, 2295]. Self-confident [1253]. self-presentation [2295]. self-registering [883]. self-taught [2269]. selfish [1893]. Selling [2090, 2403]. semen [2285]. Seminar [1910]. send [2488]. Sengers [1712]. sensing [2526]. September [795, 363, 1849, 591, 2923, 1201]. Sequeira [2257]. Sergay [1508]. Series [1786]. sermon [2070]. Sertić [2482]. Servant [1487]. Service [579, 615, 79, 16, 2707, 2745, 417, 580, 591, 776, 219, 1185, 142, 577, 707, 831, 708, 1487, 820, 1715]. Services [147, 1180]. session [2042, 2041]. Sessions [1273]. Seth [160]. settlement [1415]. Seven [2621, 1731]. Seven-year [2621, 1731]. Seventeenth [2270, 2317, 786, 1814, 2714, 2905, 2783, 1333, 2507, 1520, 766, 1343, 1427, 2213, 1860, 2324, 2181, 1453, 1967, 804, 2453, 886, 2709, 2750, 2388]. Seventeenth-Century [2507, 2270, 2317, 2714, 2905, 1333, 1520, 1343, 2213, 1860, 1453, 2453, 2750, 2388]. Sex [2837, 1138]. sextant [845]. Sexton [1205]. Sexuality [1800]. Shadows [1184, 1551, 2835, 2572]. Shale [1551, 2672]. shall [2463]. Shape [2750, 1333]. Shaped [2765, 2766]. shapes [1368]. Shaping [2726, 1384, 1667, 2829]. Shapley [2500]. shared [2408]. Sharpey [1496]. shear [2439]. Shedding [1702]. Sheffield [995]. shell [2593, 416]. shell-shocked [2593]. Shelley [1152]. Sherrington [317, 585, 782, 781, 458, 130, 244, 318, 131, 780, 961, 1935]. Shifting [1669]. Shining [1493]. Shipborne [1776]. Shirley [1160]. shock [2075]. shocked [2593]. Short [612, 562, 611, 466, 1660]. Short-writing [466]. shorter [2800]. should [2354]. Shout [2546]. Showcase [1870]. showmen [2008]. Shrewsbury [1188, 779, 584]. Sibbald [2411]. sickness [729]. siècle [2368, 2371, 1887]. Siegfried [1769, 1202]. sigils [1956]. Sigma [1475]. sign [952]. significance [1810, 2365]. silent [2376]. silkworm [1262]. Silver [2669, 1710]. Simms [1273, 1273, 2787, 2791]. Simon [1116, 1138, 1276, 1551]. simple [2502]. Since [1881, 2779, 1502]. Sinclair [1159, 2405, 2427]. Sinclair-Stevenson [1159]. Singapore [1637]. singular [1957, 1306]. Singularities [1410]. Sino [2591]. Sino-British [2591]. Sir [1267, 1071, 874, 1251, 1673, 1653, 1593, 1704, 1142, 1205, 2819, 1708, 1474, 1144, 1646, 1662, 317, 615, 1068, 1293, 39, 130, 236, 244, 1965, 2261, 2304, 2339, 2363, 2392, 1466, 271, 1177, 1226, 1311, 2878, 2630, 2632, 2633, 1203, 2091, 561, 988, 1317, 986, 1400, 777, 1917, 752, 2233, 768, 2656, 2657, 85, 2885, 1444, 2661, 1798, 2390, 2674, 185, 255, 1347, 641, 86, 2616, 388, 447, 2711, 959, 1230, 577, 968, 627, 398, 707, 659, 256, 1160, 2752, 325, 708, 2214, 1129, 2909, 1417, 1351, 1405, 1452, 1500, 1573, 493, 548, 701, 778, 281, 273]. Sir [392, 616, 854, 2793, 248, 2149, 2391, 796, 1921, 1228, 327, 1772, 2826, 1958, 2831, 268, 1708, 2838, 1647, 448, 1110, 131, 1038, 1740, 887, 1130, 580, 2845, 257, 950, 2852, 1443, 436, 776, 820, 1109, 1111, 1375, 2871, 2925, 199, 1163]. site [2337, 2492]. site-specific [2337]. Six [883, 883]. sixpence [1816]. Skara [741]. sketches [2541, 1887]. Skies [2853, 1525, 2471, 2257]. skin [2458]. skulls [2389]. Sky [2765, 2766, 2809, 2473, 1522, 1549]. Slare [1127]. slavery [1465]. Slawka [1434]. Sleeping [729]. Sloan [1248]. Sloane [255, 887, 641, 1880, 256, 257]. small [1026, 2323]. Smeaton [642, 674, 888]. Smeatonian [568]. Smith [1069, 1117, 743, 912, 1706, 1101, 1527, 1473, 1295]. Smithies [1457]. Smithson [1915, 449]. Smithsonian [2573]. smoke [1314, 2137]. Sneinton [1252]. snuff [83]. snuff-box [83]. Soane [1588]. soaring [2439]. Soc [963, 975, 941, 805]. Social [2817, 2319, 2017, 802, 632, 1138, 1740]. Socialism [1800]. Societas [13]. Société [150]. Societies [106, 144, 164, 180, 129, 2309, 119, 105, 128, 88, 89]. Society [524, 568, 666, 100, 170, 303, 786, 85, 2121, 86, 1831, 149, 1024, 1180, 2722, 2734, 98, 1316, 1613, 1160, 1070, 2908, 1270, 1506, 1171, 2409, 1716, 1098, 1855, 616, 1154, 1157, 371, 2793, 1056, 1408, 1769, 546, 197, 804, 2865, 1431, 1185, 1918, 567, 2607, 1833, 1049, 2907, 2353, 1950, 1750, 844, 2093, 2270, 2317, 739, 1134, 1448, 876, 2623, 1909, 478, 46, 50, 27, 41, 66, 53, 65, 75, 80, 96, 109, 90, 112, 115, 120, 143, 132, 145, 168, 163, 181, 192, 203, 206, 221, 240, 251, 262, 263]. Society [275, 276, 291, 290, 305, 304, 311, 321, 331, 340, 330, 358, 367, 357, 366, 441, 421, 425, 422, 424, 440, 452, 451, 474, 487, 473, 486, 498, 508, 497, 507, 517, 528, 516, 527, 540, 551, 539, 550, 570, 569, 588, 601, 587, 600, 619, 618, 634, 646, 657, 633, 645, 656, 671, 683, 670, 682, 704, 703, 718, 733, 734, 717, 732, 747, 761, 746, 760, 772, 785, 771, 784, 799, 807, 798, 806, 818, 829, 817, 828, 842, 852, 841, 851, 861, 860, 870]. Society [877, 2071, 2070, 926, 1304, 1344, 1579, 1721, 183, 269, 259, 296, 1509, 2017, 1128, 1875, 2091, 2360, 295, 329, 335, 2039, 2069, 832, 1917, 484, 1170, 1240, 738, 1560, 1572, 2326, 768, 1367, 1285, 886, 1622, 566, 1918, 313, 28, 1165, 1245, 2596, 567, 2030, 1558, 1778, 1738, 212, 816, 845, 2037, 504, 928, 1793, 8, 78, 491, 30, 103, 173, 178, 229, 512, 472, 1347, 583, 574, 2272, 179, 1369, 2040, 1796, 1603, 1499, 1562, 1559, 1220, 2345, 2615, 1760, 289, 905, 1465, 783, 1121, 2092, 857, 594, 766, 920, 945, 1082, 2720, 1303, 2726, 617, 427, 426, 455]. Society [456, 2072, 58, 2034, 2206, 1882, 37, 92, 267, 280, 430, 593, 339, 398, 764, 2459, 556, 2595, 1661, 1683, 2152, 481, 514, 2029, 2035, 2172, 802, 907, 946, 2747, 2749, 2109, 2357, 2050, 972, 1742, 1282, 1586, 2550, 195, 606, 1582, 1881, 1570, 2193, 2123, 485, 1345, 1246, 21, 20, 43, 34, 35, 69, 153, 1828, 2038, 400, 1667, 554, 716, 1112, 1150, 1195, 850, 2478, 822, 855, 724, 1398, 854, 1887, 1868, 1949, 1959, 2213, 2033, 890, 1463, 937, 966, 1052, 1133, 1393, 423, 753, 1216, 1089, 2032, 1259, 1791, 1819, 1860, 592, 1974, 1623, 2329, 956, 990]. Society [2455, 36, 104, 160, 2505, 2828, 1847, 836, 7, 5, 166, 538, 159, 702, 744, 11, 70, 947, 2579, 565, 2031, 1305, 610, 272, 1197, 126, 184, 176, 117, 471, 1538, 320, 1680, 2161, 55, 258, 2257, 2618, 444, 2434, 282, 2271, 435, 1898, 74, 1843, 1645, 191, 205, 220, 239, 2527, 969, 531]. Society-London [593]. Sociey [45]. Soddy [1626]. Soft [1273]. softback [1433, 1160, 1098, 1527, 1489, 1473, 1596, 1430]. softened [2382]. Soho [2450]. Solar [1883]. Soldier [2626, 1355, 2848, 1672]. soldiers [2593]. Solla [2256]. solution [1306]. solvent [2093]. Solving [1504]. Some [599, 350, 271, 1004, 1585, 175, 318, 1032, 1024, 479, 756, 693, 757, 855, 881, 1003, 689, 796, 1289, 1647, 663, 184, 176, 962, 199, 345, 299, 833, 2124, 566, 1776, 492, 1007, 1175, 1644, 209, 270, 1533, 230, 237, 927, 1217, 1737, 1080, 436, 973, 460]. someone [2331]. Son [2630]. sonar [2159]. sonobuoys [2454]. Sons [1593, 825]. Soo [1505]. Sorrenson [1506]. Soshua [1129]. sound [2526, 1825]. sound-ranging [1825]. source [1513, 2382, 1604, 2561]. Sources [2737, 2656, 2254, 2458, 1674, 1068]. South [2310, 1562, 2423, 2450]. Southeast [2197]. Southern [1547, 2503, 2914]. Soviet [361, 1508, 2887, 1508, 2483, 2577]. Sowerby [2238]. Space [2729, 1332, 2032]. Spain [2547, 2367]. Spanish [1024, 2784]. spatiality [2100]. speake [2413]. speaking [2410]. Spearman [1309]. Special [419, 418, 420, 2396]. Species [2687, 2879, 1117, 1595, 2502, 2546, 1890, 1913, 460, 2425]. specific [2337]. Specimen [2910, 1591, 1060]. Specimens [2510, 897, 1197]. spectro [960]. spectro-chemical [960]. spectrometry [1851]. spectroscopist [1585]. Spectroscopy [1732, 1378, 2063, 2162]. spectrum [896]. specula [927]. speculum [2058]. Speech [349, 661, 777, 752, 308, 360, 722, 821, 637]. speed [679]. Spencer [1982]. sphere [2361]. Spheres [2891]. Spider [1182]. Spider-man [1182]. Spiders [1182, 2814]. Spilman [1218]. Spinning [2768]. spirit [2383]. spiritualism [1891, 1890]. Spokesman [1300]. Spontaneous [2728, 1685]. spots [1183]. Spreadsheets [2883, 1780]. Spring [2049, 1157, 2376, 273]. Spring-tide [1157]. Springer [1103]. Springer-Verlag [1103]. springtime [2192]. Square [2508, 1739, 2450]. squares [1956]. Squaring [1534]. St [1068, 1097, 707, 2299, 2070, 1721, 591, 820]. St. [417]. Staatsbibliothek [2618]. Stability [2662, 1189]. staff [104]. Stafford [1142]. stage [2367]. Stahl [1281]. Stalbridge [480]. Stalin [2577]. Standard [1411]. Standardization [1884]. Stanhope [285]. Stanislao [535]. Stanley [1907, 1925, 1902]. Starling [1698]. starry [1660]. stars [1455, 846, 2924]. State [2707, 1206, 2785, 1487, 12, 722, 466]. States [1028, 2578]. Statesman [1300, 1361, 598, 2923]. statically [681]. statically-indeterminate [681]. Station [1997, 759]. stations [187]. Statistical [2664, 1189, 1800, 2821]. Statistics [1800]. status [1794]. Stazione [201]. Steacie [1102, 2762]. steam [663]. Stedall [1746]. Steeds [1163]. steel [58, 981]. steering [1624]. Stellar [2665, 1073, 921]. stem [2135]. Stennis [741]. Step [1275, 2856, 1918, 2551]. Stephen [1777, 968, 833, 57]. Stevenson [1159]. Stewart [107, 1232]. stipes [2570]. Stochastic [2664, 1189]. stock [2465]. Stoicheff [1732]. Stokes [938]. Stone [2101, 2270, 1346, 1370]. stones [1331]. Stoney [770]. Storia [1523]. Stories [1542, 2730, 2436, 1635]. storm [1094]. Story [2646, 2669, 1190, 2770, 2775, 2791, 1273, 2113, 2533, 1872, 624, 1164, 1710, 1495, 1802, 2425, 1967]. Stoughton [1207]. Stoye [1355]. Strange [2670, 214, 1777, 1331, 2751, 1231, 2574]. Strangelove [1319, 2732]. strategy [2576]. stratification [2102]. Street [2534]. streets [1394]. Stronsay [2594]. Stroud [1275, 1671, 1544]. struck [2258]. structural [2575]. Structure [2665, 2100, 1081, 847, 2454, 1073]. structures [681]. struggle [1343]. Strumenti [2631, 1456]. Stuart [1209, 1086]. student [1259]. students [1931, 2600]. Studi [1379]. Studies [1248, 1883, 2717, 2908, 209, 227, 246, 270, 283, 480, 513, 1786, 1353, 1947, 2903, 1143, 1442, 1248, 1116]. Study [2510, 2727, 2771, 2840, 2489, 1097, 839, 2404, 2556, 1346, 1101, 790, 2170, 2444, 409, 1943, 2523]. stuffed [1243]. stuffer [849]. Stukeley [1707, 2731, 191]. Stulz [450]. Stumbling [1489]. stumps [2570]. Stutchbury [929, 911]. style [2459, 1126]. Styles [1288, 1324]. sùance [2574]. sub [2209]. sub-Antarctic [2209]. subjects [756, 2006]. submarine [2159]. subscription [1993]. substances [2161]. substantiated [1181]. successful [2223]. Such [2669, 1710]. Suffolk [1082, 1613, 1484, 1116]. suffrage [2291]. suggestion [2370]. sulphur [1919]. Sumer [1780, 2883]. Summary [2631, 1456]. summer [1422, 1198]. sundial [1908]. sundial-clock [1908]. sundials [1908]. sunglasses [1917]. suns [1038]. Superior [189]. support [783, 1079]. Supposed [2568, 2357]. supposition [805]. suppuration [2222]. supremacy [1398]. Supreme [2652, 1615]. surface [2063, 2001, 1985, 927, 2162]. surface-enhanced [2063, 2001, 1985, 2162]. Surgeons [405, 411]. Surgery [2826, 2219, 1160, 2224, 2221, 1160]. Survey [2436, 2337, 2429, 2559, 2430]. surveying [2430]. Surveyor [1230, 139, 1366, 1394, 1415]. surveys [1366]. survival [1629]. Sussex [1332, 1315, 1424]. Sustainability [2498, 1787]. Sutton [1275, 1671, 1544]. Swade [1594]. Swammerdam [2185]. Swann [1541]. Sweden [2090]. Swedish [1291]. sweeper [1026]. swine [1110]. Swinton [827, 923, 1479]. sword [2238]. Sydney [1069, 1117, 1637, 1188, 2610, 2705, 899]. Sylva [531]. Sylvester [1583, 1373, 1176]. Symbolic [2687, 1595]. Symbolism [2696, 1428]. symmetry [2508]. Sympathetic [2607, 2696, 1428]. Symposium [795, 2908, 1450, 1228]. Symptomatology [778]. syndrome [1916]. synthesis [1203, 2617]. synthetic [2325]. system [1128, 795, 667, 2611]. systemic [1348]. systems [1561]. Szörényi [2577].

T [1528, 472, 891]. T. [2122]. tabella [2322]. table [19, 2595, 423, 1780]. Tables [2883, 1780]. Tail [1882]. taken [1394]. taking [2520]. Talbot [1184, 1409, 2835, 2836]. tale [824, 2332, 2284]. Tan [1383]. Tansey [1255]. Tartar [1344]. Tashjian [1141]. tastes [2595]. taught [2263, 2245, 2269]. Tauris [1430]. Taxis [2308]. Taylor [1410, 725, 1208, 1489, 2517, 2518, 2522, 2520, 2521, 2519, 2524, 2523]. teacher [2263, 2245]. teaching [1196, 1263, 922]. teacup [1094]. Technical [1932, 2549]. technician [1933]. Technicians [1929, 2292, 1930]. techniques [2443, 2543]. Technological [2031, 2374, 557]. Technology [1212, 1361, 2722, 2730, 1688, 1316, 2912, 1232, 2847, 2884, 2030, 1566, 1848, 826, 1947, 1635, 2768, 1761, 1786, 2488, 679, 2798, 2799, 2801, 2324, 1442]. technoscience [2373]. Ted [1767]. teeth [2542]. Tehran [2488]. telegrams [2406]. telegraphic [2406]. telescope [1168, 1944, 1260, 1303, 854, 1261, 947, 1327]. telescopes [927, 651, 642, 611]. telescopic [2058, 2453]. television [827, 923]. Telling [1542]. Temperature [1870, 2236]. temperatures [613]. Tenerife [2586]. Tennant [449, 1915]. Tennyson [1132]. Tensor [2662, 1189]. tentative [338, 560]. Tercentenary [421, 424, 399, 271, 402, 397, 403, 398, 405, 400, 409, 2851, 411, 422, 313, 1659, 1141, 455, 456, 137, 789, 401]. Tereshkova [496]. term [1941]. terms [434]. Ternate [2259]. Terrace [1204, 574, 575, 2066, 555]. terraces [2248]. Terrence [1595]. terrible [1132]. tertiary [2103]. Test [1784, 2059]. testimony [2594]. Testing [2079]. text [2061, 2800, 1943]. textbooks [1955]. Texte [150]. Textile [2768]. Texts [2205, 1425, 1860]. Thackray [1248]. Thallium [960]. thankless [2582]. Thanksgiving [580, 1715]. Thant [1833]. Thatcher [2120, 2274]. Theatrical [1766]. Their [2640, 1379, 2715, 524, 568, 2248, 2405, 2254, 2533, 1605, 1129, 2575, 1885, 2007]. Them [2730, 1270, 1635]. themselves [2111]. Thenard [970]. Theobald [1325]. Theodolites [2559]. Theodore [1800, 2600]. theological [2418]. theology [1516]. Theophilus [2476]. Theophobia [664]. Theorem [2662, 949, 1536, 1172, 1189]. Theoretical [1474]. theories [938, 2602]. theorizing [2563]. Theory [1410, 2667, 2680, 402, 2692, 1189, 694, 1053, 1339, 1117, 484, 1807, 797, 681, 1039, 2092, 2282, 2206, 1313, 72, 939, 1733, 1700, 2612, 2316, 678, 1801, 2843, 1172, 2864, 432, 1029, 1906, 1457]. theosophical [1250]. therapy [2482, 2370, 2075]. there [2580]. Thermometer [2433, 883, 613]. thesis [2059]. thicknesses [1605]. Things [2621, 2728, 1003, 1731, 2323, 1685]. think [2489, 1516]. Thinking [2151, 1586, 2332]. Third [2716, 2473, 285, 1287, 1234, 1507]. thistles [2468]. Thomas [1746, 824, 2005, 1865, 1144, 91, 643, 564, 2605, 175, 1347, 469, 2272, 2004, 353, 470, 945, 1180, 756, 2746, 2574, 226, 225, 793, 1028, 2754, 1201, 631, 906, 1695, 1186, 744, 2846, 2539, 377, 2011]. Thompson [1443]. Thomson [1489, 1459, 680, 2682, 2686, 910, 938, 980, 631, 1695, 2047, 316, 680].

Thornes [1525]. thorns [2468]. those [2566]. Thought [2721, 2724, 2794, 1353, 2606, 1231, 1126, 2094, 1231, 1382, 1088]. thought-style [1126]. Thoughts [2694, 1598, 2546, 1647, 1383]. threat [1978, 1608]. Threats [1859]. Three [1248, 2717, 2824, 677, 1589, 824, 2618, 1415]. thrift [2543]. Thurn [2308]. Thursday [666]. Tibet [2492]. Tidal [1504]. Tide [2873, 1157, 1597]. Tides [2658, 1694, 45, 1504]. Tierra [2444]. tiles [2488]. Tilli [1255]. Time [2635, 2542, 1558, 2729, 2745, 1332, 1784, 1205, 2838, 1639, 2102, 945, 1040, 2578, 2101, 2505, 803, 2876, 1670]. Times [1185, 2695, 2716, 2745, 1714, 2779, 1355, 2813, 1187, 2848, 1234, 228, 2893, 1100, 1120, 1885, 1502]. Timothy [1508]. tirade [1261]. title [2425, 1893]. Toast [401, 400]. Tobin [1768]. Today [690, 2173]. Together [2623, 2463, 1056]. told [2533]. Tomes [2091]. Tomlinson [1761]. tomorrow [1859]. Tonkin [1694]. Tony [1589]. Too [1777, 2751, 2413]. tool [1943]. tools [2101, 2103]. Top [1209]. Top-drawer [1209]. Topham [1458]. Toronto [1102, 1186]. toscana [1379]. Tour [513, 1028, 1496, 1815]. tower [2544]. Townes [1713]. Townsend [1626]. tracts [1259, 1505, 2844]. trade [2308, 2408]. trades [966]. tradition [812, 2011]. tragedy [2473]. Tragic [2674, 1593]. Trailblazing [2050]. train [2094]. training [1566]. Traité [1054]. trajectories [2007]. Trans [805]. Transactions [969, 531, 1769, 1822, 511, 1115, 2335, 2309, 72, 2478, 1967]. transatlantic [1739]. Transfer [2663, 1073]. Transferring [2549]. transformation [1065]. transformations [2344]. Transformed [2621, 1731, 1774]. Transforming [2795]. transformism [2617, 2301]. Transit [2791, 607, 1273]. Transition [2628, 1507]. transitions [2531]. Translated [1182, 1508]. Translating [1581]. Translation [1065, 2049, 970, 2530, 2206]. translations [1967]. transmission [970, 492, 795]. transmitter [931]. transmutation [2879, 2822]. transnational [2369]. transplantation [408, 2391]. Travaux [1599]. travel [1479]. traveller [216]. Travellers [2447]. Travels [2547, 1817, 837, 1478, 1326]. Treas [707, 591]. Treasure [1750, 1453, 2270]. treasures [1272]. Treasury [400]. Treatise [2846, 1746, 1919]. Treatises [547]. treatment [2541]. Tree [2890, 1546]. trees [2072, 1919]. Trend [1317]. Trends [403]. Trevor [1180, 1122]. Trial [2174]. triboelectric [676]. Tribute [2885, 1163]. tricentenaire [137]. Trifolium [2398]. trigonometrical [2559]. Trilobite [2704, 1638]. Trimen [1649]. Trinity [654, 577, 2745, 467]. trio [2033]. trip [1915]. triumphant [2260]. triumphs [1858]. Tropical [1866, 1863]. tropinone [1203]. Troughton [1273, 2787, 712]. trousers [2292]. Troves [1453]. True [1129, 2770, 2169, 1802]. Truly [1614]. Trust [2, 1056, 1079, 59]. Trusted [2660, 1488]. Trustees [1212]. truth [2366]. Tsukerman [1508]. Tsunami [1941]. Tube [2076]. Tucci [1456]. Tuesday [401, 398, 400]. Turbine [549, 1095]. Turing [2173]. Turkey [2598, 2549]. Turn [2547, 2570, 1608]. Turning [1333]. Tuscan [2300]. Tweeddale [1037]. Tweedie [2157]. twentieth [1995, 2281, 2615, 2503, 2034, 2535, 2029, 2767, 2135, 2033, 1617, 2324, 1632, 2570, 1933]. twentieth-century [2535, 2033, 1632]. Twenty [2908, 1084, 2350, 1797, 1228, 2341, 2027]. twenty-eighth [1228]. twenty-first [1797, 2341, 2027]. Twenty-fourth [2908]. Two [883, 2631, 1456, 1626, 2597, 1693, 2691, 585, 728, 33, 1474, 1048, 2303, 1406, 1766, 997, 2381, 2194, 980, 1483, 1394, 1352, 1423, 1481]. Twort [1275]. Tycho [1549]. Tyndall [2318, 2258, 2457, 1467, 1481]. type [2577]. typers [2481]. typeset [1314]. typing [2481]. Tyson [2170]. Tywardreath [1251].

U [821, 1833, 1608]. U-boat [1608]. U.S.A. [1268]. U.S.S.R [538]. Ubiquitous [1296]. Uglow [1730]. UK [2192, 2376, 2036, 2032, 2619, 2081]. Ulf [1765]. ultimate [1701]. Unauthorised [2901]. Uncertain [1229]. Uncertainty [2659, 1229]. unchampioned [989]. Uncovering [1393]. Understanding [2039, 2529, 2248, 1941, 2484, 2519]. Undescrib’d [2517]. Uneasy [2620, 2122, 1138]. unequalled [1524, 2761]. UNESCO [112, 135]. unfortunate [1681, 655]. Uniformity [1514, 2433]. Union [1339]. Unions [439, 92, 101]. Unit [1762, 1424]. United [1212, 2578, 2740, 1028]. Unity [2550, 2326]. universal [2093, 1908, 2284]. Universe [1233, 2765, 2766, 407, 847]. universities [2550, 2081]. University [1272, 995, 1338, 1656, 1088, 1300, 1269, 1069, 1117, 1071, 1729, 1184, 1273, 1550, 2286, 1145, 1528, 1388, 1379, 1501, 1542, 1411, 1428, 1673, 1520, 1782, 1458, 1591, 1119, 1746, 1780, 1121, 1085, 1066, 1141, 1206, 1337, 1525, 1386, 1297, 1457, 1685, 1688, 1706, 1387, 1213, 1505, 1120, 1384, 1179, 1427, 1101, 1118, 1359, 1207, 1548, 1784, 1356, 1800, 1687, 1335, 1667, 1138, 1157, 1250, 1355, 1073, 1099, 1102, 1551, 1487, 1189, 1633, 1142, 2795, 1504, 1290, 1549, 1353, 1357, 1471, 1205, 1319, 1598, 409, 1768, 1617, 1432, 1488, 1526, 1507, 1632, 1801, 1087, 1186, 1425, 1409, 1436, 1485]. University [1631, 1670, 1713, 1543, 1362, 1358, 1383, 1276, 1651, 1711, 2923, 1522, 1382, 1339, 1083, 2400, 1153, 238, 1268, 142, 1947, 982, 1201, 2911, 1997, 457, 1424, 2299, 282]. unknown [479]. unlucky [1518]. unmaking [2046]. unmasked [2833]. unmix [2913, 1712]. unnoticed [2618]. unorthodoxy [2582]. Unpathed [2745, 1100]. unpredictability [1021]. unprincipled [1080]. Unpublished [1905, 345, 350, 833, 566, 2381, 2061, 907, 226, 652, 1773, 1868, 160, 2519, 1037]. unrecognized [585, 1420]. Unrecorded [195, 2124]. untrained [2336]. Unusual [1528, 1809]. Unveiling [1187]. update [1829]. upon [1004]. Upper [2247]. Uppsala [1291]. urban [2465]. Ure [715]. URFs [1552]. US$10.95 [1086]. US$15 [1766]. USA [1763, 1637, 1639, 1593]. USA/£ [1593]. use [2540, 1239, 1494, 2048, 131]. Useful [2706, 1337, 2400]. usefulness [1376]. Uses [2707, 2779, 1487, 1502, 1442]. USSR [2613]. utility [2396].

V [1117, 1361, 2210, 1096, 825, 2137, 2860, 411, 283]. V. [1361]. V.P.R.S. [268]. vaccination [1372]. Valence [1955]. Valentina [496]. Valley [1290]. Valltravers [188, 108]. Values [1185, 2695, 2426]. Valve [281, 1329]. Vancouver [1639]. Vane [2214]. Varieties [1602]. Variorum [1637, 1316, 1786]. Varley [1891, 1374]. Vase [1417, 1233]. Vase-maker [1233]. Vaughan [1253]. Vectia [1383]. vegetable [1421]. Veniamin [1508]. Venice [1331, 2574]. Venkata [1740]. Venkataraman [1096, 1340]. Venki [2363, 2392]. Venus [610, 607]. verified [2533]. Verifying [797]. Verlag [1103]. Vernon [1186]. versatile [1942]. versions [2124]. versus [484]. vertebrates [2122]. Verwer [2477]. Very [2788, 1145, 1219, 409, 1958, 2579]. vessels [566]. Vesso [1541]. VI [1117, 223, 480]. via [2541]. Vicarage [1251]. Victor [1119, 1975, 1119]. Victorian [1338, 1528, 1785, 1637, 1297, 1142, 1473, 1526, 2195, 1185, 2675, 2695, 2406, 2708, 1654, 2534, 1761, 632, 2781, 667, 679, 1891, 2101, 1890, 2830, 2832, 2841, 2845, 2858, 2325, 2874]. View [1572, 1523, 2678, 2765, 2766, 1123]. Viewed [2742, 1798]. Views [2691, 868, 2064, 2458, 1766]. Vignettes [1340]. Vignoles [1199, 83]. Vigo [2586]. vii [1780, 1427, 1360, 513]. viii [1209, 1235, 1069, 1671, 1122, 1474, 1651]. vilest [2468]. Vincent [1844]. Vinci [2285, 2439]. virial [2662, 1189]. virology [2482]. Virtuoso [2848, 1355]. virus [2485]. Viruses [1741, 2484]. Viscount [383, 668, 1029]. Visible [1931]. vision [2576, 2150, 2560, 1997]. Visions [2875, 2403, 1688]. Visit [54, 87, 361, 438, 669, 2613, 496, 237, 261, 1969, 1195, 320, 74]. visitors [209, 270, 855]. Visits [90, 414]. visual [2553, 2169]. vital [2602]. vitalism [2608]. vitro [1262]. Vividness [2183]. Vladimir [2482]. vocabulary [296, 2169]. voices [2505]. Vol [133, 154, 182, 207, 241, 264, 292, 322, 341, 368, 1769, 2768, 2798]. Vol. [2896, 2897, 2899]. voltage [1449]. voltaic [2532]. Voltaire [1208, 198, 261, 1241, 1904, 1208]. Volume [51, 111, 121, 169, 442, 464, 453, 489, 476, 509, 529, 552, 542, 572, 602, 635, 658, 684, 719, 748, 773, 787, 800, 830, 843, 853, 862, 878, 892, 913, 933, 953, 974, 991, 1033, 1057, 1389, 1438, 1553, 1618, 1676, 1734, 1788, 1069, 1117, 1406, 2680, 2230, 1119, 1118, 1313, 1073, 1334, 1189, 1408, 2834, 1083, 475, 1689, 1458, 1545, 1162, 1222, 1117]. Volumes [819, 1009, 2262]. Volumme [2650]. Voyage [2876, 1285, 835, 2300, 996, 610, 1670, 1068]. Voyager [2729, 1332]. Voyages [2824, 1547, 2914, 1589]. Voyaging [2650, 1313, 1231]. Vulgarly [2685, 1425].

W [143, 287, 1248, 1039, 2262, 1323, 1779, 1163, 1188, 1229, 1714, 1208, 2762, 1102, 1611, 1056, 2128, 1339, 1455]. W. [973, 232, 938, 982, 1748, 1432, 2112, 891]. Waals [1712, 2913]. Wali [1213]. walks [2876]. wall [2541]. Wallace [1807, 346, 351, 333, 2286, 2561, 1806, 2259, 2046]. Waller [948, 1291, 2638, 2049, 1421, 2123, 2207, 62]. Wallis [374, 560, 639, 38, 2414, 2415, 2144, 1324, 1343, 2417, 2416, 2419, 2418, 2413, 1534, 1535]. Walter [1088]. Walters [1208]. Walther [2628, 1507]. War [868, 1776, 2037, 2291, 2506, 1190, 2738, 2775, 1178, 1775, 2373, 1959, 1825, 2137, 2260, 2499, 229, 890, 1921, 49, 2374, 2558, 1179, 850]. Warburton [1774]. Ward [160]. Warren [1411]. wars [2588, 2430]. wartime [1808, 1758]. Warwick [1801]. was [2258, 1775, 1997, 2833, 1827, 1262]. wasn’t [2612]. Watchers [2765, 2766, 1549]. water [227, 919, 941, 2324]. Waters [973, 1300, 2745, 2209, 1100, 2079, 2316]. waterwheel [2476]. Watson [299, 1598, 1598]. Watson-Wentworth [299]. Watt [564, 979, 626, 1397, 2125, 624, 652, 1786, 625, 663]. Waugh [1103]. wave [1992, 2209, 1992, 2209]. Waves [2667, 1189]. way [2358]. ways [2581]. weak [2060]. weapons [1757]. Wear [1651]. weather [1564, 230, 2589]. Web [1666]. Webster [1727, 2357]. Wedgwood [1233, 1417, 2823, 471]. Wednesday [402, 405, 409]. weeds [2468]. Weekly [2272]. Wegg [836]. Weidenfeld [1138]. Weimar [2481]. Welcome [1045, 92]. welcomes [1745]. Weld [1949]. Well [1705]. Wellcome [1331, 1056, 2623]. Wentworth [299]. Were [2528, 229]. Werer [1229]. Werner [2659]. Wess [1272]. West [238, 1572, 701, 834]. Western [2598, 1141, 2732, 1319, 1513, 1941]. Westminster [579, 615, 831, 708, 1864, 789, 1832, 1863, 580, 776]. Wetenschappen [1712]. Whatever [2734, 1950]. wheel [1333]. Wheeler [1611]. Wheels [1594]. Wheldale [902]. Where [2744, 1026, 1290, 1983]. Whewell [1338, 505, 1338, 2134, 2874]. Which [1837, 1919, 1442]. Whig [1398]. Whirlpool [1943]. Whitaker [1705]. Whitby [1318]. White [2292, 1767, 2681, 2517, 2518, 2522, 997, 2520, 2521, 94, 2519, 2524, 2523]. Whitehead [1119, 783, 1119, 2776, 2777]. Whitehurst [2677, 1385]. Whiteside [1085]. Whittles [1527]. Who [917, 2660, 2362, 1777, 2751, 2753, 2758, 2759, 2159, 2789, 2567, 2854, 2857, 1254, 424, 1616, 2566, 969, 971, 1730, 2915, 1749, 1768, 1488, 2580]. Whytt [2607]. Wide [1675, 1650]. Wide-Ranging [1675, 1650]. Wider [2818, 1436]. Wife [2733, 1655]. Wigglesworth [411]. Wight [1308]. Wilberforce [2382]. Wild [2503]. Wiley [2385, 1361, 1593, 1434, 1749]. Wilhelm [28, 728]. Wilkins [2384, 1437, 373, 826, 2472, 1751, 624, 1126, 2442, 2069, 754, 1751, 1404, 1040, 1644, 2388, 1587, 985, 872]. Wilkinson [2616]. will [1217]. Willem [2336]. William [437, 814, 988, 459, 1400, 768, 1273, 1798, 1275, 86, 959, 1292, 1497, 2534, 1129, 1417, 1881, 376, 2793, 1710, 95, 1110, 1234, 1111, 2856, 2330, 385, 1656, 1875, 2641, 630, 1097, 1048, 792, 2669, 629, 1103, 791, 1100, 2506, 2060, 971, 2716, 2092, 2282, 250, 1264, 2731, 968, 2900, 1707, 481, 1934, 1986, 337, 2748, 960, 980, 846, 1221, 1098, 757, 363, 2791, 2138, 940, 2491, 364, 389, 234, 1768, 989, 505, 2047, 383, 1827, 1930, 2074, 680, 2075, 1981, 677, 665, 432, 2134, 2874, 191]. Williams [1122, 2070]. Willis [1180, 1733, 377, 2605, 2746, 1700]. Willmoth [1455, 1406, 1689, 1230, 1334]. Willughby [2184]. Wilson [1360, 2500, 562, 472]. Winchester [1068, 1097]. Winckworth [204]. Wind [1883, 2744, 1290, 2439, 2471]. window [1066, 202]. Windows [1995]. Wing [469]. Winners [2128, 2112, 1757]. Winter [2632, 2633, 1708]. Winthorp [68]. wisdom [2545]. Wise [1101, 1516]. wisely [2382]. Wish [2816, 1542]. Witchcraft [2357]. within [1623]. witnessing [2459]. Wittgenstein [1984, 1894, 1890]. Wollaston [232]. Wolterstorff [1353]. Woman [1539, 1138]. Women [2620, 2622, 2640, 2698, 2291, 2708, 2513, 2290, 2760, 1138, 2807, 2293, 2817, 2837, 2447, 80, 2585, 2583, 1528, 2040, 2292, 1112, 1154, 1266, 1783, 1138, 1255, 2361, 1276]. wonder [1998]. wonderful [1335]. wonders [1751]. Wood [2524]. Woodbridge [1082, 1230, 1314, 1707, 1613, 1484, 1116, 1333, 1453]. Woods [1597, 2501]. Woolsthorpe [895, 116]. Woolwich [2432]. Words [1546, 1027]. Work [2637, 2643, 2653, 1777, 2746, 2776, 2777, 1182, 285, 2489, 1269, 1455, 1366, 1394, 1415, 492, 512, 1007, 626, 1119, 2144, 1002, 1762, 2494, 1604, 1931, 226, 960, 1761, 2331, 847, 889, 2493, 1774, 1463, 1133, 979, 1695, 1758, 663, 2519, 1283, 2434, 2924, 1850, 1180]. work-diaries [1604]. Worker [1917]. Working [1936, 1935]. Works [2755, 2580, 2880, 711, 1394, 2405, 2427, 2163, 1748, 248, 1349, 950, 1474]. World [2621, 2385, 1273, 1765, 2708, 1386, 1714, 2787, 1633, 2820, 1474, 1293, 1995, 1971, 1501, 2415, 1932, 2475, 2234, 1859, 2611, 823, 1664, 1923, 868, 1776, 2291, 2490, 1178, 1775, 2466, 1825, 2137, 2260, 1528, 1592, 1731, 1714]. worlds [2004, 2449]. Worldwide [1577]. Worm [1611]. Worm-Holes [1611]. Woronzow [1447]. Wortley [2582]. Wotton [1292, 2748]. wound [2075]. Wren [561, 1505, 699, 521, 1704, 1505, 2752, 325, 2589, 2844, 597, 2110, 379]. Wright [1673, 2122, 1221]. writes [328]. writing [2410, 2361, 2388, 466]. Writings [2890, 2892, 1314, 1505, 2771, 2844, 1003, 1097, 1546]. wrong [2489, 1262]. Wrottesley [598]. Wroxhall [1720]. Wyhe [1785]. Wytham [2501].

X [1300, 1185, 1410, 1379, 1765, 1100, 1331, 1096, 1356, 1687, 1252, 1138, 1294, 1142, 1290, 1353, 1382, 1339, 1248, 1594, 1655, 1524, 1654, 1213, 1316, 1548, 1116, 1748, 1529, 1228, 1769, 1087, 1609, 1088, 1248, 1251, 1388, 1433, 2680, 1766, 2560, 1457, 1207, 632, 1786, 1099, 2917, 1539, 1615, 1749, 1598, 1473, 1507, 1750, 1485, 1382, 1434]. X-Club [632]. X-ray [1339, 1609, 2680, 2560]. X-rays [2917, 1434]. xi [1639, 1119, 1746, 1337, 1120, 1613, 999, 1138, 1633, 1142, 1290, 1408, 1711]. xii [1248, 1184, 1275, 1388, 1163, 1673, 1520, 1731, 1229, 1593, 1316, 1233, 1667, 1333, 1102, 1504, 1549, 1211, 1473, 1546, 1507, 1436, 1631, 1382]. Xiii [1157, 1251, 1428, 1707, 1434, 1313, 1138, 1210, 1189, 1950, 2151, 1453, 1597, 1162, 1732]. XIV [1523, 1338, 2262, 1379, 1361, 1082, 1206, 1386, 1706, 1096, 1073, 1099, 1189, 1527, 1357, 1471, 1617, 1801, 1485, 1708, 1086, 1383]. xix [1547, 1189]. xl [1727]. xlii [1431]. xliii [1653]. xlix [1334]. xlvii [1433]. xv [1182, 1231, 1294, 1432, 1712]. xvi [2385, 1299, 1103, 1180, 1429, 1784, 1138, 1073, 1750, 1502, 1669, 1276, 1382]. xvii [1097, 1763, 1298, 1187, 1488, 1459]. xviii [1410, 1141, 1704, 1271, 1250, 1672]. xx [1616, 1767, 1730, 1786, 1189, 1749, 1598, 1526]. xxi [1190, 1085, 1688]. xxii [1551]. xxiii [1474]. xxiv [1592, 1782]. xxix [1069, 1117]. xxv [1069]. xxvi [1101, 1359, 1252]. xxvii [1118, 1489]. xxxii [1069]. xxxiii [1117]. xxxiv [1508]. xxxv [1409]. xxxvi [1425]. xxxvii [1458].

Yale [1184, 1520, 1297, 1355, 1083]. Yankee [2330]. Year [310, 2621, 1523, 1731, 2139, 533, 73]. years [883, 1884, 1071, 1960, 900, 2232, 2306, 1951, 1084, 2350, 1015, 1082, 1699, 2050, 1948, 1241, 2409, 42, 33, 2917, 2126, 890, 1822, 2831, 362, 1083, 1906, 2526, 1434]. Yeo [1338, 1688]. York [1069, 1273, 1103, 1673, 1229, 1593, 324, 1611, 1087, 2380, 651, 642]. Yorkshire [2570]. Youings [1317]. young [1115, 1004, 1419, 1032, 643, 1310]. youth [1217]. Yuri [438].

Zachary [2496]. Zealand [1615, 1245, 623, 1615, 1399]. Zealanders [1245]. Zeeman [1402]. Zeitung [2308]. Zénith [2473]. Zenographic [1561]. Zewail [1670]. Ziani [492]. Zilch [1634]. Ziman [1687]. Zinaida [1508]. Zollman [1150]. zone [2499]. Zoo [2089, 2573]. Zoologica [201]. Zoology [1591, 2910]. zum [1849].

References

Anonymous:1938:FMa

[1] Anonymous. Front matter. Notes and Records of the Royal Society of London, 1(1):1–7, April 1938. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

530928.

H:1938:PTL

[2] A. V. H. The Pilgrim Trust Lecture. Notes and Records of the Royal Society of London, 1(1):8, April 1938. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

530929.

Anonymous:1938:NAS

[3] Anonymous. The National Academy of Sciences. Notes and Records of the Royal Society of London, 1(1):9–10, April 1938. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/530930.

Simon:1938:AD

[4] John Simon. The anniversary dinner. Notes and Records of the Royal Society of London, 1(1):11–20, April 1938. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530931.

S:1938:SL

[5] C. S. S. The Society’s library. Notes and Records of the Royal Society of London, 1(1):21–27, April 1938. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

530932.

L:1938:FH

[6] H. G. L. A fragment of history. Notes and Records of the Royal Society of London, 1(1):28–31, April 1938. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

530933.

S:1938:NFH

[7] A. C. S. Notes on the foundation and history of the Royal Society. Notes and Records of the Royal Society of London, 1(1):32–36, April 1938. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/530934.

D:1938:AS

[8] J. D. G. D. The arms of the Society. Notes and Records of the Royal Society of London, 1(1):37–39, April 1938. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530935.

L:1938:GF

[9] H. G. L. The growth of the Fellowship. Notes and Records of the Royal Society of London, 1(1):40–48, April 1938. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530936.

Anonymous:1938:BM

[10] Anonymous. Back matter. Notes and Records of the Royal Society of London, 1(1):??, April 1938. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/530937.

Seward:1938:SN

[11] Albert C. Seward. Society’s notes. Notes and Records of the Royal Society of London, 1(2):49–55, October 1938. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/3087168.

Seward:1938:DFI

[12] Albert C. Seward. The dedication of the Franklin Institute of the State of Pennsylvania in honour of Benjamin Franklin. Notes and Records of the Royal Society of London, 1(2):55–60, October 1938. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/3087169.

Kerr:1938:SSF

[13] J. Graham Kerr. Societas Scientiarum Fennica. Notes and Records of the Royal Society of London, 1(2):60–62, October 1938. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/3087170.

Close:1938:IGC

[14] C. F. Arden Close. The International Geographical Congress at Amsterdam, July 1938. Notes and Records of the Royal Society of London, 1(2):62–64, October 1938. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/3087171.

Fleure:1938:ICA

[15] H. J. Fleure. The International Congress of Anthropological and Ethnological Sciences at Copenhagen, August 1938. Notes and Records of the Royal Society of London, 1(2):65, October 1938. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/3087172.

Bennett:1938:PAC

[16] G. T. Bennett. Pepys’ annual commemoration service. Notes and Records of the Royal Society of London, 1(2):65–67, October 1938. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/3087173.

Anonymous:1938:C

[17] Anonymous. The conversaziones in 1938. Notes and Records of the Royal Society of London, 1(2):67–69, October 1938. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/3087174.

Gordon:1938:LEJ

[18] M. H. Gordon. Letter to the Editor: Jenner and Napoleon. Notes and Records of the Royal Society of London, 1(2):69, October 1938. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/3087175.

Greenwood:1938:FLT

[19] M. Greenwood. The first life table. Notes and Records of the Royal Society of London, 1(2):70–72, October 1938. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/3087176.

Lyons:1938:SFP

[20] H. G. Lyons. The Society’s finances — Part I — 1662–1830. Notes and Records of the Royal Society of London, 1(2):73–87, October 1938. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/3087177.

Lyons:1938:LPR

[21] H. G. Lyons. The landed property of the Royal Society. Notes and Records of the Royal Society of London, 1(2):88–91, October 1938. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/3087178.

Robinson:1938:GL

[22] H. W. Robinson. Gleanings from the library — I. Notes and Records of the Royal Society of London, 1(2):92–95, October 1938. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/3087179.

Lyons:1938:AD

[23] H. G. Lyons. The Anniversary Dinner. Notes and Records of the Royal Society of London, 1(2):96–103, October 1938. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/3087180.

Kemp:1938:BOC

[24] S. W. Kemp. The Bermuda Oceanographical Committee. Notes and Records of the Royal Society of London, 1(2):104–112, October 1938. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/3087181.

Anonymous:1938:FMb

[25] Anonymous. Front matter. Notes and Records of the Royal Society of London, 1(2):??, October 1938. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

3087167.

Anonymous:1939:FM

[26] Anonymous. Front matter. Notes and Records of the Royal Society of London, 2(1):1, April 1939. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/3087241.

Anonymous:1939:SNa

[27] Anonymous. Society’s notes. Notes and Records of the Royal Society of London, 2(1):2–10, April 1939. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

3087242.

Clark:1939:KWG

[28] A. J. Clark and F. G. Donnan. Kaiser-Wilhelm Gesellschaft: Report to the Royal Society. Notes and Records of the Royal Society of London, 2 (1):11, April 1939. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/3087243.

Wood:1939:AD

[29] Kingsley Wood, Irving Langmuir, and D’Arcy Thompson. The Anniversary Dinner. Notes and Records of the Royal Society of London, 2(1): 12–24, April 1939. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/3087244.

Darwin:1939:RPM

[30] C. G. Darwin. The ‘reading’ of papers at meetings of Royal Society. Notes and Records of the Royal Society of London, 2(1):25–27, April 1939. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/3087245.

Moir:1939:PPD

[31] J. Reid Moir. A pioneer in palaeolithic discovery. Notes and Records of the Royal Society of London, 2(1):28–31, April 1939. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/3087246.

Anonymous:1939:PJP

[32] Anonymous. A portrait of Joseph Priestley, F.R.S., by James Millar, 1789. Notes and Records of the Royal Society of London, 2(1):32–33, April 1939. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/3087247.

Lyons:1939:THY

[33] H. G. Lyons. Two hundred years ago, 1739. Notes and Records of the Royal Society of London, 2(1):34–42, April 1939. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/3087248.

Lyons:1939:SFB

[34] H. G. Lyons. The Society’s first bequest. Notes and Records of the Royal Society of London, 2(1):43–46, April 1939. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/3087249.

Lyons:1939:SFP

[35] H. G. Lyons. The Society’s finances. Part II — 1831–1938. Notes and Records of the Royal Society of London, 2(1):47–67, April 1939. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/3087250.

Robinson:1939:GLI

[36] H. W. Robinson. Gleanings from the library — II: The board of longitude and the Royal Society. Notes and Records of the Royal Society of London, 2(1):68–70, April 1939. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/3087251.

Hill:1939:AER

[37] A. V. Hill. Age of election to the Royal Society. Notes and Records of the Royal Society of London, 2(1):71–73, April 1939. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/3087252.

Yule:1939:JWD

[38] G. Undy Yule. John Wallis, D.D., F.R.S., 1616–1703. Notes and Records of the Royal Society of London, 2(1):74–82, April 1939. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/3087253.

Anonymous:1939:PSJ

[39] Anonymous. Portrait of Sir Joseph Banks. Notes and Records of the Royal Society of London, 2(1):83, April 1939. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/3087254.

Anonymous:1939:RPC

[40] Anonymous. Recommendation from the Physics Committee. Notes and Records of the Royal Society of London, 2(1):84, April 1939. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/3087255.

Anonymous:1939:SNb

[41] Anonymous. Society’s notes. Notes and Records of the Royal Society of London, 2(2):85–91, November 1939. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531203.

Lyons:1939:OHY

[42] H. G. Lyons. One hundred years ago. 1839. Notes and Records of the Royal Society of London, 2(2):92–107, November 1939. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531204.

Lyons:1939:CFC

[43] H. G. Lyons. The composition of the Fellowship and the Council of the Society. Notes and Records of the Royal Society of London, 2(2):108–126, November 1939. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531205.

Anonymous:1939:C

[44] Anonymous. The conversaziones of 1939. Notes and Records of the Royal Society of London, 2(2):127–139, November 1939. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/531206.

Proudman:1939:ARS

[45] J. Proudman. The association of the Royal sociey with progress in knowledge of oceanic tides. Notes and Records of the Royal Society of London, 2(2):140–143, November 1939. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531207.

Anonymous:1939:RRS

[46] Anonymous. The record of the Royal Society. Notes and Records of the Royal Society of London, 2(2):144–148, November 1939. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531208.

Gardiner:1939:BE

[47] J. Stanley Gardiner. Biological expeditions. Notes and Records of the Royal Society of London, 2(2):149–159, November 1939. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531209.

Oliver:1939:LF

[48] F. W. Oliver. Libyan flowers. Notes and Records of the Royal Society of London, 2(2):160–172, November 1939. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531210.

Trotter:1939:MW

[49] Wilfred Trotter. The mind in war. Notes and Records of the Royal Society of London, 2(2):173–175, November 1939. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531211.

Anonymous:1939:RSC

[50] Anonymous. The Royal Society and the Central Register. Notes and Records of the Royal Society of London, 2(2):176–178, November 1939. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531212.

Anonymous:1939:INR

[51] Anonymous. Index to Notes and Records, volume 2. Notes and Records of the Royal Society of London, 2(2):179–182, November 1939. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531201.

Anonymous:1939:BM

[52] Anonymous. Back matter. Notes and Records of the Royal Society of London, 2(2):??, November 1939. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531213.

Anonymous:1940:SNa

[53] Anonymous. Society’s notes. Notes and Records of the Royal Society of London, 3(1):2–10, April 1940. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:1940:VFS

[54] Anonymous. Visit of French scientists. Notes and Records of the Royal Society of London, 3(1):11–21, April 1940. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Turnball:1940:ESR

[55] H. W. Turnball. Early Scottish relations with the Royal Society I. — James Gregory, F.R.S. (1635–1675). Notes and Records of the Royal Society of London, 3(1):22–38, April 1940. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Plummer:1940:JHH

[56] H. C. Plummer. Jeremiah Horrocks and his Opera Posthuma. Notes and Records of the Royal Society of London, 3(1):39–52, April 1940. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Thorpe:1940:SHD

[57] Jocelyn Thorpe. Stephen Hales, D.D., F.R.S. 1677–1761. Notes and Records of the Royal Society of London, 3(1):53–63, April 1940. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Hatfield:1940:ARS

[58] W. H. Hatfield. The association of the Royal Society with the iron and steel industry. Notes and Records of the Royal Society of London, 3(1): 64–79, April 1940. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Lyons:1940:FT

[59] H. G. Lyons. The Fairchild Trust. Notes and Records of the Royal Society of London, 3(1):80–84, April 1940. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Davies:1940:BF

[60] J. D. Griffith Davies. The Banks family. Notes and Records of the Royal Society of London, 3(1):85–87, April 1940. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:1940:BNF

[61] Anonymous. Biographical notes — Francis Ashton, 1645–1715. Notes and Records of the Royal Society of London, 3(1):88–92, April 1940. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Lyons:1940:BNR

[62] H. G. Lyons. Biographical notes — Richard Waller (about 1650–1715). Notes and Records of the Royal Society of London, 3(1):92–94, April 1940. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Lyons:1940:BNJ

[63] H. G. Lyons. Biographical notes — John Lewis Guillemard (1764–1844). Notes and Records of the Royal Society of London, 3(1):95–96, April 1940. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:1940:LE

[64] Anonymous. Letter to the editor. Notes and Records of the Royal Society of London, 3(1):96, April 1940. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:1940:SNb

[65] Anonymous. Society’s notes. Notes and Records of the Royal Society of London, 3(1):97–104, April 1940. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:1940:GAP

[66] Anonymous. Gift from the American Philosophical Society. Notes and Records of the Royal Society of London, 3(1):105–106, April 1940. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Aydelotte:1940:SP

[67] Dr Frank Aydelotte. Science and politics. Notes and Records of the Royal Society of London, 3(1):107–109, April 1940. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Lyons:1940:JWJ

[68] H. G. Lyons. John Winthorp (Junior), F.R.S., Governor of Connecticut 1660 to 1676. Notes and Records of the Royal Society of London, 3(1): 110–115, April 1940. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Lyons:1940:OS

[69] H. G. Lyons. The officers of the Society (1662–1860). Notes and Records of the Royal Society of London, 3(1):116–140, April 1940. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Seward:1940:CHR

[70] Albert C. Seward. Christ’s hospital and the Royal Society. Notes and Records of the Royal Society of London, 3(1):141–145, April 1940. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Lyons:1940:CBO

[71] H. G. Lyons. Charles Babbage and the ophthalmoscope. Notes and Records of the Royal Society of London, 3(1):146–148, April 1940. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531152.

Lloyd:1940:MTE

[72] Ll. S. Lloyd. Musical theory in the early Philosophical Transactions. Notes and Records of the Royal Society of London, 3(1):149–157, April 1940. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:1940:PNY

[73] Anonymous. The President’s New Year message to French scientists. Notes and Records of the Royal Society of London, 3(1):158, April 1940. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Young:1940:VCL

[74] R. Fitzgibbon Young. The visit of Comenius to London in 1641–1642 and its bearing on the origins of the Royal Society. Notes and Records of the Royal Society of London, 3(1):159–160, April 1940. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:1946:AS

[75] Anonymous. The activities of the Society. Notes and Records of the Royal Society of London, 4(1):1–15, April 1946. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531228.

Anonymous:1946:C

[76] Anonymous. Conversazione. Notes and Records of the Royal Society of London, 4(1):16–18, April 1946. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531229.

Attlee:1946:AD

[77] Attlee. The Anniversary Dinner. Notes and Records of the Royal Society of London, 4(1):19–35, April 1946. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531230.

D:1946:ORS

[78] J. D. G. D. Officers of the Royal Society in the House of Commons. Notes and Records of the Royal Society of London, 4(1):36–37, April 1946. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531231.

Anonymous:1946:PCS

[79] Anonymous. Pepys’s commemoration service. Notes and Records of the Royal Society of London, 4(1):38, April 1946. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531232.

Anonymous:1946:AWF

[80] Anonymous. Admission of women into the Fellowship of the Royal Society. Notes and Records of the Royal Society of London, 4(1):39–40, April 1946. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531233.

A:1946:PP

[81] E. N. da C. A. The Presidential portraits. Notes and Records of the Royal Society of London, 4(1):41–42, April 1946. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531234.

A:1946:RCN

[82] E. N. da C. A. The Röntgen celebrations, November 1945. Notes and Records of the Royal Society of London, 4(1):43–44, April 1946. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531235. See corrigendum [110].

Anonymous:1946:VSB

[83] Anonymous. The Vignoles snuff-box. Notes and Records of the Royal Society of London, 4(1):45–48, April 1946. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531236.

Banks:1946:LSJ

[84] Sir Joseph Banks. A letter of Sir Joseph Banks. Notes and Records of the Royal Society of London, 4(1):49–50, April 1946. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/531237.

Casey:1946:PRA

[85] Mr R. G. Casey. Presentation to the Royal Asiatic Society of Bengal of letters from Sir Joseph Banks. Notes and Records of the Royal Society of London, 4(1):51–57, April 1946. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531238.

F:1946:BBS

[86] L. L. F. The bicentenary of the birth of Sir William Jones, F.R.S., founder of the Royal Asiatic Society of Bengal. Notes and Records of the Royal Society of London, 4(1):58–62, April 1946. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531239.

Anonymous:1946:VIB

[87] Anonymous. Visit to India by the Biological Secretary. Notes and Records of the Royal Society of London, 4(1):63–64, April 1946. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531240.

Anonymous:1946:AFA

[88] Anonymous. Anniversaries of foreign Academies and Societies. Notes and Records of the Royal Society of London, 4(1):65–68, April 1946. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531241.

Anonymous:1946:CFA

[89] Anonymous. Correspondence with foreign Academies, Societies and other bodies. Notes and Records of the Royal Society of London, 4(1):69– 81, April 1946. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531242.

Anonymous:1946:VLC

[90] Anonymous. Visits to liberated countries by representatives of the Royal Society. Notes and Records of the Royal Society of London, 4(1):82– 99, April 1946. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531243.

Anonymous:1946:GTH

[91] Anonymous. A gift by Thomas Hunt Morgan, For. Mem. R.S. Notes and Records of the Royal Society of London, 4(1):100, April 1946. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531244.

Hill:1946:AWM

[92] Professor A. V. Hill. Address of welcome to the members of the Executive Commitee of the International Council of Scientific Unions meeting in the rooms of the Royal Society on 4 December 1945. Notes and Records of the Royal Society of London, 4(1):101–102, April 1946. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531245.

Tizard:1946:LR

[93] Sir Henry Tizard. Lord Rutherford. Notes and Records of the Royal Society of London, 4(1):103–108, April 1946. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/ stable/531246. Transcript of a BBC radio talk on 16 December 1945 by Sir Henry Tizard about Lord Rutherford.

R:1946:AGH

[94] H. W. R. Alfred George Hastings White (1859–1945). Notes and Records of the Royal Society of London, 4(1):109–112, April 1946. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531247.

R:1946:RWF

[95] P. D. R. Robert William Frederick Harrison (1858–1945). Notes and Records of the Royal Society of London, 4(1):113–120, April 1946. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531248.

Anonymous:1946:SNa

[96] Anonymous. Society’s notes. Notes and Records of the Royal Society of London, 4(1):121–132, April 1946. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531249.

Anonymous:1946:FM

[97] Anonymous. Front matter. Notes and Records of the Royal Society of London, 4(1):??, April 1946. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531227.

Hill:1946:NRR

[98] A. V. Hill. Notes and records of the Royal Society of London. Notes and Records of the Royal Society of London, 4(2):133–145, October 1946. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/530939.

Dale:1946:IN

[99] Henry Dale. Isaac Newton: 1642–1727. Notes and Records of the Royal Society of London, 4(2):146–161, October 1946. CODEN NOREAY.

ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/530940.

Anonymous:1946:RSE

[100] Anonymous. The Royal Society Empire Scientific Conference. Notes and Records of the Royal Society of London, 4(2):162–167, October 1946. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/530941.

Stratton:1946:ICS

[101] F. J. M. Stratton. International Council of Scientific Unions. Notes and Records of the Royal Society of London, 4(2):168–173, October 1946. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/530942.

Dobell:1946:KP

[102] Clifford Dobell, F.R.S. The Kincardine papers. Notes and Records of the Royal Society of London, 4(2):174–178, October 1946. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/530943.

Davies:1946:HS

[103] John D. Griffith Davies. The homes of the Society. Notes and Records of the Royal Society of London, 4(2):179–192, October 1946. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/530944.

Robinson:1946:ASR

[104] H. W. Robinson. The administrative staff of the Royal Society. Notes and Records of the Royal Society of London, 4(2):193–205, October 1946. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/530945.

Robinson:1946:AOA

[105] President Robert Robinson. Anniversaries of other academies and societies. Notes and Records of the Royal Society of London, 4(2):206–211, October 1946. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/530946.

Anonymous:1946:NOS

[106] Anonymous. News of other scientific academies and societies. Notes and Records of the Royal Society of London, 4(2):212–215, October 1946.

CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/530947.

Anonymous:1946:CSM

[107] Anonymous. Charles Stewart Middlemiss. Notes and Records of the Royal Society of London, 4(2):215, October 1946. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530948.

DeBeer:1946:RVF

[108] G. R. De Beer. Rodolph Valltravers, F.R.S. Notes and Records of the Royal Society of London, 4(2):216–226, October 1946. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/530949. See addendum [188].

Anonymous:1946:SNb

[109] Anonymous. Society’s notes. Notes and Records of the Royal Society of London, 4(2):227–228, October 1946. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530950.

Anonymous:1946:RCC

[110] Anonymous. Röntgen celebrations: Corrigendum. Notes and Records of the Royal Society of London, 4(2):228, October 1946. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/530951. See [82].

Anonymous:1946:INR

[111] Anonymous. Index to Notes and Records volume, 4, 1946. Notes and Records of the Royal Society of London, 4(2):229–230, October 1946. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/530938.

Anonymous:1947:ASB

[112] Anonymous. The activities of the Society: British National Committee for Co-operation with UNESCO in Scientific Affairs. Notes and Records of the Royal Society of London, 5(1):1–4, October 1947. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531171.

Cripps:1947:AD

[113] Stafford Cripps. Anniversary Dinner 1946. Notes and Records of the Royal Society of London, 5(1):5–26, October 1947. CODEN NOREAY.

ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/531172.

Anonymous:1947:ISC

[114] Anonymous. The Indian Science Congress — Delhi Meeting, 1–8 January 1947. Notes and Records of the Royal Society of London, 5(1):27–31, October 1947. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531173.

Anonymous:1947:CSC

[115] Anonymous. The Chemical Society centenary celebrations. Notes and Records of the Royal Society of London, 5(1):32–33, October 1947. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531174.

Richardson:1947:WMH

[116] A. E. Richardson and H. W. Robinson. Woolsthorpe Manor House. Notes and Records of the Royal Society of London, 5(1):34–36, October 1947. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531175.

Taylor:1947:ESP

[117] F. Sherwood Taylor. An early satirical poem on the Royal Society. Notes and Records of the Royal Society of London, 5(1):37–46, October 1947. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531176.

Robinson:1947:DP

[118] H. W. Robinson. Denis Papin (1647–1712). Notes and Records of the Royal Society of London, 5(1):47–50, October 1947. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/531177.

Goodwin:1947:NOS

[119] A. J. H. Goodwin, Hon.General Secretary. News of other scientific academies and societies. Notes and Records of the Royal Society of London, 5(1):51–62, October 1947. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531178.

Anonymous:1947:SN

[120] Anonymous. Society’s notes. Notes and Records of the Royal Society of London, 5(1):63–64, October 1947. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531179.

Anonymous:1947:VI

[121] Anonymous. Volume information. Notes and Records of the Royal Society of London, 5(1):??, October 1947. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531170.

Robinson:1948:MHM

[122] Robert Robinson. Message to H. M. the King on the occasion of the marriage of Her Royal Highness the Princess Elizabeth, F.R.S. Notes and Records of the Royal Society of London, 5(2):65–66, April 1948. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531302.

Anonymous:1948:C

[123] Anonymous. Conversazione. Notes and Records of the Royal Society of London, 5(2):67–71, April 1948. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531303.

Anonymous:1948:RJ

[124] Anonymous. Reception on 17 July 1947. Notes and Records of the Royal Society of London, 5(2):72–73, April 1948. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531304.

Anonymous:1948:INO

[125] Anonymous. Isaac Newton Observatory. Notes and Records of the Royal Society of London, 5(2):74, April 1948. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531305.

Syfret:1948:ORS

[126] Miss R. H. Syfret. The origins of the Royal Society. Notes and Records of the Royal Society of London, 5(2):75–137, April 1948. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531306.

Robinson:1948:PNB

[127] H. W. Robinson. A preliminary note on the Blagden manuscripts. Notes and Records of the Royal Society of London, 5(2):137–139, April 1948.

CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531307.

Robinson:1948:AOA

[128] President Robert Robinson. Anniversaries of other academies and societies. Notes and Records of the Royal Society of London, 5(2):140–143, April 1948. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531308.

Eccles:1948:NOS

[129] J. C. Eccles. News of other scientific academies and societies. Notes and Records of the Royal Society of London, 5(2):144–155, April 1948. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531309.

Anonymous:1948:SCS

[130] Anonymous. Sir Charles Sherrington’s ninetieth birthday. Notes and Records of the Royal Society of London, 5(2):156, April 1948. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531310.

Sherrington:1948:SCS

[131] C. S. Sherrington. Sir Charles Sherrington’s first use of diphteria antitoxin made in England. Notes and Records of the Royal Society of London, 5(2):156–159, April 1948. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531311.

Anonymous:1948:SNa

[132] Anonymous. Society’s notes. Notes and Records of the Royal Society of London, 5(2):160–161, April 1948. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531312.

Anonymous:1948:INR

[133] Anonymous. Index to Notes and Records, vol. 5, 1947–1948. Notes and Records of the Royal Society of London, 5(2):162–163, April 1948. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531301.

Anonymous:1948:SIC

[134] Anonymous. The Scientific Information Conference. Notes and Records of the Royal Society of London, 6(1):1–7, December 1948. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531251.

Anonymous:1948:U

[135] Anonymous. UNESCO. Notes and Records of the Royal Society of London, 6(1):8–11, December 1948. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531252.

Fisher:1948:RSF

[136] James Fisher. ‘Rockall’ and ‘Seal’ flights 1947. Notes and Records of the Royal Society of London, 6(1):12–17, December 1948. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531253.

Humbert:1948:PTE

[137] Pierre Humbert. Pour le tricentenaire de l’expérience du Puy de Dôme. (French) [For the tercentenary of the Puy de Døme experiment]. Notes and Records of the Royal Society of London, 6(1):18–27, December 1948. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531254.

Douglas:1948:AIH

[138] McKie Douglas. The arrest and imprisonment of Henry Oldenburg. Notes and Records of the Royal Society of London, 6(1):28–47, December 1948. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531255.

Robinson:1948:RHS

[139] H. W. Robinson. Robert Hooke as a surveyor and architect. Notes and Records of the Royal Society of London, 6(1):48–55, December 1948. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531256.

DeBeer:1948:JGS

[140] G. R. De Beer. Johann Gaspar Scheuchzer, F.R.S. 1702–1729. Notes and Records of the Royal Society of London, 6(1):56–66, December 1948. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531257.

Anonymous:1948:RCP

[141] Anonymous. Rutherford Commemoration, Paris, 7 and 8 November 1947. Notes and Records of the Royal Society of London, 6(1):67–68, December 1948. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL https://royalsocietypublishing.org/journal/
rsnrcontent/6/1/67;   http://www.jstor.org/stable/531258.

Darwin:1948:AGS

[142] Sir Charles Darwin and Professor E. N. da C. Andrade. Address given by Sir Charles Darwin at the Max Planck memorial service held on 23 April 1948 in the aula of the University of Göttingen. Notes and Records of the Royal Society of London, 6(1):69–70, December 1948. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

URL http://www.jstor.org/stable/531259;   http://www.jstor.
org/stable/view/10.2307/531259.

Anonymous:1948:RMH

[143] Anonymous. Retirement of Mr. H. W. Robinson, Librarian to the Society. Notes and Records of the Royal Society of London, 6(1):71–72, December 1948. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531260.

Anonymous:1948:NOS

[144] Anonymous. News of other scientific academies and societies. Notes and Records of the Royal Society of London, 6(1):73–77, December 1948. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531261.

Anonymous:1948:SNb

[145] Anonymous. Society’s notes. Notes and Records of the Royal Society of London, 6(1):78–80, December 1948. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531262.

Anonymous:1948:FM

[146] Anonymous. Front matter. Notes and Records of the Royal Society of London, 6(1):??, December 1948. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531250.

Anonymous:1949:DSR

[147] Anonymous. Defence Services Research Facilities Committee. Notes and Records of the Royal Society of London, 6(2):81, May 1949. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531277.

Morrison:1949:AD

[148] Herbert Morrison. Anniversary Dinner 1948. Notes and Records of the Royal Society of London, 6(2):82–103, May 1949. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/531278.

Gagnebin:1949:BRC

[149] Bernard Gagnebin. Book review: De la cause de la pesanteur. Mémoire de Nicolas Fatio de Duillier présenté à la Royal Society le 26 février 1690. Notes and Records of the Royal Society of London, 6(2):105–124, May 1949. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531279.

DeDuillier:1949:TCP

[150] Nicolas Fatio De Duillier. Texte: De la cause de la pesanteur par Nicolas Fatio De Duillier, de la Société Royale D’Angleterre. Notes and Records of the Royal Society of London, 6(2):125–160, May 1949. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531280.

Hartley:1949:BCC

[151] Harold Hartley. The Berzelius centenary celebrations. Notes and Records of the Royal Society of London, 6(2):161–165, May 1949. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531281.

Robinson:1949:LPC

[152] Robert Robinson. Langevin–Perrin commemoration. Notes and Records of the Royal Society of London, 6(2):166, May 1949. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/531282.

M:1949:SN

[153] D. C. M. Society’s notes. Notes and Records of the Royal Society of London, 6(2):167–170, May 1949. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531283.

Anonymous:1949:INR

[154] Anonymous. Index to Notes and Records, vol. 6, 1948-1949. Notes and Records of the Royal Society of London, 6(2):171, May 1949. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531276.

Mckie:1949:ALL

[155] Douglas Mckie. Antoine Laurent Lavoisier, F.R.S., 1743–1794. Notes and Records of the Royal Society of London, 7(1):1–41, December 1949. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/530719.

Anonymous:1949:PP

[156] Anonymous. A problem picture. Notes and Records of the Royal Society of London, 7(1):42, December 1949. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

530720.

Cameron:1949:EJF

[157] G. R. Cameron, F.R.S. Edward Jenner, F.R.S., 1749–1823. Notes and Records of the Royal Society of London, 7(1):43–53, December 1949. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/530721.

Anonymous:1949:C

[158] Anonymous. Conversaziones 1949. Notes and Records of the Royal Society of London, 7(1):54–60, December 1949. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530722.

Salaman:1949:JRS

[159] Redcliffe N. Salaman, F.R.S. Jews in The Royal Society: A problem in ecology. Notes and Records of the Royal Society of London, 7(1):61–67, December 1949. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/530723.

Robinson:1949:ULD

[160] H. W. Robinson. An unpublished letter of Dr Seth Ward relating to the early meetings of the Oxford Philosophical Society. Notes and Records of the Royal Society of London, 7(1):68–70, December 1949. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/530724.

DeBeer:1949:MEG

[161] G. R. De Beer, F.R.S. The malady of Edward Gibbon, F.R.S. Notes and Records of the Royal Society of London, 7(1):71–80, December 1949. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/530725.

Kerr:1949:SMB

[162] Sir John Graham Kerr, F.R.S., M.P. The Scottish marine biological association. Notes and Records of the Royal Society of London, 7(1): 81–96, December 1949. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/530726.

Anonymous:1949:SP

[163] Anonymous. The Society’s portraits. Notes and Records of the Royal Society of London, 7(1):97–107, December 1949. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530727.

Anonymous:1949:NOS

[164] Anonymous. News of other scientific academies and societies — National Institute of Sciences of India. Notes and Records of the Royal Society of London, 7(1):108–112, December 1949. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530728.

Hinshelwood:1949:RCM

[165] Sir Cyril Hinshelwood, F.R.S., Professor W. E. Garner, F.R.S., and Professor H. D. Kay, F.R.S. Reports of celebrations, meetings and conferences. Notes and Records of the Royal Society of London, 7(1):113–117, December 1949. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/530729.

S:1949:NPR

[166] D. A. S., G. R. de B., and D. Mck. Notes of publications relating to the history of the Society. Notes and Records of the Royal Society of London, 7(1):118–125, December 1949. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

530730.

Robinson:1949:AOI

[167] Robert Robinson. Anniversaries of other institutions. Notes and Records of the Royal Society of London, 7(1):126–127, December 1949. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/530731.

Anonymous:1949:SN

[168] Anonymous. Society’s notes. Notes and Records of the Royal Society of London, 7(1):128–131, December 1949. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530732.

Anonymous:1949:VI

[169] Anonymous. Volume information. Notes and Records of the Royal Society of London, 7(1):??, December 1949. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

530718.

Anonymous:1950:PCR

[170] Anonymous. The pilot catalogue of the Royal Society Archives. Notes and Records of the Royal Society of London, 7(2):133–134, April 1950. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531314.

Trevelyan:1950:AD

[171] G. M. Trevelyan. Anniversary Dinner 1949. Notes and Records of the Royal Society of London, 7(2):135–157, April 1950. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/531315.

Donnan:1950:SRF

[172] F. G. Donnan, F.R.S. The Scientific Relief Fund and its Committee. Notes and Records of the Royal Society of London, 7(2):158–171, April 1950. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531316.

DeBeer:1950:EFR

[173] E. S. De Beer. The earliest Fellows of the Royal Society. Notes and Records of the Royal Society of London, 7(2):172–192, April 1950. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531317.

Anonymous:1950:JEP

[174] Anonymous. John Evelyn’s plan for a library. Notes and Records of the Royal Society of London, 7(2):193–194, April 1950. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531318.

DeBeer:1950:SLT

[175] G. R. De Beer, F.R.S. Some letters of Thomas Hobbes. Notes and Records of the Royal Society of London, 7(2):195–206, April 1950. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531319.

Syfret:1950:SER

[176] Miss R. H. Syfret. Some early reactions to the Royal Society. Notes and Records of the Royal Society of London, 7(2):207–258, April 1950. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531320.

Gibbs:1950:CMF

[177] F. W. Gibbs. Cromwell Mortimer, F.R.S. Notes and Records of the Royal Society of London, 7(2):259–263, April 1950. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531321.

DeBeer:1950:HBS

[178] G. R. De Beer, F.R.S. H.-B. De Saussure’s election into the Royal Society. Notes and Records of the Royal Society of London, 7(2):264–267, April 1950. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531322.

F:1950:NPR

[179] L. L. F. Notices of publications relating to the history of the Society. Notes and Records of the Royal Society of London, 7(2):268–271, April 1950. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531323.

Anonymous:1950:NOS

[180] Anonymous. News of other scientific academies and societies. Notes and Records of the Royal Society of London, 7(2):272–274, April 1950. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531324.

Anonymous:1950:SNa

[181] Anonymous. Society’s notes. Notes and Records of the Royal Society of London, 7(2):275, April 1950. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531325.

Anonymous:1950:INR

[182] Anonymous. Index to Notes and Records, vol. 7, 1949–1950. Notes and Records of the Royal Society of London, 7(2):276–277, April 1950. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531313.

Armitage:1950:RDE

[183] Angus Armitage. René Descartes (1596–1650) and the early Royal Society. Notes and Records of the Royal Society of London, 8(1):1–19, October 1950. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/3087231.

Syfret:1950:SEC

[184] Miss R. H. Syfret. Some early critics of the Royal Society. Notes and Records of the Royal Society of London, 8(1):20–64, October 1950. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/3087232.

DeB:1950:DSC

[185] G. R. De B. The diary of Sir Charles Blagden. Notes and Records of the Royal Society of London, 8(1):65–89, October 1950. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/3087233.

Anonymous:1950:RMA

[186] Anonymous. The Rumford medallists of the American Academy of Arts and Sciences. Notes and Records of the Royal Society of London, 8(1): 90–94, October 1950. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/3087234.

Caullery:1950:SFB

[187] M. Caullery, For.Mem.R.S. Les stations françaises de biologie marine. (French) [French stations of marine biology]. Notes and Records of the Royal Society of London, 8(1):95–115, October 1950. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/3087235.

DeB:1950:RVF

[188] G. R. De B. Rodolph Valltravers, F.R.S. (addendum). Notes and Records of the Royal Society of London, 8(1):116–119, October 1950. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/3087236. See [108].

Adrian:1950:CCS

[189] E. D. Adrian, For.Sec.R.S. Celebrations: Consejo Superior de Investigaciones Cientificas. Notes and Records of the Royal Society of London, 8 (1):120–121, October 1950. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/3087237.

Hill:1950:DLL

[190] A. V. Hill and Edwin G. Conklin, President. Darwin’s letters to Lyell. Notes and Records of the Royal Society of London, 8(1):122–124, October 1950. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/3087238.

deB:1950:NPR

[191] G. R. de B. Notice of a publication relating ro the history of the Society, William Stukeley. Notes and Records of the Royal Society of London, 8 (1):125–126, October 1950. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/3087239.

Anonymous:1950:SNb

[192] Anonymous. Society’s notes. Notes and Records of the Royal Society of London, 8(1):127–130, October 1950. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/3087240.

Anonymous:1950:FM

[193] Anonymous. Front matter. Notes and Records of the Royal Society of London, 8(1):??, October 1950. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

3087230.

Simon:1951:AD

[194] Viscount Simon and E. D. Adrian. Anniversary Dinner 1950. Notes and Records of the Royal Society of London, 8(2):131–148, April 1951. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/3087196.

Kaye:1951:UEM

[195] I. Kaye. Unrecorded early meetings of the Royal Society. Notes and Records of the Royal Society of London, 8(2):149–166, April 1951. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/3087197.

Hall:1951:RHH

[196] A. R. Hall. Robert Hooke and horology. Notes and Records of the Royal Society of London, 8(2):167–177, April 1951. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/3087198.

Stearns:1951:CFR

[197] Raymond Phineas Stearns. Colonial Fellows of the Royal Society of London, 1661–1788. Notes and Records of the Royal Society of London, 8 (2):178–246, April 1951. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/3087199.

DeBeer:1951:VFR

[198] G. R. De Beer, F.R.S. Voltaire, F.R.S. Notes and Records of the Royal Society of London, 8(2):247–252, April 1951. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/3087200.

deB:1951:SLS

[199] G. R. de B. Some letters of Sir Charles Blagden. Notes and Records of the Royal Society of London, 8(2):253–260, April 1951. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/3087201.

Swinton:1951:GMM

[200] W. E. Swinton. Gideon Mantell and the Maidstone Iguanodon. Notes and Records of the Royal Society of London, 8(2):261–276, April 1951. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/3087202.

Dohrn:1951:SZN

[201] Reinhard Dohrn. Stazione Zoologica Napoli. Notes and Records of the Royal Society of London, 8(2):277–282, April 1951. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/3087203.

Anonymous:1951:DPM

[202] Anonymous. Dedication of the Parsons Memorial Window. Notes and Records of the Royal Society of London, 8(2):283–290, April 1951. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/3087204.

Anonymous:1951:CNC

[203] Anonymous. Congress: National Congress of Chemistry organized by the Italian Chemical Society jointly with the International Congress of Applied Chemistry. Notes and Records of the Royal Society of London, 8(2):291–292, April 1951. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/3087205.

Davies:1951:RW

[204] John D. Griffith Davies. Ronald Winckworth 1884–1950. Notes and Records of the Royal Society of London, 8(2):293–296, April 1951. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/3087206.

deB:1951:NPRa

[205] G. R. de B. Notices of publications relating to the history of the Society. Notes and Records of the Royal Society of London, 8(2):297–304, April 1951. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/3087207.

Anonymous:1951:SNa

[206] Anonymous. Society’s notes. Notes and Records of the Royal Society of London, 8(2):305, April 1951. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

3087208.

Anonymous:1951:INR

[207] Anonymous. Index to Notes and Records, vol. 8, 1950–1951. Notes and Records of the Royal Society of London, 8(2):306, April 1951. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/3087194.

Anonymous:1951:FMa

[208] Anonymous. Front matter. Notes and Records of the Royal Society of London, 8(2):??, April 1951. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/3087195.

Maddison:1951:SLR

[209] R. E. W. Maddison. Studies in the life of Robert Boyle, F.R.S. Part I. Robert Boyle and some of his foreign visitors. Notes and Records of the Royal Society of London, 9(1):1–35, October 1951. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/3087257.

vanSeters:1951:AVL

[210] Dr W. H. van Seters. Antoni van Leeuwenhoek in Amsterdam. Notes and Records of the Royal Society of London, 9(1):36–45, October 1951. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/3087258.

Mckie:1951:NA

[211] D. Mckie and G. R. De Beer. Newton’s apple. Notes and Records of the Royal Society of London, 9(1):46–54, October 1951. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/3087259.

Cope:1951:CMJ

[212] Thomas D. Cope and H. W. Robinson. Charles Mason, Jeremiah Dixon and the Royal Society. Notes and Records of the Royal Society of London, 9(1):55–78, October 1951. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/3087260.

Aubert:1951:AAF

[213] Théodore Aubert. Alexander Aubert, F.R.S. astronome, 1730–1805. Notes and Records of the Royal Society of London, 9(1):79–95, October 1951. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/3087261.

DeBeer:1951:JSF

[214] G. R. De Beer, F.R.S. John Strange, F.R.S., 1732–1799. Notes and Records of the Royal Society of London, 9(1):96–108, October 1951. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/3087262.

Cameron:1951:LA

[215] H. Charles Cameron, M.D., F.R.C.P. The last of the alchemists. Notes and Records of the Royal Society of London, 9(1):109–114, October 1951. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/3087263.

Gibbs:1951:JGM

[216] F. W. Gibbs. John Gillies, M.D., traveller and botanist, 1792–1834. Notes and Records of the Royal Society of London, 9(1):115–136, October 1951. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/3087264.

Anonymous:1951:EHR

[217] Anonymous. Election of his Royal Highness the Duke of Edinburgh, K.G. Notes and Records of the Royal Society of London, 9(1):137, October 1951. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/3087265.

Anonymous:1951:C

[218] Anonymous. Conversaziones 1951. Notes and Records of the Royal Society of London, 9(1):138–143, October 1951. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/3087266.

Anonymous:1951:PCS

[219] Anonymous. Pepys’s commemoration service. Notes and Records of the Royal Society of London, 9(1):143, October 1951. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/3087267.

deB:1951:NPRb

[220] G. R. de B. Notices of publications relating to the history of the Royal Society. Notes and Records of the Royal Society of London, 9(1):144–147, October 1951. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/3087268.

Anonymous:1951:SNb

[221] Anonymous. Society’s notes. Notes and Records of the Royal Society of London, 9(1):148–151, October 1951. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/3087269.

Anonymous:1951:FMb

[222] Anonymous. Front matter. Notes and Records of the Royal Society of London, 9(1):??, October 1951. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

3087256.

Adrian:1952:DHM

[223] E. D. Adrian and O. P. Dawnay. Death of his Majesty King George VI. Notes and Records of the Royal Society of London, 9(2):153–154, May 1952. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/3087211.

Portal:1952:AD

[224] Portal and E. D. Adrian. Anniversary Dinner 1951. Notes and Records of the Royal Society of London, 9(2):155–163, May 1952. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/3087212.

Jacquot:1952:THR

[225] Jean Jacquot. Thomas Harriot’s reputation for impiety. Notes and Records of the Royal Society of London, 9(2):164–187, May 1952. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/3087213.

Jacquot:1952:NUW

[226] Jean Jacquot. Notes on an unpublished work of Thomas Hobbes. Notes and Records of the Royal Society of London, 9(2):188–195, May 1952. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/3087214.

Maddison:1952:SLRa

[227] R. E. W. Maddison. Studies in the life of Robert Boyle, F.R.S. Part II. Salt water freshened. Notes and Records of the Royal Society of London, 9(2):196–216, May 1952. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/3087215.

Cameron:1952:LTG

[228] G. R. Cameron, F.R.S. The life and times of Giambattista Morgagni, F.R.S. 1682–1771. Notes and Records of the Royal Society of London, 9(2):217–243, May 1952. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/3087216.

DeBeer:1952:RBF

[229] G. R. De Beer, F.R.S. The relations between Fellows of the Royal Society and French men of science when France and Britain were at war. Notes and Records of the Royal Society of London, 9(2):244–299, May 1952. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/3087217.

Manley:1952:WDS

[230] Gordon Manley. The weather and diseases: some eighteenth-century contributions to observational meteorology. Notes and Records of the Royal Society of London, 9(2):300–307, May 1952. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/3087218.

Sweet:1952:BFP

[231] Jessie M. Sweet. Benjamin Franklin’s purse. Notes and Records of the Royal Society of London, 9(2):308–309, May 1952. CODEN NOREAY.

ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/3087219.

Gilbert:1952:WHW

[232] L. F. Gilbert. W. H. Wollaston MSS. at Cambridge. Notes and Records of the Royal Society of London, 9(2):311–332, May 1952. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/3087220.

Mckie:1952:NAA

[233] D. Mckie and G. R. De Beer, F.R.S. Newton’s apple — an addendum. Notes and Records of the Royal Society of London, 9(2):333–335, May 1952. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/3087221.

Punnett:1952:WBM

[234] R. C. Punnett, F.R.S. William Bateson and Mendel’s principles of heredity. Notes and Records of the Royal Society of London, 9(2):336–347, May 1952. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/3087222.

McCance:1952:IFP

[235] R. A. McCance, F.R.S. An interesting but forgotten publication of Frederick Gowland Hopkins, F.R.S. Notes and Records of the Royal Society of London, 9(2):348–352, May 1952. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

3087223.

Anonymous:1952:AFM

[236] Anonymous. Award of the Franklin Medal to Sir James Chadwick, F.R.S. Notes and Records of the Royal Society of London, 9(2):353–355, May 1952. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/3087224.

Martin:1952:VAS

[237] D. C. Martin. Visit by the Assistant Secretary to some North American scientific institutions. Notes and Records of the Royal Society of London, 9(2):356–364, May 1952. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/3087225.

Anonymous:1952:FRU

[238] Anonymous. Facilities for research in the University College of the West Indies. Notes and Records of the Royal Society of London, 9(2):365– 377, May 1952. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/3087226.

deB:1952:NPR

[239] G. R. de B. Notices of publications relating to The Society. Notes and Records of the Royal Society of London, 9(2):378–380, May 1952. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/3087227.

Anonymous:1952:SNa

[240] Anonymous. Society’s notes. Notes and Records of the Royal Society of London, 9(2):381–384, May 1952. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

3087228.

Anonymous:1952:INR

[241] Anonymous. Index to Notes and Records, vol. 9, 1951–52. Notes and Records of the Royal Society of London, 9(2):386–387, May 1952. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/3087209.

Anonymous:1952:BM

[242] Anonymous. Back matter. Notes and Records of the Royal Society of London, 9(2):??, May 1952. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/3087229.

Anonymous:1952:C

[243] Anonymous. Conversaziones 1952. Notes and Records of the Royal Society of London, 10(1):1–5, October 1952. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531072.

Anonymous:1952:PBM

[244] Anonymous. The President’s broadcast on 6 March 1952 on Sir Charles Sherrington (1857–1952). Notes and Records of the Royal Society of London, 10(1):6–7, October 1952. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531073.

DeBeer:1952:AJP

[245] G. R. De Beer, F.R.S. Andreas and Joseph Planta, FF.R.S. Notes and Records of the Royal Society of London, 10(1):8–14, October 1952. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531074.

Maddison:1952:SLRb

[246] R. E. W. Maddison. Studies in the life of Robert Boyle, F.R.S. Part III. The charitable disposal of Robert Boyle’s residuary estate. Notes and Records of the Royal Society of London, 10(1):15–27, October 1952. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531075.

Munby:1952:DFE

[247] A. N. L. Munby. The distribution of the first edition of Newton’s Principia. Notes and Records of the Royal Society of London, 10(1):28–39, October 1952. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531076.

Munby:1952:KCW

[248] A. N. L. Munby. The Keynes Collection of the works of Sir Isaac Newton at King’s College, Cambridge. Notes and Records of the Royal Society of London, 10(1):40–50, October 1952. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531077.

McKie:1952:NPA

[249] D. McKie. A note on Priestley in America. Notes and Records of the Royal Society of London, 10(1):51–59, October 1952. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/531078.

Hartley:1952:ROW

[250] Sir Harold Hartley, F.R.S. Richard Owen and William Clift. Notes and Records of the Royal Society of London, 10(1):60–62, October 1952. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531079.

Anonymous:1952:SNb

[251] Anonymous. Society’s notes. Notes and Records of the Royal Society of London, 10(1):63, October 1952. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531080.

Anonymous:1952:FM

[252] Anonymous. Front matter. Notes and Records of the Royal Society of London, 10(1):??, October 1952. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531071.

Robinson:1953:AD

[253] Robert Robinson. Anniversary Dinner 1952. Notes and Records of the Royal Society of London, 10(2):65–70, April 1953. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/530801.

Hartley:1953:RC

[254] Sir Harold Hartley, F.R.S. The Ramsay centenary. Notes and Records of the Royal Society of London, 10(2):71–80, April 1953. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/530802.

DeBeer:1953:SHS

[255] G. R. De Beer, F.R.S. Sir Hans Sloane, F.R.S. 1660–1753. Notes and Records of the Royal Society of London, 10(2):81–84, April 1953. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/530803.

Jacquot:1953:SHS

[256] Jean Jacquot. Sir Hans Sloane and French men of science. Notes and Records of the Royal Society of London, 10(2):85–98, April 1953. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/530804.

Sweet:1953:SHS

[257] Jessie M. Sweet. Sir Hans Sloane’s metalline cubes. Notes and Records of the Royal Society of London, 10(2):99–100, April 1953. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/530805.

Turnbull:1953:SHI

[258] G. H. Turnbull. Samuel Hartlib’s influence on the early history of The Royal Society. Notes and Records of the Royal Society of London, 10 (2):101–130, April 1953. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/530806.

Atkinson:1953:DJR

[259] A. D. Atkinson. Dr Johnson and The Royal Society. Notes and Records of the Royal Society of London, 10(2):131–138, April 1953. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/530807.

Price:1953:CLA

[260] Derek J. Price. The Cavendish laboratory archives. Notes and Records of the Royal Society of London, 10(2):139–147, April 1953. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/530808.

DeBeer:1953:JMV

[261] G. R. De Beer, F.R.S. John Morgan’s visit to Voltaire. Notes and Records of the Royal Society of London, 10(2):148–158, April 1953. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/530809.

Anonymous:1953:BRB

[262] Anonymous. Bibliography of recent books and articles dealing with the history of the Royal Society or its Fellows. Notes and Records of the Royal Society of London, 10(2):159–160, April 1953. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530810.

Anonymous:1953:SN

[263] Anonymous. Society’s notes. Notes and Records of the Royal Society of London, 10(2):161–164, April 1953. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

530811.

Anonymous:1953:INR

[264] Anonymous. Index to Notes and Records, vol. io, 1952–53. Notes and Records of the Royal Society of London, 10(2):165–166, April 1953. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/530800.

Anonymous:1954:C

[265] Anonymous. Conversaziones 1953. Notes and Records of the Royal Society of London, 11(1):1–5, January 1954. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531094.

Salisbury:1954:AD

[266] Marquess of Salisbury. Anniversary Dinner 1953. Notes and Records of the Royal Society of London, 11(1):6–13, January 1954. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531095.

Hill:1954:AER

[267] A. V. Hill, F.R.S. Age of election to the Royal Society. Notes and Records of the Royal Society of London, 11(1):14–16, January 1954. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531096.

Salisbury:1954:PBM

[268] Edward Salisbury. The place of biology in modern education. Address by Sir Edward Salisbury, C.B.E., D.Sc., LL.D., V.P.R.S., at the opening of the new Hatherly Biological Laboratories, Exeter, on 23 April 1953. Notes and Records of the Royal Society of London, 11(1):17–21, January 1954. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531097.

Armytage:1954:RSA

[269] W. H. G. Armytage. The Royal Society and the Apothecaries 1660– 1722. Notes and Records of the Royal Society of London, 11(1):22–37, January 1954. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531098.

Maddison:1954:SLR

[270] R. E. W. Maddison. Studies in the life of Robert Boyle, F.R.S.: Part IV. Robert Boyle and some of his foreign visitors. Notes and Records of the Royal Society of London, 11(1):38–53, January 1954. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531099.

Armytage:1954:SGC

[271] W. H. G. Armytage. Sir Godfrey Copley, F.R.S., 1653–1709: Some tercentenary glimpses through letters to his friends. Notes and Records of the Royal Society of London, 11(1):54–74, January 1954. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531100.

Stearns:1954:FRS

[272] Raymond Phineas Stearns. Fellows of The Royal Society in North Africa and the Levant, 1662–1800. Notes and Records of the Royal Society of London, 11(1):75–90, January 1954. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531101.

Maddison:1954:SGC

[273] R. E. W. Maddison and Raymond E. Maddison. Spring Grove, the country house of Sir Joseph Banks, Bart, P.R.S. Notes and Records of the Royal Society of London, 11(1):91–99, January 1954. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531102.

Maddison:1954:NCM

[274] R. E. W. Maddison. A note on the correspondence of Martin Folkes, P.R.S. Notes and Records of the Royal Society of London, 11(1):100–109, January 1954. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531103.

Anonymous:1954:BRB

[275] Anonymous. Bibliography of recent books and articles dealing with the history of the Royal Society or its Fellows. Notes and Records of the Royal Society of London, 11(1):110, January 1954. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531104.

Anonymous:1954:SN

[276] Anonymous. Society’s notes. Notes and Records of the Royal Society of London, 11(1):111, January 1954. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531105.

Anonymous:1954:FM

[277] Anonymous. Front matter. Notes and Records of the Royal Society of London, 11(1):??, January 1954. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531093.

Anonymous:1955:C

[278] Anonymous. Conversaziones 1954. Notes and Records of the Royal Society of London, 11(2):113–119, March 1955. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530953.

Butler:1955:AD

[279] R. A. Butler. Anniversary Dinner 1954. Notes and Records of the Royal Society of London, 11(2):120–128, March 1955. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530954.

Hill:1955:JDG

[280] A. V. Hill, F.R.S. J. D. Griffith Davies 1899–1953 (Assistant Secretary of the Royal Society, 1937–1946). Notes and Records of the Royal Society of London, 11(2):129–133, March 1955. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530955.

MacGregor-Morris:1955:SAF

[281] J. T. MacGregor-Morris. Sir Ambrose Fleming (jubilee of the valve). Notes and Records of the Royal Society of London, 11(2):134–144, March 1955. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/530956.

Wightman:1955:AUR

[282] William P. D. Wightman. Aberdeen University and the Royal Society. Notes and Records of the Royal Society of London, 11(2):145–158, March 1955. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/530957.

Maddison:1955:SLR

[283] R. E. W. Maddison. Studies in the life of Robert Boyle, F.R.S. — Part V. Boyle’s operator: Ambrose Godfrey Hanckwitz, F.R.S. Notes and Records of the Royal Society of London, 11(2):159–188, March 1955. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/530958.

Dwight:1955:CR

[284] Professor C. Harrison Dwight. Count Rumford. Notes and Records of the Royal Society of London, 11(2):189–201, March 1955. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/530959.

Beatty:1955:SWT

[285] F. M. Beatty. The scientific work of the third Earl Stanhope. Notes and Records of the Royal Society of London, 11(2):202–221, March 1955. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/530960.

Cole:1955:BL

[286] F. J. Cole, F.R.S. Bell’s law. Notes and Records of the Royal Society of London, 11(2):222–227, March 1955. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530961.

Armytage:1955:GWF

[287] W. H. G. Armytage. G. W. Featherstonhaugh, F.R.S., 1780–1866, Anglo–American scientist. Notes and Records of the Royal Society of London, 11(2):228–235, March 1955. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

530962.

deBeer:1955:FAM

[288] Sir Gavin de Beer, F.R.S. and Max H. Hey, M.A., D.Sc. The first ascent of Mont Blanc. Notes and Records of the Royal Society of London, 11 (2):236–255, March 1955. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/530963.

Gilbert:1955:EPR

[289] L. F. Gilbert. The election to the presidency of the Royal Society in 1820. Notes and Records of the Royal Society of London, 11(2):256–279, March 1955. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/530964.

Anonymous:1955:SN

[290] Anonymous. Society’s notes. Notes and Records of the Royal Society of London, 11(2):280–283, March 1955. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530965.

Anonymous:1955:BRB

[291] Anonymous. Bibliography of recent books and articles dealing with the history of the Royal Society or its Fellows. Notes and Records of the Royal Society of London, 11(2):284–285, March 1955. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530966.

Anonymous:1955:INR

[292] Anonymous. Index to Notes and Records, vol. 11, 1954–5. Notes and Records of the Royal Society of London, 11(2):287–288, March 1955. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/530952.

Aldrich:1956:AD

[293] Winthrop W. Aldrich. Anniversary Dinner 1955. Notes and Records of the Royal Society of London, 12(1):1–8, August 1956. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/530787.

Anonymous:1956:C

[294] Anonymous. Conversaziones 1955. Notes and Records of the Royal Society of London, 12(1):9–20, August 1956. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530788.

Bluhm:1956:GAR

[295] R. K. Bluhm. A guide to the archives of the Royal Society and to other manuscripts in its possession. Notes and Records of the Royal Society of London, 12(1):21–39, August 1956. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

530789.

Atkinson:1956:RSE

[296] A. D. Atkinson. The Royal Society and English vocabulary. Notes and Records of the Royal Society of London, 12(1):40–43, August 1956. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/530790.

Taylor:1956:GBF

[297] Sir Geoffrey Taylor, F.R.S. George Boole, F.R.S., 1815–1864. Notes and Records of the Royal Society of London, 12(1):44–52, August 1956. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/530791.

Kneale:1956:BAL

[298] W. Kneale. Boole and the algebra of logic. Notes and Records of the Royal Society of London, 12(1):53–63, August 1956. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530792.

Armytage:1956:CWW

[299] W. H. G. Armytage. Charles Watson-Wentworth, second Marquess of Rockingham, F.R.S. (1730–1782): some aspects of his scientific interests. Notes and Records of the Royal Society of London, 12(1):64–76, August 1956. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/530793.

DeBeer:1956:JTF

[300] Sir Gavin De Beer, F.R.S. and R. M. Turton. John Turton, F.R.S., 1735– 1806. Notes and Records of the Royal Society of London, 12(1):77–97, August 1956. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/530794.

Maddison:1956:JPB

[301] R. E. W. Maddison and Francis R. Maddison. Joseph Priestley and the Birmingham riots. Notes and Records of the Royal Society of London, 12 (1):98–113, August 1956. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/530795.

McKie:1956:PLL

[302] Douglas McKie. Priestley’s laboratory and library and other of his effects. Notes and Records of the Royal Society of London, 12(1):114–136, August 1956. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/530796.

Anonymous:1956:RSA

[303] Anonymous. The Royal Society Antarctic expedition. Notes and Records of the Royal Society of London, 12(1):137–138, August 1956. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/530797.

Anonymous:1956:SN

[304] Anonymous. Society’s notes. Notes and Records of the Royal Society of London, 12(1):139–145, August 1956. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530798.

Anonymous:1956:BRB

[305] Anonymous. Bibliography of recent books and articles dealing with the history of the Royal Society or its Fellows. Notes and Records of the Royal Society of London, 12(1):146–148, August 1956. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/530799.

Anonymous:1956:FM

[306] Anonymous. Front matter. Notes and Records of the Royal Society of London, 12(1):??, August 1956. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

530786.

Anonymous:1957:HMQa

[307] Anonymous. Her Majesty Queen Elizabeth the Queen Mother, F.R.S. Notes and Records of the Royal Society of London, 12(2):149, December 1957. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/530826.

Hinshelwood:1957:ADS

[308] Sir Cyril Hinshelwood, F.R.S. Anniversary Dinner 1956: Speech by the President, Sir Cyril Hinshelwood, at the Anniversary Dinner, Dorchester Hotel, 30 November 1956. Notes and Records of the Royal Society of London, 12(2):150–153, December 1957. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

530827.

Anonymous:1957:C

[309] Anonymous. Conversaziones 1956 and 1957. Notes and Records of the Royal Society of London, 12(2):154–159, December 1957. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/530828.

M:1957:IIG

[310] D. C. M. The inauguration of the international geophysical year. Notes and Records of the Royal Society of London, 12(2):160–162, December 1957. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/530829.

Anonymous:1957:HMQb

[311] Anonymous. Her Majesty the Queen honours the Society’s I.G.Y. Antarctic Expedition. Notes and Records of the Royal Society of London, 12(2): 163–165, December 1957. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/530830.

Bullard:1957:ECE

[312] Sir Edward Bullard, F.R.S. and Colin A. Ronan. The exhibition to commemorate Edmond Halley 1656–1742. Notes and Records of the Royal Society of London, 12(2):166–167, December 1957. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/530831.

Chapman:1957:EHF

[313] S. Chapman, F.R.S. Edmund Halley, F.R.S. 1656–1742 (a commemorative lecture given on 21 November 1956, at the Royal Society’s celebration of Halley’s tercentenary). Notes and Records of the Royal Society of London, 12(2):168–174, December 1957. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

530832.

Jones:1957:HA

[314] Sir Harold Spencer Jones, F.R.S. Halley as an astronomer. Notes and Records of the Royal Society of London, 12(2):175–192, December 1957. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/530833.

McKie:1957:BBF

[315] Douglas McKie. Bernard Le Bovier De Fontenelle, F.R.S., 1657–1757. Notes and Records of the Royal Society of London, 12(2):193–200, December 1957. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/530834.

Thomson:1957:JJT

[316] George Thomson and Joan Thomson. J. J. Thomson as we remember him. Notes and Records of the Royal Society of London, 12(2):201–210, December 1957. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/530835.

Adrian:1957:SCS

[317] Lord Adrian, O.M., F.R.S. Sir Charles Scott Sherrington, O.M. 1857– 1952. Notes and Records of the Royal Society of London, 12(2):211–215, December 1957. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/530836.

Eccles:1957:SAS

[318] J. C. Eccles, F.R.S. Some aspects of Sherrington’s contribution to neurophysiology. Notes and Records of the Royal Society of London, 12(2): 216–225, December 1957. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/530837.

Armytage:1957:SEN

[319] W. H. G. Armytage. Science and education: a note. Notes and Records of the Royal Society of London, 12(2):226–229, December 1957. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/530838.

Thornton:1957:NVR

[320] For. Sec. R. S. H. G. Thornton. A note on the visit to Russia of the Royal Society delegation in 1956. Notes and Records of the Royal Society of London, 12(2):230–236, December 1957. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530839.

Anonymous:1957:SN

[321] Anonymous. Society’s notes. Notes and Records of the Royal Society of London, 12(2):237–241, December 1957. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530840.

Anonymous:1957:INR

[322] Anonymous. Index to Notes and Records, vol. 12, 1956–57. Notes and Records of the Royal Society of London, 12(2):243–244, December 1957. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/530825.

Hinshelwood:1958:AD

[323] Cyril Hinshelwood. Anniversary Dinner 1957. Notes and Records of the Royal Society of London, 13(1):1–5, June 1958. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531122.

McKie:1958:JDY

[324] Douglas McKie. James, Duke of York, F.R.S. Notes and Records of the Royal Society of London, 13(1):6–18, June 1958. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/531123.

Jones:1958:SCW

[325] H. W. Jones. Sir Christopher Wren and natural philosophy: with a checklist of his scientific activities. Notes and Records of the Royal Society of London, 13(1):19–37, June 1958. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531124.

Whyte:1958:RJB

[326] L. L. Whyte. R. J. Boscovich, S.J., F.R.S. (1711–1787), and the mathematics of atomism. Notes and Records of the Royal Society of London, 13(1):38–48, June 1958. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531125.

Pumphrey:1958:FMS

[327] R. J. Pumphrey, F.R.S. The forgotten man — Sir John Lubbock, F.R.S. Notes and Records of the Royal Society of London, 13(1):49–58, June 1958. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531126.

Duveen:1958:LWF

[328] Denis I. Duveen. Lavoisier writes to Fourcroy from prison. Notes and Records of the Royal Society of London, 13(1):59–60, June 1958. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531127.

Bluhm:1958:NOS

[329] R. K. Bluhm. A note on the origin of the Society’s conversaziones. Notes and Records of the Royal Society of London, 13(1):61–63, June 1958. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531128.

Anonymous:1958:SN

[330] Anonymous. Society’s notes. Notes and Records of the Royal Society of London, 13(1):64–68, June 1958. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531129.

Anonymous:1958:BRBa

[331] Anonymous. Bibliography of recent books and articles dealing with the history of the Royal Society or its Fellows. Notes and Records of the Royal Society of London, 13(1):69–72, June 1958. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531130.

Anonymous:1958:BMa

[332] Anonymous. Back matter. Notes and Records of the Royal Society of London, 13(1):??, June 1958. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531131.

Anonymous:1958:DWC

[333] Anonymous. The Darwin–Wallace conversazione. Notes and Records of the Royal Society of London, 13(2):73–74, November 1958. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/530969.

Anonymous:1958:C

[334] Anonymous. Conversaziones 1958. Notes and Records of the Royal Society of London, 13(2):75–81, November 1958. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530970.

Bluhm:1958:RRS

[335] R. K. Bluhm. Remarks on the Royal Society’s finances 1660–1768. Notes and Records of the Royal Society of London, 13(2):82–103, November 1958. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/530971.

Birley:1958:RBH

[336] Robert Birley. Robert Boyle’s Head Master at Eton. Notes and Records of the Royal Society of London, 13(2):104–114, November 1958. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/530972.

Hunter:1958:WHR

[337] Richard A. Hunter and Ida MacAlpine. William Harvey and Robert Boyle. Notes and Records of the Royal Society of London, 13(2):115–127, November 1958. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/530973.

Maddison:1958:TIC

[338] R. E. W. Maddison. A tentative index of the correspondence of the honourable Robert Boyle, F.R.S. Notes and Records of the Royal Society of London, 13(2):128–201, November 1958. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530974.

Hinshelwood:1958:SN

[339] C. N. Hinshelwood. Society’s notes. Notes and Records of the Royal Society of London, 13(2):202, November 1958. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530975.

Anonymous:1958:BRBb

[340] Anonymous. Bibliography of recent books and articles dealing with the history of the Royal Society or its Fellows. Notes and Records of the Royal Society of London, 13(2):203–205, November 1958. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/530976.

Anonymous:1958:INR

[341] Anonymous. Index to Notes and Records, vol. 13, 1958. Notes and Records of the Royal Society of London, 13(2):207, November 1958. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/530967.

Anonymous:1958:BMb

[342] Anonymous. Back matter. Notes and Records of the Royal Society of London, 13(2):??, November 1958. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

530977.

Anonymous:1959:FM

[343] Anonymous. Front matter. Notes and Records of the Royal Society of London, 14(1):1, June 1959. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/3087182.

Heathcoat-Amory:1959:AD

[344] D. Heathcoat-Amory. Anniversary Dinner 1958. Notes and Records of the Royal Society of London, 14(1):2–11, June 1959. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/3087183.

deBeer:1959:SUL

[345] Gavin de Beer. Some unpublished letters of Charles Darwin. Notes and Records of the Royal Society of London, 14(1):12–66, June 1959. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/3087184.

Pantin:1959:ARW

[346] C. F. A. Pantin. Alfred Russel Wallace, F.R.S., and his essays of 1858 and 1855. Notes and Records of the Royal Society of London, 14(1):67– 84, June 1959. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/3087185.

Barlow:1959:EDF

[347] Nora Barlow. Erasmus Darwin, F.R.S. (1731–1802). Notes and Records of the Royal Society of London, 14(1):85–98, June 1959. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/3087186.

Anonymous:1959:BM

[348] Anonymous. Back matter. Notes and Records of the Royal Society of London, 14(1):??, June 1959. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/3087193.

Anonymous:1960:ADS

[349] Anonymous. Anniversary Dinner 1958: Speech by the Rt Hon. D. Heathcoat-Amory, Chancellor of the Exchequer, at the Anniversary Dinner, Dorchester Hotel, 1 December 1958. Notes and Records of the Royal Society of London, 14(1):2–11, November 1960. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:1960:SUL

[350] Anonymous. Some unpublished letters of Charles Darwin. Notes and Records of the Royal Society of London, 14(1):12–66, November 1960. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Pantin:1960:ARW

[351] C. F. A. Pantin, F.R.S. Alfred Russel Wallace, F.R.S., and his essays of 1858 and 1855. Notes and Records of the Royal Society of London, 14 (1):67–84, November 1960. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Barlow:1960:EDF

[352] Nora Barlow. Erasmus Darwin, F.R.S. (1731–1802). Notes and Records of the Royal Society of London, 14(1):85–98, November 1960. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Habakkuk:1960:TRM

[353] H. J. Habakkuk. Thomas Robert Malthus, F.R.S. (1766–1834). Notes and Records of the Royal Society of London, 14(1):99–108, November 1960. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/3087187.

Turrill:1960:JDH

[354] W. B. Turrill. Joseph Dalton Hooker, F.R.S. (1817–1911). Notes and Records of the Royal Society of London, 14(1):109–120, November 1960. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/3087188.

Bailey:1960:CLF

[355] Sir Edward Bailey, F.R.S. Charles Lyell, F.R.S. (1797–1875). Notes and Records of the Royal Society of London, 14(1):121–138, November 1960. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/3087189.

Salisbury:1960:E

[356] Sir Edward Salisbury, F.R.S. Epilogue. Notes and Records of the Royal Society of London, 14(1):139–141, November 1960. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/3087190.

Anonymous:1960:SNa

[357] Anonymous. Society’s notes. Notes and Records of the Royal Society of London, 14(1):142–146, November 1960. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/3087191.

Anonymous:1960:BRBa

[358] Anonymous. Bibliography of recent books and articles dealing with the history of the Royal Society or its Fellows. Notes and Records of the Royal Society of London, 14(1):147–150, November 1960. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/3087192.

Chandos:1960:AD

[359] Viscount Chandos of Aldershot. Anniversary Dinner 1959. Notes and Records of the Royal Society of London, 14(2):151–155, June 1960. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531083.

Hinshelwood:1960:PS

[360] Cyril Hinshelwood. President’s speech, 1959. Notes and Records of the Royal Society of London, 14(2):156–159, June 1960. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/531084.

Anonymous:1960:VHE

[361] Anonymous. Visit of His Excellency the Soviet Ambassador, 19 November 1959. Notes and Records of the Royal Society of London, 14(2):160– 162, June 1960. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531085.

Wilson:1960:RME

[362] Charles Thomas Rees Wilson, C.H., F.R.S. Reminiscences of my early years. Notes and Records of the Royal Society of London, 14(2):163– 173, June 1960. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531086.

Maddison:1960:AWB

[363] R. E. W. Maddison. The accompt of William Balle from 28 November 1660 to 11 September 1663. Notes and Records of the Royal Society of London, 14(2):174–183, June 1960. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531087.

Page:1960:WNF

[364] R. I. Page. William Nicolson, F.R.S. and the runes of the Bewcastle Cross. Notes and Records of the Royal Society of London, 14(2):184– 190, June 1960. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531088.

Birley:1960:RBE

[365] Robert Birley. Robert Boyle at Eton. Notes and Records of the Royal Society of London, 14(2):191, June 1960. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531089.

Anonymous:1960:SNb

[366] Anonymous. Society’s notes. Notes and Records of the Royal Society of London, 14(2):192–195, June 1960. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531090.

Anonymous:1960:BRBb

[367] Anonymous. Bibliography of recent books and articles dealing with the history of the Royal Society. Notes and Records of the Royal Society of London, 14(2):196–198, June 1960. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531091.

Anonymous:1960:INR

[368] Anonymous. Index to Notes and Records, vol. 14, 1959. Notes and Records of the Royal Society of London, 14(2):199, June 1960. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531081.

Anonymous:1960:BM

[369] Anonymous. Back matter. Notes and Records of the Royal Society of London, 14(2):??, June 1960. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531092.

H:1960:P

[370] H. H. Preface. Notes and Records of the Royal Society of London, 15 (1):0, July 1960. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

McKie:1960:OFR

[371] Douglas McKie. The origins and foundation of the Royal Society of London. Notes and Records of the Royal Society of London, 15(1):1– 37, July 1960. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

deBeer:1960:KCI

[372] E. S. de Beer. King Charles II, fundator et patronus (1630–1685). Notes and Records of the Royal Society of London, 15(1):39–45, July 1960. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Bowen:1960:RRJ

[373] Dr E. J. Bowen, F.R.S. and Sir Harold Hartley, F.R.S. The Right Reverend John Wilkins, F.R.S. Notes and Records of the Royal Society of London, 15(1):47–56, July 1960. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Scott:1960:RJW

[374] J. F. Scott. The Reverend John Wallis, F.R.S. (1616–1703). Notes and Records of the Royal Society of London, 15(1):57–67, July 1960. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Copeman:1960:DJG

[375] W. S. C. Copeman. Dr Jonathan Goddard, F.R.S. Notes and Records of the Royal Society of London, 15(1):69–77, July 1960. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Masson:1960:SWP

[376] Sir Irvine Masson, F.R.S. and A. J. Youngson. Sir William Petty, F.R.S. (1623–1687). Notes and Records of the Royal Society of London, 15(1): 79–90, July 1960. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Symonds:1960:TWF

[377] Sir Charles Symonds. Thomas Willis, F.R.S. (1621–1675). Notes and Records of the Royal Society of London, 15(1):91–97, July 1960. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Summerson:1960:SCW

[378] Sir John Summerson and P. R. S. Sir Christopher Wren. Sir Christopher Wren, P. R. S. (1632–1723). Notes and Records of the Royal Society of London, 15(1):99–105, July 1960. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Whiteside:1960:WM

[379] Derek T. Whiteside. Wren the mathematician. Notes and Records of the Royal Society of London, 15(1):107–111, July 1960. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Ronan:1960:LR

[380] C. A. Ronan. Laurence Rooke (1622–1662). Notes and Records of the Royal Society of London, 15(1):113–118, July 1960. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Fulton:1960:HRB

[381] John F. Fulton. The Honourable Robert Boyle, F.R.S. (1627–1692). Notes and Records of the Royal Society of London, 15(1):119–135, July 1960. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Andrade:1960:RHF

[382] E. N. da C. Andrade, F.R.S. Robert Hooke, F.R.S. (1635–1703). Notes and Records of the Royal Society of London, 15(1):137–145, July 1960. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Scott:1960:WVB

[383] J. F. Scott and Sir Harold Hartley F. R. S. William, Viscount Brouncker, P.R.S. (1620–1684). Notes and Records of the Royal Society of London, 15(1):147–157, July 1960. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Ronan:1960:SPN

[384] C. A. Ronan and F. R. S. Sir Harold Hartley. Sir Paul Neile, F.R.S. (1613–1686). Notes and Records of the Royal Society of London, 15(1): 159–165, July 1960. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Armitage:1960:WBF

[385] Angus Armitage. William Ball. F.R.S. (1627–1690). Notes and Records of the Royal Society of London, 15(1):167–172, July 1960. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Maddison:1960:AHF

[386] R. E. W. Maddison. Abraham Hill, F.R.S. (1635–1722). Notes and Records of the Royal Society of London, 15(1):173–182, July 1960. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Bluhm:1960:HOF

[387] R. K. Bluhm. Henry Oldenburg, F.R.S. (c. 1615–1677). Notes and Records of the Royal Society of London, 15(1):183–197, July 1960. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Fulton:1960:SKD

[388] John F. Fulton. Sir Kenelm Digby, F.R.S. (1603–1665). Notes and Records of the Royal Society of London, 15(1):199–210, July 1960. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Payne:1960:WCF

[389] L. M. Payne, Leonard G. Wilson, and F. R. S. Sir Harold Hartley. William Croone, F.R.S. (1633–1684). Notes and Records of the Royal Society of London, 15(1):211–219, July 1960. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Josten:1960:EAF

[390] C. H. Josten. Elias Ashmole, F.R.S. Notes and Records of the Royal Society of London, 15(1):221–230, July 1960. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

DeBeer:1960:JEF

[391] E. S. De Beer. John Evelyn, F.R.S. Notes and Records of the Royal Society of London, 15(1):231–238, July 1960. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Martin:1960:SRM

[392] D. C. Martin. Sir Robert Moray, F.R.S. (1608?–1673). Notes and Records of the Royal Society of London, 15(1):239–250, July 1960. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Youngson:1960:ABF

[393] A. J. Youngson. Alexander Bruce, F.R.S., second Earl of Kincardine (1629–1681). Notes and Records of the Royal Society of London, 15(1): 251–258, July 1960. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Editor:1960:E

[394] The Editor. Epilogue. Notes and Records of the Royal Society of London, 15(1):259–264, July 1960. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:1961:FM

[395] Anonymous. Front matter. Notes and Records of the Royal Society of London, 16(1):i–v, April 1961. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

530753.

Florey:1961:FP

[396] H. W. Florey. Foreword by the President. Notes and Records of the Royal Society of London, 16(1):vii, April 1961. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530754.

Elizabeth:1961:FOT

[397] Queen Elizabeth II, Cyril Hinshelwood, D. Graffi, M. Jean Lecomte, and John Eccles. The formal opening of the tercentenary celebrations by Her Majesty The Queen, Patron, at the Royal Albert Hall on 19 July 1960. Notes and Records of the Royal Society of London, 16(1):1– 12, April 1961. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/530755.

Hinshelwood:1961:TAP

[398] Cyril Hinshelwood. The tercentenary address by the President of the Royal Society, Sir Cyril Hinshelwood, O.M., at the formal opening ceremony in the Royal Albert Hall, Tuesday 19 July 1960. Notes and Records of the Royal Society of London, 16(1):13–24, April 1961. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/530756.

Anonymous:1961:TCJ

[399] Anonymous. The tercentenary conversazione, 23 July 1960. Notes and Records of the Royal Society of London, 16(1):25–30, April 1961. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/530757.

Macmillan:1961:TBG

[400] Harold Macmillan and Cyril Hinshelwood. Tercentenary banquet, Grosvenor House, Tuesday 26 July 1960. The toast of the Royal Society proposed by The Rt. Hon. Harold Macmillan, M. P. Prime Minister and First Lord of the Treasury. Notes and Records of the Royal Society of London, 16(1):31–37, April 1961. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

530758.

Adrian:1961:TBG

[401] Lord Adrian, F. Cyril James, and A. H. T. Theorell. Tercentenary Banquet, Grosvenor House, Tuesday 26 July 1960: The Toast of the Guests Proposed by Lord Adrian F.R.S. Notes and Records of the Royal Society of London, 16(1):38–43, April 1961. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

530759.

Darlington:1961:CTH

[402] C. D. Darlington, F.R.S. The chromosomes and the theory of heredity: Tercentenary lecture delivered by C. D. Professor Darlington, F.R.S., at 10.15 a.m. on Wednesday 20 July 1960 at the Royal Institution. Notes and Records of the Royal Society of London, 16(1):44–48, April 1961. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/530760.

Hall:1961:TAS

[403] Sir Arnold Hall, F.R.S. Trends in aeronautical science and engineering: Tercentenary lecture delivered by Sir Arnold Hall, F.R.S., at 10.15 a.m. on Monday 25 July 1960 at the Royal Institution. Notes and Records of the Royal Society of London, 16(1):49–50, April 1961. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/530761.

Hinton:1961:ENP

[404] Sir Christopher Hinton, F.R.S. The evolution of nuclear power plant design. Notes and Records of the Royal Society of London, 16(1):51– 56, April 1961. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/530762.

Hodgkin:1961:PCN

[405] Professor A. L. Hodgkin, F.R.S. The physics and chemistry of nervous conduction: Tercentenary lecture delivered by Professor L. Hodgkin, F.R.S., at 2.30 p.m. on Wednesday 20 July 1960 at the Royal College of Surgeons. Notes and Records of the Royal Society of London, 16(1):57– 59, April 1961. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/530763.

Hodgkin:1961:MC

[406] Dr Dorothy Hodgkin, F.R.S. Molecules in crystals. Notes and Records of the Royal Society of London, 16(1):60–64, April 1961. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/530764.

Lovell:1961:IUR

[407] Professor A. C. B. Lovell, F.R.S. The investigation of the Universe by radio astronomy. Notes and Records of the Royal Society of London, 16(1):65–68, April 1961. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/530765.

Medawar:1961:PT

[408] Professor P. B. Medawar, F.R.S. The problems of transplantation. Notes and Records of the Royal Society of London, 16(1):69–71, April 1961. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/530766.

Powell:1961:SNI

[409] C. F. Powell, F.R.S. The study of nuclear interactions at very great energies: Tercentenary lecture delivered by Professor C. F. Powell, F.R.S., at 2.30 p.m. on Wednesday 20 July 1960 at Beveridge Hall, University of London. Notes and Records of the Royal Society of London, 16(1):72– 76, April 1961. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/530767.

Todd:1961:NHO

[410] Sir Alexander Todd, F.R.S. New horizons in organic chemistry. Notes and Records of the Royal Society of London, 16(1):77–80, April 1961. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/530768.

Wigglesworth:1961:MIT

[411] Professor V. B. Wigglesworth, F.R.S. The metamorphosis of insects: Tercentenary lecture delivered by Professor V. B. Wigglesworth, F.R.S., at 10.15 a.m. on Saturday 23 July 1960 at the Royal College of Surgeons. Notes and Records of the Royal Society of London, 16(1):81–84, April 1961. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/530769.

Anonymous:1961:R

[412] Anonymous. Receptions. Notes and Records of the Royal Society of London, 16(1):85, April 1961. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/530770.

Mother:1961:HDC

[413] Queen Elizabeth The Queen Mother and King Gustav VI Adolf. Honorary degree ceremonies. Notes and Records of the Royal Society of London, 16(1):86–89, April 1961. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

530771.

Anonymous:1961:V

[414] Anonymous. Visits. Notes and Records of the Royal Society of London, 16(1):90–91, April 1961. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/530772.

Anonymous:1961:LLC

[415] Anonymous. Luncheon at the London County Council. Notes and Records of the Royal Society of London, 16(1):92, April 1961. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/530773.

Hinshelwood:1961:PFA

[416] Cyril Hinshelwood. Première of films arranged by the Shell International Petroleum Company. Notes and Records of the Royal Society of London, 16(1):93–94, April 1961. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/530774.

Matthews:1961:SSP

[417] W. R. Matthews. Service at St. Paul’s Cathedral. Notes and Records of the Royal Society of London, 16(1):95–99, April 1961. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/530775.

Anonymous:1961:SE

[418] Anonymous. Special exhibitions. Notes and Records of the Royal Society of London, 16(1):100, April 1961. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

530776.

Anonymous:1961:SBP

[419] Anonymous. Special broadcast programmes. Notes and Records of the Royal Society of London, 16(1):101, April 1961. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/530777.

Anonymous:1961:SP

[420] Anonymous. Special publications. Notes and Records of the Royal Society of London, 16(1):102, April 1961. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

530778.

Anonymous:1961:CAP

[421] Anonymous. Congratulatory addresses presented to the Royal Society on the occasion of its tercentenary. Notes and Records of the Royal Society of London, 16(1):103–113, April 1961. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

530779.

Anonymous:1961:GPR

[422] Anonymous. Gifts presented to the Royal Society on the occasion of its tercentenary. Notes and Records of the Royal Society of London, 16 (1):114–116, April 1961. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/530780.

Penfield:1961:CTM

[423] Wilder Penfield. A Canadian table for the mace of the Royal Society. Notes and Records of the Royal Society of London, 16(1):117, April 1961. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/530781.

Anonymous:1961:ORW

[424] Anonymous. The official representatives who attended the Royal Society tercentenary celebrations. Notes and Records of the Royal Society of London, 16(1):118–122, April 1961. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

530782.

Anonymous:1961:CRS

[425] Anonymous. The Council of the Royal Society 1960. Notes and Records of the Royal Society of London, 16(1):123–124, April 1961. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/530783.

Hartley:1961:GCR

[426] Sir Harold Hartley, F.R.S. and Sir Cyril Hinshelwood, P.R.S. Gresham College and the Royal Society. Notes and Records of the Royal Society of London, 16(1):125–135, April 1961. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

530784.

Hartley:1961:DEF

[427] Sir Harold Hartley, F.R.S. The debt of engineering to Fellows of the Royal Society. Notes and Records of the Royal Society of London, 16 (1):136–140, April 1961. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/530785.

Anonymous:1961:I

[428] Anonymous. Index. Notes and Records of the Royal Society of London, 16(1):141–142, April 1961. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/530752.

Slim:1961:AD

[429] Viscount Slim. Anniversary Dinner 1960. Notes and Records of the Royal Society of London, 16(2):143–150, November 1961. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/530882.

Hill:1961:AER

[430] A. V. Hill, F.R.S. Age of election to the Royal Society. Notes and Records of the Royal Society of London, 16(2):151–153, November 1961. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/530883.

Wolbarsht:1961:CIR

[431] M. L. Wolbarsht and D. S. Sax. Charles II, a royal martyr. Notes and Records of the Royal Society of London, 16(2):154–157, November 1961. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/530884.

Wilson:1961:WCT

[432] L. G. Wilson. William Croone’s theory of muscular contraction. Notes and Records of the Royal Society of London, 16(2):158–178, November 1961. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/530885.

Bell:1961:SPH

[433] H. E. Bell. The Savilian professors’ houses and Halley’s observatory at Oxford. Notes and Records of the Royal Society of London, 16(2): 179–186, November 1961. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/530886.

Ross:1961:FCS

[434] S. Ross. Faraday consults the scholars: the origins of the terms of electrochemistry. Notes and Records of the Royal Society of London, 16(2): 187–220, November 1961. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/530887.

Williams:1961:RSF

[435] L. Pearce Williams. The Royal Society and the founding of the British association for the advancement of science. Notes and Records of the Royal Society of London, 16(2):221–233, November 1961. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/530888.

Threlfall:1961:SRT

[436] R. E. Threlfall. Sir Richard Threlfall, G.B.E., F.R.S. (1861–1932): some personal memories. Notes and Records of the Royal Society of London, 16(2):234–242, November 1961. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

530889.

Andrade:1961:HWR

[437] E. N. da C. Andrade, F.R.S. Henry William Robinson. Notes and Records of the Royal Society of London, 16(2):243–246, November 1961. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/530890.

Anonymous:1961:VMY

[438] Anonymous. Visit of Major Yuri Gagarin. Notes and Records of the Royal Society of London, 16(2):247, November 1961. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530891.

Anonymous:1961:IGA

[439] Anonymous. IX General Assembly of the International Council of Scientific Unions. Notes and Records of the Royal Society of London, 16(2): 248–250, November 1961. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/530892.

Anonymous:1961:SN

[440] Anonymous. Society’s notes. Notes and Records of the Royal Society of London, 16(2):251–255, November 1961. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530893.

Anonymous:1961:BRB

[441] Anonymous. Bibliography of recent books and articles dealing with the history of the Royal Society. Notes and Records of the Royal Society of London, 16(2):256–260, November 1961. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530894.

Anonymous:1961:INR

[442] Anonymous. Index to Notes and Records, volume 16, number 2, 1961. Notes and Records of the Royal Society of London, 16(2):261, November 1961. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/530881.

James:1962:AD

[443] Lord James. Anniversary Dinner 1961. Notes and Records of the Royal Society of London, 17(1):1–8, May 1962. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531009.

Wagner:1962:RSC

[444] Sir Anthony Wagner, K.C.V.O. The Royal Society’s coat of arms. Notes and Records of the Royal Society of London, 17(1):9–14, May 1962. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531010.

Linstead:1962:PCF

[445] Sir Patrick Linstead, Rector, For.Sec.R.S. The Prince Consort, F.R.S. and the founding of the Imperial College. Notes and Records of the Royal Society of London, 17(1):15–31, May 1962. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531011.

Hall:1962:WHL

[446] Marie Boas Hall. What happened to the Latin edition of Boyle’s History of Cold? Notes and Records of the Royal Society of London, 17(1):32– 35, May 1962. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531012.

Gibbs-Smith:1962:SGC

[447] Charles H. Gibbs-Smith. Sir George Cayley ‘Father of aerial navigation’ (1773–1857). Notes and Records of the Royal Society of London, 17(1): 36–56, May 1962. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531013.

Sharrock:1962:CPS

[448] Roger Sharrock. The chemist and the poet: Sir Humphry Davy and the preface to Lyrical Ballads. Notes and Records of the Royal Society of London, 17(1):57–76, May 1962. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531014.

McDonald:1962:STF

[449] Donald McDonald. Smithson Tennant, F.R.S. (1761–1815). Notes and Records of the Royal Society of London, 17(1):77–94, May 1962. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531015.

Anonymous:1962:DBS

[450] Anonymous. Dr Bruno Stulz, Assistant Librarian, 1943–46. Notes and Records of the Royal Society of London, 17(1):95, May 1962. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531016.

Anonymous:1962:SN

[451] Anonymous. Society’s notes. Notes and Records of the Royal Society of London, 17(1):96–99, May 1962. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531017.

Anonymous:1962:BRB

[452] Anonymous. Bibliography of recent books and articles dealing with the history of the Royal Society. Notes and Records of the Royal Society of London, 17(1):100–104, May 1962. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531018.

Anonymous:1962:VI

[453] Anonymous. Volume information. Notes and Records of the Royal Society of London, 17(1):??, May 1962. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531008.

G:1962:C

[454] J. A. G. and Sir Harold Hartley, F.R.S. Conversaziones, 1962. Notes and Records of the Royal Society of London, 17(2):105–109, December 1962. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531215.

Hartley:1962:RMT

[455] Sir Harold Hartley, F.R.S. Reception to mark the tercentenary of the granting of the Society’s first charter. Notes and Records of the Royal Society of London, 17(2):110, December 1962. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531216.

Hartley:1962:TRS

[456] Sir Harold Hartley, F.R.S. The tercentenary of the Royal Society’s charter. Notes and Records of the Royal Society of London, 17(2):111–116, December 1962. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531217. See corri-

gendum [475].

Needham:1962:SFG

[457] Dr Joseph Needham, F.R.S. and Sir Harold Hartley, F.R.S. Sir Frederick Gowland Hopkins, O.M., F.R.S. (1861–1947) Centenary Lecture held on 20 November 1961 in the University of Cambridge. Notes and Records of the Royal Society of London, 17(2):117–162, December 1962. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531218.

Penfield:1962:SCS

[458] Dr Wilder Penfield, O.M., F.R.S. and Sir Harold Hartley, F.R.S. Sir Charles Sherrington, O.M., F.R.S. (1857–1952): An appreciation. Notes and Records of the Royal Society of London, 17(2):163–168, December 1962. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531219.

Bragg:1962:SWB

[459] Sir Lawrence Bragg, F.R.S., Mrs G. M. Caroe (Gwendolen Bragg), and Sir Harold Hartley, F.R.S. Sir William Bragg, F.R.S. (1862–1942). Notes and Records of the Royal Society of London, 17(2):169–182, December 1962. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531220.

Somkin:1962:CSJ

[460] Fred. Somkin and Sir Harold Hartley, F.R.S. The contributions of Sir John Lubbock, F.R.S. to The Origin of Species. Some annotations to Darwin. Notes and Records of the Royal Society of London, 17(2):183– 191, December 1962. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/531221.

Brauer:1962:PJB

[461] Dr Adalbert J. Brauer and Sir Harold Hartley, F.R.S. Professor Johann Burchard Mencke, F.R.S. (1674–1732). Notes and Records of the Royal Society of London, 17(2):192–197, December 1962. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531222.

Dewhurst:1962:LCB

[462] Kenneth Dewhurst and Sir Harold Hartley, F.R.S. Locke’s contribution to Boyle’s researches on the air and on human blood. Notes and Records of the Royal Society of London, 17(2):198–206, December 1962. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531223.

Hartley:1962:AE

[463] Sir Harold Hartley, F.R.S. Addendum and errata. Notes and Records of the Royal Society of London, 17(2):207, December 1962. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531225.

Anonymous:1962:INR

[464] Anonymous. Index to Notes and Records, volume 17, 1962. Notes and Records of the Royal Society of London, 17(2):208, December 1962. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531214.

Mountbatten:1963:AD

[465] The Earl Mountbatten of Burma. Anniversary Dinner 1962. Notes and Records of the Royal Society of London, 18(1):1–9, June 1963. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/530896.

Westfall:1963:SWS

[466] Richard S. Westfall. Short-writing and the state of Newton’s conscience, 1662. Notes and Records of the Royal Society of London, 18(1):10–16, June 1963. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/530897.

Adrian:1963:NRT

[467] The Lord Adrian, O.M., F.R.S. Newton’s rooms in Trinity. Notes and Records of the Royal Society of London, 18(1):17–24, June 1963. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/530898.

Watson:1963:DSI

[468] George Watson. Dryden and the scientific image. Notes and Records of the Royal Society of London, 18(1):25–35, June 1963. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/530899.

Ellis:1963:AWC

[469] Frank H. Ellis. The author of Wing C6727: Daniel Coxe, F.R.S., or Thomas Coxe, F.R.S. Notes and Records of the Royal Society of London, 18(1):36–38, June 1963. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/530900.

Hall:1963:HMF

[470] Marie Boas Hall. Henry Miles, F.R.S. (1698–1763) and Thomas Birch, F.R.S. (1705–66). Notes and Records of the Royal Society of London, 18(1):39–44, June 1963. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/530901.

Thomas:1963:JWP

[471] Dr John Thomas. Josiah Wedgwood’s portrait medallions of Fellows of the Royal Society. Notes and Records of the Royal Society of London, 18(1):45–53, June 1963. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/530902.

Dee:1963:ICR

[472] P. I. Dee, F.R.S. and T. W. Wormell. An index to C. T. R. Wilson’s laboratory records and notebooks in the Library of the Royal Society. Notes and Records of the Royal Society of London, 18(1):54–66, June 1963. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/530903.

Anonymous:1963:SNa

[473] Anonymous. The Society’s notes. Notes and Records of the Royal Society of London, 18(1):67–71, June 1963. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

530904.

Anonymous:1963:BRBa

[474] Anonymous. Bibliography of recent books and articles dealing with the history of the Royal Society. Notes and Records of the Royal Society of London, 18(1):72–76, June 1963. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

530905.

Anonymous:1963:CNR

[475] Anonymous. Corrigendum: Notes and Records, volume 17, number 2, page 111. Notes and Records of the Royal Society of London, 18(1): 76, June 1963. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/530906. See [456].

Anonymous:1963:VI

[476] Anonymous. Volume information. Notes and Records of the Royal Society of London, 18(1):??, June 1963. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

530895.

Anonymous:1963:CR

[477] Anonymous. Conversaziones and receptions, 1963. Notes and Records of the Royal Society of London, 18(2):77–81, December 1963. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531264.

Andrade:1963:SPR

[478] Professor E. N. da C. Andrade, F.R.S. Samuel Pepys and the Royal Society. Notes and Records of the Royal Society of London, 18(2):82–93, December 1963. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531265.

Hall:1963:SHU

[479] A. Rupert Hall and Marie Boas Hall. Some hitherto unknown facts about the private career of Henry Oldenburg. Notes and Records of the Royal Society of London, 18(2):94–103, December 1963. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531266.

Maddison:1963:SLR

[480] R. E. W. Maddison. Studies in the life of Robert Boyle, F.R.S. — Part VI. The Stalbridge Period, 1645–1655, and the Invisible College. Notes and Records of the Royal Society of London, 18(2):104–124, December 1963. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531267.

Hoppen:1963:RSI

[481] K. Theodore Hoppen. The Royal Society and Ireland — William Molyneux, F.R.S. (1656–1698). Notes and Records of the Royal Society of London, 18(2):125–135, December 1963. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531268.

Craig:1963:INC

[482] Sir John Craig. Isaac Newton and the counterfeiters. Notes and Records of the Royal Society of London, 18(2):136–145, December 1963. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531269.

Quill:1963:JHC

[483] Colonel H. Quill. John Harrison, Copley Medallist, and the £20 000 longitude prize. Notes and Records of the Royal Society of London, 18(2): 146–160, December 1963. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531270.

Brush:1963:RSF

[484] Stephen G. Brush. The Royal Society’s first rejection of the kinetic theory of gases (1821), John Herapath versus Humphry Davy. Notes and Records of the Royal Society of London, 18(2):161–180, December 1963. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531271.

LeGrand:1963:SSR

[485] N. A. W. Le Grand. The Society’s Scientific Research in Schools Committee. Notes and Records of the Royal Society of London, 18(2):181–184, December 1963. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531272.

Anonymous:1963:SNb

[486] Anonymous. The Society’s notes. Notes and Records of the Royal Society of London, 18(2):185–186, December 1963. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531273.

Anonymous:1963:BRBb

[487] Anonymous. Bibliography of recent books and articles dealing with the history of the Royal Society. Notes and Records of the Royal Society of London, 18(2):187–191, December 1963. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531274.

G:1963:A

[488] C. G. Addendum. Notes and Records of the Royal Society of London, 18(2):192, December 1963. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531275.

Anonymous:1963:INR

[489] Anonymous. Index to Notes and Records, volume 18, 1963. Notes and Records of the Royal Society of London, 18(2):193–194, December 1963. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531263.

Robens:1964:ADD

[490] The Rt Hon.The Lord Robens of Woldingham. Anniversary Dinner, Dorchester Hotel, 30 November 1963. Notes and Records of the Royal Society of London, 19(1):1–9, June 1964. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/3519857.

Dale:1964:SMF

[491] Sir Henry Dale, O.M., G.B.E., F.R.S. Sir Michael Foster, K.C.B., F.R.S. — a Secretary of the Royal Society. Notes and Records of the Royal Society of London, 19(1):10–32, June 1964. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/3519858.

DeFerranti:1964:PEP

[492] S. Z. De Ferranti. Pioneer of electric power transmission: An account of some of the early work of Sebastian Ziani De Ferranti, D.Sc., F.R.S. Notes and Records of the Royal Society of London, 19(1):33–41, June 1964. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/3519859.

LeFanu:1964:SBB

[493] William R. Le Fanu. Sir Benjamin Brodie, F.R.S. (1783–1862). Notes and Records of the Royal Society of London, 19(1):42–52, June 1964. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/3519860.

Whiteside:1964:INB

[494] D. T. Whiteside. Isaac Newton: birth of a mathematician. Notes and Records of the Royal Society of London, 19(1):53–62, June 1964. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/3519861.

Glass:1964:JGH

[495] D. V. Glass. John Graunt and his natural and political observations. Notes and Records of the Royal Society of London, 19(1):63–100, June 1964. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/3519862.

Lawrie:1964:VMV

[496] J. J. Lawrie. Visit of Mrs Valentina Nikolaeva-Tereshkova. Notes and Records of the Royal Society of London, 19(1):101, June 1964. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/3519863.

Anonymous:1964:SNa

[497] Anonymous. The Society’s notes. Notes and Records of the Royal Society of London, 19(1):102–106, June 1964. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/3519864.

Anonymous:1964:BRBa

[498] Anonymous. Bibliography of recent books and articles dealing with the history of the Royal Society. Notes and Records of the Royal Society of London, 19(1):107–111, June 1964. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

3519865.

Anonymous:1964:FM

[499] Anonymous. Front matter. Notes and Records of the Royal Society of London, 19(1):??, June 1964. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

3519856.

Anonymous:1964:CR

[500] Anonymous. Conversaziones and receptions, 1964. Notes and Records of the Royal Society of London, 19(2):113–118, December 1964. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/3519846.

Anonymous:1964:CMQ

[501] Anonymous. Conversazione to mark the quater-centenary of the birth of Galileo. Notes and Records of the Royal Society of London, 19(2):119, December 1964. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/3519847.

Andrade:1964:G

[502] Professor E. N. da C. Andrade, F.R.S. Galileo. Notes and Records of the Royal Society of London, 19(2):120–130, December 1964. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/3519848.

Cohen:1964:QSE

[503] I. Bernard Cohen. ‘Quantum in Se Est’: Newton’s concept of inertia in relation to Descartes and Lucretius. Notes and Records of the Royal Society of London, 19(2):131–155, December 1964. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/3519849.

Craig:1964:RSR

[504] Sir John Craig. The Royal Society and the Royal mint. Notes and Records of the Royal Society of London, 19(2):156–167, December 1964. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/3519850.

Robson:1964:WWF

[505] R. Robson and Walter F. Cannon. William Whewell, F.R.S. (1794– 1866). Notes and Records of the Royal Society of London, 19(2):168–191, December 1964. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/3519851.

DeBeer:1964:MDF

[506] Sir Gavin De Beer. Mendel, Darwin, and Fisher (1865–1965). Notes and Records of the Royal Society of London, 19(2):192–226, December 1964. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/3519852. See erratum [537].

Anonymous:1964:SNb

[507] Anonymous. The Society’s notes. Notes and Records of the Royal Society of London, 19(2):227, December 1964. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/3519853.

Anonymous:1964:BRBb

[508] Anonymous. Bibliography of recent books and articles dealing with the history of the Royal Society. Notes and Records of the Royal Society of London, 19(2):228–231, December 1964. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/3519854.

Anonymous:1964:INR

[509] Anonymous. Index to Notes and Records, volume 19, 1964. Notes and Records of the Royal Society of London, 19(2):232–233, December 1964. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/3519845.

Wolfenden:1965:AD

[510] John Wolfenden. Anniversary Dinner 1964. Notes and Records of the Royal Society of London, 20(1):1–8, June 1965. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/3519879.

Andrade:1965:BED

[511] E. N. da C. Andrade, F.R.S. The birth and early days of the Philosophical Transactions. Notes and Records of the Royal Society of London, 20(1): 9–27, June 1965. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/3519880.

Deacon:1965:FMS

[512] Margaret Deacon. Founders of marine science in Britain: the work of the early Fellows of the Royal Society. Notes and Records of the Royal Society of London, 20(1):28–50, June 1965. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

3519881.

Maddison:1965:SLR

[513] R. E. W. Maddison. Studies in the life of Robert Boyle, F.R.S. Part VII. The grand tour. Notes and Records of the Royal Society of London, 20(1):51–77, June 1965. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/3519882.

Hoppen:1965:RSI

[514] K. Theodore Hoppen. The Royal Society and Ireland. II. Notes and Records of the Royal Society of London, 20(1):78–99, June 1965. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/3519883.

Levene:1965:NMF

[515] John R. Levene. Nevil Maskelyne, F.R.S., and the discovery of night myopia. Notes and Records of the Royal Society of London, 20(1):100– 108, June 1965. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/3519884.

Anonymous:1965:SNa

[516] Anonymous. The Society’s notes. Notes and Records of the Royal Society of London, 20(1):109–113, June 1965. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/3519885.

Anonymous:1965:BRBa

[517] Anonymous. Bibliography of recent books and articles dealing with the history of the Royal Society. Notes and Records of the Royal Society of London, 20(1):114–118, June 1965. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

3519886.

Anonymous:1965:FM

[518] Anonymous. Front matter. Notes and Records of the Royal Society of London, 20(1):??, June 1965. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

3519878.

Anonymous:1965:C

[519] Anonymous. Conversaziones, 1965. Notes and Records of the Royal Society of London, 20(2):119–124, December 1965. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/3519867.

Lohne:1965:INR

[520] J. A. Lohne. Isaac Newton: the rise of a scientist 1661–1671. Notes and Records of the Royal Society of London, 20(2):125–139, December 1965. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/3519868.

Hall:1965:WP

[521] A. Rupert Hall. Wren’s problem. Notes and Records of the Royal Society of London, 20(2):140–144, December 1965. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/3519869.

Middleton:1965:FHB

[522] W. E. Knowles Middleton. A footnote to the history of the barometer. Notes and Records of the Royal Society of London, 20(2):145–151, December 1965. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/3519870.

Axtell:1965:LRP

[523] James L. Axtell. Locke’s review of the Principia. Notes and Records of the Royal Society of London, 20(2):152–161, December 1965. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/3519871.

Allibone:1965:DJB

[524] T. E. Allibone, F.R.S. The diaries of John Byrom, M.A., F.R.S., and their relation to the pre-history of the Royal Society Club. Notes and Records of the Royal Society of London, 20(2):162–183, December 1965. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/3519872.

Ross:1965:SEI

[525] Sydney Ross. The search for electromagnetic induction, 1820–1831. Notes and Records of the Royal Society of London, 20(2):184–219, December 1965. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/3519873.

Hartley:1965:LRP

[526] Sir Harold Hartley, F.R.S. A letter from Richard Phillips, F.R.S. (1778– 1857) to Michael Faraday, F.R.S. (1791–1867). Notes and Records of the Royal Society of London, 20(2):220–223, December 1965. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/3519874.

Anonymous:1965:SNb

[527] Anonymous. The Society’s notes. Notes and Records of the Royal Society of London, 20(2):224, December 1965. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/3519875.

Anonymous:1965:BRBb

[528] Anonymous. Bibliography of recent books and articles dealing with the history of the Royal Society. Notes and Records of the Royal Society of London, 20(2):225–230, December 1965. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/3519876.

Anonymous:1965:INR

[529] Anonymous. Index to Notes and Records, volume 20, 1965. Notes and Records of the Royal Society of London, 20(2):231–232, December 1965. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/3519866.

Wilson:1966:AD

[530] Harold Wilson. Anniversary Dinner 1965. Notes and Records of the Royal Society of London, 21(1):1–11, June 1966. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530813.

P:1966:CMA

[531] W. D. M. P. Conversazione to mark the 300th anniversary of the publication of the Philosophical Transactions of the Royal Society, Hooke’s Micrographia, and Evelyn’s Sylva. Notes and Records of the Royal Society of London, 21(1):12–19, June 1966. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

530814.

Faure-Fremiet:1966:OLS

[532] For. Mem. R. S. E. Fauré-Fremiet. Les origines de L’Académie des sciences de Paris. Notes and Records of the Royal Society of London, 21(1): 20–31, June 1966. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/530815.

Whiteside:1966:NMY

[533] D. T. Whiteside. Newton’s marvellous year: 1666 and all that. Notes and Records of the Royal Society of London, 21(1):32–41, June 1966. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/530816.

Feisenberger:1966:LNH

[534] H. A. Feisenberger. The libraries of Newton, Hooke and Boyle. Notes and Records of the Royal Society of London, 21(1):42–55, June 1966. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/530817.

Hartley:1966:SCF

[535] Sir Harold Hartley, F.R.S. Stanislao Cannizzaro, F.R.S. (1826–1910) and the first International Chemical Conference at Karlsruhe in 1860. Notes and Records of the Royal Society of London, 21(1):56–63, June 1966. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/530818. See corrigenda [571].

deBeer:1966:MDF

[536] Sir Gavin de Beer, F.R.S. Mendel, Darwin, and Fisher: Addendum. Notes and Records of the Royal Society of London, 21(1):64–71, June 1966. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/530819.

Anonymous:1966:EGB

[537] Anonymous. Erratum: Gavin de Beer, Mendel, Darwin, and Fisher (1865–1965), Notes and Records 19(2) 192–226 (p. 212). Notes and Records of the Royal Society of London, 21(1):71, June 1966. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/530820. See [506].

S:1966:EDB

[538] N. S. Exchange of delegations between the Royal Society and The U.S.S.R. Academy of Sciences, 1965. Notes and Records of the Royal Society of London, 21(1):72–79, June 1966. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530821.

Anonymous:1966:SNa

[539] Anonymous. The Society’s notes. Notes and Records of the Royal Society of London, 21(1):80–85, June 1966. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

530822.

Anonymous:1966:BRBa

[540] Anonymous. Bibliography of recent books and articles dealing with the history of the Royal Society. Notes and Records of the Royal Society of London, 21(1):86–92, June 1966. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

530823.

Anonymous:1966:BM

[541] Anonymous. Back matter. Notes and Records of the Royal Society of London, 21(1):93, June 1966. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/530824.

Anonymous:1966:VI

[542] Anonymous. Volume information. Notes and Records of the Royal Society of London, 21(1):??, June 1966. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

530812.

Anonymous:1966:CR

[543] Anonymous. Conversaziones and reception, 1966. Notes and Records of the Royal Society of London, 21(2):95–101, December 1966. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531062.

Andrade:1966:PSM

[544] E. N. da C. Andrade, F.R.S. The preservation of scientific manuscripts and records. Notes and Records of the Royal Society of London, 21(2): 102–107, December 1966. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531063.

McGuire:1966:NPP

[545] J. E. McGuire and P. M. Rattansi. Newton and the ‘pipes of pan’. Notes and Records of the Royal Society of London, 21(2):108–143, December 1966. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531064. See corrigenda

[571].

Schofield:1966:LSB

[546] Robert E. Schofield. The Lunar Society of Birmingham; a bicentenary appraisal. Notes and Records of the Royal Society of London, 21(2): 144–161, December 1966. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531065.

Brock:1966:SAB

[547] W. H. Brock. The selection of the authors of the Bridgewater Treatises. Notes and Records of the Royal Society of London, 21(2):162–179, December 1966. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531066.

Levene:1966:SGB

[548] John R. Levene. Sir George Biddell Airy, F.R.S. (1801–1892) and the discovery and correction of astigmatism. Notes and Records of the Royal Society of London, 21(2):180–199, December 1966. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/531067.

Smith:1966:HLG

[549] D. M. Smith, F.R.S. Henry Lewis Guy, F.R.S., 1887–1956, turbine designer and engineering administrator. Notes and Records of the Royal Society of London, 21(2):200–206, December 1966. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/531068.

Anonymous:1966:SNb

[550] Anonymous. The Society’s notes. Notes and Records of the Royal Society of London, 21(2):207, December 1966. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531069.

Anonymous:1966:BRBb

[551] Anonymous. Bibliography of recent books and articles dealing with the history of the Royal Society. Notes and Records of the Royal Society of London, 21(2):207–211, December 1966. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531070.

Anonymous:1966:INR

[552] Anonymous. Index to Notes and Records, volume 21, 1966. Notes and Records of the Royal Society of London, 21(2):212–213, December 1966. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531061.

Anonymous:1967:AD

[553] Anonymous. Anniversary Dinner 1966. Notes and Records of the Royal Society of London, 22(1–2):1–11, September 1967. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/531182.

Martin:1967:FHR

[554] D. C. Martin. Former homes of the Royal Society. Notes and Records of the Royal Society of London, 22(1–2):12–19, September 1967. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531183.

Summerson:1967:CHT

[555] Sir John Summerson. Carlton House Terrace. Notes and Records of the Royal Society of London, 22(1–2):20–22, September 1967. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531184.

Holford:1967:NHR

[556] Lord Holford. The new home of the Royal Society. Notes and Records of the Royal Society of London, 22(1–2):23–36, September 1967. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531185.

Editor:1967:BHS

[557] The Editor. Britain’s heritage of scientific and technological records and manuscripts and of historic scientific instruments. Notes and Records of the Royal Society of London, 22(1–2):37–39, September 1967. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531186.

Thackray:1967:PFM

[558] Arnold Thackray. ‘In praise of famous men’ — the John Dalton bicentenary celebrations, 1966. Notes and Records of the Royal Society of London, 22(1–2):40–44, September 1967. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531187.

Guerlac:1967:NOA

[559] Henry Guerlac. Newton’s optical aether: his draft of a proposed addition to his Optiks. Notes and Records of the Royal Society of London, 22(1– 2):45–57, September 1967. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531188.

Scriba:1967:TIC

[560] Christoph J. Scriba. A tentative index of the correspondence of John Wallis, F.R.S. Notes and Records of the Royal Society of London, 22(1– 2):58–93, September 1967. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531189.

Biswas:1967:ARG

[561] Asit K. Biswas. The automatic rain-gauge of Sir Christopher Wren, F.R.S. Notes and Records of the Royal Society of London, 22(1–2):94– 104, September 1967. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/531190.

Turner:1967:PJS

[562] G. L’e. Turner. A portrait of James Short, F.R.S., attributable to Benjamin Wilson, F.R.S. Notes and Records of the Royal Society of London, 22(1–2):105–112, September 1967. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531191.

Smeaton:1967:LBG

[563] W. A. Smeaton. Louis Bernard Guyton de Morveau, F.R.S. (1737– 1816) and his relations with British scientists. Notes and Records of the Royal Society of London, 22(1–2):113–130, September 1967. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531192.

Cartwright:1967:ATB

[564] F. F. Cartwright. The association of Thomas Beddoes, M.D. with James Watt, F.R.S. Notes and Records of the Royal Society of London, 22(1–2): 131–143, September 1967. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531193.

Smith:1967:JKL

[565] Barbara M. D. Smith and J. L. Moilliet. James Keir of the Lunar Society. Notes and Records of the Royal Society of London, 22(1–2):144–154, September 1967. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531194.

Chaldecott:1967:CFR

[566] J. A. Chaldecott. Contributions of Fellows of the Royal Society to the fabrication of platinum vessels: some unpublished manuscripts. Notes and Records of the Royal Society of London, 22(1–2):155–172, September 1967. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531195.

Coley:1967:ACC

[567] N. G. Coley. The Animal Chemistry Club; assistant society to the Royal Society. Notes and Records of the Royal Society of London, 22(1–2):

173–185, September 1967. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531196.

Allibone:1967:CRC

[568] T. E. Allibone, F.R.S. The Club of the Royal College of Physicians, the Smeatonian Society of Civil Engineers and their relationship to the Royal Society Club. Notes and Records of the Royal Society of London, 22(1–2): 186–192, September 1967. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531197.

Anonymous:1967:SN

[569] Anonymous. The Society’s notes. Notes and Records of the Royal Society of London, 22(1–2):193–197, September 1967. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531198.

Anonymous:1967:BRB

[570] Anonymous. Bibliography of recent books and articles dealing with the history of the Royal Society. Notes and Records of the Royal Society of London, 22(1–2):198–206, September 1967. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531199.

Anonymous:1967:CHM

[571] Anonymous. Corrigenda for Hartley, McGuire and Rattansi. Notes and Records of the Royal Society of London, 22(1–2):207, September 1967. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531200. See [535, 545].

Anonymous:1967:INR

[572] Anonymous. Index to Notes and Records, volume 22, 1967. Notes and Records of the Royal Society of London, 22(1/2):209–210, September 1967. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531180.

Anonymous:1967:FM

[573] Anonymous. Front matter. Notes and Records of the Royal Society of London, 22(1/2):??, September 1967. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531181.

Elizabeth:1968:FOS

[574] Queen Elizabeth and Lord Florey. The formal opening of the Society’s new home at 6 Carlton House Terrace by Her Majesty The Queen, Patron, on 21 November 1967. Notes and Records of the Royal Society of London, 23(1):1–10, June 1968. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

530843.

Holford:1968:NDM

[575] Lord Holford. Notes on the decoration of the main features of 6 to 9 Carlton House Terrace. Notes and Records of the Royal Society of London, 23(1):11–13, June 1968. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

530844.

Wheare:1968:AD

[576] Kenneth Wheare. Anniversary Dinner 1967. Notes and Records of the Royal Society of London, 23(1):14–20, June 1968. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/530845.

Hartley:1968:AMS

[577] Harold Hartley. Address at the memorial service for Sir Cyril Hinshelwood, O.M., F.R.S. at Holy Trinity Church, Brompton Road, London on 20 November 1967. Notes and Records of the Royal Society of London, 23(1):21–22, June 1968. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/530846.

Hartley:1968:ESC

[578] Sir Harold Hartley, F.R.S. The exhibition of Sir Cyril Hinshelwood’s paintings at Goldsmiths’ Hall on 20 March, 1968. Notes and Records of the Royal Society of London, 23(1):23–28, June 1968. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/530847.

Adrian:1968:AMS

[579] Lord Adrian. Address at the memorial service for Lord Florey, of Adelaide and Marston, O.M., F.R.S., at Westminster Abbey on 28 March 1968. Notes and Records of the Royal Society of London, 23(1):29–30, June 1968. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/530848.

Spence:1968:ASM

[580] Robert Spence. Address at the service of memorial and thanksgiving for Sir John Cockcroft, O.M., K.C.B., F.R.S. at Westminster Abbey on 17 October 1967. Notes and Records of the Royal Society of London, 23(1):31–32, June 1968. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/530849.

Hall:1968:FNH

[581] A. Rupert Hall and Marie Boas Hall. Further notes on Henry Oldenburg. Notes and Records of the Royal Society of London, 23(1):33–42 + 2, June 1968. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/530850.

Holland:1968:ECP

[582] J. D. Holland. An eighteenth-century pioneer Richard Price, D.D., F.R.S. (1723–1791). Notes and Records of the Royal Society of London, 23(1): 43–64, June 1968. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/530851.

Eales:1968:SRS

[583] Nellie B. Eales. A satire on the Royal Society, dated 1743, attributed to Henry Fielding. Notes and Records of the Royal Society of London, 23(1):65–67, June 1968. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/530852.

deBeer:1968:DLS

[584] G. de Beer. The Darwin letters at Shrewsbury School. Notes and Records of the Royal Society of London, 23(1):68–85, June 1968. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/530853.

Eccles:1968:THU

[585] Sir John Eccles, F.R.S. Two hitherto unrecognized publications by Sir Charles Sherrington, O.M., F.R.S. Notes and Records of the Royal Society of London, 23(1):86–100, June 1968. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530854.

Hartley:1968:PDM

[586] Sir Harold Hartley, F.R.S. Professor Douglas McKie, F.R.S.E. Notes and Records of the Royal Society of London, 23(1):101–103, June 1968.

CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/530855.

Anonymous:1968:SNa

[587] Anonymous. The Society’s notes. Notes and Records of the Royal Society of London, 23(1):104–108, June 1968. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530856.

Anonymous:1968:BRBa

[588] Anonymous. Bibliography of recent books and articles dealing with the history of the Royal Society. Notes and Records of the Royal Society of London, 23(1):109–118, June 1968. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

530857.

Anonymous:1968:BMa

[589] Anonymous. Back matter. Notes and Records of the Royal Society of London, 23(1):??, June 1968. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/530858.

Anonymous:1968:C

[590] Anonymous. Conversaziones, 1968. Notes and Records of the Royal Society of London, 23(2):119–126, December 1968. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530980.

Todd:1968:AMS

[591] Lord Todd. Address at the memorial service for Lord Fleck, of Saltcoats, K.B.E., Treas. R.S., at St Columba’s Church of Scotland, London, on 23 September 1968. Notes and Records of the Royal Society of London, 23(2):127–128, December 1968. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

530981.

Rattansi:1968:IOR

[592] P. M. Rattansi. The intellectual origins of the Royal Society. Notes and Records of the Royal Society of London, 23(2):129–143, December 1968. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/530982.

Hill:1968:IOR

[593] Christopher Hill. The intellectual origins of the Royal Society-London or Oxford? Notes and Records of the Royal Society of London, 23(2):

144–156, December 1968. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/530983.

Hall:1968:IOR

[594] A. Rupert Hall and Marie Boas Hall. The intellectual origins of the Royal Society — London and Oxford. Notes and Records of the Royal Society of London, 23(2):157–168, December 1968. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530984.

Lohne:1968:EC

[595] J. A. Lohne. Experimentum crucis. Notes and Records of the Royal Society of London, 23(2):169–199, December 1968. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530985.

Hawes:1968:NRA

[596] Joan L. Hawes. Newton’s revival of the aether hypothesis and the explanation of gravitational attraction. Notes and Records of the Royal Society of London, 23(2):200–212, December 1968. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530986.

VanHelden:1968:CWC

[597] Albert Van Helden. Christopher Wren’s De corpore Saturni. Notes and Records of the Royal Society of London, 23(2):213–229, December 1968. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/530987.

Layton:1968:LWF

[598] David Layton. Lord Wrottesley, F.R.S., pioneer statesman of science. Notes and Records of the Royal Society of London, 23(2):230–246, December 1968. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/530988.

Andrade:1968:SRE

[599] E. N. da C. Andrade, F.R.S. Some reminiscences of Ernest Marsden’s days with Rutherford at Manchester. Notes and Records of the Royal Society of London, 23(2):247–250, December 1968. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

URL https://royalsocietypublishing.org/journal/rsnrcontent/
23/2/247;   http://www.jstor.org/stable/530989. Prepared for

Sir Ernest Marsden’s 80th Birthday booklet, to be published privately in Wellington, New Zealand, in February 1969.

Anonymous:1968:SNb

[600] Anonymous. The Society’s notes. Notes and Records of the Royal Society of London, 23(2):251–252, December 1968. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530990.

Anonymous:1968:BRBb

[601] Anonymous. Bibliography of recent books and articles dealing with the history of the Royal Society. Notes and Records of the Royal Society of London, 23(2):253–259, December 1968. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530991.

Anonymous:1968:INR

[602] Anonymous. Index to Notes and Records, volume 23, 1968. Notes and Records of the Royal Society of London, 23(2):260–262, December 1968. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/530978.

Anonymous:1968:BMb

[603] Anonymous. Back matter. Notes and Records of the Royal Society of London, 23(2):??, December 1968. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

530992.

Blackett:1969:OCG

[604] P. M. S. Blackett. Opening of the Cook Gallery at the National Maritime Museum, Greenwich, on 17 July 1968. Notes and Records of the Royal Society of London, 24(1):1–4, June 1969. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530735.

Anonymous:1969:CCB

[605] Anonymous. Captain Cook bicentenary commemorative presentations. Notes and Records of the Royal Society of London, 24(1):5–6, June 1969. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/530736.

Kaye:1969:CJC

[606] I. Kaye. Captain James Cook and The Royal Society. Notes and Records of the Royal Society of London, 24(1):7–18, June 1969. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/530737.

Woolley:1969:CCT

[607] Sir Richard Woolley, F.R.S. Captain Cook and the transit of Venus of 1769. Notes and Records of the Royal Society of London, 24(1):19– 32, June 1969. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/530738.

Deacon:1969:CCN

[608] G. E. R. Deacon, F.R.S. and Margaret Deacon. Captain Cook as a navigator. Notes and Records of the Royal Society of London, 24(1):33– 42, June 1969. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/530739.

Kodicek:1969:CCS

[609] Egon H. Kodicek and Frank G. Young, F.R.S. Captain Cook and scurvy. Notes and Records of the Royal Society of London, 24(1):43–63, June 1969. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/530740.

Stearn:1969:RSA

[610] William T. Stearn. A Royal Society appointment with Venus in 1769: the voyage of Cook and Banks in the Endeavour in 1768–1771 and its botanical results. Notes and Records of the Royal Society of London, 24(1):64–90, June 1969. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/530741.

Turner:1969:JSF

[611] G. L’E. Turner. James Short, F.R.S., and his contribution to the construction of reflecting telescopes. Notes and Records of the Royal Society of London, 24(1):91–108, June 1969. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

530742.

Bryden:1969:NFP

[612] D. J. Bryden. Note on a further portrait of James Short, F.R.S. Notes and Records of the Royal Society of London, 24(1):109–112, June 1969. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/530743.

Chaldecott:1969:CMF

[613] J. A. Chaldecott. Cromwell Mortimer, F.R.S. (c. 1698–1752) and the invention of the metalline thermometer for measuring high temperatures. Notes and Records of the Royal Society of London, 24(1):113–135, June 1969. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/530744.

Short:1969:AD

[614] Edward Short. Anniversary Dinner 1968. Notes and Records of the Royal Society of London, 24(1):136–144, June 1969. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530745.

Adrian:1969:AMS

[615] Lord Adrian. Address at the memorial service for Sir Henry Dale, O.M., G.B.E., F.R.S. at Westminster Abbey on 11 October 1968. Notes and Records of the Royal Society of London, 24(1):145, June 1969. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/530746.

Martin:1969:RSC

[616] D. C. Martin and Harold Hartley. The Royal Society Club Dinner on 14 November 1968 in celebration of the 90th birthday of Sir Harold Hartley, F.R.S. Notes and Records of the Royal Society of London, 24(1):146– 155, June 1969. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/530747.

Hammill:1969:SPB

[617] Gillian Hammill. The Society’s portraits and busts. Notes and Records of the Royal Society of London, 24(1):156–168, June 1969. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/530748.

Anonymous:1969:SN

[618] Anonymous. The Society’s notes. Notes and Records of the Royal Society of London, 24(1):169–174, June 1969. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530749.

Anonymous:1969:BRB London, 24(1):175–180, June 1969. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

530750.

Anonymous:1969:BM

[620] Anonymous. Back matter. Notes and Records of the Royal Society of London, 24(1):??, June 1969. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/530751.

Anonymous:1970:Ca

[621] Anonymous. Conversaziones, 1969. Notes and Records of the Royal Society of London, 24(2):181–186, April 1970. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531286.

Anonymous:1970:RHM

[622] Anonymous. Reception in honour of the memory of Captain James Cook, R.N., F.R.S. Notes and Records of the Royal Society of London, 24(2): 187–188, April 1970. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/531287.

Fleming:1970:JCB

[623] C. A. Fleming, F.R.S. James Cook bicentenary celebrations in New Zealand. Notes and Records of the Royal Society of London, 24(2):189– 193, April 1970. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531288.

Jones:1970:PSJ

[624] R. V. Jones, F.R.S. The ‘plain story’ of James Watt. The Wilkins Lecture, 1969. Notes and Records of the Royal Society of London, 24(2): 194–220, April 1970. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/531289.

Robinson:1970:JWE

[625] Eric Robinson. James Watt, engineer and man of science. Notes and Records of the Royal Society of London, 24(2):221–232, April 1970. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531290.

Fox:1970:WEP

[626] Robert Fox. Watt’s expansive principle in the work of Sadi Carnot and Nicolas Clément. Notes and Records of the Royal Society of London, 24 (2):233–253, April 1970. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531291.

Hills:1970:SRA

[627] Richard L. Hills. Sir Richard Arkwright and his patent granted in 1769. Notes and Records of the Royal Society of London, 24(2):254–260, April 1970. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531292.

Bryden:1970:JOC

[628] D. J. Bryden. The Jamaican observatories of Colin Campbell, F.R.S. and Alexander Macfarlane, F.R.S. Notes and Records of the Royal Society of London, 24(2):261–272, April 1970. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531293.

Copeman:1970:WPM

[629] W. S. C. Copeman. William Prout, M.D., F.R.S., physician and chemist (1785–1850). Notes and Records of the Royal Society of London, 24(2): 273–280, April 1970. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/531294.

Brock:1970:WPB

[630] W. H. Brock. William Prout and barometry. Notes and Records of the Royal Society of London, 24(2):281–294, April 1970. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531295.

Larder:1970:TTA

[631] David F. Larder. Thomas Thomson’s activities in Edinburgh, 1791– 1811. Notes and Records of the Royal Society of London, 24(2):295– 304, April 1970. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531296.

MacLeod:1970:XCS

[632] Roy M. MacLeod. The X-Club. A social network of science in late-Victorian England. Notes and Records of the Royal Society of London, 24(2):305–322, April 1970. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531297.

Anonymous:1970:SNa

[633] Anonymous. The Society’s notes. Notes and Records of the Royal Society of London, 24(2):323, April 1970. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531298.

Anonymous:1970:BRBa

[634] Anonymous. Bibliography of recent books and articles dealing with the history of the Royal Society. Notes and Records of the Royal Society of London, 24(2):324–329, April 1970. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531299.

Anonymous:1970:INRa

[635] Anonymous. Index to Notes and Records, volume 24. 1969. Notes and Records of the Royal Society of London, 24(2):330–332, April 1970. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531284.

Anonymous:1970:BM

[636] Anonymous. Back matter. Notes and Records of the Royal Society of London, 24(2):??, April 1970. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531300.

Casimir:1970:ADS

[637] H. B. G. Casimir. Anniversary Dinner 1969. Speech by Professor H. B. G. Casimir at the Anniversary Dinner, The Dorchester, 1 December 1969. Notes and Records of the Royal Society of London, 25(1):1–7, June 1970. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/530860.

Maddox:1970:NCD

[638] John Maddox and Harold Macmillan. The Nature centenary dinner. Notes and Records of the Royal Society of London, 25(1):9–15, June 1970. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/530861.

Scriba:1970:AJW

[639] Christoph J. Scriba. The autobiography of John Wallis, F.R.S. Notes and Records of the Royal Society of London, 25(1):17–46, June 1970. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/530862.

Biswas:1970:EHF

[640] Asit K. Biswas. Edmond Halley, F.R.S., hydrologist extraordinary. Notes and Records of the Royal Society of London, 25(1):47–57, June 1970.

CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/530863.

Egerton:1970:RBR

[641] Frank N. Egerton. Richard Bradley’s relationship with Sir Hans Sloane. Notes and Records of the Royal Society of London, 25(1):59–77, June 1970. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/530864.

Setchell:1970:FJS

[642] J. R. M. Setchell. The friendship of John Smeaton, F.R.S., with Henry Hindley, instrument and clockmaker of York and the development of equatorial mounting telescopes. Notes and Records of the Royal Society of London, 25(1):79–86, June 1970. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

530865.

Cantor:1970:TYL

[643] G. N. Cantor. Thomas Young’s lectures at the Royal Institution. Notes and Records of the Royal Society of London, 25(1):87–112, June 1970. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/530866.

Levere:1970:FIM

[644] Trevor H. Levere. Friendship and influence — Martinus Van Marum, F.R.S. Notes and Records of the Royal Society of London, 25(1):113– 120, June 1970. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/530867.

Anonymous:1970:SNb

[645] Anonymous. The Society’s notes. Notes and Records of the Royal Society of London, 25(1):121–125, June 1970. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530868.

Anonymous:1970:BRBb

[646] Anonymous. Bibliography of recent books and articles dealing with the history of the Royal Society. Notes and Records of the Royal Society of London, 25(1):127–133, June 1970. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

530869.

Anonymous:1970:FM

[647] Anonymous. Front matter. Notes and Records of the Royal Society of London, 25(1):??, June 1970. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

530859.

Anonymous:1970:Cb

[648] Anonymous. Conversaziones, 1970. Notes and Records of the Royal Society of London, 25(2):135–142, December 1970. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530871.

Lawrence:1970:DGI

[649] P. D. Lawrence and A. G. Molland. David Gregory’s inaugural lecture at Oxford. Notes and Records of the Royal Society of London, 25(2): 143–178, December 1970. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/530872.

Underwood:1970:EHF

[650] T. L. Underwood. Edward Haistwell, F.R.S. Notes and Records of the Royal Society of London, 25(2):179–187, December 1970. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/530873.

Setchell:1970:FIT

[651] J. R. M. Setchell. Further information on the telescopes of Hindley of York. Notes and Records of the Royal Society of London, 25(2):189–192, December 1970. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/530874.

Larder:1970:UCE

[652] David F. Larder. An unpublished chemical essay of James Watt. Notes and Records of the Royal Society of London, 25(2):193–210, December 1970. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/530875.

Lloyd:1970:BJM

[653] J. T. Lloyd. Background to the Joule–Mayer controversy. Notes and Records of the Royal Society of London, 25(2):211–225, December 1970. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/530876.

Bowen:1970:BTL

[654] Edmund J. Bowen, F.R.S. The Balliol–Trinity Laboratories, Oxford, 1853–1940. Notes and Records of the Royal Society of London, 25(2): 227–236, December 1970. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/530877.

Thomson:1970:UEH

[655] Sir George Thomson, F.R.S. An unfortunate experiment, Hertz and the nature of cathode rays. Notes and Records of the Royal Society of London, 25(2):237–242, December 1970. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

530878.

Anonymous:1970:SNc

[656] Anonymous. The Society’s notes. Notes and Records of the Royal Society of London, 25(2):243, December 1970. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530879.

Anonymous:1970:BRBc

[657] Anonymous. Bibliography of recent books and articles dealing with the history of the Royal Society. Notes and Records of the Royal Society of London, 25(2):245–250, December 1970. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530880.

Anonymous:1970:INRb

[658] Anonymous. Index to Notes and Records, volume 25, 1970. Notes and Records of the Royal Society of London, 25(2):251–252, December 1970. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/530870.

J:1971:SHH

[659] R. V. J. and W. D. M. P. Sir Harold Hartley, F.R.S. an appreciation on his retirement from the editorship of Notes and Records. Notes and Records of the Royal Society of London, 26(1):1–2, June 1971. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531048.

Anonymous:1971:FMa

[660] Anonymous. Front matter. Notes and Records of the Royal Society of London, 26(1):3, June 1971. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531047.

Boyle:1971:ADS

[661] The Lord Boyle. Anniversary Dinner 1970, speech by The Rt Hon. The Lord Boyle at the Anniversary Dinner, The Dorchester, 30 November 1970. Notes and Records of the Royal Society of London, 26(1):5–14, June 1971. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531049.

Tunbridge:1971:JAL

[662] Paul A. Tunbridge. Jean André De Luc, F.R.S. (1727–1817). Notes and Records of the Royal Society of London, 26(1):15–33, June 1971. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531050.

Smeaton:1971:SCJ

[663] W. A. Smeaton. Some comments on James Watt’s published account of his work on steam and steam engines. Notes and Records of the Royal Society of London, 26(1):35–42, June 1971. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531051.

Morrell:1971:PRP

[664] J. B. Morrell. Professors Robison and Playfair, and the Theophobia Gallica: Natural philosophy, religion and politics in Edinburgh, 1789–1815. Notes and Records of the Royal Society of London, 26(1):43–63, June 1971. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531052.

Wilkinson:1971:WBF

[665] Lise Wilkinson. William Brockedon, F.R.S. (1787–1854). Notes and Records of the Royal Society of London, 26(1):65–72, June 1971. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531053.

Allibone:1971:TCC

[666] T. E. Allibone, F.R.S. The Thursday’s Club called the Club of the Royal Philosophers, and its relation to the Royal Society Club. Notes and Records of the Royal Society of London, 26(1):73–80, June 1971. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531054.

Macleod:1971:MMR

[667] Roy M. Macleod. Of medals and men: a reward system in Victorian science, 1826–1914. Notes and Records of the Royal Society of London, 26(1):81–105, June 1971. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531055.

Hartley:1971:VGF

[668] Sir Harold Hartley, F.R.S. Viscount Grey of Fallodon, K.G., F.R.S. Notes and Records of the Royal Society of London, 26(1):107–110, June 1971. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531056.

Anonymous:1971:VSD

[669] Anonymous. Visit of a scientific delegation from Japan. Notes and Records of the Royal Society of London, 26(1):111, June 1971. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531057.

Anonymous:1971:SNa

[670] Anonymous. The Society’s notes. Notes and Records of the Royal Society of London, 26(1):113–118, June 1971. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531058.

Anonymous:1971:BRBa

[671] Anonymous. Bibliography of recent books and articles dealing with the history of the Royal Society. Notes and Records of the Royal Society of London, 26(1):119–123, June 1971. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531059.

Anonymous:1971:BMa

[672] Anonymous. Back matter. Notes and Records of the Royal Society of London, 26(1):??, June 1971. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531060.

Anonymous:1971:C

[673] Anonymous. Conversaziones, 1971. Notes and Records of the Royal Society of London, 26(2):125–133, December 1971. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531158.

Skempton:1971:PSR

[674] A. W. Skempton, F.R.S. The publication of Smeaton’s reports. Notes and Records of the Royal Society of London, 26(2):135–155, December 1971.

CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531159.

Cilento:1971:SJB

[675] Sir Raphael Cilento. Sir Joseph Banks, F.R.S. and the naming of the kangaroo. Notes and Records of the Royal Society of London, 26(2): 157–161, December 1971. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531160.

Hackmann:1971:DTG

[676] W. D. Hackmann. The design of the triboelectric generators of Martinus van Marum, F.R.S.: a case history of the interaction between England and Holland in the field of instrument design in the eighteenth century. Notes and Records of the Royal Society of London, 26(2):163–181, December 1971. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531161.

Wilkinson:1971:TDF

[677] Lise Wilkinson. Three drawings of Fellows by William Brockedon, F.R.S. Notes and Records of the Royal Society of London, 26(2):183–187, December 1971. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531162.

Rowlinson:1971:TG

[678] J. S. Rowlinson, F.R.S. The theory of glaciers. Notes and Records of the Royal Society of London, 26(2):189–204, December 1971. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531163.

Mayr:1971:VPS

[679] Otto Mayr. Victorian physicists and speed regulation: An encounter between science and technology. Notes and Records of the Royal Society of London, 26(2):205–228, December 1971. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531164.

Tunbridge:1971:LWT

[680] Paul A. Tunbridge. A letter by William Thomson, F.R.S., on the ‘Thomson effect’. Notes and Records of the Royal Society of London, 26(2): 229–232, December 1971. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531165.

Charlton:1971:MJC

[681] T. M. Charlton. Maxwell, Jenkin and Cotterill and the theory of statically-indeterminate structures. Notes and Records of the Royal Society of London, 26(2):233–246, December 1971. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531166.

Anonymous:1971:SNb

[682] Anonymous. The Society’s notes. Notes and Records of the Royal Society of London, 26(2):247, December 1971. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531167.

Anonymous:1971:BRBb

[683] Anonymous. Bibliography of recent books and articles dealing with the history of the Royal Society. Notes and Records of the Royal Society of London, 26(2):249–254, December 1971. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531168.

Anonymous:1971:INR

[684] Anonymous. Index to Notes and Records, volume 26, 1971. Notes and Records of the Royal Society of London, 26(2):255–258, December 1971. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531156.

Anonymous:1971:BMb

[685] Anonymous. Back matter. Notes and Records of the Royal Society of London, 26(2):??, December 1971. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531169.

Anonymous:1971:FMb

[686] Anonymous. Front matter. Notes and Records of the Royal Society of London, 26(2):??, December 1971. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531157.

Anonymous:1972:E

[687] Anonymous. Editorial. Notes and Records of the Royal Society of London, 27(1):1–3, August 1972. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/530909.

Anonymous:1972:RCC

[688] Anonymous. Rutherford Centenary celebrations. Notes and Records of the Royal Society of London, 27(1):5, August 1972. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

URL https://royalsocietypublishing.org/journal/rsnrcontent/
27/1/5;   http://www.jstor.org/stable/530910.

Oliphant:1972:SPR

[689] Sir Mark Oliphant, F.R.S. Some personal recollections of Rutherford, the man. Notes and Records of the Royal Society of London, 27(1):7–23, August 1972. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL https://royalsocietypublishing.org/journal/
rsnrcontent/27/1/7;   http://www.jstor.org/stable/530911.

Massey:1972:NPT

[690] Sir Harrie Massey, Sec.R.S. Nuclear physics today and in Rutherford’s day. Notes and Records of the Royal Society of London, 27(1):25–44, August 1972. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL https://royalsocietypublishing.org/journal/
rsnrcontent/27/1/25;   http://www.jstor.org/stable/530912.

Feather:1972:RFN

[691] N. Feather, F.R.S. Rutherford — Faraday — Newton. Notes and Records of the Royal Society of London, 27(1):45–55, August 1972. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

URL https://royalsocietypublishing.org/journal/rsnrcontent/
27/1/45;   http://www.jstor.org/stable/530913.

Blackett:1972:R

[692] Lord Blackett, F.R.S. Rutherford. Notes and Records of the Royal Society of London, 27(1):57–59, August 1972. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:

//royalsocietypublishing.org/journal/rsnrcontent/27/1/57;
http://www.jstor.org/stable/530914.

Lewis:1972:SRR

[693] W. Bennett Lewis, F.R.S. Some recollections and reflections on Rutherford. Notes and Records of the Royal Society of London, 27(1):61–63, August 1972. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL https://royalsocietypublishing.org/journal/
rsnrcontent/27/1/61;   http://www.jstor.org/stable/530915.

Mott:1972:RT

[694] Sir Nevill Francis Mott, F.R.S. Rutherford and theory. Notes and Records of the Royal Society of London, 27(1):65–66, August 1972. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

URL https://royalsocietypublishing.org/journal/rsnrcontent/
27/1/65;   http://www.jstor.org/stable/530916.

OShea:1972:ERH

[695] P. P. O’Shea. Ernest Rutherford. His honours and distinctions. Notes and Records of the Royal Society of London, 27(1):67–74, August 1972. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/530917.

Adams:1972:FGN

[696] J. B. Adams, F.R.S. Four generations of nuclear physicists. Notes and Records of the Royal Society of London, 27(1):75–94, August 1972. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/530918.

Denning:1972:AD

[697] The Lord Denning. Anniversary Dinner 1971. Notes and Records of the Royal Society of London, 27(1):95–102, August 1972. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/530919.

Hetherington:1972:HAC

[698] N. S. Hetherington. The Hevelius–Auzout controversy. Notes and Records of the Royal Society of London, 27(1):103–106, August 1972. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/530920.

Bennett:1972:WLB

[699] J. A. Bennett. Wren’s last building? Notes and Records of the Royal Society of London, 27(1):107–118, August 1972. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530921.

Slater:1972:LHF

[700] A. W. Slater. Luke Howard, F.R.S. (1772–1864) and his relations with Goethe. Notes and Records of the Royal Society of London, 27(1):119– 140, August 1972. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/530922.

Levene:1972:BFF

[701] John R. Levene. Benjamin Franklin, F.R.S., Sir Joshua Reynolds, F.R.S., P.R.A., Benjamin West, P.R.A. and the invention of bifocals. Notes and Records of the Royal Society of London, 27(1):141–163, August 1972. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/530923.

Salisbury:1972:RMK

[702] Edward Salisbury. The retirement of Mr I. Kaye, Librarian to the Society. Notes and Records of the Royal Society of London, 27(1):165, August 1972. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/530924.

Anonymous:1972:SN

[703] Anonymous. The Society’s notes. Notes and Records of the Royal Society of London, 27(1):167–171, August 1972. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530925.

Anonymous:1972:BRB

[704] Anonymous. Bibliography of recent books and articles dealing with the history of the Royal Society. Notes and Records of the Royal Society of London, 27(1):173–176, August 1972. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530926.

Anonymous:1972:BM

[705] Anonymous. Back matter. Notes and Records of the Royal Society of London, 27(1):??, August 1972. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

530927.

Anonymous:1972:FM

[706] Anonymous. Front matter. Notes and Records of the Royal Society of London, 27(1):??, August 1972. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

530908.

Hodgkin:1973:AMS

[707] Alan Hodgkin. Address at the memorial service for Sir Frederick Bawden, Treas. R.S. at the Church of St James, Piccadilly on 16 March 1972. Notes and Records of the Royal Society of London, 27(2):177–180, February 1973. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/530995.

Jones:1973:AMS

[708] R. V. Jones. Address at the memorial service for Sir Harold Hartley, F.R.S. at Westminster Abbey on 17 October 1972. Notes and Records of the Royal Society of London, 27(2):181–184, February 1973. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/530996.

Anonymous:1973:C

[709] Anonymous. Conversaziones, 1972. Notes and Records of the Royal Society of London, 27(2):185–192, February 1973. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/530997.

Frank:1973:JAF

[710] Robert G. Frank, Jr. John Aubrey, F.R.S., John Lydall, and science at Commonwealth Oxford. Notes and Records of the Royal Society of London, 27(2):193–217, February 1973. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

530998.

Bolam:1973:BWN

[711] Jeanne Bolam. The botanical works of Nehemiah Grew, F.R.S. (1641– 1712). Notes and Records of the Royal Society of London, 27(2):219–231, February 1973. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/530999.

Skempton:1973:JET

[712] A. W. Skempton, F.R.S. and Joyce Brown. John and Edward Troughton, mathematical instrument makers. Notes and Records of the Royal Society of London, 27(2):233–249, February 1973. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/ stable/531000. See addendum [713].

Skempton:1973:A

[713] A. W. Skempton, F.R.S. and Joyce Brown. Addendum. Notes and Records of the Royal Society of London, 27(2):249–262, February 1973. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531001. See [712].

Tunbridge:1973:FGF

[714] Paul A. Tunbridge. Faraday’s Genevese friends. Notes and Records of the Royal Society of London, 27(2):263–298, February 1973. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531002.

Farrar:1973:AUF

[715] W. V. Farrar. Andrewy Ure, F.R.S. and the philosophy of manufactures. Notes and Records of the Royal Society of London, 27(2):299–324, February 1973. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531003.

Martin:1973:RMJ

[716] David Martin. The retirement of Mr J. C. Graddon, Assistant Editor to the Society. Notes and Records of the Royal Society of London, 27(2): 325–326, February 1973. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531004.

Anonymous:1973:SNa

[717] Anonymous. The Society’s notes. Notes and Records of the Royal Society of London, 27(2):327, February 1973. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531005.

Anonymous:1973:BRBa

[718] Anonymous. Bibliography of recent books and articles dealing with the history of the Royal Society. Notes and Records of the Royal Society of London, 27(2):329–334, February 1973. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531006.

Anonymous:1973:INR

[719] Anonymous. Index to Notes and Records, volume 27, 1972–3. Notes and Records of the Royal Society of London, 27(2):335–337, February 1973. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/530993.

Anonymous:1973:BMa

[720] Anonymous. Back matter. Notes and Records of the Royal Society of London, 27(2):??, February 1973. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531007.

Anonymous:1973:FMa

[721] Anonymous. Front matter. Notes and Records of the Royal Society of London, 27(2):??, February 1973. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

530994.

Thatcher:1973:ADS

[722] Margaret Thatcher. Anniversary Dinner 1972: Speech by the Rt Hon. Margaret Thatcher, M.P., Secretary of State for Education and Science, at the Anniversary Dinner, The Dorchester, 30 November 1972. Notes and Records of the Royal Society of London, 28(1):1–9, June 1973. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531108.

Anonymous:1973:CQB

[723] Anonymous. Celebration of the quincentenary of the birth of Nicolaus Copernicus. Notes and Records of the Royal Society of London, 28(1):11– 13, June 1973. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531109.

Millburn:1973:BMR

[724] John R. Millburn. Benjamin Martin and the Royal Society. Notes and Records of the Royal Society of London, 28(1):15–23, June 1973. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531110.

Gold:1973:STC

[725] V. Gold, F.R.S. Samuel Taylor Coleridge and the appointment of J. F. Daniell, F.R.S., as Professor of Chemistry at King’s College London. Notes and Records of the Royal Society of London, 28(1):25–29, June 1973. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531111.

Coley:1973:HBJ

[726] N. G. Coley. Henry Bence-Jones, M.D., F.R.S. (1813–1873). Notes and Records of the Royal Society of London, 28(1):31–56, June 1973. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531112.

Jones:1973:JCM

[727] R. V. Jones, F.R.S. James Clerk Maxwell at Aberdeen, 1856–1860. Notes and Records of the Royal Society of London, 28(1):57–81, June 1973.

CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531113. See erratum [749].

Krebs:1973:TLW

[728] Sir Hans Krebs, F.R.S. Two letters by Wilhelm Conrad Röntgen. Notes and Records of the Royal Society of London, 28(1):83–92, June 1973. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531114.

Boyd:1973:SSC

[729] Sir John Boyd, F.R.S. Sleeping sickness. The Castellani–Bruce controversy. Notes and Records of the Royal Society of London, 28(1):93–110, June 1973. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531115.

Paton:1973:HJS

[730] W. D. M. Paton, F.R.S. and C. G. Phillips, F.R.S. E. H. J. Schuster (1897–1969). Notes and Records of the Royal Society of London, 28(1): 111–117, June 1973. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/531116.

Forbes:1973:LRJ

[731] Eric G. Forbes. The library of the Rev. John Flamsteed, F.R.S., first Astronomer Royal. Notes and Records of the Royal Society of London, 28(1):119–143, June 1973. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531117.

Anonymous:1973:SNb

[732] Anonymous. The Society’s notes. Notes and Records of the Royal Society of London, 28(1):145–149, June 1973. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531118.

Anonymous:1973:BRBb

[733] Anonymous. Bibliography of recent books and articles dealing with the history of the Royal Society. Notes and Records of the Royal Society of London, 28(1):151–154, June 1973. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:1973:BRBc London, 28(1):151–155, June 1973. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531119.

Anonymous:1973:BMb

[735] Anonymous. Back matter. Notes and Records of the Royal Society of London, 28(1):??, June 1973. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531120.

Anonymous:1973:FMb

[736] Anonymous. Front matter. Notes and Records of the Royal Society of London, 28(1):??, June 1973. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531107.

Anonymous:1974:C

[737] Anonymous. Conversaziones, 1973. Notes and Records of the Royal Society of London, 28(2):157–166, April 1974. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531328.

Buchanan-Brown:1974:BPR

[738] J. Buchanan-Brown. The books presented to the Royal Society by John Aubrey, F.R.S. Notes and Records of the Royal Society of London, 28 (2):167–193, April 1974. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531329.

Allibone:1974:EHC

[739] T. E. Allibone, F.R.S. Edmond Halley and the Clubs of the Royal Society. Notes and Records of the Royal Society of London, 28(2):195–205, April 1974. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531330.

Tunbridge:1974:FPL

[740] Paul A. Tunbridge. Franklin’s pointed lightning conductor. Notes and Records of the Royal Society of London, 28(2):207–219, April 1974. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531331.

Lysaght:1974:JBS

[741] Averil M. Lysaght. Joseph Banks at Skara Brae and Stennis, Orkney, 1772. Notes and Records of the Royal Society of London, 28(2):221– 234, April 1974. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531332.

Scott:1974:EJF

[742] E. L. Scott. Edward Jenner, F.R.S. and the cuckoo. Notes and Records of the Royal Society of London, 28(2):235–240, April 1974. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531333.

Gibson:1974:RAS

[743] A. Gibson and W. V. Farrar. Robert Angus Smith, F.R.S. and ‘sanitary science’. Notes and Records of the Royal Society of London, 28(2):241– 262, April 1974. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531334.

Scala:1974:IPN

[744] Gail Ewald Scala. An index of proper names in Thomas Birch, the history of the Royal Society (London, 1756–1757). Notes and Records of the Royal Society of London, 28(2):263–329, April 1974. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531335.

Anonymous:1974:LCG

[745] Anonymous. Lt Cdr George R. Lush, M.B.E., R.N., 1 April 1919 to 14 July 1973. Notes and Records of the Royal Society of London, 28(2):331– 332, April 1974. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531336.

Anonymous:1974:SNa

[746] Anonymous. The Society’s notes. Notes and Records of the Royal Society of London, 28(2):333–334, April 1974. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531337.

Anonymous:1974:BRBa

[747] Anonymous. Bibliography of recent books and articles dealing with the history of the Royal Society. Notes and Records of the Royal Society of London, 28(2):335–339, April 1974. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531338.

Anonymous:1974:INR

[748] Anonymous. Index to Notes and Records, volume 28. 1973–4. Notes and Records of the Royal Society of London, 28(2):341–343, April 1974. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531326.

Anonymous:1974:EJC

[749] Anonymous. Errata: James Clerk Maxwell at Aberdeen, 1856–1860. Notes and Records of the Royal Society of London, 28(2):344, April 1974. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531339. See [727].

Anonymous:1974:BMa

[750] Anonymous. Back matter. Notes and Records of the Royal Society of London, 28(2):??, April 1974. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531340.

Anonymous:1974:FMa

[751] Anonymous. Front matter. Notes and Records of the Royal Society of London, 28(2):??, April 1974. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531327.

Bullock:1974:ADS

[752] Alan Bullock. Anniversary Dinner 1973: Speech by Sir Alan Bullock at the Anniversary Dinner, The Dorchester 30 November 1973. Notes and Records of the Royal Society of London, 29(1):1–9, October 1974. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531560.

Philip:1974:RPI

[753] H. R. H. The Prince Philip. Research and prediction: The inaugural Hartley lecture delivered at the Royal Society, 21 May 1974. Notes and Records of the Royal Society of London, 29(1):11–27, October 1974. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531561.

Hall:1974:NHE

[754] A. Rupert Hall. Newton and his editors: The Wilkins Lecture, 1973. Notes and Records of the Royal Society of London, 29(1):29–52 + 4, October 1974. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531562. The Wilkins

Lecture, 1973.

Turner:1974:HBF

[755] G. L’E Turner. Henry Baker, F.R.S.: Founder of the Bakerian Lecture. Notes and Records of the Royal Society of London, 29(1):53–79, October 1974. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531563.

Home:1974:SME

[756] R. W. Home. Some manuscripts on electrical and other subjects attributed to Thomas Bayes, F.R.S. Notes and Records of the Royal Society of London, 29(1):81–90, October 1974. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531564.

Lysaght:1974:SEL

[757] Averil M. Lysaght. Some early letters from Joseph Banks to William Phelp Perrin. Notes and Records of the Royal Society of London, 29 (1):91–99, October 1974. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531565.

Charlton:1974:PBH

[758] T. M. Charlton. Professor Bertram Hopkinson, C.M.G., M.A., B.Sc., F.R.S. (1874–1918). Notes and Records of the Royal Society of London, 29 (1):101–109, October 1974. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531566.

Griffin:1974:ARS

[759] D. J. H. Griffin. The Aldabra Research Station. Notes and Records of the Royal Society of London, 29(1):111–119, October 1974. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531567.

Anonymous:1974:SNb

[760] Anonymous. The Society’s notes. Notes and Records of the Royal Society of London, 29(1):121–125, October 1974. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531568.

Anonymous:1974:BRBb

[761] Anonymous. Bibliography of recent books and articles dealing with the history of the Royal Society. Notes and Records of the Royal Society of London, 29(1):127–133, October 1974. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531569.

Anonymous:1974:BMb

[762] Anonymous. Back matter. Notes and Records of the Royal Society of London, 29(1):??, October 1974. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531570.

Anonymous:1974:FMb

[763] Anonymous. Front matter. Notes and Records of the Royal Society of London, 29(1):??, October 1974. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531559.

Hodgkin:1975:MML

[764] Alan Hodgkin, Harrie Massey, David Martin, G. P. S. Occhialini, Bernard Lovell, C. H. Waddington, C. C. Butler, S. K. Runcorn, and M. G. K. Menon. Memorial meeting for Lord Blackett, O.M., C.H., F.R.S. at the Royal Society on 31 October 1974. Notes and Records of the Royal Society of London, 29(2):135–162, March 1975. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531462.

Anonymous:1975:C

[765] Anonymous. Conversaziones 1974. Notes and Records of the Royal Society of London, 29(2):163–171, March 1975. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531463.

Hall:1975:RSR

[766] Marie Boas Hall. The Royal Society’s role in the diffusion of information in the seventeenth century. I. Notes and Records of the Royal Society of London, 29(2):173–192, March 1975. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/ stable/531464. See errata [786].

Philip:1975:SJA

[767] J. R. Philip, F.R.S. Samuel Johnson as antiscientist. Notes and Records of the Royal Society of London, 29(2):193–203, March 1975. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531465.

Cannon:1975:SWJ

[768] Garland Cannon. Sir William Jones, Sir Joseph Banks, and the Royal Society. Notes and Records of the Royal Society of London, 29(2):205– 230, March 1975. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531466.

Rudwick:1975:CLF

[769] Martin Rudwick. Charles Lyell, F.R.S. (1797–1875) and his London lectures on geology, 1832–33. Notes and Records of the Royal Society of London, 29(2):231–263, March 1975. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531467.

OHara:1975:GJS

[770] J. G. O’Hara. George Johnstone Stoney, F.R.S. and the concept of the electron. Notes and Records of the Royal Society of London, 29(2):265– 276, March 1975. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531468.

Anonymous:1975:SNa

[771] Anonymous. The Society’s notes. Notes and Records of the Royal Society of London, 29(2):277, March 1975. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531469.

Anonymous:1975:BRBa

[772] Anonymous. Bibliography of recent books and articles dealing with the history of the Royal Society. Notes and Records of the Royal Society of London, 29(2):279–284, March 1975. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531470.

Anonymous:1975:INR

[773] Anonymous. Index to Notes and Records, volume. Notes and Records of the Royal Society of London, 29(2):285–286, March 1975. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531460.

Anonymous:1975:BM

[774] Anonymous. Back matter. Notes and Records of the Royal Society of London, 29(2):??, March 1975. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531471.

Anonymous:1975:FM

[775] Anonymous. Front matter. Notes and Records of the Royal Society of London, 29(2):??, March 1975. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531461.

Todd:1975:AMS

[776] Lord Todd. Address at the memorial service for Sir Robert Robinson, O.M., F.R.S. at Westminster Abbey on 29 April 1975. Notes and Records of the Royal Society of London, 30(1):1–3, July 1975. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531735.

Brimelow:1975:ADS

[777] Thomas Brimelow. Anniversary Dinner 1974: Speech by Sir Thomas Brimelow at the Anniversary Dinner, The Dorchester, 30 November 1974. Notes and Records of the Royal Society of London, 30(1):5–14, July 1975. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531736.

Levene:1975:SGB

[778] John R. Levene. Sir G. B. Airy, F.R.S. and the symptomatology of migraine. Notes and Records of the Royal Society of London, 30(1):15– 23, July 1975. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531737.

Gunther:1975:DLS

[779] A. E. Gunther. The Darwin letters at Shrewsbury School. Notes and Records of the Royal Society of London, 30(1):25–43, July 1975. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531738.

Sherrington:1975:CSS

[780] C. E. R. Sherrington. Charles Scott Sherrington (1857–1952). Notes and Records of the Royal Society of London, 30(1):45–63, July 1975. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531739.

Hill:1975:JMA

[781] A. V. Hill, F.R.S. Jewels in my acquaintance with C. S. Sherrington, F.R.S. Notes and Records of the Royal Society of London, 30(1):65– 68, July 1975. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531740.

Eccles:1975:LCS

[782] Sir John Eccles, F.R.S. Letters from C. S. Sherrington, F.R.S., to Angelo Ruffini between 1896 and 1903. Notes and Records of the Royal Society of London, 30(1):69–88, July 1975. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531741.

Grattan-Guinness:1975:RSF

[783] I. Grattan-Guinness. The Royal Society’s financial support of the publication of Whitehead and Russell’s Principia mathematica. Notes and Records of the Royal Society of London, 30(1):89–104, July 1975. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531742.

Anonymous:1975:SNb

[784] Anonymous. The Society’s notes. Notes and Records of the Royal Society of London, 30(1):105–109, July 1975. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531743.

Anonymous:1975:BRBb

[785] Anonymous. Bibliography of recent books and articles dealing with the history of the Royal Society. Notes and Records of the Royal Society of London, 30(1):111–116, July 1975. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531744.

Anonymous:1975:ERS

[786] Anonymous. Errata: The Royal Society’s Role in the Diffusion of Information in the Seventeenth Century (1). Notes and Records of the Royal Society of London, 30(1):116, July 1975. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531745. See [766].

Anonymous:1975:VI

[787] Anonymous. Volume information. Notes and Records of the Royal Society of London, 30(1):??, July 1975. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531734.

Anonymous:1976:CR

[788] Anonymous. Conversaziones and reception, 1975. Notes and Records of the Royal Society of London, 30(2):117–126, January 1976. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531747.

Lovell:1976:AWA

[789] Bernard Lovell. Address at Westminster Abbey, on the occasion of the tercentenary of the Royal Greenwich Observatory. Notes and Records of the Royal Society of London, 30(2):127–132, January 1976. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531748.

McCrea:1976:ROS

[790] W. H. McCrea, F.R.S. The Royal Observatory and the study of gravitation. Notes and Records of the Royal Society of London, 30(2):133–140 + 2, January 1976. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531749. Address at

a reception on 17 July 1975 in celebration of the tercentenary of the Royal Greenwich Observatory.

Edmonds:1976:WBF

[791] J. M. Edmonds and J. A. Douglas. William Buckland, F.R.S. (1784– 1856) and an Oxford geological lecture, 1823. Notes and Records of the Royal Society of London, 30(2):141–167, January 1976. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531750.

Charlton:1976:CSB

[792] T. M. Charlton. Contributions to the science of bridge-building in the nineteenth century by Henry Moseley, Hon. Ll.D., F.R.S. and William Pole, D.Mus., F.R.S. Notes and Records of the Royal Society of London, 30(2):169–179, January 1976. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531751.

Jensen:1976:THH

[793] J. Vernon Jensen. Thomas Henry Huxley’s ‘baptism into oratory’. Notes and Records of the Royal Society of London, 30(2):181–207, January 1976. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531752.

Bartholomew:1976:ACM

[794] M. J. Bartholomew. The award of the Copley Medal to Charles Darwin. Notes and Records of the Royal Society of London, 30(2):209–218, January 1976. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531753.

Eccles:1976:ECT

[795] Sir John Eccles, F.R.S. From electrical to chemical transmission in the central nervous system: The closing address of the Sir Henry Dale Centennial Symposium Cambridge, 19 September 1975. Notes and Records of the Royal Society of London, 30(2):219–230, January 1976. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531754.

Paton:1976:SHD

[796] W. D. M. Paton, F.R.S. Sir Henry Dale (1875–1968). some letters and papers. Notes and Records of the Royal Society of London, 30(2):231–248, January 1976. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531755.

Chandrasekhar:1976:VTR

[797] S. Chandrasekhar, F.R.S. Verifying the theory of relativity. Notes and Records of the Royal Society of London, 30(2):249–260, January 1976. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531756.

Anonymous:1976:SNa

[798] Anonymous. The Society’s notes. Notes and Records of the Royal Society of London, 30(2):261, January 1976. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531757.

Anonymous:1976:BRBa

[799] Anonymous. Bibliography of recent books and articles dealing with the history of the Royal Society. Notes and Records of the Royal Society of London, 30(2):263–266, January 1976. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531758.

Anonymous:1976:INR

[800] Anonymous. Index to Notes and Records, volume 30, 1975–6. Notes and Records of the Royal Society of London, 30(2):267–269, January 1976. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531746.

Lst:1976:AD

[801] R. L˘ st. Anniversary Dinner 1975. Notes and Records of the Royal Society of London, 31(1):1–8, July 1976. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531551.

Hunter:1976:SBC

[802] Michael Hunter. The social basis and changing fortunes of an early scientific institution: an analysis of the membership of the Royal Society, 1660–1685. Notes and Records of the Royal Society of London, 31(1):9– 114, July 1976. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531552.

Singer:1976:RHM

[803] B. R. Singer. Robert Hooke on memory, association and time perception (i). Notes and Records of the Royal Society of London, 31(1):115–131, July 1976. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531553.

Underwood:1976:QRS

[804] T. L. Underwood. Quakers and the Royal Society of London in the seventeenth century. Notes and Records of the Royal Society of London, 31(1):133–150, July 1976. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531554.

Norton:1976:NBR

[805] Bernard Norton and E. S. Pearson, F.R.S. A note on the background to, and refereeing of, R. A. Fisher’s 1918 paper: “On the correlation between relatives on the supposition of Mendelian inheritance” (Trans. Roy. Soc. Edinburgh 52 (1918/19), 399–434). Notes and Records of the Royal Society of London, 31(1):151–162, July 1976. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531555.

Anonymous:1976:SNb

[806] Anonymous. The Society’s notes. Notes and Records of the Royal Society of London, 31(1):163–168, July 1976. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531556.

Anonymous:1976:BRBb

[807] Anonymous. Bibliography of recent books and articles dealing with the history of the Royal Society. Notes and Records of the Royal Society of London, 31(1):169–173, July 1976. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531557.

Anonymous:1976:FM

[808] Anonymous. Front matter. Notes and Records of the Royal Society of London, 31(1):??, July 1976. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531550.

Anonymous:1977:EAB

[809] Anonymous. Editorial: The American bicentenary and Anglo–American intellectual relations. Notes and Records of the Royal Society of London, 31(2):177, January 1977. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531825.

Bailyn:1977:BD

[810] Bernard Bailyn. 1776: The British dimension. Notes and Records of the Royal Society of London, 31(2):179–199, January 1977. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531826.

Jones:1977:BF

[811] R. V. Jones, F.R.S. Benjamin Franklin. Notes and Records of the Royal Society of London, 31(2):201–225, January 1977. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/531827.

Plumb:1977:BAC

[812] J. H. Plumb, F.B.A. Britain and America — the cultural tradition. Notes and Records of the Royal Society of London, 31(2):227–243, January 1977. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531828.

Anonymous:1977:C

[813] Anonymous. Conversaziones 1976. Notes and Records of the Royal Society of London, 31(2):245–253, January 1977. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531829.

Beckett:1977:DWB

[814] J. V. Beckett. Dr William Brownrigg, F.R.S.: Physician, chemist and country gentleman. Notes and Records of the Royal Society of London, 31 (2):255–271, January 1977. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531830.

Farrar:1977:ESF

[815] W. V. Farrar. Edward Schunck, F.R.S., a pioneer of natural-product chemistry. Notes and Records of the Royal Society of London, 31(2): 273–296, January 1977. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531831.

Cotter:1977:RSD

[816] Charles H. Cotter. The Royal Society and the deviation of the compass. Notes and Records of the Royal Society of London, 31(2):297–309, January 1977. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531832.

Anonymous:1977:SNa

[817] Anonymous. The Society’s notes. Notes and Records of the Royal Society of London, 31(2):311, January 1977. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531833.

Anonymous:1977:BRBa

[818] Anonymous. Bibliography of recent books and articles dealing with the history of the Royal Society. Notes and Records of the Royal Society of London, 31(2):313–322, January 1977. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531834.

Anonymous:1977:INR

[819] Anonymous. Index to Notes and Records, volumes 21 to 30 (1966 to 1975). Notes and Records of the Royal Society of London, 31(2):323–356, January 1977. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531824.

Todd:1977:AMS

[820] P. R. S. Lord Todd. Address at the memorial service for Sir David Christie Martin (1914–1976) at St Columba’s Church of Scotland, London on 9 February 1977. Notes and Records of the Royal Society of London, 32(1):1–3, July 1977. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531760.

Handler:1977:ADS

[821] Philip Handler. Anniversary Dinner 1976: speech by Dr Philip Handler, President of the U. S. National Academy of Sciences, The Dorchester, 30 November 1976. Notes and Records of the Royal Society of London, 32(1): 5–11, July 1977. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531761.

Middleton:1977:WDC

[822] W. E. Knowles Middleton. What did Charles II call the Fellows of the Royal Society? Notes and Records of the Royal Society of London, 32(1):

13–16, July 1977. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531762.

Schaffer:1977:HAE

[823] Simon Schaffer. Halley’s atheism and the end of the world. Notes and Records of the Royal Society of London, 32(1):17–40, July 1977. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531763.

Levere:1977:DTB

[824] Trevor H. Levere. Dr Thomas Beddoes and the establishment of his Pneumatic Institution: a tale of three Presidents. Notes and Records of the Royal Society of London, 32(1):41–49, July 1977. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/531764.

Moore:1977:EDS

[825] James R. Moore. On the education of Darwin’s sons: the correspondence between Charles Darwin and the Reverend G. V. Reed, 1857–1864. Notes and Records of the Royal Society of London, 32(1):51–70, July 1977. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531765.

Gowing:1977:STE

[826] Margaret Gowing. Science, technology and education: England in 1870. The Wilkins Lecture, 1976. Notes and Records of the Royal Society of London, 32(1):71–90, July 1977. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531766.

McGee:1977:CCS

[827] J. D. McGee, F.R.S. The contribution of A. A. Campbell Swinton, F.R.S. to television. Notes and Records of the Royal Society of London, 32(1): 91–105, July 1977. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/531767.

Anonymous:1977:SNb

[828] Anonymous. The Society’s notes. Notes and Records of the Royal Society of London, 32(1):107–112, July 1977. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531768.

Anonymous:1977:BRBb

[829] Anonymous. Bibliography of recent books and articles dealing with the history of the Royal Society. Notes and Records of the Royal Society of London, 32(1):113–117, July 1977. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531769.

Anonymous:1977:VI

[830] Anonymous. Volume information. Notes and Records of the Royal Society of London, 32(1):??, July 1977. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531759.

Hodgkin:1978:AMS

[831] Alan Hodgkin. Address at the memorial service for Lord Adrian, O.M., F.R.S., at Westminster Abbey on 18 October 1977. Notes and Records of the Royal Society of London, 32(2):119–121, March 1978. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531722.

Brimblecombe:1978:IAP

[832] Peter Brimblecombe. Interest in air pollution among early Fellows of the Royal Society. Notes and Records of the Royal Society of London, 32 (2):123–129, March 1978. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531723.

Boss:1978:CSO

[833] Jeffrey M. N. Boss. A collection of some observations on bills of mortality & parish registers: An unpublished manuscript by Stephen Hales, F.R.S. (1677–1761). Notes and Records of the Royal Society of London, 32(2): 131–147, March 1978. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/531724.

Rauschenberg:1978:JEF

[834] Roy A. Rauschenberg. John Ellis, F.R.S. eighteenth century naturalist and Royal agent to West Florida. Notes and Records of the Royal Society of London, 32(2):149–164, March 1978. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531725.

Chapin:1978:LLL

[835] Seymour L. Chapin. Lalande and the longitude: A little known London voyage of 1763. Notes and Records of the Royal Society of London, 32 (2):165–180, March 1978. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531726.

Ruggles:1978:GSW

[836] R. I. Ruggles. Governor Samuel Wegg, intelligent layman of the Royal Society, 1753–1802. Notes and Records of the Royal Society of London, 32(2):181–199, March 1978. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531727.

Bayliss:1978:TJB

[837] Robert A. Bayliss. The travels of Joseph Beete Jukes, F.R.S. Notes and Records of the Royal Society of London, 32(2):201–212, March 1978. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531728.

Bolton:1978:DBJ

[838] H. C. Bolton and W. C. Price, F.R.S. The date of birth of James Clerk Maxwell. Notes and Records of the Royal Society of London, 32(2):213– 214, March 1978. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531729.

Buchanan:1978:SEC

[839] R. A. Buchanan. Science and engineering: a case study in British experience in the mid-nineteenth century. Notes and Records of the Royal Society of London, 32(2):215–223, March 1978. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531730.

Feather:1978:IIF

[840] N. Feather, F.R.S. Isotopes, isomers and the fundamental law of radioactive change. Notes and Records of the Royal Society of London, 32 (2):225–231, March 1978. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531731.

Anonymous:1978:SNa

[841] Anonymous. The Society’s notes. Notes and Records of the Royal Society of London, 32(2):233, March 1978. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531732.

Anonymous:1978:BRBa

[842] Anonymous. Bibliography of recent books and articles dealing with the history of the Royal Society. Notes and Records of the Royal Society of London, 32(2):235–239, March 1978. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531733.

Anonymous:1978:INR

[843] Anonymous. Index to Notes and Records, volume 32, 1977. Notes and Records of the Royal Society of London, 32(2):241–243, March 1978. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531721.

Adamson:1978:RSG

[844] Ian Adamson. The Royal Society and Gresham College 1660–1711. Notes and Records of the Royal Society of London, 33(1):1–21, August 1978. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531673.

Cotter:1978:MSR

[845] Charles H. Cotter. The mariner’s sextant and the Royal Society. Notes and Records of the Royal Society of London, 33(1):23–36, August 1978. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531674.

Jones:1978:TMS

[846] R. V. Jones, F.R.S. Through music to the stars. William Herschel, 1738– 1822. Notes and Records of the Royal Society of London, 33(1):37–56, August 1978. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531675.

Lovell:1978:HWS

[847] Sir Bernard Lovell, F.R.S. Herschel’s work on the structure of the Universe. Notes and Records of the Royal Society of London, 33(1):57–75, August 1978. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531676.

Ross:1978:JHF

[848] Sydney Ross. John Herschel on Faraday and on science. Notes and Records of the Royal Society of London, 33(1):77–82, August 1978. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531677.

Freeman:1978:DNB

[849] R. B. Freeman. Darwin’s negro bird-stuffer. Notes and Records of the Royal Society of London, 33(1):83–86, August 1978. CODEN NOREAY.

ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/531678.

McGucken:1978:RSG

[850] William McGucken. The Royal Society and the genesis of the Scientific Advisory Committee to Britain’s War Cabinet, 1939–1940. Notes and Records of the Royal Society of London, 33(1):87–115, August 1978. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531679.

Anonymous:1978:SNb

[851] Anonymous. The Society’s notes. Notes and Records of the Royal Society of London, 33(1):117–122, August 1978. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531680.

Anonymous:1978:BRBb

[852] Anonymous. Bibliography of recent books and articles dealing with the history of the Royal Society. Notes and Records of the Royal Society of London, 33(1):123–132, August 1978. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531681.

Anonymous:1978:VI

[853] Anonymous. Volume information. Notes and Records of the Royal Society of London, 33(1):??, August 1978. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531672.

Mills:1979:NTE

[854] A. A. Mills and P. J. Turvey. Newton’s telescope, an examination of the reflecting telescope attributed to Sir Isaac Newton in the possession of the Royal Society. Notes and Records of the Royal Society of London, 33 (2):133–155, March 1979. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531664.

Middleton:1979:SIV

[855] W. E. Knowles Middleton. Some Italian visitors to the early Royal Society. Notes and Records of the Royal Society of London, 33(2):157–173, March 1979. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531665.

Fryer:1979:JFF

[856] Geoffrey Fryer, F.R.S. John Fryer, F.R.S. and his scientific observations, made chiefly in India and Persia between 1672 and 1682. Notes and Records of the Royal Society of London, 33(2):175–206, March 1979. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531666.

Gunther:1979:RSF

[857] A. E. Gunther. The Royal Society and the foundation of the British Museum, 1753–1781. Notes and Records of the Royal Society of London, 33(2):207–216, March 1979. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531667.

Oldroyd:1979:CLO

[858] D. R. Oldroyd and D. W. Hutchings. The chemical lectures at Oxford (1822–1854) of Charles Daubeny, M.D., F.R.S. Notes and Records of the Royal Society of London, 33(2):217–259, March 1979. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531668.

Shorter:1979:HOJ

[859] John Shorter. Humphrey Owen Jones, F.R.S. (1878–1912) chemist and mountaineer. Notes and Records of the Royal Society of London, 33 (2):261–277, March 1979. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531669.

Anonymous:1979:SNa

[860] Anonymous. The Society’s notes. Notes and Records of the Royal Society of London, 33(2):279, March 1979. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531670.

Anonymous:1979:BRB

[861] Anonymous. Bibliography of recent books and articles dealing with the history of the Royal Society. Notes and Records of the Royal Society of London, 33(2):281–285, March 1979. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531671.

Anonymous:1979:INR

[862] Anonymous. Index to Notes and Records, volume 33, 1978–9. Notes and Records of the Royal Society of London, 33(2):287–288, March 1979.

CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531663.

Johnson:1979:MPP

[863] L. W. Johnson and M. L. Wolbarsht. Mercury poisoning: A probable cause of Isaac Newton’s physical and mental ills. Notes and Records of the Royal Society of London, 34(1):1–9, July 1979. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531510.

Spargo:1979:NDI

[864] P. E. Spargo and C. A. Pounds. Newton’s ‘derangement of the intellect’. New light on an old problem. Notes and Records of the Royal Society of London, 34(1):11–32, July 1979. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531511.

Edmonds:1979:FOR

[865] J. M. Edmonds. The founding of the Oxford Readership in Geology, 1818. Notes and Records of the Royal Society of London, 34(1):33–51, July 1979. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531512.

Gunther:1979:JGC

[866] A. E. Gunther. J. E. Gray, Charles Darwin, and the Cirripedes, 1846– 1851. Notes and Records of the Royal Society of London, 34(1):53–63, July 1979. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531513.

Jones:1979:AER

[867] R. V. Jones, F.R.S. Alfred Ewing and ‘Room 40’. Notes and Records of the Royal Society of London, 34(1):65–90, July 1979. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/531514.

Badash:1979:BAV

[868] Lawrence Badash. British and American views of the German menace in World War I. Notes and Records of the Royal Society of London, 34(1):91–121, July 1979. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531515.

Gowing:1979:CSA

[869] Margaret Gowing. The contemporary scientific archives centre. Notes and Records of the Royal Society of London, 34(1):123–131, July 1979. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531516.

Anonymous:1979:SNb

[870] Anonymous. The Society’s notes. Notes and Records of the Royal Society of London, 34(1):133–137, July 1979. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531517.

Anonymous:1979:FM

[871] Anonymous. Front matter. Notes and Records of the Royal Society of London, 34(1):??, July 1979. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531509.

Whitteridge:1980:LMA

[872] Gweneth Whitteridge. Of the local movement of animals: The Wilkins Lecture, 1979. Notes and Records of the Royal Society of London, 34(2): 139–153, February 1980. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531806.

Cumming:1980:JMF

[873] David A. Cumming. John MacCulloch, F.R.S., at Addiscombe: The lectureships in chemistry and geology. Notes and Records of the Royal Society of London, 34(2):155–183, February 1980. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531807.

Brown:1980:MCS

[874] Joyce Brown. A memoir of Colonel Sir Proby Cautley, F.R.S., 1802–1871, engineer and palaeontologist. Notes and Records of the Royal Society of London, 34(2):185–225, February 1980. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531808.

Brown:1980:DEH

[875] G. Burniston Brown. David Edward Hughes, F.R.S., 1831–1900. Notes and Records of the Royal Society of London, 34(2):227–239, February 1980. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531809.

Alter:1980:RSI

[876] Peter Alter. The Royal Society and the International Association of Academies 1897–1919. Notes and Records of the Royal Society of London, 34(2):241–264, February 1980. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531810.

Anonymous:1980:SN

[877] Anonymous. The Society’s notes. Notes and Records of the Royal Society of London, 34(2):265, February 1980. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531811.

Anonymous:1980:INRa

[878] Anonymous. Index to Notes and Records, volume 34, 1979–80. Notes and Records of the Royal Society of London, 34(2):267–268, March 1980. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531804.

Anonymous:1980:FMa

[879] Anonymous. Front matter. Notes and Records of the Royal Society of London, 34(2):??, March 1980. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531805.

Ditchburn:1980:NI

[880] R. W. Ditchburn, F.R.S. Newton’s illness of 1692–3. Notes and Records of the Royal Society of London, 35(1):1–16, July 1980. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/531598.

Oldroyd:1980:SPS

[881] D. R. Oldroyd. Some ‘Philosophicall Scribbles’ attributed to Robert Hooke. Notes and Records of the Royal Society of London, 35(1):17– 32, July 1980. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531599.

Bennett:1980:RHM

[882] J. A. Bennett. Robert Hooke as mechanic and natural philosopher. Notes and Records of the Royal Society of London, 35(1):33–48, July 1980. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531600.

Austin:1980:JSF

[883] Jillian F. Austin and Anita McConnell. James Six F.R.S. — two hundred years of the Six’s self-registering thermometer. Notes and Records of the Royal Society of London, 35(1):49–65, July 1980. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/531601.

Domb:1980:JCM

[884] C. Domb, F.R.S. James Clerk Maxwell in London: 1860–1865. Notes and Records of the Royal Society of London, 35(1):67–103 + 1, July 1980. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531602.

Anonymous:1980:FMb

[885] Anonymous. Front matter. Notes and Records of the Royal Society of London, 35(1):??, July 1980. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531597.

Cavazza:1980:BRS

[886] Marta Cavazza. Bologna and the Royal Society in the Seventeenth Century. Notes and Records of the Royal Society of London, 35(2):105–123, December 1980. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531502.

Sloan:1980:SHS

[887] W. R. Sloan. Sir Hans Sloane, F.R.S. legend and lineage. Notes and Records of the Royal Society of London, 35(2):125–129, December 1980. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531503.

Woolrich:1980:PSR

[888] A. P. Woolrich. The printing of Smeaton’s reports. Notes and Records of the Royal Society of London, 35(2):131–133, December 1980. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531504.

Manuel:1980:MHF

[889] Diana E. Manuel. Marshall Hall, F.R.S. (1790–1857), a conspectus of his life and work. Notes and Records of the Royal Society of London, 35(2): 135–166, December 1980. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531505.

Moseley:1980:GSR

[890] Russell Moseley. Government science and the Royal Society: the control of the national physical laboratory in the inter-war years. Notes and Records of the Royal Society of London, 35(2):167–193, December 1980. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531506.

Witkowski:1980:WAR

[891] J. A. Witkowski. W. T. Astbury and Ross G. Harrison: the search for the molecular determination of form in the developing embryo. Notes and Records of the Royal Society of London, 35(2):195–219, December 1980. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531507.

Anonymous:1980:INRb

[892] Anonymous. Index to Notes and Records, volume 35, 1980. Notes and Records of the Royal Society of London, 35(2):221–222, December 1980. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531500.

Anonymous:1980:FMc

[893] Anonymous. Front matter. Notes and Records of the Royal Society of London, 35(2):??, December 1980. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531501.

Anonymous:1981:ENN

[894] Anonymous. Editorial note: Newton’s rainbow. Notes and Records of the Royal Society of London, 36(1):1, August 1981. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531654.

Bishop:1981:RWM

[895] Roy L. Bishop and Isaac Newton. Rainbow over Woolsthorpe Manor. Notes and Records of the Royal Society of London, 36(1):2–11, August 1981. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531654. Page 1 is a color

plate of the double rainbow, and pages 4–11 are reprinted from Newton’s Opticks, pages 126–132.

Mills:1981:NPH

[896] A. A. Mills. Newton’s prisms and his experiments on the spectrum. Notes and Records of the Royal Society of London, 36(1):13–36, August 1981.

CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531655.

Ford:1981:VLS

[897] Brian J. Ford. The van Leeuwenhoek specimens. Notes and Records of the Royal Society of London, 36(1):37–59, August 1981. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531656.

Cotter:1981:CEH

[898] Charles H. Cotter. Captain Edmond Halley R.N., F.R.S. Notes and Records of the Royal Society of London, 36(1):61–77, August 1981. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531657.

Lysaght:1981:LSP

[899] Averil Lysaght. A letter from Sydney Parkinson in Batavia to Dr John Fothergill. Notes and Records of the Royal Society of London, 36(1):79– 81, August 1981. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531658.

Flinn:1981:JMM

[900] Derek Flinn. John MacCulloch, M.D., F.R.S., and his geological map of Scotland: His years in the Ordnance. 1795–1826. Notes and Records of the Royal Society of London, 36(1):83–101, August 1981. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531659.

Bayliss:1981:NAF

[901] Robert A. Bayliss and C. William Ellis. Neil Arnott, F.R.S. reformer, innovator and popularizer of science 1788–1874. Notes and Records of the Royal Society of London, 36(1):103–123, August 1981. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531660.

Scott-Moncrieff:1981:CPC

[902] Rose Scott-Moncrieff, (Mrs O.M.Meares). The classical period in chemical genetics, recollections of Muriel Wheldale Onslow, Robert and Gertrude Robinson and J. B. S. Haldane. Notes and Records of the Royal Society of London, 36(1):125–154, August 1981. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/531661.

Anonymous:1981:BM

[903] Anonymous. Back matter. Notes and Records of the Royal Society of London, 36(1):??, August 1981. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531662.

Anonymous:1981:FM

[904] Anonymous. Front matter. Notes and Records of the Royal Society of London, 36(1):??, August 1981. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531653.

Gouk:1982:AER

[905] Penelope M. Gouk. Acoustics in the early Royal Society 1660–1680. Notes and Records of the Royal Society of London, 36(2):155–175, February 1982. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531772.

Pasmore:1982:THF

[906] Stephen Pasmore. Thomas Henshaw, F.R.S. (1618–1700). Notes and Records of the Royal Society of London, 36(2):177–188, February 1982. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531773.

Hunter:1982:EPP

[907] Michael Hunter. Early problems in professionalizing scientific research: Nehemiah Grew (1641–1712) and the Royal Society, with an unpublished letter to Henry Oldenburg. Notes and Records of the Royal Society of London, 36(2):189–209, February 1982. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531774.

Wallis:1982:JBM

[908] Ruth Wallis. John Bevis, M.D., F.R.S., (1695–1771) Astronomer Loyal. Notes and Records of the Royal Society of London, 36(2):211–225, February 1982. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531775.

Wilkinson:1982:OJH

[909] Lise Wilkinson. ‘The other’ John Hunter, M.D., F.R.S. (1754–1809): His contributions to the medical literature, and to the introduction of animal experiments into infectious disease research. Notes and Records of the Royal Society of London, 36(2):227–241, February 1982. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531776.

Gooding:1982:COD

[910] David Gooding. A convergence of opinion on the divergence of lines: Faraday and Thomson’s discussion of diamagnetism. Notes and Records of the Royal Society of London, 36(2):243–259, February 1982. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531777.

Matthews:1982:SSD

[911] L. Harrison Matthews, F.R.S. Samuel Stutchbury and Darwin’s cirripedes. Notes and Records of the Royal Society of London, 36(2):261–266, February 1982. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531778.

Gorham:1982:RAS

[912] Eville Gorham. Robert Angus Smith, F.R.S., and ‘chemical climatology’. Notes and Records of the Royal Society of London, 36(2):267–272, February 1982. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531779.

Anonymous:1982:INR

[913] Anonymous. Index to Notes and Records, volume 36, 1981–82. Notes and Records of the Royal Society of London, 36(2):273–275, February 1982. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531770.

Anonymous:1982:BMa

[914] Anonymous. Back matter. Notes and Records of the Royal Society of London, 36(2):??, February 1982. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531780.

Anonymous:1982:FMa

[915] Anonymous. Front matter. Notes and Records of the Royal Society of London, 36(2):??, February 1982. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531771.

Cornforth:1982:PDH

[916] Sir John Cornforth, F.R.S. Portrait of Dorothy Hodgkin, O.M., F.R.S. Notes and Records of the Royal Society of London, 37(1):1–4, August 1982. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531473.

Bowen:1982:WFR

[917] E. J. Bowen, F.R.S. Who founded the R.S.? Notes and Records of the Royal Society of London, 37(1):5, August 1982. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531474.

Hall:1982:FNC

[918] A. Rupert Hall. Further Newton correspondence. Notes and Records of the Royal Society of London, 37(1):7–34, August 1982. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531475.

Mills:1982:NWC

[919] A. A. Mills. Newton’s water clocks and the fluid mechanics of clepsydrae. Notes and Records of the Royal Society of London, 37(1):35–61, August 1982. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531476. See errata

[941].

Hall:1982:RSI

[920] Marie Boas Hall. The Royal Society and Italy 1667–1795. Notes and Records of the Royal Society of London, 37(1):63–81, August 1982. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531477.

Williams:1982:JBE

[921] Mari Williams. James Bradley and the Eighteenth Century ‘gap’ in attempts to measure annual stellar parallax. Notes and Records of the Royal Society of London, 37(1):83–100, August 1982. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531478.

Miles:1982:RLP

[922] Sir Ashley Miles, F.R.S. Reports by Louis Pasteur and Claude Bernard on the organization of scientific teaching and research. Notes and Records of the Royal Society of London, 37(1):101–118, August 1982. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531479.

McGee:1982:CCS

[923] J. D. McGee, F.R.S. The contribution of A. A. Campbell Swinton, F.R.S., to television — a correction. Notes and Records of the Royal Society of London, 37(1):119, August 1982. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531480.

Anonymous:1982:BMb

[924] Anonymous. Back matter. Notes and Records of the Royal Society of London, 37(1):??, August 1982. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531481.

Anonymous:1982:FMb

[925] Anonymous. Front matter. Notes and Records of the Royal Society of London, 37(1):??, August 1982. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531472.

Appleby:1983:GRS

[926] John H. Appleby. Ginseng and the Royal Society. Notes and Records of the Royal Society of London, 37(2):121–145, March 1983. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531484.

Mills:1983:PPS

[927] A. A. Mills and R. Hall. The production of a plane surface, as illustrated by specula from some early Newtonian telescopes. Notes and Records of the Royal Society of London, 37(2):147–166, March 1983. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531485.

Crosland:1983:EQC

[928] Maurice Crosland. Explicit qualifications as a criterion for membership of the Royal Society: a historical review. Notes and Records of the Royal Society of London, 37(2):167–187, March 1983. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531486.

Crane:1983:SSN

[929] Michael D. Crane. Samuel Stutchbury (1798–1859), naturalist and geologist. Notes and Records of the Royal Society of London, 37(2):189–200, March 1983. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531487.

Baker:1983:LCM

[930] R. A. Baker and R. A. Bayliss. Louis Compton Miall, F.R.S.: scientist and educator, 1842–1921. Notes and Records of the Royal Society of London, 37(2):201–234, March 1983. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531488.

Blaschko:1983:FHS

[931] H. Blaschko, F.R.S. Frederick Hughes Scott and his contribution to the early history of the transmitter concept. Notes and Records of the Royal Society of London, 37(2):235–247, March 1983. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531489.

Cock:1983:CIS

[932] A. G. Cock. Chauvinism and internationalism in science: the International Research Council, 1919–1926. Notes and Records of the Royal Society of London, 37(2):249–288, March 1983. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/ stable/531490. See note [963].

Anonymous:1983:INR

[933] Anonymous. Index to Notes and Records, volume 37, 1982–1983. Notes and Records of the Royal Society of London, 37(2):289–290, March 1983. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531482.

Anonymous:1983:BMa

[934] Anonymous. Back matter. Notes and Records of the Royal Society of London, 37(2):??, March 1983. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531491.

Anonymous:1983:FMa

[935] Anonymous. Front matter. Notes and Records of the Royal Society of London, 37(2):??, March 1983. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531483.

Griffith:1983:PL

[936] William P. Griffith. Priestley in London. Notes and Records of the Royal Society of London, 38(1):1–16, August 1983. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531343.

OHara:1983:GRS

[937] James Gabriel O’Hara. Gauss and the Royal Society: the reception of his ideas on magnetism in Britain (1832–1842). Notes and Records of the Royal Society of London, 38(1):17–78, August 1983. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531344.

James:1983:CET

[938] Frank A. J. L. James. The conservation of energy, theories of absorption and resonating molecules, 1851–1854: G. G. Stokes, A. J. Ångström and W. Thomson. Notes and Records of the Royal Society of London, 38 (1):79–107, August 1983. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531345.

Mills:1983:DIT

[939] William Mills. Darwin and the iceberg theory. Notes and Records of the Royal Society of London, 38(1):109–127, August 1983. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531346.

Morgan:1983:WDH

[940] Neil Morgan. William Dobinson Halliburton, F.R.S. (1860–1931) pioneer of British biochemistry? Notes and Records of the Royal Society of London, 38(1):129–145, August 1983. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531347.

Mills:1983:ENW

[941] A. A. Mills. Errata: “Newton’s water clocks and the fluid mechanics of clepsydrae” [Notes and Records Roy. Soc. London 37 (1982), no. 1, 35–61; MR0684478 (84j:01041)]. Notes and Records of the Royal Society of London, 38(1):146, August 1983. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/ 531348. See [919].

Anonymous:1983:BMb

[942] Anonymous. Back matter. Notes and Records of the Royal Society of London, 38(1):??, August 1983. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531349.

Anonymous:1983:FMb

[943] Anonymous. Front matter. Notes and Records of the Royal Society of London, 38(1):??, August 1983. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531342.

Huxley:1984:GG

[944] Sir Andrew Huxley, O.M., P.R.S. Grandfather and grandson. Notes and Records of the Royal Society of London, 38(2):147–151, March 1984. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531814.

Hall:1984:RST

[945] Marie Boas Hall. The Royal Society in Thomas Henry Huxley’s time. Notes and Records of the Royal Society of London, 38(2):153–158, March 1984. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531815.

Hunter:1984:CRS

[946] Michael Hunter. A ‘College’ for the Royal Society: The abortive plan of 1667–1668. Notes and Records of the Royal Society of London, 38 (2):159–186, March 1984. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531816.

Simpson:1984:NTC

[947] A. D. C. Simpson. Newton’s telescope and the cataloguing of the Royal Society’s repository. Notes and Records of the Royal Society of London, 38(2):187–214, March 1984. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531817.

Ezell:1984:RWR

[948] Margaret J. M. Ezell. Richard Waller, S.R.S.: ‘In the pursuit of nature’. Notes and Records of the Royal Society of London, 38(2):215–233, March 1984. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531818.

Mills:1984:NBM

[949] Stella Mills. Note on the Braikenridge–Maclaurin theorem. Notes and Records of the Royal Society of London, 38(2):235–240, March 1984. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531819.

Tansey:1984:LWS

[950] E. M. Tansey. The life and works of Sir Alexander Crichton, F.R.S. (1763–1856): a Scottish physician to the Imperial Russian Court. Notes and Records of the Royal Society of London, 38(2):241–259, March 1984. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531820.

Deacon:1984:GHF

[951] Margaret B. Deacon. G. Herbert Fowler (1861–1940): The forgotten oceanographer. Notes and Records of the Royal Society of London, 38 (2):261–296, March 1984. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531821.

Shoppee:1984:SER

[952] C. W. Shoppee, F.R.S. The sign of the electron-releasing effect of the methyl group. Notes and Records of the Royal Society of London, 38(2): 297, March 1984. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531822.

Anonymous:1984:INR

[953] Anonymous. Index to Notes and Records, volume 38, 1983–1984. Notes and Records of the Royal Society of London, 38(2):299–300, March 1984. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531812.

Anonymous:1984:BMa

[954] Anonymous. Back matter. Notes and Records of the Royal Society of London, 38(2):??, March 1984. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531823.

Anonymous:1984:FMa

[955] Anonymous. Front matter. Notes and Records of the Royal Society of London, 38(2):??, March 1984. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531813.

Rivington:1984:EPR

[956] Charles A. Rivington. Early printers to the Royal Society 1663–1708. Notes and Records of the Royal Society of London, 39(1):1–27, September 1984. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531572. See addendum

[990].

Cardwell:1984:FLJ

[957] Donald Cardwell and Joan Mottram. Fresh light on John Dalton. Notes and Records of the Royal Society of London, 39(1):29–40, September 1984. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531573.

Katritzky:1984:GCL

[958] Linde Katritzky. Georg Christoph Lichtenberg, F.R.S. Notes and Records of the Royal Society of London, 39(1):41–49, September 1984. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531574. See corrigendum [975].

Gorman:1984:SWB

[959] Mel Gorman. Sir William Brooke O’Shaughnessy, F.R.S. (1809–1889) Anglo–Indian forensic chemist. Notes and Records of the Royal Society of London, 39(1):51–64, September 1984. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531575.

James:1984:MMC

[960] Frank A. J. L. James. Of ‘medals and muddles’ the context of the discovery of thallium: William Crookes’s early spectro-chemical work. Notes and Records of the Royal Society of London, 39(1):65–90, September 1984. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531576.

Sinclair:1984:SIF

[961] H. M. Sinclair. Sherrington and industrial fatigue. Notes and Records of the Royal Society of London, 39(1):91–104, September 1984. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531577.

Waters:1984:SCD

[962] W. A. Waters, F.R.S. Some comments on the development of free radical chemistry. Notes and Records of the Royal Society of London, 39(1):

105–124, September 1984. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531578.

See corrigendum [973].

Cartwright:1984:NCI

[963] Dame Mary Cartwright, F.R.S. Note on: “Chauvinism and internationalism in science: the International Research Council, 1919–1926” [Notes and Records Roy. Soc. London 37 (1983), no. 2, 249–288; MR0782423 (86h:01097)] by A. G. Cock. Notes and Records of the Royal Society of London, 39(1):125–128, September 1984. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531579. See [932].

Anonymous:1984:BMb

[964] Anonymous. Back matter. Notes and Records of the Royal Society of London, 39(1):??, September 1984. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531580.

Anonymous:1984:FMb

[965] Anonymous. Front matter. Notes and Records of the Royal Society of London, 39(1):??, September 1984. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531571.

Ochs:1985:RSL

[966] Kathleen H. Ochs. The Royal Society of London’s history of trades programme: An early episode in applied science. Notes and Records of the Royal Society of London, 39(2):129–158, April 1985. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531622.

Mendyk:1985:RPB

[967] S. Mendyk. Robert Plot: Britain’s ‘genial father of county natural histories’. Notes and Records of the Royal Society of London, 39(2):159–177, April 1985. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531623.

Heberden:1985:CWH

[968] Ernest Heberden. Correspondence of William Heberden, F.R.S. with The Reverend Stephen Hales and Sir Charles Blagden. Notes and Records of the Royal Society of London, 39(2):179–189, April 1985. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531624.

Freemon:1985:ACS

[969] Frank R. Freemon. American colonial scientists who published in the Philosophical Transactions of the Royal Society. Notes and Records of the Royal Society of London, 39(2):191–206, April 1985. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531625.

Arthur:1985:DDG

[970] C. W. P. Mac Arthur. Davy’s differences with Gay-Lussac and Thenard: new light on events in Paris and on the transmission and translation of Davy’s papers in 1810. Notes and Records of the Royal Society of London, 39(2):207–228, April 1985. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531626.

Gooding:1985:WPD

[971] David Gooding. ‘He who proves, discovers’: John Herschel, William Pepys and the Faraday effect. Notes and Records of the Royal Society of London, 39(2):229–244, April 1985. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531627.

Hutchison:1985:RSF

[972] Sir Kenneth Hutchison, F.R.S. The Royal Society and the foundation of the British gas industry. Notes and Records of the Royal Society of London, 39(2):245–270, April 1985. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531628.

Anonymous:1985:CWW

[973] Anonymous. Corrigendum: W. A. Waters: Some comments on the development of Free Radical Chemistry. Notes and Records of the Royal Society of London, 39(2):271, April 1985. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531629. See [962].

Anonymous:1985:INR

[974] Anonymous. Index to Notes and Records, volume 39, 1984–1985. Notes and Records of the Royal Society of London, 39(2):272–274, April 1985. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531620.

Katritzky:1985:CGC

[975] Linde Katritzky. Corrigendum: “Georg Christoph Lichtenberg, F.R.S.” [Notes and Records Roy. Soc. London 39 (1984), no. 1, 41–49; MR0782424 (86i:01049)]. Notes and Records of the Royal Society of London, 39(2):274, April 1985. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531630.

See [958].

Anonymous:1985:BMa

[976] Anonymous. Back matter. Notes and Records of the Royal Society of London, 39(2):??, April 1985. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531631.

Anonymous:1985:FMa

[977] Anonymous. Front matter. Notes and Records of the Royal Society of London, 39(2):??, April 1985. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531621.

Cohen:1985:CBE

[978] Edward H. Cohen, John S. Ross, and F. R. S. Sir William Paton. The commonplace book of Edmond Halley. Notes and Records of the Royal Society of London, 40(1):1–40, November 1985. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531533.

Pugh:1985:CWJ

[979] Jennifer S. Pugh, John Hudson, and F. R. S. Sir William Paton. The chemical work of James Watt, F.R.S. Notes and Records of the Royal Society of London, 40(1):41–52, November 1985. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531534.

James:1985:BTS

[980] Frank A. J. L. James and F. R. S. Sir William Paton. Between two scientific generations: John Herschel’s rejection of the conservation of energy in his 1864 correspondence with William Thomson. Notes and Records of the Royal Society of London, 40(1):53–62, November 1985. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531535.

Tweedale:1985:SRA

[981] Geoffrey Tweedale and F. R. S. Sir William Paton. Sir Robert Abbott Hadfield F.R.S. (1858–1940) and the discovery of manganese steel. Notes and Records of the Royal Society of London, 40(1):63–74, November 1985. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531536.

Jenkin:1985:AWH

[982] John G. Jenkin and Sir William Paton, F.R.S. The appointment of W. H. Bragg, F.R.S. to the University of Adelaide. Notes and Records of the Royal Society of London, 40(1):75–99, November 1985. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531537.

Anonymous:1985:BMb

[983] Anonymous. Back matter. Notes and Records of the Royal Society of London, 40(1):??, November 1985. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531538.

Anonymous:1985:FMb

[984] Anonymous. Front matter. Notes and Records of the Royal Society of London, 40(1):??, November 1985. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531532.

Stearn:1986:WLJ

[985] William T. Stearn. The Wilkins Lecture, 1985 John Wilkins, John Ray and Carl Linnaeus. Notes and Records of the Royal Society of London, 40(2):101–123, May 1986. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531684.

Booth:1986:SSG

[986] C. C. Booth. Sir Samuel Garth, F.R.S.: The dispensary poet. Notes and Records of the Royal Society of London, 40(2):125–145, May 1986. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531685.

Appleby:1986:HJF

[987] John H. Appleby. Humphrey Jackson, F.R.S., 1717–1801: A pioneering chemist. Notes and Records of the Royal Society of London, 40(2):147– 168, May 1986. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531686.

Blouet:1986:SWR

[988] Olwyn Mary Blouet. Sir William Reid, F.R.S., 1791–1858: Governor of Bermuda, Barbados and Malta. Notes and Records of the Royal Society of London, 40(2):169–191, May 1986. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531687.

Richards:1986:CDC

[989] Stewart Richards. Conan Doyle’s ‘challenger’ unchampioned: William Rutherford, F.R.S. (1839–99), and the origins of practical physiology in Britain. Notes and Records of the Royal Society of London, 40(2):193– 217, May 1986. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531688.

Rivington:1986:AEP

[990] Charles A. Rivington. Addendum: Early printers to the Royal Society, 1663–1708. Notes and Records of the Royal Society of London, 40(2):219– 220, May 1986. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531689. See [956].

Anonymous:1986:INR

[991] Anonymous. Index to Notes and Records, volume 40, 1985–1986. Notes and Records of the Royal Society of London, 40(2):221–222, May 1986. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531682.

Anonymous:1986:BMa

[992] Anonymous. Back matter. Notes and Records of the Royal Society of London, 40(2):??, May 1986. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531690.

Anonymous:1986:FMa

[993] Anonymous. Front matter. Notes and Records of the Royal Society of London, 40(2):??, May 1986. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531683.

Brodsley:1986:PRD

[994] Laurel Brodsley, Sir Charles Frank, F.R.S., and John W. Steeds. Prince Rupert’s drops. Notes and Records of the Royal Society of London, 41 (1):1–26, October 1986. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531493.

See erratum [1008].

Applebaum:1986:DCM

[995] Wilbur Applebaum. A descriptive catalogue of the manuscripts of Nicolaus Mercator, F.R.S. (1620–87), in Sheffield University Library. Notes and Records of the Royal Society of London, 41(1):27–37, October 1986. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531494.

McConnell:1986:FMV

[996] A. McConnell. L. F. Marsigli’s voyage to London and Holland, 1721– 1722. Notes and Records of the Royal Society of London, 41(1):39–76, October 1986. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531495.

Foster:1986:HDB

[997] P. G. M. Foster. The Hon. Daines Barrington, F.R.S. — annotations on two journals compiled by Gilbert White. Notes and Records of the Royal Society of London, 41(1):77–93, October 1986. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531496.

Ford:1986:LCD

[998] Brian J. Ford. The Leeuwenhoekiana of Clifford Dobell. Notes and Records of the Royal Society of London, 41(1):95–105, October 1986. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531497.

J:1986:BRR

[999] R. V. J. Book review: The Royal Institution: An Informal History by Gwendy Caroe, (pp. xi+ 180). Published by John Murray, London 1985, £13.95. Notes and Records of the Royal Society of London, 41(1):107–110, October 1986. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531498.

Anonymous:1986:BMb

[1000] Anonymous. Back matter. Notes and Records of the Royal Society of London, 41(1):??, October 1986. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531499.

Anonymous:1986:FMb

[1001] Anonymous. Front matter. Notes and Records of the Royal Society of London, 41(1):??, October 1986. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531492.

Hall:1987:BMW

[1002] Marie Boas Hall. Boyle’s method of work: Promoting his corpuscular philosophy. Notes and Records of the Royal Society of London, 41(2): 111–143, June 1987. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/531541.

Oldroyd:1987:SWR

[1003] D. R. Oldroyd. Some writings of Robert Hooke on procedures for the prosecution of scientific inquiry, including his ‘Lectures of Things Requisite to a Ntral History’. Notes and Records of the Royal Society of London, 41(2):145–167, June 1987. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531542.

Baird:1987:SIU

[1004] K. A. Baird. Some influences upon the young Isaac Newton. Notes and Records of the Royal Society of London, 41(2):169–179, June 1987. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531543.

Berman:1987:LP

[1005] R. Berman. I. Lindemann in physics. Notes and Records of the Royal Society of London, 41(2):181–189, June 1987. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531544.

Jones:1987:LBL

[1006] R. V. Jones, F.R.S. 2. Lindemann beyond the laboratory. Notes and Records of the Royal Society of London, 41(2):191–210, June 1987. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531545. See corrigendum [1191].

Dronamraju:1987:SAL

[1007] Krishna R. Dronamraju. On some aspects of the life and work of John Burdon Sanderson Haldane, F.R.S., in India. Notes and Records of the Royal Society of London, 41(2):211–237, June 1987. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531546.

Frank:1987:EPR

[1008] Sir Charles Frank, F.R.S. and John W. Steeds. Erratum: Prince Rupert’s drops. Notes and Records of the Royal Society of London, 41(2): 239, June 1987. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531547. See [994].

Anonymous:1987:INR

[1009] Anonymous. Index to Notes and Records, volumes 31 to 40 (1976–1986). Notes and Records of the Royal Society of London, 41(2):241–262, June 1987. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531539.

Anonymous:1987:BM

[1010] Anonymous. Back matter. Notes and Records of the Royal Society of London, 41(2):??, June 1987. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531548.

Anonymous:1987:FM

[1011] Anonymous. Front matter. Notes and Records of the Royal Society of London, 41(2):??, June 1987. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531540.

Anonymous:1988:FMa

[1012] Anonymous. Front matter. Notes and Records of the Royal Society of London, 42(1):1, January 1988. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531362.

King-Hele:1988:NPL

[1013] D. G. King-Hele and A. R. Hall. Newton’s Principia and its legacy. Notes and Records of the Royal Society of London, 42(1):1–122, January 1988. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

King-Hele:1988:IR

[1014] D. G. King-Hele, F.R.S. Introductory remarks. Notes and Records of the Royal Society of London, 42(1):3, January 1988. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www. jstor.org/stable/531364. Newton’s ıt Principia and its legacy.

Hall:1988:NAY

[1015] A. R. Hall, F.B.A. Newtonianism after 300 years. Notes and Records of the Royal Society of London, 42(1):5–9, January 1988. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:/ /www.jstor.org/stable/531365. Newton’s ıt Principia and its legacy.

Whiteside:1988:EP

[1016] D. T. Whiteside, F.B.A. The evolution of the Principia from 1665 to 1686. Notes and Records of the Royal Society of London, 42(1):11, January 1988. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531366. Newton’s ıt

Principia and its legacy.

Fellmann:1988:PCM

[1017] E. A. Fellmann. The Principia and continental mathematicians. Notes and Records of the Royal Society of London, 42(1):13–34, January 1988. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531367. Newton’s ıt Principia and its

legacy.

Casini:1988:NPP

[1018] P. Casini. Newton’s Principia and the philosophers of the Enlightenment. Notes and Records of the Royal Society of London, 42(1):35–52, January 1988. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531368. Newton’s ıt

Principia and its legacy.

Hughes:1988:PC

[1019] D. W. Hughes. The Principia and comets. Notes and Records of the Royal Society of London, 42(1):53–74, January 1988. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/ stable/531369. Newton’s ıt Principia and its legacy.

Harman:1988:NMP

[1020] P. M. Harman. Newton to Maxwell: The ‘Principia’ and British physics. Notes and Records of the Royal Society of London, 42(1):75–96, January 1988. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531370. Newton’s

ıt Principia and its legacy.

Thompson:1988:PCM

[1021] J. M. T. Thompson, F.R.S. The Principia and contemporary mechanics: Chaotic dynamics and the new unpredictability. Notes and Records of the Royal Society of London, 42(1):97–122, January 1988. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:/ /www.jstor.org/stable/531371. Newton’s ıt Principia and its legacy.

Anonymous:1988:BMa

[1022] Anonymous. Back matter. Notes and Records of the Royal Society of London, 42(1):??, January 1988. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531372.

Anonymous:1988:IIN

[1023] Anonymous. [illustration]: Isaac Newton by Godfrey Kneller, reproduced by kind permission of the Earl of Portsmouth. Notes and Records of the Royal Society of London, 42(1):NP, January 1988. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Grice-Hutchinson:1988:SSC

[1024] Marjorie Grice-Hutchinson. Some Spanish contributions to the early activities of the Royal Society of London. Notes and Records of the Royal Society of London, 42(2):123–132, July 1988. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531352.

Nuttall:1988:CCS

[1025] R. H. Nuttall. That curious curiosity: The scotoscope. Notes and Records of the Royal Society of London, 42(2):133–138, July 1988. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531353.

Bullard:1988:MSN

[1026] Margaret Bullard. My small Newtonian sweeper — where is it now? Notes and Records of the Royal Society of London, 42(2):139–148, July 1988. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531354.

King-Hele:1988:EDM

[1027] D. G. King-Hele, F.R.S. Erasmus Darwin, man of ideas and inventor of words. Notes and Records of the Royal Society of London, 42(2):149– 180, July 1988. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531355.

Jensen:1988:THH

[1028] J. V. Jensen. Thomas Henry Huxley’s lecture tour of the United States, 1876. Notes and Records of the Royal Society of London, 42(2):181– 195, July 1988. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531356.

Wright:1988:NTO

[1029] E. M. Wright. Number theory and other reminiscences of Viscount Cherwell. Notes and Records of the Royal Society of London, 42(2):197–204, July 1988. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531357.

Boag:1988:LKH

[1030] J. W. Boag and D. Shoenberg. Letters from Kapitza to his mother, 1921–27. Notes and Records of the Royal Society of London, 42(2):205– 228, July 1988. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531358.

Peierls:1988:RB

[1031] Sir Rudolf Peierls, F.R.S. Rutherford and Bohr. Notes and Records of the Royal Society of London, 42(2):229–241, July 1988. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531359.

Fellgett:1988:SIY

[1032] Mary Fellgett. Some influences on the young Isaac Newton. Notes and Records of the Royal Society of London, 42(2):243, July 1988. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531360.

Anonymous:1988:INR

[1033] Anonymous. Index to Notes and Records, volume 42, 1988. Notes and Records of the Royal Society of London, 42(2):245–246, July 1988. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531350.

Anonymous:1988:BMb

[1034] Anonymous. Back matter. Notes and Records of the Royal Society of London, 42(2):??, July 1988. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531361.

Anonymous:1988:FMb

[1035] Anonymous. Front matter. Notes and Records of the Royal Society of London, 42(2):??, July 1988. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531351.

Brand:1989:DRE

[1036] J. C. D. Brand. The discovery of the Raman effect. Notes and Records of the Royal Society of London, 43(1):1–23, January 1989. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531714.

Wood:1989:HBU

[1037] P. B. Wood. Hevelius’s business: an unpublished letter from Henry Oldenburg to the Earl of Tweeddale. Notes and Records of the Royal Society of London, 43(1):25–29, January 1989. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531715.

Simms:1989:BTH

[1038] D. L. Simms and P. L. Hinkley. Brighter than how many suns? Sir Isaac Newton’s burning mirror. Notes and Records of the Royal Society of London, 43(1):31–51, January 1989. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531716.

Charlton:1989:EMT

[1039] T. M. Charlton. An extension of Maxwell’s theory of pin-jointed frameworks by M. W. Crofton, F.R.S. Notes and Records of the Royal Society of London, 43(1):53–56, January 1989. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531717.

Landes:1989:WLH

[1040] D. S. Landes. The Wilkins Lecture, 1988 hand and mind in time measurement: the contributions of art and science. Notes and Records of the Royal Society of London, 43(1):57–69, January 1989. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531718.

Whitrow:1989:NRH

[1041] G. J. Whitrow. Newton’s role in the history of mathematics. Notes and Records of the Royal Society of London, 43(1):71–92, January 1989. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531719.

Anonymous:1989:BMa (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531720.

Anonymous:1989:FMa

[1043] Anonymous. Front matter. Notes and Records of the Royal Society of London, 43(1):??, January 1989. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531713.

Anonymous:1989:FMb

[1044] Anonymous. Front matter. Notes and Records of the Royal Society of London, 43(2):93, July 1989. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531374.

Porter:1989:WI

[1045] Sir George Porter, P.R.S. Welcome and introduction. Notes and Records of the Royal Society of London, 43(2):95–96, July 1989. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531375.

Hall:1989:CR

[1046] A. R. Hall, F.B.A. Chairman’s remarks. Notes and Records of the Royal Society of London, 43(2):97–98, July 1989. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531376.

Hunter:1989:CPN

[1047] M. C. W. Hunter. The Crown, the public and the new science, 1689– 1702. Notes and Records of the Royal Society of London, 43(2):99–116, July 1989. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531377.

Bruijn:1989:WIH

[1048] J. R. Bruijn. William III and his two navies. Notes and Records of the Royal Society of London, 43(2):117–132, July 1989. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531378.

Dunn:1989:BEA

[1049] J. M. Dunn. ‘Bright enough for all our purposes’: John Locke’s conception of a civilized society. Notes and Records of the Royal Society of London, 43(2):133–153, July 1989. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531379.

Leopold:1989:CBN

[1050] J. H. Leopold. Clockmaking in Britain and the Netherlands. Notes and Records of the Royal Society of London, 43(2):155–165, July 1989. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531380.

Schaffer:1989:GRM

[1051] Simon Schaffer. The glorious revolution and medicine in Britain and the Netherlands. Notes and Records of the Royal Society of London, 43 (2):167–190, July 1989. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531381.

Palm:1989:LOD

[1052] L. C. Palm. Leeuwenhoek and other Dutch correspondents of the Royal Society. Notes and Records of the Royal Society of London, 43(2):191– 207, July 1989. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531382.

Snelders:1989:CHN

[1053] H. A. M. Snelders. Christiaan Huygens and Newton’s theory of gravitation. Notes and Records of the Royal Society of London, 43(2):209–222, July 1989. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531383. With a discus-

sion by B. Hughes and a reply by the author.

Shapiro:1989:HTL

[1054] A. E. Shapiro. Huygens’ Traité de la lumière and Newton’s Opticks: pursuing and eschewing hypotheses. Notes and Records of the Royal Society of London, 43(2):223–247, July 1989. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/ stable/531384. With a discussion by J. H. van der Waals and a reply by the author.

Hall:1989:LLA

[1055] A. R. Hall, F.B.A. The Leeuwenhoek Lecture, 1988, Antoni Van Leeuwenhoek 1632–1723. Notes and Records of the Royal Society of London, 43(2):249–273, July 1989. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531385.

Rowlinson:1989:BRS

[1056] J. S. Rowlinson. Book review — Science, medicine and dissent: Joseph Priestley (1733–1804) (papers celebrating the 250th anniversary of the birth of Joseph Priestley together with a catalogue of an exhibition held at the Royal Society and the Wellcome Institute for the History of Medicine), edited by R. G. W. Anderson & C. Lawrence (pp. ix + 105). Published by Wellcome Trust and Science Museum, London, 1987, £9.95. Notes and Records of the Royal Society of London, 43(2):275–276, July 1989. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531386.

Anonymous:1989:INR

[1057] Anonymous. Index to Notes and Records, volume 43, 1989. Notes and Records of the Royal Society of London, 43(2):277–279, July 1989. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531373.

Anonymous:1989:BMb

[1058] Anonymous. Back matter. Notes and Records of the Royal Society of London, 43(2):??, July 1989. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531387.

Anonymous:1990:FMa

[1059] Anonymous. Front matter. Notes and Records of the Royal Society of London, 44(1):1, January 1990. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531581.

Tozer:1990:DAS

[1060] E. T. Tozer. Discovery of an ammonoid specimen described by Robert Hooke. Notes and Records of the Royal Society of London, 44(1):3–12, January 1990. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531582.

Hall:1990:BFD

[1061] A. Rupert Hall, F.B.A. Beyond the fringe: diffraction as seen by Grimaldi, Fabri, Hooke and Newton. Notes and Records of the Royal Society of London, 44(1):13–23, January 1990. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531583.

Appleby:1990:DDD

[1062] John H. Appleby. Daniel Dumaresq, D.D., F.R.S. (1712–1805) as a promoter of Anglo–Russian science and culture. Notes and Records of the Royal Society of London, 44(1):25–50, January 1990. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531584.

Griffin:1990:BRM

[1063] Nicholas Griffin and Albert C. Lewis. Bertrand Russell’s mathematical education. Notes and Records of the Royal Society of London, 44(1): 51–71, January 1990. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/531585.

Jeffreys:1990:JAG

[1064] Bertha Swirles Jeffreys. John Arthur Gaunt (1904–1944). Notes and Records of the Royal Society of London, 44(1):73–79, January 1990. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531586.

Beer:1990:TTR

[1065] Gillian Beer. Translation or transformation? The relations of literature and science. Notes and Records of the Royal Society of London, 44(1): 81–99, January 1990. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/531587.

Hall:1990:ERW

[1066] A. Rupert Hall. Essay reviews: A window on the East [Science and civilisation in China by Joseph Needham and others. Cambridge University Press]. Notes and Records of the Royal Society of London, 44(1):101–110, January 1990. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531588.

Whiteside:1990:LNL

[1067] D. T. Whiteside. The latest on Newton …. Notes and Records of the Royal Society of London, 44(1):111–117, January 1990. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531589.

Allen:1990:BFF

[1068] David E. Allen. Banks in full flower [Harold B. Carter, Sir Joseph Banks 1743–1820. British Museum (Natural History), London, 1988. Pp. 671.

£45.00. ISBN 0-565-00993-1. Harold B. Carter, Sir Joseph Banks (1743– 1820): a guide to biographical and bibliographical sources. St Paul’s Bibliographies, Winchester, 1987. Pp. 328. £45.00. ISBN 0-906795-45-1. Judith Diment and C.J. Humphries (ed.), Banks’ Florilegium: a publication in 34 parts of 738 copperplate engravings of plants collected on Captain Cook’s first voyage. Alecto Historical Editions, London, 1980–88. About £147,000. Patrick O’Brian, Joseph Banks: a life. Collins Harvill, London, 1988. Pp. 328. £15.00. ISBN 0-00-217350-6. £6.95 paperback. ISBN 0-00-272340-9]. Notes and Records of the Royal Society of London, 44 (1):119–124, January 1990. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531590.

Bowler:1990:ERC

[1069] Peter J. Bowler. Essay reviews — Charles Darwin: Imprimatur! [Charles Darwin’s Beagle diary, ed. Richard Darwin Keynes. Cambridge University Press, 1988. Pp. xxix + 464. £35. ISBN 0-521-23503-0. Charles Darwin’s notebooks, 1836–1844, ed. Paul H. Barrett et al. Cambridge University Press, 1987. Pp. viii + 747. £65. ISBN 0-521-35055-7. A calendar of the correspondence of Charles Darwin, 1821–1882, ed. Frederick Burkhardt & Sydney Smith. New York and London: Garland Publishing Inc., 1985. Pp. 690. £37.50. ISBN 0-521-35055-7. The correspondence of Charles Darwin, ed. Frederick Burkhardt & Sydney Smith. Cambridge: Cambridge University Press: Volume 1 (1821–1836), 1985. Pp. xxix + 702. £32.50. ISBN 0-521-25587-2; Volume 2 (1837–1843), 1986. Pp. xxv + 603. £32.50. ISBN 0-521-25588-0; Volume 3 (1844–1846), 1987. Pp. xxxii + 523. £32.50. ISBN 0-521-25589-9]. Notes and Records of the Royal Society of London, 44(1):125–131, January 1990. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531591.

James:1990:BRA

[1070] Frank A. J. L. James. Book review: All scientists now: the Royal Society in the nineteenth century, by Marie Boas Hall. Notes and Records of the Royal Society of London, 44(1):133–134, January 1990. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531592.

Brooke:1990:BRC

[1071] John Hedley Brooke. Book reviews: Colin A. Russell, Lancastrian chemist: the early years of Sir Edward Frankland. Open University Press, Milton Keynes and Philadelphia, 1986. Pp. ix + 187. £35.00. ISBN 0-335-15175-2. Notes and Records of the Royal Society of London, 44(1):

135–136, January 1990. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531593.

Burgen:1990:HFP

[1072] Arnold Burgen. Howard Florey — penicillin and after. Notes and Records of the Royal Society of London, 44(1):136–137, January 1990. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531594.

McCrea:1990:BRC

[1073] William McCrea. Book reviews: S. Chandrasekhar. Selected papers. Volume 1: Stellar structure and stellar atmospheres. Pp. xiv + 516; Volume 2: Radiative transfer and negative ion of hydrogen, Pp. xvi + 622. University of Chicago Press, 1989. Notes and Records of the Royal Society of London, 44(1):137–139, January 1990. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531595.

Anonymous:1990:BMa

[1074] Anonymous. Back matter. Notes and Records of the Royal Society of London, 44(1):141–142, January 1990. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531596.

Hall:1990:ACC

[1075] A. Rupert Hall and Marie Boas Hall. Additions and corrections to the correspondence of Henry Oldenburg. Notes and Records of the Royal Society of London, 44(2):143–150, July 1990. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531604.

Thomas:1990:JRF

[1076] David J. Thomas and Judith M. Smith. Joseph Raphson, F.R.S. Notes and Records of the Royal Society of London, 44(2):151–167, July 1990. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531605.

Andrews:1990:RMD

[1077] Jonathan Andrews. A respectable mad — doctor? Dr Richard Hale, F.R.S. (1670–1728). Notes and Records of the Royal Society of London, 44(2):169–204, July 1990. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531606.

Wilkes:1990:HPB

[1078] M. V. Wilkes, F.R.S. Herschel, Peacock, Babbage and the development of the Cambridge curriculum. Notes and Records of the Royal Society of London, 44(2):205–219, July 1990. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531607.

Tansey:1990:GES

[1079] E. M. Tansey. George Eliot’s support for physiology: the George Henry Lewes trust 1879–1939. Notes and Records of the Royal Society of London, 44(2):221–240, July 1990. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531608.

Smith:1990:MCA

[1080] Meg Weston Smith. E. A. Milne and the creation of air defence: some letters from an unprincipled brigand, 1916–1919. Notes and Records of the Royal Society of London, 44(2):241–255, July 1990. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531609.

Lipson:1990:RDI

[1081] H. Lipson, F.R.S. Reminiscences and discoveries, the introduction of Fourier methods into crystal-structure determination. Notes and Records of the Royal Society of London, 44(2):257–264, July 1990. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531610.

Hall:1990:ERE

[1082] Marie Boas Hall. Essay reviews: The early years of the Royal Society [Michael Hunter, Establishing the new science: the experience of the early Royal Society. Boydell Press, Woodbridge, Suffolk, 1989. Pp. xiv + 382. £45. ISBN 0-85115-506-5]. Notes and Records of the Royal Society of London, 44(2):265–268, July 1990. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531611.

Wright:1990:BRB

[1083] Esmond Wright. Book reviews: Benjamin Franklin after 200 years [The papers of Benjamin Franklin, ed. Claude A. Lopez. New Haven: Yale University Press, Volume 27, 1988. Pp. 727. £45. ISBN 0-300-04177-2]. Notes and Records of the Royal Society of London, 44(2):269–279, July 1990. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531612.

Grattan-Guinness:1990:BRA

[1084] Ivor Grattan-Guinness. Bertrand Russell (1872–1970) after twenty years. Notes and Records of the Royal Society of London, 44(2):280–306, July 1990. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531613.

Hall:1990:BRP

[1085] A. Rupert Hall. Book reviews: The preliminary manuscripts for Isaac Newton’s 1687 1684–1685. Facsimiles of the original autographs, now in Cambridge University, with an introduction by D.T. Whiteside. Cambridge University Press, 1989. Pp. xxi + 246. £60. ISBN 0-521-33499-3. Notes and Records of the Royal Society of London, 44(2):307–309, July 1990. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531614.

Smithies:1990:BRS

[1086] Frank Smithies. Book reviews: Stuart Hollingdale, Makers of mathematics [Penguin Books, 1989. Pp. xiv + 433. £7.99; US$10.95 ISBN 0-14-022732-6]. Notes and Records of the Royal Society of London, 44 (2):309–311, July 1990. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531615.

Russell:1990:BRD

[1087] Colin A. Russell. Book reviews: Donald S. L. Cardwell, James Joule: a biography. Manchester University Press, Manchester and New York, 1989. Pp. x + 333. £35.00. ISBN 0-7190-3025-0. Notes and Records of the Royal Society of London, 44(2):311–313, July 1990. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531616.

Bondi:1990:BRW

[1088] Hermann Bondi. Book reviews: Walter Moore, Schrödinger: life and thought. Cambridge University Press, 1989. Pp. 513, £25.00. ISBN 0-521-35434-X. Notes and Records of the Royal Society of London, 44(2): 313–315, July 1990. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/531617.

Porter:1990:BPC

[1089] Roy Porter. Books and papers concerned with the history of the Royal Society (1988). Notes and Records of the Royal Society of London, 44 (2):317–318, July 1990. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531618.

Anonymous:1990:BMb

[1090] Anonymous. Back matter. Notes and Records of the Royal Society of London, 44(2):??, July 1990. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531619.

Anonymous:1990:FMb

[1091] Anonymous. Front matter. Notes and Records of the Royal Society of London, 44(2):??, July 1990. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531603.

Morris:1991:IJD

[1092] G. C. R. Morris. On the identity of Jaques Du Moulin, F.R.S. 1667. Notes and Records of the Royal Society of London, 45(1):1–10, January 1991. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531519.

Whiteside:1991:PP

[1093] D. T. Whiteside. The prehistory of the Principia from 1664 to 1686. Notes and Records of the Royal Society of London, 45(1):11–61, January 1991. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531520.

Hamilton:1991:GBS

[1094] Beryl M. Hamilton. ‘A geological blunder’, 1893: A scientific storm in a journalistic teacup. Notes and Records of the Royal Society of London, 45(1):63–77, January 1991. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531521.

Hawthorne:1991:EHA

[1095] Sir William Hawthorne, F.R.S. The early history of the aircraft gas turbine in Britain. Notes and Records of the Royal Society of London, 45 (1):79–108, January 1991. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531522.

Herzberg:1991:ERL

[1096] Gerhard Herzberg. Essay review: A leader of Indian science [G. Venkataraman, Journey into light: life and science V. Raman. Indian Academy of Sciences in cooperation with Indian National Science Academy, 1988. Pp. xiv + 570. ISBN 81-85324-00-X]. Notes and Records of the Royal Society of London, 45(1):109–113, January 1991. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531523.

Browne:1991:BRW

[1097] Janet Browne. Book reviews: William LeFanu, Nehemiah Grew, M.D., F.R.S.: a study and bibliography of his writings [Winchester: St Paul’s Bibliographies, 1990. Pp. xvii + 182, £42.00. ISBN 0-906795-43-5]. Notes and Records of the Royal Society of London, 45(1):115–116, January 1991. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531524.

Lawrence:1991:BRE

[1098] Christopher Lawrence. Book reviews: Ernest Heberden, William Heberden, physician of the Age of Reason [London: Royal Society of Medicine, 1990. Pp. 246, £12.95 (hardback), £7.95 (softback). ISBN 1-85315-116-5]. Notes and Records of the Royal Society of London, 45(1):116–118, January 1991. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531525.

McCrea:1991:BRD

[1099] William McCrea. Book reviews: Derek Howse, Nevil Maskelyne: the seaman’s astronomer. Cambridge: Cambridge University Press, 1989. Pp. xiv + 280, £40.00. ISBN 0-521-36261-X. Notes and Records of the Royal Society of London, 45(1):118–120, January 1991. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531526.

Fisher:1991:BRW

[1100] Susanna Fisher. Book reviews: William Huddart, Unpathed waters: account of the life and times of Joseph Huddart, F.R.S. [London: Quiller Press, 1989. Pp. 200, £12.95. ISBN 1-870948-36-X]. Notes and Records of the Royal Society of London, 45(1):120–122, January 1991. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531527.

James:1991:BRE

[1101] Frank A. J. L. James. Book reviews: Energy and Empire: a biographical study of Lord Kelvin [Crosbie Smith & M. Norton Wise, Cambridge: Cambridge University Press, 1989. Pp. xxvi + 866, £60.00. ISBN 0-521-26173-2]. Notes and Records of the Royal Society of London, 45(1): 122–124, January 1991. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531528.

McLaren:1991:BRW

[1102] Digby J. McLaren. Book reviews: E. W. R. Steacie and science in Canada [M. Christine King, Toronto: Toronto University Press, 1989. Pp. xii + 243, $35.00, £24.00. ISBN 0-8020-2667-2]. Notes and Records of the Royal Society of London, 45(1):124–126, January 1991. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531529.

Dickson:1991:BRW

[1103] J. W. Dickson. Book reviews: William Waugh, John Charnley, the man and the hip. Springer-Verlag, London, Berlin, Heidelberg, New York, 1990. Pp. xvi + 268. £24.50. ISBN 3-540-19587-4. Notes and Records of the Royal Society of London, 45(1):127, January 1991. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531530.

Anonymous:1991:BMa

[1104] Anonymous. Back matter. Notes and Records of the Royal Society of London, 45(1):??, January 1991. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531531.

Anonymous:1991:FMa

[1105] Anonymous. Front matter. Notes and Records of the Royal Society of London, 45(1):??, January 1991. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531518.

Cook:1991:EHN

[1106] Sir Alan Cook, F.R.S. Edmond Halley and Newton’s Principia. Notes and Records of the Royal Society of London, 45(2):129–138, July 1991. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531693.

Appleby:1991:RDF

[1107] John H. Appleby. Robert Dingley, F.R.S. (1710–1781), merchant, architect and pioneering philanthropist. Notes and Records of the Royal Society of London, 45(2):139–154, July 1991. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531694.

Mason:1991:JHM

[1108] Stephen F. Mason, F.R.S. Jean Hyacinthe De Magellan, F.R.S., and the chemical revolution of the eighteenth century. Notes and Records of the Royal Society of London, 45(2):155–164, July 1991. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531695.

Turvey:1991:SJH

[1109] Peter J. Turvey. Sir John Herschel and the abandonment of Charles Babbage’s Difference Engine No. 1. Notes and Records of the Royal Society of London, 45(2):165–176, July 1991. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531696.

Sheets-Pyenson:1991:PBS

[1110] Susan Sheets-Pyenson. ‘Pearls before swine’: Sir William Dawson’s Bakerian Lecture of 1870. Notes and Records of the Royal Society of London, 45(2):177–191, July 1991. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531697.

Tweedale:1991:NDE

[1111] Geoffrey Tweedale and Timothy Procter. New documentary evidence on the career of Sir William Boyd Dawkins, F.R.S. Notes and Records of the Royal Society of London, 45(2):193–200, July 1991. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531698.

Mason:1991:HAA

[1112] Joan Mason. Hertha Ayrton (1854–1923) and the admission of women to the Royal Society of London. Notes and Records of the Royal Society of London, 45(2):201–220, July 1991. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531699.

Devons:1991:RSH

[1113] S. Devons, F.R.S. Rutherford and the science of his day. Notes and Records of the Royal Society of London, 45(2):221–242, July 1991. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

URL https://royalsocietypublishing.org/journal/rsnrcontent/
45/2/221;   http://www.jstor.org/stable/531700.  Rutherford

memorial lecture, 1989, delivered in June 1989 at Canberra, Sydney, and Melbourne, Australia, and at Auckland, New Zealand.

Thomas:1991:RDB

[1114] J. M. Thomas, F.R.S. Reminiscences and discoveries: Bragg reflections. Notes and Records of the Royal Society of London, 45(2):243–252, July 1991. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531701.

Anonymous:1991:YMP

[1115] Anonymous. The young Mozart in the Philosophical Transactions. Notes and Records of the Royal Society of London, 45(2):253–256, July 1991. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531702.

Mason:1991:ERH

[1116] Stephen F. Mason. Essay reviews: Hooke redivived [Robert Hooke: new studies. Edited by Michael Hunter & Simon Schaffer. Woodbridge, Suffolk: The Boydell Press, 1989. Pp. x + 310, £39.50. ISBN 0-85115-523-5]. Notes and Records of the Royal Society of London, 45(2):257–261, July 1991. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531703.

Bowler:1991:ERS

[1117] Peter J. Bowler. Essay reviews: The species theory evolves [The correspondence of Charles Darwin. Volume IV (1847–1850) (ed. Frederick Burkhardt & Sydney Smith) Cambridge University Press, 1988. Pp. xxxiii + 711, £32.50. ISBN 0-521-25590-2. Volume V (1851–1855). 1989. Pp. xxix + 705. £32.50. ISBN 0-521-25591-0. Volume VI (1856–1857). 1990. Pp. xxix + 673. £35.00 ISBN 0-521-25586-4]. Notes and Records of the Royal Society of London, 45(2):262–265, July 1991. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531704.

James:1991:ERM

[1118] Frank A. J. L. James. Essay reviews: Maxwell’s early career [P. M. Harman, The scientific letters and papers of James Clerk Maxwell. Volume 1. 1846–1862. Cambridge University Press, 1990. Pp. xxvii + 748, £125.00. ISBN 0-521-25625-9]. Notes and Records of the Royal Society of London, 45(2):266–270, July 1991. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531705.

Grattan-Guinness:1991:ERF

[1119] Ivor Grattan-Guinness. Essay reviews: At the feet of Whitehead [Victor Lowe, Alfred North Whitehead. The man and his work. Volume I (1861– 1910). Baltimore and London: The Johns Hopkins University Press, 1986. Pp. xi + 352, £27.00. ISBN 0-8018-2488-5 Victor Lowe with L.B. McHenry, Volume II (1910–1947), ed. J. B. Schneewind, 1990. Pp. xi + 389. ISBN 0-8018-3960-2]. Notes and Records of the Royal Society of London, 45(2):271–276, July 1991. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531706.

Hollingdale:1991:BRB

[1120] Stuart Hollingdale. Book reviews: Before Newton: the life and times of Isaac Barrow. Edited by Mordechai Feingold. Cambridge: Cambridge University Press, 1990. Pp. xi + 380, £35.00. ISBN 0-521-30694-9. Notes and Records of the Royal Society of London, 45(2):277–279, July 1991. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531707.

Green:1991:BRP

[1121] Leslie Green. Book reviews: Peter Rowlands, Oliver Lodge and the Liverpool Physical Society. Liverpool: Liverpool University Press, 1990. Pp. 336, £15.00. ISBN 0-85323-027-7. Notes and Records of the Royal Society of London, 45(2):279–281, July 1991. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531708.

Mason:1991:BRR

[1122] Stephen F. Mason. Book reviews — Robert Robinson: chemist extraordinary [Trevor I. Williams, Robert Robinson: chemist extraordinary. Oxford: Clarendon Press, 1990. Pp. viii + 201, £25.00. ISBN 0-19-858180-7]. Notes and Records of the Royal Society of London, 45(2):281–284, July 1991. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531709.

Postgate:1991:BRW

[1123] John Postgate. Book reviews: What mad pursuit: a personal view of scientific discovery [Francis Crick, What mad pursuit: a personal view of scientific discovery. London: Penguin Books, 1990. Pp. 182, £4.99. ISBN 0-14-011973-6]. Notes and Records of the Royal Society of London, 45(2):284–286, July 1991. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531710.

Anonymous:1991:BMb

[1124] Anonymous. Back matter. Notes and Records of the Royal Society of London, 45(2):??, July 1991. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531711.

Anonymous:1991:FMb

[1125] Anonymous. Front matter. Notes and Records of the Royal Society of London, 45(2):??, July 1991. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531692.

Mason:1992:BJW

[1126] Stephen F. Mason, F.R.S. Bishop John Wilkins, F.R.S. (1614–72): Analogies of thought-style in the Protestant reformation and early modern science. Notes and Records of the Royal Society of London, 46(1): 1–21, January 1992. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/531438.

Hall:1992:FSF

[1127] Marie Boas Hall. Frederick Slare, F.R.S. (1648–1727). Notes and Records of the Royal Society of London, 46(1):23–41, January 1992. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531439.

Bektas:1992:CME

[1128] M. Yakup Bektas and Maurice Crosland. The Copley Medal: The establishment of a reward system in the Royal Society, 1731–1839. Notes and Records of the Royal Society of London, 46(1):43–76, January 1992. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531440.

Kemp:1992:TTN

[1129] Martin Kemp, F.B.A. True to their natures: Sir Soshua Reynolds and Dr William Hunter at the Royal Academy of Arts. Notes and Records of the Royal Society of London, 46(1):77–88, January 1992. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531441.

Smith:1992:SGE

[1130] James R. Smith. Sir George Everest, F.R.S. (1790–1866). Notes and Records of the Royal Society of London, 46(1):89–102, January 1992. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531442.

Chapman:1992:GBA

[1131] Allan Chapman. George Biddell Airy, F.R.S. (1801–1892): a centenary commemoration. Notes and Records of the Royal Society of London, 46 (1):103–110, January 1992. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531443.

Meadows:1992:AGT

[1132] A. J. Meadows. Astronomy and geology, terrible muses! Tennyson and 19th-century science. Notes and Records of the Royal Society of London, 46(1):111–118, January 1992. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531444.

Paton:1992:FRS

[1133] Sir William Paton, F.R.S. The Fellowship of the Royal Society and its work. Notes and Records of the Royal Society of London, 46(1):119–128, January 1992. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531445.

Allibone:1992:PRS

[1134] T. E. Allibone, F.Eng., F.R.S. Philately and the Royal Society. Notes and Records of the Royal Society of London, 46(1):129–154, January 1992. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531446.

Atiyah:1992:APS

[1135] Sir Michael Atiyah. Address of the President, Sir Michael Atiyah, given at the Anniversary Meeting on 29 November 1991. Notes and Records of the Royal Society of London, 46(1):155–169, January 1992. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531447.

Moon:1992:RDL

[1136] P. B. Moon, F.R.S. Reminiscences and discoveries the (London) physics club, 1928–1953. Notes and Records of the Royal Society of London, 46 (1):171–174, January 1992. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531448.

Fellgett:1992:RCL

[1137] Peter Fellgett, F.R.S. A reminiscence of the Cavendish laboratory. Notes and Records of the Royal Society of London, 46(1):175–176, January 1992. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531449.

Mason:1992:BRW

[1138] Joan Mason. Book review: Women in science: Breaking out of the circle [Marilyn Bailey Ogilvie, Women in science, antiquity through the nineteenth century. Cambridge, Massachusetts: MIT Press, 1986. Pp. xi + 254, £24.75. ISBN 0-262-15031-X. Margaret Alic, Hypatia’s heritage: a history of women in science from antiquity to the late nineteenth century. London: The Women’s Press, 1986. Pp. ix + 230, £4.95. ISBN 0-7043-3954-4. Londa Schiebinger, The mind has no sex? Women in the origins of modem science. Cambridge, Massachusetts: Harvard University Press, 1989. Pp. xi + 355, £23.50. ISBN 0-674-57623-3. Patricia Phillips, The scientific lady: a social history of woman’s scientific interests 1520–1918. London: Weidenfeld and Nicolson, 1990. Pp. xiii + 279, £25.00. ISBN 0-297-82043-5. Uneasy careers and intimate lives: women in science, 1789–1979. Edited by Pnina G. Abir-Am & Dorinda Outram. New Brunswick: Rutgers University Press, 1987. Pp. xiii + 365, £11.00. ISBN 0-8135-1255-7. Women of science: righting the record. Edited by G. Kass-Simon & Patricia Fames. Bloomington, Indiana: Indiana University Press, 1990. Pp. xvi + 398, $39.95. ISBN 0-253-33264-8]. Notes and Records of the Royal Society of London, 46(1):177–182, January 1992. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531450.

Hoyle:1992:AD

[1139] F. Hoyle. The achievement of Dirac. Notes and Records of the Royal Society of London, 46(1):183–187, January 1992. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531451.

Jones:1992:BRL

[1140] R. V. Jones. Book reviews — the Leeuwenhoek legacy [Brian J. Ford, The Leeuwenhoek legacy. London: Farrand Press, 1991. Pp. 185, £25.00. ISBN 1-85083-016-9]. Notes and Records of the Royal Society of London, 46 (1):189–190, January 1992. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531452.

Hall:1992:BRR

[1141] A. Rupert Hall. Book reviews: Richard Rawlinson: a tercentenary memorial. Edited by Georgian R. Tashjian, David R. Tashjian and Brian J. Enright. New Issues Press, Western Michigan University, 1990. Pp. xviii + 221. ISBN 0-932826-23-7. Notes and Records of the Royal Society of London, 46(1):190–191, January 1992. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531453.

Morrell:1992:BRR

[1142] Jack Morrell. Book reviews: Robert A. Stafford, Scientist of empire: Sir Roderick Murchison, scientific exploration and Victorian imperialism [Cambridge University Press, 1989. Pp. xi + 293, £30.00. ISBN 0-521-33537-X]. Notes and Records of the Royal Society of London, 46(1): 191–193, January 1992. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531454.

Postgate:1992:BRE

[1143] John Postgate. Book reviews: Evolutionary studies, a centenary celebration of the life of Julian Huxley. Edited by Milo Keynes & G. Ainsworth Harrison. London: Macmillan, 1989. Pp. 256, £63.00. ISBN 0-333-45723-4. Notes and Records of the Royal Society of London, 46(1):193–196, January 1992. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531455.

Woolfson:1992:BRS

[1144] Michael M. Woolfson. Book reviews: Selections and reflections: the legacy of Sir Lawrence Bragg. Edited by John M. Thomas and Sir David Phillips. Northwood, Middlesex: Science Reviews Ltd, 1990. Pp. 308, £27.00. ISBN 0-905927-43-5. Notes and Records of the Royal Society of London, 46(1):196–198, January 1992. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531456.

Clarke:1992:BRV

[1145] Patricia H. Clarke, F.R.S. Book reviews: A very decided preference [Jean Medawar, A very decided preference: Life with Peter Medawar. Oxford University Press, 1990. Pp. 256, £15.00. ISBN 0-19-217779-6]. Notes and Records of the Royal Society of London, 46(1):198–200, January 1992. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531457.

Brown:1992:BRB

[1146] R. Hanbury Brown. Book reviews: Bernard Lovell, Astronomer by Chance. London: Macmillan, 1991. Pp. 380, £18.99. ISBN 0-333-55195-8. Notes and Records of the Royal Society of London, 46(1):200–202, January 1992. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531458.

Anonymous:1992:BMa (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531459.

Anonymous:1992:FMa

[1148] Anonymous. Front matter. Notes and Records of the Royal Society of London, 46(1):??, January 1992. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531437.

Kirsanov:1992:ECR

[1149] V. S. Kirsanov. The earliest copy in Russia of Newton’s Principia: is it David Gregory’s annotated copy? Notes and Records of the Royal Society of London, 46(2):203–218, July 1992. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531634.

Massarella:1992:PHZ

[1150] Derek Massarella. Philip Henry Zollman, the Royal Society’s First Assistant Secretary for Foreign Correspondence. Notes and Records of the Royal Society of London, 46(2):219–234, July 1992. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531635.

Johnson:1992:BRF

[1151] W. Johnson, F.Eng., F.R.S. Benjamin Robins, F.R.S. (1707–1751): new details of his life. Notes and Records of the Royal Society of London, 46(2):235–252, July 1992. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531636.

King-Hele:1992:SS

[1152] D. G. King-Hele, F.R.S. Shelley and science. Notes and Records of the Royal Society of London, 46(2):253–265, July 1992. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/531637.

Synge:1992:JDB

[1153] A. Synge. J. D. Bernal, F.R.S.: family, school and university. Notes and Records of the Royal Society of London, 46(2):267–278, July 1992. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531638.

Mason:1992:AFW

[1154] Joan Mason. The admission of the first women to the Royal Society of London. Notes and Records of the Royal Society of London, 46(2):279– 300, July 1992. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531639.

Swirles:1992:RDH

[1155] Bertha Swirles, (Lady Jeffreys). Reminiscences and discoveries: Harold Jeffreys from 1891 to 1940. Notes and Records of the Royal Society of London, 46(2):301–308, July 1992. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531640.

Synge:1992:HRN

[1156] R. L. M. Synge, F.R.S. How the Robinsons nearly invented partition chromatography in 1934. Notes and Records of the Royal Society of London, 46(2):309–311, July 1992. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531641.

Mason:1992:ERS

[1157] Stephen F. Mason. Essay reviews: The spring-tide of experimental philosophy [Marie Boas Hall, Promoting Experimental Learning: Experiment and the Royal Society, 1660–1727. Cambridge University Press, 1991. Pp. Xiii + 207, £35.00 ISBN 0-521-40503-3]. Notes and Records of the Royal Society of London, 46(2):313–316, July 1992. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531642.

Smith:1992:GMF

[1158] Crosbie Smith. The genius of Michael Faraday. Notes and Records of the Royal Society of London, 46(2):317–323, July 1992. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531643.

Hall:1992:ROP

[1159] A. R. Hall and M. B. Hall. Richard Ollard, Pepys: a biography. London: Sinclair-Stevenson, 1991. Pp. 411, £25. ISBN 1-85619-066-8. Notes and Records of the Royal Society of London, 46(2):325–326, July 1992. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531644.

Jacyna:1992:BRS

[1160] L. S. Jacyna. Book reviews: Sir James Paget: the rise of clinical surgery [Shirley Roberts, Sir James Paget: the rise of clinical surgery. Royal Society of Medicine, 1990. Pp. 223. £12.95 (hardback), £7.95 (softback).

ISBN 0-905958-91-8]. Notes and Records of the Royal Society of London, 46(2):326–327, July 1992. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531645.

Postgate:1992:BRD

[1161] John Postgate. Book reviews: Darwin [Adrian Desmond & James Moore, Darwin. London: Michael Joseph, 1991. Pp. 808, £20.00. ISBN 0-3403-3]. Notes and Records of the Royal Society of London, 46(2):327–328, July 1992. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531646.

Teich:1992:BRE

[1162] Mikuláš Teich. Book reviews: Essays on science: felicitation volume in honour of Dr Joseph Needham. Edited by Hakim Mohammed Said. Karachi: Hamdard Foundation Pakistan, 1990. Pp. xiii + 267. Pak.Rs. 400.00, $40.00, £20.00. Notes and Records of the Royal Society of London, 46(2):329–330, July 1992. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531647.

Edwards:1992:BRS

[1163] Sam Edwards. Book review: Sir Charles Frank O.B.E., F.R.S.; an eightieth birthday tribute [Edited by R. G. Chambers, J. E. Enderby, A. Keller, A. R. Lang & J. W. Steeds. Bristol: Adam Hilger, 1991. Pp. xii + 448, £27.50. ISBN 0-7503-0100-7]. Notes and Records of the Royal Society of London, 46(2):330–331, July 1992. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531648.

Lovell:1992:BRR

[1164] Bernard Lovell. Book reviews: R. Hanbury Brown, Boffin: a personal story of the early days of radar, radio astronomy and quantum optics [Bristol: Adam Hilger, 1991, £17.50. ISBN 0-7503-0130-9]. Notes and Records of the Royal Society of London, 46(2):332–334, July 1992. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531649.

Clark:1992:HRS

[1165] Alan J. Clark. The history of the Royal Society: a chronological checklist of books in English. Notes and Records of the Royal Society of London, 46(2):335–345, July 1992. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531650.

Anonymous:1992:BMb

[1166] Anonymous. Back matter. Notes and Records of the Royal Society of London, 46(2):??, July 1992. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531651.

Anonymous:1992:FMb

[1167] Anonymous. Front matter. Notes and Records of the Royal Society of London, 46(2):??, July 1992. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531633.

Cohen:1993:NDR

[1168] I. Bernard Cohen. Newton’s description of the reflecting telescope. Notes and Records of the Royal Society of London, 47(1):1–9, July 1993. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531389.

Meynell:1993:AMF

[1169] Guy Meynell. André De Monceaux, F.R.S. 1670. Notes and Records of the Royal Society of London, 47(1):11–15, July 1993. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/531390.

Bryden:1993:MIN

[1170] D. J. Bryden. Magnetic inclinatory needles: approved by the Royal Society? Notes and Records of the Royal Society of London, 47(1):17– 31, July 1993. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531391.

Kochhar:1993:ACF

[1171] R. K. Kochhar. Ardaseer Cursetjee (1808–1877), the first Indian Fellow of the Royal Society of London. Notes and Records of the Royal Society of London, 47(1):33–47, July 1993. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531392.

Warwick:1993:FTF

[1172] Andrew Warwick. Frequency, theorem and formula: remembering Joseph Larmor in electromagnetic theory. Notes and Records of the Royal Society of London, 47(1):49–60, July 1993. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531393.

McCrea:1993:SRH

[1173] Sir William McCrea, F.R.S. Sir Ralph Howard Fowler, 1889–1944: a centenary lecture. Notes and Records of the Royal Society of London, 47(1):61–78, July 1993. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531394.

Gowing:1993:JCA

[1174] Margaret Gowing, F.R.S., F.B.A. James Chadwick and the atomic bomb. Notes and Records of the Royal Society of London, 47(1):79–92, July 1993. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531395.

Johnston:1993:JDB

[1175] Roy H. W. Johnston. J. D. Bernal, F.R.S.: some Irish influences. Notes and Records of the Royal Society of London, 47(1):93–103, July 1993. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531396.

Grattan-Guinness:1993:SMO

[1176] Ivor Grattan-Guinness. The Sylvester Medal: origins, and recipients 1901–1949. Notes and Records of the Royal Society of London, 47(1): 105–108, July 1993. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/531397.

Atiyah:1993:APS

[1177] Michael Atiyah. Address of the President, Sir Michael Atiyah, O. M., given at the Anniversary Meeting on 30 November 1992. Notes and Records of the Royal Society of London, 47(1):109–118, July 1993. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531398.

Lovell:1993:RDB

[1178] Sir Bernard Lovell, F.R.S. Reminiscences and discoveries: The Blackett– Eckersley–Lovell correspondence of World War II and the origin of Jodrell Bank. Notes and Records of the Royal Society of London, 47(1):119– 131, July 1993. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531399.

Huxley:1993:ERL

[1179] Andrew Huxley. Essay review: Living with electrical impulses [Alan Hodgkin, Chance and Design. Reminiscences of Science in Peace and War. Cambridge University Press, 1992. Pp. 412, £40.00. ISBN 0-521-40099-6]. Notes and Records of the Royal Society of London, 47(1):133– 139, July 1993. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531400.

Hall:1993:BRD

[1180] Marie Boas Hall. Book reviews — dispenser of restoration medicine [J. Trevor Hughes, Thomas Willis 1621–1675: His Life and Work. London, Royal Society of Medicine Services, 1991. Pp. xvi + 151, £12.95 (hdbk), £7.95 (pbk). ISBN 1-85315-162-9]. Notes and Records of the Royal Society of London, 47(1):141–142, July 1993. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531401.

Brooke:1993:BRI

[1181] John Hedley Brooke. Book reviews: Immaterial beings or substantiated experiments? [A. Rupert Hall, Henry More. Magic, Religion and Experiment. Oxford: Blackwell, 1990. Pp. 304, £30.00. ISBN 0-631-17295-5]. Notes and Records of the Royal Society of London, 47(1):142–144, July 1993. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531402.

Allen:1993:BRS

[1182] David E. Allen. Book reviews: Spider-man at work [Martin Lister’s English Spiders 1678. Translated by Malcolm Davies & Basil Harley. Edited by John Parker & Basil Harley. Colchester, Harley Books, 1992. Pp. xv + 208, £49.95. ISBN 0-946589-27-5]. Notes and Records of the Royal Society of London, 47(1):144–145, July 1993. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531403.

Wilkinson:1993:BRD

[1183] Lise Wilkinson. Book reviews: Out, damned spots! [Richard B. Fisher, Edward Jenner 1749–1823. London: André Deutsch, 1991. Pp. 361, £20.00. ISBN 0-233-98681-2]. Notes and Records of the Royal Society of London, 47(1):145–147, July 1993. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531404.

Cantor:1993:BRL

[1184] Geoffrey Cantor. Book reviews: In love with light: [Larry J. Schaaf, Out of the Shadows. Herschel, Talbot and the Invention of Photography. New Haven and London: Yale University Press, 1992. Pp. xii + 188, £26.50.

ISBN 0-300-05705-9]. Notes and Records of the Royal Society of London, 47(1):147–149, July 1993. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531405.

Bowler:1993:BRS

[1185] Peter J. Bowler. Book reviews: Science in the service of society [Mary P. English, Victorian Values: The Life and Times of Dr Edwin Lankester, M.D., F.R.S. Bristol: Biopress Ltd., 1990. Pp. 187, £33.95. ISBN 0-948737-14-X]. Notes and Records of the Royal Society of London, 47 (1):149–150, July 1993. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531406.

Russell:1993:BRE

[1186] Colin A. Russell. Book reviews: Evangelist for science [J. Vernon Jensen, Thomas Henry Huxley: Communicating for Science. London and Toronto: Associated University Presses, 1991. Pp. 253. £29.95. ISBN 0-87413-379-3]. Notes and Records of the Royal Society of London, 47(1): 150–152, July 1993. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/531407.

Redhead:1993:BRU

[1187] Michael Redhead. Book reviews: Unveiling the atom [Abraham Pais, Niels Bohr’s Times, in Physics, Philosophy and Polity. Clarendon Press, Oxford, 1991. Pp. xvii + 565, £25.00. ISBN 0-19-852049-2]. Notes and Records of the Royal Society of London, 47(1):152–154, July 1993. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531408.

Feilden:1993:BRM

[1188] G. B. R. Feilden. Book review: Master of the airs [Sydney Camm and the Hurricane: Perspectives on the Master Fighter Designer and his finest Achievement. Edited by Dr John W. Fozard, F.R.S. Airlife Publishing, Shrewsbury, 1992. Pp. 256, £25.00. ISBN 1-85310-270-9]. Notes and Records of the Royal Society of London, 47(1):154–155, July 1993. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531409.

Mestel:1993:BRM

[1189] L. Mestel. Book reviews: ‘mathematics of the city of heaven’ [S. Chandrasekhar, Selected Papers. Volume 3: Stochastic Statistical and Hydromagnetic Problems in Physics and Astronomy Pp. xiv + 642; Volume 4: Plasma Physics, Hydrodynamic and Hydromagnetic Stability, and Applications of the Tensor–virial Theorem. Pp. xiii + 586; Volume 5: Relativistic Astrophysics. Pp. xx + 588; Volume 6: The Mathematical Theory of Black Holes and of Colliding Plane Waves. Pp. xix + 740. University of Chicago Press, 1989–91]. Notes and Records of the Royal Society of London, 47(1):156–158, July 1993. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531410.

Graham-Smith:1993:BRS

[1190] Sir Francis Graham-Smith, F.R.S. Book reviews: Seeing through the clouds [Bernard Lovell, Echoes of War: The Story of H_{2}S Radar. Bristol: Adam Hilger, 1992. Pp. xxi + 287, £17.50. ISBN 0-85274-317-3]. Notes and Records of the Royal Society of London, 47(1):159–160, July 1993. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531411.

Jones:1993:CLB

[1191] R. V. Jones. Corrigendum: Lindemann beyond the laboratory. Notes and Records of the Royal Society of London, 47(1):161, July 1993. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531412. See [1006].

Anonymous:1993:BMa

[1192] Anonymous. Back matter. Notes and Records of the Royal Society of London, 47(1):??, January 1993. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531413.

Anonymous:1993:FMa

[1193] Anonymous. Front matter. Notes and Records of the Royal Society of London, 47(1):??, January 1993. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531388.

Cook:1993:HL

[1194] Sir Alan Cook, F.R.S. Halley the Londoner. Notes and Records of the Royal Society of London, 47(2):163–177, July 1993. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/531783.

McConnell:1993:FMV

[1195] Anita McConnell. L. F. Marsigli’s visit to London in 1721, and his report on the Royal Society. Notes and Records of the Royal Society of London, 47(2):179–204, July 1993. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531784.

Chapman:1993:PRP

[1196] Allan Chapman. Pure research and practical teaching: the astronomical career of James Bradley, 1693–1762. Notes and Records of the Royal Society of London, 47(2):205–212, July 1993. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531785.

Stungo:1993:RSS

[1197] Ruth Stungo. The Royal Society specimens from the Chelsea Physic Garden, 1722–1799. Notes and Records of the Royal Society of London, 47(2):213–224, July 1993. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531786.

Johnson:1993:RJH

[1198] W. Johnson, F.Eng., F.R.S. Richard Jack and Henry Baker, F.R.S., in the late Summer of 1746. Notes and Records of the Royal Society of London, 47(2):225–231, July 1993. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531787.

Hall-Patch:1993:CBV

[1199] Anthony Hall-Patch. Charles Blacker Vignoles, F.R.S. Notes and Records of the Royal Society of London, 47(2):233–242, July 1993. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531788.

Anderson:1993:RAF

[1200] Ronald Anderson. The referees’ assessment of Faraday’s electromagnetic induction paper of 1831. Notes and Records of the Royal Society of London, 47(2):243–256, July 1993. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531789.

Jensen:1993:THH

[1201] J. Vernon Jensen. Thomas Henry Huxley’s address at the opening of the Johns Hopkins University in September 1876. Notes and Records of the Royal Society of London, 47(2):257–269, July 1993. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531790.

Pippard:1993:SRF

[1202] Sir Brian Pippard, F.R.S. Siegfried Ruhemann (1859–1943), F.R.S. 1914– 1923. Notes and Records of the Royal Society of London, 47(2):271– 276, July 1993. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531791.

Birch:1993:ISL

[1203] Arthur J. Birch, F.R.S. Investigating a scientific legend: the tropinone synthesis of Sir Robert Robinson, F.R.S. Notes and Records of the Royal Society of London, 47(2):277–296, July 1993. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531792.

Clark:1993:NHN

[1204] Alan J. Clark. Notes on the history of nos. 6, 7, 8 & 9 Carlton House Terrace, with a list of known occupants, 1829–1967. Notes and Records of the Royal Society of London, 47(2):297–303, July 1993. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531793.

Pirie:1993:ERB

[1205] N. W. Pirie. Essay review: Broad-minded biologist [Christopher Sexton, The Seeds of Time: the Life of Sir Macfarlane Burnet. Oxford University Press, 1991. Pp. 301, £19.95. ISBN 0-19-553274-0]. Notes and Records of the Royal Society of London, 47(2):305–310, July 1993. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531794.

Hall:1993:BRM

[1206] A. Rupert Hall. Book reviews: Ministry of science? [Julian Martin, Francis Bacon, the State, and the Reform of Natural Philosophy. Cambridge University Press, 1992. Pp. xiv + 236, £35 (hdbk). ISBN 0-521-38249-1]. Notes and Records of the Royal Society of London, 47(2):311, July 1993. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531795.

King-Hele:1993:BRO

[1207] D. G. King-Hele. Book reviews: Open Eurobook [The Rise of Scientific Europe, 1500–1800, edited by David Goodman and Colin A. Russell. Hodder & Stoughton and the Open University, 1991. Pp. ix + 437, £14.99 ISBN 0-340-55861-X]. Notes and Records of the Royal Society of London, 47(2):312–313, July 1993. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531796.

Johnson:1993:BRN

[1208] W. Johnson. Book reviews: Newtonianism pour la France [Éléments de la Philosophie de Newton, by Voltaire. Edited by Robert L. Walters and W. H. Barber. Oxford: the Voltaire Foundation, Taylor Institution, 1992]. Notes and Records of the Royal Society of London, 47(2):313– 315, July 1993. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531797.

Allen:1993:BRT

[1209] David E. Allen. Book reviews: Top-drawer birdman [A. Stuart Mason, George Edwards: the Bedell and his Birds. London: Royal College of Physicians, 1992. Pp. viii + 71, £10.00. ISBN 1-873240-48-1]. Notes and Records of the Royal Society of London, 47(2):316–317, July 1993. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531798.

Mason:1993:BRP

[1210] Stephen F. Mason. Book reviews: Poet, chemist — and sage? [David Knight, Humphry Davy: Science and Power. Oxford: Blackwells, 1992. Pp. 218 + xiii, £30.00. ISBN 0-631-16816-8]. Notes and Records of the Royal Society of London, 47(2):317–319, July 1993. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/531799.

Postgate:1993:BRE

[1211] John Postgate. Book reviews: Evolution in perspective [Peter J. Bowler, Charles Darwin — The Man and his Influence. Oxford: Basil Blackwell, 1990. Pp. xii + 250, £19.95. ISBN 0-631-16818-4]. Notes and Records of the Royal Society of London, 47(2):320–321, July 1993. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531800.

Buchanan:1993:BRC

[1212] R. Angus Buchanan. Book reviews: A cornucopia of instruments [Science Preserved: A Directory of Scientific Instruments in Collections in the United Kingdom and Eire. Compiled by Mary Holbrook, with additions and revisions by R.G.W. Anderson and D.J. Bryden. Published by the Trustees of the Science Museum on behalf of the British National Committee for the History of Science, Technology and Medicine. London: HMSO, 1992. Pp. 271, £35. ISBN 0-11-290060-7]. Notes and Records of the Royal Society of London, 47(2):321–322, July 1993. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531801.

Hide:1993:BRM

[1213] Raymond Hide, F.R.S. Book review: Mathematical physicist par excellence: Kameshwar C. Wali, Chandra: A Biography of S. Chandrasekhar. Chicago: University of Chicago Press, 1991. Pp. x + 341, £13.50, ISBN 0-226-87055-3. Notes and Records of the Royal Society of London, 47 (2):322–324, July 1993. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531802.

Anonymous:1993:BMb

[1214] Anonymous. Back matter. Notes and Records of the Royal Society of London, 47(2):??, July 1993. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531803.

Anonymous:1993:FMb

[1215] Anonymous. Front matter. Notes and Records of the Royal Society of London, 47(2):??, July 1993. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531782.

Philip:1994:IPR

[1216] J. R. Philip, F.R.S. An innumerate President of the Royal Society? Notes and Records of the Royal Society of London, 48(1):1–10, January 1994. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531415.

Nakajima:1994:RHF

[1217] Hideto Nakajima. Robert Hooke’s family and his youth: some new evidence from the will of the Rev. John Hooke. Notes and Records of the Royal Society of London, 48(1):11–16, January 1994. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531416.

Appleby:1994:JSF

[1218] John H. Appleby. James Spilman, F.R.S. (1680–1763), and Anglo– Russian commerce. Notes and Records of the Royal Society of London, 48(1):17–29, January 1994. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531417.

Johnson:1994:BRD

[1219] W. Johnson, F.Eng., F.R.S. Benjamin Robins during 1739–1742: ‘called to a publick employment … a very honorable post’. Notes and Records of the Royal Society of London, 48(1):31–42, January 1994. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531418.

Fraser:1994:JHR

[1220] Kevin J. Fraser. John Hill and the Royal Society in the eighteenth century. Notes and Records of the Royal Society of London, 48(1):43–67, January 1994. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531419.

Keynes:1994:PDE

[1221] Milo Keynes. Portraits of Dr Erasmus Darwin, F.R.S., by Joseph Wright, James Rawlinson and William Coffee. Notes and Records of the Royal Society of London, 48(1):69–84, January 1994. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531420.

White:1994:PMV

[1222] J. Edmund White. The Priestley memorial volume. Notes and Records of the Royal Society of London, 48(1):85–96, January 1994. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531421.

Edwards:1994:RFK

[1223] A. W. F. Edwards. R. A. Fisher on Karl Pearson. Notes and Records of the Royal Society of London, 48(1):97–106, January 1994. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531422.

Young:1994:RI

[1224] D. A. B. Young. Ramanujan’s illness. Notes and Records of the Royal Society of London, 48(1):107–119, January 1994. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531423.

Appleby:1994:NLL

[1225] John H. Appleby. A new lease of life for 71 missing Fellows. Notes and Records of the Royal Society of London, 48(1):121–125, January 1994. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531424.

Atiyah:1994:APS

[1226] Michael Atiyah. Address of the President, Sir Michael Atiyah, O. M., given at the Anniversary Meeting on 30 November 1993. Notes and Records of the Royal Society of London, 48(1):127–133, January 1994. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531425.

Peierls:1994:RDR

[1227] Sir Rudolf Peierls, F.R.S. Reminiscences and discoveries: Recollections of James Chadwick. Notes and Records of the Royal Society of London, 48 (1):135–141, January 1994. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531426.

Postgate:1994:ERE

[1228] John Postgate. Essay reviews: Eugenics revisited [Sir Francis Galton, F.R.S. — The Legacy of his Ideas. Proceedings of the twenty-eighth annual symposium of the Galton Institute, 1991. Edited by Milo Keynes. London: Macmillan, 1993. Pp x + 237, £40. ISBN 0-333-54695-4]. Notes and Records of the Royal Society of London, 48(1):143–146, January 1994. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531427.

Frank:1994:BRU

[1229] Sir Charles Frank, F.R.S. Book review: An uncertain life. David C. Cassidy, Uncertainty: The Life and Science of Werer Heisenberg. W. H. Freeman and Co., New York, 1991. $29.95, pp. xii + 669. ISBN 0-7167-2243-7. Notes and Records of the Royal Society of London, 48(1): 147–151, January 1994. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531428.

Hall:1994:BRS

[1230] Marie Boas Hall. Book review: Surveyor of the Restoration. Frances Willmoth, Sir Jonas Moore: Practical Mathematics and Restoration Science. Woodbridge, The Boydell Press, 1993. Pp.xi + 244. £35. ISBN 0-85115-321-6. Notes and Records of the Royal Society of London, 48 (1):153–154, January 1994. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531429.

Hollingdale:1994:BRV

[1231] Stuart Hollingdale. Book review: ‘Voyaging through strange seas of thought, alone’: A. Rupert Hall. Isaac Newton: Adventurer in Thought. Blackwells, Oxford, 1992. Pp. xv + 468, £19.99. ISBN 0-631-17906-2.

Notes and Records of the Royal Society of London, 48(1):154–156, January 1994. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531430.

Morton:1994:BRP

[1232] Alan Q. Morton. Book review: Previous public perceptions of science: Larry Stewart, The Rise of Public Science: Rhetoric, Technology and Natural Philosophy in Newtonian Britain, 1660–1750. Cambridge UP, 1992. Pp. 34 + 453, £45, ISBN 0-521-41700-7 Jan Golinski, Science as Public Culture: Chemistry and Enlightenment in Britain, 1760–1820. Cambridge UP, 1992. Pp. 330, £32.50. ISBN 0-521-39414-7. Notes and Records of the Royal Society of London, 48(1):157–159, January 1994. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531431.

King-Hele:1994:BRV

[1233] D. G. King-Hele. Book review: ‘Vase-maker general to the Universe’: Robin Reilly, Josiah Wedgwood, 1730–1795. Macmillan London, 1992. pp. xii + 412, £20. ISBN 0-333-51041-0. Notes and Records of the Royal Society of London, 48(1):159–161, January 1994. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/531432.

Torrens:1994:BRF

[1234] Hugh S. Torrens. Book review: First-class third man? John Griffiths, The Third Man — the Life and Times of William Murdoch (1754–1839)

— the Inventor of Gas Lighting. London: André Deutsch, 1992. Pp. 373, £20. ISBN 0-233-98778-9. Notes and Records of the Royal Society of London, 48(1):161–163, January 1994. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531433.

Bone:1994:BRP

[1235] Quentin Bone. Book review: Pictures from the past: Brian J. Ford, Images of Science — A History of Scientific Illustration, London: The British Library, 1993. Pp. viii + 208, ISBN 0-7123-0267-0. Notes and Records of the Royal Society of London, 48(1):163–164, January 1994. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531434.

Jones:1994:A

[1236] R. V. Jones. Addendum. Notes and Records of the Royal Society of London, 48(1):165, January 1994. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531435.

Anonymous:1994:BMa

[1237] Anonymous. Back matter. Notes and Records of the Royal Society of London, 48(1):??, January 1994. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531436.

Anonymous:1994:FMa

[1238] Anonymous. Front matter. Notes and Records of the Royal Society of London, 48(1):??, January 1994. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531414.

Chapman:1994:EHU

[1239] Allan Chapman. Edmond Halley’s use of historical evidence in the advancement of science. Notes and Records of the Royal Society of London, 48(2):167–191, July 1994. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/532161.

Bryden:1994:JGB

[1240] D. J. Bryden. John Gedde’s bee-house and the Royal Society. Notes and Records of the Royal Society of London, 48(2):193–213, July 1994. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/532162.

Johnson:1994:VAY

[1241] W. Johnson, F.Eng., F.R.S. Voltaire after 300 years. Notes and Records of the Royal Society of London, 48(2):215–220, July 1994. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/532163.

Johnson:1994:EGG

[1242] W. Johnson, F.Eng., F.R.S. Edward Gibbon and George Grote: a bicentenary in common. Notes and Records of the Royal Society of London, 48(2):221–226, July 1994. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/532164.

Mandelstam:1994:DCS

[1243] Joel Mandelstam, F.R.S. Du Chaillu’s stuffed gorillas and the savants from the British Museum. Notes and Records of the Royal Society of London, 48(2):227–245, July 1994. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

532165.

Bleaney:1994:PSO

[1244] Brebis Bleaney, F.R.S. The physical sciences in Oxford, 1918–1939 and earlier. Notes and Records of the Royal Society of London, 48(2):247– 261, July 1994. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/532166.

Clark:1994:NZN

[1245] Robin J. H. Clark, F.R.S. and Michael J. Kelly, F.R.S. New Zealand, New Zealanders and the Royal Society. Notes and Records of the Royal Society of London, 48(2):263–281, July 1994. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532167.

Lovell:1994:RSR

[1246] Sir Bernard Lovell, F.R.S. The Royal Society, the Royal Greenwich Observatory and the Astronomer Royal. Notes and Records of the Royal Society of London, 48(2):283–297, July 1994. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532168.

Holt:1994:RDJ

[1247] J. R. Holt, F.R.S. Reminiscences and discoveries: James Chadwick at Liverpool. Notes and Records of the Royal Society of London, 48(2):299– 308, July 1994. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/532169.

Browne:1994:ERC

[1248] Janet Browne. Essay reviews: Cataloguing for Empire [Richard Owen, The Hunterian Lectures in Comparative Anatomy. May and June 1837. Edited with an Introduction by Phillip Reid Sloan. Natural History Museum Publications, 1992. Pp. xii + 340, £37.50 hardback, £15.95 paper. ISBN 0-565-01106-5, 0-565-01144-8. Jacob W. Gruber and John C. Thackray, Richard Owen Commemoration: Three Studies. Historical Studies in the Life and Earth Sciences No. 1. Natural History Museum Publications, 1992. Pp. x + 181, £29.95. ISBN 0-565-01109-X]. Notes and Records of the Royal Society of London, 48(2):309–311, July 1994. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/532170.

Bondi:1994:ERE

[1249] H. Bondi. Essay reviews: Einstein illuminated [Roger Highfield and Paul Carter, The Private Lives of Albert Einstein. Faber & Faber, 1993. £15.99. ISBN 0-571-16744-6]. Notes and Records of the Royal Society of London, 48(2):313–316, July 1994. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

532171.

Mason:1994:BRT

[1250] Stephen F. Mason. Book reviews: A theosophical Boyle? [Robert Boyle Reconsidered. Edited by Michael Hunter. Cambridge University Press, 1994. Pp. xviii + 231, £30.00. ISBN 0-521-44205-2]. Notes and Records of the Royal Society of London, 48(2):317–319, July 1994. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/532172.

Cartwright:1994:BRN

[1251] David E. Cartwright. Book reviews: Naval maverick [Hugh Popham, A Damned Cunning Fellow — The Eventful Life of Rear Admiral Sir Home Popham, K.C.B., K.C.H., K.M., F.R.S., 1762–1820. The Old Ferry Press, 53 Vicarage Road, Tywardreath, Cornwall, 1991. Pp. xiii + 256, £27.50. ISBN 0-9516758-0-X]. Notes and Records of the Royal Society of London, 48(2):319–320, July 1994. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532173.

Mansfield:1994:BRM

[1252] Eric H. Mansfield. Book reviews: The miller of Sneinton [D. M. Cannell, George Green, Mathematician & Physicist 1793–1841. Athlone Press, 1993. Pp. xxvi + 265, £35 (hdbk). ISBN 0-485-11433-X]. Notes and Records of the Royal Society of London, 48(2):321–322, July 1994. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/532174.

Wood:1994:BRS

[1253] Alan Muir Wood. Book reviews: Self-confident engineer [Adrian Vaughan, Isambard Kingdom Brunel: Engineering Knight Errant. John Murray, paperback, 1993, £12.99 (first published 1991). ISBN 0-7195-5282-6]. Notes and Records of the Royal Society of London, 48(2):323– 325, July 1994. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/532175.

Zucker:1994:BRI

[1254] I. J. Zucker. Book reviews: India’s mathematical prodigy [Robert Kanigel, The Man Who Knew Infinity. A Life of the Genius Ramanujan. Abacus, 1991. Pp. 438. £6.99. ISBN 0-349-10452-2]. Notes and Records of the Royal Society of London, 48(2):325–326, July 1994. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/532176.

Muir:1994:BRO

[1255] Helen Muir. Book reviews: Outrunning the dominant males [Women Physiologists, edited by Lynn Bindman, Alison Bruding and Tilli Tansey. Portland Press, 1993. £16.95. ISBN 1-85578-021-6]. Notes and Records of the Royal Society of London, 48(2):327–328, July 1994. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/532177.

Anonymous:1994:BMb

[1256] Anonymous. Back matter. Notes and Records of the Royal Society of London, 48(2):??, July 1994. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/532178.

Anonymous:1994:FMb

[1257] Anonymous. Front matter. Notes and Records of the Royal Society of London, 48(2):??, July 1994. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/532160.

Keynes:1995:PIN

[1258] Milo Keynes. The personality of Isaac Newton. Notes and Records of the Royal Society of London, 49(1):1–56, July 1995. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/531881.

Principe:1995:NDB

[1259] Lawrence M. Principe. Newly discovered Boyle documents in the Royal Society archive: alchemical tracts and his student notebook. Notes and Records of the Royal Society of London, 49(1):57–70, July 1995. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531882.

Hall:1995:JCN

[1260] A. Rupert Hall, F.B.A. John Collins on Newton’s telescope. Notes and Records of the Royal Society of London, 49(1):71–78, July 1995. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531883.

Simcock:1995:RDT

[1261] A. V. Simcock. ‘Reason’s dim telescope’: a poetic tirade against Joseph Priestley, F.R.S. Notes and Records of the Royal Society of London, 49 (1):79–84, July 1995. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/531884.

Clarke:1995:VFB

[1262] Cyril A. Clarke, F.R.S. In vitro fertilization in Bombyx mori, the silkworm. Was John Hunter right or wrong in 1792? Notes and Records of the Royal Society of London, 49(1):85–92, July 1995. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/531885.

Johnson:1995:CIT

[1263] W. Johnson, F.Eng., F.R.S. Contributors to improving the teaching of calculus in early 19th-century England. Notes and Records of the Royal Society of London, 49(1):93–103, July 1995. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531886.

Harvey:1995:PGC

[1264] R. D. Harvey. Pioneers of genetics: a comparison of the attitudes of William Bateson and Erwin Baur to eugenics. Notes and Records of the Royal Society of London, 49(1):105–117, July 1995. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531887.

Sarkar:1995:JBH

[1265] Sahotra Sarkar. J. B. S. Haldane and R. A. Fisher’s draft life of Karl Pearson. Notes and Records of the Royal Society of London, 49(1):119– 124, July 1995. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531888.

Mason:1995:WFJ

[1266] Joan Mason. The women Fellows’ jubilee. Notes and Records of the Royal Society of London, 49(1):125–140, July 1995. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531889.

Atiyah:1995:APS

[1267] Sir Michael Atiyah, O.M. Address of the President, Sir Michael Atiyah, O. M., given at the Anniversary Meeting on 30 November 1994. Notes and Records of the Royal Society of London, 49(1):141–151, July 1995. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531890.

Chandrasekhar:1995:RDR

[1268] S. Chandrasekhar. Reminiscences and discoveries on Ramanujan’s bust: University of Chicago, Chicago, Illinois 60637, U.S.A. Notes and Records of the Royal Society of London, 49(1):153–157, July 1995. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531891.

Bone:1995:ERE

[1269] Quentin Bone. Essay review: Excavator extraordinary [Mark Bowden, Pitt Rivers: The life and archaeological work of Lieutenant-General Augustus Henry Lane Fox Pitt Rivers, D.C.L., F.R.S., F.S.A. Cambridge University Press 1991. Pp. 182, £27.95. ISBN 0-521-40077-5]. Notes and Records of the Royal Society of London, 49(1):159–162, July 1995. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531892.

King-Hele:1995:BRF

[1270] D. G. King-Hele. Book review: Fine Fellows, most of them: Michael Hunter, The Royal Society and its Fellows 1660–1700, 2nd edition. British Society for the History of Science, 1994. Pp. ix + 291, £10 (ppbk). ISBN 0-906450-09-8. Notes and Records of the Royal Society of London, 49(1):163–164, July 1995. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531893.

Howse:1995:BRF

[1271] Derek Howse. Book review: Filling naval roles: N. A. M. Rodger, The Insatiable Earl: A Life of John Montagu, 4th Earl of Sandwich 1718–1792. HarperCollins Publishers, London, 1993. Pp. xviii + 425, 19 illustrations, £20. ISBN 0-00-215784-5. Notes and Records of the Royal Society of London, 49(1):165–166, July 1995. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531894.

Anderson:1995:BRR

[1272] Robert Anderson. Book review: Royal treasures of science: Alan Q. Morton and Jane Wess, Public and Private Science: The King George III Collection. Oxford University Press in association with the Science Museum, 1993. Pp. 710. £75. ISBN 0-19-856392-2. Notes and Records of the Royal Society of London, 49(1):167–168, July 1995. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531895.

Chapman:1995:BRP

[1273] Allan Chapman. Book review: Pre-eminently precise: Anita McConnell, Instrument Makers to the World. A history of Cooke, Troughton, and Simms. University of York, William Sessions, York, 1992. Soft covers, 116 pp. £16.00. Eleanor Mennim, Transit Circle. The story of William Simms, 1793–1860. William Sessions, York, 1992. Soft covers. Notes and Records of the Royal Society of London, 49(1):168–170, July 1995. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531896.

Smith:1995:MLF

[1274] Crosbie Smith. More light on Faraday. Notes and Records of the Royal Society of London, 49(1):170–172, July 1995. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531897.

Clarke:1995:BRD

[1275] Patricia H. Clarke, F.R.S. Book review: Dissolving microbes in dissension [Antony Twort, In Focus, Out of Step: A Biography of Frederick William Twort F.R.S. Alan Sutton Publishing Ltd. Stroud, 1993. pp. xii + 340. £25. ISBN 0-7509-0327-9]. Notes and Records of the Royal Society of London, 49(1):173–174, July 1995. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531898.

Tutin:1995:BRI

[1276] Winifred Tutin, F.R.S. Book review: The invisible company. [G. Kass-Simon and Patricia Fames (eds). Women of Science. Indiana University Press, 1993. Pp. xvi + 398. £32.50, paperback £12.99. ISBN 0-253-33264-8]. Notes and Records of the Royal Society of London, 49(1):175–176, July 1995. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531899.

Clark:1995:A

[1277] R. J. H. Clark and M. J. Kelly. Addendum. Notes and Records of the Royal Society of London, 49(1):177, July 1995. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531900.

Anonymous:1995:BMa

[1278] Anonymous. Back matter. Notes and Records of the Royal Society of London, 49(1):??, January 1995. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531901.

Anonymous:1995:FMa

[1279] Anonymous. Front matter. Notes and Records of the Royal Society of London, 49(1):??, January 1995. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531880.

Hall:1995:FAR

[1280] A. Rupert Hall, F.B.A. The Francis affair reconsidered. Notes and Records of the Royal Society of London, 49(2):179–184, July 1995. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/532008.

Meynell:1995:LBP

[1281] Guy Meynell. Locke, Boyle and Peter Stahl. Notes and Records of the Royal Society of London, 49(2):185–192, July 1995. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/532009.

Jagger:1995:JMF

[1282] Graham Jagger. Joseph Moxon, F.R.S., and the Royal Society. Notes and Records of the Royal Society of London, 49(2):193–208, July 1995. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/532010.

Unwin:1995:PMS

[1283] Robert W. Unwin. A provincial man of science at work: Martin Lister, F.R.S., and his illustrators 1670–1683. Notes and Records of the Royal Society of London, 49(2):209–230, July 1995. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532011.

King-Hele:1995:EDL

[1284] Desmond King-Hele, F.R.S. Erasmus Darwin’s life at Lichfield: fresh evidence. Notes and Records of the Royal Society of London, 49(2):231– 243, July 1995. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/532012.

Carter:1995:RSV

[1285] Harold B. Carter. The Royal Society and the voyage of HMS Endeavour 1768–71. Notes and Records of the Royal Society of London, 49(2):245– 260, July 1995. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/532013.

Katritzky:1995:CLL

[1286] Linde Katritzky. Coleridge’s links with leading men of science. Notes and Records of the Royal Society of London, 49(2):261–276, July 1995. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/532014.

Mackintosh:1995:TMC

[1287] A. R. Mackintosh, F.R.S. The third man: Charles Drummond Ellis, 1895–1980. Notes and Records of the Royal Society of London, 49(2): 277–293, July 1995. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/532015.

Bolton:1995:CLS

[1288] H. C. Bolton and Alan Roberts. On the comparison of literary and scientific styles: The letters and articles of Max Born, F.R.S. Notes and Records of the Royal Society of London, 49(2):295–302, July 1995. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

URL http://www.jstor.org/stable/532016;   http://www.jstor.
org/stable/pdfplus/532016.pdf.

Plesch:1995:SRA

[1289] Janos Plesch and Peter H. Plesch. Some reminiscences of Albert Einstein. Notes and Records of the Royal Society of London, 49(2):303–328, July 1995. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/532017.

Pagel:1995:ERL

[1290] B. E. J. Pagel, F.R.S. Essay review: A life of achievement and controversy [Fred Hoyle, Home is where the Wind Blows. University Science Books, Mill Valley, California, 1994. Pp. xi + 443, $28.50. ISBN 0-935702-27-X]. Notes and Records of the Royal Society of London, 49 (2):329–332, July 1995. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/532018.

Torrens:1995:BRN

[1291] Hugh S. Torrens. Book reviews: A neglected ‘Mecca’ for historians of science [Marco Beretta, A History of Non-Printed Science: a select catalogue of the Waller Collection. Uppsala, 1993. Pp. 199, 178 Swedish Krona. ISBN 91-554-3070-8 (ISSN 0346-7465)]. Notes and Records of the Royal Society of London, 49(2):333–334, July 1995. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/532019.

Hall:1995:BRB

[1292] Marie Boas Hall. Book reviews: Boyle anatomized [Robert Boyle by Himself and his Friends with a fragment of William Wotton’s lost ‘Life of Boyle’. Edited with an introduction by Michael Hunter. London, William Pickering, 1994. Pp. ciii + 188, £9.95. ISBN 1-85196-085-6]. Notes and Records of the Royal Society of London, 49(2):334–335, July 1995. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/532020.

Allen:1995:BRM

[1293] David E. Allen. Book reviews: Man of the world [Sir Joseph Banks: a global perspective. Edited by R. E. R. Banks, B. Elliott, J. G. Hawkes, D. King-Hele & G. Ll. Lucas (eds), Kew: Royal Botanic Gardens, 1994. Pp. ii + 235. £12.00. ISBN 0-947643-61-3]. Notes and Records of the Royal Society of London, 49(2):336–337, July 1995. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532021.

Mason:1995:BRP

[1294] Stephen F. Mason, F.R.S. Book review: From phlogiston to caloric and oxygen. [Arthur Donovan, Antoine Lavoisier: Science, Administration, and Revolution. Blackwell, Oxford, 1993. Pp. xv + 351. £35.00. ISBN 0-521-56672-X]. Notes and Records of the Royal Society of London, 49 (2):338–340, July 1995. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/532022.

Woolrich:1995:BRE

[1295] A. P. Woolrich. Book review: Engineering myths. [Perceptions of Great Engineers: fact and fantasy. Edited by Denis Smith. Science Museum, 1994. Pp. 115, paperback, illus., Price £9.95 ISBN 0-901805-65-3]. Notes and Records of the Royal Society of London, 49(2):340–341, July 1995. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/532023.

Wilkes:1995:BRU

[1296] Maurice V. Wilkes. Book review: The ubiquitous Mr Babbage: Passages from the Life of a Philosopher [Charles Babbage, Passages from the Life of a Philosopher (ed. M. Campbell-Kelly). London. Pickering and Chatto, 1994. Pp. 392, £12.95. ISBN 1-85196-040-6]. Notes and Records of the Royal Society of London, 49(2):341–342, July 1995. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/532024.

Harvey:1995:BRR

[1297] Joy Harvey. Book reviews: Rise and fall of the chief anatomist [Nicolaas A. Rupke, Richard Owen: Victorian Naturalist. New Haven & London: Yale University Press, 1994. Pp. 480, £35.00. ISBN 0-300-05820-9]. Notes and Records of the Royal Society of London, 49(2):343–344, July 1995. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/532025.

Keynes:1995:BRD

[1298] Richard D. Keynes. Book review: Darwin’s bulldog. [Adrian Desmond, Huxley: The Devil’s Disciple. London, Michael Joseph, 1994. Pp. xvii + 475, 38 illustrations, £20. ISBN 0-7181-3641-1]. Notes and Records of the Royal Society of London, 49(2):345–347, July 1995. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/532026.

Clarke:1995:BRI

[1299] Patricia H. Clarke. Book review: Inborn errors of metabolism. [Alexander G. Beam, Archibald Garrod and the Individuality of Man. Oxford, Clarendon Press 1993. Pp. xvi + 227, £20. ISBN 0-19-262145-9]. Notes and Records of the Royal Society of London, 49(2):348–349, July 1995. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/532027.

Bone:1995:BRS

[1300] Quentin Bone. Book review: Spokesman for science. [Julian Huxley: Biologist and Statesman of Science. Edited by C.K. Waters and A. Van Helden. Rice University Press, 1993. Pp. 344 $32.50 ISBN 0-89263-314-X]. Notes and Records of the Royal Society of London, 49(2):350–352, July 1995. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/532028.

Anonymous:1995:BMb

[1301] Anonymous. Back matter. Notes and Records of the Royal Society of London, 49(2):??, July 1995. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/532029.

Anonymous:1995:FMb

[1302] Anonymous. Front matter. Notes and Records of the Royal Society of London, 49(2):??, July 1995. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/532007.

Hall:1996:ARS

[1303] A. Rupert Hall, F.B.A. and A. D. C. Simpson. An account of the Royal Society’s Newton telescope. Notes and Records of the Royal Society of London, 50(1):1–11, January 1996. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531836.

Appleby:1996:HCR

[1304] John H. Appleby. Human curiosities and the Royal Society, 1699–1751. Notes and Records of the Royal Society of London, 50(1):13–27, January 1996. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531837.

Sorrenson:1996:THR

[1305] Richard Sorrenson. Towards a history of the Royal Society in the eighteenth century. Notes and Records of the Royal Society of London, 50 (1):29–46, January 1996. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531838.

Stungo:1996:RAC

[1306] Ruth Stungo. Recording the Aloes at Chelsea — a singular solution to a difficult problem. Notes and Records of the Royal Society of London, 50(1):47–57, January 1996. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531839.

Cook:1996:MDI

[1307] Gordon C. Cook. Mary Darwin’s illness. Notes and Records of the Royal Society of London, 50(1):59–63, January 1996. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531840.

Bowers:1996:MFG

[1308] Brian Bowers and Keith Bowers. Michael Faraday’s geological notes on the Isle of Wight. Notes and Records of the Royal Society of London, 50 (1):65–74, January 1996. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531841.

Lovie:1996:CES

[1309] P. Lovie and A. D. Lovie. Charles Edward Spearman, F.R.S. (1863– 1945). Notes and Records of the Royal Society of London, 50(1):75–88, January 1996. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531842.

Rothman:1996:GCY

[1310] Patricia Rothman. Grace Chisholm Young and the division of laurels. Notes and Records of the Royal Society of London, 50(1):89–100, January 1996. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531843.

Atiyah:1996:APS

[1311] Michael Atiyah, Joseph Rotblat, and Patricia A. Jacobs. Address of the President, Sir Michael Atiyah, O. M., given at the Anniversary Meeting on 30 November 1995. Notes and Records of the Royal Society of London, 50(1):101–113, January 1996. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531844.

Ramaseshan:1996:RDD

[1312] S. Ramaseshan. Reminiscences and discoveries: Dorothy Hodgkin and the Indian connection. Notes and Records of the Royal Society of London, 50(1):115–127, January 1996. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531845.

Keynes:1996:BRE

[1313] Richard D. Keynes. Book review: The evolution of a theory. [Janet Browne, Charles Darwin Voyaging. Volume 1 of a Biography. Jonathan Cape, London, 1995. Pp. xiii + 605, 86 illustrations and maps, £25. ISBN 0-224-04202-5]. Notes and Records of the Royal Society of London, 50 (1):129–136, January 1996. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531846.

Hall:1996:BRO

[1314] Marie Boas Hall. Book review: Of oak and smoke, inter alia. [The Writings of John Evelyn (edited and typeset by Guy de la Bedoyere). Woodbridge, The Boydell Press, 1995. Pp. 435, £39.50. ISBN 0-85115-631-2]. Notes and Records of the Royal Society of London, 50(1):137–138, January 1996. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531847.

Mason:1996:HRR

[1315] Stephen F. Mason. Hooke recommemorated. [Richard Nichols, The Diaries of Robert Hooke, The Leonardo of London, 1635–1703. Lewes, Sussex: The Book Guild, 1994, Pp. 185, £15.00. ISBN 0-86332-930-6]. Notes and Records of the Royal Society of London, 50(1):139–140, January 1996. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531848.

Hunter:1996:RPR

[1316] Michael Hunter. The realm of Prince Rupert. [A. Rupert Hall, Newton, his Friends and Foes and Science and Society: Historical Essays on the Relations of Science, Technology and Medicine. Aldershot: Variorum, 1993 and 1994; pp. xii + 330 and x + 324, £49.50 and £47.50. ISBN 0-86078-347-2 and 0-86078-400-2]. Notes and Records of the Royal Society of London, 50(1):140–142, January 1996. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531849.

Bone:1996:BRF

[1317] Quentin Bone. Book review: Favourite or outsider? [George Bartle, An Old Radical and his Brood, Janus Publishing Co. London, 1994, £9.95. ISBN 1-85756-132-5. Sir John Bowring, 1792–1872: Aspects of his Life and Career. Edited by Joyce Youings, Latimer Trend Plymouth for the Devonshire Association, 1993, £9.00. ISBN 0-85214-052-5]. Notes and Records of the Royal Society of London, 50(1):142–144, January 1996. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531850.

Cartwright:1996:BRP

[1318] David E. Cartwright. Book review: Pedestrian extraordinary. [Robert L. Richards, Dr John Rae. Caedmon of Whitby, Publishers, 1994. Pp. 231, £10.50 (Paperback). ISBN 0-905355-29-6]. Notes and Records of the Royal Society of London, 50(1):144–146, January 1996. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531851.

Postgate:1996:BRS

[1319] John Postgate, F.R.S. Book review: Saint or demon? Roslynn D. Haynes, From Faust to Strangelove, Representations of the Scientist in Western Literature, Baltimore and London. The Johns Hopkins University Press, 1994, Pp.417, paperback £16.50, (Hardback £45.50). ISBN 0-8018-4801-6. Notes and Records of the Royal Society of London, 50 (1):146–147, January 1996. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531852.

King-Hele:1996:FE

[1320] D. G. King-Hele. Farewell editorial. Notes and Records of the Royal Society of London, 50(1):149, January 1996. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531853.

Anonymous:1996:FM

[1321] Anonymous. Front matter. Notes and Records of the Royal Society of London, 50(1):??, January 1996. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531835.

Cook:1996:PNE

[1322] Sir Alan Cook, F.R.S. Preface [from the new Editor]. Notes and Records of the Royal Society of London, 50(2):151, July 1996. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531903.

Cook:1996:ABG

[1323] Sir Alan Cook, F.R.S. The 350th anniversary of the birth of G. W. Leibniz, F.R.S. Notes and Records of the Royal Society of London, 50 (2):153–163, July 1996. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531904.

Hill:1996:NAN

[1324] Katherine Hill. Neither ancient nor modern: Wallis and Barrow on the composition of continua. I. Mathematical styles and the composition of continua. Notes and Records of the Royal Society of London, 50(2):165– 178, July 1996. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531905.

Appleby:1996:JTF

[1325] John H. Appleby. James Theobald, F.R.S. (1688–1759), merchant and natural historian. Notes and Records of the Royal Society of London, 50(2):179–189, July 1996. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531906.

Edwards:1996:FGT

[1326] A. W. F. Edwards. Is the frontispiece of Gulliver’s Travels a likeness of Newton? Notes and Records of the Royal Society of London, 50(2):191– 194, July 1996. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531907.

Willach:1996:NLI

[1327] R. Willach. New light on the invention of the achromatic telescope objective. Notes and Records of the Royal Society of London, 50(2):195–210, July 1996. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531908.

Brown:1996:HRP

[1328] Malcolm Brown. How not to ‘regain paradise’: Henry Bellenden Ker, F.R.S. from 1819 to 1831. Notes and Records of the Royal Society of London, 50(2):211–215, July 1996. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531909.

Bradley:1996:CEA

[1329] J. K. Bradley and E. M. Tansey. The coming of the electronic age to the Cambridge Physiological Laboratory: E. D. Adrian’s valve amplifier in 1921. Notes and Records of the Royal Society of London, 50(2):217– 228, July 1996. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531910.

Kendall:1996:RDB

[1330] David Kendall, F.R.S. and T. D. Kenneth Post, C.B.E. Reminiscences and discoveries: The British 3-inch anti-aircraft rocket. Part one: Divebombers. Notes and Records of the Royal Society of London, 50(2):229– 239, July 1996. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531911. See also

Part 2 [1352].

Hamblyn:1996:ERS

[1331] Richard Hamblyn. Essay reviews: The stones of Venice [A Calendar of the Correspondence of John Strange, F.R.S. (1732–1799). Edited with an introduction by Luca Ciancio. London, The Wellcome Institute for the History of Medicine, 1995. Pp. 152, £8.00 (Paperback). ISBN 0-85484-063-X]. Notes and Records of the Royal Society of London, 50 (2):241–244, July 1996. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531912.

Hughes:1996:BRJ

[1332] David W. Hughes. Book reviews: J. C. Adams, Cambridge and Neptune [H. M. Harrison, Voyager in Time and Space: The life of John Couch Adams, Cambridge Astronomer. The Book Guild Ltd, Sussex, 1994. Pp. 282, £15.00 (Hardbound ISBN 0-86332-918-7)]. Notes and Records of the Royal Society of London, 50(2):245–248, July 1996. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531913.

Mason:1996:BRT

[1333] Stephen F. Mason. Book reviews: Turning the wheel of history [Michael Hunter, Science and the Shape of Orthodoxy: Intellectual Change in Late Seventeenth-century Britain. Woodbridge, The Boydell Press, 1995. Pp. xii + 345, £55.00. ISBN 0-85115-594-4]. Notes and Records of the Royal Society of London, 50(2):249–250, July 1996. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531914.

Meadows:1996:BRJ

[1334] A. J. Meadows. Book reviews: John Flamsteed, our astronomical observator [The Correspondence of John Flamsteed, First Astronomer Royal. Volume I (1666–1682). Compiled and edited by Eric G. Forbes and (for Maria Forbes) by Lesley Murdin and Frances Willmoth. Bristol, Institute of Physics Publishing, 1995. Pp. xlix + 955. ISBN 0-7503-0147-3]. Notes and Records of the Royal Society of London, 50(2):250–253, July 1996. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531915.

Lynden-Bell:1996:BRW

[1335] D. Lynden-Bell. Book reviews: The wonderful geometry of dynamics [S. Chandrasekhar, Newton’s Principia for the Common Reader: Oxford University Press, 1995. ISBN 0-19-851744-0]. Notes and Records of the Royal Society of London, 50(2):253–255, July 1996. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531916.

King-Hele:1996:BRM

[1336] Desmond King-Hele. Book reviews: Man of many mysteries [Joy Hancox, The Queen’s Chameleon: the Life of John Byrom. London, Jonathan Cape, 1994. Pp. 276, £18.99. ISBN 0-224-03047-7]. Notes and Records of the Royal Society of London, 50(2):256–258, July 1996. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531917.

Hall:1996:BRE

[1337] Marie Boas Hall. Book reviews: The enlightening of science [John Gascoigne, Joseph Banks and the English Enlightenment: Useful Knowledge and Polite Culture. Cambridge University Press, 1994. Pp. xi + 324, 111., £35.00. ISBN 0-521-45077-2]. Notes and Records of the Royal Society of London, 50(2):258–260, July 1996. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531918.

Ashworth:1996:BRD

[1338] William J. Ashworth. Book reviews: Defining Whewell [Richard Yeo, Defining Science: William Whewell, Natural Knowledge, and Public Debate in Early Victorian Britain. Cambridge University Press, 1993. Pp. xiv + 280, £35.00. ISBN 0-521-43182-4]. Notes and Records of the Royal Society of London, 50(2):260–262, July 1996. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531919.

Woolfson:1996:BRG

[1339] M. M. Woolfson. Book reviews: The gentle pioneer [P. P. Ewald and his Dynamical Theory of X-ray Diffraction. Edited by D. W. J. Cruickshank, H. J. Juretschke & N. Kato. International Union of Crystallography, Oxford University Press, 1992. Pp. x + 161, £40.00. ISBN 0-19-855379-X]. Notes and Records of the Royal Society of London, 50(2):263–264, July 1996. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531920.

Wolfendale:1996:BRM

[1340] Arnold Wolfendale. Book reviews: A magnificent obsession [G. Venkataraman, Bhabha and his Magnificent Obsession Vignettes in Physics, Sangam Books, London, 1994. £6.95 (Paperback) ISBN 0-86311-555-1]. Notes and Records of the Royal Society of London, 50(2):265–266, July 1996. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531921.

Anonymous:1996:BM

[1341] Anonymous. Back matter. Notes and Records of the Royal Society of London, 50(2):??, July 1996. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531922.

Cook:1997:LSR

[1342] Alan Cook, F.R.S. Ladies in the scientific revolution. Notes and Records of the Royal Society of London, 51(1):1–12, January 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/532031.

Hill:1997:NAN

[1343] Katherine Hill. Neither ancient nor modern: Wallis and Barrow on the composition of continua. II. The seventeenth-century context: the struggle between ancient and modern. Notes and Records of the Royal Society of London, 51(1):13–22, January 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532032.

Appleby:1997:RST

[1344] John H. Appleby. The Royal Society and the Tartar lamb. Notes and Records of the Royal Society of London, 51(1):23–34, January 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/532033.

Lopez:1997:RSP

[1345] Susana Gómez López. The Royal Society and post-Galilean science in Italy. Notes and Records of the Royal Society of London, 51(1):35–44, January 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/532034.

Hull:1997:RHF

[1346] Derek Hull, F.R.S., F.Eng. Robert Hooke: A fractographic study of Kettering-stone. Notes and Records of the Royal Society of London, 51 (1):45–55, January 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/532035.

Dickson:1997:THS

[1347] Donald R. Dickson. Thomas Henshaw and Sir Robert Paston’s pursuit of the red elixir: An early collaboration between Fellows of The Royal Society. Notes and Records of the Royal Society of London, 51(1):57–76, January 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/532036.

Young:1997:DIS

[1348] D. A. B. Young. Darwin’s illness and systemic Lupus erythematosus. Notes and Records of the Royal Society of London, 51(1):77–86, January 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/532037.

Packer:1997:LOO

[1349] Kathryn Packer. A laboratory of one’s own: the life and works of Agnes Arber, F.R.S. (1879–1960). Notes and Records of the Royal Society of London, 51(1):87–104, January 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532038.

Tanner:1997:PLN

[1350] P. G. Tanner and T. E. Allibone, F.Eng., F.R.S. The patent literature of Nobel laureate Dennis Gabor (1900–1979). Notes and Records of the Royal Society of London, 51(1):105–120, January 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/532039.

Klug:1997:APS

[1351] Aaron Klug, O.M., P.R.S. Address of the President, Sir Aaron Klug, O.M., P.R.S., given at the Anniversary Meeting on 29 November 1996. Notes and Records of the Royal Society of London, 51(1):121–131, January 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/532040.

Kendall:1997:RDB

[1352] D. Kendall, F.R.S. and T. D. K.Post, C.B.E. Reminiscences and discoveries: The British 3-inch anti-aircraft rocket. Part two: high–flying bombers. Notes and Records of the Royal Society of London, 51(1):133– 140, January 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/532041.

See also Part 1 [1330].

Parkin:1997:BRL

[1353] Jon Parkin. Book review: Locke and modernity: Nicholas Wolterstorff, John Locke and the Ethics of Belief. Cambridge Studies in Religion and Critical Thought, Cambridge University Press, 1996. Pp. 248, £14.95. ISBN 0-521-55909-X, paperback. Notes and Records of the Royal Society of London, 51(1):141–142, January 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532042.

Parkin:1997:BR

[1354] Jon Parkin. Book reviews. Notes and Records of the Royal Society of London, 51(1):141–157, January 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

McConnell:1997:BRD

[1355] Anita McConnell. Book review: The Danube displayed: John Stoye, Marsigli’s Europe 1680–1730, the Life and Times of Luigi Ferdinando Marsigli, Soldier and Virtuoso. New Haven and London, Yale University Press, 1994. Pp. 356 + maps and illust., £29.95. ISBN 0-300-05542-0. Notes and Records of the Royal Society of London, 51(1):142–144, January 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/532043.

Kurtz-Newell:1997:BRL

[1356] Nicole Kurtz-Newell. Book review: Lavoisier, scientist and public figure: Arthur Donovan, Antoine Lavoisier: Science, Administration and Revolution, Cambridge University press, 1996. ISBN 0-521-56218-X, hardback. Notes and Records of the Royal Society of London, 51(1):144–146, January 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/532044.

Pippard:1997:BRG

[1357] Sir Brian Pippard, F.R.S. Book review: A great man, after all: James R. Hofmann, André-Marie Ampère. Cambridge University Press, 1996. Pp. xiv + 406, £40.00. ISBN 0-521-56220-1, hardback. Notes and Records of the Royal Society of London, 51(1):146–147, January 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/532045.

Torrens:1997:BRA

[1358] H. S. Torrens. Book review: An American in Glasgow (and elsewhere): Patsy Gerstner, Henry Darwin Rogers, 1808–1866 — American Geologist. University of Alabama Press, 1995. Pp. 313, £44.95. ISBN 0-8173-0735-4. Notes and Records of the Royal Society of London, 51(1):147–148, January 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/532046.

Keynes:1997:BRB

[1359] Richard D. Keynes, C.B.E., F.R.S. Book review: From the Beagle to Down House: Charles Darwin’s Letters. A selection 1825–1859. Edited by Frederick Burkhardt. Cambridge University Press, 1996. Pp. xxvi + 249, 2 illustrations and map, £14.95. ISBN 0-521-56212-0, hardback. Notes and Records of the Royal Society of London, 51(1):149–150, January 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/532047.

Lovell:1997:BRI

[1360] Sir Bernard Lovell, F.R.S. Book review: The invention of radio communication: Oliver Lodge and the Invention of Radio. Edited by Peter Rowlands and J. Patrick Wilson. PD Publications,1994. Pp. vii + 241. ISBN 1-873694-02-4, paperback. Notes and Records of the Royal Society of London, 51(1):151–153, January 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532048.

Dainton:1997:BRR

[1361] Lord Dainton, F.R.S. Book review: Reflections of an Indian scientist– statesman: V. V. Krishna, S. S. Bhatnagar on Science, Technology and Development 1938–54. Wiley Eastern, 1993. Pp xiv + 290. ISBN 81-224-0501-0. Notes and Records of the Royal Society of London, 51(1): 153–155, January 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/532049.

Taylor:1997:BRF

[1362] John C. Taylor. Book review: Feynman: Jagdish Mehra, The Beat of a Different Drum. The Life and Science of Richard Feynman. Oxford University Press, 1994. Pp. 630, £25.00. ISBN 0-19-853948-7, hardback. Notes and Records of the Royal Society of London, 51(1):156–157, January 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/532050.

Anonymous:1997:BMa

[1363] Anonymous. Back matter. Notes and Records of the Royal Society of London, 51(1):??, January 1997. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

532051.

Anonymous:1997:FMa

[1364] Anonymous. Front matter. Notes and Records of the Royal Society of London, 51(1):??, January 1997. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

532030.

Cook:1997:I

[1365] Alan Cook, F.R.S. In this issue. Notes and Records of the Royal Society of London, 51(2):159–160, July 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531982.

Cooper:1997:RHW

[1366] M. A. R. Cooper. Robert Hooke’s work as surveyor for the City of London in the aftermath of the Great Fire. Part one: Robert Hooke’s first surveys for the City of London. Notes and Records of the Royal Society of London, 51(2):161–174, July 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/ stable/531983. See also Parts 2 and 3 [1394, 1415].

Cantor:1997:QRS

[1367] G. Cantor. Quakers in the Royal Society, 1660–1750. Notes and Records of the Royal Society of London, 51(2):175–193, July 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531984.

Lynden-Bell:1997:SNR

[1368] D. Lynden-Bell and R. M. Lynden-Bell. On the shapes of Newton’s revolving orbits. Notes and Records of the Royal Society of London, 51 (2):195–198, July 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531985.

Fara:1997:RSP

[1369] P. Fara. The Royal Society’s portrait of Joseph Banks. Notes and Records of the Royal Society of London, 51(2):199–210, July 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531986.

Pierpoint:1997:ESE

[1370] William S. Pierpoint. Edward Stone (1702–1768) and Edmund Stone (1700–1768): confused identities resolved. Notes and Records of the Royal Society of London, 51(2):211–217, July 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/531987.

Ross:1997:DE

[1371] S. Ross. The Dalton entail. Notes and Records of the Royal Society of London, 51(2):219–233, July 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531988.

Henderson:1997:MV

[1372] D. A. Henderson. The miracle of vaccination. Notes and Records of the Royal Society of London, 51(2):235–245, July 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531989.

James:1997:JJS

[1373] I. M. James. James Joseph Sylvester, F.R.S. (1814–1897). Notes and Records of the Royal Society of London, 51(2):247–261, July 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531990.

Jeffery:1997:VFE

[1374] J. V. Jeffery. The Varley family: engineers and artists. Notes and Records of the Royal Society of London, 51(2):263–279, July 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531991.

Wilkes:1997:SEA

[1375] M. V. Wilkes. Sir Edward Appleton and early ionosphere research. Notes and Records of the Royal Society of London, 51(2):281–290, July 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531992.

Debru:1997:UHS

[1376] C. Debru. On the usefulness of the history of science for scientific education. Notes and Records of the Royal Society of London, 51(2):291–307, July 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531993.

Mackintosh:1997:CE

[1377] A. R. Mackintosh. The crocodile and the elephant. Notes and Records of the Royal Society of London, 51(2):309–316, July 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531994.

Bleaney:1997:JRF

[1378] B. Bleaney. Jubilees of radio-frequency spectroscopy. Notes and Records of the Royal Society of London, 51(2):317–326, July 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531995.

Cook:1997:ERS

[1379] Alan Cook, F.R.S. Essay review: Seeds of the scientific revolution [Edward Grant, The Foundations of Modern Science in the Middle Ages, Their Religious, Institutional and Intellectual Contexts, Cambridge University Press, 1996. Pp. xiv + 247. ISBN 0-521-56137-X. Benedino Gemelli, Aspetti dell’Atomismo classico nella Filosofia di Francis Bacon e nel Seicento, Accademia toscana di Scienze e Lettere ‘La Colombaria’, ‘Studi’ CLII, Olschki, Firenze, 1996. Pp. 433, Lit 86 000. ISBN 88-222-4460-5]. Notes and Records of the Royal Society of London, 51 (2):327–334, July 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531996.

Cook:1997:BRA

[1380] Sir Alan Cook, F.R.S. Book review: An astronomical patriarch: Anna Cassini, Gio: Domenico Cassini Uno scienzato del Seicento. Comune di Perinaldo, 1994. Pp. 315, Lit 35 000. Notes and Records of the Royal Society of London, 51(2):335–337, July 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531997.

Cook:1997:BR

[1381] Alan Cook, F.R.S. Book reviews. Notes and Records of the Royal Society of London, 51(2):335–348, July 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Willmoth:1997:BRR

[1382] Frances Willmoth. Book review: Reissued: Hall’s More and Newton: A. Rupert Hall, Henry More and the Scientific Revolution. Cambridge University Press, 1996. Pp. xii + 304, £40 (hardback). ISBN 0-521-56223-6. A. Rupert Hall, Isaac Newton: Adventurer in Thought. Cambridge University Press, 1996. Pp. xvi + 468, £37.50 (hardback) ISBN 0-521-56221-X; £13.95 (paperback) ISBN 0-521-56669-X. Notes and Records of the Royal Society of London, 51(2):337, July 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531998.

Torrens:1997:BRN

[1383] H. S. Torrens. Book review: New insights from Vectia: Ellen Tan Drake, Restless Genius: Robert Hooke and his Earthly Thoughts. Oxford University Press, 1996. Pp. xiv + 386, $65. ISBN 0-19-506695-2. Notes and Records of the Royal Society of London, 51(2):338–339, July 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531999.

Hull:1997:BRE

[1384] Derek Hull, F.R.S., F.Eng. Book review: Early microscopists: Edward G. Ruestow, The Microscope in the Dutch Republic: The Shaping of Discovery. Cambridge University Press, 1996. Pp. 348, £40.00 (hardback). ISBN 0-521-47078-1. Notes and Records of the Royal Society of London, 51(2):339–340, July 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

532000.

Torrens:1997:BRF

[1385] H. S. Torrens. Book review: The father of Derbyshire geology: Maxwell Craven, John Whitehurst of Derby: Clockmaker and Scientist 1713–88, Ashbourne, Derbyshire, Mayfield Books, 1997. Pp. 272, £24.99 (hardback). ISBN 0-9523270-3-1. Notes and Records of the Royal Society of London, 51(2):341–342, July 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532001.

Harman:1997:BRC

[1386] P. M. Harman. Book review: The chemist in the world: David Knight, Humphry Davy, Science and Power. Cambridge University Press, 1996. Pp. xiv + 218, c. £30.00. ISBN 0-521-56309-7. Notes and Records of the Royal Society of London, 51(2):343–344, July 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/532002.

Hesse:1997:BRP

[1387] Mary Hesse, F.B.A. Book review: A philosopher of science and humanism: Karl Popper: Philosophy and Problems. Edited by Anthony O’Hear. Cambridge University Press, 1995. Pp. 297, £14.95. ISBN 0-521-55815-8. Notes and Records of the Royal Society of London, 51(2):344–346, July 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/532003.

Coley:1997:BRC

[1388] Noel Coley. Book review: Creators of modern science: David, Ian, John and Margaret Millar, The Cambridge Dictionary of Scientists. Cambridge University Press, 1996. Pp. 387 + xii, £30.00 (hardback). ISBN 0-521-56185-X. Notes and Records of the Royal Society of London, 51 (2):346–348, July 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/532004.

Anonymous:1997:AIN

[1389] Anonymous. Author index to Notes and Records, volume 51, 1997. Notes and Records of the Royal Society of London, 51(2):349–350, July 1997. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531980.

Anonymous:1997:BMb

[1390] Anonymous. Back matter. Notes and Records of the Royal Society of London, 51(2):??, July 1997. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/532005.

Anonymous:1997:FMb

[1391] Anonymous. Front matter. Notes and Records of the Royal Society of London, 51(2):??, July 1997. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531981.

Cook:1998:SI

[1392] Alan Cook. Science from India. Notes and Records of the Royal Society of London, 52(1):1–2, January 22, 1998. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532072.

Peck:1998:UAL

[1393] L. V. Peck. Uncovering the Arundel Library at the Royal Society: changing meanings of science and the fate of the Norfolk donation. Notes and Records of the Royal Society of London, 52(1):3–24, January 22, 1998. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/532073.

Cooper:1998:RHWa

[1394] M. A. R. Cooper. Robert Hooke’s work as surveyor for the city of London in the aftermath of the Great Fire. Part two: certification of areas of ground taken away for streets and other new works. Notes and Records of the Royal Society of London, 52(1):25–38, January 22, 1998. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http: //www.jstor.org/stable/532074. See also Parts 1 and 3 [1366, 1415].

Heyman:1998:HCP

[1395] Jacques Heyman. Hooke’s cubico–parabolical conoid. Notes and Records of the Royal Society of London, 52(1):39–50, January 22, 1998. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/532075.

Hall:1998:INA

[1396] A. R. Hall. Isaac Newton and the aerial nitre. Notes and Records of the Royal Society of London, 52(1):51–61, January 22, 1998. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/532076.

Hills:1998:JWM

[1397] R. L. Hills. John Watt’s map of the Clyde. Notes and Records of the Royal Society of London, 52(1):63–71, January 22, 1998. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/532077.

Miller:1998:HCW

[1398] D. P. Miller. The ‘Hardwicke circle’: the Whig supremacy and its demise in the 18th-century Royal Society. Notes and Records of the Royal Society of London, 52(1):73–91, January 22, 1998. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532078.

Tee:1998:RDF

[1399] G. J. Tee. Relics of Davy and Faraday in New Zealand. Notes and Records of the Royal Society of London, 52(1):93–102, January 22, 1998. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/532079.

Bridge:1998:SWB

[1400] J. A. Bridge. Sir William Brooke O’Shaughnessy, M.D., F.R.S., F.R.C.S., F.S.A.: a biographical appreciation by an electrical engineer. Notes and Records of the Royal Society of London, 52(1):103–120, January 22, 1998. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/532080.

Bridges:1998:LDL

[1401] P. Bridges. A long-distance Londoner. Notes and Records of the Royal Society of London, 52(1):121–129, January 22, 1998. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/532081.

Bleaney:1998:CZE

[1402] B. Bleaney. Centenary of the Zeeman effect. Notes and Records of the Royal Society of London, 52(1):131–136, January 22, 1998. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/532082.

Watt:1998:LPFa

[1403] D. C. Watt. Lionel Penrose, F.R.S. (1898–1972) and eugenics: Part one. Notes and Records of the Royal Society of London, 52(1):137–151, January 22, 1998. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/532083. See

also Part 2 [1423].

King-Hele:1998:WLE

[1404] D. King-Hele. The 1997 Wilkins Lecture: Erasmus Darwin, the Lunaticks and evolution. Notes and Records of the Royal Society of London, 52(1): 153–180, January 22, 1998. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/532084.

See erratum [1437].

Klug:1998:APS

[1405] A. Klug. Address of the President, Sir Aaron Klug, O.M., P.R.S., given at the Anniversary Meeting on 1 December 1997. Notes and Records of the Royal Society of London, 52(1):181–190, January 22, 1998. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/532085.

Cook:1998:BRG

[1406] Alan Cook, F.R.S. Book review: Greenwich comes of age: Eric G. Forbes, Lesley Murdin and Frances Willmoth, The correspondence of John Flamsteed, The first Astronomer Royal. Volume two, 1682–1703. Institute of Physics Publishing, 1997. Pp.xlvii + 1095, £140/$280. ISBN 0-7503-0391-3, hardback. Notes and Records of the Royal Society of London, 52(1):191–193, January 22, 1998. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

532086.

Cook:1998:BR

[1407] Alan Cook, F.R.S. Book reviews. Notes and Records of the Royal Society of London, 52(1):191–201, January 22, 1998. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Rowlinson:1998:BRA

[1408] J. S. Rowlinson, F.Eng., Sec.R.S. Book review: Aristocratic physics: Christa Jungnickel and Russell McCormmach, Cavendish. Memoirs of the American Philosophical Society, Philadelphia, Volume 220, 1996. Pp xi + 414. $50.00. ISBN 0-87169-220-1. Notes and Records of the Royal Society of London, 52(1):194–196, January 22, 1998. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532087.

Russell:1998:BRP

[1409] Colin A. Russell. Book review: Photographic pioneer: Larry J. Schaaf, Records of the Dawn of Photography, Talbot’s Notebooks P & Q, Cambridge University Press, 1996. Pp. xxxv + 413, £90. ISBN 0-521-44051-3, hardback. Notes and Records of the Royal Society of London, 52(1):196– 197, January 22, 1998. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/532088.

Brush:1998:BRQ

[1410] Stephen G. Brush. Book review: Quantifying singularities: Cyril Domb, The Critical Point: A Historical Introduction to the Modern Theory of Critical Phenomena. Foreword by Michael E. Fisher. Taylor & Francis, 1996. Pp. xviii + 376. ISBN 0-7484-0435-X, hardback. Notes and Records of the Royal Society of London, 52(1):198–200, January 22, 1998. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/532089.

Cutland:1998:BRD

[1411] Nigel Cutland. Book review: Discovery of non-standard analysis: Joseph Warren Dauben, Abraham Robinson: The Creation of Nonstandard Analysis; A Personal and Mathematical Odyssey. Princeton University Press, 1995, ISBN 0-691-03745-0. Notes and Records of the Royal Society of London, 52(1):200–201, January 22, 1998. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532090.

Anonymous:1998:BMa

[1412] Anonymous. Back matter. Notes and Records of the Royal Society of London, 52(1):??, January 1998. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

532091.

Anonymous:1998:FMa

[1413] Anonymous. Front matter. Notes and Records of the Royal Society of London, 52(1):??, January 1998. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

532071.

Cook:1998:I

[1414] Alan Cook, F.R.S. In this issue. Notes and Records of the Royal Society of London, 52(2):203, July 22, 1998. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531856.

Cooper:1998:RHWb

[1415] M. A. R. Cooper. Robert Hooke’s work as surveyor for the city of London in the aftermath of the Great Fire. Part three: settlement of disputes and complaints arising from rebuilding. Notes and Records of the Royal Society of London, 52(2):205–220, July 22, 1998. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/ stable/531857. See also Parts 1 and 2 [1366, 1394].

Rowlinson:1998:OCR

[1416] J. S. Rowlinson. ‘Our common room in Geneva’ and the early exploration of the Alps of Savoy. Notes and Records of the Royal Society of London, 52(2):221–235, July 22, 1998. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531858.

Keynes:1998:PVS

[1417] M. Keynes. The Portland vase: Sir William Hamilton, Josiah Wedgwood and the Darwins. Notes and Records of the Royal Society of London, 52 (2):237–259, July 22, 1998. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531859.

Cook:1998:DED

[1418] G. C. Cook and D. King-Hele. Doctor Erasmus Darwin’s death in 1802. Notes and Records of the Royal Society of London, 52(2):261–265, July 22, 1998. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531860.

Crilly:1998:YAC

[1419] Tony Crilly. The young Arthur Cayley. Notes and Records of the Royal Society of London, 52(2):267–282, July 22, 1998. CODEN NOREAY.

ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/531861.

Millett:1998:IRU

[1420] D. Millett. Illustrating a revolution: an unrecognized contribution to the ‘golden era’ of cerebral localization. Notes and Records of the Royal Society of London, 52(2):283–305, July 22, 1998. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531862.

Dasgupta:1998:JBA

[1421] S. Dasgupta. Jagadis Bose, Augustus Waller and the discovery of ‘vegetable electricity’. Notes and Records of the Royal Society of London, 52 (2):307–322, July 22, 1998. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531863.

Boney:1998:SDB

[1422] A. D. Boney. The summer of 1914: diary of a botanist. Notes and Records of the Royal Society of London, 52(2):323–338, July 22, 1998. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531864.

Watt:1998:LPFb

[1423] D. C. Watt. Lionel Penrose, F.R.S. (1898–1972) and eugenics. Part two. Notes and Records of the Royal Society of London, 52(2):339–354, July 22, 1998. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531865. See also Part 1

[1403].

Postgate:1998:OUN

[1424] J. Postgate. The origins of the unit of nitrogen fixation at the University of Sussex. Notes and Records of the Royal Society of London, 52(2): 355–362, July 22, 1998. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531866.

Russell:1998:BRD

[1425] Colin A. Russell. Book review: Debunking ‘nature’: Robert Boyle. A free enquiry into the vulgarly received notion of nature. Edited by Edward B. Davis and Michael Hunter. Cambridge Texts in the History of Philosophy, Cambridge University Press, 1996. Pp. xxxvi + 171, £13.95 (paperback), ISBN 0-521-56796-3; £37.50 (hardback). ISBN 0-521-56100-0. Notes and Records of the Royal Society of London, 52(2):363–364, July 22, 1998.

CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531867.

Russell:1998:BR

[1426] Colin A. Russell. Book reviews. Notes and Records of the Royal Society of London, 52(2):363–379, July 22, 1998. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Iliffe:1998:BRM

[1427] Rob Iliffe. Book review: Mathematics as a science: P. Mancosu, Philosophy of mathematics and mathematical practice in the seventeenth century, Oxford University Press, 1996. Pp. vii + 275, £45.00 (hardback). ISBN 0-19-508463-2. Notes and Records of the Royal Society of London, 52(2): 364–365, July 22, 1998. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531868.

Dolan:1998:BRH

[1428] Brian P. Dolan. Book review: How opposite forces attract: Patricia Fara, Sympathetic attractions: magnetic practices, beliefs, and symbolism in eighteenth-century England. Princeton University Press, 1996. Pp. xiii + 326, £35 (hardback). ISBN 0-691-01099-4. Notes and Records of the Royal Society of London, 52(2):366–369, July 22, 1998. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531869.

Hall:1998:BRP

[1429] A. Rupert Hall, F.B.A. Book review: A practical natural philosopher: Alan Cook, Edmond Halley: charting the heavens and the seas. Clarendon Press, Oxford, 1998. Pp. 540 + xvi, £29.95 (hardback). ISBN 0-19-850031-9. Notes and Records of the Royal Society of London, 52(2): 369–371, July 22, 1998. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531870.

Taub:1998:BRC

[1430] L. Taub. Book review: Collecting curiosities: Jan Bondeson, A cabinet of medical curiosities. I. B. Tauris Publishers, 1997. Pp. ix + 250, £10.95 (softback). ISBN 1-86064-228-4. Notes and Records of the Royal Society of London, 52(2):371–372, July 22, 1998. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531871.

Weatherall:1998:BRD

[1431] D. J. Weatherall, F.R.S. Book review: Distinguished men of medicine: J. M. H. Moll, Presidents of The Royal Society of Medicine. Royal Society of Medicine Press, 1996. Pp. xlii + 552, £65.00 (hardback). ISBN 1-85315-264-1. Notes and Records of the Royal Society of London, 52(2): 372–374, July 22, 1998. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531872.

Rowlinson:1998:BRE

[1432] J. S. Rowlinson, F.Eng., Sec.R.S. Book review: An entrepreneurial chemist: W. H. Brock, Justus von Liebig: the chemical gatekeeper. Cambridge University Press, 1997. Pp. xv + 374, £50, $79.95 (hardback). ISBN 0-521-56224-4. Notes and Records of the Royal Society of London, 52(2):374–375, July 22, 1998. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531873.

Craig:1998:BRE

[1433] G. Y. Craig. Book review: The essence of Lyell: Charles Lyell’s Principles of Geology (1830–33). Edited by J. Secord. Penguin Classics, London, 1997. Pp. 472 + xlvii, £9.99 (softback). ISBN 0-14-043528-X. Notes and Records of the Royal Society of London, 52(2):375, July 22, 1998. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531874.

Howie:1998:BRR

[1434] Archie Howie, F.R.S. Book review: Roentgen’s legacy: X-rays — the first hundred years. Edited by Alan Michette and Slawka Pfauntsch. J. Wiley, 1996. Pp. xiii + 262, £37.50 (hardback). ISBN 0-471-96502-2. Notes and Records of the Royal Society of London, 52(2):376–377, July 22, 1998. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531875.

Pippard:1998:BRP

[1435] Sir Brian Pippard, F.R.S. Book review: The past is a foreign country: Jack Morrell, Science at Oxford 1914–1939. Clarendon Press, Oxford, 1997. Pp. 473, £55 (hardback). ISBN 0-19-820657-7. Notes and Records of the Royal Society of London, 52(2):377–378, July 22, 1998. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531876.

Russell:1998:BRS

[1436] Colin A. Russell. Book review: Science in its human context: John Polkinghorne, Beyond science. The wider human context. Cambridge University Press, 1996. Pp. xii + 131, £13.95 (hardback). ISBN 0-521-57212-6. Notes and Records of the Royal Society of London, 52(2):378–379, July 22, 1998. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531877.

Anonymous:1998:EWL

[1437] Anonymous. Erratum: The 1997 Wilkins Lecture: Erasmus Darwin, the Lunaticks and Evolution. Notes and Records of the Royal Society of London, 52(2):381–382, July 1998. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/ 531878. See [1404].

Anonymous:1998:AIN

[1438] Anonymous. Author index to Notes and Records, volume 52, 1998. Notes and Records of the Royal Society of London, 52(2):383–384, July 1998. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531854.

Anonymous:1998:BMb

[1439] Anonymous. Back matter. Notes and Records of the Royal Society of London, 52(2):??, July 1998. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531879.

Anonymous:1998:FMb

[1440] Anonymous. Front matter. Notes and Records of the Royal Society of London, 52(2):??, July 1998. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531855.

Cook:1999:MFI

[1441] Alan Cook, F.R.S. More frequent issues. Notes and Records of the Royal Society of London, 53(1):1–2, January 22, 1999. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/531924.

Wilkes:1999:HSS

[1442] M. V. Wilkes. Historical studies in science and technology and the uses to which they can be put. Notes and Records of the Royal Society of London, 53(1):3–10, January 22, 1999. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531925.

Thomas:1999:SBT

[1443] J. M. Thomas. Sir Benjamin Thompson, count Rumford and the Royal Institution. Notes and Records of the Royal Society of London, 53(1): 11–25, January 22, 1999. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531926.

Chambers:1999:LPC

[1444] N. Chambers. Letters from the President: the correspondence of Sir Joseph Banks. Notes and Records of the Royal Society of London, 53(1): 27–57, January 22, 1999. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531927.

Elliott:1999:ABF

[1445] P. Elliott. Abraham Bennet, F.R.S. (1749–1799): a provincial electrician in eighteenth-century England. Notes and Records of the Royal Society of London, 53(1):59–78, January 22, 1999. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531928.

Ross:1999:JDL

[1446] S. Ross. John Dalton’s Lakeland excursions. Notes and Records of the Royal Society of London, 53(1):79–94, January 22, 1999. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531929.

Appleby:1999:WGF

[1447] J. H. Appleby. Woronzow Greig (1805–1865), F.R.S., and his scientific interests. Notes and Records of the Royal Society of London, 53(1):95– 106, January 22, 1999. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531930.

Allibone:1999:PRS

[1448] T. E. Allibone, C.B.E., F.Eng., F.R.S. Philately and the Royal Society. II. Notes and Records of the Royal Society of London, 53(1):107–119, January 22, 1999. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/531931.

Burcham:1999:CHV

[1449] W. E. Burcham. The Cavendish high-voltage laboratory 1935–39. Notes and Records of the Royal Society of London, 53(1):121–134, January 22, 1999. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531932.

Plesch:1999:SFC

[1450] P. H. Plesch. Symposium on Friedel–Crafts catalysts and polymerization. Notes and Records of the Royal Society of London, 53(1):135–141, January 22, 1999. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531933.

Butler:1999:RPB

[1451] C. Butler. Recollections of Patrick Blackett 1945–70. Notes and Records of the Royal Society of London, 53(1):143–156, January 22, 1999. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531934.

Klug:1999:APS

[1452] A. Klug. Address of the President, Sir Aaron Klug, O.M., P.R.S., given at the Anniversary Meeting on 30 November 1998. Notes and Records of the Royal Society of London, 53(1):157–167, January 22, 1999. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531935.

Shea:1999:BRT

[1453] William Shea. Book review: Treasure troves: Archives of the Scientific Revolution. The formation and exchange of ideas in seventeenthcentury Europe, Michael Hunter (ed.). Woodbridge, Boydell Press, 1998. Pp. xiii + 216. £45.00. ISBN 0-85115-553-7. Notes and Records of the Royal Society of London, 53(1):169–170, January 22, 1999. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531936.

Shea:1999:BR

[1454] W. Shea. Book reviews. Notes and Records of the Royal Society of London, 53(1):169–177, January 22, 1999. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Chapman:1999:BRO

[1455] Allan Chapman. Book review: Our astronomical observator: Flamsteed’s stars, new perspectives on the life and work of the first Astronomer Royal (1646–1719), Frances Willmoth (ed.). Boydell Press and National Maritime Museum, London, 1997. Pp. 271 + 14 b/w plates, catalogue of Flamsteed papers, index. £45.00. ISBN 0-85115-706-8. Notes and Records of the Royal Society of London, 53(1):170–171, January 22, 1999. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531937.

Bennett:1999:BRG

[1456] Jim Bennett. Book review: Geomagnetic instruments in Italy: M. Basso Ricci, L. Cafaraella, A. Meloni and P. Tucci, Due Secoli di Strumenti Geomagnetici in Italia (1740–1971). Two Centuries of Geomagnetic Instruments in Italy (with a summary in English). Bologna, Editrice Compositori, 1997. Pp. 234 pp, ISBN 88-7794-091-3. Notes and Records of the Royal Society of London, 53(1):172, January 22, 1999. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531938.

Hayman:1999:BRC

[1457] W. K. Hayman, F.R.S. Book review: Complex integration: Frank Smithies, Cauchy and the creation of Complex Function Theory. Cambridge University Press, 1997. Pp. 216, £35 (hardback). ISBN 0-521-59278-X. Notes and Records of the Royal Society of London, 53(1):173– 174, January 22, 1999. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531939.

Grafen:1999:BRD

[1458] Alan Grafen. Book review: Darwin in 1862: The correspondence of Charles Darwin, volume 10, 1862. Frederick Burkhardt, Duncan M. Porter, Joy Harvey and Jonathan R. Topham (eds). Cambridge University Press, 1997. Pp. xxxvii + 936. £50. ISBN 0-521-59032-9. Notes and Records of the Royal Society of London, 53(1):175–176, January 22, 1999. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531940.

Squires:1999:BRH

[1459] Gordon Squires. Book review: A history of the electron: Per F. Dahl, Flash of the cathode rays: A history of J. J. Thomson’s electron. Institute of Physics Publishing, Bristol and Philadelphia, 1997. Pp. xvii + 526. £29.50 (hardback). ISBN 0-7503-0453-7. Notes and Records of the Royal Society of London, 53(1):176–177, January 22, 1999. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531941.

Anonymous:1999:BMa (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531942.

Anonymous:1999:FMa

[1461] Anonymous. Front matter. Notes and Records of the Royal Society of London, 53(1):??, January 1999. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531923.

Cook:1999:OFC

[1462] Alan Cook. Our foreign correspondence. Notes and Records of the Royal Society of London, 53(2):179–182, May 22, 1999. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/532203.

Nixon:1999:RBS

[1463] M. Nixon. The Raymond and Beverly Sackler archive resource at The Royal Society: a work in progress. Notes and Records of the Royal Society of London, 53(2):183–186, May 22, 1999. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532204.

Avramov:1999:ASC

[1464] I. Avramov. An apprenticeship in scientific communication: the early correspondence of Henry Oldenburg (1656–63). Notes and Records of the Royal Society of London, 53(2):187–201, May 22, 1999. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/532205.

Govier:1999:RSS

[1465] M. Govier. The Royal Society, slavery and the island of Jamaica: 1660– 1700. Notes and Records of the Royal Society of London, 53(2):203–217, May 22, 1999. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/532206.

Appleby:1999:SAC

[1466] R. H. Appleby. Sir Alexander Crichton, F.R.S. (1763–1856), imperial Russian physician at large. Notes and Records of the Royal Society of London, 53(2):219–230, May 22, 1999. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532207.

Yamalidou:1999:JTRa

[1467] M. Yamalidou. John Tyndall, the rhetorician of molecularity. Part one. Crossing the boundary towards the invisible. Notes and Records of the Royal Society of London, 53(2):231–242, May 22, 1999. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/532208. See also part two [1481].

Harris:1999:HFD

[1468] H. Harris. Howard Florey and the development of penicillin. Notes and Records of the Royal Society of London, 53(2):243–252, May 22, 1999. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/532209.

Anderson:1999:PBIa

[1469] R. S. Anderson. Patrick Blackett in India: military consultant and scientific intervenor, 1947–72. Part one. Notes and Records of the Royal Society of London, 53(2):253–273, May 22, 1999. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/ stable/532210. See also Part 2 [1483].

Hamblyn:1999:ERA

[1470] R. Hamblyn. Essay review. the alum-maker’s secret [Roger Osborne, The floating egg: episodes in the making of geology. London, Jonathan Cape, 1998. Pp. 372, £15.99 (hardback). ISBN 0-224-05028-1]. Notes and Records of the Royal Society of London, 53(2):275–278, May 22, 1999. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/532211.

Pippard:1999:BRG

[1471] Brian Pippard, F.R.S. Book review: Genius in context: P. M. Harman, The natural philosophy of James Clerk Maxwell. Cambridge University Press, 1998. Pp. xiv + 232, £35/US$59.95 (hardback). ISBN 0-521-56102-7. Notes and Records of the Royal Society of London, 53(2): 279–281, May 22, 1999. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/532212.

Pippard:1999:BR

[1472] Brian Pippard, F.R.S. Book reviews. Notes and Records of the Royal Society of London, 53(2):279–289, May 22, 1999. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Rowlinson:1999:BRE

[1473] J. S. Rowlinson, F.Eng., Sec.R.S. Book review: Energy in the 19th Century: Crosbie Smith, The science of energy. A cultural history of energy physics in Victorian Britain. London, Athlone Press, 1998. Pp. xii + 404, £19.95 (softback). ISBN 0-485-12145-X. Notes and Records of the Royal Society of London, 53(2):281–282, May 22, 1999. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/532213.

Thouless:1999:BRT

[1474] David Thouless, F.R.S. Book review: Two pioneers of modern theoretical physics: H. A. Bethe, Selected works of Hans A. Bethe, with commentary. World Scientific Publishing Company, 1997. Pp. viii + 605, £47.00 (hardback). ISBN 981-02-2876-7. R. H. Dalitz and Sir Rudolf Peierls, Selected scientific papers of Sir Rudolf Peierls, with commentary. World Scientific Publishing Company and Imperial College Press, 1997. Pp. xxiii + 805, £67.00 (hardback). ISBN 981-02-2692-6. Notes and Records of the Royal Society of London, 53(2):282–288, May 22, 1999. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/532214.

Bleaney:1999:BRA

[1475] B. Bleaney, F.R.S. Book review: Atomic energy in Argentina: Mario A. J. Mariscotti, El secreto atómico de Humuel. Cronica del origen de la energia atómica en la Argentina. Estudio Sigma S.R.L., Buenos Aires, 1996. Pp 286. ISBN 950-9446-24-6. Notes and Records of the Royal Society of London, 53(2):288–289, May 22, 1999. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532215.

Anonymous:1999:BMb

[1476] Anonymous. Back matter. Notes and Records of the Royal Society of London, 53(2):??, May 1999. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/532216.

Anonymous:1999:FMb

[1477] Anonymous. Front matter. Notes and Records of the Royal Society of London, 53(2):??, May 1999. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

532202.

Cook:1999:FT

[1478] A. Cook, F.R.S. Foreign travels. Notes and Records of the Royal Society of London, 53(3):291–293, September 22, 1999. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/41236969.

Shaw:1999:JSF

[1479] T. Shaw. John Swinton, F.R.S., identified as the author of a 1734 travel journal. Notes and Records of the Royal Society of London, 53 (3):295–304, September 22, 1999. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

41236970.

McConnell:1999:NEP

[1480] A. McConnell and A. Brech. Nathaniel and Edward Pigott, itinerant astronomers. Notes and Records of the Royal Society of London, 53 (3):305–318, September 22, 1999. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

41236971.

Yamalidou:1999:JTRb

[1481] M. Yamalidou. John Tyndall, the rhetorician of molecularity. Part two. Questions put to nature. Notes and Records of the Royal Society of London, 53(3):319–331, September 22, 1999. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/ stable/41236972. See also part one [1467].

Bleaney:1999:CPO

[1482] B. Bleaney, F.R.S. A century of physics in Oxford. Notes and Records of the Royal Society of London, 53(3):333–343, September 22, 1999. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/41236973.

Anderson:1999:PBIb

[1483] R. S. Anderson. Patrick Blackett in India: Military consultant and scientific intervenor, 1947–72. Part Two. Notes and Records of the Royal Society of London, 53(3):345–360, September 22, 1999. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/41236974. See also Part 1 [1469].

Keynes:1999:ERL

[1484] M. Keynes. Essay review. Lancelot Hogben, F.R.S. (1895–1975): A review of his autobiography [Adrian and Anne Hogben, Lancelot Hogben — scientific humanist. Merlin Press, Woodbridge, Suffolk, 1998. Pp. 254, £14.95 (hardback). ISBN 0-85036-470-1]. Notes and Records of the Royal Society of London, 53(3):361–369, September 22, 1999. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/41236975.

Russell:1999:BRS

[1485] C. A. Russell. Book review: The secret alchemist: L. M. Principe, The aspiring adept: Robert Boyle and his alchemical quest, Princeton University Press, Princeton, 1998. Pp. xiv + 339, £32.50/US$45.00 (hardback). ISBN 0-691-01678-X. Notes and Records of the Royal Society of London, 53(3):371–372, September 22, 1999. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

41236976.

Russell:1999:BR

[1486] C. A. Russell. Book reviews. Notes and Records of the Royal Society of London, 53(3):371–376, September 22, 1999. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Meer:1999:BRS

[1487] Sarah Meer. Book review: Servant of the state, promoter of empire: John Gascoigne, Science in the service of Empire: Joseph Banks, the British State and the uses of science in the Age of Revolution, Cambridge University Press, 1998. Pp. iv + 252, £40.00 (hardback). ISBN 0-521-55069-6. Notes and Records of the Royal Society of London, 53 (3):372–373, September 22, 1999. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

41236977.

Rowlinson:1999:BRP

[1488] J. S. Rowlinson, F.R.Eng., Sec.R.S. Book review: The paradox of irreversibility: Carlo Cercignani, Ludwig Boltzmann: the man who trusted atoms. Oxford University Press, 1998. Pp. xvii + 329, £29.50 (hardback). ISBN 0-19-850154-4. Notes and Records of the Royal Society of London, 53(3):374, September 22, 1999. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

41236978.

Pippard:1999:BRS

[1489] Sir Brian Pippard, F.R.S. Book review: Stumbling towards the light: E. A. Davis and I. J. Falconer, J. J. Thomson and the discovery of the electron. Taylor & Francis, London, 1997. Pp. xxvii + 243, £16.95 (softback). ISBN 0-7484-0720-0. Notes and Records of the Royal Society of London, 53(3):375–376, September 22, 1999. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

41236979.

Anonymous:1999:BMc

[1490] Anonymous. Back matter. Notes and Records of the Royal Society of London, 53(3):??, September 1999. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

41236980.

Anonymous:1999:FMc

[1491] Anonymous. Front matter. Notes and Records of the Royal Society of London, 53(3):??, September 1999. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

41236968.

Cook:2000:Ia

[1492] A. Cook. In this issue. Notes and Records of the Royal Society of London, 54(1):1–3, January 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/532053.

Ford:2000:STC

[1493] B. J. Ford. Shining through the centuries: John Ray’s life and legacy. Notes and Records of the Royal Society of London, 54(1):5–22, January 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/532054.

Koinm:2000:CMU

[1494] A. J. Koinm. Christopher Merret’s use of experiment. Notes and Records of the Royal Society of London, 54(1):23–32, January 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/532055.

Rothman:2000:LHO

[1495] P. Rothman. By ‘the light of his own mind’: The story of James Ferguson, astronomer. Notes and Records of the Royal Society of London, 54(1): 33–45, January 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/532056.

Sykes:2000:FST

[1496] A. H. Sykes. Foster and Sharpey’s tour of Europe. Notes and Records of the Royal Society of London, 54(1):47–52, January 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/532057.

Hass:2000:RDW

[1497] Jr J. W. Hass. The Reverend Dr William Henry Dallinger, F.R.S. (1839– 1909). Notes and Records of the Royal Society of London, 54(1):53–65, January 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/532058.

Mallik:2000:REK

[1498] D. C. V. Mallik. The Raman effect and Krishnan’s diary. Notes and Records of the Royal Society of London, 54(1):67–83, January 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/532059.

Fogg:2000:RSA

[1499] G. E. Fogg. The Royal Society and the Antarctic. Notes and Records of the Royal Society of London, 54(1):85–98, January 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/532060.

Klug:2000:APS

[1500] A. Klug. Address of the President, Sir Aaron Klug, O.M., P.R.S., given at the Anniversary Meeting on 30 November 1999. Notes and Records of the Royal Society of London, 54(1):99–108, January 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/532061.

Cook:2000:ERG

[1501] A. Cook. Essay review: The gatekeeper of the modern world: [Niccolo Guicciardini, Reading the Principia; The debate on Newton’s mathematical methods for natural philosophy from 1687 to 1736. Cambridge University Press, 1999. Pp. 285, £50.00. ISBN 0-521-64066-0]. Notes and Records of the Royal Society of London, 54(1):109–113, January 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/532062.

Russell:2000:BRA

[1502] Colin A. Russell. Book review: Antimony’s curious history: R. Ian Mc-Callum, Antimony in medical history: an account of the medical uses of antimony and its compounds since early times to the present. The Pentland Press, Edinburgh, Cambridge, Durham, 1999. Pp. xvi + 125, £15.00 (hardback). ISBN 1-85821-642-7. Notes and Records of the Royal Society of London, 54(1):115–116, January 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532063.

Russell:2000:BR

[1503] Colin A. Russell. Book reviews. Notes and Records of the Royal Society of London, 54(1):115–125, January 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Munk:2000:BRS

[1504] Walter Munk, For.Mem.R.S. Book review: Solving the tidal problem: David E. Cartwright, Tides: A scientific history. Cambridge University Press, 1998. Pp. xii + 292, £45.00. ISBN 0-521-62145-3. Notes and Records of the Royal Society of London, 54(1):116–119, January 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/532064.

Higgott:2000:BRW

[1505] Gordon Higgott. Book review: Wren’s architectural writings: Lydia M. Soo, Wren’s ‘Tracts’ on architecture and other writings. Cambridge University Press, 1998. Pp. 320, 63 illus. £45.00 (hardback). ISBN 0-521-57369-6. Notes and Records of the Royal Society of London, 54(1): 119–120, January 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/532065.

King-Hele:2000:BRD

[1506] Desmond King-Hele, F.R.S. Book review: Not so dumb, after all? ‘Did the Royal Society matter in the eighteenth century?’ Edited by R. Sorrenson. British Journal for the History of Science 32(113), 130–221 (June 1999) ISSN 0007-0874. Notes and Records of the Royal Society of London, 54(1):120–122, January 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

532066.

Rowlinson:2000:BRT

[1507] J. S. Rowlinson, F.R.Eng., F.R.S. Book review: The Third Law: D. K. Barkan, Walther Nernst and the transition to modern physical science. Cambridge University Press, 1999. Pp. xii + 288, £45.00 (hardback). ISBN 0-521-44456-X. Notes and Records of the Royal Society of London, 54(1):123–124, January 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

532067.

Dixon:2000:BRS

[1508] Henry B. F. Dixon. Book review: The Soviet bomb: Veniamin Tsukerman and Zinaida Azarkh, Arzamas-16. Soviet scientists in the Nuclear Age: a memoir. Translated by Timothy Sergay. Edited by Michael Pursglove. Bramcote Press, Nottingham, 1999. Pp. xxxiv + 182, £13.95 (paperback). ISBN 1-900405-04-0. Notes and Records of the Royal Society of London, 54(1):124–125, January 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532068.

Baker:2000:RPC

[1509] R. Baker. Recent publications concerning the Royal Society. Notes and Records of the Royal Society of London, 54(1):127–128, January 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/532069.

Anonymous:2000:BMa

[1510] Anonymous. Back matter. Notes and Records of the Royal Society of London, 54(1):??, January 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

532070.

Anonymous:2000:FMa

[1511] Anonymous. Front matter. Notes and Records of the Royal Society of London, 54(1):??, January 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

532052.

Cook:2000:Ib

[1512] A. Cook. In this issue. Notes and Records of the Royal Society of London, 54(2):129–130, May 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531962.

Beretta:2000:SWS

[1513] M. Beretta. At the source of Western science: The organization of experimentalism at the Accademia del Cimento (1657–1667). Notes and Records of the Royal Society of London, 54(2):131–151, May 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531963.

McMahon:2000:JRA

[1514] S. McMahon. John Ray (1627–1705) and the Act of Uniformity 1662. Notes and Records of the Royal Society of London, 54(2):153–178, May 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531964.

Kusukawa:2000:HP

[1515] S. Kusukawa. The Historia Piscium (1686). Notes and Records of the Royal Society of London, 54(2):179–197, May 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531965.

Brooke:2000:WMN

[1516] J. H. Brooke. ‘Wise men nowadays think otherwise’: John Ray, natural theology and the meanings of anthropocentrism. Notes and Records of the Royal Society of London, 54(2):199–213, May 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531966.

Blouet:2000:BEF

[1517] Olwyn M. Blouet. Bryan Edwards, F.R.S., 1743–1800. Notes and Records of the Royal Society of London, 54(2):215–222, May 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531967.

Craik:2000:JIF

[1518] Alex D. D. Craik. James Ivory, F.R.S., mathematician: “the most unlucky person that ever existed”. Notes and Records of the Royal Society of London, 54(2):223–247, May 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531968.

Schroder:2000:JKK

[1519] W. Schröder and K.-H. Wiederkehr. Johann Kiessling, the Krakatoa event and the development of atmospheric optics after 1883. Notes and Records of the Royal Society of London, 54(2):249–258, May 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531969.

Fara:2000:BRM

[1520] Patricia Fara. Book review: The mystery of music: Penelope Gouk, Music, science and natural magic in seventeenth-century England. Yale University Press, New Haven and London, 1999. Pp. xii + 308, £30.00 (hardback). ISBN 0-300-07383-6. Notes and Records of the Royal Society of London, 54(2):259–260, May 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531970.

Fara:2000:BR

[1521] Patricia Fara. Book reviews. Notes and Records of the Royal Society of London, 54(2):259–272, May 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Williams:2000:BRP

[1522] Iwan Williams. Book review: Portraits of portents: J. M. Olson and J.M. Pasachoff, Fire in the sky: comets and meteors, the decisive centuries, in British art and science. Cambridge University Press, 1999. Pp 383, £18.95. ISBN 0-521-66359-8. Notes and Records of the Royal Society of London, 54(2):260–261, May 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531971.

Cook:2000:BRV

[1523] Sir Alan Cook, F.R.S. Book review: A view from Florence: Nuncius, Annali di Storia della Scienza. Anno XIV, fasc. 1. Olschki, Firenze, 1999. Pp. 417, Lit. 200 000 per year outside Italy. Notes and Records of the Royal Society of London, 54(2):261–262, May 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531972.

Ford:2000:BRC

[1524] Brian J. Ford. Book review: Crossing boundaries: Desmond King-Hele, Erasmus Darwin, a life of unequalled achievement. Giles de la Mere, London, 1999. Pp. x + 422, £24.99. ISBN 1-900357-08-9. Notes and Records of the Royal Society of London, 54(2):262–265, May 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531973.

Hamblyn:2000:BRJ

[1525] Richard Hamblyn. Book review: John Constable, meteorologist: John E. Thornes, John Constable’s skies: a fusion of art and science. Birmingham, University Press, 1999. Pp. 288, £40.00 (paperback). ISBN 1-902459-02-4. Notes and Records of the Royal Society of London, 54(2): 265–266, May 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531974.

Rowlinson:2000:BRN

[1526] J. S. Rowlinson, F.R.Eng., F.R.S. Book review: A northern chemist: Colin A. Russell, Edward Frankland: chemistry, controversy and conspiracy in Victorian England. Cambridge University Press, 1996. Pp. xx + 535, £65.00. ISBN 0-521-49636-5. Notes and Records of the Royal Society of London, 54(2):266–267, May 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531975.

Olliver:2000:BRE

[1527] Joe Olliver. Book review: Everest: J. R. Smith, Everest: the man and the mountain. Whittles Publishing, Caithness, 1999. Pp. xiv + 306, £37.50 (softback). ISBN 1-870325-72-9. Notes and Records of the Royal Society of London, 54(2):267–268, May 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531976.

Clarke:2000:BRU

[1528] Patricia H. Clarke, F.R.S. Book review: Unusual and enterprising: Barbara T. Gates, Kindred nature: Victorian and Edwardian women embrace the living world. University of Chicago Press, 1998. Pp. 239, £14.00. IBSN 0-226-28443-3. Notes and Records of the Royal Society of London, 54(2):269–270, May 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531977.

Oswald:2000:BRR

[1529] Philip H. Oswald. Book review: Rum on rum: Karl Sabbagh, A rum affair. London, Allen Lane, The Penguin Press, 1999. Pp. x + 224, £16.99 (hardback). ISBN 0-7139-9277-8. Notes and Records of the Royal Society of London, 54(2):270–272, May 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531978.

Anonymous:2000:BMb

[1530] Anonymous. Back matter. Notes and Records of the Royal Society of London, 54(2):??, May 2000. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531979.

Anonymous:2000:FMb (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531961.

Cook:2000:EM

[1532] A. Cook. The end of the millennium. Notes and Records of the Royal Society of London, 54(3):273, September 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532181.

Malcolm:2000:PJP

[1533] Noel Malcolm. The publications of John Pell, F.R.S. (1611–1685): some new light and some old confusions. Notes and Records of the Royal Society of London, 54(3):275–292, September 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/532182.

Stedall:2000:CPCa

[1534] Jacqueline A. Stedall. Catching Proteus: the collaborations of Wallis and Brouncker. I. Squaring the circle. Notes and Records of the Royal Society of London, 54(3):293–316, September 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532183.

Stedall:2000:CPCb

[1535] Jacqueline A. Stedall. Catching Proteus: the collaborations of Wallis and Brouncker. II. Number problems. Notes and Records of the Royal Society of London, 54(3):317–331, September 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532184.

Smithies:2000:FPF

[1536] Frank Smithies. A forgotten paper on the fundamental theorem of algebra. Notes and Records of the Royal Society of London, 54(3):333–341, September 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/532185.

Estes:2000:DGH

[1537] G. Estes, K. T. Grant, and P. R. Grant. Darwin in the Galápagos: his footsteps through the archipelago. Notes and Records of the Royal Society of London, 54(3):343–368, September 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532186.

Thomas:2000:PRS

[1538] J. M. Thomas. Peterhouse, The Royal Society and molecular biology. Notes and Records of the Royal Society of London, 54(3):369–385, September 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/532187.

Perutz:2000:BRW

[1539] Max Perutz, O.M., C.H., C.B.E., F.R.S. Book review: Woman scientist’s role model: Georgina Ferry, Dorothy Hodgkin: a life. London, Granta Books, 1999. Pp. 432, £9.99. ISBN 1-86207-285-X. Notes and Records of the Royal Society of London, 54(3):387–389, September 22, 2000. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/532188.

Perutz:2000:BR

[1540] Max Perutz, O.M., C.H., C.B.E., F.R.S. Book reviews. Notes and Records of the Royal Society of London, 54(3):387–408, September 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Ogilvie:2000:BRG

[1541] Bridget M. Ogilvie, D.B.E. Book review: The great sage: J. D. Bernal: a life in science and politics. Edited by Brenda Swann and Francis Aprahamian. Vesso Books, London, 1998. Pp. 339, £25. ISBN 1-85984-854-0. Notes and Records of the Royal Society of London, 54(3):389–390, September 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/532189.

Cooke:2000:BRT

[1542] Anne Cooke. Book review: Telling stories of science: Max Perutz, I wish I’d made you angry earlier. Essays on science and scientists. Oxford University Press, Oxford, 1999. 370 Pp., £20.00. ISBN 0-19-850531-0. Notes and Records of the Royal Society of London, 54(3):390–392, September 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/532190.

Speiser:2000:BRL

[1543] David Speiser. Book review: Lives of Newton: A. Rupert Hall, Isaac Newton: eighteenth-century perspectives. Oxford University Press, 1999. Pp. 224, £60. ISBN 0-19-850364-4. Notes and Records of the Royal Society of London, 54(3):392–393, September 22, 2000. CODEN NOREAY.

ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/532191.

Thomas:2000:BRG

[1544] Sir John Meurig Thomas, F.R.Eng., F.R.S. Book review: A great character in science: G.I. Brown, Count Rumford. The extraordinary life of a scientific genius. Sutton Publishing, Stroud, 1999. Pp. ix + 182, £12.99 (hardback). ISBN 0-7509-2184-6. Notes and Records of the Royal Society of London, 54(3):393–395, September 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532192.

Pippard:2000:BRF

[1545] Sir Brian Pippard, F.R.S. Book review: Faraday’s letters: The correspondence of Michael Faraday, volume 4. Edited by Frank A. J. L. James. Institution of Electrical Engineers, London, 1999, Pp. lxv + 1003 (hardback), £75.00. ISBN 0-86341-251-3. Notes and Records of the Royal Society of London, 54(3):395–396, September 22, 2000. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/532193.

Rowlinson:2000:BRF

[1546] Sir John Rowlinson, F.R.Eng., F.R.S. Book review: Faraday in his own words: The philosopher’s tree: A selection of Michael Faraday’s writings. Compiled, with commentary, by Peter Day. Institute of Physics, Bristol, 1999. Pp. xii + 211. £16.00 (paperback). ISBN 0-7503-0571-1. Notes and Records of the Royal Society of London, 54(3):397, September 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/532194.

Headland:2000:BRN

[1547] R. K. Headland. Book review: The new Endeavour: Science and exploration in the Pacific. European voyages to the Southern Oceans in the 18th century. Edited by Margarette Lincoln. Boydell Press in association with the National Maritime Museum, London, 1998. Pp. xix + 228 pages, £35. ISBN 0-85115-721-1. Notes and Records of the Royal Society of London, 54(3):398–399, September 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532195.

King-Hele:2000:BRI

[1548] Desmond King-Hele, F.R.S. Book review: Are ideas free to roam? Gillian Beer, Open fields: science in cultural encounter. Oxford University Press, 1999. Pp. x + 341, £17 (paperback). ISBN 0-19-818369-0. Notes and Records of the Royal Society of London, 54(3):400–402, September 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/532196.

Pagel:2000:BRT

[1549] B. E. J. Pagel, F.R.S. Book review: From Tycho to Big Bang: Rocky Kolb, Blind watchers of the sky. Oxford University Press, 1999. Pp xii + 338, £8.99 (paperback). ISBN 0-19-286203-0. Notes and Records of the Royal Society of London, 54(3):403, September 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/532197.

Chapman:2000:BRA

[1550] Allan Chapman. Book review: Anthropomorphic nature? Jan Golinski, Making natural knowledge. Constructivism and the history of science. Cambridge University Press, 1998. Pp. 136, £37.50 (hardback), £13.95 (paperback). ISBN 0-521-44913-8. Notes and Records of the Royal Society of London, 54(3):404–407, September 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532198.

McMenamin:2000:BRS

[1551] Mark A. S. McMenamin. Book review: Out of the shadows: Simon Conway Morris, The crucible of Creation: the Burgess Shale and the rise of animals. Oxford University Press, 1999. Pp. xxii + 242, £8.99 (paperback). ISBN 0-19-286202-2. Notes and Records of the Royal Society of London, 54(3):407–408, September 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532199.

Cook:2000:UBF

[1552] A. Cook. URFs become FRS. Notes and Records of the Royal Society of London, 54(3):409–411, September 22, 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532200.

Anonymous:2000:AIN

[1553] Anonymous. Author index to Notes and Records, volume 54, 2000. Notes and Records of the Royal Society of London, 54(3):413–414, September 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/532179.

Anonymous:2000:BMc

[1554] Anonymous. Back matter. Notes and Records of the Royal Society of London, 54(3):??, September 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

532201.

Anonymous:2000:FMc

[1555] Anonymous. Front matter. Notes and Records of the Royal Society of London, 54(3):??, September 2000. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

532180.

Cook:2001:NM

[1556] A. Cook. The new millennium. Notes and Records of the Royal Society of London, 55(1):1–2, January 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532140.

Shea:2001:HSI

[1557] W. R. Shea. The history of science and the image of science. Notes and Records of the Royal Society of London, 55(1):3–7, January 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/532141.

Cook:2001:TRS

[1558] Alan Cook, F.R.S. Time and The Royal Society. Notes and Records of the Royal Society of London, 55(1):9–27, January 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/532142.

Ford:2001:RSM

[1559] B. J. Ford. The Royal Society and the microscope. Notes and Records of the Royal Society of London, 55(1):29–49, January 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/532143.

Budd:2001:RFR

[1560] G. E. Budd. Royal fossils: The Royal Society and progress in palaeontology. Notes and Records of the Royal Society of London, 55(1):51–67, January 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/532144.

Hide:2001:ZLS

[1561] Raymond Hide, C.B.E., F.R.S. Zenographic longitude systems and Jupiter’s differential rotation. Notes and Records of the Royal Society of London, 55(1):69–79, January 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532145.

Fogg:2001:RSS

[1562] G. E. Fogg, C.B.E., F.R.S. The Royal Society and the South Seas. Notes and Records of the Royal Society of London, 55(1):81–103, January 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/532146.

Thomas:2001:P

[1563] J. M. Thomas, F.R.Eng., F.R.S. Predictions. Notes and Records of the Royal Society of London, 55(1):105–117, January 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/532147.

Cook:2001:RW

[1564] Alan Cook, F.R.S. Royal weather. Notes and Records of the Royal Society of London, 55(1):119–127, January 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532148.

Cook:2001:PIE

[1565] Alan Cook, F.R.S. Plants illustrating exotic collections. Notes and Records of the Royal Society of London, 55(1):129–144, January 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/532149.

Debru:2001:HST

[1566] Claude Debru. History of science and technology in education and training in Europe. Notes and Records of the Royal Society of London, 55(1): 145–146, January 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/532150.

Cook:2001:I

[1567] Alan Cook. [Introduction]. Notes and Records of the Royal Society of London, 55(1):147, January 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

532151.

Cook:2001:CN

[1568] Alan Cook, F.R.S. Centenary of the NPL. Notes and Records of the Royal Society of London, 55(1):147–160, January 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Sainsbury:2001:OA

[1569] Lord Sainsbury. Opening address. Notes and Records of the Royal Society of London, 55(1):148–149, January 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532152.

Klug:2001:RSN

[1570] Aaron Klug. The Royal Society and the NPL. Notes and Records of the Royal Society of London, 55(1):150–152, January 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/532153.

Anonymous:2001:ACM

[1571] Anonymous. Abstracts from the centenary meeting of the National Physical Laboratory, 7 November 2000. Notes and Records of the Royal Society of London, 55(1):153–160, January 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532154.

Cahn:2001:BRM

[1572] Robert W. Cahn, F.R.S. Book review: A material view of The Royal Society: D. R. F. West and J. E. Harris, F.R.S., Metals and the Royal Society. IoM Communications, London, 1999. Pp. 787, £40/$80. ISBN 1-86125-028-2. Notes and Records of the Royal Society of London, 55(1): 161–163, January 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/532155.

Klug:2001:APS

[1573] Aaron Klug, O.M., P.R.S. Address of the President, Sir Aaron Klug, O.M., P.R.S., given at the Anniversary Meeting on 30 November 2000. Notes and Records of the Royal Society of London, 55(1):165–177, January 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/532156.

Taylor:2001:JEA

[1574] J. Taylor. JSTOR: an electronic archive from 1665. Notes and Records of the Royal Society of London, 55(1):179–181, January 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/532157.

Anonymous:2001:BMa

[1575] Anonymous. Back matter. Notes and Records of the Royal Society of London, 55(1):??, January 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

532158.

Anonymous:2001:FMa

[1576] Anonymous. Front matter. Notes and Records of the Royal Society of London, 55(1):??, January 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

532138.

Cook:2001:WC

[1577] A. Cook. Worldwide connections. Notes and Records of the Royal Society of London, 55(2):183, May 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

532093.

Simms:2001:DGN

[1578] D. L. Simms and P. L. Hinkley. David Gregory on Newton’s burning mirror. Notes and Records of the Royal Society of London, 55(2):185–190, May 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/532094.

Appleby:2001:MRF

[1579] J. H. Appleby. Mapping Russia: Farquharson, Delisle and the Royal Society. Notes and Records of the Royal Society of London, 55(2):191– 204, May 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/532095.

Hall:2001:CNL

[1580] A. Rupert Hall. Cambridge: Newton’s legacy. Notes and Records of the Royal Society of London, 55(2):205–226, May 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/532096.

Zinsser:2001:TNP

[1581] Judith P. Zinsser. Translating Newton’s Principia: the Marquise Du Châtelet’s revisions and additions for a French audience. Notes and Records of the Royal Society of London, 55(2):227–245, May 22, 2001.

CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/532097.

Keynes:2001:BNC

[1582] M. Keynes. A bogus Newtonian ‘curiosity’ at The Royal Society. Notes and Records of the Royal Society of London, 55(2):247–252, May 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/532098.

Grattan-Guinness:2001:CJJ

[1583] I. Grattan-Guinness. The contributions of J. J. Sylvester, F.R.S., to mechanics and mathematical physics. Notes and Records of the Royal Society of London, 55(2):253–265, May 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532099.

Singh:2001:NPP

[1584] R. Singh and F. Riess. The 1930 Nobel Prize for Physics: A close decision? Notes and Records of the Royal Society of London, 55(2):267–283, May 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/532100.

Bleaney:2001:DAJ

[1585] B. Bleaney. Derek Ainslie Jackson (1906–1982): Some recollections of a great European spectroscopist. Notes and Records of the Royal Society of London, 55(2):285–287, May 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532101.

Jardine:2001:MMS

[1586] L. Jardine. Monuments and microscopes: Scientific thinking on a grand scale in the early Royal Society. Notes and Records of the Royal Society of London, 55(2):289–308, May 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532102.

Porter:2001:WLM

[1587] R. S. Porter. The Wilkins Lecture 2000: Medical futures. Notes and Records of the Royal Society of London, 55(2):309–323, May 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/532103.

Honer:2001:ERG

[1588] J. Honer. Essay review: The genius of Soane. Notes and Records of the Royal Society of London, 55(2):325–329, May 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/532104.

Fogg:2001:BRC

[1589] G. E. Fogg, C.B.E., F.R.S. Book review: Cargoes of knowledge: Tony Rice, Voyages of discovery: three centuries of natural history exploration. The Natural History Museum, London; Scriptum Editions, London & Hong Kong, 2000. Pp. 335, £45.00 (hardback). ISBN 1-902686-02-0. Notes and Records of the Royal Society of London, 55(2):331–332, May 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/532105;   http:/
/www.jstor.org/stable/532109.

Fogg:2001:BR

[1590] G. E. Fogg, C.B.E., F.R.S. Book reviews. Notes and Records of the Royal Society of London, 55(2):331–346, May 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Grant:2001:BRD

[1591] Peter R. Grant, F.R.S. Book review: Darwin the naturalist: R. Keynes (ed.), Charles Darwin’s Zoology Notes & specimen lists from H.M.S. Beagle, Cambridge University Press, 2000. Pp. 430, £95.00 (hardback). ISBN 0-521-46569-9. Notes and Records of the Royal Society of London, 55(2): 332–335, May 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/532106;
 http://www.jstor.org/stable/532109.

Fara:2001:BRE

[1592] Patricia Fara. Book review: Enlightenment or enlightenment? Roy Porter, Enlightenment: Britain and the creation of the modern world. London, Penguin Press, 2000. Pp. xxiv + 728, £25.00 (hardback). ISBN 0-7139-9152-6. Notes and Records of the Royal Society of London, 55(2): 335–337, May 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/532107;
 http://www.jstor.org/stable/532109.

Hattersley-Smith:2001:BRL

[1593] G. F. Hattersley-Smith. Book review: Lost in the ice: Scott Cookman, Ice blink: the tragic fate of Sir John Franklin’s lost polar expedition.

John Wiley & Sons, New York, 2000. Pp. xii + 244, $24.95 USA/£17.95 (hardback). ISBN 0-471-37790-2. Notes and Records of the Royal Society of London, 55(2):337–338, May 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532108;   http://www.jstor.org/stable/532109.

Campbell-Kelly:2001:BRB

[1594] Martin Campbell-Kelly. Book review: Bits and wheels: Doron Swade, The cogwheel brain: Charles Babbage and the quest to build the first computer. Little, Brown & Co., London, 2000. Pp. x + 342, £14.99 (hardback). ISBN 0-316-64847-7. Notes and Records of the Royal Society of London, 55(2):339–340, May 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532109.

Briscoe:2001:BRC

[1595] Ted Briscoe. Book review: Children evolve languages. Terrence Deacon, The symbolic species: the co-evolution of language and the human brain. Penguin Books Ltd, London, 1998. Pp. 528, £8.99 (paperback). ISBN 0-14-026405-1. Notes and Records of the Royal Society of London, 55 (2):340–341, May 2001. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/532110.

Rowlinson:2001:BRP

[1596] Sir John Rowlinson, F.R.Eng., F.R.S. Book review: Portraits of scientists: Ludmilla Jordanova, Defining features: scientific and medical portraits, 1660–2000. Reaktion Books in association with the National Portrait Gallery, 2000. Pp. 192, £14.95 (softback). ISBN 1-86189-059-1. Notes and Records of the Royal Society of London, 55(2):342–343, May 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/532109;   http:/
/www.jstor.org/stable/532111.

Stuart:2001:BRE

[1597] J. T. Stuart, F.R.S. Book review: Not elected: Leslie C. Woods, Against the tide: an autobiographical account of a professional outsider. Institute of Physics, 2000. Pp. xiii + 319. Notes and Records of the Royal Society of London, 55(2):343–344, May 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532109;   http://www.jstor.org/stable/532112.

Postgate:2001:BRT

[1598] John Postgate, F.R.S. Book review: The thoughts of Laureate Watson: James D. Watson, A passion for DNA. Oxford University Press, 2000, Pp. xx + 250, £18.99 (hardback). ISBN 0-19-850697-X. Notes and Records of the Royal Society of London, 55(2):344–346, May 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

URL http://www.jstor.org/stable/532109;   http://www.jstor.
org/stable/532113.

Cook:2001:EDC

[1599] Alan Cook, F.R.S. Éditions du Comité des Travaux Historiques et Scientifiques. Notes and Records of the Royal Society of London, 55(2):347, May 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/532115.

Anonymous:2001:BMb

[1600] Anonymous. Back matter. Notes and Records of the Royal Society of London, 55(2):??, May 2001. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/532116.

Anonymous:2001:FMb

[1601] Anonymous. Front matter. Notes and Records of the Royal Society of London, 55(2):??, May 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

532092.

Cook:2001:VS

[1602] A. Cook. Varieties of science. Notes and Records of the Royal Society of London, 55(3):349, September 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531945.

Fisher:2001:RBM

[1603] N. R. R. Fisher. Robert Balle, merchant of Leghorn and Fellow of the Royal Society (ca. 1640–ca. 1734). Notes and Records of the Royal Society of London, 55(3):351–371, September 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531946.

Hunter:2001:WDR

[1604] M. Hunter and C. Littleton. The work-diaries of Robert Boyle: a newly discovered source and its Internet publication. Notes and Records of the Royal Society of London, 55(3):373–390, September 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531947.

Hammond:2001:NNM

[1605] C. Hammond. A note on Newton’s measurements of film thicknesses and their corresponding (interference) colours. Notes and Records of the Royal Society of London, 55(3):391–394, September 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531948.

Coley:2001:GFM

[1606] N. G. Coley. George Fordyce M.D., F.R.S. (1736–1802): physician– chemist and eccentric. Notes and Records of the Royal Society of London, 55(3):395–409, September 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531949.

Grattan-Guinness:2001:IGH

[1607] I. Grattan-Guinness. The interest of G. H. Hardy, F.R.S., in the philosophy and the history of mathematics. Notes and Records of the Royal Society of London, 55(3):411–424, September 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531950.

Wilson:2001:SLG

[1608] D. A. H. Wilson. Sea lions, greasepaint and the U-boat threat: Admiralty scientists turn to the music hall in 1916. Notes and Records of the Royal Society of London, 55(3):425–455, September 22, 2001. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531951.

Arndt:2001:IXR

[1609] U. W. Arndt. Instrumentation in X-ray crystallography: Past, present and future. Notes and Records of the Royal Society of London, 55(3):457– 472, September 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531952.

Cook:2001:EHM

[1610] A. Cook. Edmond Halley and the magnetic field of the Earth. Notes and Records of the Royal Society of London, 55(3):473–490, September 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531953.

Polkinghorne:2001:BRB

[1611] John Polkinghorne, F.R.S. Book review: Bombs, bits and worm-holes: John Archibald Wheeler and Kenneth Ford, Geons, black holes and quantum foam — a life in physics. W. W. Norton, New York, 1998. Pp. 380, $14.95 (paperback). ISBN 0-393-31991-1. Notes and Records of the Royal Society of London, 55(3):491, September 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/531954.

Anonymous:2001:BR

[1612] Anonymous. Book reviews. Notes and Records of the Royal Society of London, 55(3):491–497, September 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Hunter:2001:BRN

[1613] Michael Hunter. Book review: Natural law and moral law: Jon Parkin, Science, religion and politics in Restoration England: Richard Cumberland’s De Legibus Naturae. Boydell & Brewer (Woodbridge, Suffolk), for the Royal Historical Society, 1999. Pp. xi + 251, £40.00 (hardback). ISBN 0-86193-241-2. Notes and Records of the Royal Society of London, 55(3):492–493, September 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531955.

Watts:2001:BRF

[1614] Fraser Watts. Book review: Firmly I believe and truly …: Robert A. Hinde, Why gods persist: a scientific approach to religion. Routledge, London, 1999. Pp. 248, £45.00 (hardback). ISBN 0-415-20825-4. Notes and Records of the Royal Society of London, 55(3):493–494, September 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531956.

Pippard:2001:BRR

[1615] Sir Brian Pippard, F.R.S. Book review: Rutherford, from New Zealand: John Campbell, Rutherford, scientist supreme. ASS Publications, Christchurch, New Zealand. Pp. 494, $50 (hardback). ISBN 0-473-05700-X. Notes and Records of the Royal Society of London, 55(3):494– 495, September 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/531957.

Bleaney:2001:BRG

[1616] Brebis Bleaney, C.B.E., F.R.S. Book review: A German exodus: Jean Medawar and David A. Pyke, Hitler’s gift. Scientists who fled Nazi Germany. Richard Cohen Books, London, 2000. Pp. xx + 268 (hardback). ISBN 1-86066-172-6. Notes and Records of the Royal Society of London, 55(3):495–496, September 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531958.

Redhead:2001:BRC

[1617] Michael Redhead. Book review: A century of physics: Helge Kragh, The quantum generations: a history of physics in the twentieth century. Princeton University Press, 1999. Pp. xiv + 494, £18.95 (hardback). ISBN 0-691-01206-7. Notes and Records of the Royal Society of London, 55(3):496–497, September 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531959.

Anonymous:2001:AIN

[1618] Anonymous. Author index to Notes and Records, volume 55, 2001. Notes and Records of the Royal Society of London, 55(3):499–500, September 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531943.

Anonymous:2001:BMc

[1619] Anonymous. Back matter. Notes and Records of the Royal Society of London, 55(3):??, September 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531960.

Anonymous:2001:FMc

[1620] Anonymous. Front matter. Notes and Records of the Royal Society of London, 55(3):??, September 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

531944.

Cook:2002:AAL

[1621] A. Cook. Across the Alps: London and Bologna in the eighteenth century. Notes and Records of the Royal Society of London, 56(1):1–2, January 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/532118.

Cavazza:2002:ISB

[1622] M. Cavazza. The Institute of Science of Bologna and The Royal Society in the eighteenth century. Notes and Records of the Royal Society of London, 56(1):3–25, January 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

532119.

Richardson:2002:NNW

[1623] G. Richardson. A Norfolk network within The Royal Society. Notes and Records of the Royal Society of London, 56(1):27–39, January 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/532120.

King-Hele:2002:EDI

[1624] Desmond King-Hele. Erasmus Darwin’s improved design for steering carriages — and cars. Notes and Records of the Royal Society of London, 56(1):41–62, January 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

532121.

Pippard:2002:SFC

[1625] B. Pippard. Schoolmaster-Fellows and the campaign for science education. Notes and Records of the Royal Society of London, 56(1):63–81, January 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/532122.

Bleaney:2002:TOS

[1626] B. Bleaney. Two Oxford science professors, F. Soddy and J. S. E. Townsend. Notes and Records of the Royal Society of London, 56(1): 83–88, January 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/532123.

Murrell:2002:CEF

[1627] John N. Murrell and Nicole Grobert. The centenary of Einstein’s first scientific paper. Notes and Records of the Royal Society of London, 56(1): 89–94, January 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/532124.

Day:2002:MMP

[1628] P. Day. Molecular magnets: the prehistory. Notes and Records of the Royal Society of London, 56(1):95–103, January 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/532125.

Ogilvie:2002:BRP

[1629] Bridget Ogilvie. Book review: Pushing it! Frances Ashcroft, Life at the extremes: the science of survival. HarperCollins, London, 2000. Pp. 337, £17.99. ISBN 0-00-255946-3. Notes and Records of the Royal Society of London, 56(1):105–106, January 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532126.

Anonymous:2002:BR

[1630] Anonymous. Book reviews. Notes and Records of the Royal Society of London, 56(1):105–120, January 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Shea:2002:BRE

[1631] William Shea. Book review: Evolution not revolution: Rethinking the Scientific Revolution. Edited by Margaret J. Osler. Cambridge University Press, 2000. Pp. xii + 340, £48.86. ISBN 0-521-66790-9. Notes and Records of the Royal Society of London, 56(1):106–107, January 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/532127.

Rowlinson:2002:BRP

[1632] J. S. Rowlinson, F.R.Eng., F.R.S. Book review: A portrait gallery: Abraham Pais, The genius of science. A portrait gallery of twentieth-century physicists. Oxford University Press, 2000. Pp. 355. £26.50 (hardback). ISBN 0-19-850614-7. Notes and Records of the Royal Society of London, 56(1):107–108, January 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

532128.

Milanesi:2002:BRR

[1633] Marica Milanesi. Book review: This round world: Elly Dekker, Globes at Greenwich. Oxford University Press and the National Maritime Museum, 1999. Pp. xi + 592, £95.00 (hardback). ISBN 0-19-856559-3. Notes and Records of the Royal Society of London, 56(1):108–110, January 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/532129.

Liddle:2002:BRZ

[1634] Andrew Liddle. Book review: Zilch: John D. Barrow, The book of nothing. Jonathan Cape, London, 2000. Pp. 380, £16.99 (hardback). ISBN 0-224-05962-9. Notes and Records of the Royal Society of London, 56 (1):111, January 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/532130.

King-Hele:2002:BRP

[1635] Desmond King-Hele, F.R.S. Book review: Putting faces to the names: Adam Hart-Davis, Chain reactions: pioneers of British science and technology and the stories that link them. National Portrait Gallery, 2000. Pp. 192, £9.50 (paperback). ISBN 1-85514-291-0. Notes and Records of the Royal Society of London, 56(1):112–113, January 22, 2002. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/532131.

Brooke:2002:BRI

[1636] John Hedley Brooke. Book review: Isaac’s faith: Newton and religion: context, nature, and influence. Edited by James E. Force and Richard H. Popkin. Kluwer, Dordrecht, 1999. Pp. 325, £98.00, $168.00. ISBN 0-7923-5744-2. Notes and Records of the Royal Society of London, 56(1): 113–115, January 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/532132.

Fara:2002:BRS

[1637] Patricia Fara. Book review: Science and progress: Roy M. MacLeod, The ‘creed of science’ in Victorian England. Ashgate/Variorum, Aldershot, Burlington, USA, Singapore and Sydney, 2000. Pp. 346, £57.50 (hardback). ISBN 0-86078-669-2. Notes and Records of the Royal Society of London, 56(1):115–116, January 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532133.

Budd:2002:BRL

[1638] Graham E. Budd. Book review: In love with Olenus: R. A. Fortey, F.R.S., Trilobite! Eyewitness to evolution. HarperCollins, London, 2000. Pp. 269, £15.99 (hardback). ISBN 0-00-257012-2. Notes and Records of the Royal Society of London, 56(1):117–118, January 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/532134.

Fogg:2002:BRL

[1639] G. E. Fogg, F.R.S. Book review: Lead astray: Owen Beattie and John Geiger, Frozen in time: the fate of the Franklin expedition. Greystone Books, Vancouver, 1998. Pp. xi + 185, $19.95 Canada, $14.95 USA (paperback). ISBN 1-55054-616-3. Notes and Records of the Royal Society of London, 56(1):119–120, January 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532135.

May:2002:APL

[1640] R. M. May. Address of the President, Lord May of Oxford OM AC Kt FRS, given at the Anniversary Meeting on 30 November 2001. Notes and Records of the Royal Society of London, 56(1):121–129, January 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/532136.

Anonymous:2002:BMa

[1641] Anonymous. Back matter. Notes and Records of the Royal Society of London, 56(1):??, January 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

532137.

Anonymous:2002:FMa

[1642] Anonymous. Front matter. Notes and Records of the Royal Society of London, 56(1):??, January 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

532117.

Cook:2002:MB

[1643] Alan Cook. Max Born. Notes and Records of the Royal Society of London, 56(2):131–132, May 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

3557662;   http://www.jstor.org/stable/pdfplus/3557662.pdf.

Lewis:2002:PJW

[1644] R. Lewis. The publication of John Wilkins’s essay (1668): some contextual considerations. Notes and Records of the Royal Society of London, 56 (2):133–146, May 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/3557663.

daCosta:2002:CCR

[1645] P. Fontes da Costa. The culture of curiosity at The Royal Society in the first half of the eighteenth century. Notes and Records of the Royal Society of London, 56(2):147–166, May 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/3557664.

Ackroyd:2002:SGCa

[1646] J. A. D. Ackroyd. Sir George Cayley, the father of aeronautics. I. The invention of the aeroplane. Notes and Records of the Royal Society of London, 56(2):167–181, May 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

3557665.

Sharp:2002:SFT

[1647] R. Sharp. Some final thoughts of Sir John Herschel, F.R.S. Notes and Records of the Royal Society of London, 56(2):183–186, May 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/3557666.

Craik:2002:JIL

[1648] Alex D. D. Craik. James Ivory’s last papers on the “Figure of the Earth” (with biographical additions). Notes and Records of the Royal Society of London, 56(2):187–204, May 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/3557667.

Cohen:2002:RTM

[1649] A. Cohen. Roland Trimen and the Merope harem. Notes and Records of the Royal Society of London, 56(2):205–218, May 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/3557668.

Born:2002:WRF

[1650] G. V. R. Born. The wide-ranging family history of Max Born. Notes and Records of the Royal Society of London, 56(2):219–262, May 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/3557669;   http://
www.jstor.org/stable/pdfplus/3557669.pdf. See errata [1675].

Wallis:2002:BRK

[1651] Patrick Wallis. Book review: Knowledge not practice? Andrew Wear, Knowledge and practice in English medicine, 1550–1680. Cambridge University Press, 2000. Pp. viii + 496. £45.00, ISBN 0-521-55226-5 (hardback); £16.95, ISBN 0-521-55827-1 (paperback). Notes and Records of the Royal Society of London, 56(2):263–264, May 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/3557670.

Wallis:2002:BR

[1652] Patrick Wallis. Book reviews. Notes and Records of the Royal Society of London, 56(2):263–271, May 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Farr:2002:BRL

[1653] Patrick Farr. Book review: Letters from a scientific emperor: The letters of Sir Joseph Banks: a selection, 1768–1820. Edited by Neil Chambers. Imperial College Press, London, 2000. Pp. xliii + 420, £33.00 (hardback). ISBN 1-86094-204-0. Notes and Records of the Royal Society of London, 56(2):264–265, May 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

3557671.

Gooday:2002:BRV

[1654] Graeme Gooday. Book review: A Victorian scientist and engineer: Gillian Cookson and Colin Hempstead, Fleeming Jenkin and the birth of electrical engineering. Ashgate, Aldershot, 2000. Pp. x + 217, £45.00. ISBN 0-7546-0079-3. Notes and Records of the Royal Society of London, 56(2):265–266, May 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

3557672.

Fara:2002:BRD

[1655] Patricia Fara. Book review: Darwin’s origin of inspiration: Edna Healey, Emma Darwin: the inspirational wife of a genius. London, Headline, 2001. Pp. x + 372, £20.00 (hardback). ISBN 0-7472-7579-3. Notes and Records of the Royal Society of London, 56(2):267–268, May 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/3557673.

Bearn:2002:BRF

[1656] Alexander G. Bearn. Book review: Fading icon: Michael Bliss, William Osler: a life in medicine. Oxford University Press, 1999. Pp. 504, £31.25. ISBN 0-19-512346-8. Notes and Records of the Royal Society of London, 56(2):268–271, May 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

3557674.

Anonymous:2002:BMb

[1657] Anonymous. Back matter. Notes and Records of the Royal Society of London, 56(2):??, May 2002. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/3557675.

Anonymous:2002:FMb (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

3557661.

Cook:2002:RT

[1659] A. Cook. A Roman tercentenary. Notes and Records of the Royal Society of London, 56(3):273, September 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/3557733.

Glyn:2002:ILS

[1660] Lynn B. Glyn. Israel Lyons: a short but starry career. The life of an eighteenth-century Jewish botanist and astronomer. Notes and Records of the Royal Society of London, 56(3):275–305, September 22, 2002. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/3557734.

Home:2002:RSE

[1661] R. W. Home. The Royal Society and the Empire: the colonial and Commonwealth Fellowship. Part 1. 1731–1847. Notes and Records of the Royal Society of London, 56(3):307–332, September 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/3557735.

Ackroyd:2002:SGCb

[1662] J. A. D. Ackroyd. Sir George Cayley, the father of aeronautics. II. Cayley’s aeroplanes. Notes and Records of the Royal Society of London, 56 (3):333–348, September 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

3557736.

James:2002:EF

[1663] F. A. J. L. James. Editing Faraday. Notes and Records of the Royal Society of London, 56(3):349–352, September 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/3557737.

Wilkes:2002:CBH

[1664] Maurice V. Wilkes. Charles Babbage and his world. Notes and Records of the Royal Society of London, 56(3):353–365, September 22, 2002. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/3557738.

Schwarz:2002:FB

[1665] K. K. Schwarz. Faraday and Babbage. Notes and Records of the Royal Society of London, 56(3):367–381, September 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/3557739.

Corden:2002:RWS

[1666] J. Corden. Report. ‘web of science history’ project. Notes and Records of the Royal Society of London, 56(3):383–388, September 22, 2002. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/3557740.

Mandelbrote:2002:BRB

[1667] Scott Mandelbrote. Book review: From Bremen to Pall Mall: Marie Boas Hall, Henry Oldenburg. Shaping The Royal Society. Oxford University Press, 2002. Pp. xii + 369, £70.00, $100.00. ISBN 0-19-851053-5. Notes and Records of the Royal Society of London, 56(3):389–390, September 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/3557741.

Mandelbrote:2002:BR

[1668] Scott Mandelbrote. Book reviews. Notes and Records of the Royal Society of London, 56(3):389–401, September 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Shea:2002:BRS

[1669] William R. Shea. Book review: The shifting image of Newton: Patricia Fara, Newton: the making of genius. Macmillan, London, 2002. Pp. xvi + 347, £20.00. ISBN 0-333-90735-3. Notes and Records of the Royal Society of London, 56(3):391–392, September 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/3557742.

Sibbett:2002:BRP

[1670] Wilson Sibbett. Book review: Pioneering the femtoworld of molecules: Ahmed Zewail, Voyage through time. American University in Cairo Press, 2002. Pp. 288, $22.95 (hardback). ISBN 977-424-677-2. Notes and Records of the Royal Society of London, 56(3):392–395, September 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/3557743.

Fara:2002:BRB

[1671] Patricia Fara. Book review: Big bangs: G. I. Brown, Invisible rays: a history of radioactivity. Sutton Publishing, Stroud, 2002. Pp. viii + 248, £19.99 (hardback). ISBN 0-7509-2667-8. Notes and Records of the Royal Society of London, 56(3):395–396, September 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/3557744.

Nabarro:2002:BRS

[1672] Frank R. N. Nabarro, F.R.S. Book review: A scientist of the Commonwealth: Brian Austin, Schonland, scientist and soldier, Institute of Physics Publishing, 2001. Pp. xviii + 620, £49.00 (hardback). ISBN 0-7503-0501-0. Notes and Records of the Royal Society of London, 56 (3):396–398, September 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

3557745.

Edwards:2002:BRA

[1673] A. W. F. Edwards. Book review: From African exploration to the birth of eugenics: Nicholas Wright Gillham, A life of Sir Francis Galton: from African exploration to the birth of eugenics. Oxford University Press, New York, 2001. Pp. xii + 416. £22.50 (hardback). ISBN 0-19-514365-5. Notes and Records of the Royal Society of London, 56(3):399–400, September 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/3557746.

Watts:2002:BRE

[1674] Fraser Watts. Book review: Explaining morality: Robert Hinde, Why good is good: the sources of morality, Routledge, London, 2002. Pp. 241, £15.99 (paperback). ISBN 0-415-27753-1. Notes and Records of the Royal Society of London, 56(3):400–401, September 22, 2002. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/3557747.

Anonymous:2002:EWR

[1675] Anonymous. Errata: “The Wide-Ranging Family History of Max Born”. Notes and Records of the Royal Society of London, 56(3):403, September 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/3557748;   http:
//www.jstor.org/stable/pdfplus/3557748.pdf. See [1650].

Anonymous:2002:AIN

[1676] Anonymous. Author index to Notes and Records, volume 56, 2002. Notes and Records of the Royal Society of London, 56(3):405–406, September 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/3557731.

Anonymous:2002:BMc

[1677] Anonymous. Back matter. Notes and Records of the Royal Society of London, 56(3):??, September 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

3557749.

Anonymous:2002:FMc

[1678] Anonymous. Front matter. Notes and Records of the Royal Society of London, 56(3):??, September 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

3557732.

Cook:2003:OA

[1679] A. Cook. 1703 and other anniversaries. Notes and Records of the Royal Society of London, 57(1):1–2, January 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/3557677.

Thrower:2003:SPF

[1680] N. J. W. Thrower. Samuel Pepys FRS (1633–1703) and The Royal Society. Notes and Records of the Royal Society of London, 57(1):3–13, January 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/3557678.

Jones:2003:DUC

[1681] R. W. Jones. Dalton’s unfortunate choice. Notes and Records of the Royal Society of London, 57(1):15–33, January 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/3557679.

Rowlinson:2003:SEC

[1682] J. S. Rowlinson. Le Sage’s Essai de chymie méchanique. Notes and Records of the Royal Society of London, 57(1):35–45, January 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/3557680.

Home:2003:RSE

[1683] R. W. Home. The Royal Society and the Empire: The colonial and Commonwealth Fellowship. Part 2. After 1847. Notes and Records of the Royal Society of London, 57(1):47–84, January 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/3557681.

Greenwood:2003:ALR

[1684] N. N. Greenwood and J. A. Spink. An Antipodean laboratory of remarkable distinction. Notes and Records of the Royal Society of London, 57(1): 85–105, January 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/3557682.

Heap:2003:BRD

[1685] Sir Brian Heap, FRS. Book review: From the dust of the Earth? Henry Harris, Things come to life: spontaneous generation revisited. Oxford University Press, 2002. Pp. 178, £20.00 (hardback). ISBN 0-19-851538-3. Notes and Records of the Royal Society of London, 57(1):107–108, January 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/3557683.

Heap:2003:BR

[1686] B. Heap, FRS. Book reviews. Notes and Records of the Royal Society of London, 57(1):107–116, January 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Lipton:2003:BRS

[1687] Peter Lipton. Book review: The science of science: John Ziman, Real science: what it is, and what it means. Cambridge University Press, 2000. Pp. 412, £25.00 (hardback). ISBN 0-521-77229-X; £16.95 (paperback). ISBN 0-521-89310-0. Notes and Records of the Royal Society of London, 57(1):108–111, January 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

3557684.

Heilbron:2003:BRI

[1688] J. L. Heilbron. Book review: Information technology: Richard Yeo. Encyclopaedic visions. Scientific dictionaries and Enlightenment culture. Cambridge University Press, 2001. Pp. xxi + 336, £40.00 (hardback). ISBN 0-521-65191-3. Notes and Records of the Royal Society of London, 57(1):111–112, January 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

3557685.

Cook:2003:BRE

[1689] Sir Alan Cook, FRS. Book review: The end of the affair: Eric G. Forbes, Lesley Murdin and Frances Willmoth, The correspondence of John Flamsteed, the first Astronomer Royal, volume 3 (1703–1719). Institute of Physics Publishing, Bristol, 2002. Pp. lxvi + 1038, £199.00 (hardback). ISBN 0-7503-0763-3. Notes and Records of the Royal Society of London, 57(1):113–116, January 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

3557686.

May:2003:APL

[1690] R. M. May. Address of the President, Lord May of Oxford OM AC KT FRS, given at the Anniversary Meeting on 29 November 2002. Notes and Records of the Royal Society of London, 57(1):117–132, January 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/3557687.

Anonymous:2003:BMa

[1691] Anonymous. Back matter. Notes and Records of the Royal Society of London, 57(1):??, January 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

3557688.

Anonymous:2003:FMa

[1692] Anonymous. Front matter. Notes and Records of the Royal Society of London, 57(1):??, January 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

3557676.

Cook:2003:TA

[1693] A. Cook. Two anniversaries. Notes and Records of the Royal Society of London, 57(2):133, May 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

3557690.

Cartwright:2003:TTR

[1694] David E. Cartwright. The Tonkin tides revisited. Notes and Records of the Royal Society of London, 57(2):135–142, May 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/3557691.

Rowlinson:2003:WTA

[1695] J. S. Rowlinson. The work of Thomas Andrews and James Thomson on the liquefaction of gases. Notes and Records of the Royal Society of London, 57(2):143–159, May 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/3557692.

Coley:2003:GPM

[1696] N. G. Coley. George Pearson MD, FRS (1751–1828): ‘the greatest chemist in England’? Notes and Records of the Royal Society of London, 57(2):161–175, May 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

3557693.

Ingram:2003:AYC

[1697] N. R. Ingram. ‘All that you can’t leave behind’. Notes and Records of the Royal Society of London, 57(2):177–184, May 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/3557694.

VanderKloot:2003:ESA

[1698] W. Van der Kloot. Ernest Starling’s analysis of the energy balance of the German people during the blockade, 1914–19. Notes and Records of the Royal Society of London, 57(2):185–193, May 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/3557695.

Holliday:2003:EYM

[1699] R. Holliday. The early years of molecular biology: personal recollections. Notes and Records of the Royal Society of London, 57(2):195–208, May 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/3557696.

Moon:2003:RWF

[1700] F. C. Moon. Robert Willis and Franz Reuleaux: pioneers in the theory of machines. Notes and Records of the Royal Society of London, 57(2): 209–230, May 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/3557697.

See correction [1733].

Zinsser:2003:ERU

[1701] J. P. Zinsser. Essay review: The ultimate commentary: a consideration of I. Bernard Cohen’s Guide to Newton’s Principia. Notes and Records of the Royal Society of London, 57(2):231–238, May 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/3557698.

Cooke:2003:BRS

[1702] Anne Cooke. Book review: Shedding light on the enigmatic Rosalind Franklin: Brenda Maddox, Rosalind Franklin the dark lady of DNA. HarperCollins, London, 2002. Pp. 380, £20.00 (hardback). ISBN 0-00-257149-8. Notes and Records of the Royal Society of London, 57(2): 239–243, May 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/3557699.

Cooke:2003:BR

[1703] A. Cooke. Book reviews. Notes and Records of the Royal Society of London, 57(2):239–260, May 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Heyman:2003:BRB

[1704] Jacques Heyman, FREng. Book review: The birth of the modern architect: Lisa Jardine, On a grander scale: the outstanding career of Sir Christopher Wren. HarperCollins, London, 2002. Pp. xviii + 600, £25 (hardback). ISBN 0-00-710775-7. Notes and Records of the Royal Society of London, 57(2):243–244, May 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/3557700.

Hayes:2003:BRP

[1705] W. Hayes. Book review: Physics as well as poetry: Physicists of Ireland. Edited by Mark McCartney and Andrew Whitaker. Institute of Physics, Bristol and Philadelphia, 2003. Pp. 298, £40.00. ISBN 0-7503-0866-4. Notes and Records of the Royal Society of London, 57(2):245–246, May 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/3557701.

Henry:2003:BRN

[1706] John Henry. Book review: Newton for philosophers: The Cambridge Companion to Newton. Edited by I. Bernard Cohen and George E. Smith. Cambridge University Press, 2002. Pp. xiv + 500, £45.00 (hardback), ISBN 0-521-65177-8; £15.95 (paperback), ISBN 0-521-65696-6. Notes and Records of the Royal Society of London, 57(2):246–247, May 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/3557702.

Hesse:2003:BRE

[1707] Mary Hesse. Book review: An Eighteenth-Century polymath: David Boyd Haycock, William Stukeley: science, religion and archaeology in eighteenth-century England. Boydell Press, Woodbridge, 2002. Pp. xiii + 290, £40.00 (hardback). ISBN 0-85115-864-1. Notes and Records of the Royal Society of London, 57(2):248–249, May 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/3557703.

Savours:2003:BRN

[1708] Ann Savours. Book review: A new life of Sir John Franklin FRS: Martyn Beardsley. Deadly winter: the life of Sir John Franklin. Chatham Publishing, London, 2002. Pp. xiv + 272, £20.00 (hardback). ISBN 1-86176-187-2. Notes and Records of the Royal Society of London, 57(2): 249–250, May 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/3557704.

Morris:2003:BRW

[1709] Simon Conway Morris, FRS. Book review: When the penny dropped: J. Browne, Charles Darwin: the power of place. Jonathan Cape, London, 2002. Pp. 591, £25.00 (hardback). ISBN 0-224-04212-2. Notes and Records of the Royal Society of London, 57(2):250–251, May 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/3557705.

Pippard:2003:BRE

[1710] Brian Pippard, FRS. Book review: Euclid and other popular romances: M. Chisholm, Such silver currents: the story of William and Lucy Clifford 1845–1929. The Lutterworth Press, Cambridge, 2002, Pp. ix + 198, £17.50. ISBN 0-7188-3017-2. Notes and Records of the Royal Society of London, 57(2):252–253, May 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/3557706.

Warwick:2003:BRR

[1711] Andrew Warwick. Book review: Righting the history of electrodynamics: Olivier Darrigol, Electrodynamics from Ampère to Einstein. Oxford University Press, 2000. Pp. xi + 532, £79.50 (hardback). ISBN 0-19-850594-9. Notes and Records of the Royal Society of London, 57(2): 253–254, May 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/3557707.

Rowlinson:2003:BRG

[1712] J. S. Rowlinson, FRS, FREng. Book review: Gases and liquids: Johanna Levelt Sengers, How fluids unmix: discoveries by the school of Van der Waals and Kamerlingh Onnes. EDITA, Koninklijke Nederlandse Akademie van Wetenschappen, Amsterdam, 2002. Pp. xv + 302, EUR 45.00. ISBN 90-6984-357-9. Notes and Records of the Royal Society of London, 57(2):255–256, May 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/3557708.

Sibbett:2003:BRL

[1713] Wilson Sibbett, CBE, FRS. Book review: Lasers and masers: Charles H. Townes, How the laser happened: adventures of a scientist. Oxford University Press, 2002. Pp. 208, £9.99 (paperback). ISBN 0-19-515376-6. Notes and Records of the Royal Society of London, 57(2):257–258, May 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/3557709.

Hull:2003:BRC

[1714] Derek Hull, FRS, FREng. Book review: Changing times in a material world: A. Macfarlane and G. Martin, The glass bathyscaphe: how glass changed the world. Profile Books Ltd, London, 2002. Pp. 255, £15.00 (hardback). ISBN 1-86197-400-0. R. W. Cahn, The coming of materials science. Elsevier Science Ltd, Oxford, 2001. Pp. 568, £40.00 (hardback). ISBN 0-08-042679-4. Notes and Records of the Royal Society of London, 57(2):259–260, May 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

3557710.

Klug:2003:AGS

[1715] Aaron Klug, OM, FRS. Address given at the Service of Thanksgiving for the life of George Porter OM FRS. Notes and Records of the Royal Society of London, 57(2):261–264, May 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/3557711.

Lambert:2003:NRS

[1716] F. Lambert. News: The Royal Society Rosalind Franklin Award. Notes and Records of the Royal Society of London, 57(2):265–266, May 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/3557712.

Anonymous:2003:BMb

[1717] Anonymous. Back matter. Notes and Records of the Royal Society of London, 57(2):??, May 2003. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/3557713.

Anonymous:2003:FMb

[1718] Anonymous. Front matter. Notes and Records of the Royal Society of London, 57(2):??, May 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

3557689.

Cook:2003:CC

[1719] A. Cook. Continental connections. Notes and Records of the Royal Society of London, 57(3):267–268, September 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/3557716.

Breeze:2003:ARH

[1720] P. Breeze. The ancestry of Robert Hooke: John Hooke of Wroxhall. Notes and Records of the Royal Society of London, 57(3):269–271, September 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/3557717.

Appleby:2003:FSP

[1721] J. H. Appleby. The founding of St Petersburg in the context of The Royal Society’s relationship with Russia. Notes and Records of the Royal Society of London, 57(3):273–284, September 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/3557718.

Croarken:2003:ALM

[1722] M. Croarken. Astronomical labourers: Maskelyne’s assistants at the Royal Observatory, Greenwich, 1765–1811. Notes and Records of the Royal Society of London, 57(3):285–298, September 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/3557719.

Griffith:2003:CHF

[1723] W. P. Griffith and P. J. T. Morris. Charles Hatchett FRS (1765– 1847), chemist and discoverer of niobium. Notes and Records of the Royal Society of London, 57(3):299–316, September 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/3557720.

Bleaney:2003:NONa

[1724] B. Bleaney and O. V. Lounasmaa. Nuclear orientation and nuclear cooling experiments in Oxford and Helsinki. Part 1. Progress before 1940. Notes and Records of the Royal Society of London, 57(3):317– 322, September 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/3557721.

Bleaney:2003:NONb

[1725] B. Bleaney and O. V. Lounasmaa. Nuclear orientation and nuclear cooling experiments in Oxford and Helsinki. Part 2. Progress from 1945 to 1970. Notes and Records of the Royal Society of London, 57(3):323–330, September 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/3557722.

Bleaney:2003:NONc

[1726] B. Bleaney and O. V. Lounasmaa. Nuclear orientation and nuclear cooling experiments in Oxford and Helsinki. Part 3. Progress from 1975 to 2001. Notes and Records of the Royal Society of London, 57(3):331–344, September 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/3557723.

Anderson:2003:BRP

[1727] Robert Anderson. Book review: Puritan founding fathers? Charles Webster: The great instauration: science, medicine and reform 1626–1660. Peter Lang, Bern, 2002. Pp. xl + 630, £47.00. ISBN 3-906768-37-6. Notes and Records of the Royal Society of London, 57(3):345–346, September 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/3557724.

Anderson:2003:BR

[1728] Robert Anderson. Book reviews. Notes and Records of the Royal Society of London, 57(3):345–352, September 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Browne:2003:BRD

[1729] Janet Browne. Book review: Darwin on Darwin: Desmond King-Hele (editor), Charles Darwin’s The Life of Erasmus Darwin. Cambridge University Press, 2003. Pp 172, £17.50. ISBN 0-521-81526-6. Notes and Records of the Royal Society of London, 57(3):346–348, September 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/3557725.

King-Hele:2003:BRM

[1730] Desmond King-Hele, FRS. Book review: Makers of the future: Jenny Uglow, The Lunar men: the friends who made the future. Faber and Faber, London, 2002. Pp. xx + 588, £25 (hardback). ISBN 0-571-19647-0. Notes and Records of the Royal Society of London, 57(3):348–350, September 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/3557726.

Fara:2003:BRM

[1731] Patricia Fara. Book review: Measuring mistakes: K. Alder, The measure of all things: the seven-year odyssey that transformed the world. Little Brown, London, 2002. Pp. xii + 466, £15.99 (hardback). ISBN 0-316-85989-3. Notes and Records of the Royal Society of London, 57(3):350– 351, September 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/3557727.

Thrush:2003:BRM

[1732] Brian Thrush, FRS. Book review: MR spectroscopy: Boris Stoicheff, Gerhard Herzberg: an illustrious life in science. NRC Press, Ottawa, Canada, 2002. Pp. xiii + 468, £37.95 (hardback). ISBN 0-660-18757-4. Notes and Records of the Royal Society of London, 57(3):351–352, September 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/3557728.

Moon:2003:CFC

[1733] F. C. Moon. Correction for F. C. Moon, Robert Willis and Franz Reuleaux: pioneers in the theory of machines, Notes and Records 57(2) 209–230 (2003). Notes and Records of the Royal Society of London, 57(3): 353, September 22, 2003. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/3557729.

See [1700].

Anonymous:2003:AIN

[1734] Anonymous. Author index to Notes and Records, volume 57, 2003. Notes and Records of the Royal Society of London, 57(3):355–356, September 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/3557714.

Anonymous:2003:BMc

[1735] Anonymous. Back matter. Notes and Records of the Royal Society of London, 57(3):??, September 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

3557730.

Anonymous:2003:FMc

[1736] Anonymous. Front matter. Notes and Records of the Royal Society of London, 57(3):??, September 2003. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

3557715.

Quinn:2004:RNR

[1737] T. Quinn. The role of Notes and Records and some future anniversaries. Notes and Records of the Royal Society of London, 58(1):1–2, January 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/4142029.

Cook:2004:RRS

[1738] A. Cook. Rome and the Royal Society, 1660–1740. Notes and Records of the Royal Society of London, 58(1):3–19, January 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/4142030.

Agnew:2004:RFM

[1739] Duncan Carr Agnew. Robert Fitzroy and the myth of the “Marsden square”: transatlantic rivalries in early marine meteorology. Notes and Records of the Royal Society of London, 58(1):21–46, January 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/4142031.

Singh:2004:NLS

[1740] Rajinder Singh and Falk Riess. The Nobel laureate Sir Chandrasekhara Venkata Raman FRS and his contacts with the British scientific community in a social and political context. Notes and Records of the Royal Society of London, 58(1):47–64, January 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/4142032.

Crowther:2004:VDQ

[1741] R. A. Crowther. Viruses and the development of quantitative biological electron microscopy. Notes and Records of the Royal Society of London, 58(1):65–81, January 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

4142033.

Iliffe:2004:RNN

[1742] R. Iliffe and J. Young. REPORT. Newton on the net: first and prospective fruits of a Royal Society grant. Notes and Records of the Royal Society of London, 58(1):83–88, January 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/4142034.

Hunter:2004:RRH

[1743] M. Hunter. REPORT. Robert Hooke revivified. Notes and Records of the Royal Society of London, 58(1):89–91, January 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/4142035.

Cook:2004:RRH

[1744] A. Cook. REPORT. Robert Hooke at Christ Church, Oxford. Notes and Records of the Royal Society of London, 58(1):93–95, January 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/4142036.

McConnell:2004:ROW

[1745] A. McConnell. Report. OxDNB welcomes new old Fellows. Notes and Records of the Royal Society of London, 58(1):97–99, January 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/4142037.

Gray:2004:BRH

[1746] Jeremy Gray. Book review: Harriot restored: Jacqueline A. Stedall, The greate invention of algebra: Thomas Harriot’s treatise on equations. Oxford University Press, 2003. Pp. xi + 322, £65.00, ISBN 0-19-852602-4. Notes and Records of the Royal Society of London, 58(1):101–102, January 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/4142038.

Anonymous:2004:BRa

[1747] Anonymous. Book reviews. Notes and Records of the Royal Society of London, 58(1):101–106, January 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

McConnell:2004:BRA

[1748] Anita McConnell. Book review: All guns blazing: W. Johnson, Collected works on Benjamin Robins and Charles Hutton. Phoenix Press, New Delhi, 2001. Pp. x + 540 + errata. ISBN 81-7484-054-0. Notes and Records of the Royal Society of London, 58(1):102–103, January 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/4142039.

Pippard:2004:BRP

[1749] Brian Pippard, FRS. Book review: A paragon of physics: Basil Mahon, The man who changed everything — the life of James Clerk Maxwell. Wiley, Chichester, 2003. Pp. xx + 226, £18.99 (hardback). ISBN 0-470-86088-X. Notes and Records of the Royal Society of London, 58(1):103– 104, January 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/4142040.

Rowlinson:2004:BRN

[1750] J. S. Rowlinson, FRS, FREng. Book review: ‘a national treasure house’: ‘The common purposes of life’: science and society at the Royal Institution of Great Britain. Edited by Frank A. J. L. James. Ashgate, Aldershot, 2002. Pp. xvi + 456, £55.00 (hardback). ISBN 0-7546-0960-X. Notes and Records of the Royal Society of London, 58(1):105–106, January 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/4142041.

Jardine:2004:WLD

[1751] L. Jardine. The 2003 Wilkins Lecture: Dr Wilkins’s boy wonders. Notes and Records of the Royal Society of London, 58(1):107–129, January 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/4142042.

May:2004:APL

[1752] R. M. May. Address of the President, Lord May of Oxford OM AC KT FRS, given at the Anniversary Meeting on 1 December 2003. Notes and Records of the Royal Society of London, 58(1):131–145, January 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/4142043.

Anonymous:2004:BMa

[1753] Anonymous. Back matter. Notes and Records of the Royal Society of London, 58(1):??, January 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

4142044.

Anonymous:2004:FMa

[1754] Anonymous. Front matter. Notes and Records of the Royal Society of London, 58(1):??, January 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

4142028.

Quinn:2004:MOQ

[1755] T. Quinn. Microscopes and the observation of quantum dots. Notes and Records of the Royal Society of London, 58(2):147, May 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/4142046.

Anonymous:2004:FMb

[1756] Anonymous. Front matter. Notes and Records of the Royal Society of London, 58(2):148, May 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

4142045.

VanderKloot:2004:AFF

[1757] W. Van der Kloot. April 1915: Five future Nobel Prize-winners inaugurate weapons of mass destruction and the academic–industrial–military complex. Notes and Records of the Royal Society of London, 58(2):149– 160, May 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/4142047.

Rowlinson:2004:WWH

[1758] J. S. Rowlinson. The wartime work of Hinshelwood and his colleagues. Notes and Records of the Royal Society of London, 58(2):161–175, May 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/4142048.

Blow:2004:PBC

[1759] D. Blow and S. Wallwork. Prehistory of the British crystallographic association. Notes and Records of the Royal Society of London, 58(2): 177–186, May 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/4142049.

Gest:2004:DMR

[1760] H. Gest. The discovery of microorganisms by Robert Hooke and Antoni van Leeuwenhoek, Fellows of The Royal Society. Notes and Records of the Royal Society of London, 58(2):187–201, May 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/4142050.

Kurzer:2004:LWC

[1761] F. Kurzer. The life and work of Charles Tomlinson FRS: a career in Victorian science and technology. Notes and Records of the Royal Society of London, 58(2):203–226, May 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/4142051.

Harper:2004:RPS

[1762] P. Harper. Report. preserving scientific archives: the work of the national cataloguing unit for the archives of contemporary scientists (NCUACS). Notes and Records of the Royal Society of London, 58(2):227–232, May 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/4142052.

Coley:2004:BRM

[1763] Noel Coley. Book review: The making of a hero of science: Marco Beretta, Imaging a career in science: the iconography of Antoine Laurent Lavoisier. Science History Publications, Canton, MA, USA, 2001. Pp. xvii + 126, $29.95. ISBN 0-88135-294-2. Notes and Records of the Royal Society of London, 58(2):233–234, May 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/4142053.

Anonymous:2004:BRb

[1764] Anonymous. Book reviews. Notes and Records of the Royal Society of London, 58(2):233–243, May 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Cooke:2004:BRM

[1765] Anne Cooke. Book review: From microbes to man: Ulf Lagerkvist, Pioneers of microbiology and the Nobel Prize. World Scientific, 2003. Pp. 192, £33.00 (hardback). ISBN 981-238-233-X. £16.00 (paperback). ISBN 981-238-234-8. Notes and Records of the Royal Society of London, 58(2): 234–236, May 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/4142054.

Fara:2004:BRT

[1766] Patricia Fara. Book review: Theatrical science. C. Djerassi and D. Pinner, Newton’s darkness: two dramatic views. Imperial College Press, London, 2003. Pp. 184, $24 (hardback), US$15 (paperback). ISBN 1-86094-390-X. Notes and Records of the Royal Society of London, 58(2):237–238, May 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/4142055.

Hunter:2004:BRF

[1767] Michael Hunter. Book review: The first ethologist? Ted Dadswell, The Selborne pioneer. Gilbert White as naturalist and scientist: a reexamination. Ashgate, Aldershot, 2003. Pp. xx + 238, £45.00. ISBN 0-7546-0749-6. Notes and Records of the Royal Society of London, 58(2): 238–239, May 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/4142056.

Quinn:2004:BRF

[1768] Terry Quinn, FRS. Book review: Foucault, a man of many parts: William Tobin, The life and science of Léon Foucault: the man who proved the Earth rotates. Cambridge University Press, 2003. Pp. 352, £40.00 (hardback). ISBN 0-521-80855-3. Notes and Records of the Royal Society of London, 58(2):239–242, May 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/4142057.

Rowlinson:2004:BRD

[1769] J. S. Rowlinson, FRS, FREn. Book review: Displacing phlogiston: Robert Siegfried, From elements to atoms: a history of chemical composition. American Philosophical Society, Philadelphia, PA, 2002 (from Transactions of the American Philosophical Society, vol. 92, pt. 4). Pp. x + 278, $24.00 (paperback). ISBN 0-87169-924-9. Notes and Records of the Royal Society of London, 58(2):242–243, May 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/4142058.

Anonymous:2004:BMb

[1770] Anonymous. Back matter. Notes and Records of the Royal Society of London, 58(2):??, May 2004. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/4142059.

Quinn:2004:I

[1771] T. Quinn. In this issue. Notes and Records of the Royal Society of London, 58(3):245–247, September 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

4142062.

Quinn:2004:SAC

[1772] Terry Quinn. Sir Alan Cook FRS 1922–2004. Notes and Records of the Royal Society of London, 58(3):247, September 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/4142063.

Malcolm:2004:ULH

[1773] Noel Malcolm. An unpublished letter from Henry Oldenburg to Johann Heinrich Rahn. Notes and Records of the Royal Society of London, 58 (3):249–266, September 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

4142064.

Murphy:2004:MHT

[1774] E. Murphy. A mad house transformed: the lives and work of Charles James Beverly FRS (1788–1868) and John Warburton MD FRS (1795– 1847). Notes and Records of the Royal Society of London, 58(3):267–281, September 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/4142065.

Lovell:2004:CMW

[1775] B. Lovell. The cavity magnetron in World War II: was the secrecy justified? Notes and Records of the Royal Society of London, 58(3):283–294, September 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/4142066.

Cook:2004:SRW

[1776] A. Cook. Shipborne radar in World War II: some recollections. Notes and Records of the Royal Society of London, 58(3):295–298, September 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/4142067.

Friesen:2004:ERR

[1777] J. Friesen. Essay review. Robert Hooke under the microscope: London’s Leonardo. The Life and Work of Robert Hooke by Jim Bennett, Michael Cooper, Michael Hunter, Lisa Jardine; ‘A More Beautiful City.’ Robert Hooke and the Rebuilding of London after the Great Fire by Michael Cooper; The Man Who Knew Too Much. The Strange and Inventive Life of Robert Hooke (1635–1703) by Stephen Inwood; The Curious Life of Robert Hooke. The Man Who Measured London by Lisa Jardine. Notes and Records of the Royal Society of London, 58(3):299–302, September 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/4142068.

Cook:2004:RMR

[1778] Alan Cook. Report. migrants to the Royal Society, 1930–1940. Notes and Records of the Royal Society of London, 58(3):303–304, September 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/4142069.

Cordon:2004:RAW

[1779] Joanna Cordon. Report. The Andrew W. Mellon Cataloguing Project. Notes and Records of the Royal Society of London, 58(3):305–310, September 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/4142070.

Gray:2004:BRT

[1780] Jeremy Gray. Book review: Table Mountain: M. Campbell-Kelly, M. Croarken, R. Flood and E. Robson (eds), The history of mathematical tables: from Sumer to Spreadsheets. Oxford University Press, 2003. Pp. vii + 361, £45.00. ISBN 0-19-850841-7. Notes and Records of the Royal Society of London, 58(3):311–312, September 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/4142071.

Anonymous:2004:BRc

[1781] Anonymous. Book reviews. Notes and Records of the Royal Society of London, 58(3):311–320, September 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Fara:2004:BRL

[1782] Patricia Fara. Book review: Lucasian Professors: Kevin C. Knox and Richard Noakes (eds), From Newton to Hawking: a history of Cambridge University’s Lucasian professors of mathematics. Cambridge University Press, 2003. Pp. xxiv + 486, £27.50 (hardback). ISBN 0-521-66310-5. Notes and Records of the Royal Society of London, 58(3):312–313, September 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/4142072.

Ogilvie:2004:BRH

[1783] Marilyn Bailey Ogilvie. Book review: In her place: Patricia Fara, Pandora’s breeches: women, science and power in the Enlightenment. Pimlico, London, 2004. Pp. 274, £12.50 (paperback). ISBN 1-84413-082-7. Notes and Records of the Royal Society of London, 58(3):314–315, September 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/4142073.

Knight:2004:BRT

[1784] David Knight. Book review: Test of time: Maurice Crosland, Gay Lussac: scientist and bourgeois. Cambridge University Press, 2004. Pp. xvi + 333, £33.00 (paperback). ISBN 0-521-52483-0. Notes and Records of the Royal Society of London, 58(3):315–316, September 22, 2004. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/4142074.

Cooter:2004:BRS

[1785] Roger Cooter. Book review: Scratching heads: John van Wyhe, Phrenology and the origins of Victorian scientific naturalism. Ashgate Publishing Ltd, Aldershot, 2004. Pp. 282, £49.50. ISBN 0-7546-3408-6. Notes and Records of the Royal Society of London, 58(3):316–317, September 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/4142075.

Marsden:2004:BRW

[1786] Ben Marsden. Book review: From Watt to Cardwell: Donald Cardwell, The development of science and technology in nineteenth-century Britain: the importance of Manchester. Edited by Richard L. Hills. Variorum Collected Studies Series, CS765. Ashgate, Aldershot, 2003. Pp. xx + 320. £29.50 (hardback). ISBN 0-86078-908-X. Notes and Records of the Royal Society of London, 58(3):318–320, September 22, 2004. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/4142076.

Lawton:2004:JPC

[1787] J. H. Lawton. Japan Prize commemorative lecture: biodiversity, conservation and sustainability. Notes and Records of the Royal Society of London, 58(3):321–333, September 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/4142077.

Anonymous:2004:AIN

[1788] Anonymous. Author index to Notes and Records, volume 58, 2004. Notes and Records of the Royal Society of London, 58(3):335–336, September 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/4142060.

Anonymous:2004:BMc

[1789] Anonymous. Back matter. Notes and Records of the Royal Society of London, 58(3):??, September 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

4142078.

Anonymous:2004:FMc (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

4142061.

Quinn:2005:EIC

[1791] Terry Quinn. Editorial. An Italian correspondence, an Italian earthquake and the homes of The Royal Society. Notes and Records of the Royal Society of London, 59(1):1–4, January 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/30041467.

Cook:2005:RCG

[1792] Alan Cook. A Roman correspondence: George Ent and Cassiano dal Pozzo, 1637–55. Notes and Records of the Royal Society of London, 59(1):5–23, January 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

30041468.

Crosland:2005:RBR

[1793] Maurice Crosland. Relationships between the Royal Society and the Académie des Sciences in the late eighteenth century. Notes and Records of the Royal Society of London, 59(1):25–34, January 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/30041469.

Cook:2005:RDK

[1794] G. C. Cook. Richard Dobson Kt MRCS FRS (1773–1847) and the inferior status of naval medicine in the early nineteenth century: end of the fleet physicians. Notes and Records of the Royal Society of London, 59(1):35– 43, January 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/30041470.

Ferrari:2005:RMG

[1795] Graziano Ferrari and Anita McConnell. Robert Mallet and the ‘Great Neapolitan earthquake’ of 1857. Notes and Records of the Royal Society of London, 59(1):45–64, January 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/30041471.

Fischer:2005:RRS

[1796] Stefanie Fischer. Report. The Royal Society redevelopment. Notes and Records of the Royal Society of London, 59(1):65–85, January 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/30041472.

Hunter:2005:RRB

[1797] Michael Hunter. Report. Robert Boyle for the twenty-first century. Notes and Records of the Royal Society of London, 59(1):87–90, January 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/30041473.

Chapman:2005:BRS

[1798] Allan Chapman. Book review: Sir William and Miss Caroline Herschel: Michael Hoskin, The Herschel partnership, as viewed by Caroline. Science History Publications Ltd, Cambridge, 2003. Pp. 182. £25.00. ISBN 0-905193-05-9. Michael Hoskin, Caroline Herschel’s autobiographies. Science History Publications Ltd, Cambridge, 2003. Pp. 146. £25.00. ISBN 0-905193-06-7. Notes and Records of the Royal Society of London, 59 (1):91–92, January 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

30041474.

Anonymous:2005:BR

[1799] Anonymous. Book reviews. Notes and Records of the Royal Society of London, 59(1):91–97, January 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Lee:2005:BRS

[1800] Peter M. Lee. Book review: Saxony, sexuality, science, socialism and statistics: Theodore M. Porter, Karl Pearson: the scientific life in a statistical age. Princeton University Press, 2004. Pp.viii + 342, £22.95 (hardback). ISBN 0-691-11445-5. Notes and Records of the Royal Society of London, 59(1):92–93, January 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/30041475.

Rowlinson:2005:BRM

[1801] J. S. Rowlinson, FRS, FREng. Book review: Mixed mathematics: Andrew Warwick, Masters of theory: Cambridge and the rise of mathematical physics. University of Chicago Press, 2003. Pp. xiv + 572, $29.00 (paperback). ISBN 0-226-87375-7. Notes and Records of the Royal Society of London, 59(1):94–96, January 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/30041476.

Sidebottom:2005:BRP

[1802] Eric Sidebottom. Book review: Penicillin revisited: From mould to miracle: Eric Lax, The mould in Dr Florey’s coat: the remarkable true story of the penicillin miracle. Little, Brown & Co., 2004. Pp. 389, £16.99. ISBN 0-316-85925-7. Notes and Records of the Royal Society of London, 59(1):96–97, January 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

30041477.

May:2005:APL

[1803] R. M. May. Address of the President Lord May of Oxford OM AC Kt FRS given at the anniversary meeting on 30 November 2004: Global problems and global science. Notes and Records of the Royal Society of London, 59(1):99–119, January 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

30041478.

Anonymous:2005:BMa

[1804] Anonymous. Back matter. Notes and Records of the Royal Society of London, 59(1):??, January 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

30041479.

Anonymous:2005:FMa

[1805] Anonymous. Front matter. Notes and Records of the Royal Society of London, 59(1):??, January 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

30041466.

Quinn:2005:DWD

[1806] Terry Quinn. From Darwin and Wallace to the discovery of Invar. Notes and Records of the Royal Society of London, 59(2):121–123, May 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/30041482.

Bulmer:2005:TNS

[1807] Michael Bulmer. The theory of natural selection of Alfred Russel Wallace FRS. Notes and Records of the Royal Society of London, 59(2):125–136, May 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/30041483.

Ord:2005:OBD

[1808] Margery G. Ord and Lloyd A. Stocken. The Oxford biochemistry department in wartime, 1939–45. Notes and Records of the Royal Society of London, 59(2):137–143, May 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/30041484.

Cahn:2005:UNP

[1809] Robert W. Cahn. An unusual Nobel prize. Notes and Records of the Royal Society of London, 59(2):145–153, May 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/30041485.

Higgitt:2005:PPF

[1810] Rebekah Higgitt. President, patron, friend and lover: Charles Montagu’s significance to the history of science. Notes and Records of the Royal Society of London, 59(2):155–170, May 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/30041486.

Ke:2005:CPC

[1811] Jie Ke and Martyn Poliakoff. The critical point of CO_{2} + N_{2}: an experiment inspired by Notes and Records. Notes and Records of the Royal Society of London, 59(2):171–174, May 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/30041487.

Seitz:2005:HCC

[1812] Frederick Seitz. Henry Cavendish: the catalyst for the chemical revolution. Notes and Records of the Royal Society of London, 59(2):175–199, May 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/30041488.

Broad:2005:LLP

[1813] Cathy Broad. The Linnaeus link project. Notes and Records of the Royal Society of London, 59(2):201–203, May 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/30041489.

Friesen:2005:HSC

[1814] John Friesen. Hooke and seventeenth century science. Notes and Records of the Royal Society of London, 59(2):205–206, May 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/30041490.

Thomas:2005:GLT

[1815] John Meurig Thomas. A good lecture is a tour de force …. Notes and Records of the Royal Society of London, 59(2):207–208, May 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/30041491.

Russell:2005:SSG

[1816] Colin A. Russell. Science for sixpence (in green covers). Notes and Records of the Royal Society of London, 59(2):209–210, May 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/30041492.

Williams:2005:TAP

[1817] Glyn Williams. Travels of an artist: from Polynesia to Bengal. Notes and Records of the Royal Society of London, 59(2):211–212, May 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/30041493.

Wykes:2005:DI

[1818] David L. Wykes. Dissenting influences. Notes and Records of the Royal Society of London, 59(2):213–214, May 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/30041494.

Quinn:2005:PRS

[1819] Terry Quinn. Plates from Royal Society publications: Experiments of Henry Cavendish. Notes and Records of the Royal Society of London, 59(2):215–218, May 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

30041495.

Anonymous:2005:BMb

[1820] Anonymous. Back matter. Notes and Records of the Royal Society of London, 59(2):??, May 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

30041496.

Anonymous:2005:FMb

[1821] Anonymous. Front matter. Notes and Records of the Royal Society of London, 59(2):??, May 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

30041481.

Quinn:2005:NCE

[1822] Terry Quinn, FRS. Newton’s calendar, Einstein and 340 years of Philosophical Transactions. Notes and Records of the Royal Society of London, 59(3):219–222, September 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

30041500.

Belenkiy:2005:HOD

[1823] Ari Belenkiy and Eduardo Vila Echagüe. History of one defeat: reform of the Julian calendar as envisaged by Isaac Newton. Notes and Records of the Royal Society of London, 59(3):223–254, September 22, 2005. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/30041501.

Rowlinson:2005:ECP

[1824] J. S. Rowlinson. Einstein: The classical physicist. Notes and Records of the Royal Society of London, 59(3):255–271, September 22, 2005. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/30041502.

VanderKloot:2005:LBR

[1825] William Van der Kloot. Lawrence Bragg’s role in the development of sound-ranging in World War I. Notes and Records of the Royal Society of London, 59(3):273–284, September 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/30041503.

Crilly:2005:ACF

[1826] Tony Crilly. Arthur Cayley, FRS and the four-colour map problem. Notes and Records of the Royal Society of London, 59(3):285–304, September 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/30041504.

Sen:2005:WWR

[1827] Siddhartha Sen. Why William Rowan Hamilton was not an FRS. Notes and Records of the Royal Society of London, 59(3):305–308, September 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/30041505.

MacFarlane:2005:RSM

[1828] Ross MacFarlane. Royal Society miscellaneous manuscripts. Notes and Records of the Royal Society of London, 59(3):309–312, September 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/30041506.

Young:2005:NNU

[1829] John T. Young. Newton on the net: an update. Notes and Records of the Royal Society of London, 59(3):313–316, September 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/30041507.

Noakes:2005:SNC

[1830] Richard Noakes. Science in the nineteenth-century periodical: an electronic index. Notes and Records of the Royal Society of London, 59 (3):317–318, September 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

30041508.

Fox:2005:ESH

[1831] Robert Fox. The European Society for the History of Science. Notes and Records of the Royal Society of London, 59(3):319–320, September 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/30041509.

May:2005:ALM

[1832] Robert M. May. Address by Lord May PRS, given at the dedication of the memorial to Dr Robert Hooke at Westminster Abbey on 3 March 2005. Notes and Records of the Royal Society of London, 59(3):321–323, September 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/30041510.

Doherty:2005:UTL

[1833] Peter C. Doherty. U Thant Lecture: science, society and the challenge of the future. Notes and Records of the Royal Society of London, 59 (3):325–331, September 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

30041511.

Bolt:2005:CH

[1834] Marvin Bolt. Colonizing the heavens. Notes and Records of the Royal Society of London, 59(3):333–334, September 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/30041512.

Fara:2005:PR

[1835] Patricia Fara. Physics rules. Notes and Records of the Royal Society of London, 59(3):335–336, September 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/30041513.

James:2005:AK

[1836] Frank A. J. L. James. Anthologizing kinetics. Notes and Records of the Royal Society of London, 59(3):337–338, September 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/30041514.

Rowlinson:2005:WK

[1837] J. S. Rowlinson. Which Kelvin? Notes and Records of the Royal Society of London, 59(3):339–341, September 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/30041515.

Quinn:2005:E

[1838] Terry Quinn. Editorial. Notes and Records of the Royal Society of London, 59(3):343–347, September 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/30041516.

Anonymous:2005:BMc

[1839] Anonymous. Back matter. Notes and Records of the Royal Society of London, 59(3):??, September 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

30041517.

Anonymous:2005:FMc

[1840] Anonymous. Front matter. Notes and Records of the Royal Society of London, 59(3):??, September 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

30041499.

Quinn:2006:LPP

[1841] Terry Quinn, FRS. Libraries: past, present and future. Notes and Records of the Royal Society of London, 60(1):1–3, January 22, 2006. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/20462546.

Quarrie:2006:SLE

[1842] Paul Quarrie. The scientific library of the Earls of Macclesfield. Notes and Records of the Royal Society of London, 60(1):5–24, January 22, 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/20462547.

Young:2006:INA

[1843] John T. Young. Isaac Newton’s alchemical notes in the Royal Society. Notes and Records of the Royal Society of London, 60(1):25–34, January 22, 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/20462548.

Jenkinson:2006:NVC

[1844] Matt Jenkinson. Nathanael Vincent and Confucius’s ‘great learning’ in restoration England. Notes and Records of the Royal Society of London, 60(1):35–47, January 22, 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20462549.

Ambrose:2006:HKL

[1845] Douglas Ambrose. A history of Kaye and Laby. Notes and Records of the Royal Society of London, 60(1):49–57, January 22, 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/20462550.

Thomas:2006:MPC

[1846] John Meurig Thomas. Max Perutz: chemist, molecular biologist, human rights activist. Notes and Records of the Royal Society of London, 60(1):59–65, January 22, 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20462551.

Rowlinson:2006:RSP

[1847] J. S. Rowlinson. Royal Society purchases from the Macclesfield Library. Notes and Records of the Royal Society of London, 60(1):67–68, January 22, 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/20462552.

Fox:2006:HSM

[1848] Robert Fox. The history of science, medicine and technology at Oxford. Notes and Records of the Royal Society of London, 60(1):69–83, January 22, 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/20462553.

Nutton:2006:JLB

[1849] Vivian Nutton. Johann Laurentius Bausch zum 400. Geburtstag: Die Gründung der Leopoldina (Academia Naturae Curiosorum) im historischen Kontext. Schweinfurt, 30 September–1 October 2005. Notes and Records of the Royal Society of London, 60(1):85, January 22, 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/20462554.

vanWyhe:2006:CWC

[1850] John van Wyhe. The complete work of Charles Darwin online. Notes and Records of the Royal Society of London, 60(1):87–89, January 22, 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/20462555.

Fellgett:2006:NOM

[1851] P. B. Fellgett. The nature and origin of multiplex Fourier spectrometry. Notes and Records of the Royal Society of London, 60(1):91–93, January 22, 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/20462556.

Chance:2006:CMA

[1852] Britton Chance. The 3cm cavity magnetron arrives in Germany. Notes and Records of the Royal Society of London, 60(1):95–96, January 22, 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/20462557.

Fara:2006:PN

[1853] Patricia Fara. NOT PLATO TO NATO. Notes and Records of the Royal Society of London, 60(1):97–98, January 22, 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/20462558.

Hughes:2006:LSL

[1854] Jeff Hughes. A life on the scientific left. Notes and Records of the Royal Society of London, 60(1):99–100, January 22, 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/20462559.

Leaney:2006:LSS

[1855] Enda Leaney. Into the light: Science & society in Ireland. Notes and Records of the Royal Society of London, 60(1):101–102, January 22, 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/20462560.

Rowlinson:2006:BCC

[1856] J. S. Rowlinson. Bringing chemistry to Cambridge. Notes and Records of the Royal Society of London, 60(1):103–104, January 22, 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/20462561.

Marsden:2006:DD

[1857] Ben Marsden. Deconstructing discovery. Notes and Records of the Royal Society of London, 60(1):105–106, January 22, 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/20462562.

Nutton:2006:HRT

[1858] Vivian Nutton. Heralds, roots and triumphs. Notes and Records of the Royal Society of London, 60(1):107–108, January 22, 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/20462563.

May:2006:APL

[1859] Robert M. May. Address of the President, Lord May of Oxford OM AC FRS, given at the Anniversary Meeting on 30 November 2005: Threats to tomorrow’s world. Notes and Records of the Royal Society of London, 60(1):109–130, January 22, 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20462564.

Quinn:2006:PSC

[1860] Terry Quinn. Plates from seventeenth-century medical and alchemical texts in the Royal Society library. Notes and Records of the Royal Society of London, 60(1):131–134, January 22, 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/20462565.

Anonymous:2006:BMa

[1861] Anonymous. Back matter. Notes and Records of the Royal Society of London, 60(1):??, January 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20462566.

Anonymous:2006:FMa

[1862] Anonymous. Front matter. Notes and Records of the Royal Society of London, 60(1):??, January 22, 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20462545.

Quinn:2006:PWF

[1863] Terry Quinn. The Palace of Westminster frescoes, Kelvin’s secretary and the Calcutta School of Tropical Medicine. Notes and Records of the Royal Society of London, 60(2):135–137, May 22, 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/20462570.

Kurzer:2006:AHC

[1864] Frederick Kurzer. Arthur Herbert Church FRS and the Palace of Westminster frescoes. Notes and Records of the Royal Society of London, 60(2):139–159, May 22, 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20462571.

Rowlinson:2006:DTC

[1865] J. S. Rowlinson. Dr Thomas Carver and Lord Kelvin. Notes and Records of the Royal Society of London, 60(2):161–170, May 22, 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/20462572.

Cook:2006:LRK

[1866] G. C. Cook. Leonard Rogers KCSI FRCP FRS (1868–1962) and the founding of the Calcutta school of tropical medicine. Notes and Records of the Royal Society of London, 60(2):171–181, May 22, 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/20462573.

Gooday:2006:HPS

[1867] Graeme Gooday. History and philosophy of science at Leeds. Notes and Records of the Royal Society of London, 60(2):183–192, May 22, 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/20462574.

Moore:2006:ULC

[1868] Keith Moore. An unpublished letter by Charles Darwin in the Royal Society’s archives. Notes and Records of the Royal Society of London, 60(2):193–197, May 22, 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20462575.

Fara:2006:EHL

[1869] Patricia Fara. Edmond Halley’s last portrait. Notes and Records of the Royal Society of London, 60(2):199–201, May 22, 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/20462576.

Quinn:2006:TSC

[1870] Terry Quinn. Temperature: A showcase for complementary science? Notes and Records of the Royal Society of London, 60(2):203–214, May 22, 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/20462577.

King-Hele:2006:MII

[1871] Desmond King-Hele. The many images of Isaac Newton. Notes and Records of the Royal Society of London, 60(2):215–216, May 22, 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/20462578.

Home:2006:BS

[1872] R. W. Home. A beguiling story. Notes and Records of the Royal Society of London, 60(2):217–218, May 22, 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/20462579.

VanDamme:2006:DGG

[1873] Stephane Van Damme. The Dutch Galileo? The greatness of Huygens’ science. Notes and Records of the Royal Society of London, 60(2):219– 220, May 22, 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/20462580.

Rowlinson:2006:LM

[1874] J. S. Rowlinson. Laplace: the man. Notes and Records of the Royal Society of London, 60(2):221–223, May 22, 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/20462581.

Bennett:2006:PRS

[1875] Jim Bennett. Plates from Royal Society publications: Illustrating William Roy’s baseline on Hounslow Heath. Notes and Records of the Royal Society of London, 60(2):225–230, May 22, 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/20462582.

Anonymous:2006:BMb

[1876] Anonymous. Back matter. Notes and Records of the Royal Society of London, 60(2):??, May 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20462583.

Anonymous:2006:FMb

[1877] Anonymous. Front matter. Notes and Records of the Royal Society of London, 60(2):??, May 22, 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20462568.

Quinn:2006:HF

[1878] Terry Quinn, FRS. The Hooke folio. Notes and Records of the Royal Society of London, 60(3):231–233, September 22, 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/20462587.

Adams:2006:RHF

[1879] Robyn Adams and Lisa Jardine. The return of the Hooke folio. Notes and Records of the Royal Society of London, 60(3):235–239, September 22, 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/20462588.

Ibanez:2006:PME

[1880] Neus Ibáñez, Josep M. Montserrat, Ignasi Soriano, and Josep M. Camarasa. Plant material exchanged between James Petiver (ca. 1663–1718) and Joan Salvador I Riera (1683–1725). I. The Balearic plants conserved in the Bc-Salvador and Bm-Sloane herbaria. Notes and Records of the Royal Society of London, 60(3):241–248, September 22, 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/20462589.

Keynes:2006:PDW

[1881] Milo Keynes. The portrait of Dr William Harvey in the Royal Society since 1683. Notes and Records of the Royal Society of London, 60 (3):249–252, September 22, 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20462590.

Higgitt:2006:WDF

[1882] Rebekah Higgitt. Why i don’t FRS my tail: Augustus DE Morgan and the Royal Society. Notes and Records of the Royal Society of London, 60 (3):253–259, September 22, 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20462591.

Durham:2006:RHS

[1883] Ian T. Durham. Rethinking the history of solar wind studies: Eddington’s analysis of Comet Morehouse. Notes and Records of the Royal Society of London, 60(3):261–270, September 22, 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/20462592.

Bristow:2006:SBM

[1884] Adrian F. Bristow, Trevor Barrowcliffe, and Derek R. Bangham. Standardization of biological medicines: the first hundred years, 1900– 2000. Notes and Records of the Royal Society of London, 60(3):271–289, September 22, 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/20462593.

Osler:2006:HTT

[1885] Margaret J. Osler. A hero for their times: early biographies of Newton. Notes and Records of the Royal Society of London, 60(3):291–305, September 22, 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/20462594.

Smallwood:2006:CC

[1886] John Smallwood. Circumscribing the circle? Notes and Records of the Royal Society of London, 60(3):307–309, September 22, 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/20462595.

Moore:2006:PRS

[1887] Keith Moore. Plates from Royal Society publications: Fin-de-siÈcle newspaper sketches. Notes and Records of the Royal Society of London, 60(3):311–320, September 22, 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20462596.

Anonymous:2006:BMc

[1888] Anonymous. Back matter. Notes and Records of the Royal Society of London, 60(3):??, September 2006. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20462597.

Anonymous:2006:FMc (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20462586.

Quinn:2007:MAD

[1890] Terry Quinn. More about Darwin and the evolution of species, Victorian spiritualism and Wittgenstein’s aeronautical researches. Notes and Records of the Royal Society of London, 61(1):1–3, January 22, 2007. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/20462601.

Noakes:2007:CVF

[1891] Richard Noakes. Cromwell Varley FRS, electrical discharge and Victorian spiritualism. Notes and Records of the Royal Society of London, 61(1):5– 21, January 22, 2007. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/20462602.

Orrego:2007:DIF

[1892] Fernando Orrego and Carlos Quintana. Darwin’s illness: a final diagnosis. Notes and Records of the Royal Society of London, 61(1):23–29, January 22, 2007. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/20462603.

deChadarevian:2007:SGO

[1893] Soraya de Chadarevian. The selfish gene at 30: the origin and career of a book and its title. Notes and Records of the Royal Society of London, 61(1):31–38, January 22, 2007. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20462604.

Lemco:2007:WAI

[1894] Ian Lemco. Wittgenstein’s aeronautical investigation. Notes and Records of the Royal Society of London, 61(1):39–51, January 22, 2007. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/20462605.

Taylor:2007:JHC

[1895] J. C. Taylor and A. W. Wolfendale. John Harrison: clockmaker and Copley Medalist. a public memorial at last. Notes and Records of the Royal Society of London, 61(1):53–62, January 22, 2007. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/20462606.

vanWyhe:2007:LRD

[1896] John van Wyhe and Antranig Basman. The launch and reception of Darwin online. Notes and Records of the Royal Society of London, 61(1):63– 65, January 22, 2007. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/20462607.

Rowlinson:2007:FE

[1897] J. S. Rowlinson. Flattening the earth. Notes and Records of the Royal Society of London, 61(1):67–68, January 22, 2007. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/20462608.

Woollett:2007:MPP

[1898] Christine Woollett. The Maull Photographic Portrait Collection held at the Royal Society. Notes and Records of the Royal Society of London, 61(1):69–74, January 22, 2007. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20462609.

Rees:2007:APL

[1899] M. J. Rees. Address of the President, Lord Rees of Ludlow Kt FRS, given at the Anniversary Meeting on 30 November 2006. Notes and Records of the Royal Society of London, 61(1):75–83, January 22, 2007. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/20462610.

Anonymous:2007:BMa

[1900] Anonymous. Back matter. Notes and Records of the Royal Society of London, 61(1):??, January 2007. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20462611.

Anonymous:2007:FMa

[1901] Anonymous. Front matter. Notes and Records of the Royal Society of London, 61(1):??, January 22, 2007. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20462599.

Quinn:2007:EDC

[1902] Terry Quinn. Émilie Du Châtelet, John Freind, Robert Hooke, Charles Darwin and John Stanley Gardiner. Notes and Records of the Royal Society of London, 61(2):85–88, May 22, 2007. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/ stable/20462615. See correction [1925].

Zinsser:2007:MMD

[1903] Judith P. Zinsser. Mentors, the Marquise Du Châtelet and historical memory. Notes and Records of the Royal Society of London, 61(2):89–108, May 22, 2007. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/20462616.

Rowlinson:2007:JFP

[1904] J. S. Rowlinson. John Freind: physician, chemist, Jacobite, and friend of Voltaire’s. Notes and Records of the Royal Society of London, 61(2):109– 127, May 22, 2007. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/20462617.

Henderson:2007:UMM

[1905] Felicity Henderson. Unpublished material from the memorandum book of Robert Hooke, Guildhall Library MS 1758. Notes and Records of the Royal Society of London, 61(2):129–175, May 22, 2007. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/20462618.

vanWyhe:2007:MGD

[1906] John van Wyhe. Mind the gap: did Darwin avoid publishing his theory for many years? Notes and Records of the Royal Society of London, 61(2):177–205, May 22, 2007. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20462619.

Brown:2007:LPJ

[1907] Barbara E. Brown. The legacy of Professor John Stanley Gardiner FRS to reef science. Notes and Records of the Royal Society of London, 61(2):207–217, May 22, 2007. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20462620.

Mills:2007:RHU

[1908] Allan Mills. Robert Hooke’s ‘universal joint’ and its application to sundials and the sundial-clock. Notes and Records of the Royal Society of London, 61(2):219–236, May 22, 2007. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20462621.

Anderson:2007:BPR

[1909] Clara Anderson. Blagden Papers at the Royal Society. Notes and Records of the Royal Society of London, 61(2):237–238, May 22, 2007. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/20462622.

Kohn:2007:PBK

[1910] Ralph Kohn. The pharmacological basis of the Kohn Foundation: lecture to commemorate the presentation of Dr Ralph Kohn FRS at the New Fellows’ Seminar 2006. Notes and Records of the Royal Society of London, 61(2):239–245, May 22, 2007. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20462623.

Anonymous:2007:BMb

[1911] Anonymous. Back matter. Notes and Records of the Royal Society of London, 61(2):??, May 2007. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20462624.

Anonymous:2007:FMb

[1912] Anonymous. Front matter. Notes and Records of the Royal Society of London, 61(2):??, May 22, 2007. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20462613.

Quinn:2007:OES

[1913] Terry Quinn. Origin and evolution of the species. Notes and Records of the Royal Society of London, 61(3):247–249, September 22, 2007. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/20462629.

Poole:2007:AFP

[1914] William Poole. Antoine-François Payen, the 1666 Selenelion, and a rediscovered letter to Robert Hooke. Notes and Records of the Royal Society of London, 61(3):251–263, September 22, 2007. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/20462630.

Grossman:2007:STM

[1915] Mark I. Grossman. Smithson Tennant: meteorites and the final trip to France. Notes and Records of the Royal Society of London, 61(3):265–283, September 22, 2007. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/20462631.

Sheehan:2007:JCA

[1916] William Sheehan and Steven Thurber. John couch Adams’s Asperger syndrome and the British non-discovery of Neptune. Notes and Records of the Royal Society of London, 61(3):285–299, September 22, 2007. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/20462632.

Brock:2007:RSG

[1917] William H. Brock. The Royal Society’s Glass Workers’ Cataract Committee; Sir William Crookes and the development of sunglasses. Notes and Records of the Royal Society of London, 61(3):301–312, September 22, 2007. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/20462633.

Chambers:2007:SAJ

[1918] N. A. Chambers. The Society of Arts and Joseph Banks: a first step in London learned society. Notes and Records of the Royal Society of London, 61(3):313–325, September 22, 2007. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/20462634.

Santer:2007:HIH

[1919] Melvin Santer. How it happened that a portion of a treatise entitled new improvements of planting and gardening both philosophical and practical by Richard Bradley FRS, which dealt with blights of trees and plants, provided the first report of an environment that contained green sulphur photosynthetic bacteria. Notes and Records of the Royal Society of London, 61(3):327–332, September 22, 2007. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/20462635.

Singh:2007:IPC

[1920] Rajinder Singh. India’s Physics and Chemistry Nobel Prize nominators and nominees in colonial and international context. Notes and Records of the Royal Society of London, 61(3):333–345, September 22, 2007. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/20462636.

Pohancenik:2007:PWP

[1921] Rebecca Pohancenik. Pharmacology and war: the papers of Sir John Henry Gaddum (1900–65). Notes and Records of the Royal Society of London, 61(3):347–348, September 22, 2007. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/20462637.

Elliott:2007:BB

[1922] Clark A. Elliott. On becoming Ben. Notes and Records of the Royal Society of London, 61(3):349–350, September 22, 2007. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/20462638.

Wulf:2007:BRC

[1923] Andrea Wulf. Book review: Collecting the Empire: Neil Chambers, Joseph Banks and the British Museum: the world of collecting. 1770– 1830, Pickering & Chatto, London, 2007. pp. 195 £60.00 (hardback). ISBN 978-1-85196-858-9. Notes and Records of the Royal Society of London, 61(3):351–352, September 22, 2007. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/20462639.

Heilbron:2007:BFE

[1924] J. L. Heilbron. Benjamin Franklin in Europe: electrician, academician, politician. Notes and Records of the Royal Society of London, 61(3): 353–373, September 22, 2007. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20462640.

Quinn:2007:CQE

[1925] Terry Quinn. Correction for Quinn, Émilie Du Châtelet, John Freind, Robert Hooke, Charles Darwin and John Stanley Gardiner. Notes and Records of the Royal Society of London, 61(3):375, September 22, 2007. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/20462641. See [1902].

Anonymous:2007:BMc

[1926] Anonymous. Back matter. Notes and Records of the Royal Society of London, 61(3):??, September 2007. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20462642.

Anonymous:2007:FMc

[1927] Anonymous. Front matter. Notes and Records of the Royal Society of London, 61(3):??, September 2007. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20462627.

Fox:2008:Ea

[1928] Robert Fox. Editorial. Notes and Records of the Royal Society of London, 62(1):1, March 20, 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20462646.

Iliffe:2008:T

[1929] Rob Iliffe. Technicians. Notes and Records of the Royal Society of London, 62(1):3–16, March 20, 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20462647.

Stewart:2008:AEW

[1930] Larry Stewart. Assistants to enlightenment: William Lewis, Alexander Chisholm and invisible technicians in the industrial revolution. Notes and Records of the Royal Society of London, 62(1):17–29, March 20, 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/20462648.

Jackson:2008:VWR

[1931] Catherine M. Jackson. Visible work: the role of students in the creation of Liebig’s Giessen research school. Notes and Records of the Royal Society of London, 62(1):31–49, March 20, 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/20462649.

Gay:2008:TAW

[1932] Hannah Gay. Technical assistance in the world of London science, 1850– 1900. Notes and Records of the Royal Society of London, 62(1):51–75, March 20, 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/20462650.

Tansey:2008:KCA

[1933] E. M. Tansey. Keeping the culture alive: the laboratory technician in mid-twentieth-century British medical research. Notes and Records of the Royal Society of London, 62(1):77–95, March 20, 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/20462651.

Hughes:2008:WKS

[1934] Jeff Hughes. William Kay, Samuel Devons and memories of practice in Rutherford’s Manchester laboratory. Notes and Records of the Royal Society of London, 62(1):97–121, March 20, 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/20462652.

Tansey:2008:WCS

[1935] E. M. Tansey. Working with C. S. Sherrington, 1918–24. Notes and Records of the Royal Society of London, 62(1):123–130, March 20, 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/20462653.

Tansey:2008:WCP

[1936] E. M. Tansey. Working with Cambridge physiologists. Notes and Records of the Royal Society of London, 62(1):131–137, March 20, 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/20462654.

Martin:2008:CNP

[1937] John Martin. A career at the National Physical Laboratory. Notes and Records of the Royal Society of London, 62(1):139–148, March 20, 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/20462655.

Anonymous:2008:BMa

[1938] Anonymous. Back matter. Notes and Records of the Royal Society of London, 62(1):??, March 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20462656.

Anonymous:2008:FMa

[1939] Anonymous. Front matter. Notes and Records of the Royal Society of London, 62(1):??, March 20, 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20462644.

Fox:2008:Eb

[1940] Robert Fox. Editorial. Notes and Records of the Royal Society of London, 62(2):149–150, June 20, 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20462660.

Cartwright:2008:THT

[1941] Julyan H. E. Cartwright and Hisami Nakamura. Tsunami: a history of the term and of scientific understanding of the phenomenon in Japanese and Western culture. Notes and Records of the Royal Society of London, 62(2):151–166, June 20, 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20462661.

Fryer:2008:JCS

[1942] Geoffrey Fryer. Jacob Christian Schäffer FRS, a versatile eighteenthcentury naturalist, and his remarkable pioneering researches on microscopic crustaceans. Notes and Records of the Royal Society of London, 62(2):167–185, June 20, 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20462662.

Tobin:2008:FTS

[1943] William Tobin. Full-text search capability: a new tool for researching the development of scientific language. The ‘Whirlpool Nebula’ as a case study. Notes and Records of the Royal Society of London, 62(2):187–196, June 20, 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/20462663.

Graham-Smith:2008:DRT

[1944] Sir Francis Graham-Smith and Sir Bernard Lovell. Diversions of a radio telescope. Notes and Records of the Royal Society of London, 62(2):197– 204, June 20, 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/20462664.

Sheehan:2008:MDI

[1945] William Sheehan, William H. Meller, and Steven Thurber. More on Darwin’s illness: comment on the final diagnosis of Charles Darwin. Notes and Records of the Royal Society of London, 62(2):205–209, June 20, 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/20462665.

Anonymous:2008:APL

[1946] Anonymous. Address of the President, Lord Rees of Ludlow OM Kt PRS, given at the Anniversary Meeting on 30 November 2007. Notes and Records of the Royal Society of London, 62(2):211–221, June 20, 2008.

CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/20462666.

Henry:2008:HOS

[1947] John Henry. Historical and other studies of science, technology and medicine in the University of Edinburgh. Notes and Records of the Royal Society of London, 62(2):223–235, June 20, 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/20462667.

Jedlitschka:2008:AGA

[1948] Karsten Jedlitschka. The Archive of the German Academy of Sciences Leopoldina in Halle (Saale): more than 350 years of the history of science. Notes and Records of the Royal Society of London, 62(2):237–244, June 20, 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/20462668.

Moore:2008:GCW

[1949] Keith Moore. A grangerized copy of Weld’s history of the Royal Society. Notes and Records of the Royal Society of London, 62(2):245–250, June 20, 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/20462669.

Moorbath:2008:BRO

[1950] S. Moorbath. Book review: The oldest geological society. Gordon L. Herries Davies, Whatever is under the Earth: The Geological Society of London 1807–2007. The Geological Society London, 2007. pp. xiii + 356, £50.00 (hardback). ISBN 978-1-86239-214-4. Notes and Records of the Royal Society of London, 62(2):251–252, June 20, 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/20462670.

Glynn:2008:RFY

[1951] Jenifer Glynn. Rosalind Franklin: 50 years on. Notes and Records of the Royal Society of London, 62(2):253–255, June 20, 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/20462671.

Anonymous:2008:BMb

[1952] Anonymous. Back matter. Notes and Records of the Royal Society of London, 62(2):??, June 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20462672.

Anonymous:2008:FMb

[1953] Anonymous. Front matter. Notes and Records of the Royal Society of London, 62(2):??, June 20, 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20462658.

Fox:2008:Ec

[1954] Robert Fox. Editorial. Notes and Records of the Royal Society of London, 62(3):257–258, September 20, 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20462676.

Simoes:2008:QCD

[1955] Ana Simões. A quantum chemical dialogue mediated by textbooks: Pauling’s The nature of the chemical bond and Coulson’s Valence. Notes and Records of the Royal Society of London, 62(3):259–269, September 20, 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/20462677.

Roos:2008:MCM

[1956] Anna Marie Roos. ‘Magic coins’ and ‘magic squares’: the discovery of astrological sigils in the Oldenburg Letters. Notes and Records of the Royal Society of London, 62(3):271–288, September 20, 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/20462678.

Keynes:2008:BNM

[1957] Milo Keynes. Balancing Newton’s mind: his singular behaviour and his madness of 1692–93. Notes and Records of the Royal Society of London, 62(3):289–300, September 20, 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20462679.

Robinson:2008:VCA

[1958] Pamela Robinson. A ‘very curious Almanack’: the gift of Sir Robert Moray FRS, 1668. Notes and Records of the Royal Society of London, 62 (3):301–314, September 20, 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20462680.

Moore:2008:RSW

[1959] Keith Moore. The Royal Society’s war committee on engineering 1914– 19. Notes and Records of the Royal Society of London, 62(3):315–319, September 20, 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/20462681.

Fenner:2008:AAS

[1960] Frank Fenner. The Australian Academy of Science: the first 50 years. Notes and Records of the Royal Society of London, 62(3):321–330, September 20, 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/20462682.

Dunn:2008:IG

[1961] Richard Dunn. An infuriating genius. Notes and Records of the Royal Society of London, 62(3):331–334, September 20, 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/20462683.

Anderson:2008:P

[1962] Stuart Anderson. Penicillin. Notes and Records of the Royal Society of London, 62(3):335–336, September 20, 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/20462684.

Anonymous:2008:BMc

[1963] Anonymous. Back matter. Notes and Records of the Royal Society of London, 62(3):??, September 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20462685.

Anonymous:2008:FMc

[1964] Anonymous. Front matter. Notes and Records of the Royal Society of London, 62(3):??, September 20, 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/20462674.

Anonymous:2008:ISC

[1965] Anonymous. [Illustration]: Sir Charles Blagden FRS (ca. 1748–1820). Notes and Records of the Royal Society of London, 62(4):i–ii, December 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/20462699.

Fox:2008:Ed

[1966] Robert Fox. Editorial. Notes and Records of the Royal Society of London, 62(4):337–339, December 20, 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20462689.

Turner:2008:ISF

[1967] Anthony Turner. An interrupted story: French translations from Philosophical Transactions in the seventeenth and eighteenth centuries. Notes and Records of the Royal Society of London, 62(4):341–354, December 20, 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/20462690.

Martin:2008:JOP

[1968] Jean-Pierre Martin and Anita McConnell. Joining the observatories of Paris and Greenwich. Notes and Records of the Royal Society of London, 62(4):355–372, December 20, 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20462691.

Fauque:2008:EAC

[1969] Danielle M. E. Fauque. An Englishman abroad: Charles Blagden’s visit to Paris in 1783. Notes and Records of the Royal Society of London, 62 (4):373–390, December 20, 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20462692.

Carrell:2008:BML

[1970] Robin Carrell. Bringing molecules to life. Notes and Records of the Royal Society of London, 62(4):391–395, December 20, 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/20462693.

Chambers:2008:WPC

[1971] Neil A. Chambers. The world of plant collecting: an eighteenth-century obsession. Notes and Records of the Royal Society of London, 62(4): 397–399, December 20, 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20462694.

Elliott:2008:FP

[1972] Roger Elliott. Friendship and physics. Notes and Records of the Royal Society of London, 62(4):401–403, December 20, 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/20462695.

Roos:2008:ELS

[1973] Anna Marie Roos. Elizabethan London and scientific practice. Notes and Records of the Royal Society of London, 62(4):405–406, December 20, 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/20462696.

Rattansi:2008:RLE

[1974] Piyo M. Rattansi. Rediscovery of lost early Royal Society papers on the alkahest. Notes and Records of the Royal Society of London, 62(4):407, December 20, 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/20462697.

Moore:2008:PFS

[1975] Keith Moore. Photographs by the French scientist Henri Victor Regnault. Notes and Records of the Royal Society of London, 62(4):409–413, December 20, 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/20462698.

Anonymous:2008:BMd

[1976] Anonymous. Back matter. Notes and Records of the Royal Society of London, 62(4):??, December 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20462700.

Anonymous:2008:FMd

[1977] Anonymous. Front matter. Notes and Records of the Royal Society of London, 62(4):??, December 2008. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20462688.

Fox:2009:JUT

[1978] Robert Fox (Notes, Records of the Royal Society), Hanne Andersen (Centaurus), Roger Ariew, Moti Feingold (Perspectives on Science), A. K. Bag (Indian Journal of History of Science), June Barrow-Green, Benno van Dalen (Historia mathematica), Keith Benson (History, Philosophy of the Life Sciences), Marco Beretta (Nuncius), Michel Blay (Revue d’Histoire des Sciences), Cornelius Borck (Berichte zur Wissenschaftsgeschichte), Geof Bowker, Technology Susan Leigh Star (Science, Human Values), Massimo Bucciantini, Michele Camerota (Galilaeana: Journal of Galilean Studies), Jed Buchwald, Jeremy Gray (Archive for History of Exact Sciences), Vincenzo Cappelletti, Guido Cimino (Physis), Mark Clark, Alex Keller (ICON), Roger Cline (International Journal for the History of Engineering, Technology), Stephen Clucas, Stephen Gaukroger (Intellectual History Review), Hal Cook, Anne Hardy (Medical History), Leo Corry, Alexandre Métraux, Jürgen Renn (Science in Context), Brian Dolan, Bill Luckin (Social History of Medicine), Hilmar Duerbeck, Wayne Orchiston (Journal of Astronomical History, Heritage), Moritz Epple, Mikael Hard, Hans-Jorg Rheinberger, Technik und Medizin) Volker Roelcke (NTM: Zeitschrift für Geschichte der Wissenschaften, Steven French (Metascience), Paul Farber (Journal of the History of Biology), Mary Fissell, Randall Packard (Bulletin of the History of Medicine), Jim Good (History of the Human Sciences), Willem Hackmann (Bulletin of the Scientific Instrument Society), Robert Halleux (Archives Internationales d’Histoire des Sciences), Bosse Holmqvist (Lychnos), Michael Hoskin (Journal for the History of Astronomy), Ian Inkster (History of Technology), Marina Frasca Spada (Studies in History, Philosophy of Science), Nick Jardine (Studies in History, Philosophy of Biological, Biomedical Sciences), Trevor Levere (Annals of Science), Bernard Lightman (Isis), Christoph Luthy (Early Science, Medicine), Michael Lynch (Social Studies of Science), Stephen McCluskey, Clive Ruggles (Archaeostronomy: the Journal of Astronomy in Culture), Peter Morris (Ambix), E. Charles Nelson (Archives of Natural History), Ian Nicholson (Journal of the History of the Behavioural Sciences), Kathy Olesko (Osiris), Liliane Perez (Documents pour l’Histoire des Techniques), Iwan Rhys Morus (History of Science), John Rigden, Roger H. Stuewer (Physics in Perspective), Julio Samsó (Suhayl: Journal for the History of the Exact, Natural Sciences in Islamic Civilisation), Simon Schaffer (British Journal for the History of Science), Norbert Shappacher (Revue d’Histoire des Mathématiques), Claire Strom (Agricultural History), Paul Unschuld (Sudhoffs Archiv), Peter Weingart (Minerva), Stefan Zamecki (Kwartalnik Historii Nauki i Techniki), and Huib Zuidervaart (Studium. Tijdschrift Voor Wetenschaps en Universiteitsgeschiedenis/ Revue de l’Histoire des Sciences et des Universités). Journals under threat: a joint response from history of science, technology and medicine editors. Notes and Records of the Royal Society of London, 63(1):1–3, March 20, 2009. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/40647189.

Fox:2009:Ea

[1979] Robert Fox. Editorial. Notes and Records of the Royal Society of London, 63(1):5–6, March 20, 2009. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/ 40647190. See correction [2000].

Unwin:2009:HDR

[1980] Patrick R. Unwin and Robert W. Unwin. Humphry Davy and the Royal Institution of Great Britain. Notes and Records of the Royal Society of London, 63(1):7–33, March 20, 2009. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/40647191.

Werrett:2009:WCR

[1981] Simon Werrett. William Congreve’s rational rockets. Notes and Records of the Royal Society of London, 63(1):35–56, March 20, 2009. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/40647192.

Kenworthy:2009:CMS

[1982] Joan M. Kenworthy and Margaret S. McCollum. A contribution to meteorology by Spencer Cowper, Dean of Durham 1746–74. Notes and Records of the Royal Society of London, 63(1):57–80, March 20, 2009. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/40647193.

Baldwin:2009:WYI

[1983] Melinda Baldwin. ‘Where are your intelligent mothers to come from?’: marriage and family in the scientific career of Dame Kathleen Lonsdale FRS (1903–71). Notes and Records of the Royal Society of London, 63(1):81–94, March 20, 2009. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

40647194.

Cater:2009:WCC

[1984] John Cater and Ian Lemco. Wittgenstein’s combustion chamber. Notes and Records of the Royal Society of London, 63(1):95–104, March 20, 2009. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/40647195.

McQuillan:2009:DSE

[1985] A. James McQuillan. The discovery of surface-enhanced Raman scattering. Notes and Records of the Royal Society of London, 63(1):105–109, March 20, 2009. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/40647196. See cor-

rections [2001].

Hughes:2009:FAW

[1986] David W. Hughes. The family of the astronomers William and Caroline Herschel. Notes and Records of the Royal Society of London, 63(1):111– 112, March 20, 2009. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/40647197.

Bruyns:2009:MC

[1987] Willem F. J. Mörzer Bruyns. Meeting cultures. Notes and Records of the Royal Society of London, 63(1):113–114, March 20, 2009. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/40647198.

Anderson:2009:GBS

[1988] Robert G. W. Anderson. Gentleman-benefactor of science. Notes and Records of the Royal Society of London, 63(1):115–116, March 20, 2009. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/40647199.

Anonymous:2009:BMa

[1989] Anonymous. Back matter. Notes and Records of the Royal Society of London, 63(1):??, March 2009. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

40647200.

Anonymous:2009:FMa

[1990] Anonymous. Front matter. Notes and Records of the Royal Society of London, 63(1):??, March 20, 2009. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

40647188.

Fox:2009:Eb

[1991] Robert Fox. Editorial. Notes and Records of the Royal Society of London, 63(2):117–118, June 20, 2009. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

40647252.

Cartwright:2009:WKW

[1992] Julyan H. E. Cartwright and Hisami Nakamura. What kind of a wave is Hokusai’s Great Wave off Kanagawa? Notes and Records of the Royal Society of London, 63(2):119–135, June 20, 2009. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/40647253.

Bellhouse:2009:MKC

[1993] D. R. Bellhouse, E. M. Renouf, R. Raut, and M. A. Bauer. De Moivre’s knowledge community: an analysis of the subscription list to the miscellanea analytica. Notes and Records of the Royal Society of London, 63(2):137–162, June 20, 2009. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

40647254.

Pumfrey:2009:HMM

[1994] Stephen Pumfrey. Harriot’s maps of the Moon: new interpretations. Notes and Records of the Royal Society of London, 63(2):163–168, June 20, 2009. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/40647255.

Andrews:2009:WLW

[1995] Elizabeth B. Andrews. Windows on a Lilliputian world: a personal perspective on the development of electron microscopy in the twentieth century. Notes and Records of the Royal Society of London, 63(2):169–181, June 20, 2009. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/40647256.

Anonymous:2009:APL

[1996] Anonymous. Address of the President, Lord Rees of Ludlow OM Kt PRS, given at the Anniversary Meeting on 1 December 2008. Notes and Records of the Royal Society of London, 63(2):183–190, June 20, 2009. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/40647257.

Moore:2009:UMB

[1997] P. G. Moore. The University Marine Biological Station Millport: in the beginning was the vision (1970). Notes and Records of the Royal Society of London, 63(2):191–202, June 20, 2009. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/40647258.

Knight:2009:AW

[1998] David Knight. The age of wonder. Notes and Records of the Royal Society of London, 63(2):203–206, June 20, 2009. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/40647259.

Bryden:2009:MEL

[1999] D. J. Bryden. Made in England — but not in London! Notes and Records of the Royal Society of London, 63(2):207–208, June 20, 2009. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/40647260.

Fox:2009:CE

[2000] Robert Fox. Correction to editorial. Notes and Records of the Royal Society of London, 63(2):209, June 20, 2009. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/40647262. See [1979].

McQuillan:2009:CRD

[2001] A. James McQuillan. Corrections: Recollection. The discovery of surfaceenhanced Raman scattering. Notes and Records of the Royal Society of London, 63(2):209, June 20, 2009. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/ 40647262. See [1985].

Anonymous:2009:BMb

[2002] Anonymous. Back matter. Notes and Records of the Royal Society of London, 63(2):??, June 2009. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

40647263.

Anonymous:2009:FMb

[2003] Anonymous. Front matter. Notes and Records of the Royal Society of London, 63(2):??, June 20, 2009. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

40647250.

Fox:2009:MWT

[2004] Robert Fox. The many worlds of Thomas Beddoes. Notes and Records of the Royal Society of London, 63(3):211–213, September 20, 2009. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/40647274.

Levere:2009:DTB

[2005] Trevor H. Levere. Dr Thomas Beddoes: chemistry, medicine, and the perils of democracy. Notes and Records of the Royal Society of London, 63(3):215–229, September 20, 2009. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

40647275.

Stewart:2009:HMS

[2006] Larry Stewart. His Majesty’s subjects: from laboratory to human experiment in pneumatic chemistry. Notes and Records of the Royal Society of London, 63(3):231–245, September 20, 2009. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/40647276.

Pancaldi:2009:HOT

[2007] Giuliano Pancaldi. On hybrid objects and their trajectories: Beddoes, Davy and the battery. Notes and Records of the Royal Society of London, 63(3):247–262, September 20, 2009. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

40647277.

Morus:2009:RRE

[2008] Iwan Rhys Morus. Radicals, romantics and electrical showmen: placing galvanism at the end of the English Enlightenment. Notes and Records of the Royal Society of London, 63(3):263–275, September 20, 2009. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/40647278.

Rousseau:2009:PGD

[2009] George Rousseau. Political gout: dissolute patients, deceitful physicians, and other blue devils. Notes and Records of the Royal Society of London, 63(3):277–296, September 20, 2009. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

40647279.

Jay:2009:AHN

[2010] Mike Jay. The atmosphere of heaven: the 1799 nitrous oxide researches reconsidered. Notes and Records of the Royal Society of London, 63 (3):297–309, September 20, 2009. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

40647280.

Vickers:2009:TBG

[2011] Neil Vickers. Thomas Beddoes and the German psychological tradition. Notes and Records of the Royal Society of London, 63(3):311–321, September 20, 2009. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/40647281.

Anonymous:2009:BMc

[2012] Anonymous. Back matter. Notes and Records of the Royal Society of London, 63(3):??, September 2009. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

40647282.

Anonymous:2009:Fa

[2013] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 63(3):NP, September 20, 2009. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2009:FMc

[2014] Anonymous. Front matter. Notes and Records of the Royal Society of London, 63(3):??, September 20, 2009. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/40647273.

Fox:2009:Ec

[2015] Robert Fox. Editorial. Notes and Records of the Royal Society of London, 63(4):323–324, December 20, 2009. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

40647307.

Wardhaugh:2009:MEP

[2016] Benjamin Wardhaugh. Mathematics in English printed books, 1473– 1800: a bibliometric analysis. Notes and Records of the Royal Society of London, 63(4):325–338, December 20, 2009. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/40647308.

Balmer:2009:RSB

[2017] Brian Balmer, Matthew Godwin, and Jane Gregory. The Royal Society and the ‘brain drain’: natural scientists meet social science. Notes and Records of the Royal Society of London, 63(4):339–353, December 20, 2009. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/40647309.

Guagnini:2009:JFM

[2018] Anna Guagnini. John Fletcher Moulton and Guglielmo Marconi: bridging science, law and industry. Notes and Records of the Royal Society of London, 63(4):355–363, December 20, 2009. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/40647310.

Grattan-Guinness:2009:BRM

[2019] Ivor Grattan-Guinness. Bertrand Russell (1872–1970), man of dissent. Notes and Records of the Royal Society of London, 63(4):365–379, December 20, 2009. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/40647311.

Moorbath:2009:DEO

[2020] Stephen Moorbath. The discovery of the Earth’s oldest rocks. Notes and Records of the Royal Society of London, 63(4):381–392, December 20, 2009. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/40647312.

Bowler:2009:DWN

[2021] Peter J. Bowler. Do we need a non-Darwinian industry? Notes and Records of the Royal Society of London, 63(4):393–398, December 20, 2009. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/40647313.

Morrell:2009:HG

[2022] Jack Morrell. Homeric geologists. Notes and Records of the Royal Society of London, 63(4):399–406, December 20, 2009. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/40647314.

Anonymous:2009:BMd

[2023] Anonymous. Back matter. Notes and Records of the Royal Society of London, 63(4):??, December 2009. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

40647315.

Anonymous:2009:Fb

[2024] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 63(4):NP, December 20, 2009. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2009:FMd

[2025] Anonymous. Front matter. Notes and Records of the Royal Society of London, 63(4):??, December 20, 2009. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/40647306.

Collins:2010:E

[2026] Peter Collins. Editorial. Notes and Records of the Royal Society of London, 64(S1):S1–S3, September 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Weatherall:2010:MMR

[2027] D. J. Weatherall. Molecular medicine: the road to the better integration of the medical sciences in the twenty-first century. Notes and Records of the Royal Society of London, 64(S1):S5–S15, September 20, 2010. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Bud:2010:AMB

[2028] Robert Bud. From applied microbiology to biotechnology: science, medicine and industrial renewal. Notes and Records of the Royal Society of London, 64(S1):S17–S29, September 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Horrocks:2010:RSF

[2029] Sally M. Horrocks. The Royal Society, its Fellows and industrial R&D in the mid twentieth century. Notes and Records of the Royal Society of London, 64(S1):S31–S41, September 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Collins:2010:RST

[2030] Peter Collins. A Royal Society for technology. Notes and Records of the Royal Society of London, 64(S1):S43–S54, September 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Snowden:2010:TII

[2031] Christopher M. Snowden. Technological innovation in industry and the role of the Royal Society. Notes and Records of the Royal Society of London, 64(S1):S55–S63, September 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Pounds:2010:RSF

[2032] Ken Pounds. The Royal Society’s formative role in UK space research. Notes and Records of the Royal Society of London, 64(S1):S65–S76, September 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Morris:2010:DMR

[2033] Peter J. T. Morris. Does membership of the Royal Society affect careers? A trio in twentieth-century organic chemistry. Notes and Records of the Royal Society of London, 64(S1):S77–S87, September 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Hemmen:2010:RSE

[2034] George E. Hemmen. Royal Society expeditions in the second half of the twentieth century. Notes and Records of the Royal Society of London, 64 (S1):S89–S99, September 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Hughes:2010:DRD

[2035] Jeff Hughes. ‘Divine right’ or democracy? The Royal Society ‘revolt’ of 1935. Notes and Records of the Royal Society of London, 64(S1):S101– S117, September 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Collins:2010:RRU

[2036] Peter Collins. A role in running UK science? Notes and Records of the Royal Society of London, 64(S1):S119–S130, September 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Cox:2010:RSC

[2037] Stephen Cox. The Royal Society in Cold War Europe. Notes and Records of the Royal Society of London, 64(S1):S131–S136, September 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

MacLeod:2010:RSC

[2038] Roy MacLeod. The Royal Society and the Commonwealth: old friendships, new frontiers. Notes and Records of the Royal Society of London, 64(S1):S137–S149, September 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Bodmer:2010:PUS

[2039] Walter Bodmer. Public understanding of science: The BA, the Royal Society and COPUS. Notes and Records of the Royal Society of London, 64(S1):S151–S161, September 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Ferry:2010:ERW

[2040] Georgina Ferry. The exception and the rule: women and the Royal Society 1945–2010. Notes and Records of the Royal Society of London, 64(S1):S163–S172, September 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Hughes:2010:ICF

[2041] Jeff Hughes. Introductory comments to final discussion session. Notes and Records of the Royal Society of London, 64(S1):S173–S176, September 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Howard:2010:ICF

[2042] Judith Howard. Introductory comments to final discussion session. Notes and Records of the Royal Society of London, 64(S1):S177–S179, September 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2010:Fc

[2043] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 64(S1):NP, September 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Fox:2010:Ea

[2044] Robert Fox. Editorial. Notes and Records of the Royal Society of London, 64(1):1–3, March 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

40647329.

Lightman:2010:DPE

[2045] Bernard Lightman. Darwin and the popularization of evolution. Notes and Records of the Royal Society of London, 64(1):5–24, March 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/40647330.

Vetter:2010:UAW

[2046] Jeremy Vetter. The unmaking of an anthropologist: Wallace returns from the field, 1862–70. Notes and Records of the Royal Society of London, 64(1):25–42, March 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

40647331.

Rowlinson:2010:JJW

[2047] J. S. Rowlinson. James Joule, William Thomson and the concept of a perfect gas. Notes and Records of the Royal Society of London, 64(1):43– 57, March 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/40647332.

Santer:2010:JLF

[2048] Melvin Santer. Joseph Lister: first use of a bacterium as a ‘model organism’ to illustrate the cause of infectious disease of humans. Notes and Records of the Royal Society of London, 64(1):59–65, March 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/40647333.

Boschiero:2010:TES

[2049] Luciano Boschiero. Translation, experimentation and the spring of the air: Richard Waller’s essayes of natural experiments. Notes and Records of the Royal Society of London, 64(1):67–83, March 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/40647334.

Hurst:2010:TYR

[2050] Phil Hurst. Trailblazing — 350 years of Royal Society publishing. Notes and Records of the Royal Society of London, 64(1):85–89, March 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/40647335.

Hughes:2010:CC

[2051] Jeff Hughes. A Crookes charter. Notes and Records of the Royal Society of London, 64(1):91–93, March 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/40647336.

Robinson:2010:PEH

[2052] Derek A. Robinson. A political and evolutionary history of Oxford chemistry. Notes and Records of the Royal Society of London, 64(1):95–96, March 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/40647337.

Noble:2010:FPD

[2053] Denis Noble. Funding the pink diamonds: a historical perspective. Notes and Records of the Royal Society of London, 64(1):97–102, March 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/40647338.

Anonymous:2010:BMa

[2054] Anonymous. Back matter. Notes and Records of the Royal Society of London, 64(1):??, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/40647339.

Anonymous:2010:Fa

[2055] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 64(1):NP, March 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2010:FMa

[2056] Anonymous. Front matter. Notes and Records of the Royal Society of London, 64(1):??, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

40647327.

Fox:2010:Eb

[2057] Robert Fox. Editorial. Notes and Records of the Royal Society of London, 64(2):103–104, June 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Roos:2010:SCP

[2058] Anna Marie Roos. A speculum of chymical practice: Isaac Newton, Martin Lister (1639–1712), and the making of telescopic mirrors. Notes and Records of the Royal Society of London, 64(2):105–120, June 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Cook:2010:LCP

[2059] Alexandra Cook. Linnaeus and Chinese plants: A test of the linguistic imperialism thesis. Notes and Records of the Royal Society of London, 64 (2):121–138, June 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Forsdyke:2010:GRW

[2060] Donald R. Forsdyke. George Romanes, William Bateson, and Darwin’s ‘weak point’. Notes and Records of the Royal Society of London, 64(2): 139–154, June 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Freire:2010:DMM

[2061] Olival Freire and Christoph Lehner. ‘Dialectical materialism and modern physics’, an unpublished text by Max Born. Notes and Records of the Royal Society of London, 64(2):155–162, June 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2010:APL

[2062] Anonymous. Address of the President, Lord Rees of Ludlow OM Kt PRS, given at the Anniversary Meeting on 30 November 2009. Notes and Records of the Royal Society of London, 64(2):163–173, June 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Creighton:2010:CED

[2063] J. A. Creighton. Contributions to the early development of surfaceenhanced Raman spectroscopy. Notes and Records of the Royal Society of London, 64(2):175–183, June 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Gull:2010:BCP

[2064] Keith Gull. Boveri and cancer: prescient views of molecular mechanisms. Notes and Records of the Royal Society of London, 64(2):185–187, June 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Insley:2010:KME

[2065] Jane Insley. Kettle as myth, engineer as chemist. Notes and Records of the Royal Society of London, 64(2):189–190, June 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Moore:2010:NPC

[2066] Keith Moore. New portraits at Carlton House Terrace. Notes and Records of the Royal Society of London, 64(2):191–196, June 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2010:Fb

[2067] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 64(2):NP, June 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Fox:2010:Ec

[2068] Robert Fox. Editorial. Notes and Records of the Royal Society of London, 64(3):197–199, September 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20753899.

Bragg:2010:WMB

[2069] Melvyn Bragg. Wilkins Medawar Bernal Lecture 2010: Observations of an amateur on the history of the Royal Society 1660–2010. Notes and Records of the Royal Society of London, 64(3):201–212, September 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/20753900.

Anonymous:2010:SGM

[2070] Anonymous. A sermon given by the Most Reverend & Right Honourable Rowan Williams, Archbishop of Canterbury, at an Evensong attended by Fellows of the Royal Society on the occasion of the 350th anniversary, St Paul’s Cathedral, 9 June 2010. Notes and Records of the Royal Society of London, 64(3):213–215, September 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/20753901.

Anonymous:2010:CFR

[2071] Anonymous. Convocation of the Fellowship of the Royal Society at the Royal Festival Hall, 23 June 2010. Notes and Records of the Royal Society of London, 64(3):217–227, September 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/20753902.

Hartley:2010:ECK

[2072] Beryl Hartley. Exploring and communicating knowledge of trees in the early Royal Society. Notes and Records of the Royal Society of London, 64(3):229–250, September 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20753903.

Hunter:2010:HFF

[2073] Matthew C. Hunter. Hooke’s figurations: a figural drawing attributed to Robert Hooke. Notes and Records of the Royal Society of London, 64 (3):251–260, September 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20753904.

Tomory:2010:WBP

[2074] Leslie Tomory. William Brownrigg’s papers on fire-damps. Notes and Records of the Royal Society of London, 64(3):261–270, September 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/20753905.

VanderKloot:2010:WMB

[2075] William Van der Kloot. William Maddock Bayliss’s therapy for wound shock. Notes and Records of the Royal Society of London, 64(3):271–286, September 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/20753906.

Melford:2010:TIG

[2076] David Melford, Peter Duncumb, Mike Stowell, and Bill Graham. Tube Investments Group Research Laboratory, Hinxton Hall (1954–88). Notes and Records of the Royal Society of London, 64(3):287–301, September 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/20753907.

Schemmel:2010:BC

[2077] Matthias Schemmel. Before calculus. Notes and Records of the Royal Society of London, 64(3):303–304, September 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/20753908.

Ayres:2010:AC

[2078] Peter Ayres. The artful collectors. Notes and Records of the Royal Society of London, 64(3):305–306, September 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/20753909.

Jay:2010:TW

[2079] Mike Jay. Testing the waters. Notes and Records of the Royal Society of London, 64(3):306–307, September 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/20753910.

Fyfe:2010:DHO

[2080] Aileen Fyfe. Darwin and human origins. Notes and Records of the Royal Society of London, 64(3):307–309, September 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/20753911.

Smith:2010:PAU

[2081] Robert C. Smith. Practical astronomy at UK universities. Notes and Records of the Royal Society of London, 64(3):309–311, September 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/20753912.

Hurst:2010:NRE

[2082] Phil Hurst. The Notes and Records essay award. Notes and Records of the Royal Society of London, 64(3):313, September 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/20753913.

Anonymous:2010:BMb

[2083] Anonymous. Back matter. Notes and Records of the Royal Society of London, 64(3):??, September 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

20753914.

Anonymous:2010:Fd

[2084] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 64(3):NP, September 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2010:FMb

[2085] Anonymous. Front matter. Notes and Records of the Royal Society of London, 64(3):??, September 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/20753898.

Fox:2010:Ed

[2086] Robert Fox. Editorial. Notes and Records of the Royal Society of London, 64(4):315–317, December 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

25802122.

Mitchell:2010:RNC

[2087] Daniel Jon Mitchell. Reflecting nature: chemistry and comprehensibility in Gabriel Lippmann’s ‘physical’ method of photographing colours. Notes and Records of the Royal Society of London, 64(4):319–337, December 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/25802123.

Lightman:2010:MLC

[2088] Bernard Lightman. The many lives of Charles Darwin: early biographies and the definitive evolutionist. Notes and Records of the Royal Society of London, 64(4):339–358, December 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/25802124.

Cain:2010:JHG

[2089] Joe Cain. Julian Huxley, general biology and the London zoo, 1935– 42. Notes and Records of the Royal Society of London, 64(4):359–378, December 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/25802125.

Bjorkman:2010:SEC

[2090] Maria Björkman and Sven Widmalm. Selling eugenics: the case of Sweden. Notes and Records of the Royal Society of London, 64(4):379–400, December 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/25802126.

Bishop:2010:SJT

[2091] Malcolm Bishop and Melanie Parker. Sir John Tomes FRS, Fellows of the Royal Society, and dental reform in the nineteenth century. Notes and Records of the Royal Society of London, 64(4):401–416, December 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/25802127.

Grossman:2010:WHD

[2092] Mark I. Grossman. William Higgins at the Dublin Society, 1810–20: the loss of a professorship and a claim to the atomic theory. Notes and Records of the Royal Society of London, 64(4):417–434, December 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/25802128.

Alfonso-Goldfarb:2010:LRS

[2093] Ana Maria Alfonso-Goldfarb, Márcia Helena Mendes Ferraz, and Piyo M. Rattansi. Lost Royal Society documents on ‘alkahest’ (universal solvent) rediscovered. Notes and Records of the Royal Society of London, 64 (4):435–456, December 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

25802129.

Wess:2010:CRT

[2094] Jane Wess. Crookes’s radiometers: a train of thought manifest. Notes and Records of the Royal Society of London, 64(4):457–470, December 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/25802130.

Anderson:2010:SMS

[2095] Robert G. W. Anderson. Science museums and the Science Museum. Notes and Records of the Royal Society of London, 64(4):471–476, December 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/25802131.

Anonymous:2010:BMc

[2096] Anonymous. Back matter. Notes and Records of the Royal Society of London, 64(4):??, December 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

25802132.

Anonymous:2010:Fe

[2097] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 64(4):NP, December 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2010:FMc

[2098] Anonymous. Front matter. Notes and Records of the Royal Society of London, 64(4):??, December 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/25802120.

Eddy:2011:PMH

[2099] Matthew D. Eddy. The prehistoric mind as a historical artefact. Notes and Records of the Royal Society of London, 65(1):1–8, March 20, 2011. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/25802136.

Eddy:2011:LRH

[2100] Matthew D. Eddy. The line of reason: Hugh Blair, spatiality and the progressive structure of language. Notes and Records of the Royal Society of London, 65(1):9–24, March 20, 2011. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/25802137.

Pettitt:2011:CMS

[2101] Paul B. Pettitt and Mark J. White. Cave men: Stone tools, Victorian science, and the ‘primitive mind’ of deep time. Notes and Records of the Royal Society of London, 65(1):25–42, March 20, 2011. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/25802138.

Gamble:2011:TRS

[2102] Clive Gamble and Theodora Moutsiou. The time revolution of 1859 and the stratification of the primeval mind. Notes and Records of the Royal Society of London, 65(1):43–63, March 20, 2011. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/25802139.

Sommer:2011:HTE

[2103] Marianne Sommer. Human tools of the European tertiary? Artefacts, brains and minds in evolutionist reasoning, 1870–1920. Notes and Records of the Royal Society of London, 65(1):65–82, March 20, 2011. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/25802140.

Kjaergaard:2011:HML

[2104] Peter C. Kjærgaard. ‘Hurrah for the missing link!’: a history of apes, ancestors and a crucial piece of evidence. Notes and Records of the Royal Society of London, 65(1):83–98, March 20, 2011. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/25802141.

Anonymous:2011:BMa

[2105] Anonymous. Back matter. Notes and Records of the Royal Society of London, 65(1):??, March 2011. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

25802142.

Anonymous:2011:Fa

[2106] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 65(1):NP, March 20, 2011. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2011:FMa

[2107] Anonymous. Front matter. Notes and Records of the Royal Society of London, 65(1):??, March 20, 2011. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

25802134.

Fox:2011:Ea

[2108] Robert Fox. Editorial. Notes and Records of the Royal Society of London, 65(2):99–101, June 20, 2011. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

23056900.

Hunter:2011:RSD

[2109] Michael Hunter. The Royal Society and the decline of magic. Notes and Records of the Royal Society of London, 65(2):103–119, June 20, 2011. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/23056901.

Walker:2011:LCR

[2110] Matthew F. Walker. The limits of collaboration: Robert Hooke, Christopher Wren and the designing of the monument to the Great Fire of London. Notes and Records of the Royal Society of London, 65(2):121–143, June 20, 2011. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/23056902.

Barrett:2011:ETB

[2111] Katy Barrett. ‘Explaining’ themselves: The Barrington papers, the Board of Longitude, and the fate of John Harrison. Notes and Records of the Royal Society of London, 65(2):145–162, June 20, 2011. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/23056903.

Thomas:2011:PWW

[2112] Sir John Meurig Thomas. Picking winners: W. H. and W. L. Bragg at the Royal Institution. Notes and Records of the Royal Society of London, 65(2):163–182, June 20, 2011. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/ 23056904. See erratum [2128].

Collins:2011:SCH

[2113] Peter Collins and Stefanie Fischer. The story of Chicheley Hall. Notes and Records of the Royal Society of London, 65(2):183–196, June 20, 2011. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/23056905.

Rees:2011:APL

[2114] M. J. Rees. Address of the President, Lord Rees of Ludlow OM KT PRS, given at the Anniversary Meeting on 30 November 2010. Notes and Records of the Royal Society of London, 65(2):197–205, June 20, 2011. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/23056906.

Knights:2011:SM

[2115] Francis Knights. The science of musicke. Notes and Records of the Royal Society of London, 65(2):207–208, June 20, 2011. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/23056907.

Anonymous:2011:BMb

[2116] Anonymous. Back matter. Notes and Records of the Royal Society of London, 65(2):??, June 2011. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

23056908.

Anonymous:2011:Fb

[2117] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 65(2):NP, June 20, 2011. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2011:FMb

[2118] Anonymous. Front matter. Notes and Records of the Royal Society of London, 65(2):??, June 20, 2011. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

23056898.

Fox:2011:Eb

[2119] Robert Fox. Editorial. Notes and Records of the Royal Society of London, 65(3):211–213, September 20, 2011. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

23056319.

Agar:2011:TS

[2120] Jon Agar. Thatcher, scientist. Notes and Records of the Royal Society of London, 65(3):215–232, September 20, 2011. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/23056320.

Despeaux:2011:FPR

[2121] Sloan Evans Despeaux. Fit to print? Referee reports on mathematics for the nineteenth-century journals of the Royal Society of London. Notes and Records of the Royal Society of London, 65(3):233–252, September 20, 2011. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/23056321.

DeArce:2011:UCB

[2122] Miguel DeArce, Nigel T. Monaghan, and Patrick N. Wyse Jackson. The uneasy correspondence between T. H. Huxley and E. P. Wright on fossil vertebrates found in Jarrow, Co. Kilkenny (1865–67). Notes and Records of the Royal Society of London, 65(3):253–271, September 20, 2011. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/23056322.

Kusukawa:2011:PKE

[2123] Sachiko Kusukawa. Picturing knowledge in the early Royal Society: the examples of Richard Waller and Henry hunt. Notes and Records of the Royal Society of London, 65(3):273–294, September 20, 2011. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/23056323.

Burns:2011:PUD

[2124] Duncan Thorburn Burns. Previously unrecorded Dutch citations and versions of some of the scientific publications of Robert Boyle FRS. Notes and Records of the Royal Society of London, 65(3):295–299, September 20, 2011. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/23056324.

Insley:2011:JWC

[2125] Jane Insley. James Watt’s cookbook chemistry. Notes and Records of the Royal Society of London, 65(3):301–308, September 20, 2011. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/23056325.

Moore:2011:PBY

[2126] Keith Moore. Portugal–Britain: 250 years of scientific exchange (1660– 1910). Notes and Records of the Royal Society of London, 65(3):309–310, September 20, 2011. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/23056326.

Roos:2011:HRB

[2127] Anna Marie Roos. Hunting Robert Boyle: Michael Hunter and Boyle’s life and letters. Notes and Records of the Royal Society of London, 65 (3):311–315, September 20, 2011. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

23056327.

Thomas:2011:EPW

[2128] Sir John Meurig Thomas. Erratum: Picking Winners: W. H. and W. L. Bragg at the Royal Institution. Notes and Records of the Royal Society of London, 65(3):317, September 20, 2011. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/23056328. See [2112].

Anonymous:2011:BMc

[2129] Anonymous. Back matter. Notes and Records of the Royal Society of London, 65(3):??, September 2011. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

23056329.

Anonymous:2011:Fc

[2130] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 65(3):NP, September 20, 2011. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2011:FMc

[2131] Anonymous. Front matter. Notes and Records of the Royal Society of London, 65(3):??, September 20, 2011. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/23056318.

Fox:2011:Ec

[2132] Robert Fox. Editorial. Notes and Records of the Royal Society of London, 65(4):321–323, December 20, 2011. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

23056301.

Collins:2011:PPC

[2133] Peter Collins. Presidential politics: The controversial election of 1945. Notes and Records of the Royal Society of London, 65(4):325–342, December 20, 2011. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/23056302.

Wilson:2011:WWG

[2134] David B. Wilson. William Whewell, Galileo, and reconceptualizing the history of science and religion. Notes and Records of the Royal Society of London, 65(4):343–358, December 20, 2011. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/23056303.

Maehle:2011:ACE

[2135] Andreas-Holger Maehle. Ambiguous cells: the emergence of the stem cell concept in the nineteenth and twentieth centuries. Notes and Records of the Royal Society of London, 65(4):359–378, December 20, 2011. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/23056304.

Tansey:2011:EEN

[2136] E. M. Tansey. The early education of a Nobel laureate: Henry Dale’s schooldays. Notes and Records of the Royal Society of London, 65(4): 379–391, December 20, 2011. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

23056305.

VanderKloot:2011:MSV

[2137] William Van der Kloot. Mirrors and smoke: A. V. Hill, his Brigands, and the science of anti-aircraft gunnery in World War I. Notes and Records of the Royal Society of London, 65(4):393–410, December 20, 2011. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/23056306.

Mills:2011:WGM

[2138] Allan Mills. William Gilbert and ‘magnetization by percussion’. Notes and Records of the Royal Society of London, 65(4):411–416, December 20, 2011. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/23056307.

Moore:2011:AYP

[2139] Keith Moore. Anniversary year portraits. Notes and Records of the Royal Society of London, 65(4):417–422, December 20, 2011. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/23056308.

Anonymous:2011:BMd

[2140] Anonymous. Back matter. Notes and Records of the Royal Society of London, 65(4):??, December 2011. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

23056309.

Anonymous:2011:Fd

[2141] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 65(4):NP, December 20, 2011. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2011:FMd

[2142] Anonymous. Front matter. Notes and Records of the Royal Society of London, 65(4):??, December 20, 2011. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/23056300.

Fox:2012:Ea

[2143] Robert Fox. Editorial. Notes and Records of the Royal Society of London, 66(1):1–2, March 20, 2012. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

41723281.

Guicciardini:2012:JWE

[2144] Niccolò Guicciardini. John Wallis as editor of Newton’s mathematical work. Notes and Records of the Royal Society of London, 66(1):3–17, March 20, 2012. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/41723282.

Roos:2012:ASR

[2145] Anna Marie Roos. The art of science: a ‘rediscovery’ of the Lister copperplates. Notes and Records of the Royal Society of London, 66(1):19–40, March 20, 2012. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/41723283.

deAlmeida:2012:PCM

[2146] Maria Antónia Pires de Almeida. The Portuguese cholera morbus epidemic of 1853–56 as seen by the press. Notes and Records of the Royal Society of London, 66(1):41–53, March 20, 2012. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/41723284.

Cahan:2012:HBS

[2147] David Cahan. Helmholtz and the British scientific elite: From force conservation to energy conservation. Notes and Records of the Royal Society of London, 66(1):55–68, March 20, 2012. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/41723285.

Finnegan:2012:JCM

[2148] Diarmid A. Finnegan. James croll, metaphysical geologist. Notes and Records of the Royal Society of London, 66(1):69–88, March 20, 2012. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/41723286.

Nurse:2012:APS

[2149] P. Nurse. Address of the President, Sir Paul Nurse, given at the Anniversary Meeting on 30 November 2011. Notes and Records of the Royal Society of London, 66(1):89–93, March 20, 2012. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/41723287.

Chambers:2012:PPF

[2150] Neil A. Chambers. Plants and the presidents: A founding vision for America. Notes and Records of the Royal Society of London, 66(1):95–99, March 20, 2012. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/41723288.

Muller-Wille:2012:TPP

[2151] Staffan Müller-Wille. Thinking plants: Peter Ayres, The aliveness of plants: the Darwins at the dawn of plant science [Pickering & Chatto, London, 2007. Pp. xiii + 227, £60.00 (hardback). ISBN 978-1-85196-970-8]. Notes and Records of the Royal Society of London, 66(1):101–103, March 20, 2012. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/41723289.

Hopkins:2012:RSP

[2152] Joanna Hopkins. The Royal Society picture library database. Notes and Records of the Royal Society of London, 66(1):105–110, March 20, 2012. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/41723290.

Anonymous:2012:BMa

[2153] Anonymous. Back matter. Notes and Records of the Royal Society of London, 66(1):??, March 2012. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

41723291.

Anonymous:2012:Fa

[2154] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 66(1):NP, March 20, 2012. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2012:FMa

[2155] Anonymous. Front matter. Notes and Records of the Royal Society of London, 66(1):??, March 20, 2012. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

41723279.

Fox:2012:Eb

[2156] Robert Fox. Editorial. Notes and Records of the Royal Society of London, 66(2):111–113, June 20, 2012. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

41723294.

Ollerton:2012:JTC

[2157] Jeff Ollerton, Gordon Chancellor, and John van Wyhe. John Tweedie and Charles Darwin in Buenos Aires. Notes and Records of the Royal Society of London, 66(2):115–124, June 20, 2012. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/41723295.

Cahan:2012:ACM

[2158] David Cahan. The awarding of the Copley Medal and the ‘discovery’ of the law of conservation of energy: Joule, Mayer and Helmholtz revisited. Notes and Records of the Royal Society of London, 66(2):125–139, June 20, 2012. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/41723296.

Katzir:2012:WKP

[2159] Shaul Katzir. Who knew piezoelectricity? Rutherford and Langevin on submarine detection and the invention of sonar. Notes and Records of the Royal Society of London, 66(2):141–157, June 20, 2012. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

URL https://royalsocietypublishing.org/journal/rsnrcontent/
66/2/141;   http://www.jstor.org/stable/41723297.

Bradley:2012:NGR

[2160] Bruce Bradley. Newton’s gift to Roger Cotes. Notes and Records of the Royal Society of London, 66(2):159–167, June 20, 2012. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/41723298.

Todd:2012:RCR

[2161] Neil Todd. The Radium Committee of the Royal Society and the fate of the substances purchased by it. Notes and Records of the Royal Society of London, 66(2):169–194, June 20, 2012. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/41723299.

Moskovits:2012:HLS

[2162] Martin Moskovits. How the localized surface plasmon became linked with surface-enhanced Raman spectroscopy. Notes and Records of the Royal Society of London, 66(2):195–203, June 20, 2012. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/41723300.

Hunter:2012:WFL

[2163] Michael Hunter. The works of Francis Lodwick. Notes and Records of the Royal Society of London, 66(2):205–206, June 20, 2012. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/41723301.

Anonymous:2012:BMb

[2164] Anonymous. Back matter. Notes and Records of the Royal Society of London, 66(2):??, June 2012. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

41723302.

Anonymous:2012:Fb

[2165] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 66(2):NP, June 20, 2012. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2012:FMb

[2166] Anonymous. Front matter. Notes and Records of the Royal Society of London, 66(2):??, June 20, 2012. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

41723292.

Fox:2012:MFS

[2167] Robert Fox. The many faces of science. Notes and Records of the Royal Society of London, 66(3):207–209, September 20, 2012. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Fox:2012:EMF

[2168] Robert Fox. Editorial: the many faces of science. Notes and Records of the Royal Society of London, 66(3):207–210, September 2012. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/41723305.

Doherty:2012:DTF

[2169] Meghan C. Doherty. Discovering the ‘true form:’ Hooke’s micrographia and the visual vocabulary of engraved portraits. Notes and Records of the Royal Society of London, 66(3):211–234, September 20, 2012. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/41723306.

Moxham:2012:ETP

[2170] Noah Moxham. Edward Tyson’s Phocaena: a case study in the institutional context of scientific publishing. Notes and Records of the Royal Society of London, 66(3):235–252, September 20, 2012. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/41723307.

Brundtland:2012:FHH

[2171] Terje Brundtland. Francis Hauksbee and his air pump. Notes and Records of the Royal Society of London, 66(3):253–272, September 20, 2012. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/41723308.

Hughes:2012:DDD

[2172] Jeff Hughes. Doing diaries: David Martin, the Royal Society and scientific London, 1947–1950. Notes and Records of the Royal Society of London, 66(3):273–294, September 20, 2012. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/41723309.

Sumner:2012:TT

[2173] James Sumner. Turing today. Notes and Records of the Royal Society of London, 66(3):295–300, September 20, 2012. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/41723310.

Carrell:2012:TM

[2174] Robin W. Carrell. Trial by media. Notes and Records of the Royal Society of London, 66(3):301–306, September 20, 2012. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/41723311.

Turner:2012:NPB

[2175] Anthony Turner. Naturalist, physician and book-maker. Notes and Records of the Royal Society of London, 66(3):307–308, September 20, 2012. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/41723312.

Anderson:2012:AC

[2176] Robert G. W. Anderson. The ascent of a chemist. Notes and Records of the Royal Society of London, 66(3):309–310, September 20, 2012. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/41723313.

Burgis:2012:ACO

[2177] Mary J. Burgis and Ian Dunn. Archives from central Office of the international biological programme, 1964–74. Notes and Records of the Royal Society of London, 66(3):311–312, September 20, 2012. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/41723314.

Anonymous:2012:BMc

[2178] Anonymous. Back matter. Notes and Records of the Royal Society of London, 66(3):??, September 2012. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

41723315.

Anonymous:2012:Fc

[2179] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 66(3):NP, September 20, 2012. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2012:FMc

[2180] Anonymous. Front matter. Notes and Records of the Royal Society of London, 66(3):??, September 20, 2012. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/41723303.

Roos:2012:NHN

[2181] Anna Marie Roos. Naturalia: the history of natural history and medicine in the seventeenth century. Notes and Records of the Royal Society of London, 66(4):313–321, December 20, 2012. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/41723318.

Lewis:2012:DJR

[2182] Gillian Lewis. The debt of John Ray and Martin Lister to Guillaume Rondelet of Montpellier. Notes and Records of the Royal Society of London, 66(4):323–339, December 20, 2012. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

41723319.

Wragge-Morley:2012:VEN

[2183] Alexander Wragge-Morley. ‘Vividness’ in English natural history and anatomy, 1650–1700. Notes and Records of the Royal Society of London, 66(4):341–356, December 20, 2012. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

41723320.

Ogilvie:2012:AIF

[2184] Brian W. Ogilvie. Attending to insects: Francis Willughby and John Ray. Notes and Records of the Royal Society of London, 66(4):357–372, December 20, 2012. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/41723321.

Sleigh:2012:JSF

[2185] Charlotte Sleigh. Jan Swammerdam’s frogs. Notes and Records of the Royal Society of London, 66(4):373–392, December 20, 2012. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/41723322.

Guerrini:2012:PBN

[2186] Anita Guerrini. Perrault, Buffon and the natural history of animals. Notes and Records of the Royal Society of London, 66(4):393–409, December 20, 2012. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/41723323.

Anonymous:2012:BMd

[2187] Anonymous. Back matter. Notes and Records of the Royal Society of London, 66(4):??, December 2012. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

41723324.

Anonymous:2012:Fd

[2188] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 66(4):NP, December 20, 2012. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2012:FMd 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/41723316.

Anonymous:2013:Ia

[2190] Anonymous. [Illustration]. Notes and Records of the Royal Society of London, 67(1):1–2, March 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

43287655.

Fox:2013:NR

[2191] Robert Fox. Notes and Records at 75. Notes and Records of the Royal Society of London, 67(1):3–5, March 20, 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/43287656.

Agar:2013:SSR

[2192] Jon Agar. ‘It’s springtime for science’: renewing China–UK scientific relations in the 1970s. Notes and Records of the Royal Society of London, 67(1):7–24, March 20, 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

43287657.

Kuehn:2013:KNR

[2193] Daniel Kuehn. Keynes, Newton and the Royal Society: the events of 1942 and 1943. Notes and Records of the Royal Society of London, 67(1):25–36, March 20, 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/43287658.

Greaves:2013:PRT

[2194] G. Neville Greaves. Poisson’s ratio over two centuries: challenging hypotheses. Notes and Records of the Royal Society of London, 67(1):37–58, March 20, 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/43287659.

Arapostathis:2013:EIA

[2195] Stathis Arapostathis. Electrical innovations, authority and consulting expertise in late Victorian Britain. Notes and Records of the Royal Society of London, 67(1):59–76, March 20, 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/43287660.

Anonymous:2013:APP

[2196] Anonymous. Address of the President, Paul Nurse, given at the Anniversary Meeting on 30 November 2012. Notes and Records of the Royal Society of London, 67(1):77–81, March 20, 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/43287661.

Quinn:2013:FBS

[2197] Terry Quinn. Fitzroy, Bailey, Southeast Iceland. Notes and Records of the Royal Society of London, 67(1):83–85, March 20, 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/43287662.

Hutchins:2013:BPE

[2198] Roger Hutchins. Battling for priority in early astrophysics. Notes and Records of the Royal Society of London, 67(1):87–88, March 20, 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/43287663.

Brock:2013:EH

[2199] William H. Brock. Exploring the hyperarctic. Notes and Records of the Royal Society of London, 67(1):89–90, March 20, 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/43287664.

Mobbs:2013:ABB

[2200] Tessa Mobbs and Robert William Unwin. Additions to the Babbage bibliography. Notes and Records of the Royal Society of London, 67(1):91–93, March 20, 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/43287665.

Anonymous:2013:BMa

[2201] Anonymous. Back matter. Notes and Records of the Royal Society of London, 67(1):??, March 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

43287666.

Anonymous:2013:Fa

[2202] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 67(1):NP, March 20, 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2013:FMa

[2203] Anonymous. Front matter. Notes and Records of the Royal Society of London, 67(1):??, March 20, 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

43287654.

Anonymous:2013:Ib

[2204] Anonymous. [Illustration]. Notes and Records of the Royal Society of London, 67(2):95–96, June 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

43287668.

Fox:2013:TI

[2205] Robert Fox. Texts and images. Notes and Records of the Royal Society of London, 67(2):97–99, June 20, 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

43287669.

Henderson:2013:FIT

[2206] Felicity Henderson. Faithful interpreters? Translation theory and practice at the early Royal Society. Notes and Records of the Royal Society of London, 67(2):101–122, June 20, 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/43287670.

Kusukawa:2013:DFR

[2207] Sachiko Kusukawa. Drawings of fossils by Robert Hooke and Richard Waller. Notes and Records of the Royal Society of London, 67(2):123–138, June 20, 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/43287671.

Launder:2013:HLC

[2208] Brian Launder. Horace Lamb and the circumstances of his appointment at Owens College. Notes and Records of the Royal Society of London, 67(2):139–158, June 20, 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

43287672.

Dudley:2013:HGW

[2209] J. M. Dudley, V. Sarano, and F. Dias. On Hokusai’s Great Wave off Kanagawa: localization, linearity and a rogue wave in sub-Antarctic waters. Notes and Records of the Royal Society of London, 67(2):159–164, June 20, 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/43287673.

Fox:2013:RFC

[2210] Robert Fox. On receiving a first copy of Notes and Records: George Sarton to A. V. Hill, 24 February 1942. Notes and Records of the Royal Society of London, 67(2):165–168, June 20, 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/43287674.

Cetto:2013:NDN

[2211] Ana María Cetto and Octavio Miramontes. Noble dreams never die. Notes and Records of the Royal Society of London, 67(2):169–170, June 20, 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/43287675.

Crease:2013:PRP

[2212] Robert P. Crease. From platinum rods to Planck’s constant. Notes and Records of the Royal Society of London, 67(2):171–172, June 20, 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/43287676.

Moore:2013:RSL

[2213] Keith Moore. The Royal Society’s lost seventeenth-century portraits. Notes and Records of the Royal Society of London, 67(2):173–178, June 20, 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/43287677.

Keates:2013:PSJ

[2214] Fiona Keates. Papers of Sir John Robert vane FRS (1927–2004). Notes and Records of the Royal Society of London, 67(2):179–183, June 20, 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/43287678.

Anonymous:2013:BMb

[2215] Anonymous. Back matter. Notes and Records of the Royal Society of London, 67(2):??, June 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

43287679.

Anonymous:2013:Fb

[2216] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 67(2):NP, June 20, 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2013:FMb

[2217] Anonymous. Front matter. Notes and Records of the Royal Society of London, 67(2):??, June 20, 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

43287667.

Anonymous:2013:Ic

[2218] Anonymous. [Illustration]. Notes and Records of the Royal Society of London, 67(3):185–186, September 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/43287681.

Hurwitz:2013:LLA

[2219] Brian Hurwitz and Marguerite W. Dupree. Learning from Lister: antisepsis, safer surgery and global health. Notes and Records of the Royal Society of London, 67(3):187–190, September 20, 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/43287682.

Howard:2013:JLH

[2220] Edward R. Howard. Joseph Lister: his contributions to early experimental physiology. Notes and Records of the Royal Society of London, 67 (3):191–198, September 20, 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

43287683.

Worboys:2013:JLP

[2221] Michael Worboys. Joseph Lister and the performance of antiseptic surgery. Notes and Records of the Royal Society of London, 67(3): 199–209, September 20, 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

43287684.

Richardson:2013:ISP

[2222] Ruth Richardson. Inflammation, suppuration, putrefaction, fermentation: Joseph Lister’s microbiology. Notes and Records of the Royal Society of London, 67(3):211–229, September 20, 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/43287685.

Carpenter:2013:LRP

[2223] Mary Wilson Carpenter. Lister’s relationship with patients: ‘a successful case’. Notes and Records of the Royal Society of London, 67(3):231–244, September 20, 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/43287686.

Schlich:2013:FIL

[2224] Thomas Schlich. Farmer to industrialist: Lister’s antisepsis and the making of modern surgery in Germany. Notes and Records of the Royal Society of London, 67(3):245–260, September 20, 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/43287687.

Dupree:2013:MSL

[2225] Marguerite Wright Dupree. From mourning to scientific legacy: commemorating Lister in London and Scotland. Notes and Records of the Royal Society of London, 67(3):261–280, September 20, 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/43287688.

Crowther:2013:LHA

[2226] M. Anne Crowther. Lister at home and abroad: a continuing legacy. Notes and Records of the Royal Society of London, 67(3):281–294, September 20, 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/43287689.

Anonymous:2013:BMc

[2227] Anonymous. Back matter. Notes and Records of the Royal Society of London, 67(3):??, September 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

43287690.

Anonymous:2013:Fc

[2228] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 67(3):NP, September 20, 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2013:FMc

[2229] Anonymous. Front matter. Notes and Records of the Royal Society of London, 67(3):??, September 20, 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/43287680.

Fox:2013:VI

[2230] Robert Fox. Volume information. Notes and Records of the Royal Society of London, 67(4):i–v, December 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/43287701.

Anonymous:2013:Id

[2231] Anonymous. [Illustration]. Notes and Records of the Royal Society of London, 67(4):295–296, December 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/43287692.

Fox:2013:NY

[2232] Robert Fox. The next 75 years. Notes and Records of the Royal Society of London, 67(4):297–300, December 20, 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/43287693.

Calver:2013:SPM

[2233] Neil Calver. Sir Peter Medawar: science, creativity and the popularization of Karl Popper. Notes and Records of the Royal Society of London, 67(4):301–314, December 20, 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

43287694.

Levitin:2013:HEW

[2234] Dmitri Levitin. Halley and the eternity of the world revisited. Notes and Records of the Royal Society of London, 67(4):315–329, December 20, 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/43287695.

Carneiro:2013:GRP

[2235] Ana Carneiro, Ana Simoes, Maria Paula Diogo, and Teresa Salomé Mota. Geology and religion in Portugal. Notes and Records of the Royal Society of London, 67(4):331–354, December 20, 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/43287696.

Papanelopoulou:2013:LPC

[2236] Faidra Papanelopoulou. Louis Paul Cailletet: The liquefaction of oxygen and the emergence of low-temperature research. Notes and Records of the Royal Society of London, 67(4):355–373, December 20, 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/43287697.

Richardson:2013:JLF

[2237] Ruth Richardson and Bryan Rhodes. Joseph Lister’s first operation. Notes and Records of the Royal Society of London, 67(4):375–385, December 20, 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/43287698.

Henderson:2013:JSM

[2238] Paul Henderson. James Sowerby: meteorites and his meteoritic sword made for the Emperor of Russia, Alexander I, in 1814. Notes and Records of the Royal Society of London, 67(4):387–401, December 20, 2013. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/43287699.

Roos:2013:HEM

[2239] Anna Marie Roos. How the early modern British economy grew. Notes and Records of the Royal Society of London, 67(4):403–404, December 20, 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/43287700.

Anonymous:2013:BMd

[2240] Anonymous. Back matter. Notes and Records of the Royal Society of London, 67(4):??, December 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

43287702.

Anonymous:2013:Fd

[2241] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 67(4):NP, December 20, 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2013:FMd

[2242] Anonymous. Front matter. Notes and Records of the Royal Society of London, 67(4):??, December 20, 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/43287691.

Anonymous:2014:Ia

[2243] Anonymous. [Illustration]. Notes and Records of the Royal Society of London, 68(1):1–2, March 2014. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

43288094.

Clark:2014:SJL

[2244] J. F. M. Clark. The science of John Lubbock. Notes and Records of the Royal Society of London, 68(1):3–6, March 20, 2014. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/43288095.

Pearn:2014:TTW

[2245] Alison Pearn. The teacher taught? What Charles Darwin owed to John Lubbock. Notes and Records of the Royal Society of London, 68(1):7–19, March 20, 2014. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/43288096. See

addendum [2263].

Owen:2014:HAJ

[2246] Janet Owen. From Down House to Avebury: John Lubbock, prehistory and human evolution through the eyes of his collection. Notes and Records of the Royal Society of London, 68(1):21–34, March 20, 2014. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/43288097.

Pettitt:2014:JLC

[2247] Paul Pettitt and Mark White. John Lubbock, caves, and the development of Middle and Upper Palaeolithic archaeology. Notes and Records of the Royal Society of London, 68(1):35–48, March 20, 2014. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/43288098.

Bridgland:2014:JLE

[2248] David R. Bridgland. John Lubbock’s early contribution to the understanding of river terraces and their importance to geography, archaeology and earth science. Notes and Records of the Royal Society of London, 68(1):49–63, March 20, 2014. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

43288099.

Clark:2014:JLS

[2249] J. F. M. Clark. John Lubbock, science, and the liberal intellectual. Notes and Records of the Royal Society of London, 68(1):65–87, March 20, 2014. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/43288100.

Anonymous:2014:BMa

[2250] Anonymous. Back matter. Notes and Records of the Royal Society of London, 68(1):??, March 2014. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

43288101.

Anonymous:2014:Fa

[2251] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 68(1):NP, March 20, 2014. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2014:FMa

[2252] Anonymous. Front matter. Notes and Records of the Royal Society of London, 68(1):??, March 20, 2014. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

43288093.

Anonymous:2014:Ib

[2253] Anonymous. [Illustrations]. Notes and Records of the Royal Society of London, 68(2):89–90, June 2014. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

43287713.

Fox:2014:HTS

[2254] Robert Fox. Historians and their sources. Notes and Records of the Royal Society of London, 68(2):91–92, June 20, 2014. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/43287714.

Hunter:2014:JRI

[2255] Michael Hunter. John Ray in Italy: lost manuscripts rediscovered. Notes and Records of the Royal Society of London, 68(2):93–109, June 20, 2014. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/43287715.

Falk:2014:SCD

[2256] Seb Falk. The scholar as craftsman: Derek de Solla Price and the reconstruction of a medieval instrument. Notes and Records of the Royal Society of London, 68(2):111–134, June 20, 2014. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/43287716.

Vieira:2014:OSL

[2257] Carla Costa Vieira. Observing the skies of Lisbon. Isaac de Sequeira Samuda, an estrangeirado in the Royal Society. Notes and Records of the Royal Society of London, 68(2):135–149, June 20, 2014. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/43287717.

Jackson:2014:JTR

[2258] Roland Jackson. John Tyndall and the Royal Medal that was never struck. Notes and Records of the Royal Society of London, 68(2):151–164, June 20, 2014. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/43287718.

Smith:2014:WDT

[2259] Charles H. Smith. Wallace, Darwin and Ternate 1858. Notes and Records of the Royal Society of London, 68(2):165–170, June 20, 2014. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/43287719.

VanderKloot:2014:LJA

[2260] William Van der Kloot. Lord Justice of Appeal John Fletcher Moulton and explosives production in World War I: ‘the mathematical mind triumphant’. Notes and Records of the Royal Society of London, 68(2):171– 186, June 20, 2014. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/43287720.

Anonymous:2014:APS

[2261] Anonymous. Address of the President, Sir Paul Nurse, given at the Anniversary Meeting on 29 November 2013. Notes and Records of the Royal Society of London, 68(2):187–192, June 20, 2014. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/43287721.

Christie:2014:APG

[2262] John R. R. Christie. Essay review: An attenuated philosophical gentleman. Robert G. W. Anderson and Jean Jones (eds.), The correspondence of Joseph Black (2 volumes). Ashgate, Farnham, 2012. Pp. xiv + 1564. £300.00 (hardback). ISBN 978-0-7546-0131-9. Notes and Records of the Royal Society of London, 68(2):193–197, June 20, 2014. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/43287722.

Pearn:2014:ATT

[2263] Alison Pearn. Addendum: The teacher taught? What Charles Darwin owed to John Lubbock. Notes and Records of the Royal Society of London, 68(2):199, June 20, 2014. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/ 43287723. See [2245].

Anonymous:2014:BMb

[2264] Anonymous. Back matter. Notes and Records of the Royal Society of London, 68(2):??, June 2014. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

43287724.

Anonymous:2014:Fb

[2265] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 68(2):NP, June 20, 2014. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2014:FMb

[2266] Anonymous. Front matter. Notes and Records of the Royal Society of London, 68(2):??, June 20, 2014. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

43287712.

Anonymous:2014:Ic

[2267] Anonymous. [Illustration]. Notes and Records of the Royal Society of London, 68(3):201–202, September 2014. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/43287726.

Fox:2014:BA

[2268] Robert Fox. Behind the appearances. Notes and Records of the Royal Society of London, 68(3):203–205, September 20, 2014. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/43287727.

Winterburn:2014:PHM

[2269] Emily Winterburn. Philomaths, Herschel, and the myth of the self-taught man. Notes and Records of the Royal Society of London, 68(3):207–225, September 20, 2014. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/43287728.

Alfonso-Goldfarb:2014:SCT

[2270] Ana Maria Alfonso-Goldfarb, Márcia Helena Mendes Ferraz, and Piyo M. Rattansi. Seventeenth-century ‘treasure’ found in Royal Society archives: The Ludus helmontii and the stone disease. Notes and Records of the Royal Society of London, 68(3):227–243, September 20, 2014. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/43287729.

Wilkins:2014:MCR

[2271] Emma Wilkins. Margaret Cavendish and the Royal Society. Notes and Records of the Royal Society of London, 68(3):245–260, September 20, 2014. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/43287730.

Ellis:2014:TBW

[2272] Markman Ellis. Thomas Birch’s ‘weekly letter’ (1741–66): correspondence and history in the mid-eighteenth-century Royal Society. Notes and Records of the Royal Society of London, 68(3):261–278, September 20, 2014. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/43287731.

Todd:2014:BHL

[2273] Neil Todd. A brief history of Lord Rutherford’s radium. Notes and Records of the Royal Society of London, 68(3):279–300, September 20, 2014. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL https://royalsocietypublishing.org/journal/
rsnrcontent/68/3/279;   http://www.jstor.org/stable/43287732.

Guise:2014:MTI

[2274] George Guise. Margaret Thatcher’s influence on British science. Notes and Records of the Royal Society of London, 68(3):301–309, September 20, 2014. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/43287733.

McManus:2014:EBP

[2275] Joanna McManus. Einstein in Britain: a portrait. Notes and Records of the Royal Society of London, 68(3):311–315, September 20, 2014. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/43287734.

Anonymous:2014:BMc

[2276] Anonymous. Back matter. Notes and Records of the Royal Society of London, 68(3):??, September 2014. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

43287735.

Anonymous:2014:Fc

[2277] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 68(3):NP, September 20, 2014. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2014:FMc

[2278] Anonymous. Front matter. Notes and Records of the Royal Society of London, 68(3):??, September 20, 2014. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/43287725.

Anonymous:2014:Id

[2279] Anonymous. [Illustration]. Notes and Records of the Royal Society of London, 68(4):317–318, December 2014. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/43288103.

Fox:2014:ND

[2280] Robert Fox. The nature of discovery. Notes and Records of the Royal Society of London, 68(4):319–321, December 20, 2014. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/43288104.

Clarke:2014:HMR

[2281] Imogen Clarke. How to manage a revolution: Isaac Newton in the early twentieth century. Notes and Records of the Royal Society of London, 68 (4):323–337, December 20, 2014. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

43288105.

Grossman:2014:JDL

[2282] Mark I. Grossman. John Dalton and the London atomists: William and Bryan Higgins, William Austin, and new Daltonian doubts about the origin of the atomic theory. Notes and Records of the Royal Society of London, 68(4):339–356, December 20, 2014. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/43288106.

Berkowitz:2014:DDP

[2283] Carin Berkowitz. Defining a discovery: priority and methodological controversy in early nineteenth-century anatomy. Notes and Records of the Royal Society of London, 68(4):357–372, December 20, 2014. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http:

//www.jstor.org/stable/43288107.

Santos:2014:TOI

[2284] Gildo Magalhães Santos. A tale of oblivion: Ida Noddack and the ‘universal abundance’ of matter. Notes and Records of the Royal Society of London, 68(4):373–389, December 20, 2014. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/43288108.

Noble:2014:LVO

[2285] Denis Noble, Dario DiFrancesco, and Diego Zancani. Leonardo da Vinci and the origin of semen. Notes and Records of the Royal Society of London, 68(4):391–402, December 20, 2014. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/43288109.

Cicuzza:2014:RAR

[2286] Daniele Cicuzza. A rediscovery of Alfred Russel Wallace’s fern collection from Borneo at the Cambridge University Herbarium. Notes and Records of the Royal Society of London, 68(4):403–412, December 20, 2014. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/43288110.

Anonymous:2014:BMd

[2287] Anonymous. Back matter. Notes and Records of the Royal Society of London, 68(4):??, December 2014. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/stable/

43288111.

Anonymous:2014:Fd

[2288] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 68(4):NP, December 20, 2014. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2014:FMd 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/43288102.

Jones:2015:WS

[2290] Claire G. Jones and Sue Hawkins. Women and science. Notes and Records of the Royal Society of London, 69(1):5–9, March 20, 2015. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Fara:2015:WSS

[2291] Patricia Fara. Women, science and suffrage in World War I. Notes and Records of the Royal Society of London, 69(1):11–24, March 20, 2015. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Hartley:2015:WCN

[2292] J. M. Hartley and E. M. Tansey. White coats and no trousers: narrating the experiences of women technicians in medical laboratories, 1930–90. Notes and Records of the Royal Society of London, 69(1):25–36, March 20, 2015. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Orr:2015:WPS

[2293] Mary Orr. Women peers in the scientific realm: Sarah Bowdich (Lee)’s expert collaborations with Georges Cuvier, 1825–33. Notes and Records of the Royal Society of London, 69(1):37–51, March 20, 2015. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Waring:2015:MFL

[2294] Sophie Waring. Margaret Fountaine: a lepidopterist remembered. Notes and Records of the Royal Society of London, 69(1):53–68, March 20, 2015. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Winterburn:2015:CHA

[2295] Emily Winterburn. Caroline Herschel: agency and self-presentation. Notes and Records of the Royal Society of London, 69(1):69–83, March 20, 2015. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Goodman:2015:HAR

[2296] Martin Goodman. The high-altitude research of Mabel Purefoy FitzGerald, 1911–13. Notes and Records of the Royal Society of London, 69(1): 85–99, March 20, 2015. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2015:Fa

[2297] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 69(1):NP, March 20, 2015. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Marsden:2015:Ea

[2298] Ben Marsden. Editorial. Notes and Records of the Royal Society of London, 69(2):105–107, June 20, 2015. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Rawson:2015:JGU

[2299] Helen C. Rawson. James Gregory, the University observatory and the early acquisition of scientific instruments at the University of St Andrews. Notes and Records of the Royal Society of London, 69(2):109–133, June 20, 2015. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Jarvis:2015:CAJ

[2300] Charles E. Jarvis and Philip H. Oswald. The collecting activities of James Cuninghame FRS on the voyage of Tuscan to China (Amoy) between 1697 and 1699. Notes and Records of the Royal Society of London, 69 (2):135–153, June 20, 2015. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Jenkins:2015:HHC

[2301] Bill Jenkins. Henry H. Cheek and transformism: new light on Charles Darwin’s Edinburgh background. Notes and Records of the Royal Society of London, 69(2):155–171, June 20, 2015. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Kenny:2015:FJB

[2302] Robert Kenny. Freud, Jung and Boas: the psychoanalytic engagement with anthropology revisited. Notes and Records of the Royal Society of London, 69(2):173–190, June 20, 2015. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Cole:2015:IPP

[2303] Rupert Cole. The importance of picking Porter: the Royal Institution, George Porter and the two cultures, 1959–64. Notes and Records of the Royal Society of London, 69(2):191–216, June 20, 2015. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2015:APS

[2304] Anonymous. Address of the President, Sir Paul Nurse, given at the Anniversary Meeting on 1 December 2014. Notes and Records of the Royal Society of London, 69(2):217–222, June 20, 2015. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2015:Fb

[2305] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 69(2):NP, June 20, 2015. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Fyfe:2015:YSP

[2306] Aileen Fyfe, Julie McDougall-Waters, and Noah Moxham. 350 years of scientific periodicals. Notes and Records of the Royal Society of London, 69(3):227–239, September 20, 2015. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Moxham:2015:FPD

[2307] N. Moxham. Fit for print: developing an institutional model of scientific periodical publishing in England, 1665–ca. 1714. Notes and Records of the Royal Society of London, 69(3):241–260, September 20, 2015. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Broman:2015:PPP

[2308] Thomas Broman. The profits and perils of publicity: Allgemeine Literatur-Zeitung, the Thurn und Taxis Post, and the periodical trade at the end of the eighteenth century. Notes and Records of the Royal Society of London, 69(3):261–276, September 20, 2015. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Fyfe:2015:JLS

[2309] Aileen Fyfe. Journals, learned societies and money: Philosophical Transactions, ca. 1750–1900. Notes and Records of the Royal Society of London, 69(3):277–299, September 20, 2015. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

leRoux:2015:PSA

[2310] Elizabeth le Roux. Publishing south African scholarship in the global academic community. Notes and Records of the Royal Society of London, 69(3):301–320, September 20, 2015. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Clarke:2015:CAO

[2311] Imogen Clarke and James Mussell. Conservative attitudes to oldestablished organs: Oliver Lodge and philosophical magazine. Notes and Records of the Royal Society of London, 69(3):321–336, September 20, 2015. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Baldwin:2015:CPR

[2312] Melinda Baldwin. Credibility, peer review, and Nature, 1945–1990. Notes and Records of the Royal Society of London, 69(3):337–352, September 20, 2015. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2015:Fc

[2313] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 69(3):NP, September 20, 2015. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Marsden:2015:Eb

[2314] Ben Marsden. Editorial. Notes and Records of the Royal Society of London, 69(4):357–359, December 20, 2015. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Penman:2015:SHD

[2315] Leigh T. I. Penman. Samuel Hartlib on the death of Descartes: a rediscovered letter to Henry More. Notes and Records of the Royal Society of London, 69(4):361–372, December 20, 2015. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Roos:2015:MWA

[2316] Anna Marie Roos and Victor D. Boantza. Mineral waters across the Channel: matter theory and natural history from Samuel Duclos’s minerallogenesis to Martin Lister’s chymical magnetism, ca. 1666–86. Notes and Records of the Royal Society of London, 69(4):373–394, December 20, 2015. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Alfonso-Goldfarb:2015:SCE

[2317] Ana Maria Alfonso-Goldfarb, Márcia Helena Mendes Ferraz, and Piyo M. Rattansi. Seventeenth-century experimenta, magisterial formulae and the ‘animal alkahest’: new documents found in Royal Society archives.

Notes and Records of the Royal Society of London, 69(4):395–418, December 20, 2015. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Cantor:2015:JTR

[2318] Geoffrey Cantor. John Tyndall’s religion: a fragment. Notes and Records of the Royal Society of London, 69(4):419–436, December 20, 2015. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Tabery:2015:RFL

[2319] James Tabery and Sahotra Sarkar. R. A. Fisher, Lancelot Hogben, and the ‘competition’ for the Chair of Social Biology at the London School of Economics in 1930: Correcting the legend. Notes and Records of the Royal Society of London, 69(4):437–446, December 20, 2015. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2015:Fd

[2320] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 69(4):NP, December 20, 2015. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Marsden:2016:Ea

[2321] Ben Marsden. Editorial. Notes and Records of the Royal Society of London, 70(1):5–7, March 20, 2016. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Joalland:2016:INL

[2322] Michael Joalland and Scott Mandelbrote. Isaac Newton learns Hebrew: Samuel Johnson’s Nova cubi Hebræi tabella. Notes and Records of the Royal Society of London, 70(1):9–21, March 20, 2016. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Lawson:2016:CMH

[2323] Ian Lawson. Crafting the microworld: how Robert Hooke constructed knowledge about small things. Notes and Records of the Royal Society of London, 70(1):23–44, March 20, 2016. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Rollinson:2016:ETI

[2324] Andrew N. Rollinson. Engineering and technology of industrial water power at Castleford Mills from the seventeenth century to the twentieth century. Notes and Records of the Royal Society of London, 70(1):45–63, March 20, 2016. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Yates:2016:JPG

[2325] Edwin Yates and Andrew Yates. Johann Peter Griess FRS (1829– 88): Victorian brewer and synthetic dye chemist. Notes and Records of the Royal Society of London, 70(1):65–81, March 20, 2016. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Calver:2016:LSU

[2326] Neil Calver and Miles Parker. The logic of scientific unity? Medawar, the Royal Society and the Rothschild controversy 1971–72. Notes and Records of the Royal Society of London, 70(1):83–100, March 20, 2016. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2016:Fa

[2327] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 70(1):NP, March 20, 2016. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Marsden:2016:Eb

[2328] Ben Marsden. Editorial. Notes and Records of the Royal Society of London, 70(2):105–106, June 20, 2016. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Riggs:2016:AAD

[2329] Christina Riggs. An autopsic art: drawings of ‘Dr Granville’s mummy’ in the Royal Society archives. Notes and Records of the Royal Society of London, 70(2):107–133, June 20, 2016. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Ungureanu:2016:YOJ

[2330] James C. Ungureanu. A Yankee at Oxford: John William Draper at the British Association for the Advancement of Science at Oxford, 30 June 1860. Notes and Records of the Royal Society of London, 70(2):135–150, June 20, 2016. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Lalli:2016:DWS

[2331] Roberto Lalli. ‘Dirty work’, but someone has to do it: Howard P. Robertson and the refereeing practices of Physical Review in the 1930s. Notes and Records of the Royal Society of London, 70(2):151–174, June 20, 2016. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Sankaran:2016:ACF

[2332] Neeraja Sankaran and Ton van Helvoort. Andrewes’s Christmas fairy tale: atypical thinking about cancer aetiology in 1935. Notes and Records of the Royal Society of London, 70(2):175–201, June 20, 2016. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2016:Fb

[2333] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 70(2):NP, June 20, 2016. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Marsden:2016:Ec

[2334] Ben Marsden. Editorial. Notes and Records of the Royal Society of London, 70(3):207–208, September 20, 2016. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Astengo:2016:RPD

[2335] Gregorio Astengo. The rediscovery of Palmyra and its dissemination in Philosophical Transactions. Notes and Records of the Royal Society of London, 70(3):209–230, September 20, 2016. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

vanBesouw:2016:ICU

[2336] Jip van Besouw. The impeccable credentials of an untrained philosopher: Willem Jacob ’s Gravesande’s career before his Leiden professorship, 1688–1717. Notes and Records of the Royal Society of London, 70 (3):231–249, September 20, 2016. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Goodman:2016:PIC

[2337] Matthew Goodman. Proving instruments credible in the early nineteenth century: The British magnetic survey and site-specific experimentation. Notes and Records of the Royal Society of London, 70(3):251–268, September 20, 2016. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Stark:2016:ARC

[2338] James F. Stark. Anti-reductionism at the confluence of philosophy and science: Arthur Koestler and the biological periphery. Notes and Records of the Royal Society of London, 70(3):269–286, September 20, 2016. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2016:APS

[2339] Anonymous. Address of the President, Sir Paul Nurse, given at the Anniversary Meeting on 30 November 2015. Notes and Records of the Royal Society of London, 70(3):287–291, September 20, 2016. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2016:Fc

[2340] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 70(3):NP, September 20, 2016. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Shuttleworth:2016:SPN

[2341] Sally Shuttleworth and Berris Charnley. Science periodicals in the nineteenth and twenty-first centuries. Notes and Records of the Royal Society of London, 70(4):297–304, December 20, 2016. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Topham:2016:SLP

[2342] Jonathan R. Topham. The scientific, the literary and the popular: Commerce and the reimagining of the scientific journal in Britain, 1813–1825. Notes and Records of the Royal Society of London, 70(4):305–324, December 20, 2016. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Corsi:2016:WDY

[2343] Pietro Corsi. What do you mean by a periodical? Forms and functions. Notes and Records of the Royal Society of London, 70(4):325–341, December 20, 2016. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Lightman:2016:PPT

[2344] Bernard Lightman. Popularizers, participation and the transformations of nineteenth-century publishing: From the 1860s to the 1880s. Notes and Records of the Royal Society of London, 70(4):343–359, December 20, 2016. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Fyfe:2016:MPA

[2345] Aileen Fyfe and Noah Moxham. Making public ahead of print: Meetings and publications at the Royal Society, 1752–1892. Notes and Records of the Royal Society of London, 70(4):361–379, December 20, 2016. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Heggie:2016:DEC

[2346] Vanessa Heggie. Distrust and expertise: Can scientific journals continue as gatekeepers? Notes and Records of the Royal Society of London, 70 (4):381–382, December 20, 2016. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Neylon:2016:CNJ

[2347] Cameron Neylon. Communities need journals. Notes and Records of the Royal Society of London, 70(4):383–385, December 20, 2016. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Higgitt:2016:SP

[2348] Rebekah Higgitt. Science publishing 2035. Notes and Records of the Royal Society of London, 70(4):387–388, December 20, 2016. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Patterson:2016:RSC

[2349] Mark Patterson. Rescaling scientific communication. Notes and Records of the Royal Society of London, 70(4):389–390, December 20, 2016. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Gupta:2016:HPC

[2350] Sunetra Gupta. How has publishing changed in the last twenty years? Notes and Records of the Royal Society of London, 70(4):391–392, December 20, 2016. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Leonelli:2016:DPD

[2351] Sabina Leonelli. The disruptive potential of data publication. Notes and Records of the Royal Society of London, 70(4):393–395, December 20, 2016. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Butler:2016:FSP

[2352] Stella Butler. The future of scientific periodicals: A librarian’s perspective. Notes and Records of the Royal Society of London, 70(4):397–398, December 20, 2016. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Madden:2016:ECS

[2353] Mary Madden. Engaging civil society with health research. Notes and Records of the Royal Society of London, 70(4):399–401, December 20, 2016. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Gadagkar:2016:PPM

[2354] Raghavendra Gadagkar. The ‘pay-to-publish’ model should be abolished. Notes and Records of the Royal Society of London, 70(4):403–404, December 20, 2016. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2016:Fd

[2355] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 70(4):NP, December 20, 2016. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Marsden:2017:E

[2356] Ben Marsden. Editorial. Notes and Records of the Royal Society of London, 71(1):5–6, March 20, 2017. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Hunter:2017:JWR

[2357] Michael Hunter. John Webster, the Royal Society and the displaying of supposed witchcraft (1677). Notes and Records of the Royal Society of London, 71(1):7–19, March 20, 2017. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Henry:2017:EBM

[2358] John Henry. Enlarging the bounds of moral philosophy: Why did Isaac Newton conclude the Opticks the way he did? Notes and Records of the Royal Society of London, 71(1):21–39, March 20, 2017. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Wardhaugh:2017:CHD

[2359] Benjamin Sutherland Wardhaugh. Charles Hutton and the ‘Dissensions’ of 1783–84: scientific networking and its failures. Notes and Records of the Royal Society of London, 71(1):41–59, March 20, 2017. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Bishop:2017:ACR

[2360] Malcolm G. H. Bishop. The Athenæum Club, the Royal Society and the reform of dentistry in nineteenth-century Britain: a research report. Notes and Records of the Royal Society of London, 71(1):61–70, March 20, 2017. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Peters:2017:OEA

[2361] Eleanor Anne Peters. Observation, experiment or autonomy in the domestic sphere? Women’s familiar science writing in Britain, 1790–1830. Notes and Records of the Royal Society of London, 71(1):71–90, March 20, 2017. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Chang:2017:WCA

[2362] Hasok Chang. Who cares about the history of science? Notes and Records of the Royal Society of London, 71(1):91–107, March 20, 2017. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2017:APS

[2363] Anonymous. Address of the President, Sir Venki Ramakrishnan, given at the Anniversary Meeting on 30 November 2016. Notes and Records of the Royal Society of London, 71(1):109–114, March 20, 2017. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2017:Fa

[2364] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 71(1):NP, March 20, 2017. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Maehle:2017:HHE

[2365] Andreas-Holger Maehle and Heather Wolffram. History of hypnotism in Europe and the significance of place. Notes and Records of the Royal Society of London, 71(2):119–123, June 20, 2017. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Hajek:2017:PTD

[2366] Kim M. Hajek. ‘A portion of truth’: Demarcating the boundaries of scientific hypnotism in late nineteenth-century France. Notes and Records of the Royal Society of London, 71(2):125–139, June 20, 2017. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Graus:2017:HLS

[2367] Andrea Graus. Hypnosis lessons by stage magnetizers: Medical and lay hypnotists in Spain. Notes and Records of the Royal Society of London, 71 (2):141–156, June 20, 2017. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Brancaccio:2017:BCB

[2368] Maria Teresa Brancaccio. Between Charcot and Bernheim: The debate on hypnotism in fin-de-siècle Italy. Notes and Records of the Royal Society of London, 71(2):157–177, June 20, 2017. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Wils:2017:TRM

[2369] Kaat Wils. From transnational to regional magnetic fevers: The making of a law on hypnotism in late nineteenth-century Belgium. Notes and Records of the Royal Society of London, 71(2):179–196, June 20, 2017. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Maehle:2017:DMG

[2370] Andreas-Holger Maehle. A dangerous method? The German discourse on hypnotic suggestion therapy around 1900. Notes and Records of the Royal Society of London, 71(2):197–211, June 20, 2017. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Wolffram:2017:CHF

[2371] Heather Wolffram. Crime and hypnosis in fin-de-siècle Germany: the Czynski case. Notes and Records of the Royal Society of London, 71(2): 213–226, June 20, 2017. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2017:Fb

[2372] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 71(2):NP, June 20, 2017. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Marsden:2017:WNT

[2373] Ben Marsden. War, nature and technoscience. Notes and Records of the Royal Society of London, 71(3):231–232, September 20, 2017. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Bektas:2017:CWT

[2374] Yakup Bektas. The Crimean War as a technological enterprise. Notes and Records of the Royal Society of London, 71(3):233–262, September 20, 2017. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Sheppard:2017:RRE

[2375] Deri Sheppard. Robert Le Rossignol, 1884–1976: Engineer of the ‘Haber’ process. Notes and Records of the Royal Society of London, 71(3):263– 296, September 20, 2017. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Clark:2017:PPU

[2376] J. F. M. Clark. Pesticides, pollution and the UK’s silent spring, 1963– 1964: Poison in the garden of England. Notes and Records of the Royal Society of London, 71(3):297–327, September 20, 2017. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Agar:2017:OB

[2377] Jon Agar. Outward bound. Notes and Records of the Royal Society of London, 71(3):329–332, September 20, 2017. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2017:Fc

[2378] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 71(3):NP, September 20, 2017. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Marsden:2017:SC

[2379] Ben Marsden. Science communications. Notes and Records of the Royal Society of London, 71(4):333–334, December 20, 2017. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Rogers:2017:YBD

[2380] Pat Rogers. The York buildings dragons: Desaguliers, Arbuthnot and attitudes towards the scientific community. Notes and Records of the Royal Society of London, 71(4):335–360, December 20, 2017. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Frame:2017:CBR

[2381] Paul Frame. Charles Blagden in revolutionary America: two unpublished letters to John Lloyd. Notes and Records of the Royal Society of London, 71(4):361–369, December 20, 2017. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

England:2017:CHW

[2382] Richard England. Censoring Huxley and Wilberforce: A new source for the meeting that the Athenaeum ‘wisely softened down’. Notes and Records of the Royal Society of London, 71(4):371–384, December 20, 2017. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Larsen:2017:ESM

[2383] Jordan Larsen. The evolving spirit: morals and mutualism in Arabella Buckley’s evolutionary epic. Notes and Records of the Royal Society of London, 71(4):385–408, December 20, 2017. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Agar:2017:WBM

[2384] Jon Agar. 2016 Wilkins–Bernal–Medawar lecture: The curious history of curiosity-driven research. Notes and Records of the Royal Society of London, 71(4):409–429, December 20, 2017. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Andersen:2017:BRR

[2385] Casper Andersen. Book review: A riverbank of science: Bernard Lightman, A companion to the history of science. Wiley Blackwell Companions to World History. Wiley Blackwell, Chichester, 2016. Pp. xvi + 601. £120.00 (hardback). ISBN 978-1-118-62077-9. Notes and Records of the Royal Society of London, 71(4):431–433, December 20, 2017. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2017:Fd

[2386] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 71(4):NP, December 20, 2017. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Marsden:2018:LLM

[2387] Ben Marsden. Living with, learning from and managing scientific failure. Notes and Records of the Royal Society of London, 72(1):5–6, March 20, 2018. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Poole:2018:SCD

[2388] William Poole. Seventeenth-Century ‘double writing’ schemes, and a 1676 letter in the phonetic script and real character of John Wilkins. Notes and Records of the Royal Society of London, 72(1):7–23, March 20, 2018. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Wright:2018:RSM

[2389] Jonathan Jeffrey Wright and Diarmid A. Finnegan. Rocks, skulls and materialism: geology and phrenology in late-Georgian Belfast. Notes and Records of the Royal Society of London, 72(1):25–55, March 20, 2018. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Choudhuri:2018:FNS

[2390] Arnab Rai Choudhuri and Rajinder Singh. The FRS nomination of Sir Prafulla C. Ray and the correspondence of N. R. Dhar. Notes and Records of the Royal Society of London, 72(1):57–73, March 20, 2018. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Park:2018:MFS

[2391] Hyung Wook Park. Managing failure: Sir Peter Brian Medawar’s transplantation research. Notes and Records of the Royal Society of London, 72(1):75–100, March 20, 2018. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2018:APS

[2392] Anonymous. Address of the President, Sir Venki Ramakrishnan, given at the Anniversary Meeting on 30 November 2017. Notes and Records of the Royal Society of London, 72(1):101–105, March 20, 2018. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Jordanova:2018:NKP

[2393] Ludmilla Jordanova. Natural knowledge in the 1660s: probing an iconic image. Notes and Records of the Royal Society of London, 72(1):107–108, March 20, 2018. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2018:Fa

[2394] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 72(1):NP, March 20, 2018. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2018:Fb

[2395] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 72(2):NP, June 20, 2018. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Stewart:2018:EUE

[2396] Larry Stewart and Kelly J. Whitmer. Expectations and utility in eighteenth-century knowledge economies: Notes and Records special issue introduction. Notes and Records of the Royal Society of London, 72 (2):111–117, June 20, 2018. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Keller:2018:DBM

[2397] Vera Keller. Deprogramming Baconianism: The meaning of desiderata in the eighteenth century. Notes and Records of the Royal Society of London, 72(2):119–137, June 20, 2018. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Whitmer:2018:PPE

[2398] Kelly J. Whitmer. Projects and pedagogical expectations: Inside P. J. Marperger’s ‘golden clover leaf’ (trifolium), 1700–1730. Notes and Records of the Royal Society of London, 72(2):139–157, June 20, 2018. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Phillips:2018:EAE

[2399] Denise Phillips. Experimentation in the agricultural Enlightenment Place, profit and norms of knowledge-making in eighteenth-century Germany. Notes and Records of the Royal Society of London, 72(2):159–172, June 20, 2018. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Hunniger:2018:WUU

[2400] Dominik Hünniger. What is a useful university? Knowledge economies and higher education in late eighteenth-century Denmark and central Europe. Notes and Records of the Royal Society of London, 72(2):173– 194, June 20, 2018. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2018:Fc

[2401] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 72(3):NP, September 20, 2018. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2018:LHH

[2402] Anonymous. Life histories, or history comes to life. Notes and Records of the Royal Society of London, 72(3):195–198, September 20, 2018. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Bud:2018:SVK

[2403] Robert Bud. Selling visions: Kantianism, cameralism and applied science realized through encyclopaediae in Germany and beyond. Notes and Records of the Royal Society of London, 72(3):199–216, September 20, 2018. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Cantor:2018:HDS

[2404] Geoffrey Cantor. Humphry Davy: a study in narcissism? Notes and Records of the Royal Society of London, 72(3):217–237, September 20, 2018. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Craik:2018:HWG

[2405] Alex D. D. Craik. The hydrostatical works of George Sinclair (c. 1630– 1696): their neglect and criticism. Notes and Records of the Royal Society of London, 72(3):239–273, September 20, 2018. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Fava-Verde:2018:VTE

[2406] Jean-François Fava-Verde. Victorian telegrams: the early development of the telegraphic despatch and its interplay with the letter post. Notes and Records of the Royal Society of London, 72(3):275–292, September 20, 2018. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Fox:2018:EO

[2407] Robert Fox. Einstein in Oxford. Notes and Records of the Royal Society of London, 72(3):293–318, September 20, 2018. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Huang:2018:SAP

[2408] Hsiang-Fu Huang. A shared arena: the private astronomy lecturing trade and its institutional counterpart in Britain, 1817–1865. Notes and Records of the Royal Society of London, 72(3):319–341, September 20, 2018. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Kolchinsky:2018:RSR

[2409] Edouard Kolchinsky, Uwe Hossfeld, and Georgy S. Levit. Russian scientists and the Royal Society of London: 350 years of scientific collaboration. Notes and Records of the Royal Society of London, 72(3):343–363, September 20, 2018. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Merchant:2018:PPS

[2410] Paul Merchant. Particular popular science: British scientists writing, speaking and broadcasting on science and religion from the 1980s. Notes and Records of the Royal Society of London, 72(3):365–381, September 20, 2018. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Raye:2018:RSS

[2411] Lee Raye. Robert Sibbald’s Scotia Illustrata (1684): A faunal baseline for Britain. Notes and Records of the Royal Society of London, 72(3):383– 405, September 20, 2018. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2018:F

[2412] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 72(4):NP, December 20, 2018. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Richter:2018:TMM

[2413] Adam D. Richter and Stephen D. Snobelen. ‘Too much for mee to speake of’: the many facets of John Wallis’s life and legacy. Notes and Records of the Royal Society of London, 72(4):407–412, December 20, 2018. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Beeley:2018:PAM

[2414] Philip Beeley. Physical arguments and moral inducements: John Wallis on questions of antiquarianism and natural philosophy. Notes and Records of the Royal Society of London, 72(4):413–430, December 20, 2018. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Ferlier:2018:JWW

[2415] Louisiane Ferlier. John Wallis’s world of ink: from manuscripts to library. Notes and Records of the Royal Society of London, 72(4):431–445, December 20, 2018. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Kaplan:2018:ADW

[2416] Abram Kaplan. Analysis and demonstration: Wallis and Newton on mathematical presentation. Notes and Records of the Royal Society of London, 72(4):447–468, December 20, 2018. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Jesseph:2018:GRP

[2417] Douglas Jesseph. Geometry, religion and politics: context and consequences of the Hobbes–Wallis dispute. Notes and Records of the Royal Society of London, 72(4):469–486, December 20, 2018. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Richter:2018:JWC

[2418] Adam D. Richter. John Wallis and the Catholics: confessional and theological antagonism in Wallis’s mathematics and philosophy. Notes and Records of the Royal Society of London, 72(4):487–503, December 20, 2018. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Rampelt:2018:PLP

[2419] Jason M. Rampelt. Polity and liturgy in the philosophy of John Wallis. Notes and Records of the Royal Society of London, 72(4):505–525, December 20, 2018. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2019:Fa

[2420] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 73(1):??, February 6, 2019. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https://

royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2018.0070.

Ramakrishnan:2019:AAF

[2421] Venki Ramakrishnan. Anniversary address Friday 30 November 2018. Notes and Records of the Royal Society of London, 73(1):5–9, February 6, 2019. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL https://royalsocietypublishing.org/doi/pdf/
10.1098/rsnr.2018.0067.

Camerota:2019:RGO

[2422] Michele Camerota, Franco Giudice, and Salvatore Ricciardo. The reappearance of Galileo’s original letter to Benedetto Castelli. Notes and Records of the Royal Society of London, 73(1):11–28, February 6, 2019. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL https://royalsocietypublishing.org/doi/pdf/10.
1098/rsnr.2018.0053.

Odlyzko:2019:NFM

[2423] Andrew Odlyzko. Newton’s financial misadventures in the South Sea Bubble. Notes and Records of the Royal Society of London, 73(1):29–59, February 6, 2019. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL https://royalsocietypublishing.org/doi/
pdf/10.1098/rsnr.2018.0018.

Wills:2019:CBD

[2424] Hannah Wills. Charles Blagden’s diary: Information management and British science in the eighteenth century. Notes and Records of the Royal Society of London, 73(1):61–81, February 6, 2019. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:

//royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2018.0016.

Tanghe:2019:OSS

[2425] Koen B. Tanghe. On The Origin of Species: The story of Darwin’s title. Notes and Records of the Royal Society of London, 73(1):83–100, February 6, 2019. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL https://royalsocietypublishing.org/doi/
pdf/10.1098/rsnr.2018.0015.

Schell:2019:NHV

[2426] Patience A. Schell. Natural history values and meanings in nineteenthcentury Chile. Notes and Records of the Royal Society of London, 73(1): 101–124, February 6, 2019. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL https://royalsocietypublishing.org/
doi/pdf/10.1098/rsnr.2017.0051.

Craik:2019:HWG

[2427] Alex D. D. Craik and Danielle Spittle. The hydrostatical works of George Sinclair (c. 1630–1696): an addendum. Notes and Records of the Royal Society of London, 73(1):125–130, February 6, 2019. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:

//royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2018.0048.

Anonymous:2019:Fb

[2428] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 73(2):??, June 8, 2019. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL https://royalsocietypublishing.org/
doi/pdf/10.1098/rsnr.2019.0006.

Naylor:2019:NCS

[2429] Simon Naylor and Simon Schaffer. Nineteenth-century survey sciences: enterprises, expeditions and exhibitions. Notes and Records of the Royal Society of London, 73(2):135–147, June 8, 2019. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:

//royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2019.0005.

Ratcliff:2019:HHS

[2430] Jessica Ratcliff. Hand-in-hand with the survey: surveying and the accumulation of knowledge capital at India House during the Napoleonic Wars. Notes and Records of the Royal Society of London, 73(2):149– 166, June 8, 2019. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL https://royalsocietypublishing.org/doi/
pdf/10.1098/rsnr.2018.0039.

Bulstrode:2019:CCC

[2431] Jenny Bulstrode. Cetacean citations and the covenant of iron. Notes and Records of the Royal Society of London, 73(2):167–185, June 8, 2019. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL https://royalsocietypublishing.org/doi/pdf/10.
1098/rsnr.2018.0033.

Goodman:2019:FDA

[2432] Matthew Goodman. Follow the data: administering science at Edward Sabine’s magnetic department, Woolwich, 1841–57. Notes and Records of the Royal Society of London, 73(2):187–202, June 8, 2019. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:

//royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2018.0036.

Naylor:2019:TSG

[2433] Simon Naylor. Thermometer screens and the geographies of uniformity in nineteenth-century meteorology. Notes and Records of the Royal Society of London, 73(2):203–221, June 8, 2019. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:

//royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2018.0037.

Wess:2019:IPG

[2434] Jane A. Wess and Charles W. J. Withers. Instrument provision and geographical science: the work of the Royal Geographical Society, 1830– ca 1930. Notes and Records of the Royal Society of London, 73(2):223– 241, June 8, 2019. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL https://royalsocietypublishing.org/doi/
pdf/10.1098/rsnr.2018.0034.

Morrison-Low:2019:GIN

[2435] A. D. Morrison-Low. Geomagnetic instruments at National Museums Scotland. Notes and Records of the Royal Society of London, 73(2): 243–257, June 8, 2019. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL https://royalsocietypublishing.org/
doi/pdf/10.1098/rsnr.2018.0035.

Connelly:2019:SSH

[2436] Charlotte Connelly and Claire Warrior. Survey stories in the history of British polar exploration: museums, objects and people. Notes and Records of the Royal Society of London, 73(2):259–274, June 8, 2019. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL https://royalsocietypublishing.org/doi/pdf/10.
1098/rsnr.2018.0038.

Anonymous:2019:Fc

[2437] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 73(3):??, September 20, 2019. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https://

royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2019.0017.

Roos:2019:OBI

[2438] Professor Anna Marie Roos. Object biographies and interdisciplinarity. Notes and Records of the Royal Society of London, 73(3):??, September 20, 2019. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL https://royalsocietypublishing.org/doi/pdf/
10.1098/rsnr.2019.0016.

Richardson:2019:LVD

[2439] Dr Philip L. Richardson. Leonardo da Vinci’s discovery of the dynamic soaring by birds in wind shear. Notes and Records of the Royal Society of London, 73(3):??, September 20, 2019. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:

//royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2018.0024.

Bennett:2019:AIS

[2440] Jim Bennett. Adventures with instruments: science and seafaring in the precarious career of Christopher Middleton. Notes and Records of the Royal Society of London, 73(3):??, September 20, 2019. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:

//royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2018.0046.

Lu:2019:HFF

[2441] Di Lu. ‘Homoeopathy flourishes in the far East’: A forgotten history of homeopathy in late nineteenth-century China. Notes and Records of the Royal Society of London, 73(3):??, September 20, 2019. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:

//royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2018.0041.

Massimi:2019:WBM

[2442] Michela Massimi. 2017 Wilkins–Bernal–Medawar Lecture: why philosophy of science matters to science. Notes and Records of the Royal Society of London, 73(3):??, September 20, 2019. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:

//royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2018.0054.

Hendriksen:2019:CLC

[2443] Marieke M. A. Hendriksen. Casting life, casting death: connections between early modern anatomical corrosive preparations and artistic materials and techniques. Notes and Records of the Royal Society of London, 73(3):??, September 20, 2019. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https://

royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2018.0068.

Owen:2019:TMA

[2444] Janet Owen. Towards a methodology for analysing nineteenth-century collecting journeys of science and empire, with Charles Darwin’s activities in Tierra del Fuego as a case study. Notes and Records of the Royal Society of London, 73(3):??, September 20, 2019. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:

//royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2018.0012.

Anonymous:2019:Fd

[2445] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 73(4):??, December 20, 2019. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https://

royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2019.0040.

Werrett:2019:IRJ

[2446] Simon Werrett. Introduction: Rethinking Joseph Banks. Notes and Records of the Royal Society of London, 73(4):425–429, December 20, 2019. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL https://royalsocietypublishing.org/doi/pdf/10.
1098/rsnr.2018.0064.

Thompson:2019:WTR

[2447] Carl Thompson. Women travellers, romantic-era science and the Banksian empire. Notes and Records of the Royal Society of London, 73(4):431–455, December 20, 2019. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https://

royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2018.0062.

Fulford:2019:RPE

[2448] Tim Fulford. The role of patronage in early nineteenth-century science, as evidenced in letters from Humphry Davy to Joseph Banks. Notes and Records of the Royal Society of London, 73(4):457–475, December 20, 2019. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL https://royalsocietypublishing.org/doi/pdf/
10.1098/rsnr.2018.0057.

Wills:2019:JBC

[2449] Hannah Wills. Joseph Banks and Charles Blagden: cultures of advancement in the scientific worlds of late eighteenth-century London and Paris. Notes and Records of the Royal Society of London, 73(4):477–497, December 20, 2019. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL https://royalsocietypublishing.org/doi/
pdf/10.1098/rsnr.2018.0060.

Rose:2019:SSS

[2450] Edwin D. Rose. From the South Seas to Soho Square: Joseph Banks’s library, collection and Kingdom of natural history. Notes and Records of the Royal Society of London, 73(4):499–526, December 20, 2019. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:

//royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2018.0059.

Anonymous:2020:Fa

[2451] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 74(1):??, March 20, 2020. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https://

royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2019.0042.

Havard:2020:PPF

[2452] Lucy J. Havard. ‘Preserve or perish’: food preservation practices in the early modern kitchen. Notes and Records of the Royal Society of London, 74(1):5–33, March 20, 2020. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL https://royalsocietypublishing.org/
doi/pdf/10.1098/rsnr.2019.0004.

Wlodarczyk:2020:PTO

[2453] Jarosław Włodarczyk. The pre-telescopic observations of the Moon in early seventeenth-century London: The case of Edward Gresham (1565– 1613). Notes and Records of the Royal Society of London, 74(1):35–53, March 20, 2020. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL https://royalsocietypublishing.org/doi/
pdf/10.1098/rsnr.2019.0009.

Mason:2020:CRS

[2454] Melvyn Mason and Robert S. White FRS. Cambridge radio sonobuoys and the seismic structure of oceanic crust. Notes and Records of the Royal Society of London, 74(1):55–72, March 20, 2020. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:

//royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2018.0061.

Robertson:2020:GCE

[2455] Haileigh Robertson. A gunpowder controversy in the early Royal Society, 1667–70. Notes and Records of the Royal Society of London, 74(1):73– 94, March 20, 2020. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL https://royalsocietypublishing.org/doi/
pdf/10.1098/rsnr.2018.0050.

Zuber:2020:RBA

[2456] Mike A. Zuber and Leigh T. I. Penman. Robert Boyle’s anonymous ‘Crosey-Crucian’ identified: The German alchemist and religious dissenter Peter Moritz. Notes and Records of the Royal Society of London, 74 (1):95–103, March 20, 2020. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL https://royalsocietypublishing.org/
doi/pdf/10.1098/rsnr.2018.0055.

Jackson:2020:EFJ

[2457] Roland Jackson. Eunice Foote, John Tyndall and a question of priority. Notes and Records of the Royal Society of London, 74(1):105–118, March 20, 2020. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL https://royalsocietypublishing.org/doi/
pdf/10.1098/rsnr.2018.0066.

Minen:2020:AMV

[2458] Francesca Minen. Ancient Mesopotamian views on human skin and body: a cultural–historical analysis of dermatological data from cuneiform sources. Notes and Records of the Royal Society of London, 74(1):119– 130, March 20, 2020. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL https://royalsocietypublishing.org/doi/
pdf/10.1098/rsnr.2018.0056.

Hogarth:2020:RWB

[2459] Alan James Hogarth and Michael Witmore. Reflexive witnessing: Boyle, the Royal Society and scientific style. Notes and Records of the Royal Society of London, 74(1):131–148, March 20, 2020. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:

//royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2018.0051.

Skouen:2020:BCN

[2460] Tina Skouen. Brian Cox and the new enlightenment. Notes and Records of the Royal Society of London, 74(1):149–178, March 20, 2020. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:

//royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2019.0002.

Anonymous:2020:Fb

[2461] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 74(2):??, June 20, 2020. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https://

royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2020.0011.

Jarvis:2020:CLJ

[2462] Charles E. Jarvis and Richard Coulton. A chronology of the life of James Petiver (ca. 1663–1718). Notes and Records of the Royal Society of London, 74(2):183–187, June 20, 2020. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https://

royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2020.0010.

Coulton:2020:WHG

[2463] Richard Coulton. ‘What he hath gather’d together shall not be lost’: remembering James Petiver. Notes and Records of the Royal Society of London, 74(2):189–211, June 20, 2020. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https://

royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2020.0012.

Maydom:2020:JPA

[2464] Katrina Elizabeth Maydom. James Petiver’s apothecary practice and the consumption of American drugs in early modern London. Notes and Records of the Royal Society of London, 74(2):213–238, June 20, 2020. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL https://royalsocietypublishing.org/doi/pdf/10.
1098/rsnr.2019.0015.

Marples:2020:JPJ

[2465] Alice Marples. James Petiver’s ‘joynt-stock’: middling agency in urban collecting networks. Notes and Records of the Royal Society of London, 74 (2):239–258, June 20, 2020. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL https://royalsocietypublishing.org/
doi/pdf/10.1098/rsnr.2019.0013.

Murphy:2020:JPK

[2466] Kathleen Susan Murphy. James Petiver’s ‘kind friends’ and ‘curious persons’ in the Atlantic World: commerce, colonialism and collecting. Notes and Records of the Royal Society of London, 74(2):259–274, June 20, 2020. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL https://royalsocietypublishing.org/doi/pdf/
10.1098/rsnr.2019.0011.

Vane-Wright:2020:JPP

[2467] Richard I. Vane-Wright. James Petiver’s 1717 Papilionum Britanniae: an analysis of the first comprehensive account of British butterflies (Lepidoptera: Papilionoidea). Notes and Records of the Royal Society of London, 74(2):275–302, June 20, 2020. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:

//royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2019.0014.

Jarvis:2020:MCG

[2468] Charles E. Jarvis. ‘The most common grass, rush, moss, fern, thistles, thorns or vilest weeds you can find’: James Petiver’s plants. Notes and Records of the Royal Society of London, 74(2):303–328, June 20, 2020. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL https://royalsocietypublishing.org/doi/pdf/10.
1098/rsnr.2019.0012.

Jarvis:2020:JPC

[2469] Charles E. Jarvis. James Petiver (c. 1663–1718): a concise bibliography. Notes and Records of the Royal Society of London, 74(2):329–333, June 20, 2020. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL https://royalsocietypublishing.org/doi/pdf/
10.1098/rsnr.2020.0013.

Anonymous:2020:Fc

[2470] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 74(3):??, September 20, 2020. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https://

royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2020.0009.

Roos:2020:PWB

[2471] Professor Anna Marie Roos. Plague wind and blue skies. Notes and Records of the Royal Society of London, 74(3):339–343, September 20, 2020. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL https://royalsocietypublishing.org/doi/pdf/
10.1098/rsnr.2020.0033.

Jackson:2020:WBM

[2472] Mark Jackson. 2019 Wilkins–Bernal–Medawar lecture: Life begins at 40: the demographic and cultural roots of the midlife crisis. Notes and Records of the Royal Society of London, 74(3):345–364, September 20, 2020. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL https://royalsocietypublishing.org/doi/pdf/
10.1098/rsnr.2020.0008.

SullivanDeOliveira:2020:MMS

[2473] Patrick Luiz Sullivan De Oliveira. Martyrs made in the sky: the Zénith balloon tragedy and the construction of the French Third Republic’s first scientific heroes. Notes and Records of the Royal Society of London, 74(3):365–386, September 20, 2020. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:

//royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2019.0022.

Schilt:2020:MME

[2474] Cornelis J. Schilt. Of manuscripts and men: the editorial history of Isaac Newton’s Chronology and Observations. Notes and Records of the Royal Society of London, 74(3):387–408, September 20, 2020. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:

//royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2018.0069.

Gomori:2020:WMS

[2475] George Gömöri and Stephen D. Snobelen. What he may seem to the world: Isaac Newton’s autograph book epigrams. Notes and Records of the Royal Society of London, 74(3):409–452, September 20, 2020. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL https://royalsocietypublishing.org/doi/pdf/10.
1098/rsnr.2020.0006.

Morris:2020:EJT

[2476] Andrew M. A. Morris. Evaluating John Theophilus Desaguliers’ Newtonianism: the case of waterwheel knowledge in A course of experimental philosophy. Notes and Records of the Royal Society of London, 74(3):453– 477, September 20, 2020. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL https://royalsocietypublishing.org/
doi/pdf/10.1098/rsnr.2019.0023.

Ducheyne:2020:AVF

[2477] Steffen Ducheyne. Adriaen Verwer (1654/5–1717) and the first edition of Isaac Newton’s Principia in the Dutch Republic. Notes and Records of the Royal Society of London, 74(3):479–505, September 20, 2020. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:

//royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2019.0008.

Meyns:2020:DRS

[2478] Chris Meyns. ‘Data’ in the Royal Society’s Philosophical Transactions, 1665–1886. Notes and Records of the Royal Society of London, 74(3):507– 528, September 20, 2020. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL https://royalsocietypublishing.org/
doi/pdf/10.1098/rsnr.2019.0024.

Anonymous:2020:F

[2479] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 74(4):??, December 20, 2020. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https://

royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2020.0044.

Sankaran:2020:IDH

[2480] Neeraja Sankaran. Introduction: Diversifying the historiography of bacteriophages. Notes and Records of the Royal Society of London, 74(4):533–538, December 20, 2020. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https://

royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2020.0037.

Kirchhelle:2020:FTR

[2481] Claas Kirchhelle. The forgotten typers: The rise and fall of Weimar bacteriophage-typing (1921–1935). Notes and Records of the Royal Society of London, 74(4):539–565, December 20, 2020. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:

//royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2019.0020.

Lackovic:2020:VSF

[2482] Zdravko Lacković and Karlo Toljan. Vladimir Sertić: forgotten pioneer of virology and bacteriophage therapy. Notes and Records of the Royal Society of London, 74(4):567–578, December 20, 2020. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:

//royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2019.0010.

Myelnikov:2020:CFL

[2483] Dmitriy Myelnikov. Creature features: The lively narratives of bacteriophages in Soviet biology and medicine. Notes and Records of the Royal Society of London, 74(4):579–597, December 20, 2020. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:

//royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2019.0035.

Kostyrka:2020:OLE

[2484] Gladys Kostyrka and Neeraja Sankaran. From obstacle to lynchpin: the evolution of the role of bacteriophage lysogeny in defining and understanding viruses. Notes and Records of the Royal Society of London, 74(4):599–623, December 20, 2020. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:

//royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2019.0033.

Kochhar:2020:VRH

[2485] Rijul Kochhar. The virus in the rivers: histories and antibiotic afterlives of the bacteriophage at the sangam in Allahabad. Notes and Records of the Royal Society of London, 74(4):625–651, December 20, 2020. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:

//royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2020.0019.

Summers:2020:APH

[2486] William C. Summers. Afterword: Phage, history and historiography. Notes and Records of the Royal Society of London, 74(4):653–656, December 20, 2020. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL https://royalsocietypublishing.org/doi/
pdf/10.1098/rsnr.2020.0038.

Anonymous:2021:Fa

[2487] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 75(1):??, March 20, 2021. CODEN NOREAY.

ISSN 0035-9149 (print), 1743-0178 (electronic). URL https://

royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2020.0061.

Matin:2021:WMS

[2488] Moujan Matin, Mohammad Gholamnejad, and Ali Nemati Abkenar. ‘we must send you a sample’ — a Persian–European dialogue: insights into late nineteenth-century ceramic technology based on chemical analysis of tiles from the Ettehadieh house complex, Tehran, Iran. Notes and Records of the Royal Society of London, 75(1):5–37, March 20, 2021. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL https://royalsocietypublishing.org/doi/pdf/
10.1098/rsnr.2020.0015.

Bellis:2021:PWT

[2489] Richard T. Bellis. ‘As to the plan of this work … we think Dr. Baillie has done wrong’: changing the study of disease through epistemic genre in Georgian Britain. Notes and Records of the Royal Society of London, 75 (1):39–58, March 20, 2021. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL https://royalsocietypublishing.org/
doi/pdf/10.1098/rsnr.2019.0036.

Giovannetti-Singh:2021:GNW

[2490] Gianamar Giovannetti-Singh. Galenizing the New World: Joseph-François Lafitau’s ‘Galenization’ of Canadian ginseng, ca 1716–1724. Notes and Records of the Royal Society of London, 75(1):59–72, March 20, 2021. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL https://royalsocietypublishing.org/doi/pdf/
10.1098/rsnr.2019.0037.

Packham:2021:LCH

[2491] Ellen Packham. The limits of chemistry: how William Gregory contested the boundaries of ‘established science’, 1820–1850. Notes and Records of the Royal Society of London, 75(1):73–89, March 20, 2021. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:

//royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2019.0025.

Zhang:2021:QTP

[2492] Jiuchen Zhang and Jingfei Zhang. The Qinghai–Tibet plateau: pilot site for international collaboration in geoscience during China’s early period of reform and opening-up. Notes and Records of the Royal Society of London, 75(1):91–109, March 20, 2021. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:

//royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2019.0031.

Muller:2021:LWR

[2493] Wolfgang H. Müller and Richard J. Howarth. The life and work of Robert J. Adcock. Part I: the figure of the Earth. Notes and Records of the Royal Society of London, 75(1):111–140, March 20, 2021. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:

//royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2019.0027.

Howarth:2021:LWR

[2494] Richard J. Howarth and Wolfgang H. Müller. The life and work of Robert J. Adcock. Part II: biography and establishment of authorship. Notes and Records of the Royal Society of London, 75(1):141–157, March 20, 2021. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL https://royalsocietypublishing.org/doi/
pdf/10.1098/rsnr.2019.0028.

Smith:2021:LEM

[2495] Meg Weston Smith. Away from the limelight: Einstein and Milne. Notes and Records of the Royal Society of London, 75(1):159–165, March 20, 2021. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL https://royalsocietypublishing.org/doi/pdf/
10.1098/rsnr.2019.0047.

Hunting:2021:EJL

[2496] Penelope Hunting. The Endeavour journal of Lieutenant Zachary Hicks 1768–1771. Notes and Records of the Royal Society of London, 75(1): 167–175, March 20, 2021. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL https://royalsocietypublishing.org/
doi/pdf/10.1098/rsnr.2020.0021.

Anonymous:2021:Fb

[2497] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 75(2):NP, June 20, 2021. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https://

royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2021.0014.

Guerrini:2021:ISH

[2498] Anita Guerrini and Georgina M. Montgomery. Introduction: Sustainability and the history of scientific environments. Notes and Records of the Royal Society of London, 75(2):181–187, June 20, 2021. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:

//royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2021.0011.

Brady:2021:WZB

[2499] Lisa M. Brady. From war zone to biosphere reserve: the Korean DMZ as a scientific landscape. Notes and Records of the Royal Society of London, 75 (2):189–205, June 20, 2021. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL https://royalsocietypublishing.org/
doi/pdf/10.1098/rsnr.2020.0023.

Simpson:2021:AMA

[2500] Emily Simpson. Ant mazes and astronomy: Harlow Shapley’s entomological experiments at Mount Wilson Observatory and Pasadena, California. Notes and Records of the Royal Society of London, 75(2):207– 221, June 20, 2021. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL https://royalsocietypublishing.org/doi/
pdf/10.1098/rsnr.2020.0024.

Montgomery:2021:NHC

[2501] Georgina M. Montgomery. ‘Never so at home’: Charles Elton and the Woods of Wytham. Notes and Records of the Royal Society of London, 75 (2):223–238, June 20, 2021. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL https://royalsocietypublishing.org/
doi/pdf/10.1098/rsnr.2021.0007.

Brown:2021:ESL

[2502] Kevin C. Brown. ‘An exceedingly simple, little ecosystem’: Devils Hole, endangered species conservation, and scientific environments. Notes and Records of the Royal Society of London, 75(2):239–257, June 20, 2021. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL https://royalsocietypublishing.org/doi/pdf/
10.1098/rsnr.2020.0025.

Guerrini:2021:WGL

[2503] Anita Guerrini. The Wild Garden: landscaping southern California in the early twentieth century. Notes and Records of the Royal Society of London, 75(2):259–276, June 20, 2021. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:

//royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2021.0002.

Anonymous:2021:Fc

[2504] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 75(3):277–278, September ??, 2021. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https://

royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2021.0028.

Roos:2021:ELD

[2505] Anna Marie Roos. Editorial: Life and death in the archive: The Royal Society, Martin Folkes (1690–1754), and Coram’s voices through time. Notes and Records of the Royal Society of London, 75(3):281– 284, September ??, 2021. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL https://royalsocietypublishing.org/
doi/pdf/10.1098/rsnr.2021.0017.

Fleetwood:2021:SWL

[2506] Lachlan Fleetwood. Science and war at the limit of empire: William Griffith with the Army of the Indus. Notes and Records of the Royal Society of London, 75(3):285–310, September ??, 2021. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:

//royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2019.0048.

vanTrijp:2021:FFO

[2507] Didi van Trijp. Fresh fish: Observation up close in Late Seventeenth-Century England. Notes and Records of the Royal Society of London, 75(3):311–332, September ??, 2021. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https://

royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2019.0051.

Gatzia:2021:DSI

[2508] Dimitria Electra Gatzia and Rex D. Ramsier. Dimensionality, symmetry and the Inverse Square Law. Notes and Records of the Royal Society of London, 75(3):333–348, September ??, 2021. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:

//royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2019.0044.

Diniz:2021:ESC

[2509] Marcio A. Diniz and David R. Bellhouse. The early scientific connections of Richard Price. Notes and Records of the Royal Society of London, 75(3):349–358, September ??, 2021. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https://

royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2019.0052.

An:2021:DHS

[2510] Sooyoung An. Daniel Hanbury’s study of Chinese MATERIA MED-ICA: A British network of letters and specimens during the Nineteenth Century. Notes and Records of the Royal Society of London, 75(3):359– 380, September ??, 2021. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL https://royalsocietypublishing.org/
doi/pdf/10.1098/rsnr.2019.0039.

Turner:2021:JCD

[2511] Anthony Turner and John Davis. John Collins and the double horizontal dial. Notes and Records of the Royal Society of London, 75(3):381–396, September ??, 2021. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL https://royalsocietypublishing.org/doi/
pdf/10.1098/rsnr.2019.0045.

Polloni:2021:ERG

[2512] Nicola Polloni. Early Robert Grosseteste on matter. Notes and Records of the Royal Society of London, 75(3):397–414, September ??, 2021. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL https://royalsocietypublishing.org/doi/pdf/10.
1098/rsnr.2020.0017.

Hong:2021:AHA

[2513] Jiang Hong. Angel in the house, angel in the scientific empire: Women and colonial botany during the Eighteenth and Nineteenth Centuries. Notes and Records of the Royal Society of London, 75(3):415–438, September ??, 2021. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL https://royalsocietypublishing.org/doi/
pdf/10.1098/rsnr.2020.0046.

Orrman-Rossiter:2021:PID

[2514] Kevin Orrman-Rossiter. Places of ‘invention and discovery’ and the Nobel prize in physics. Notes and Records of the Royal Society of London, 75(3):439–460, September ??, 2021. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https://

royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2020.0014.

Shi:2021:GKB

[2515] John Z. Shi. George Keith Batchelor’s interaction with Chinese fluid dynamicists and inspirational influence: a historical perspective. Notes and Records of the Royal Society of London, 75(3):461–502, September ??, 2021. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL https://royalsocietypublishing.org/doi/pdf/
10.1098/rsnr.2019.0034.

Anonymous:2021:F

[2516] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 75(4):505–506, December ??, 2021. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https://

royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2021.0077.

Dickenson:2021:IUT

[2517] Victoria Dickenson. Introduction: Undescrib’d: Taylor White (1701– 1772) and his collections. Notes and Records of the Royal Society of London, 75(4):507–513, December ??, 2021. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https://

royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2021.0062.

Dickenson:2021:OCT

[2518] Victoria Dickenson. ‘Obliging and curious’: Taylor White (1701– 1772) and his remarkable collections. Notes and Records of the Royal Society of London, 75(4):515–541, December ??, 2021. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:

//royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2020.0073.

Stantina:2021:TWP

[2519] Céline M. Stantina. Taylor White’s ‘paper museum’ (1725–1772): understanding the scientific work of an unpublished naturalist. Notes and Records of the Royal Society of London, 75(4):543–557, December ??, 2021. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL https://royalsocietypublishing.org/doi/pdf/
10.1098/rsnr.2020.0069.

Greenfield:2021:PNT

[2520] Emilienne Greenfield. The practice of note-taking in Taylor White’s natural history collection. Notes and Records of the Royal Society of London, 75(4):559–579, December ??, 2021. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https://

royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2020.0067.

Javidi:2021:OIT

[2521] Vida Javidi and Robert Montgomerie. Ornithological insights from Taylor White’s birds. Notes and Records of the Royal Society of London, 75(4):581–598, December ??, 2021. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https://

royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2020.0066.

Dickenson:2021:TWP

[2522] Victoria Dickenson and Jennifer Garland. Taylor White’s ‘paper museum’. Notes and Records of the Royal Society of London, 75(4):599– 626, December ??, 2021. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL https://royalsocietypublishing.org/
doi/pdf/10.1098/rsnr.2021.0020.

Zinger:2021:JPI

[2523] Emily Zinger. ‘Just put it online’: the Taylor White project as a digitization case study. Notes and Records of the Royal Society of London, 75(4):627–642, December ??, 2021. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https://

royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2020.0068.

Williams:2021:FCT

[2524] Lauren Williams. Fruitful collaborations: the Taylor White project in the Blacker Wood Natural History Collection. Notes and Records of the Royal Society of London, 75(4):643–656, December ??, 2021. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:

//royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2021.0012.

Anonymous:2022:Fa

[2525] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 76(1):NP, March 20, 2022. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https://

royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2021.0076.

Amery:2022:DSA

[2526] Fiona Amery. The disputed sound of the aurora borealis: sensing liminal noise during the First and Second International Polar Years, 1882–3 and 1932–3. Notes and Records of the Royal Society of London, 76(1):5–26, March 20, 2022. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL https://royalsocietypublishing.org/doi/
pdf/10.1098/rsnr.2021.0031.

vanBlanken:2022:EOA

[2527] Kerrewin van Blanken. Earthquake observations in the age before Lisbon: eyewitness observation and earthquake philosophy in the Royal Society, 1665–1755. Notes and Records of the Royal Society of London, 76(1): 27–48, March 20, 2022. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL https://royalsocietypublishing.org/
doi/pdf/10.1098/rsnr.2020.0005.

Gillin:2022:CSM

[2528] Edward J. Gillin. Cornish science, mine experiments and Robert Were Fox’s Penjerrick letters. Notes and Records of the Royal Society of London, 76(1):49–66, March 20, 2022. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https://

royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2020.0030.

Ortiz:2022:UEF

[2529] Joseph D. Ortiz and Roland Jackson. Understanding Eunice Foote’s 1856 experiments: heat absorption by atmospheric gases. Notes and Records of the Royal Society of London, 76(1):67–84, March 20, 2022. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:

//royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2020.0031.

Moura:2022:FFT

[2530] Breno Arsioli Moura. The first French translation of Book II of Newton’s Opticks: omissions, abridgements and the quest for authorship. Notes and Records of the Royal Society of London, 76(1):85–106, March 20, 2022. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL https://royalsocietypublishing.org/doi/pdf/
10.1098/rsnr.2020.0029.

Mahnke:2022:LMD

[2531] Heinz-Eberhard Mahnke. Lise Meitner, β-decay and non-radiative electromagnetic transitions. Notes and Records of the Royal Society of London, 76(1):107–116, March 20, 2022. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https://

royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2020.0036.

deAndradeMartins:2022:JMP

[2532] Roberto de Andrade Martins. Joule’s 1840 manuscript on the production of heat by voltaic electricity. Notes and Records of the Royal Society of London, 76(1):117–154, March 20, 2022. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:

//royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2020.0027.

Gilmore:2022:ERV

[2533] Gerard Gilmore and Gudrun Tausch-Pebody. The 1919 eclipse results that verified general relativity and their later detractors: a story retold. Notes and Records of the Royal Society of London, 76(1):155–180, March 20, 2022. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL https://royalsocietypublishing.org/doi/
pdf/10.1098/rsnr.2020.0040.

Huang:2022:GSC

[2534] Hsiang-Fu Huang. From Grub Street to the Colony: George William Francis and an early Victorian scientific career. Notes and Records of the Royal Society of London, 76(1):181–208, March 20, 2022. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:

//royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2020.0002.

Holmes:2022:HFP

[2535] Matthew Holmes. Houseflies and fungi: the promise of an early twentieth-century biotechnology. Notes and Records of the Royal Society of London, 76(1):209–224, March 20, 2022. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https://

royalsocietypublishing.org/doi/pdf/10.1098/rsnr.2019.0043.

Anonymous:2022:F

[2536] Anonymous. Frontispiece. Notes and Records of the Royal Society of London, 76(2):NP, June ??, 2022. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https://

royalsocietypublishing.org/doi/epdf/10.1098/rsnr.2022.0003.

Blewett:2022:ICE

[2537] Morwenna Blewett and Lucy Wrapson. Introduction: Cabinet, elaboratory, gallery 1500–1800. the preservation of art and material culture in Europe. Notes and Records of the Royal Society of London, 76(2): 229–235, June ??, 2022. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsnr.2021.0079.

Beltrami:2022:DDP

[2538] Costanza Beltrami. Defence by demolition? Preserving and relocating the cloister of Segovia cathedral. Notes and Records of the Royal Society of London, 76(2):237–252, June ??, 2022. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsnr.2021.0045.

Stoddart:2022:TSH

[2539] Olivia Stoddart and Gerry Alabone. Thomas Sackville’s Hall of Fame: displaced, reinvented and preserved at Knole. Notes and Records of the Royal Society of London, 76(2):253–272, June ??, 2022. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsnr.2021.0044.

Cardinali:2022:PEL

[2540] Giacomo Cardinali. The problem of ‘extinguished letters’ and the use of chemical reagents on manuscripts (1551–1553). Notes and Records of the Royal Society of London, 76(2):273–286, June ??, 2022. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsnr.2021.0027.

Limb:2022:LSG

[2541] Alice Limb. From the life school to the gallery wall, via the portfolio: the collection, treatment, and display of oil sketches on paper produced in the contexts of the Carracci school. Notes and Records of the Royal Society of London, 76(2):287–302, June ??, 2022. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsnr.2021.0043.

Burgess:2022:TTN

[2542] W. G. Burgess. Time’s teeth: narratives of preservation in the eighteenth-century Cotton Library. Notes and Records of the Royal Society of London, 76(2):303–315, June ??, 2022. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https://

royalsocietypublishing.org/doi/epdf/10.1098/rsnr.2021.0041.

Werrett:2022:PND

[2543] Simon Werrett. Preserving nature: domestic thrift and techniques of conservation in early modern England. Notes and Records of the Royal Society of London, 76(2):317–329, June ??, 2022. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsnr.2021.0029.

Anonymous:2022:CTA

[2544] Anonymous. Cemetery tower at Atena. Notes and Records of the Royal Society of London, 76(3):NP, September ??, 2022. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsnr.2022.0024.

Roos:2022:PFC

[2545] Anna Marie Roos. Publish and flourish, or the collective wisdom of peer review. Notes and Records of the Royal Society of London, 76(3):335– 336, September ??, 2022. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsnr.2022.0022.

Griffiths:2022:SHN

[2546] Thomas A. Griffiths. ‘Shout hurrah!’ New thoughts on the origin and meaning of the bat species name Ia io, created in 1902 by Oldfield Thomas FRS. Notes and Records of the Royal Society of London, 76(3):337–350, September ??, 2022. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https://

royalsocietypublishing.org/doi/epdf/10.1098/rsnr.2022.0006.

Garcia-Pereda:2022:BBG

[2547] Ignacio García-Pereda, Ana Duarte Rodrigues, and Francisco Manuel Parejo-Moruno. The Boutelou brothers: From gardening to agronomic practices, education, and travels in Spain at the turn of the Nineteenth Century. Notes and Records of the Royal Society of London, 76(3):351– 369, September ??, 2022. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsnr.2020.0053.

Wlodarczyk:2022:GLE

[2548] Jarosław Włodarczyk. ‘Out of a greate laborinth of errors’: Lunar astronomy in London before Kepler. Notes and Records of the Royal Society of London, 76(3):371–386, September ??, 2022. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsnr.2020.0058.

Tunc:2022:TTK

[2549] Tanfer Emin Tunc and Gokhan Tunc. Transferring technical knowledge to Turkey: American engineers, scientific experts, and the Erzincan earthquake of 1939. Notes and Records of the Royal Society of London, 76(3):387–406, September ??, 2022. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https://

royalsocietypublishing.org/doi/epdf/10.1098/rsnr.2020.0054.

Jons:2022:UBG

[2550] Heike Jöns, Michael Heffernan, and Dean W. Bond. Unity in bronze: German universities and the 250th anniversary of the Royal Society. Notes and Records of the Royal Society of London, 76(3):407–443, September ??, 2022. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsnr.2020.0051.

Partridge:2022:MSD

[2551] Derek Partridge. Mind the step: did Hooker’s judgement clinch Darwin’s disenchantment? Notes and Records of the Royal Society of London, 76(3):445–461, September ??, 2022. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https://

royalsocietypublishing.org/doi/epdf/10.1098/rsnr.2020.0055.

Lopes:2022:SFU

[2552] Quintino Lopes and Elisabete J. Santos Pereira. Science funding under an authoritarian regime: Portugal’s National Education Board and the European ‘academic landscape’ in the interwar period. Notes and Records of the Royal Society of London, 76(3):463–483, September ??, 2022. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/10.
1098/rsnr.2021.0037.

Bush:2022:AFM

[2553] Martin Bush. Again with feeling: modes of visual representation of popular astronomy in the mid-nineteenth century. Notes and Records of the Royal Society of London, 76(3):485–506, September ??, 2022. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsnr.2021.0040.

Marples:2022:SMI

[2554] Alice Marples. The science of money: Isaac Newton’s mastering of the Mint. Notes and Records of the Royal Society of London, 76(3):507– 525, September ??, 2022. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsnr.2021.0033.

Lomas:2022:MIG

[2555] Benjamin Lomas. ‘A man of intrigue’: Giles Rawlins, 1631?–1662. Notes and Records of the Royal Society of London, 76(3):527–540, September ??, 2022. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/10.
1098/rsnr.2021.0065.

Guicciardini:2022:DGM

[2556] Niccolò Guicciardini. David Gregory’s manuscript ‘Isaaci Neutoni Methodus fluxionum’ (1694): a study on the early publication of Newton’s discoveries on calculus. Notes and Records of the Royal Society of London, 76(3):541–564, September ??, 2022. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsnr.2021.0013.

Gillin:2022:IES

[2557] Edward J. Gillin. The instruments of expeditionary science and the reworking of nineteenth-century magnetic experiment. Notes and Records of the Royal Society of London, 76(3):565–592, September ??, 2022. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/10.
1098/rsnr.2022.0002.

Debru:2022:EAG

[2558] Claude Debru, Wolfgang U. Eckart, Heiner Fangerau, and Robert Fox. European academies and the Great War: an inter-academy initiative, 2014–2021. Notes and Records of the Royal Society of London, 76(3):593– 602, September ??, 2022. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsnr.2021.0047.

Ohnesorge:2022:TFJ

[2559] Miguel Ohnesorge. Theodolites at 20 000 feet: justifying precision measurement during the trigonometrical survey of Kashmir, 1855–1865. Notes and Records of the Royal Society of London, 76(3):603–618, September ??, 2022. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsnr.2020.0052.

Hall:2022:FBH

[2560] Kersten Hall. Florence Bell — the ‘housewife’ with X-ray vision. Notes and Records of the Royal Society of London, 76(3):619–631, September ??, 2022. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/10.
1098/rsnr.2020.0064.

Limeira-DaSilva:2022:IAR

[2561] Victor Rafael Limeira-DaSilva. The itinerary of Alfred Russel Wallace’s Amazonian journey (1848–1852): a source for researchers and readers. Notes and Records of the Royal Society of London, 76(3):633–652, September ??, 2022. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsnr.2020.0047.

Anonymous:2022:FD

[2562] Anonymous. Frontispiece for December 2022. Notes and Records of the Royal Society of London, 76(4):NP, December ??, 2022. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsnr.2022.0023.

Pearl:2022:ITA

[2563] Sharrona Pearl. Introduction: theorizing and applying the meaningfully anecdotal patient in neurodiversity research. Notes and Records of the Royal Society of London, 76(4):657–661, December ??, 2022. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsnr.2021.0083.

Casper:2022:APB

[2564] Stephen T. Casper. The anecdotal patient: brain injury and the magnitude of harm. Notes and Records of the Royal Society of London, 76(4):663–682, December ??, 2022. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https://

royalsocietypublishing.org/doi/epdf/10.1098/rsnr.2021.0021.

Silverman:2022:HRR

[2565] Chloe Silverman. How to read ‘Reading the Mind in the Eyes’. Notes and Records of the Royal Society of London, 76(4):683–697, December ??, 2022. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/10.
1098/rsnr.2021.0058.

Cole:2022:IWL

[2566] Jonathan Cole. Insights from those who live with impairments of facial mobility. Notes and Records of the Royal Society of London, 76(4):699– 712, December ??, 2022. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsnr.2021.0018.

Sellers:2022:WY

[2567] Heather Sellers. Who are you? Notes and Records of the Royal Society of London, 76(4):713–720, December ??, 2022. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsnr.2021.0024.

Edkins:2022:SK

[2568] Jenny Edkins. Supposed to know. Notes and Records of the Royal Society of London, 76(4):721–722, December ??, 2022. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsnr.2021.0022.

Roos:2023:FPL

[2569] Anna Marie Roos. Flat places: 1 Lincolnshire and science. Notes and Records of the Royal Society of London, 77(1):1–3, March ??, 2023. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/10.
1098/rsnr.2022.0045.

Smith:2023:SSC

[2570] Nathan Smith. Of stumps and stipes: comparisons between the cultures and identities of Yorkshire cricket and mycology at the turn of the twentieth century. Notes and Records of the Royal Society of London, 77 (1):5–18, March ??, 2023. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsnr.2021.0036.

Rogers:2023:LBL

[2571] Ethan S. Rogers and Stephanie L. Canington. Lemurs before lemur: depictions of captive lemurs prior to Linnaeus. Notes and Records of the Royal Society of London, 77(1):19–48, March ??, 2023. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsnr.2021.0039.

Bush:2023:MPA

[2572] Martin Bush. Mary Proctor: an astronomical popularizer in the shadows. Notes and Records of the Royal Society of London, 77(1):49–72, March ??, 2023. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsnr.2021.0042.

Gutierrez:2023:IHN

[2573] Sara Gutierrez, Stephanie L. Canington, Andrea R. Eller, Elizabeth S. Herrelko, and Sabrina B. Sholts. The intertwined history of nonhuman primate health and human medicine at the Smithsonian’s National Zoo and Conservation Biology Institute. Notes and Records of the Royal Society of London, 77(1):73–96, March ??, 2023. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsnr.2021.0009.

Hunter:2023:THS

[2574] Michael Hunter. Thomas Henshaw’s strange sùance in Venice, circa 1648: a coda to Robert Boyle by himself and his friends. Notes and Records of the Royal Society of London, 77(1):97–112, March ??, 2023. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsnr.2021.0019.

Miccoli:2023:NHR

[2575] Stefano Miccoli, Luisa María Gil-Martín, and Enrique Hernández-Montes. New historical records about the construction of the Arch of Ctesiphon and their impact on the history of structural engineering. Notes and Records of the Royal Society of London, 77(1):113–134, March ??, 2023. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/10.
1098/rsnr.2021.0025.

Belteki:2023:GSO

[2576] Daniel Belteki. ‘The grand strategy of an observatory’: George Airy’s vision for the division of astronomical labour among observatories during the nineteenth century. Notes and Records of the Royal Society of London, 77(1):135–151, March ??, 2023. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsnr.2021.0034.

Orosz:2023:FST

[2577] Ferenc Orosz and Miklós Müller. The first ‘Soviet type’ research institute of the Hungarian Academy of Sciences and its Stalin Prizeawarded director, Imre Szörényi. Notes and Records of the Royal Society of London, 77(1):153–168, March ??, 2023. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https://

royalsocietypublishing.org/doi/epdf/10.1098/rsnr.2021.0010.

Laurence:2023:PPO

[2578] Alison Laurence. Pleistocene Park, and other designs on deep time in the Interwar United States. Notes and Records of the Royal Society of London, 77(1):169–190, March ??, 2023. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL https://

royalsocietypublishing.org/doi/epdf/10.1098/rsnr.2021.0032.

Sims:2023:YVO

[2579] Liam Sims. ‘Your very obliging correspondence’: the Royal Society and the provincial Republic of Letters in Georgian Lincolnshire. Notes and Records of the Royal Society of London, 77(1):191–212, March ??, 2023. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/10.
1098/rsnr.2021.0053.

Verburgt:2023:TRW

[2580] Lukas M. Verburgt. ‘Is there a Reader who can Handle it with any Comfort?’: A brief publication history of the works of Francis Bacon. Notes and Records of the Royal Society of London, 77(1):213–220, March ??, 2023. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/10.
1098/rsnr.2020.0072.

Maerker:2023:EFN

[2581] Anna Maerker, Elena Serrano, and Simon Werrett. Enlightened female networks: gendered ways of producing knowledge (1720–1830). Notes and Records of the Royal Society of London, 77(2):225–234, June ??, 2023. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Esfandiary:2023:TEL

[2582] Helen Esfandiary. ‘A thankless enterprise’: Lady Mary Wortley Montagu’s campaign to establish medical unorthodoxy amongst her female network. Notes and Records of the Royal Society of London, 77(2):235– 250, June ??, 2023. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Charles:2023:FEP

[2583] Loïc Charles and Christine Théré. Les femmes économistes: the place of women in the physiocratic community. Notes and Records of the Royal Society of London, 77(2):251–264, June ??, 2023. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

FontesdaCosta:2023:GBE

[2584] Palmira Fontes da Costa. Gender and botany in early nineteenth-century Portugal: the circle of the Marquise of Alorna. Notes and Records of the Royal Society of London, 77(2):265–282, June ??, 2023. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Antonelli:2023:MLW

[2585] Francesca Antonelli. Madame Lavoisier and the others: women in Marie-Anne Paulze-Lavoisier’s network (1771–1836). Notes and Records of the Royal Society of London, 77(2):283–302, June ??, 2023. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Bolufer:2023:MCK

[2586] Mónica Bolufer and Elena Serrano. Maritime crossroads: the knowledge pursuits of María de Betancourt (Tenerife, 1758–1824) and Joana de Vigo (Menorca, 1779–1855). Notes and Records of the Royal Society of London, 77(2):303–322, June ??, 2023. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Hansen:2023:QCS

[2587] Mascha Hansen. Queen Charlotte’s scientific collections and natural history networks. Notes and Records of the Royal Society of London, 77(2): 323–336, June ??, 2023. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Govoni:2023:FNB

[2588] Paola Govoni. Feminist networks beyond the science wars: the ‘female brain’ in the 1790s and the 1990s. Notes and Records of the Royal Society of London, 77(2):337–352, June ??, 2023. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Mihailescu:2023:GDS

[2589] Ion Mihailescu. Graphical details: the secret life of Christopher Wren’s drawing of the weather clock. Notes and Records of the Royal Society of London, 77(3):355–378, September ??, 2023. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

deMFigueiroa:2023:SPN

[2590] Silvia F. de M.Figueirôa. Science popularization in nineteenth century France: Nérée Boubée (1806–1862) and the journal L’Écho du Monde Savant. Notes and Records of the Royal Society of London, 77(3):379– 393, September ??, 2023. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Newman:2023:SIM

[2591] Francis Newman. Scientific ideologies on the move: Sino-British exchanges, scientific freedoms and the governance of science in Britain, 1961–1966. Notes and Records of the Royal Society of London, 77(3):399– 420, September ??, 2023. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Wess:2023:NLR

[2592] Jane Wess. New light on the role of instruments in exploration during the 1830s. Notes and Records of the Royal Society of London, 77(3):421– 442, September ??, 2023. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

JoannaPark:2023:HHI

[2593] Louise Neilson Joanna Park and Andreas K. Demetriades. Hysteria, head injuries and heredity: ‘shell-shocked’ soldiers of the Royal Edinburgh Asylum, Edinburgh (1914–24). Notes and Records of the Royal Society of London, 77(3):443–470, September ??, 2023. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Jenkins:2023:SBT

[2594] Bill Jenkins. The ‘Stronsay Beast’: testimony, evidence and authority in early nineteenth-century natural history. Notes and Records of the Royal Society of London, 77(3):471–494, September ??, 2023. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Holt:2023:AEP

[2595] Paddy Holt. An appetite for experiment: putting early Royal Society tastes back on the table. Notes and Records of the Royal Society of London, 77(3):495–512, September ??, 2023. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Clary:2023:FMR

[2596] David C. Clary. Foreign membership of the Royal Society: Schrödinger and Heisenberg? Notes and Records of the Royal Society of London, 77 (3):513–536, September ??, 2023. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Butler:2023:TNL

[2597] Stella V. F. Butler. Two Nobel laureates in conversation: Robert Robinson listens to Dorothy Hodgkin’s account of her life scientific. Notes and Records of the Royal Society of London, 77(3):537–556, September ??, 2023. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Erken:2023:LWE

[2598] Ali Erken. Localizing western expertise: İhsan Doğramaci, Ş. Raşit Hatipoğlu, and the quest for scientific development in modern Turkey. Notes and Records of the Royal Society of London, 77(3):557–574, September ??, 2023. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Jurkowlaniec:2023:BCJ

[2599] Grażyna Jurkowlaniec. Bones of contention: Johann Heinrich Merck’s palaeontological encounters with academic scholars and professional printmakers. Notes and Records of the Royal Society of London, 77 (3):575–604, September ??, 2023. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Zhihui:2023:ERI

[2600] Zhang Zhihui. Emigration or return? International mobility and Theodore von Kármán’s Chinese students and associates. Notes and Records of the Royal Society of London, 77(3):605–646, September ??, 2023. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Dickenson:2023:LGB

[2601] Victoria Dickenson. Lady Gwillim and the birds of Madras. Notes and Records of the Royal Society of London, 77(3):647–670, September ??, 2023. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Wolfe:2023:LME

[2602] Charles T. Wolfe. The life of matter: early modern vital matter theories introduction. Notes and Records of the Royal Society of London, 77 (4):673–675, December ??, 2023. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Giglioni:2023:LLF

[2603] Guido Giglioni. Large as life: Francis Bacon on the animate matter of plants. Notes and Records of the Royal Society of London, 77(4):677– 696, December ??, 2023. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Georgescu:2023:CL

[2604] Laura Georgescu. Cavendish on life. Notes and Records of the Royal Society of London, 77(4):697–715, December ??, 2023. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Clericuzio:2023:TWI

[2605] Antonio Clericuzio. Thomas Willis’ iatrochemistry and the activity of matter. Notes and Records of the Royal Society of London, 77(4):717– 732, December ??, 2023. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Garau:2023:GST

[2606] Rodolfo Garau. Gassendi’s second thought. from a materialistic picture of cognition to the defence of dualism: the lasting influence of the polemic with Descartes. Notes and Records of the Royal Society of London, 77 (4):733–751, December ??, 2023. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Demeter:2023:SOB

[2607] Tamás Demeter. Sympathetic organizations: body, mind, and society in Robert Whytt and David Hume. Notes and Records of the Royal Society of London, 77(4):753–769, December ??, 2023. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Pecere:2023:MLL

[2608] Paolo Pecere. Materialism, Lebenskraft and the limits of science: metaphysical vitalism in post-Kantian scenarios. Notes and Records of the Royal Society of London, 77(4):771–787, December ??, 2023. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Roos:2024:EJJ

[2609] Anna Marie Eleanor Roos. Eloge to James (Jim) Arthur Bennett April 1947– 28 October 2023. Notes and Records of the Royal Society of London, 78(1):3–8, March ??, 2024. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Fielding:2024:CNB

[2610] Russell Fielding. ‘The correct name for the breadfruit’: on interdisciplinarity and the artist Sydney Parkinson’s contested contributions to the botanical sciences. Notes and Records of the Royal Society of London, 78(1):9–28, March ??, 2024. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Rabinovitch:2024:SWS

[2611] Oded Rabinovitch. The ‘system of the world’ and the scientific culture of early modern France. Notes and Records of the Royal Society of London, 78(1):29–52, March ??, 2024. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Peters:2024:CRH

[2612] Winfried S. Peters. The cells of Robert Hooke: pores, fibres, diaphragms and the cell theory that wasn’t. Notes and Records of the Royal Society of London, 78(1):53–74, March ??, 2024. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Garrett:2024:SVU

[2613] Chris Garrett. A scientific visit to the USSR in 1963. Notes and Records of the Royal Society of London, 78(1):75–80, March ??, 2024. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Mackenzie:2024:NGI

[2614] Pamela Mackenzie. Nehemiah Grew, the illustrator. Notes and Records of the Royal Society of London, 78(1):81–114, March ??, 2024. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Fyfe:2024:PBE

[2615] Aileen Fyfe. From philanthropy to business: the economics of Royal Society journal publishing in the twentieth century. Notes and Records of the Royal Society of London, 78(1):115–142, March ??, 2024. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Frost:2024:GEC

[2616] Robert Frost. A geologist and an Egyptologist in conversation: Sir Charles Lyell and Sir John Gardner Wilkinson. Notes and Records of the Royal Society of London, 78(1):143–166, March ??, 2024. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Dagg:2024:PMS

[2617] Joachim L. Dagg and J. F. Derry. Patrick Matthew’s synthesis of catastrophism and transformism. Notes and Records of the Royal Society of London, 78(1):167–188, March ??, 2024. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Vozar:2024:RHI

[2618] Thomas Matthew Vozar. Robert Hooke, Isaac Newton and the Royal Society: three unnoticed letters at the Staatsbibliothek zu Berlin. Notes and Records of the Royal Society of London, 78(1):189–202, March ??, 2024. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Seddon:2024:ODF

[2619] Elaine A. Seddon and Michael W. Poole. The origins and development of free-electron lasers in the UK. Notes and Records of the Royal Society of London, 78(1):203–252, March ??, 2024. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Abir-Am:1987:UCI

[2620] Pnina G. Abir-Am and Dorinda Outram, editors. Uneasy Careers and Intimate Lives: Women in Science, 1789–1979. The Douglass series on women’s lives and the meaning of gender. Rutgers University Press, New Brunswick, NJ, USA, 1987. ISBN 0-8135-1255-7 (hardcover), 0-8135-1256-5 (paperback). xii + 365 pp. LCCN Q130 .U525 1987. URL http:

//www.loc.gov/catdir/enhancements/fy1310/87009660-d.html.

Alder:2002:MAT

[2621] Ken Alder. The Measure of All Things: the Seven-year Odyssey and Hidden Error That Transformed the World. Little, Brown and Co., Boston, MA, USA, 2002. ISBN 0-316-85989-3. xii + 466 pp. LCCN QB291 .A43 2002.

Alic:1986:HHH

[2622] Margaret Alic. Hypatia’s Heritage: a History of Women in Science from Antiquity to the Late Nineteenth Century. Women’s Press, London, UK, 1986. ISBN 0-7043-2860-7 (hardcover), 0-7043-3954-4 (paperback). ix + 230 + 8 pp. LCCN Q130 .A48 1986.

Anderson:1987:SMD

[2623] R. G. W. (Robert Geoffrey William) Anderson and Christopher Lawrence, editors. Science, Medicine, and Dissent: Joseph Priestley (1733–1804): Papers Celebrating the 250th Anniversary of the Birth of Joseph Priestley Together with a Catalogue of an Exhibition Held at the Royal Society and the Wellcome Institute for the History of Medicine. Wellcome Trust/Science Museum, London, UK, 1987. ISBN 0-901805-28-9. xi + 105 pp. LCCN QD22.P8 S35x 1987.

Anderson:2012:CJB

[2624] R. G. W. (Robert Geoffrey William) Anderson and Jean Jones, editors. The correspondence of Joseph Black. Ashgate, Farnham, Surrey, England, 2012. ISBN 0-7546-0131-5 (hardcover). xiv + 1564 (two volumes) pp. LCCN QD22.B549 A4 2012.

Ashcroft:2000:LE

[2625] Frances M. Ashcroft. Life at the Extremes. HarperCollins, London, UK, 2000. ISBN 0-00-255946-3. xxi + 326 pp. LCCN QP82 .A75 2000.

Austin:2001:SSS

[2626] Brian Austin. Schonland: Scientist and Soldier. IOP Publishing, Bristol, UK, 2001. ISBN 0-7503-0501-0 (hardcover), 1-4200-3357-3 (PDF e-book). xviii + 639 pp. LCCN QC16.S2646 A97 2001; Q143.S36 A87

2001. URL http://www.crcnetbase.com/isbn/9781420033571.

Babbage:1991:PLP

[2627] Charles Babbage. Passages from the Life of a Philosopher: the Autobiography of Charles Babbage. Pickering and Chatto, London, UK, 1991. ISBN 1-85196-040-6 (paperback). viii + 383 pp. LCCN QA29.B2 B333 1991. Edited, and with a new introduction, by Martin Campbell-Kelly.

Barkan:1999:WNT

[2628] Diana Kormos Barkan. Walther Nernst and the Transition to Modern Physical Science. Cambridge University Press, Cambridge, UK, 1999. ISBN 0-521-44456-X (hardcover), 0-521-17629-8 (paperback). xii + 288 pp. LCCN QD22.N39 B37 1999. URL

http://www.loc.gov/catdir/description/cam029/98022028.html;
 http://www.loc.gov/catdir/toc/cam026/98022028.html.

Barrow:2000:BN

[2629] John D. Barrow. The Book of Nothing. Jonathan Cape, London, UK, 2000. ISBN 0-224-05962-9. xiv + 380 pp. LCCN QA141 .B35 2000. URL

ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/;   http:/
/www.loc.gov/cat1dir/description/random041/2002421946.html;
 http://www.loc.gov/catdir/bios/random054/2002421946.html.

Bartle:1994:ORH

[2630] G. F. Bartle. An Old Radical and His Brood: a Portrait of Sir John Bowring and His Family Based Mainly on the Correspondence of Bowring and His Son, Frederick Bowring. Janus Publishing, London, UK, 1994. ISBN 1-85756-132-5 (paperback). 139 pp. LCCN DA565.B675 B37 1994.

BassoRicci:1997:DSS

[2631] M. Basso Ricci, L. Cafarella, A. Meloni, and P. Tucci. Due Secoli di Strumenti Geomagnetici in Italia (1740–1971) = Two Centuries of Geomagnetic Instruments in Italy (1740–1971): with a Summary in English. Editrice Compositori, Bologna, Italia, 1997. ISBN 88-7794-091-3. 234 pp.

LCCN QC819 .D84 1997. URL http://www.gbv.de/dms/goettingen/
241534852.pdf.

Beardsley:2002:DWLa

[2632] Martyn Beardsley. Deadly Winter: the Life of Sir John Franklin. Chatham, London, UK, 2002. ISBN 1-86176-187-2. xiv + 272 + 8

pp. LCCN G660 .B43 2002. URL http://www.loc.gov/catdir/toc/
fy051/2003430707.html.

Beardsley:2002:DWLb

[2633] Martyn Beardsley. Deadly Winter: the Life of Sir John Franklin. Naval Institute Press, Annapolis, MD, USA, 2002. ISBN 1-55750-179-3. xiv + 272 + 8 pp. LCCN G660 .B43 2002.

Bearn:1993:AGI

[2634] Alexander G. Bearn. Archibald Garrod and the Individuality of Man. Clarendon Press, Oxford, UK, 1993. ISBN 0-19-262145-9. xvi + 227 pp. LCCN QP511.8.G37 B4 1993. URL http://www.

loc.gov/catdir/enhancements/fy0638/92027504-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0638/92027504-t.html.

Beattie:1998:FTF

[2635] Owen Beattie and John Geiger. Frozen in Time: the Fate of the Franklin Expedition. Greystone Books, Vancouver, BC, Canada, 1998. ISBN 1-55054-616-3 (paperback). xi + 185 + 24 pp. LCCN G670 1981 .F73

B43 1998. URL http://www.loc.gov/catdir/description/ste021/
00688217.html.

Beer:1996:OFS

[2636] Gillian Beer. Open Fields: Science in Cultural Encounter. Clarendon Press, Oxford, UK, 1996. ISBN 0-19-818369-0 (hardcover), 0-19-818635-5 (paperback). viii + 341 pp. LCCN PR468.S34 B44 1996. URL http:

//www.loc.gov/catdir/enhancements/fy0639/95043234-d.html.

Bennett:2003:LLL

[2637] J. A. (James A.) Bennett, Michael Cooper, Michael Hunter, and Lisa Jardine. London’s Leonardo: the Life and Work of Robert Hooke. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2003. ISBN 0-19-852579-6. xii + 224 pp. LCCN Q143.H7 L66 2003. URL

http://www.e-streams.com/es0612/es0612\_2943.html;   http:/
/www.h-net.org/review/hrev-a0e0c8-aa;   http://www.loc.gov/
catdir/enhancements/fy0616/2003545106-d.html;   http://www.
loc.gov/catdir/enhancements/fy0616/2003545106-t.html;   http:
//www.loc.gov/catdir/enhancements/fy0723/2003545106-b.html.

Beretta:1993:HNP

[2638] Marco Beretta. A History of Non-printed Science: a Select Catalogue of the Waller Collection, volume 14; 31; 32 of Uppsala studies in history of science; Acta Bibliothecae R. Universitatis Upsaliensis; Uppsala universitetsbiblioteks utställninskataloger. University Library, Uppsala, Sweden, 1993. ISBN 91-554-3070-8. ISSN 0346-7465. 198 pp. LCCN Q173 .B47 1993; Z7409 .B42 1993.

Beretta:2001:ICS

[2639] Marco Beretta. Imaging a Career in Science: the Iconography of Antoine Laurent Lavoisier, volume 1 (Bologna), 29 (Uppsala) of Bologna studies in scientific heritage; Uppsala studies in history of science. Science History Publications/USA, Canton, MA, USA and Sagamore Beach, MA, USA, 2001. ISBN 0-88135-294-2. xvii + 126 pp. LCCN

QD22.L4 B29 2001. URL http://www.loc.gov/catdir/toc/fy031/
2001049615.html.

Bindman:1993:WPA

[2640] Lynn Bindman. Women Physiologists: an Anniversary Celebration of Their Contributions to British Physiology. Portland Press, London, UK, 1993. ISBN 1-85578-049-6. ix + 166 pp. LCCN QP25 .W66 1993.

Bliss:1999:WOL

[2641] Michael Bliss. William Osler: a Life in Medicine. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1999. ISBN 0-19-512346-8. xiv + 581 pp. LCCN R134 .B55 1999.

URL http://lcweb.loc.gov/catdir/toc/99032066.html;   http:
//www.loc.gov/catdir/enhancements/fy0604/99032066-d.html;
http://www.loc.gov/catdir/enhancements/fy0726/99032066-b.html.

Bondeson:1997:CMC

[2642] Jan Bondeson. A Cabinet of Medical Curiosities. I. B. Tauris, London, UK, 1997. ISBN 1-86064-228-4. ix + 250 pp. LCCN R706.B6.

Bowden:1991:PRL

[2643] Mark Bowden. Pitt Rivers: the Life and Archaeological Work of Lieutenant-General Augustus Henry Lane Fox Pitt Rivers, DCL, FRS, FSA. Cambridge University Press, Cambridge, UK, 1991. ISBN 0-521-40077-5. xv + 182 pp. LCCN DA3.P53 B69 1991. URL http://www.

loc.gov/catdir/enhancements/fy0903/90045313-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0903/90045313-t.html.

Bowler:1990:CDM

[2644] Peter J. Bowler. Charles Darwin: the Man and His Influence. Blackwell scientific biographies. Blackwell, Oxford, UK, 1990. ISBN 0-631-16818-4. xii + 250 pp. LCCN QH31.D2 B743 1991.

Brock:1997:JLC

[2645] W. H. (William Hodson) Brock. Justus Von Liebig: the Chemical Gatekeeper. Cambridge University Press, Cambridge, UK, 1997. ISBN 0-521-56224-4 (hardcover). xiv + 374 pp. LCCN QD22.L7 B76 1997. URL

http://www.loc.gov/catdir/description/cam027/96025468.html;
 http://www.loc.gov/catdir/samples/cam031/96025468.html;   http:
//www.loc.gov/catdir/toc/cam031/96025468.html.

Brown:1991:BPS

[2646] R. Hanbury (Robert Hanbury) Brown. Boffin: a Personal Story of the Early Days of Radar, Radio Astronomy, and Quantum Optics. Adam Hilger Ltd., Bristol, UK, 1991. ISBN 0-7503-0130-9. ii + 184 pp. LCCN UG612 .H26 1991.

Brown:1999:CRE

[2647] G. I. (George Ingham) Brown. Count Rumford: the Extraordinary Life of a Scientific Genius. Sutton, Stroud, Gloucestershire, UK, 1999. ISBN 0-7509-2184-6 (paperback). x + 182 + 8 pp. LCCN Q143.R8 B76 1999.

Brown:2002:IRH

[2648] George Ingham Brown. Invisible Rays: the History of Radioactivity. Sutton, Stroud, Gloucestershire, UK, 2002. ISBN 0-7509-2667-8 (hardcover). 248 pp. LCCN QC794.98 .B76 2002. URL http://www.gbv.de/dms/

bowker/toc/9780750926676.

Browne:1995:CD

[2649] E. J. (E. Janet) Browne. Charles Darwin. Alfred A. Knopf, New York, NY, USA, 1995. ISBN 0-224-04202-5 (vol. 1), 0-224-04212-2 (vol. 2). xiii + 605 + 32 pp. LCCN QH31.D2 B84 1995.

Browne:1995:CDV

[2650] Janet Browne. Charles Darwin: Voyaging: Volumme I of a Biography. Jonathan Cape, London, UK, 1995. ISBN 0-224-04202-5. xiii + 605 + 32 pp. LCCN QH31.D2 B76 1995.

Cahn:2001:CMS

[2651] R. W. (Robert W.) Cahn. The Coming of Materials Science, volume 5 of Pergamon materials series. Pergamon, New York, NY, USA, 2001. ISBN 0-08-042679-4, 1-281-07198-6, 0-08-052942-9 (ebook). xvii + 568 pp. LCCN TA403 .C225 2001. URL http://www.

loc.gov/catdir/enhancements/fy0610/00065565-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0610/00065565-t.html.

Campbell:1999:RSS

[2652] John Campbell. Rutherford: Scientist Supreme. AAS Publications, Christchurch, New Zealand, 1999. ISBN 0-473-05700-X (hardcover). xvi + 516 + 32 + 16 pp. LCCN QC16.R8 C35 1999.

Cannell:1992:GGM

[2653] Doris Mary Cannell. George Green, Mathematician and Physicist: 1793– 1841: the Background to His Life and Work. Athlone Press, London, UK, 1992. ISBN 0-485-11433-X. xxvi + 265 + 4 pp. LCCN QC16.G64 C36

Cardwell:1989:JJB

[2654] D. S. L. (Donald Stephen Lowell) Cardwell. James Joule: a Biography. Manchester University Press, Manchester, UK, 1989. ISBN 0-7190-3025-0, 0-7190-3479-5 (paperback). ix + 333 pp. LCCN QC16.J68 C368 1989.

Caroe:1985:RII

[2655] G. M. (Gwendolen Mary) Caroe. The Royal Institution: an Informal History. J. Murray, London, UK, 1985. ISBN 0-7195-4245-6. xi + 180 + 12 pp. LCCN Q41 .C27 1985. With a final chapter by Alban Caroe.

Carter:1987:SJB

[2656] Harold B. (Harold Burnell) Carter. Sir Joseph Banks (1743–1820): a Guide to Biographical and Bibliographical sources. St Paul’s Bibliographies in association with British Museum (Natural History), Winchester, UK, 1987. ISBN 0-906795-45-1. 328 pp. LCCN MLCM 91/09660 (Q).

Carter:1988:SJB

[2657] Harold B. (Harold Burnell) Carter. Sir Joseph Banks, 1743–1820. British Museum (Natural History), London, UK, 1988. ISBN 0-565-00993-1. xi + 671 pp. LCCN Q143.B3 C37 1988.

Cartwright:1999:TSH

[2658] David Edgar Cartwright. Tides: a Scientific History. Cambridge University Press, Cambridge, UK, 1999. ISBN 0-521-62145-3 (hardcover), 0-521-79746-2 (paperback). xii + 292 pp. LCCN GC301.2 .C37 1999. URL

http://www.loc.gov/catdir/description/cam029/98004660.html;
 http://www.loc.gov/catdir/samples/cam031/98004660.html;   http:
//www.loc.gov/catdir/toc/cam021/98004660.html.

Cassidy:1992:ULS

[2659] David C. Cassidy. Uncertainty: the Life and Science of Werner Heisenberg. W. H. Freeman and Company, New York, NY, USA, 1992. ISBN 0-7167-2243-7. xii + 669 + 16 pp. LCCN QC16.H35 C37 1992; QC 16 .H35C37 1992.

Cercignani:1998:LBM

[2660] Carlo Cercignani. Ludwig Boltzmann: the Man Who Trusted Atoms. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1998. ISBN 0-19-850154-4 (hardcover), 0-19-857064-3 (paperback). xvi + 329 pp. LCCN QC16.B64 C47 1998. URL http://www.

loc.gov/catdir/enhancements/fy0605/98017743-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0605/98017743-t.html. Fore-

word by Roger Penrose.

Chambers:2000:LSJ

[2661] Neil Chambers. The Letters of Sir Joseph Banks: a Selection, 1768– 1820. Imperial College Press, London, UK, 2000. ISBN 1-86094-204-0. xliii + 420 pp. LCCN Q143.B17 A4 2000.

Chandrasekhar:1989:PPH

[2662] S. (Subrahmanyan) Chandrasekhar. Plasma Physics, Hydrodynamic and Hydromagnetic Stability, and Applications of the Tensor–virial Theorem, volume 4 of Selected papers. University of Chicago Press, Chicago, IL, USA, 1989. ISBN 0-226-10096-0, 0-226-10097-9 (paperback). xiii + 585 pp. LCCN QC718 .C38 1989. URL http://www.

loc.gov/catdir/enhancements/fy0608/89004906-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0608/89004906-t.html.

Chandrasekhar:1989:RTN

[2663] S. (Subrahmanyan) Chandrasekhar. Radiative Transfer and Negative Ion of Hydrogen, volume 2 of Selected papers. University of Chicago Press, Chicago, IL, USA, 1989. ISBN 0-226-10092-8, 0-226-10093-6 (paperback). xvi + 622 pp. LCCN QB809 .C47 1989. URL http://www.

loc.gov/catdir/enhancements/fy0608/88026111-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0608/88026111-t.html.

Chandrasekhar:1989:SSH

[2664] S. (Subrahmanyan) Chandrasekhar. Stochastic, Statistical, and Hydromagnetic Problems in Physics and Astronomy, volume 3 of Selected papers. University of Chicago Press, Chicago, IL, USA, 1989. ISBN 0-226-10094-4, 0-226-10095-2 (paperback). xiv + 642 pp. LCCN QB461

.C47 1989. URL http://www.loc.gov/catdir/enhancements/fy0608/
89004883-d.html;   http://www.loc.gov/catdir/enhancements/fy0608/
89004883-t.html.

Chandrasekhar:1989:SSS

[2665] S. (Subrahmanyan) Chandrasekhar. Stellar Structure and Stellar Atmospheres, volume 1 of Selected papers / S. Chandrasekhar. University of Chicago Press, Chicago, IL, USA, 1989. ISBN 0-226-10089-8, 0-226-10090-1 (paperback). xiv + 515 pp. LCCN QB808 .C47 1989. URL http:

//www.loc.gov/catdir/enhancements/fy0608/88026110-d.html;
http://www.loc.gov/catdir/enhancements/fy0608/88026110-t.html;
 http://www.loc.gov/catdir/enhancements/fy1304/88026110-b.html.

Chandrasekhar:1990:RA

[2666] S. (Subrahmanyan) Chandrasekhar. Relativistic Astrophysics, volume 5 of Selected papers / S. Chandrasekhar. University of Chicago Press, Chicago, IL, USA, 1990. ISBN 0-226-10098-7, 0-226-10099-5 (paperback). xx + 587 pp. LCCN QB461 .C468 1990. URL http://www.

loc.gov/catdir/enhancements/fy0608/89020400-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0608/89020400-t.html.

Chandrasekhar:1991:MTB

[2667] S. (Subrahmanyan) Chandrasekhar. The Mathematical Theory of Black Holes and of Colliding Plane Waves, volume 6 of Selected papers. University of Chicago Press, Chicago, IL, USA, 1991. ISBN 0-226-10100-2 (vol. 6), 0-226-10101-0 (vol. 6, paperback). xvii + 739 pp. LCCN QB843.B55 C483 1991. URL http://www.

loc.gov/catdir/enhancements/fy0608/90011149-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0608/90011149-t.html.

Chandrasekhar:1995:NPC

[2668] S. (Subrahmanyan) Chandrasekhar. Newton’s Principia for the Common Reader. Clarendon Press, Oxford, UK, 1995. ISBN 0-19-851744-0. xix + 593 pp. LCCN QA803 .C48 1995. URL http://www.

loc.gov/catdir/enhancements/fy0606/95001800-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0606/95001800-t.html.

Chisholm:2002:SSC

[2669] Monty Chisholm. Such Silver Currents: the Story of William and Lucy Clifford, 1845–1929. Lutterworth Press, Cambridge, UK, 2002. ISBN 0-7188-3017-2 (paperback). ix + 198 pp. LCCN QA29.C56 C45 2002.

Ciancio:1995:CCJ

[2670] Luca Ciancio. A Calendar of the Correspondence of John Strange F.R.S. (1732–1799), volume 2 of Occasional publications // Wellcome Institute for the History of Medicine. Wellcome Inst. for the History of Medicine, London, UK, 1995. ISBN 0-85484-063-X. 152 pp. LCCN QH31.S783 C27 1995.

Cohen:1999:PMP

[2671] I. Bernard Cohen and Anne Miller Whitman. The Principia: Mathematical Principles of Natural Philosophy. University of California Press, Berkeley, CA, USA, 1999. ISBN 0-520-08816-6 (hardcover), 0-520-08817-4 (paperback), 0-520-96091-2 (e-book). xvii + 966 pp. LCCN

QA803 .N413 1999. URL http://www.loc.gov/catdir/bios/ucal052/
99010278.html;   http://www.loc.gov/catdir/description/ucal041/

99010278.html. A new translation by I. Bernard Cohen and Anne Whitman and assisted by Julia Budenz. Preceded by a guide to Newton’s Principia by I. Bernard Cohen.

ConwayMorris:1998:CCB

[2672] S. (Simon) Conway Morris. The Crucible of Creation: the Burgess Shale and the Rise of Animals. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1998. ISBN 0-19-850256-7 (hardcover), 0-19-286202-2 (paperback). xxiii + 242 + 4 pp. LCCN QE770 .C67 1998. URL http:/

/catdir.loc.gov/catdir/enhancements/fy0636/97027752-d.html;
 http://catdir.loc.gov/catdir/enhancements/fy0636/97027752-
t.html;   http://catdir.loc.gov/catdir/enhancements/fy0723/
97027752-b.html.

Cook:1998:EHC

[2673] Alan H. Cook. Edmond Halley: Charting the Heavens and the Seas. Clarendon Press, Oxford, UK, 1998. ISBN 0-19-850031-

9. xvi + 540 pp. LCCN QB36.H25 C66 1998. URL http:/

/www.gbv.de/dms/bowker/toc/9780198500315.pdf;   http://www.
gbv.de/dms/goettingen/232124825.pdf;   http://www.loc.gov/catdir/
enhancements/fy0635/97012596-d.html;   http://www.loc.gov/catdir/
enhancements/fy0635/97012596-t.html;   http://www.loc.gov/catdir/
enhancements/fy0723/97012596-b.html.

Cookman:2000:IBT

[2674] Scott Cookman. Ice Blink: the Tragic Fate of Sir John Franklin’s Lost Polar Expedition. John Wiley, New York, NY, USA, 2000. ISBN 0-471-37790-2 (hardcover), 0-471-40420-9. xii + 244 pp. LCCN G660 .C66 2000.

URL http://www.loc.gov/catdir/bios/wiley046/99047620.html;
 http://www.loc.gov/catdir/description/wiley0310/99047620.html;
 http://www.loc.gov/catdir/toc/onix01/99047620.html.

Cookson:2000:VSE

[2675] Gillian Cookson and Colin A. Hempstead. A Victorian Scientist and Engineer: Fleeming Jenkin and the Birth of Electrical Engineering. Ashgate, Aldershot, UK, 2000. ISBN 0-7546-0079-3. xii + 217 pp.

LCCN Q143.J5 H46 2000. URL http://opac.nebis.ch/cgi-bin/
showAbstract.pl?u20=0754600793.

Cooper:2003:MBC

[2676] Michael Cooper. “A More Beautiful City”: Robert Hooke and the Rebuilding of London After the Great Fire. Sutton, Stroud, Gloucestershire, UK, 2003. ISBN 0-7509-2959-6. vii + 271 pp. LCCN Q143.H7 C666 2003.

Craven:1996:JWD

[2677] Maxwell Craven. John Whitehurst of Derby: Clockmaker and Scientist, 1713–88. Mayfield Books, Mayfield, Ashbourne, Derbyshire, England, 1996. ISBN 0-9523270-3-1 (hardcover). 272 pp. LCCN TS544.8.W48 C73 1996.

Crick:1989:WMP

[2678] Francis Crick. What Mad Pursuit: a Personal View of Scientific Discovery. Penguin Books, Harmondsworth, UK, 1989. ISBN 0-14-011973-6 (paperback). ix + 182 + 8 pp. LCCN QH31.C85 A3 1990.

Crosland:2002:GLS

[2679] Maurice P. Crosland. Gay-Lussac: Scientist and Bourgeois. Cambridge University Press, Cambridge, UK, 2002. ISBN 0-521-21979-5 (hardcover), 0-521-52483-0 (paperback), 0-511-56439-2 (e-book). xvi + 333 pp. LCCN QD22.G35 C76 1978.

Cruickshank:1992:PPE

[2680] D. W. J. (Durward W. J.) Cruickshank, Hellmut J. Juretschke, and N. (Norio) Katō, editors. P. P. Ewald and His Dynamical Theory of X-ray Diffraction: a Memorial Volume for Paul P. Ewald, 23 January 1888–22 August 1985, volume 2 of International Union of Crystallography monographs on crystallography. International Union of Crystallography, Chester, UK, 1992. ISBN 0-19-855379-X. x + 161 pp. LCCN QD945 .P22 1992. URL http://www.

loc.gov/catdir/enhancements/fy0606/92009876-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0606/92009876-t.html.

Dadswell:2002:SPG

[2681] Ted Dadswell. The Selborne Pioneer: Gilbert White as Naturalist and Scientist, a Re-examination. Ashgate, Aldershot, Hampshire, England, 2002. ISBN 0-7546-0749-6. xviii + 238 pp. LCCN QH31.W58 D34 2002.

URL http://www.loc.gov/catdir/toc/fy037/2002018205.html.

Dahl:1997:FCR

[2682] Per F. Dahl. Flash of the Cathode Rays: a History of J. J. Thomson’s Electron. IOP Publishing, Bristol, UK, 1997. ISBN 0-7503-0453-7 (hardcover), 0-585-20853-0 (e-book). xvii + 526 pp. LCCN 99.E01389.

Darwin:1988:CDB

[2683] Charles Darwin and R. D. Keynes. Charles Darwin’s Beagle diary. Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-23503-0. xxix + 464 pp. LCCN QH365 .A1 1988. URL

http://www.loc.gov/catdir/description/cam023/87031169.html;
 http://www.loc.gov/catdir/enhancements/fy0732/87031169-b.html;
 http://www.loc.gov/catdir/toc/cam025/87031169.html.

Dauben:1995:ARC

[2684] Joseph Warren Dauben. Abraham Robinson: the Creation of Nonstandard Analysis: a Personal and Mathematical Odyssey. Princeton University Press, Princeton, NJ, USA, 1995. ISBN 0-691-03745-0. xix + 559 pp. LCCN QA29.R57 D38 1995. URL http://www.loc.gov/

catdir/description/prin031/94032715.html;   http://www.loc.
gov/catdir/toc/prin031/94032715.html.

Davis:1996:RBF

[2685] Edward Bradford Davis and Michael Cyril William Hunter. Robert Boyle: A Free Enquiry Into the Vulgarly Received Notion of Nature. Cambridge texts in the history of philosophy. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-56100-0 (hardcover), 0-521-56796-3 (paperback). xxxvi + 171 pp. LCCN Q155 .B68 1996. URL

http://www.loc.gov/catdir/description/cam027/96010997.html;
 http://www.loc.gov/catdir/toc/cam024/96010997.html.

Davis:1997:JJT

[2686] Edward Arthur Davis and Isolde J. Falconer. J. J. Thomson and the Discovery of the Electron. Taylor and Francis, Boca Raton, FL, USA, 1997. ISBN 0-7484-0696-4 (hardcover), 0-7484-0720-0 (paperback). xxvii + 243 pp. LCCN QC16.T45 D38 1997. URL http://www.gbv.de/dms/

bowker/toc/9780748407200.pdf.

Deacon:1997:SSC

[2687] Terrence William Deacon. The Symbolic Species: the Co-evolution of Language and the Human Brain. Penguin Books, Harmondsworth, UK, 1997. ISBN 0-7139-9188-7 (hardcover), 0-14-026405-1 (paperback), 0-393-03838-6, 0-393-31754-4 (paperback). 527 pp. LCCN QP399 .D433 1998.

Denis:1994:PGE

[2688] Smith Denis, editor. Perceptions of Great Engineers: Fact and Fantasy. Science Museum for the Newcomen Society, London, UK, 1994. ISBN 0-901805-65-3. 115 + 8 pp. LCCN TA139 .P47 1994.

Desmond:1991:D

[2689] Adrian J. Desmond and James R. (James Richard) Moore. Darwin. Michael Joseph, London, UK, 1991. ISBN 0-7181-3430-3. xxi + 807 pp. LCCN QH31.D2 D47 1991.

Desmond:1994:HDD

[2690] Adrian J. Desmond. Huxley: the Devil’s Disciple. M. Joseph, London, UK, 1994. ISBN 0-7181-3641-1. xvii + 475 + 24 pp. LCCN Q143.H956 D47 1994.

Djerassi:2003:NDT

[2691] Carl Djerassi and David Pinner. Newton’s Darkness: Two Dramatic Views. Imperial College Press, London, UK, 2003. ISBN 1-86094-389-6, 1-86094-390-X (paperback), 1-86094-459-0 (e-book). 184 pp. LCCN PS3554.J47 N49 2003.

Domb:1996:CPH

[2692] Cyril Domb. The Critical Point: a Historical Introduction to the Modern Theory of Critical Phenomena. Taylor and Francis, Boca Raton, FL, USA, 1996. ISBN 0-7484-0435-X. xviii + 376 pp. LCCN QC173.4.C74

D66 1996. URL http://www.loc.gov/catdir/enhancements/fy0745/
96141603-d.html.

Donovan:1996:ALS

[2693] Arthur Donovan. Antoine Lavoisier: Science, Administration, and Revolution. Cambridge science biographies series. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-56672-X (paperback), 0-521-56218-X (hardcover). xvi + 351 pp. LCCN QD22.L4 D58 1996.

URL http://www.loc.gov/catdir/description/cam021/96178046.
html;   http://www.loc.gov/catdir/toc/cam021/96178046.html.

Drake:1996:RGR

[2694] Ellen Tan Drake. Restless Genius: Robert Hooke and His Earthly Thoughts. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1996. ISBN 0-19-506695-2 (hardcover). xiv + 386 pp.

LCCN QE22.H76 D74 1996. URL http://www.loc.gov/catdir/
enhancements/fy0602/95042703-d.html;   http://www.loc.gov/catdir/
enhancements/fy0602/95042703-t.html.

English:1990:VVL

[2695] Mary P. English. Victorian Values: the Life and Times of Dr. Edwin Lankester, M.D., F.R.S. Biopress, Bristol, England, 1990. ISBN 0-948737-14-X. xv + 187 pp. LCCN R489.L33 E54 1990.

Fara:1996:SAM

[2696] Patricia Fara. Sympathetic Attractions: Magnetic Practices, Beliefs, and Symbolism in Eighteenth-Century England. Princeton University Press, Princeton, NJ, USA, 1996. ISBN 0-691-01099-4 (hardcover). x + 326 pp. LCCN QC750.4 .F37 1996. URL http:/

/www.gbv.de/dms/bowker/toc/9780691010991.pdf;   http://www.
loc.gov/catdir/description/prin031/96010785.html;   http://
www.loc.gov/catdir/toc/prin031/96010785.html.

Fara:2002:NMG

[2697] Patricia Fara. Newton: the Making of a Genius. MacMillan Publishing Company, New York, NY, USA, 2002. ISBN 0-333-90735-3 (hardcover), 0-231-12806-1 (hardcover), 0-231-12807-X (paperback). xvi + 347 + 16 pp. LCCN QC16.N7 .F37 2002.

Fara:2004:PBW

[2698] Patricia Fara. Pandora’s Breeches: Women, Science and Power in the Enlightenment, volume 637 of Pimlico. Pimlico, London, UK, 2004. ISBN 1-84413-082-7. xi + 274 pp. LCCN Q130 .F37 2004. URL http://www.

loc.gov/catdir/enhancements/fy0628/2005415948-d.html.

Ferry:1998:DHL

[2699] Georgina Ferry. Dorothy Hodgkin: a Life. Granta, London, UK, 1998. ISBN 1-86207-285-X. viii + 423 + 16 pp. LCCN QD921 .F47 1998.

Fisher:1991:EJ

[2700] Richard B. Fisher. Edward Jenner: 1749–1823. André Deutsch Ltd., London, UK, 1991. ISBN 0-233-98681-2. 361 + 8 pp. LCCN R489.J5 F57 1991x.

Forbes:1995:CJF

[2701] Eric Gray Forbes, Lesley Carol Murdin, and Frances Willmoth. The Correspondence of John Flamsteed, the First Astronomer Royal. IOP Publishing, Bristol, UK, 1995–2002. ISBN 0-7503-0819-2 (set), 0-7503-0147-3 (vol. 1), 0-7503-0391-3 (vol. 2), 0-7503-0763-3 (vol. 3). xlix + 955 (vol. 1), xlvii + 1095 (vol. 2), lxvi + 1038 (vol. 3) pp. LCCN QB36.F53 A4

1995. URL http://catdir.loc.gov/catdir/enhancements/fy0668/
95047433-d.html.

Ford:1991:LL

[2702] Brian J. (Brian John) Ford. The Leeuwenhoek Legacy. Biopress, Bristol, UK, 1991. ISBN 1-85083-016-9 (Farrand), 0-948737-10-7 (Biopress). 185 pp. LCCN QH205.2 .F67 1991.

Ford:1992:ISH

[2703] Brian John Ford. Images of Science: a History of Scientific Illustration. British Library, London, UK, 1992. ISBN 0-7123-0267-0. viii + 208 pp. LCCN Q222 .F67 1992.

Fortey:2000:TEE

[2704] Richard A. Fortey. Trilobite!: Eyewitness to Evolution. HarperCollins, London, UK, 2000. ISBN 0-00-257012-2. xiii + 269 + 16 pp. LCCN QE821 .F67 2000.

Fozard:1991:SCH

[2705] John W. Fozard. Sydney Camm and the Hurricane: Perspectives on the Master Fighter Designer and His Finest Achievement. Airlife Publishing, Shrewsbury, UK, 1991. ISBN 1-85310-270-9. 258 pp. LCCN TL540.C313 S93 1991b.

Gascoigne:1994:JBE

[2706] John Gascoigne. Joseph Banks and the English Enlightenment: Useful Knowledge and Polite Culture. Cambridge University Press, Cambridge, UK, 1994. ISBN 0-521-45077-2 (hardcover). xi + 324 pp. LCCN Q127.G7

G37 1994. URL http://www.loc.gov/catdir/description/cam025/
93031696.html;   http://www.loc.gov/catdir/toc/cam027/93031696.
html.

Gascoigne:1998:SSE

[2707] John Gascoigne. Science in the Service of Empire: Joseph Banks, the British State and the Uses of Science in the Age of Revolution. Cambridge University Press, Cambridge, UK, 1998. ISBN 0-521-55069-6. vii + 247 pp. LCCN Q127.G7 G38 1998. URL

http://www.loc.gov/catdir/description/cam029/98192120.html;
 http://www.loc.gov/catdir/toc/cam027/98192120.html.

Gates:1998:KNV

[2708] Barbara T. Gates. Kindred Nature: Victorian and Edwardian Women Embrace the Living World. University of Chicago Press, Chicago, IL, USA, 1998. ISBN 0-226-28442-5 (hardcover), 0-226-28443-3 (paperback). xv + 293 pp. LCCN HQ1595.A3 G37 1998. URL http://www.loc.gov/

catdir/description/uchi052/98034934.html;   http://www.loc.
gov/catdir/toc/uchi052/98034934.html.

Gemelli:1996:ADC

[2709] Benedino Gemelli. Aspetti dell’atomismo classico nella filosofia di Francis Bacon e nel Seicento. (Italian) [Aspects of classical atomism in the philosophy of Francis Bacon and in the Seventeenth Century], volume 152 of Studi / Accademia toscana di scienze e lettere La Colombaria. L. S. Olschki, Firenze, Italia, 1996. ISBN 88-222-4460-5 (paperback). ISSN 0065-0781. 433 pp. LCCN B1198 .G45 1996.

Gerstner:1994:HDR

[2710] Patsy Gerstner. Henry Darwin Rogers, 1808–1866: American Geologist. History of American science and technology series. University of Alabama Press, Tuscaloosa, AB, USA, 1994. ISBN 0-8173-0735-4 (hardcover). xi + 311 pp. LCCN QE22.R64 G47 1994.

Gillham:2001:LSF

[2711] Nicholas W. Gillham. A Life of Sir Francis Galton: from African Exploration to the Birth of Eugenics. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2001. ISBN 0-19-514365-5. xii + 416 pp. LCCN QH429.2.G35 G53 2001. URL http://www.loc.gov/

catdir/enhancements/fy0612/2001021612-d.html;   http://www.
loc.gov/catdir/enhancements/fy0725/2001021612-b.html.

Golinski:1992:SPC

[2712] Jan Golinski. Science as Public Culture: Chemistry and Enlightenment in Britain, 1760–1820. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-39414-7 (hardcover), 0-521-65952-3 (paperback). xi + 342 pp. LCCN QD18.G7 G65 1992. URL

http://www.loc.gov/catdir/description/cam024/91039024.html;
 http://www.loc.gov/catdir/toc/cam029/91039024.html.

Golinski:1998:MNK

[2713] Jan Golinski. Making Natural Knowledge: Constructivism and the History of Science. Cambridge history of science. Cambridge University Press, Cambridge, UK, 1998. ISBN 0-521-44471-3, 0-521-44913-8 (paperback). xiv + 236 pp. LCCN Q125 .G63 1998. URL

http://www.loc.gov/catdir/description/cam028/97024028.html;
 http://www.loc.gov/catdir/toc/cam026/97024028.html.

Gouk:1999:MSN

[2714] Penelope Gouk. Music, Science, and Natural Magic in Seventeenthcentury England. Yale University Press, New Haven, CT, USA, 1999. ISBN 0-300-07383-6 (hardcover). xii + 308 pp. LCCN ML3800 .G66 1999.

Grant:1996:FMS

[2715] Edward Grant. The Foundations of Modern Science in the Middle Ages: Their Religious, Institutional, and Intellectual Contexts. Cambridge history of science. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-56137-X (hardcover), 0-521-56762-9 (paperback). xiv + 247 pp. LCCN Q124.97 .G68 1996. URL

http://www.loc.gov/catdir/description/cam027/95026530.html;
 http://www.loc.gov/catdir/toc/cam026/95026530.html.

Griffiths:1992:TML

[2716] John Charles Griffiths. The Third Man: the Life and Times of William Murdoch, 1754–1839, the Inventor of Gas Lighting. A. Deutsch, London, UK, 1992. ISBN 0-233-98778-9. 373 + 16 pp. LCCN T40.M87 G75 1992.

Gruber:1992:ROC

[2717] Jacob W. Gruber and John C. Thackray. Richard Owen Commemoration: Three Studies, volume 1 of Historical studies in the life and earth sciences. Natural History Museum Publications, London, UK, 1992. ISBN 0-565-01109-X. ix + 181 pp. LCCN QL31.O9 G78 1992.

Guicciardini:1999:RPD

[2718] Niccolò Guicciardini. Reading the Principia: the Debate on Newton’s Mathematical Methods for Natural Philosophy from 1687 to 1736. Cambridge University Press, Cambridge, UK, 1999. ISBN 0-521-64066-0 (hardcover). v + 285 pp. LCCN QA803 .G85 1999. URL

http://www.loc.gov/catdir/description/cam029/98051726.html;
 http://www.loc.gov/catdir/toc/cam021/98051726.html.

Hall:1990:HMM

[2719] A. Rupert (Alfred Rupert) Hall. Henry More — Magic, Religion, and Experiment. Blackwell science biographies. Blackwell, Oxford, UK, 1990. ISBN 0-631-17295-5. xii + 304 pp. LCCN Q143.M77 H35 1990.

Hall:1991:PEL

[2720] Marie Boas Hall. Promoting Experimental Learning: Experiment and the Royal Society 1660–1727. Cambridge University Press, Cambridge, UK, 1991. ISBN 0-521-40503-3. xiii + 207 pp. LCCN Q41 .H33 1991. URL

http://www.loc.gov/catdir/description/cam024/91027181.html;
 http://www.loc.gov/catdir/samples/cam031/91027181.html;   http:
//www.loc.gov/catdir/toc/cam028/91027181.html.

Hall:1992:INA

[2721] A. Rupert (Alfred Rupert) Hall. Isaac Newton, Adventurer in Thought. Blackwell science biographies. Blackwell, Oxford, UK, 1992. ISBN 0-631-17906-2. xv + 468 pp. LCCN QC16.N7 H35 1992.

Hall:1994:SSH

[2722] A. Rupert (Alfred Rupert) Hall. Science and Society: Historical Essays on the Relations of Science, Technology, and Medicine, volume CS434 of Collected studies series. Variorum, Aldershot, Hampshire, UK, 1994. ISBN 0-86078-400-2. ix + 465 pp. LCCN Q175.5 .H37 1994.

Hall:1996:HMS

[2723] A. Rupert (Alfred Rupert) Hall. Henry More, and the Scientific Revolution. Cambridge science biographies series. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-56223-6. xii + 304 pp. LCCN Q143.M77 H35 1996. URL

http://www.loc.gov/catdir/description/cam028/96176251.html;
 http://www.loc.gov/catdir/toc/cam021/96176251.html.

Hall:1996:INA

[2724] A. Rupert (Alfred Rupert) Hall. Isaac Newton, Adventurer in Thought. Cambridge science biographies series. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-56221-X (hardcover), 0-521-56669-X (paperback). xvi + 468 pp. LCCN QC16.N7 H35 1996. URL

http://www.loc.gov/catdir/description/cam028/96176355.html;
 http://www.loc.gov/catdir/toc/cam021/96176355.html.

Hall:1999:INE

[2725] A. Rupert (Alfred Rupert) Hall. Isaac Newton: Eighteenth Century Perspectives. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1999. ISBN 0-19-850364-4 (hardcover). 215 pp. LCCN QC16.N7 H354

1999.  URL http://www.loc.gov/catdir/enhancements/fy0606/
98026274-d.html;   http://www.loc.gov/catdir/enhancements/fy0606/
98026274-t.html.

Hall:2002:HOS

[2726] Marie Boas Hall. Henry Oldenburg: Shaping the Royal Society. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2002. ISBN 0-19-851053-5. xii + 369 pp. LCCN Q143.O4 H35 2002. URL http://

www.loc.gov/catdir/enhancements/fy0611/2001053128-d.html;
http://www.loc.gov/catdir/enhancements/fy0611/2001053128-t.
html.

Hancox:1994:QCL

[2727] Joy Hancox. The Queen’s Chameleon: the Life of John Byrom, a Study of Conflicting Loyalties. Jonathan Cape, London, UK, 1994. ISBN 0-224-03047-7. xii + 276 + 12 pp. LCCN PR3339.B5 Z69 1994.

Harris:2002:TCL

[2728] Henry Harris. Things Come to Life: Spontaneous Generation Revisited. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2002.

ISBN 0-19-851538-3. viii + 168 pp. LCCN QH325 .H35 2002. URL http:

//www.loc.gov/catdir/enhancements/fy0636/2001054856-d.html;
 http://www.loc.gov/catdir/enhancements/fy0723/2001054856-b.
html;   http://www.loc.gov/catdir/toc/fy033/2001054856.html.

Harrison:1994:VTS

[2729] H. M. (Hilda Mary) Harrison. Voyager in Time and Space: the Life of John Couch Adams, Cambridge Astronomer. The Book Guild, Sussex, England, 1994. ISBN 0-86332-918-7. xii + 282 pp. LCCN QB36.A2 H37 1994.

Hart-Davis:2000:CRP

[2730] Adam Hart-Davis. Chain Reactions: Pioneers of British Science and Technology and the Stories That Link Them. National Portrait Gallery, London, UK, 2000. ISBN 1-85514-291-0 (paperback). 192 pp. LCCN Q141 .H32 2000.

Haycock:2002:WSS

[2731] David Boyd Haycock. William Stukeley: Science, Religion, and Archaeology in Eighteenth-century England. Boydell Press, Woodbridge, Suffolk, UK, 2002. ISBN 0-85115-864-1. xiii + 290 pp. LCCN

DA93.S85 H39 2002. URL http://www.loc.gov/catdir/toc/fy034/
2001056099.html.

Haynes:1994:FSR

[2732] Roslynn D. (Roslynn Doris) Haynes. From Faust to Strangelove: Representations of the Scientist in Western Literature. The Johns Hopkins University Press, Baltimore, MD, USA, 1994. ISBN 0-8018-4801-6 (hardcover), 0-8018-4983-7 (paperback). 417 pp. LCCN PN56.5.S35 H39 1994.

URL http://www.loc.gov/catdir/bios/jhu052/94007023.html;
http://www.loc.gov/catdir/description/jhu051/94007023.html.

Healey:2001:EDI

[2733] Edna Healey. Emma Darwin: the Inspirational Wife of a Genius. Headline, London, UK, 2001. ISBN 0-7472-7579-3. 372 + 16 pp. LCCN

CT788.D25 H43 2001. URL http://www.loc.gov/catdir/toc/fy02/
2002327528.html.

HerriesDavies:2007:WUE

[2734] Gordon L. Herries Davies. Whatever is under the earth: the Geological Society of London 1807–2007. Geological Society, London, UK, 2007. ISBN 1-86239-214-5 (hardcover). xiii + 356 pp. LCCN QE1.G475 H47

Highfield:1993:PLA

[2735] J. R. L. (John Roger Loxdale) Highfield and Paul Carter. The Private Lives of Albert Einstein. Faber, London, UK, 1993. ISBN 0-571-16744-6, 0-571-17170-2. xii + 355 + 8 pp. LCCN QC16.E5 H54 1993.

Hinde:1999:WGP

[2736] Robert A. Hinde. Why Gods Persist: a Scientific Approach to Religion. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1999. ISBN 0-415-20825-4 (hardcover), 0-415-20826-2 (paperback), 0-203-15872-5 (e-book). vi + 288 pp. LCCN BL48 .H475 1999. URL http:

//www.loc.gov/catdir/enhancements/fy0650/98049646-d.html.

Hinde:2002:WGG

[2737] Robert A. Hinde. Why Good Is Good: the Sources of Morality. Routledge, London, UK, 2002. ISBN 0-415-27752-3 (hardcover), 0-415-27753-1 (paperback). xiv + 241 pp. LCCN BJ1012 .H54 2002. URL http://

www.loc.gov/catdir/enhancements/fy0650/2001048669-d.html;
http://www.loc.gov/catdir/enhancements/fy1102/2001048669-b.
html.

Hodgkin:1992:CDR

[2738] A. L. (Alan Lloyd) Hodgkin. Chance and Design: Reminiscences of Science in Peace and War. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-40099-6. xi + 412 pp. LCCN QP355.2 .H63 1992.

URL http://www.loc.gov/catdir/description/cam024/91027737.
html;   http://www.loc.gov/catdir/toc/cam028/91027737.html.

Hofmann:1996:AMA

[2739] James R. Hofmann. André-Marie Ampère. Cambridge science biographies series. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-56220-1. xiv + 406 pp. LCCN QC16.A47 H64 1996. URL

http://www.loc.gov/catdir/description/cam028/96176376.html;
 http://www.loc.gov/catdir/toc/cam021/96176376.html.

Holbrook:1992:SPD

[2740] Mary Holbrook, R. G. W. (Robert Geoffrey William) Anderson, and D. J. Bryden. Science Preserved: a Directory of Scientific Instruments in Collections in the United Kingdom and Eire. Her Majesty’s Stationary Office, London, England, 1992. ISBN 0-11-290060-7. 271 pp. LCCN Q185.7 .H65 1992.

Hollingdale:1989:MM

[2741] Stuart H. Hollingdale. Makers of Mathematics. Penguin Books, Harmondsworth, UK, 1989. ISBN 0-14-022732-6. xiii + 433 pp. LCCN QA21 .H74 1989.

Hoskin:2003:HPV

[2742] Michael A. Hoskin. The Herschel Partnership: as Viewed by Caroline. Science History Publications, Cambridge, UK, 2003. ISBN 0-905193-05-9. vii + 182 pp. LCCN QB36.H6 H69 2003.

Howse:1989:NMS

[2743] Derek Howse. Nevil Maskelyne, the Seaman’s Astronomer. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-36261-X. xiv + 280 pp. LCCN QB36.M378 H69 1989.

Hoyle:1994:HWW

[2744] Fred Hoyle. Home is Where the Wind Blows: Chapters from a Cosmologist’s Life. University Science Books, Mill Valley, CA, USA, 1994. ISBN 0-935702-27-X (hardcover). xi + 443 + 24 pp. LCCN QB460.72.H69 A3 1994.

Huddart:1989:UWA

[2745] William Huddart. Unpathed Waters: Account of the Life and Times of Joseph Huddart, FRS, at One Time a Captain in the Service of the Honorable East India Company and Later an Elder Brother of the Corporation of Trinity House. Quiller Press, London, UK, 1989. ISBN 1-870948-36-X. viii + 200 + 4 pp. LCCN VK140.H83 H83 1989.

Hughes:1991:TWH

[2746] J. Trevor (John Trevor) Hughes. Thomas Willis 1621–1675: His Life and Work. Royal Society of Medicine Services, London, UK and New York, NY, USA, 1991. ISBN 1-85315-162-9 (hardcover). xvi + 151 pp. LCCN R489.W565 H845 1991.

Hunter:1989:ENS

[2747] Michael Cyril William Hunter. Establishing the New Science: the Experience of the Early Royal Society. Boydell Press, Woodbridge, Suffolk, UK, 1989. ISBN 0-85115-506-5. xiv + 382 pp. LCCN Q41.L85 H93 1989.

Hunter:1994:RBH

[2748] Michael Cyril William Hunter and William Wotton. Robert Boyle by Himself and His Friends, with a Fragment of William Wotton’s Lost Life of Boyle. W. Pickering, London, UK, 1994. ISBN 1-85196-085-6. cvii + 188 pp. LCCN Q143.B77 A3 1994.

Hunter:1994:RSF

[2749] Michael (Michael Cyril William) Hunter. The Royal Society and Its Fellows 1660–1700: the Morphology of an Early Scientific Institution, volume 4 of BSHS monographs. British Society for the History of Science, Chalfont St. Giles, UK, second edition, 1994. ISBN 0-906450-09-8. ix + 291 pp. LCCN Q41.L85 H95 1994.

Hunter:1995:SSO

[2750] Michael Cyril William Hunter. Science and the Shape of Orthodoxy: Intellectual Change in Late Seventeenth-Century Britain. Boydell Press, Woodbridge, Suffolk, UK, 1995. ISBN 0-85115-594-4. xii + 345 pp. LCCN Q127.G4 H853 1995.

Inwood:2002:MWK

[2751] Stephen Inwood. The Man Who Knew Too Much: the Strange and Inventive Life of Robert Hooke, 1635–1703. MacMillan Publishing Company, New York, NY, USA, 2002. ISBN 0-333-78286-0. xxix + 497 + 16 pp. LCCN Q143.H7 I58 2002.

Jardine:2002:GSO

[2752] Lisa Jardine. On a Grander Scale: the Outstanding Career of Sir Christopher Wren. HarperCollins, London, UK, 2002. ISBN 0-06-019974-1, 0-00-710775-7 (hardcover). xviii + 600 + 16 pp. LCCN NA997.W8 J37 2002b.

Jardine:2004:CLR

[2753] Lisa Jardine. The Curious Life of Robert Hooke: the Man Who Measured London. HarperCollins, London, UK, 2004. ISBN 0-06-053897-X. ix + 422 pp. LCCN Q143.H7 J37 2004. URL http://www.h-net.org/

review/hrev-a0e0o2-aa.

Jensen:1991:THH

[2754] J. Vernon (John Vernon) Jensen. Thomas Henry Huxley: Communicating for Science. University of Delaware Press, Newark, DE, USA, 1991. ISBN 0-87413-379-3. 253 pp. LCCN QH31.H9 J46 1991.

Johnson:2001:CWB

[2755] W. (William) Johnson. Collected Works on Benjamin Robins and Charles Hutton. Phoenix Publishing House, New Delhi, India, 2001. ISBN 81-7484-054-0. x + 540 pp. LCCN UF57 .J65 2001.

Jordanova:2000:DFS

[2756] Ludmilla J. Jordanova. Defining Features: Scientific and Medical Portraits, 1660–2000. Reaktion Books, London, UK, 2000. ISBN 1-86189-059-1 (paperback). 191 pp. LCCN NX180,S3 J67 2000; R153.5 .J67 2000.

Jungnickel:1996:C

[2757] Christa Jungnickel and Russell McCormmach. Cavendish, volume 220 of Memoirs of the American Philosophical Society held at Philadelphia for Promoting Useful Knowledge. American Philosophical Society, Philadelphia, PA, USA, 1996. ISBN 0-87169-220-1 (hardcover). xi + 414 pp. LCCN Q11 .P612 vol. 220 QD22.C4.

Kanigel:1991:MWK

[2758] Robert Kanigel. The Man Who Knew Infinity — a Life of the Genius Ramanujan. Collier Macmillan Canada, Toronto, Ontario, Canada, 1991. ISBN 0-684-19259-4. ix + 438 pp. LCCN QA29.R3 K36 1991.

Kanigel:1992:MWK

[2759] Robert Kanigel. Man Who Knew Infinity — a Life of the Genius Ramanujan. Abacus, London, UK, 1992. ISBN 0-349-10452-2 (paperback), 0-356-20330-1. ix + 438 + 8 pp. LCCN QA29.R3 K36 1992.

Kass-Simon:1990:WSR

[2760] G. (Gabriele) Kass-Simon, Patricia Farnes, and Deborah Nash, editors. Women of Science: Righting the Record. Indiana University Press, Bloomington, IN, USA, 1990. ISBN 0-253-20813-0 (paperback), 0-253-33264-8. xvi + 398 pp. LCCN Q130 .W67 1990.

King-Hele:1999:EDL

[2761] Desmond George King-Hele. Erasmus Darwin: a life of unequalled achievement. Giles de la Mare, London, UK, 1999. ISBN 1-900357-08-9 (paperback). x + 422 + 16 pp. LCCN QH31.D3 K5 1999. URL http:

//www.loc.gov/catdir/enhancements/fy0740/99488514-b.html;
http://www.loc.gov/catdir/enhancements/fy0740/99488514-d.html.

King:1989:WRS

[2762] M. Christine King. E. W. R. Steacie and Science in Canada. University of Toronto Press, Toronto, ON, Canada, 1989. ISBN 0-8020-2667-2. xii + 243 + 8 pp. LCCN QD22.S431 K56 1989.

Knight:1992:HDS

[2763] David M. Knight. Humphry Davy: Science and Power. Blackwell science biographies. Blackwell, Oxford, UK, 1992. ISBN 0-631-16816-8. xiii + 218 pp. LCCN QD22.D3 K65 1992.

Knight:1996:HDS

[2764] David M. Knight. Humphry Davy: Science and Power. Cambridge science biographies series. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-56309-7 (hardcover). xiv + 218 pp. LCCN QD22.D3

K65 1996. URL http://www.loc.gov/catdir/description/cam028/
96176267.html;   http://www.loc.gov/catdir/toc/cam021/96176267.
html.

Kolb:1996:BWS

[2765] Edward W. (Rocky) Kolb. Blind Watchers of the Sky: the People and Ideas That Shaped Our View of the Universe. Helix books. Addison-Wesley, Reading, MA, USA, 1996. ISBN 0-201-48992-9. xii + 338 pp. LCCN QB981 .K689 1996.

Kolb:1999:BWS

[2766] Edward W. Kolb. Blind Watchers of the Sky: the People and Ideas That Shaped Our View of the Universe. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1999. ISBN 0-19-286203-0. xii + 338 pp. LCCN QB981.

Kragh:1999:QGH

[2767] Helge Kragh. Quantum generations: a history of physics in the twentieth century. Princeton University Press, Princeton, NJ, USA, 1999. ISBN 0-691-01206-7 (hardcover), 0-691-09552-3 (paperback). xiv + 494 pp.

LCCN QC7 .K7 2002. URL http://press.princeton.edu/titles/
6683.html.

Kuhn:1988:SCC

[2768] Dieter Kuhn and Joseph Needham, editors. Science and civilisation in China. Vol. 5. Chemistry and chemical technology. Pt. 9, Textile technology: Spinning and reeling. Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-32021-6. xxxiv + 520 pp. LCCN DS721 .N3913 1975; Q127.C5 T27.C5; TS1401 K83 1988.

Lagerkvist:2003:PMN

[2769] Ulf Lagerkvist. Pioneers of Microbiology and the Nobel Prize. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2003. ISBN 1-281-92824-0, 981-238-233-X (hardcover), 981-238-234-8 (paperback), 981-277-556-0 (e-book). ix + 178

pp. LCCN QR21 .L34 2003. URL https://www.worldscientific.
com/worldscibooks/10.1142/5161.

Lax:2004:MDF

[2770] Eric Lax. The Mould in Dr Florey’s Coat: the Remarkable True Story of the Penicillin Miracle. Little, Brown and Co., Boston, MA, USA, 2004. ISBN 0-316-85925-7. 389 + 8 pp. LCCN RM666.P35 L392 2004. URL

http://www.loc.gov/catdir/toc/fy051/2004426869.html.

LeFanu:1990:NGM

[2771] William LeFanu. Nehemiah Grew M.D., F.R.S.: a Study and Bibliography of His Writings. St Paul’s Bibliographies, Winchester, UK, 1990. ISBN 0-906795-43-5. xvii + 182 pp. LCCN QK31.G819 Z994 1990; Z8369.262 .L4 1990.

Lightman:2016:CHS

[2772] Bernard V. Lightman, editor. A Companion to the History of Science. Blackwell companions to world history. John Wiley, New York, NY, USA, 2016. ISBN 1-118-62077-1 (hardcover), 1-118-62074-7. xv + 601 pp. LCCN Q124.7 .C65 2016.

Lopez:1988:PBF

[2773] Claude A. Lopez. The Papers of Benjamin Franklin. July 1 Through October 31, 1778, volume 27 of The papers of Benjamin Franklin. Yale University Press, New Haven, CT, USA, 1988. ISBN 0-300-04177-2. lxix + 727 pp. LCCN E302 .F82 1959.

Lovell:1990:AC

[2774] Sir Bernard Lovell. Astronomer by Chance. MacMillan Publishing Company, New York, NY, USA, 1990. ISBN 0-333-55195-8. xi + 381 + 8 pp. LCCN QB36 L68 1991.

Lovell:1991:EWS

[2775] Bernard Lovell. Echoes of War: the Story of H_{2}S Radar. Adam Hilger Ltd., Bristol, UK, 1991. ISBN 0-85274-317-3. xxi + 287 pp. LCCN D810.R33 L68 1991.

Lowe:1985:ANW

[2776] Victor Lowe. Alfred North Whitehead: the Man and His Work (1861– 1910), volume 1. The Johns Hopkins University Press, Baltimore, MD, USA, 1985. ISBN 0-8018-2488-5. xi + 351 pp. LCCN B1674.W354 L57

Lowe:1990:ANW

[2777] Victor Lowe and J. B. Schneewind. Alfred North Whitehead: the Man and His Work (1910–1947), volume 2. The Johns Hopkins University Press, Baltimore, MD, USA, 1990. ISBN 0-8018-3960-2. xi + 389 pp. LCCN B1674.W354 L57 1985.

Lyell:1997:PG

[2778] Sir Charles Lyell and James A. Secord. Principles of Geology. Penguin classics. Penguin Books, Harmondsworth, UK, 1997. ISBN 0-14-043528-X. xlvii + 471 pp. LCCN QE26.2 .L94 1997.

MacCallum:1999:AMH

[2779] R. Ian MacCallum. Antimony in Medical History: an Account of the Medical Uses of Antimony and Its Compounds Since Early Times to the Present. Pentland Press, Edinburgh, UK, 1999. ISBN 1-85821-642-7. xvi + 125 pp. LCCN RM666.A55 M43 1999. URL http://www.gbv.de/

dms/bs/toc/265905265.pdf.

Macfarlane:2002:GB

[2780] Alan Macfarlane and Gerry Martin. The Glass Bathyscaphe. Profile, London, UK, 2002. ISBN 1-86197-400-0. xi + 255 pp. LCCN TP849 .M34 2002.

MacLeod:2000:CSV

[2781] Roy M. MacLeod. The “Creed of Science” in Victorian England, volume CS598 of Variorum collected studies series. Ashgate/Variorum, Aldershot, Hampshire, England, 2000. ISBN 0-86078-669-2. xxiv + 318 pp. LCCN Q127.G7 M33 2000.

Maddox:2002:RFD

[2782] Brenda Maddox. Rosalind Franklin: the Dark Lady of DNA. Harper-Collins, London, UK, 2002. ISBN 0-00-257149-8. 304 + 16 pp. LCCN QH506.F72 M33 2002; QP26.F68 M33 2002.

Mancosu:1996:PMM

[2783] Paolo Mancosu. Philosophy of Mathematics and Mathematical Practice in the Seventeenth Century. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1996. ISBN 0-19-508463-2. viii + 275 pp. LCCN

QA8.4 .M36 1996. URL http://www.loc.gov/catdir/enhancements/
fy0604/94047024-d.html.

Mariscotti:1996:SAH

[2784] Mario Mariscotti. El Secreto Atómico de Huemul. (Spanish) [The Atomic Secret of Huemul]. Estudio Sigma, Buenos Aires, Argentina, third edition, 1996. ISBN 950-9446-24-6. 286 pp. LCCN QC773.3.A7 M37 1996.

Martin:1992:FBS

[2785] Julian Martin. Francis Bacon, the State and the Reform of Natural Philosophy. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-38249-1 (hardcover). xiii + 236 pp. LCCN B1198 .M36 1992.

URL http://www.loc.gov/catdir/description/cam024/90024242.
html;   http://www.loc.gov/catdir/toc/cam028/90024242.html.

Mason:1992:GEB

[2786] Adair Stuart Mason. George Edwards: the bedell and his birds. Royal College of Physicians, London, UK, 1992. ISBN 1-873240-48-1. viii + 71 pp. LCCN QL31.E39 M37 1992.

McConnell:1992:IMW

[2787] Anita McConnell. Instrument Makers to the World: a History of Cooke, Troughton and Simms. University of York, York, UK, 1992. ISBN 1-85072-096-7. 116 pp. LCCN HD9706.6 .M12 1992.

Medawar:1990:VDP

[2788] J. S. (Jean Shinglewood) Medawar. A Very Decided Preference: Life with Peter Medawar. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1990. ISBN 0-19-217779-6. 256 pp. LCCN QR180.72.M43 M43 1990b.

Medawar:1999:HGS

[2789] Jean S. Medawar and David Pyke. Hitler’s Gift: Scientists Who Fled Nazi Germany. Richard Cohen Books and the European Jewish Publication Society, London, UK, 1999. ISBN 1-86066-172-6. xx + 268 pp. LCCN Q141 .M385 1999.

Mehra:1994:BDD

[2790] Jagdish Mehra. The Beat of a Different Drum: the Life and Science of Richard Feynman. Clarendon Press, Oxford, UK, 1994. ISBN 0-19-853948-7. xxxii + 630 + 16 pp. LCCN QC16.F49 M45 1994.

Mennim:1992:TCS

[2791] Eleanor Mennim. Transit Circle: the Story of William Simms, 1793 to 1860. W. Sessions, York, UK, 1992. ISBN 1-85072-101-7. xiv + 302 pp. LCCN QB84.75.S56 M46 1992.

Millar:1996:CDS

[2792] David Millar, Ian Millar, John Millar, and Margaret Millar. The Cambridge Dictionary of Scientists. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-56185-X (hardcover), 0-521-56718-1 (paperback). xii + 387 pp. LCCN Q141 .C128 1996. URL

http://www.loc.gov/catdir/description/cam027/95038471.html;
 http://www.loc.gov/catdir/toc/cam021/95038471.html.

Moll:1996:PRS

[2793] J. M. H. (John Michael Henderson) Moll. Presidents of the Royal Society of Medicine: Illustrated Profiles, 1805–1996: Dr. William Saunders to Sir Donald Harrison. Royal Society of Medicine Press, London, UK, 1996. ISBN 1-85315-264-1 (cased edition), 1-85315-265-X (halfleather limited edition). xlii + 552 pp. LCCN R489.A1 M65 1996.

Moore:1989:SLT

[2794] Walter John Moore. Schrödinger: Life and Thought. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-35434-X. xi + 513 pp. LCCN QC16.S265 M66 1989. URL

http://www.loc.gov/catdir/description/cam023/88025807.html;
 http://www.loc.gov/catdir/toc/cam021/88025807.html.

Morrell:1997:SOT

[2795] Jack Morrell. Science at Oxford, 1914–1939: Transforming an Arts University. Clarendon Press, Oxford, UK, 1997. ISBN 0-19-820657-7, 0-19-167722-1. xvii + 473 pp. LCCN Q183.4.G73 O946 1997. URL http:

//www.loc.gov/catdir/enhancements/fy0606/97007976-d.html;
http://www.loc.gov/catdir/enhancements/fy0726/97007976-b.html.

Morton:1993:PPS

[2796] Alan Q. Morton and Jane A. Wess. Public and Private Science: the King George III Collection. Oxford University Press in association with Science Museum, Oxford, UK, 1993. ISBN 0-19-856392-2. ix + 710 pp. LCCN Q105.G72 O946 1993; Q185.7 S342 1993.

Needham:1954:SCC

[2797] Joseph Needham and Ling Wang. Science and civilisation in China. University Press, Cambridge, UK, 1954–2008. ISBN 0-521-05799-X (vol. 1), 0-521-05800-7 (vol. 2), 0-521-05801-5 (vol. 3), 0-521-07060-0 (vol. 4, pt 3), 0-521-32727-X (vol. 5, pt 6), (vol. 5, pt. 11), 0-521-65270-7 (vol. 6, pt 5), 0-521-63262-5 (vol. 6, pt 6), 0-521-57143-X (vol. 7, pt 1). xxxviii + 319 (vol. 1), xxii + 697 (vol. 2), xlvii + 886 (vol. 3), lvii + 931 (vol. 4, pt 3), xxviii + 601 (vol. 5, pt 6), xxviii + 741 (vol. 6, pt 5), xviii + 261 (vol. 6, pt 6), xxiv + 479 (vol. 7, pt 1) pp. LCCN DS721 .N39. URL

http://www.loc.gov/catdir/description/cam022/54004723.html;
 http://www.loc.gov/catdir/samples/cam031/54004723.html;   http:
//www.loc.gov/catdir/toc/cam022/54004723.html.

Needham:1984:SCC

[2798] Joseph Needham, Gwei-Djen Lu, H. T. Huang, Francesca Bray, Christian Daniels, Nicholas K. Menzies, and Nathan Sivin. Science and civilisation in China. Vol. 6, Biology and biological technology. Cambridge University Press, Cambridge, UK, 1984. ISBN 0-521-08731-7 (pt. 1), 0-521-25076-5 (pt. 2), 0-521-41999-9 (pt. 3), 0-521-65270-7 (pt. 5), 0-521-63262-5 (pt. 6). ???? pp. LCCN HD2098.NEE.

Needham:1986:CCT

[2799] Joseph Needham, Ho Ping-Yu, Lu Gwei-Djen, and Wang Ling. Chemistry and chemical technology. Part 7, Military technology: the gunpowder epic, volume 5, pt. 7 of Science and civilisation in China. Cambridge University Press, Cambridge, UK, 1986. ISBN 0-521-30358-3. xxxiii + 703 pp. LCCN DS721 .N39 vol. 5 pt. 7.

Needham:1986:SSC

[2800] Joseph Needham and C. A. (Colin Alistair) Ronan. The shorter science and civilisation in China: an abridgement of Joseph Needham’s original text. Cambridge University Press, Cambridge, UK, 1986. ISBN 0-521-25272-5, 0-521-31560-3 (paperback). ix + 298 pp. LCCN DS721 .N392; Q127.C5 N32.

Needham:1987:SCC

[2801] Joseph Needham, Tsuen hsuin Tsien, Ling Ed Wang, and Gwei Djen Lu. Science and civilisation in China. 05, 01, Chemistry and chemical technology, Paper and printing. Cambridge University Press, Cambridge, UK, third edition, 1987. ISBN 0-521-08690-6. xxv + 485 pp. LCCN DS721 .N3913 1975; Q127.C5; T27.C5; TS1083.C6 T8 1987.

Newton:1989:PMI

[2802] Sir Isaac Newton. The preliminary manuscripts for Isaac Newton’s 1687 Principia, 1684–1685, volume 2 of Newton manuscripts series. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-33499-3. xxi + 246 pp. LCCN QA803 .N413 1989.

Nichols:1994:DRH

[2803] Richard Nichols. The Diaries of Robert Hooke, the Leonardo of London, 1635–1703. Book Guild, Lewes, Sussex, UK, 1994. ISBN 0-86332-930-6. 185 pp. LCCN Q143.H7 N53 1994.

OBrian:1987:JBL

[2804] Patrick O’Brian. Joseph Banks: a Life. Collins Harvill, London, UK, 1987. ISBN 0-00-217350-6, 0-00-272340-9. 328 + 8 pp. LCCN Q143.B3; QH31.B19 O27 1987.

OBrian:1993:JBL

[2805] Patrick O’Brian. Joseph Banks: a Life. D. R. Godine, Boston, MA, USA, 1993. ISBN 0-87923-930-1. 328 + 8 pp. LCCN QH31.B19 O27 1993.

OBrian:1997:JBL

[2806] Patrick O’Brian. Joseph Banks: a Life. University of Chicago Press, Chicago, IL, USA, 1997. ISBN 0-226-61628-2. 328 + 6 pp. LCCN QH31.B19 O27 1997. URL http://www.loc.gov/

catdir/description/uchi051/97024475.html;   http://www.loc.
gov/catdir/toc/uchi052/97024475.html.

Ogilvie:1986:WSA

[2807] Marilyn Bailey Ogilvie. Women in Science: Antiquity Through the Nineteenth Century: a Biographical Dictionary with Annotated Bibliography. MIT Press, Cambridge, MA, USA, 1986. ISBN 0-262-15031-X (hardcover), 0-262-65038-X (paperback). xi + 254 pp. LCCN Q141 .O34 1986.

Ollard:1991:PB

[2808] Richard Lawrence Ollard. Pepys: a Biography. Sinclair-Stevenson, London, UK, 1991. ISBN 1-85619-066-8. xii + 411 + 16 pp. LCCN PR3618.P2 Z75 1991.

Olson:1998:FSC

[2809] Roberta J. M. Olson and Jay M. Pasachoff. Fire in the Sky: Comets and Meteors, the Decisive Centuries, in Art and Science. Cambridge University Press, Cambridge, UK, 1998. ISBN 0-521-63060-6 (hardcover), 0-521-66359-8 (paperback). xiv + 369 + 32 pp. LCCN N6766 .O45 1998.

Osborne:1998:FEE

[2810] Roger Osborne. The Floating Egg: Episodes in the Making of Geology. Jonathan Cape, London, UK, 1998. ISBN 0-224-05028-1. xii + 372 pp. LCCN QE13.G7 O83 1998.

Osler:2000:RSR

[2811] Margaret J. Osler, editor. Rethinking the Scientific Revolution. Cambridge University Press, Cambridge, UK, 2000. ISBN 0-521-66101-3 (hardcover), 0-521-66790-9 (paperback). xii + 340 pp. LCCN Q125.2

.R48 2000. URL http://www.loc.gov/catdir/description/cam029/
99021576.html;   http://www.loc.gov/catdir/toc/cam025/99021576.
html.

Owen:1992:HLC

[2812] Sir Richard Owen and Phillip R. Sloan. The Hunterian Lectures in Comparative Anatomy: May–June, 1837. Natural History Museum publications. Natural History Museum, London, UK, 1992. ISBN 0-565-01106-5 (hardcover), 0-565-01144-8 (paperback). xvi + 340 pp. LCCN M93.E02368; M93.E02370; QL808 OWE 1992.

Pais:1991:NBT

[2813] Abraham Pais. Niels Bohr’s Times: in Physics, Philosophy, and Polity. Clarendon Press, Oxford, UK, 1991. ISBN 0-19-852048-4 (paperback), 0-19-852049-2. xvii + 565 pp. LCCN QC773 .P35 1991. URL http://

www.loc.gov/catdir/enhancements/fy0635/90027248-d.html.

Parker:1992:MLE

[2814] John (John R.) Parker and Basil Harley. Martin Lister’s English Spiders: 1678. Harley Books, Colchester, Essex, England, 1992. ISBN 0-946589-41-0 (paperpack edition), 0-946589-27-5 (hardback edition). xv + 208 pp. LCCN QL453.4.G7 L5713 1992.

Parkin:1999:SRP

[2815] Jon (Jonathan Bruce) Parkin. Science, Religion, and Politics in Restoration England: Richard Cumberland’s De Legibus Naturae. Royal Historical Society studies in history. New series. Royal Historical Society/Boydell Press, Woodbridge, Suffolk, UK, 1999. ISBN 0-86193-241-2. ISSN 0269-2244. xi + 251 pp. LCCN B1201.C83 D424 1999.

Perutz:1998:WIM

[2816] Max F. (Max Ferdinand) Perutz. I Wish I’d Made You Angry Earlier: Essays on Science, Scientists, and Humanity. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1998. ISBN 0-19-850531-0 (hardcover), 0-19-859027-X (paperback). xv + 354 + 10 pp. LCCN Q175 .P387 1998.

Phillips:1990:SLS

[2817] Patricia Phillips. The Scientific Lady: a Social History of Women’s Scientific Interests 1520–1918. Weidenfeld and Nicolson, London, UK, 1990. ISBN 0-297-82043-5, 0-312-05685-0. xiii + 279 + 8 pp. LCCN Q130 .P53 1990.

Polkinghorne:1996:BSW

[2818] J. C. (John Charlton) Polkinghorne. Beyond Science: the Wider Human Context. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-57212-6 (hardcover). xii + 131 pp. LCCN Q175 .P834 1996.

URL http://www.loc.gov/catdir/description/cam027/96013177.
html;   http://www.loc.gov/catdir/toc/cam021/96013177.html.

Popham:1991:DCF

[2819] Hugh Popham. A Damned Cunning Fellow: the Eventful Life of Rear-Admiral Sir Home Popham, KCB, KCH, KM, FRS, 1762–1820. Old Ferry Press, Tywardreath, Cornwall, UK, 1991. ISBN 0-9516758-0-X. xiii + 256 pp. LCCN DA88.1.P6 P67 1991.

Porter:2000:EBC

[2820] Roy Porter. Enlightenment: Britain and the Creation of the Modern World. Allen Lane/Penguin Press, London, UK; New York, NY, USA, 2000. ISBN 0-7139-9152-6. xxiv + 727 + 16 pp. LCCN DA485 .P68 2000.

Porter:2004:KPS

[2821] Theodore M. Porter. Karl Pearson: the scientific life in a statistical age. Princeton University Press, Princeton, NJ, USA, 2004. ISBN 0-691-11445-5. viii + 342 pp. LCCN QA276.P64 P67 2004. URL http:

//www.loc.gov/catdir/description/prin051/2003054820.html;
http://www.loc.gov/catdir/toc/prin051/2003054820.html.

Principe:1998:AAR

[2822] Lawrence Principe. The aspiring adept: Robert Boyle and his alchemical quest: including Boyle’s “lost” Dialogue on the transmutation of metals. Princeton University Press, Princeton, NJ, USA, 1998. ISBN 0-691-01678-X (hardcover). xiv + 339 pp. LCCN

QD24.B685 P75 1998. URL http://www.h-net.org/review/hrev-
a0a8s4-aa;   http://www.loc.gov/catdir/description/prin021/

97041793.html;   http://www.loc.gov/catdir/toc/prin031/97041793.
html.

Reilly:1992:JW

[2823] Robin Reilly. Josiah Wedgwood 1730–1795. MacMillan Publishing Company, New York, NY, USA, 1992. ISBN 0-333-51041-0. xii + 412 + 16 pp. LCCN NK4210.W4 R4 1992.

Rice:1999:VDT

[2824] A. L. (Anthony Leonard) Rice. Voyages of Discovery: Three Centuries of Natural History Exploration. Scriptum Editions, London, UK, 1999. ISBN 1-902686-02-0. 335 pp. LCCN QH15 .R44 1999.

Richards:1985:DJR

[2825] Robert L. Richards. Dr. John Rae. Caedmon of Whitby Publishers, Whitby, North Yorkshire, UK, 1985. ISBN 0-905355-29-6. xii + 231 + 16 pp. LCCN RD273.R34 R53 1985.

Roberts:1989:SJP

[2826] Shirley Roberts. Sir James Paget: the Rise of Clinical Surgery. Eponymists in medicine. Royal Society of Medicine Services, London, UK and New York, NY, USA, 1989. ISBN 0-905958-99-3 (hardcover), 0-905958-91-8 (paperback). ISSN 0955-8640. 223 pp. LCCN

RA489.R62 R62 1989. URL http://catalog.hathitrust.org/api/
volumes/oclc/22547173.html.

Rodger:1993:IEL

[2827] N. A. M. Rodger. The Insatiable Earl: a Life of John Montagu, Fourth Earl of Sandwich, 1718–1792. HarperCollins, London, UK, 1993. ISBN 0-00-215784-5. xviii + 425 + 8 pp. LCCN DA506.S3 R63 1993.

Rowlands:1990:OLL

[2828] Peter Rowlands. Oliver Lodge and the Liverpool Physical Society, volume 4 of Liverpool historical studies. Published for the Department of History, University of Liverpool [by] Liverpool University Press, Liverpool, UK, 1990. ISBN 0-85323-027-7 (paperback). iv + 310 pp. LCCN QC16.L6 R69 1990.

Ruestow:1996:MDR

[2829] Edward G. (Edward Grant) Ruestow. The Microscope in the Dutch Republic: the Shaping of Discovery. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-47078-1 (hardcover). xii + 348 pp. LCCN

QH204 .R84 1996. URL http://www.loc.gov/catdir/description/
cam027/95045548.html;   http://www.loc.gov/catdir/samples/cam031/
95045548.html;   http://www.loc.gov/catdir/toc/cam026/95045548.
html.

Rupke:1994:ROV

[2830] Nicolaas A. Rupke. Richard Owen: Victorian Naturalist. Yale University Press, New Haven, CT, USA, 1994. ISBN 0-300-05820-9. xvii + 462 pp. LCCN QH31.O94 R86 1994.

Russell:1986:LCE

[2831] Colin Archibald Russell. Lancastrian chemist: the early years of Sir Edward Frankland. Open University Press, Milton Keynes, England, 1986. ISBN 0-335-15175-2. ix + 187 pp. LCCN QD22.F65 R87 1986.

Russell:1996:EFC

[2832] Colin Archibald Russell. Edward Frankland: chemistry, controversy, and conspiracy in Victorian England. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-49636-5 (hardcover). xx + 535 pp. LCCN QD22.F65 R86 1996. URL

http://www.loc.gov/catdir/description/cam027/95040319.html;
 http://www.loc.gov/catdir/toc/cam021/95040319.html.

Sabbagh:1999:RAH

[2833] Karl Sabbagh. A rum affair: how botany’s “Piltdown Man” was unmasked. Allen Lane, London, UK, 1999. ISBN 0-7139-9277-8. 224 + 8 pp. LCCN QK31.H26 S33 1999.

Said:1990:ESF

[2834] Hakim Mohammad Said, editor. Essays on Science: Felicitation Volume in Honour of Dr. Joseph Needham. Hamdard Foundation Pakistan, Karachi, Pakistan, 1990. xiii + 267 pp. LCCN Q126.8 .E78 1990.

Schaaf:1992:SHT

[2835] Larry J. (Larry John) Schaaf. Out of the Shadows: Herschel, Talbot and the Invention of Photography. Yale University Press, New Haven, CT, USA, 1992. ISBN 0-300-05705-9 (hardcover). xii + 188 pp. LCCN TR57 .S32 1992.

Schaaf:1996:RDP

[2836] Larry J. (Larry John) Schaaf. Records of the Dawn of Photography: Talbot’s Notebooks P and Q. Cambridge University Press in cooperation with the National Museum of Photography, Film and Television, Cambridge, UK, 1996. ISBN 0-521-44051-3 (hardcover). xxxv + 413 pp. LCCN TR144 .S35 1996. URL

http://www.loc.gov/catdir/description/cam027/95010540.html;
 http://www.loc.gov/catdir/toc/cam021/95010540.html.

Schiebinger:1989:MNS

[2837] Londa L. Schiebinger. The Mind Has No Sex?: Women in the Origins of Modern Science. Harvard University Press, Cambridge, MA, USA, 1989. ISBN 0-674-57623-3 (hardcover), 0-674-57625-X (paperback). xi + 355 pp. LCCN Q130 .S32 1989.

Sexton:1991:STL

[2838] Christopher Sexton. The Seeds of Time: a Life of Sir MacFarlane Burnet. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1991. ISBN 0-19-553274-0. x + 301 pp. LCCN QH31.B92 S49 1991; QR180.72.B87 S48 1991.

Siegfried:2002:EAH

[2839] Robert Siegfried. From Elements to Atoms: a History of Chemical Composition, volume 92 (part 4) of Transactions of the American Philosophical Society, 0065-9746. American Philosophical Society, Philadelphia, PA, USA, 2002. ISBN 0-87169-924-9 (paperback), 0-87169-931-1 (paperback). x + 278 pp. LCCN QD461 .S463 2002.

Smith:1989:EEB

[2840] Crosbie Smith and M. Norton Wise. Energy and Empire: a Biographical Study of Lord Kelvin. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-26173-2. xxvi + 866 pp. LCCN QC16.K3 S65 1989.

URL http://www.loc.gov/catdir/description/cam041/88025685.
html;   http://www.loc.gov/catdir/toc/cam041/88025685.html.

Smith:1998:SEC

[2841] Crosbie Smith. The Science of Energy: a Cultural History of Energy Physics in Victorian Britain. Athlone, London, UK, 1998. ISBN 0-485-11431-3 (hardcover), 0-485-12145-X (paperback). xi + 404 pp. LCCN QC72 .S65 1998; TJ 163.9 .S65 1998.

Smith:1999:EMM

[2842] James R. (James Raymond) Smith. Everest: the Man and the Mountain. Whittles Publishing, Dunbeath Mill, Dunbeath KW6 6EG, UK, 1999. ISBN 1-870325-72-9. xiv + 306 pp. LCCN TA533.E94 S63 1999.

Smithies:1997:CCC

[2843] F. (Frank) Smithies. Cauchy and the creation of complex function theory. Cambridge University Press, Cambridge, UK, 1997. ISBN 0-521-59278-X (hardcover). 0-511-55169-X (e-book). 216 pp. LCCN QA331.7 .S63 1997.

URL http://www.loc.gov/catdir/description/cam028/97003028.
html;   http://www.loc.gov/catdir/toc/cam027/97003028.html.

Soo:1998:WTA

[2844] Lydia M. Soo. Wren’s “Tracts” on Architecture and Other Writings. Cambridge University Press, Cambridge, UK, 1998. ISBN 0-521-57369-6 (hardcover). xv + 320 pp. LCCN NA997.W8 S66 1998.

URL http://www.loc.gov/catdir/description/cam028/97013148.
html;   http://www.loc.gov/catdir/toc/cam025/97013148.html.

Stafford:1989:SES

[2845] Robert A. Stafford. Scientist of Empire: Sir Roderick Murchison, Scientific Exploration and Victorian Imperialism. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-33537-X. xii + 293 pp. LCCN

QE22.M8 S82 1989. URL http://www.loc.gov/catdir/description/
cam023/88034134.html;   http://www.loc.gov/catdir/samples/cam031/
88034134.html;   http://www.loc.gov/catdir/toc/cam027/88034134.
html.

Stedall:2003:GIA

[2846] Jacqueline A. Stedall and Thomas Harriot. The Greate Invention of Algebra: Thomas Harriot’s Treatise on Equations. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2003. ISBN 0-19-852602-4. xi + 322 pp. LCCN QA211 .S77 2003. URL http://www.loc.gov/

catdir/enhancements/fy0613/2002042558-d.html;   http://www.
loc.gov/catdir/toc/fy042/2002042558.html.

Stewart:1992:RPS

[2847] Larry Stewart. The Rise of Public Science: Rhetoric, Technology, and Natural Philosophy in Newtonian Britain, 1660–1750. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-41700-7 (hardcover). xxxiv + 453 pp. LCCN Q127.G4 S84 1992. URL

http://www.loc.gov/catdir/description/cam024/91044172.html;
 http://www.loc.gov/catdir/toc/cam029/91044172.html.

Stoye:1994:MEL

[2848] John Stoye. Marsigli’s Europe, 1680–1730: the Life and Times of Luigi Ferdinando Marsigli, Soldier and Virtuoso. Yale University Press, New Haven, CT, USA, 1994. ISBN 0-300-05542-0. xii + 356 pp. LCCN D274.M37 S88 1994.

Swade:2000:CBC

[2849] Doron Swade. The Cogwheel Brain: Charles Babbage and the Quest to Build the First Computer. Little, Brown and Co., Boston, MA, USA, 2000. ISBN 0-316-64847-7. x + 342 + 16 pp. LCCN QA75

.S93 2000. URL http://hdl.library.upenn.edu/1017.12/366197;
http://www.gbv.de/dms/bsz/toc/bsz086755331inh.pdf.

Swann:1999:JDB

[2850] Brenda Swann and Francis Aprahamian. J. D. Bernal: a life in science and politics. Verso, London, UK, 1999. ISBN 1-85984-854-0 (hardcover). xxv + 324 + 8 pp. LCCN Q143.B47 J2 1999.

Tashjian:1990:RRT

[2851] Georgian Rawlinson Tashjian, David R. Tashjian, and Brian J. Enright. Richard Rawlinson: a Tercentenary Memorial. New Issues Press, Western Michigan University, Kalamazoo, MI, USA, 1990. ISBN 0-932826-23-7. xviii + 221 pp. LCCN DA93.R38 T37 1990.

Thomas:1990:LSL

[2852] J. M. (John Meurig) Thomas and D. (David) Phillips, editors. The Legacy of Sir Lawrence Bragg: selections and reflections. Science Reviews Ltd., Northwood, Middlesex, UK, 1990. ISBN 0-905927-43-5. x + 308 pp. LCCN QC16.B66 L44 1990.

Thornes:1999:JCS

[2853] John E. Thornes. John Constable’s Skies: a Fusion of Art and Science. University of Birmingham, University Press, Edgbaston, Birmingham, UK, 1999. ISBN 1-902459-02-4. 288 pp. LCCN ND497.C7 T48 1999.

Tobin:2003:LSL

[2854] W. (William) Tobin. The Life and Science of Léon Foucault: the Man Who Proved the Earth Rotates. Cambridge University Press, Cambridge, UK, 2003. ISBN 0-521-80855-3. xiv + 338 pp. LCCN QC16.F626 T63 2003. URL http://www.

loc.gov/catdir/description/cam032/2003043577.html;   http:/
/www.loc.gov/catdir/enhancements/fy0731/2003043577-b.html;
http://www.loc.gov/catdir/samples/cam041/2003043577.html;
http://www.loc.gov/catdir/toc/cam032/2003043577.html.

Townes:1999:HLH

[2855] Charles H. Townes. How the Laser Happened: Adventures of a Scientist. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1999. ISBN 0-19-512268-2 (hardcover), 0-19-515376-6 (paperback), 1-280-47085-2. 200 pp. LCCN QC687.2 .T68 1999. URL http://www.

loc.gov/catdir/enhancements/fy0603/98022216-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0603/98022216-t.html;
http://www.loc.gov/catdir/enhancements/fy0724/98022216-b.html.

Twort:1993:FSB

[2856] Antony Twort. In Focus, Out of Step: a Biography of Frederick William Twort F.R.S., 1877–1950. A. Sutton, Phoenix Mill, UK, 1993. ISBN 0-7509-0327-9. xi + 340 pp. LCCN QR31.T86 T86 1993.

Uglow:2002:LMF

[2857] Jennifer S. Uglow. The Lunar Men: the Friends Who Made the Future, 1730–1810. Faber, London, UK, 2002. ISBN 0-571-19647-0, 0-571-21610-2 (paperback). xx + 588 + 16 pp. LCCN T39 .U356 2003.

VanWyhe:2004:POV

[2858] John Van Wyhe. Phrenology and the Origins of Victorian Scientific Naturalism. Science, technology, and culture, 1700–1945. Ashgate, Aldershot, Hampshire, England, 2004. ISBN 0-7546-3408-6. xvii + 282 pp. LCCN BF879 .V36 2004.

Vaughan:1993:IKB

[2859] Adrian Vaughan. Isambard Kingdom Brunel, Engineering Knight-Errant. J. Murray, London, UK, 1993. ISBN 0-7195-5282-6. xiii + 285 pp. LCCN TA140.B75 V38 1993.

Venkataraman:1988:JLL

[2860] G. (Ganesan) Venkataraman. Journey Into Light: Life and Science of C. V. Raman. Published by Indian Academy of Sciences in co-operation with Indian National Science Academy, Bangalore, India, 1988. ISBN 81-85324-00-X. xviii + 570 + 24 pp. LCCN QC16.R27 V46 1988. URL http:

//www.loc.gov/catdir/enhancements/fy0639/89903531-d.html.

Venkataraman:1994:BHM

[2861] G. Venkataraman. Bhabha and His Magnificent Obsessions. Vignettes in physics. Sangam Books, London, UK, 1994. ISBN 0-86311-555-1. ix + 209 pp. LCCN QC16.B43 V46 1994.

Wali:1991:CBC

[2862] K. C. (Kameshwar C.) Wali. Chandra: a Biography of S. Chandrasekhar. University of Chicago Press, Chicago, IL, USA, 1991. ISBN 0-226-87054-5 (hardcover), 0-226-87055-3 (paperback). x + 341 + 32 pp. LCCN QB36.C46 W35 1991. URL http://www.loc.

gov/catdir/description/uchi052/90010845.html;   http://www.
loc.gov/catdir/enhancements/fy0608/90010845-t.html;   http:
//www.loc.gov/catdir/enhancements/fy0609/90010845-b.html.

Walters:2015:EPN

[2863] Robert L. Walters and W. H. Barber. Eléments de la philosophie de Newton. Les Œuvres complètes de Voltaire. Voltaire Foundation, Oxford, UK, 2015. ISBN 0-7294-0374-2. xxii + 805 pp. LCCN PQ2070 .B4 v.15.

Warwick:2003:MTC

[2864] Andrew Warwick. Masters of theory: Cambridge and the rise of mathematical physics. University of Chicago Press, Chicago, IL, USA, 2003. ISBN 0-226-87374-9 (hardcover), 0-226-87375-7 (paperback). xiv + 572 pp. LCCN QC19.6 .W37 2003. URL http:/

/www.loc.gov/catdir/bios/uchi051/2002153732.html;   http://
www.loc.gov/catdir/description/uchi051/2002153732.html;   http:
//www.loc.gov/catdir/toc/fy037/2002153732.html.

Watson:2000:PDG

[2865] James D. Watson. A Passion for DNA: Genes, Genomes, and Society. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2000. ISBN 0-19-850697-X. xx + 250 pp. LCCN QH438 .W38 2000.

Waugh:1990:JCM

[2866] William Waugh. John Charnley: the Man and the Hip. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. ISBN 3-540-19587-4, 0-387-19587-4. xvi + 268 pp. LCCN RD728.C48.

Wear:2000:KPE

[2867] A. (Andrew) Wear. Knowledge and Practice in Early Modern English Medicine, 1550–1680. Cambridge University Press, Cambridge, UK, 2000. ISBN 0-521-55226-5 (hardcover), 0-521-55827-1 (paperback), 0-511-05326-6 (e-book). viii + 496 pp. LCCN R486 .W43 2000. URL

http://www.loc.gov/catdir/description/cam021/00022600.html;
 http://www.loc.gov/catdir/samples/cam031/00022600.html;   http:
//www.loc.gov/catdir/toc/cam021/00022600.html.

Webster:2002:GIS

[2868] Charles Webster. The Great Instauration: Science, Medicine, and Reform, 1626–1660, volume 3 of Studies in the history of medicine. Peter Lang, Oxford, UK, second edition, 2002. ISBN 3-906768-37-6 (paperback), 0-8204-5616-0. ISSN 1424-7933. xl + 630 pp. LCCN Q125 .W35 2002.

Wheeler:1998:GBH

[2869] John Archibald Wheeler and Kenneth William Ford. Geons, Black Holes, and Quantum Foam: a Life in Physics. Norton, New York, NY, USA, 1998. ISBN 0-393-04642-7 (hardcover), 0-393-31991-1 (paperback), 0-393-07948-1 (e-book). 380 pp. LCCN QC16.W48A3 1998.

Williams:1990:RRC

[2870] Trevor Illtyd Williams. Robert Robinson, chemist extraordinary. Clarendon Press, Oxford, UK, 1990. ISBN 0-19-858180-7. viii + 201 + 7 pp. LCCN QD22.R63 W54 1990. URL http://www.

loc.gov/catdir/enhancements/fy0635/89022123-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0635/89022123-t.html.

Willmoth:1993:SJM

[2871] Frances Willmoth. Sir Jonas Moore: Practical Mathematics and Restoration Science. Boydell Press, Woodbridge, Suffolk, UK, 1993. ISBN 0-85115-321-6. xi + 244 pp. LCCN QA29.M65 W55 1993.

Wolterstorff:1996:JLE

[2872] Nicholas Wolterstorff. John Locke and the Ethics of Belief, volume 2 of Cambridge studies in religion and critical thought. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-55118-8 (hardcover), 0-521-55909-X (paperback). xxi + 248 pp. LCCN B1298.R4 W65 1996.

URL http://www.loc.gov/catdir/description/cam026/95007256.
html;   http://www.loc.gov/catdir/toc/cam024/95007256.html.

Woods:2000:ATA

[2873] L. C. (Leslie Colin) Woods. Against the Tide: an Autobiographical Account of a Professional Outsider. IOP Publishing, Bristol, UK, 2000. ISBN 0-7503-0690-4. xiii + 319 pp. LCCN QC16.W65 W66 2000.

URL http://www.loc.gov/catdir/enhancements/fy0653/99089496-
d.html.

Yeo:1993:DSW

[2874] Richard R. Yeo. Defining Science: William Whewell, Natural Knowledge, and Public Debate in Early Victorian Britain. Ideas in context. Cambridge University Press, Cambridge, UK, 1993. ISBN 0-521-43182-4 (hardcover). xiii + 280 pp. LCCN Q127.G4 Y46 1993.

URL http://www.loc.gov/catdir/description/cam025/92030620.
html;   http://www.loc.gov/catdir/toc/cam028/92030620.html.

Yeo:2001:EVS

[2875] Richard R. Yeo. Encyclopaedic Visions: Scientific Dictionaries and Enlightenment Culture. Cambridge University Press, Cambridge, UK, 2001. ISBN 0-521-65191-3 (hardcover). xxi + 336 pp. LCCN AE1 .Y46

2001. URL http://www.e-streams.com/es0805/es0805_2064.htm;
http://www.loc.gov/catdir/description/cam021/00063005.html;
 http://www.loc.gov/catdir/enhancements/fy0731/00063005-b.html;
 http://www.loc.gov/catdir/toc/cam021/00063005.html.

Zewail:2002:VTT

[2876] Ahmed H. Zewail. Voyage through time: walks of life to the Nobel Prize. American University in Cairo Press, Cairo, Egypt, 2002. ISBN 977-424-677-2. xii + 287 pp. LCCN QD22.Z46 A3 2002.

Ziman:2002:RSW

[2877] John Ziman. Real Science: What It Is, and What It Means. Cambridge University Press, Cambridge, UK, 2002. ISBN 0-521-77229-X (hardcover), 0-521-89310-0 (paperback), 0-511-03450-4 (e-book). xii + 399 pp. LCCN Q175 .Z547 2002.

Banks:1994:SJB

[2878] R. E. R. Banks et al., editors. Sir Joseph Banks: a Global Perspective. Published by the Royal Botanic Gardens, Kew on behalf of the cosponsors, Richmond, UK, 1994. ISBN 0-947643-61-3. LCCN QK31.B28 S57 1994.

Barrett:1987:CDN

[2879] Paul H. Barrett, Peter Jack Gautrey, Sandra Herbert, David Kohn, and Sydney Smith, editors. Charles Darwin’s notebooks, 1836–1844: geology, transmutation of species, metaphysical enquiries. Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-35055-7 (hardcover), 0-521-09975-7 (paperback). viii + 747 pp. LCCN QH31.D2; QH365.Z9 B37 1987.

Bethe:1997:SWH

[2880] Hans Albrecht Bethe. Selected works of Hans A. Bethe: with commentary, volume 18 of World Scientific series in 20th century physics. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1997. ISBN 981-02-2876-7. viii + 605 pp. LCCN QC780 .B452

1996. URL http://www.zentralblattmath.org/zmath/en/search/?

an=0916.01027. English translations (when originals are in German) of 28 of the author’s papers, accompanied by a list of 290 of his publications.

Burkhardt:1985:CCD

[2881] Frederick Burkhardt and Sydney Smith, editors. The correspondence of Charles Darwin. Cambridge University Press, Cambridge, UK, 1985– 2015. ISBN 0-521-25587-2 (vol. 1), 0-521-25588-0 (vol. 2), 0-521-25589-9 (vol. 3), 0-521-25590-2 (vol. 4), 0-521-25591-0 (vol. 5), 0-521-25586-4 (vol. 6), 0-521-38564-4 (vol. 7), 0-521-44241-9 (vol. 8), 0-521-45156-6 (vol. 9), 0-521-59032-9 (vol. 10), 0-521-59033-7 (vol. 11), 0-521-59034-5 (vol. 12), 0-521-82413-3 (vol. 13), 0-521-84459-2 (vol. 14), 0-521-85931-X (vol. 15: hardcover), 0-521-88195-1 (vol. 16, part 1), 0-521-88196-X (vol. 16, part 2), 0-521-51836-9 (vol. 16), 0-521-19030-4 (vol. 17), 0-521-76889-6 (vol. 18), 1-107-01648-7 (vol. 19), 1-107-03844-8 (vol. 20), 1-107-05214-9 (vol. 21), 1-107-08872-0 (vol. 22). ???? pp. LCCN QH31.D2 A4 1985.

URL http://www.loc.gov/catdir/description/cam022/84045347.
html;   http://www.loc.gov/catdir/toc/cam026/84045347.html.

Burkhardt:1996:CDL

[2882] Frederick Burkhardt, editor. Charles Darwin’s Letters: a Selection, 1825–1859. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-56212-0 (hardcover). xxvi + 249 pp. LCCN QH31.D2 A4 1996.

URL http://www.loc.gov/catdir/description/cam027/95051854.
html;   http://www.loc.gov/catdir/toc/cam026/95051854.html.

Campbell-Kelly:2003:HMT

[2883] Martin Campbell-Kelly, Mary Croarken, Raymond Flood, and Eleanor Robson, editors. The History of Mathematical Tables: From Sumer to Spreadsheets. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2003. ISBN 0-19-850841-7. viii + 361 pp. LCCN QA47

.H57 2003. URL http://www.loc.gov/catdir/enhancements/fy0620/
2004298837-d.html;   http://www.loc.gov/catdir/enhancements/
fy0620/2004298837-t.html.

Cardwell:2003:DST

[2884] D. S. L. (Donald Stephen Lowell) Cardwell and Richard Leslie Hills, editors. The development of science and technology in nineteenth-century Britain: the importance of Manchester, volume CS765 of Variorum collected studies series. Ashgate/Variorum, Aldershot, Hampshire, UK, 2003. ISBN 0-86078-908-X. xx + 320 pp. LCCN Q127.G4 C37 2003.

URL http://www.loc.gov/catdir/toc/fy038/2002034533.html.

Chambers:1991:SCF

[2885] Robert Guy Chambers et al., editors. Sir Charles Frank, OBE, FRS: an Eightieth Birthday Tribute. Adam Hilger Ltd., Bristol, UK, 1991. ISBN 0-7503-0100-7. xii + 448 pp. LCCN QD906.5 .S57 1991. URL http://

www.loc.gov/catdir/enhancements/fy0745/90028640-d.html.

Cohen:2002:CCN

[2886] I. Bernard Cohen and George E. (George Edwin) Smith, editors. The Cambridge Companion to Newton. Cambridge University Press, Cambridge, UK, 2002. ISBN 0-521-65177-8 (hardcover), 0-521-65696-6 (paperback). xiv + 500 pp. LCCN QC16.N7 C35 2002. URL

http://www.e-streams.com/es0601/es0601_2274.html;   http://
www.loc.gov/catdir/description/cam022/2001037836.html;   http:
//www.loc.gov/catdir/samples/cam033/2001037836.html;   http:
//www.loc.gov/catdir/toc/cam024/2001037836.html.

Cukerman:1999:ASS

[2887] Veniamin Aronovich Čukerman and Zinaida Matveevna Azarch, editors. Arzamas-16: Soviet Scientists in the Nuclear Age: a Memoir, volume 3 of Russian memoirs series. Bramcote Press, Nottingham, UK, 1999. ISBN 1-900405-04-0. xxxiv + 182 pp. LCCN QC773.3.R9 T7813 1999.

URL http://www.gbv.de/dms/bowker/toc/9781900405041.pdf.

Darrigol:2000:EAE

[2888] Olivier Darrigol. Electrodynamics from Ampère to Einstein. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2000. ISBN 0-19-850594-9. xix + 532 pp. LCCN QC630.5 .D37 2000.

Darwin:2003:LED

[2889] Charles Darwin and Desmond King-Hele, editors. The life of Erasmus Darwin. Cambridge University Press, Cambridge, UK, 2003. ISBN 0-521-81526-6 (hardcover). xx + 172 pp. LCCN QH31.D3

D37 2003. URL http://www.loc.gov/catdir/description/cam022/
2002067073.html;   http://www.loc.gov/catdir/enhancements/fy0731/
2002067073-b.html;   http://www.loc.gov/catdir/samples/cam033/
2002067073.html;   http://www.loc.gov/catdir/toc/cam025/2002067073.
html.

Day:1999:PTS

[2890] Peter Day, editor. The Philosopher’s Tree: a Selection of Michael Faraday’s Writings. IOP Publishing, Bristol, UK, 1999. ISBN 0-7503-0571-1 (paperback), 0-7503-0570-3 (hardcover). xv + 211 pp. LCCN QC16.F2

A3 1999. URL http://www.loc.gov/catdir/enhancements/fy0668/
98054683-d.html.

Dekker:1999:GGC

[2891] Elly Dekker, Silke Ackermann, Kristen Lippincott, et al., editors. Globes at Greenwich: a Catalogue of the Globes and Armillary Spheres in the National Maritime Museum, Greenwich. Oxford University Press and the National Maritime Museum, Oxford, UK, 1999. ISBN 0-19-856559-3. ix + 592 pp. LCCN GA195.G7 D45 1999. URL http://www.

loc.gov/catdir/enhancements/fy0608/00269924-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0608/00269924-t.html.

Evelyn:1995:WJE

[2892] John Evelyn and Guy De la Bédoyère, editors. The Writings of John Evelyn. Boydell Press, Woodbridge, Suffolk, UK, 1995. ISBN 0-85115-631-2 (hardcover). 435 pp. LCCN PR3433.E5 A6 1995.

Feingold:1990:BNL

[2893] Mordechai Feingold, editor. Before Newton: the life and times of Isaac Barrow. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-30694-9. xi + 380 pp. LCCN QA29.B37 B44 1990. URL

http://www.loc.gov/catdir/description/cam024/89038508.html;
 http://www.loc.gov/catdir/enhancements/fy0632/89038508-t.html.

Force:1999:NRC

[2894] James E. Force and Richard H. (Richard Henry) Popkin, editors. Newton and Religion: Context, Nature, and Influence, volume 161 of International archives of the history of ideas = Archives internationales d’histoire des idées. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1999. ISBN 0-7923-5744-2 (hardcover). xvii + 325 pp. LCCN BX4827.N48 N48 1999. URL http://www.

loc.gov/catdir/enhancements/fy0821/99015083-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0821/99015083-t.html.

Goodman:1991:RSE

[2895] David C. (David Charles) Goodman and Colin A. (Colin Archibald) Russell, editors. The Rise of Scientific Europe, 1500–1800. Hodder and Stoughton, London, UK, 1991. ISBN 0-340-55861-X. ix + 437 pp. LCCN Q127.E8 R574 1991.

Harman:1990:SLP

[2896] P. M. (Peter Michael) Harman, editor. The Scientific Letters and Papers of James Clerk Maxwell. Vol. I: 1846–1862, volume I. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-25625-9 (hardcover). xxviii + 748 pp. LCCN QC21.2; QC670 .M385 1990.

URL http://www.gbv.de/dms/bowker/toc/9780521256254.pdf;
http://www.loc.gov/catdir/description/cam023/89000452.html;
 http://www.loc.gov/catdir/enhancements/fy0732/89000452-b.html;
 http://www.loc.gov/catdir/toc/cam021/89000452.html;   http:
//www.zentralblatt-math.org/zmath/en/search/?an=0744.01034.

Harman:1995:SLP

[2897] P. M. Harman, editor. The Scientific Letters and Papers of James Clerk Maxwell. Vol. II: 1862–1873, volume II. Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-25626-7 (hardcover). xxx + 999 pp. LCCN QC21.2; QC670 .M385 1990.

Harman:1998:NPJ

[2898] P. M. (Peter Michael) Harman. The natural philosophy of James Clerk Maxwell. Cambridge University Press, Cambridge, UK, 1998. ISBN 0-521-56102-7, 0-521-00585-X (paperback). xiv + 232 pp. LCCN QC16.M4

H37 1998. URL http://www.loc.gov/catdir/description/cam028/
97035703.html;   http://www.loc.gov/catdir/toc/cam021/97035703.
html;   http://www.zentralblattmath.org/zmath/en/search/?an=
1005.01006.

Harman:2002:SLP

[2899] P. M. Harman, editor. The Scientific Letters and Papers of James Clerk Maxwell. Vol. III: 1874–1879, volume III. Cambridge University Press, Cambridge, UK, 2002. ISBN 0-521-25627-5. xxvii + 932 pp. LCCN

QC21.2; QC670 .M385 1990. URL http://www.zentralblatt-math.
org/zmath/en/search/?an=0744.01034.

Heberden:1989:WHP

[2900] Ernest Heberden and Hugh L’Étang, editors. William Heberden: physician of the age of reason. Eponymists in medicine. Royal Society of Medicine Services, London, UK and New York, NY, USA, 1989. ISBN 1-85315-117-3 (hardcover), 1-85315-116-5 (paperback). xiv + 246 pp. LCCN R489.H47 H4 1990; R489.H424 H4239 1989.

Hogben:1998:LHS

[2901] Lancelot Thomas Hogben, Adrian Hogben, and Anne Hogben, editors. Lancelot Hogben, Scientific Humanist: an Unauthorised Autobiography. Merlin Press, Woodbridge, Suffolk, UK, 1998. ISBN 0-85036-470-1. xvii + 254 + 16 pp. LCCN QH31.H644 A3 1998. URL http://www.loc.

gov/catdir/enhancements/fy1214/00500714-d.html.

Hoskin:2003:CHA

[2902] Michael A. Hoskin, editor. Caroline Herschel’s Autobiographies. Science History Publications, Cambridge, UK, 2003. ISBN 0-905193-06-7. vii + 147 pp. LCCN QB36.H6 H47 2003.

Hunter:1989:RHN

[2903] Michael Cyril William Hunter and Simon Schaffer, editors. Robert Hooke: new studies; grown out of a conference. Boydell Press, Woodbridge, Suffolk, UK, 1989. ISBN 0-85115-523-5. LCCN Q143.H7 H86 1989.

Hunter:1994:RBR

[2904] Michael Cyril William Hunter, editor. Robert Boyle Reconsidered. Cambridge University Press, Cambridge, UK, 1994. ISBN 0-521-44205-2. xviii + 231 pp. LCCN Q143.B77 R6 1994. URL

http://www.loc.gov/catdir/description/cam025/93023081.html;
 http://www.loc.gov/catdir/samples/cam031/93023081.html;   http:
//www.loc.gov/catdir/toc/cam021/93023081.html.

Hunter:1998:ASR

[2905] Michael Cyril William Hunter, editor. Archives of the Scientific Revolution: the Formation and Exchange of Ideas in Seventeenth-century Europe. Boydell Press, Woodbridge, UK, 1998. ISBN 0-85115-553-7. xii + 216 pp. LCCN Q224 .A73 1998.

James:1991:CMF

[2906] Frank A. J. L. James, editor. The correspondence of Michael Faraday. [History of technology series]. Institution of Electrical Engineers, London, UK, 1991–2012. ISBN 0-86341-248-3 (vol. 1), 0-86341-249-1 (vol. 2), 0-86341-250-5 (vol. 3), 0-86341-251-3 (vol. 4). ???? pp. LCCN QC16.F2 A4 1991.

James:2002:CPL

[2907] Frank A. J. L. James, editor. The common purposes of life: science and society at the Royal Institution of Great Britain. Ashgate, Aldershot, Hampshire, England, 2002. ISBN 0-7546-0960-X. xvi + 456 pp. LCCN Q41 .C66 2002. URL http://www.loc.gov/

catdir/toc/fy036/2002026184.html;   http://www.tandfebooks.
com/isbn/9781315264141.

Keynes:1989:ESC

[2908] W. Milo (William Milo) Keynes and G. A. (Geoffrey Ainsworth) Harrison, editors. Evolutionary Studies: a Centenary Celebration of the Life of Julian Huxley: Proceedings of the Twenty-fourth Annual Symposium of the Eugenics Society, London, 1987, Studies in biology, economy, and society. Macmillan, Houndmills, Basingstoke, Hampshire, 1989. ISBN 0-333-45723-4. LCCN GN281 .E84 1989.

Keynes:1993:SFG

[2909] W. Milo (William Milo) Keynes, editor. Sir Francis Galton, FRS: the Legacy of His Ideas. Studies in biology, economy, and society. MacMillan Publishing Company, New York, NY, USA, 1993. ISBN 0-333-54695-4. ix + 237 pp. LCCN Q143.G3 G35 1991.

Keynes:2000:CDZ

[2910] R. D. Keynes, editor. Charles Darwin’s Zoology Notes and Specimen Lists from H.M.S. Beagle. Cambridge University Press, Cambridge, UK, 2000. ISBN 0-521-46569-9 (hardcover), 0-521-67350-X (paperback). xxxiv + 430 pp. LCCN QH365.Z9 K4 2000. URL http://www.

loc.gov/catdir/description/cam022/2001269276.html;   http:/
/www.loc.gov/catdir/toc/cam026/2001269276.html.

Knox:2003:NHH

[2911] Kevin C. Knox and Richard Noakes, editors. From Newton to Hawking: a History of Cambridge University’s Lucasian Professors of Mathematics. Cambridge University Press, Cambridge, UK, 2003. ISBN 0-521-66310-5. xxiv + 486 pp. LCCN QA28 .F76 2003. URL http://www.

loc.gov/catdir/description/cam032/2003043584.html;   http:/
/www.loc.gov/catdir/enhancements/fy0731/2003043584-b.html;
http://www.loc.gov/catdir/toc/cam032/2003043584.html.

Krishna:1993:BST

[2912] V. V. Krishna, editor. S. S. Bhatnagar on Science, Technology and Development, 1938–54. Wiley Eastern, New Delhi, India, 1993. ISBN 81-224-0501-0. xiv + 290 pp. LCCN T177.I4 B47 1993.

LeveltSengers:2002:HFU

[2913] J. M. H. Levelt Sengers, editor. How fluids unmix: discoveries by the school of van der Waals and Kamerlingh Onnes, volume 4 of History of science and scholarship in the Netherlands. Koninklijke Nerlandse Akademie van Wetenschappen, Amsterdam, The Netherlands, 2002. ISBN 90-6984-357-9 (hardcover). xv + 302 pp. LCCN QC307

.L49 2002. URL http://www.loc.gov/catdir/description/uchi052/
2004422975.html.

Lincoln:1998:SEP

[2914] Margarette Lincoln, editor. Science and exploration in the Pacific: European voyages to the southern oceans in the eighteenth century. Boydell Press, Woodbridge, Suffolk, UK, 1998. ISBN 0-85115-721-1 (hardcover). LCCN G420.C65 S35 1998. Published in association with the National Maritime Museum.

Mahon:2003:MWC

[2915] Basil Mahon. The man who changed everything: the life of James Clerk Maxwell. John Wiley, New York, NY, USA, 2003. ISBN 0-470-86088-X, 0-470-86171-1 (paperback). xx + 226 + 16 pp.

LCCN QC16.M4 M34 2003. URL http://www.loc.gov/catdir/
bios/wiley047/2004426225.html;   http://www.loc.gov/catdir/
description/wiley042/2004426225.html;   http://www.loc.gov/
catdir/toc/wiley041/2004426225.html.

McCartney:2003:PIP

[2916] Mark McCartney and Andrew Whitaker, editors. Physicists of Ireland: passion and precision. Institute of Physics Publishers, Bristol and Philadelphia, 2003. ISBN 0-7503-0866-4 (hardback). xiv + 298 pp. LCCN QC15 .P47 2003. URL http://www.loc.gov/

catdir/enhancements/fy0806/2007298520-d.html;   http://www.
loc.gov/catdir/toc/fy0802/2007298520.html.

Michette:1996:XRF

[2917] Alan G. Michette and Sławka Pfauntsch, editors. X-rays: the first hundred years. John Wiley, New York, NY, USA, 1996. ISBN 0-471-96502-2. xi + 262 pp. LCCN QC481 .X24 1996.

OHear:1995:KPP

[2918] Anthony O’Hear, editor. Karl Popper: Philosophy and Problems, volume 39 of Royal Institute of Philosophy supplement. Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-55815-8 (paperback). ISSN 1358-2461. iv + 297 pp. LCCN B1649.P64 K36 1995. URL

http://www.loc.gov/catdir/description/cam027/95033090.html;
 http://www.loc.gov/catdir/toc/cam025/95033090.html.

Pais:2000:GSP

[2919] Abraham Pais. The Genius of Science: a Portrait Gallery. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2000. ISBN 0-19-850614-7 (hardcover). 356 pp. LCCN Q141 .P29 2000. URL

ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/;   http:/
/www.loc.gov/catdir/toc/fy02/99046603.html.

Peierls:1997:SSP

[2920] Sir Rudolf Ernst Peierls and Richard Henry Dalitz, editors. Selected scientific papers of Sir Rudolf Peierls: with commentary, volume 19 of World Scientific series in 20th century physics. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1997. ISBN 981-02-2692-6. xxiii + 805 pp. LCCN QC21.2 .P42 1997.

Rowlands:1994:OLI

[2921] Peter Rowlands and J. P. Wilson, editors. Oliver Lodge and the invention of radio. PD Publications, Liverpool, UK, 1994. ISBN 1-873694-02-4. vii + 241 + 58 pp. LCCN QC16.L6 O45 1994.

Stoicheff:2002:GHI

[2922] B. P. (Boris Peter) Stoicheff. Gerhard Herzberg: an illustrious life in science. NRC press biography series. NRC Press, Ottawa, ON, Canada, 2002. ISBN 0-660-18757-4, 0-660-19258-6 (e-book). ISSN 1701-1833. xiii + 468 pp. LCCN QC16.H47 S86 2002.

Waters:1992:JHB

[2923] C. Kenneth Waters and Albert Van Helden, editors. Julian Huxley, biologist and statesman of science: proceedings of a conference held at Rice University, 25–27 September 1987. Rice University Press, Houston, TX, USA, 1992. ISBN 0-89263-314-X. LCCN QH31.H88 J85 1992.

URL http://www.loc.gov/catdir/enhancements/fy1003/92050136-
b.html;   http://www.loc.gov/catdir/enhancements/fy1011/92050136-
d.html.

Willmoth:1997:FSN

[2924] Frances Willmoth, editor. Flamsteed’s stars: new perspectives on the life and work of the first Astronomer Royal, 1646–1719. Boydell Press, Woodbridge, Suffolk, UK, 1997. ISBN 0-85115-706-8 (hardcover). xiv + 271 pp. LCCN QB36.F6 F53 1997.

Youings:1993:SJB

[2925] Joyce A. Youings, editor. Sir John Bowring 1792-1872: aspects of his life and career. Devonshire Association, Plymouth, UK, 1993. ISBN 0-85214-052-5. LCCN CT788.B764 S57 1993.